A Complete Bibliography of Publications in Technovation

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

12 June 2024 Version 1.31

Title word cross-reference

$160 [Sap10]. $19.75 [Ros89b]. $19.95 [Hal97a]. $24.00 [Lin90]. $24.95 [Gil97, Bes95a]. $29.95 [Ros90]. 3 [ALFM23, CB19, CNN23, DTI^{+}23, SA22, UF22, WAdN19, ZYSL21]. $34.95 [Buc91]. $37.95 [Ros89c]. $80.00 [Giu09]. E = mc^{2} [Lin17a].

0 [Ada07, Aug09, Aug06a, Aug06b, Bes88, Bes02a, Bes02c, Bes03a, Bon02, Box01, Coo07, Cou90, Cre04, Del97, Dut07, Fit00, For90, Fra02c, Fra02a, Hay97a, Hay01a, Hay01c, Hob03, Hop04, Kap01, Koc09, Lin90, Lop06, New04, Nuv07, OH04, Poy97, Rot88, Saa98, Sco89, Tat04, Tra00, Ver90a, Ver90b, Wan04]. 0-08-044214-5 [Lop06]. 0-09174-1181 [Ver90b]. 0-13-0437875 [Bes95b]. 0-13-145947-3 [Min94]. 0-13-149626-3 [Per97b]. 0-13-295304-8 [Per96]. 0-13-556507-3 [Her92b]. 0-13-781824-6 [Gar94]. 0-19-268255-5 [But96]. 0-19-514097-4 [Bes02a]. 0-19-828904-9 [Poy97]. 0-19-829462-X [Fra02b].

0-19-8295375 [Wan04]. 0-19-829542-1 [Hob03]. 0-19-926322-1 [Edw04]. 0-19-926943-2 [dC07]. 0-19-926944-0 [dC07]. 0-19-929619-7 [Tan08]. 0-226-85175-3 [Min96]. 0-230-00201-3 [Coo07]. 0-230-61173-7 [Koc09]. 0-262-01198-0 [OH04]. 0-262-02327-X [Hol93]. 0-262-18226-2 [Par03a]. 0-262-19514-3 [Aug06a]. 0-273-64322-3 [Hay01b]. 0-330-32046-7 [Joh92d]. 0-335-19158-4 [Sha95]. 0-408-028940-4 [Rot88]. 0-408-03957-4 [Hay92d]. 0-412-48600-8 [Hay93d]. 0-412-991217 [Col97]. 0-415-02379-3 [Hay92e]. 0-415-03480-9 [Hay92c]. 0-415-07080-5 [Bes93]. 0-415-07385-5 [Gar95]. 0-415-07926-8 [Har93b]. 0-415-1162-7 [Ric95]. 0-415-125030 [Hay97a]. 0-415-12609-6 [Bes97]. 0-415-12610-X [Bes97]. 0-415-14624-0 [Hay97b]. 0-415-16828-7 [Wya00]. 0-415-19210-2 [Tra00]. 0-415-22884-0 [Aug06b]. 0-41513-323-8 [Saa98]. 0-419-16920-2 [Hay93a]. 0-419-21350-3 [Hob97]. 0-442-00243-2 [Joh92b]. 0-442-02246-8 [Bul97]. 0-442-31493-0 [Hay93a]. 0-444-42865-8 [Bes88]. 0-471-31096-4 [Gil97]. 0-471-31546-X [Pre01]. 0-471-39283-9 [Fra02d]. 0-471-48557-8 [Fra02a]. 0-471-54674-7 [Bes94b]. 0-471-61024-0 [Ver91b]. 0-471-61997-3 [Mur91]. 0-471-92169-6 [Mor91, Sco89]. 0-471-92499-7 [Joh92g]. 0-471-92858-5 [Her92c]. 0-471-93418-6 [Bes94c]. 0-471-94451-3 [Per95]. 0-471-95490-X [Del97]. 0-471-96452-2 [Gri98]. 0-521-55393-8 [Hay01c]. 0-521-55641-4 [Hay01c]. 0-521-63327-3 [Hay01a]. 0-521-63361-3 [Hay01a]. 0-521-66269-9 [Kap01]. 0-521-68415-3 [Nuv07]. 0-521-81086-8 [Cre04]. 0-566-07836-8 [O’H99]. 0-566-07961-5 [Fit00]. 0-566-08027-3 [Fit00]. 0-566-08364-7 [Mer02]. 0-566-08567-4 [Tat04]. 0-631-15304-7 [Her91]. 0-631-15392-6 [Cou91, Cou90]. 0-631-17104-5 [Ver91a, Ver90a]. 0-669-1277-X [For91, For90]. 0-669-21128-1 [Lin91, Lin90]. 0-7123-0788-5 [Hay93b]. 0-7425-1846-9 [Wya04]. 0-7456-2456-1 [Dut07]. 0-7494-0201-6 [Joh92f]. 0-7494-0527-9 [Joh92e]. 0-7494-1007-8 [Gil96]. 0-7494-3686-7 [New04]. 0-7506-4999-2 [Bes02c]. 0-7619-2181-8 [Fra03a]. 0-7619-2468-X [Fra03a]. 0-7619-3477-4 [Ada07]. 0-7619-4874-0 [Kuc07]. 0-7619-4875-9 [Kuc07]. 0-7619-6948-9 [Fra03b]. 0-7619-6949-7 [Fra03b]. 0-7619-9670-2 [Bas04]. 0-7923-9051-2 [Joh92a]. 0-7923-9758-4 [Ano97b]. 0-7923-9760-6 [Hay97d]. 0-8039-5942-7 [Box01]. 0-8039-7048-X [Hay97c]. 0-8039-7049-8 [Hay97c]. 0-8122-8205-1 [Buc91]. 0-8493-3741-0 [Bon02]. 0-85296-768-3 [Bes02d]. 0-86341-265-3 [Rus94]. 0-86341-285-8 [Rus94]. 0-87584-226-7 [Dav92]. 0-87584-342-5 [Bes96]. 0-87584-413-8 [Bes95a]. 0-87584-865-6 [Bes01b]. 0-88427-170-6 [Fra02c]. 0-939246-53-8 [Pov97]. 0-948061-72-3 [Pot93]. 0-9523328-4-1 [Bes02b]. 0-9662697-1-3 [Bes03a]. 000 [Ano93b, Ano93i, Ano93l, Ano93n, Ano93s].

1 [Arr04, Aug09, Ban04, Bes01a, CF07, Col08, Dut07, Giu09, Har93a, Hob03, Lae04, Lin90, Per04, Que06, Rad97, Sap05, Szw07, ZS04a]. 1-220 [Koc09]. 1-4020-7000-4 [Per04]. 1-4039-4245-5 [Giu09]. 1-4129-1903-7 [Szw07]. 1-4805-7747-2 [Kra15]. 1-57851-903-2 [Bes01a]. 1-84064-431-1 [Arr04]. 1-84064-469-9 [Lae03]. 1-84064-640-3 [BH03]. 1-84064-922-4 [Kap03].

1-84064-957-7 [Ban04]. 1-84376-018-5 [De 04]. 1-84376-210-2 [Mor04]. 1-84376-231-5 [Que06]. 1-84376-376-1 [Sap05]. 1-84376-497-0 [Wya07]. 1-84376-799-6 [Her04]. 1-84376-855-0 [Aug09]. 1-84542-119-1 [Col08]. 1-84542-199X [CF07]. 1-85278-655-8 [Hay95]. 1-85396-274-0 [Gar96]. 1-85396-339-9 [Per97a]. 1-85423-106-5 [Bes95c]. 1-85567-082-8 [Har93a]. 1-85898-717-2 [Lal00]. 1-85907-002-7 [Bes94d]. 1-85907-004-3 [BB95]. 1-85907-005-1 [Bes94a]. 1-85972-184-2 [Rad97]. 1-85972-330-0 [Har97c]. 1-86094-359-4 [Lae04]. 1-86094-367-5 [Dre04]. 1-861-52023-9 [Coo97]. 1-871786-04-5 [Hay93c]. 1-89872312-5 [Hal97a]. 10 [Pra10]. 10-0-415-43177-8 [Pra10]. 10.95 [Hay86e]. 10.99 [Gar95]. 110 [Ano94-42]. 1181 [Ver90b]. 12 [II92b, JWS92b, Swi92b, Zho92b]. 120 [Ano94-33]. 125030 [Hay97a]. 128 [Ano94c]. 13 [Ano93a, Ano93b, Ano93f, Ano93e, Ano93j, Ano93i, Ano93k, Ano93l, Ano93m, Ano93o, Ano93n, Ano93p, Ano93s]. 13.25 [Hay85b]. 13.75 [Hay86d]. 14 [Ano94c, Ano94f, Ano94k, Ano94l, Ano94m, Ano94n, Ano94z, Ano94-27, Ano94-30, Ano94-31, Ano94-34, Ano94-37, Ano94-33, Ano94-38, Ano94-42, BM94b, Hol94b, Paw94, Szá94b]. 14.99 [Pov97, Ric95]. 15.00 [Hay85c]. 15.95 [Hay86a]. 159 [Ano94f]. 16.50 [Gar85]. 16.99/$29.95 [Dut07]. 16th [Ano07k]. 17-99 [Fra02a]. 171 [Ano94l]. 175 [Rot87a]. 179 [Ano94-31]. 18.50 [Ram86]. 19 [ABAHP^{+}23, BKA^{+}23, BEE24, DGR23, EHSB24, FGB^{+}24, GP22, GMB^{+}23, KRK^{+}22, Kra22b, MWJ23, Rea21, SJP23, SSKB24, TFS^{+}22, YVG24]. 19.50 [Mil88]. 19.95 [Bes95b, Gar94, Kap03]. 195 [Ano94-30]. 1981/1982 [Ano82a]. 1983/1984 [Ano84b]. 1985 [Hob87, Mil85]. 1988 [Ano88h]. 1990s [ALRSD04, Bes94c, Gar96, Her92a, Hay83]. 1993 [Mer94b, Par03a]. 1997 [Fra00b]. 1999 [Bes01b, Eva02, Fra99, Fra01a]. 199X [CF07]. 1AW [Hay93b]. 1LA [Per97a]. 1UD [Gri98, Per95, Pov97]. 1XW [Sha95].

2 [Bes01a, Bes02c, Rad97, Tra00]. 2-stroke [WTN98a]. 20.00 [Hay88d]. 2000 [Gra01, Pap02a]. 2001 [Bes03b, Gra03, Pap03]. 2002 [Arc03, Gra04, Pap04]. 2005 [Gof07]. 2007 [Ano07k]. 2016 [Coo16]. 2017 [Bai18]. 2018 [CR19]. 2019 [HdWN21]. 2020 [Ano20a, Ano20d, Ano21b, Ano21s]. 2021 [Ano21a, Ano21c, Ano21d, Ano21t, Ano21u, Ano21v, Ano21w]. 2022 [Ano22a, Ano22b, Ano22c, Ano22n, Ano22q, Ano22p, Ano22r, Ano22s, Ano22t]. 2023 [Ano23a, Ano23b, Ano23c, Ano23n, Ano23o, Ano23q, Ano23p, Ano23r, Ano23s, Ano23t]. 2024 [Ano24a, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l]. 203 [Ano94-34]. 21st [Ano88m, Hit96, JMK05]. 22 [Joh92b]. 24.99 [Hay01b]. 25.00 [Gil96]. 26 [YGW08]. 27.00 [Hay88a]. 29 [BP09a]. 29.00 [Hay88c]. 29.95 [Bes94d, Bes95c, Dod88, O’H99]. 2nd [Ros89c].

3 [Aug06a, Bes03a, Coo07, Hay01a, Hop04, Nuv07, Won08, dC07]. 3-540-40717-0 [Hop04]. 3-8244-6253-2 [Hay96a]. 3.0 [CMH14]. 300p [Wya07]. 31 [CDD^{+}14]. 32.00 [Hay86c, Hay88b]. 33.50 [Par03a]. 35 [Har97c]. 35.00 [Hay88g, Hay88e]. 35.95 [Fra02d]. 380 [Ano94k]. 39.95 [Gar96]. 394 [Ano94z]. 3G [KY06]. 3GPP [JKF19, LBG14]. 3HR [Tat04]. 3rd [Hay84].

4 [Ada07, Bes02a, Lae04, Per04, Rot88, Tan08, Tat04]. 4.0 [CFP^{+}23, CCBP22, KRM^{+}24, TFS^{+}22]. 40 [Oak97]. 40.00 [Bes02b, Rus94]. 406 [Ano94-37]. 418 [Ano94-38]. 424 [Ano93a]. 43 [Mou17]. 43-44 [Mou17]. 431 [Hay93d]. 438 [Ano93j]. 44 [Mou17]. 44.50 [Pap02b]. 447 [Ano93f]. 45 [Lin11f, Lin10c]. 469 [Ano93e]. 48 [Dre04]. 495 [Ano93o]. 4BJ [Hay93c]. 4EE [Gar95]. 4th [II21].

5 [Fit00, Giu09, Lop06, Que06, Ver90a]. 5.0 [BK24]. 50 [Edw04]. 507 [II92b]. 514 [JWS92b]. 518 [Ano93p]. 520 [Par06, Zho92b]. 532 [Ano93m]. 533 [Swi92b]. 544 [Ano93k]. 55/$77.95 [Dut07]. 59.95 [Lae03]. 5th [Hay93c].

6 [Cou90, Fra02c, Sco89]. 66.00 [Hay87b]. 69 [Ano94-27]. 69.95 [Her04].

7 [Ban04, Box01, Koc09, New04, Szw07]. 74.75 [Ano97b, Hay97d]. 75.00 [Hay95]. 7EZ [Per97b]. 7PW [BB95, Bes94a, Bes95c]. 7th [Ano90o].

8 [Bes88, Cre04, Fra02a, Har93a, Hay01c, Pra10, Saa98, Sap10]. 82.00 [Hay87a, Hay87c]. 8295375 [Wan04]. 83 [Ano94n].

9 [Kap01, Poy97]. 9.95 [Hay85a, Hay86f]. 92 [Ano94m, Hay96a]. 95.00 [Bes02d]. 978 [Sap10, Tan08, Won08, dC07]. 978-0-19-929619-4 [Tan08]. 978-0-415-41975-8 [Sap10]. 978-019-926943-3 [dC07]. 978-019-926944-0 [dC07]. 978-1-4805-7747-3 [Kra15]. 978-1-84720-056-3 [Won08]. 98 [YJLL21].

A. [Ano93f]. Åke [Dre06]. Abandoning [LYY^{+}22]. Abdullah [AHAAEA03]. Abilities [WVG^{+}98, WGV^{+}98]. above [JWY23]. abroad [ZIB23]. Absorbing [RvdSdE22]. absorption [GL22]. Absorptive [MJ24, MP09b, NQ05, SC17, ALMW21, BBS22, CJL22, DLTQ23, EH14, EHHW17, FMFMA16, dCHDAALMM07, HLO14, JBGMM11, LWC^{+}12, PS21, PMT23, PA15, SBP21, SCK10, SCK11, WSH23, YMMYK24, XR08]. abstract [Ano95o, Ano95p, Ano95q, Ano97q, HAH84]. abstracts [Ano92h, Ano93q, Ano93r, Ano94-39, Ano94-40, Ano95r, Ano95x, Ano95s, Ano95t, Ano95u, Ano95v, Ano95w, Ano96o, Ano96u, Ano96p, Ano96q, Ano96r, Ano96s, Ano96t, Ano97r, Ano97s, Ano97t, Ano97u, Ano97v, Ano97w, Ano97x, Ano97y, Ano98o]. academia [AG24, Ahn95, AC04, Bje10, BJ05, DMNR12, Fuj98, GKSW19, HKW^{+}13, MI07, MR19]. academia-industry [Fuj98]. Academic [Ano97b, Ano07l, EKY24, Hal97a, Hay97d, HW17, Kuc06, Lin14, Lin17c, Par03b, Per04, QWM92, Sha91b, vGS09, AKKS12, ABS24, BRH^{+}97, CDMV24, DV23, GvBR10, JFCFMC^{+}12, KA15, KM97, LPMV22, LM94a, LM94b, LL05b, LP07, Mas23, NHS24, NPS02, OIOA01, OIN96, PDOL11, PD10, RV21, RGFLRP21, SOKE15, SSC23, SSS17, SJ16, SLJ13, TKT13, Wec15, Wür17, ZGSG23]. academic-entrepreneurial [PD10]. Academic-industry [QWM92, LM94a, LM94b]. academics [MT08, SWW12, VM97, vBDK21]. accelerate [TM06]. acceleration [Mer22, SC20]. Accelerator [Hed89, BEE24, CB22, MKL19, PCWV16]. Accelerators [MNL22, CvZ05, DID21, DCD22]. acceptance [ASGC09, BKA^{+}23, CT06, CC14, CSL06, Dan19, FGB^{+}24, HO11, HJM08, VM11]. access [ABS24, CT12, CCY05, GGJ22, Pal00, SSF12]. accompany [WBM22]. according [HN08, PSdLC02]. Accounting [Ano03q, ON88, HGL98, LL05a, SD95]. Accreditation [Ano07l]. accumulated [KK23]. accumulation [Fig02, Se14]. accumulative [CG11]. achievable [Wil14]. achieve [BSG^{+}24, GCH96]. achievement [AHAAEA03, HKW^{+}13]. achievements [XX21]. Achieving [Mac88, ZGSG23, GMB^{+}23]. acquire [WVG^{+}98, WGV^{+}98]. acquirer [SH22]. acquiring [SS07]. acquisition [Aka98b, Aka98a, CY00, DFBC98a, DFBC98b, FC95, GR23, GL22, HT08, Hus10, JGM98b, JGM98a, KLL87, KKG07, LPN22, OTA01, SH22, Sik96, VJ23a]. acquisitions [ALRSD04, BC82, dCHDAALMM07, JBL23, YNS18, dMD05]. across [BRCG12, CD92a, CD92b, DQ06, GZXX23, LS05, LM24, PBOR23, Ver90b, YHL22]. act [Ano93a, GT93]. Action [RS06, Sér96, Ano94-44, Ham94, Ken13, Ott03b, TAC96, Fra01a]. Action-based [RS06]. Actions [Cur89, DFFM22, RFB24]. active [CKM13, Fox86a, Kap99, LWP14, New04, Tei22, VLVH23, ZWZ^{+}23]. Activities [HL24, CPP09, CBK03, Dag04, DM91, Fuk09, GC10a, HLO14, JD09, KA15, KTE04, KL24, KFL11, LLB89, LMMPVO09, ML08a, Mar01, MI07, MK96, Rie89, SMW21, Sun96, SD89a, Tam16, Uzu01, VB05, VRS^{+}22, WW21, WH21, dlHBR22]. Activity [GG03, ABS24, Bal10, Clo91, CR08, Kaa09, Kih03, LH14, MROS09, Miy00, MG08, Szá92, Tit90b, WC06b, Ano81a, Gar87]. Activity-based [GG03]. actor [Bre23]. actors [ABGS13, BBTW00, FLYTB24, Hu13, SLA^{+}23a]. acts [JLWS23]. actualizing [JLL10]. Adaptation [LGZ22, BLQS00b, BLQS00a, EKT^{+}22, Ken13, TCMS22, Wal85]. adapted [BBK23, Kra22b]. adaption [Fox86b, GK91, Mud95]. adaption-innovation [Fox86b, GK91, Mud95]. Adaptive [AAK^{+}24, DDOR08, Shi24, CG04a, FC95]. Adaptors [Hay90a, Ric95]. added [JMT03, LDB13, PSG22, Rya96, Soh97b, WJK15]. additionalities [WSF13]. additionality [HST21, HHH09]. additives [MG21]. address [Smi85]. Addressing [DGMR21]. adequate [SIOO01]. ADHD [ASOB23]. adjusting [Wal85]. adopt [Ban04, SU11]. adopters [MTB01]. Adopting [Ano93a, FGS07, GT93, KRK^{+}22, Ket09, SSS17, ZHY22]. Adoption [FRWT06, KRA^{+}22c, AH84, Alc99, AGM89, AAZ22, AR99, BMLL07, BL97, BF21, BR16, BM07, BWS11, dABKS14, CUF13, CF01b, CJS00, CGBV21, DSS05, HO05, JMPR01, JMPR04, JPBL^{+}23, KBC22, KH06, LD05, LKCK22, LLEB05, MRGTDG21, MMC^{+}23, MWC22, Mud90, Mus09, ND15b, OCA09, OOA04, PSdLCJ05, QMP24, RH21, RFMC21, SO02, SSM09, SMD23, TS00, TPBM23, UF22, WNLY06, YT09, ZSX24, ZYSL21, ZS04a, ZS04b]. adoptions [NG09]. adoptive [HM11]. advance [CD00, RGB23, Thu05]. Advanced [ETOA00, HO05, Pap06, Alc99, Che12, ETAO98, Hay88c, HW18, Kim16, LWP14, LD01, LfSS19, MG98b, MG98a, MS98, Sgo95, Sun00, WH04a, ETA02, Par00, VS06]. advancement [KET22]. advances [Gol86, Gol90]. advancing [Sto21]. Advantage [Nuv07, And21, Ano90o, ARR23, BS01, Bar90a, Bes95a, CG11, FCM21, GL22, GCS01, Hun00, Kar93a, Kar93b, Lak08, LXODT15, LH07, NGK08, YGM99, ZL01, Bes03b, Bro90, Har93a]. advantages [Kha97, TCT05]. adversity [SMGPP24, Fra00b]. advice [Lin10b, Lin07a, Lin14]. advisors [MCT06]. Aero [Eri00]. aeronautics [ADNP01]. aerospace [CWBG03, CH01, GTW02, KK05, LL98, MOM08, PV00]. aerospace-challenging [CWBG03]. aesthetic [CS08a]. affect [Ano94m, CJ23, FBBBM22, LWC09, LQ94, LLZ23, Paf04, WSF13, Won05a, vWvRN17]. affecting [HT08, dLP06, LL93a, LL93b, MMC^{+}23, NDP07, OWC03, PS97a, PS97b, WHP85, WWH15]. affects [DLTQ23, ZWZ^{+}23, vRE21]. affiliates [ZMH22]. affiliation [Boa08, GC10a]. affordance [LDMS23, PRID23]. affordances [BKP23]. Africa [AN16, Cha22, Fab96, HB03, KK04, MBA^{+}12, OKvB23, SPO08, Ven14, BH03]. African [BBDM01, LP07, Moo03, OOA04]. after [ND15b, PPMC24, SSP04, Szá94a, Szá94b]. AG [Car98a, OS98]. again [Mow90]. against [CAdNFP04, Lin10c, McB13]. Age [Kra15, ARR23, FN21, MFH16, MT05, Tan87, Wan04, dlHBR22, Arr04]. agencies [SJCT24, CNPV04]. Agency [FCH01, PV00]. agenda [AM23, AGFdMCM24, AHS^{+}23, CKV^{+}23, Da 01, DMRS23, GLÖS23, Hew99, LAC24, MGAP23, OVD21, SWW12, WML^{+}23, ZKW^{+}23, ZAI23]. agendas [CJS00, IH07]. agent [Cha08, LGY05, PPMC24, ZZL22]. agent-based [PPMC24, ZZL22]. agents [Ano94x, Aut94, Car98c, Ved97]. ages [vGS09, Kap03]. AGFA [Car98a]. AGFA/BAYER [Car98a]. agglomeration [OWCvdB15, dSSD09, TKSM08]. aggregation [PPM24]. aggregations [CT97]. Agile [VBA06, AAM24, GTW02, Ket09]. agility [LW21, PRID23, TKYM24, WA04a]. agility-control-nexus [LW21]. agility-mediated [TKYM24]. Agnelli [Agn89]. ago [Hit93]. agreements [CCB20, Cab19, NH99]. agri [ARvT04]. agri-food [ARvT04]. agricultural [KL08a, PLG24, Ree99, RS03]. agro [SO02]. agro-biotechnology [SO02]. ahead [LD02]. AHP [CPXC02]. AI [ZAI23, AHS^{+}23, DBR24, HMFS24, KBC22, LKCK22, MM24, MWC22, PK23, PDHW23, SCM24, TRR^{+}24]. AI-adoption [LKCK22]. AI-embedded [PDHW23]. AI-enabled [HMFS24]. AI-powered [AHS^{+}23, MWC22]. aid [Sma01]. aided [Pur92, Tan87]. Aiding [Mal02]. aimed [YNS18]. aircraft [Eri00, MY02, MTU23, Mow90]. airline [Rue88, Sch92]. Airy [Hay88f]. al [Fra02d]. Alan [Gil96]. Albert [Lae03, Per04, Ros89c, Ano94m]. Alberta [Mer94b, HK02]. alcoholic [IIA97b, IIA97a]. Aldershot [Fit00, Gra03, Har97c, Hay95, Rad97, Tat04]. Alec [Bes02a]. alertness [Fel22, Fel23]. Alex [Par03a]. Alice [OH04]. Aligning [GBV23]. alignment [DNS^{+}03, HKR^{+}22, Mol97b, Mol97a, Pil04, SH02, VM12]. Alison [Szw07]. Alliance [LL04b, OKP13, RG06, TWGB02b, BA10, KM06b, LWC^{+}12, LH21, MZG17, PSS23, Vel15, LL04b]. Alliances [AP08, ALRSD04, Ano94-32, Ano94-37, BCS06, CKR00, CS00, CH01, GM23, HF18, How94, JBL23, KS07, KS04a, LPN22, LWC^{+}12, LH21, NH99, NB92, OM88, PG94, SK23, SJ21, VD04, XR08, Yas05, YI05, ZKW^{+}23, dMD05]. alliancing [LW91]. Allinson [Min94]. allocation [NJ04]. Allworth [Why05]. alone [FBBBM22]. along [CBK03]. alta [Ano94-28]. alternate [Paf04]. Altman [Har93b]. altruism [NK22]. aluminium [HS02]. Alvey [Hob88, Hob90]. always [BZIZ21, GdWNPC03, SCH19, YK08, vWvRN17]. ALZA [Zaf81]. Amazon [Sap10, WT19]. Ambidexterity [FN21, RC23, ASA^{+}22, CYM^{+}16, DTD13, Der14, GMWK23, dVdWNF^{+}10]. Ambidextrous [LH21, YXH21]. ambulatory [BV24]. améliorisation [Ano94-36]. America [Ros89b, HHA84, Ram81]. American [Buc91, Ros89b, BYCS00, FS84, Mas86, May86, MS98, Rue86, ZYSL21, Joh92a]. Amex [Won04]. amidst [WH04c]. Amit [Bul97]. Among [Koc09, CW06, DMW13, FDVB10, GS10, JJS^{+}18, KDH02, KFEC15, MT08, MNL22, Neu12, SLA^{+}23a, SSGN23, SG24, ZCF19]. amongst [BS97]. Amsden [OH04]. Amsterdam [Bro07, Dod87a, Gof07, Leh07]. AMT [ETA02, Par00, BR16, Soh96]. Anaerobic [Ano94a, RGSK94]. analog [WKN03]. analogy [Aab09]. analyse [ML05]. analyses [Bro92]. analysing [Tõn15, WSH23, DJ05]. Analysis [AM10, CLT06, CLZ18, CLZ21, Gal06b, MRGT11, OOKBD14, PSS00, wtcoLSDD91, AH85, AM01, AT99, AT93, AW12, Ano93j, Ano93i, Ano93p, Ano94p, AS18, AdA08, BMV02, BH16, BF03, BF05, BG10, BBA15, BMLL07, BCV21, BCC^{+}11, BNV05a, BLLEM23, CA01, CT12, Cav16, CS04, CS05, CYM^{+}16, CYJ07, CCY05, CY00, CbLjJL12, CR02, Col81, Dav93, DID21, DHT22, DTI^{+}23, DPH^{+}22, DGS23, DGR23, DLR07, DB19, EKMS12, Fel23, FC06, FRWT06, FMV00, FCS21, FR15, FGS07, GKK23, Gra24, GKB12, Gup89, HBS07, HS14, Han94, HPB03, Har97c, HL24, Hit04, Hit06, Ho09, Hos00, HS19, HT08, HKJ07, IKL01, IM10, JLY23, JWL24, JFCFMC^{+}12, KTSM10, KZ18, KT02, KHKL23, KBS01, KK07, Kum05, Kum06, Lay93, LS07, LJAK21, Lev93, LB87, LFU23, LL05a]. analysis [MKK04, ML08b, MRGTDG21, MW14, Mat22, MOWGB02, MLC24, MRC14, Miy96, Miy00, MI07, MR00, MYR^{+}21, MII90, Nay88, NLC98, NB92, Nwo02, OS08, Pet98, Pet11, PMN^{+}24, Phi99, Pri17, RDAV23, RG10, Red14, RGJ13, RPASLM23, SGB12, SRDS06, SLA^{+}23a, SF05, SOA^{+}24, Sna01, STS88, SKL17, SH10b, TW02, TWGB02a, TKSM08, TRR^{+}24, TDC11, TDA05, Tsu02, UBCF19, Vit90, WTAW97, Wan24, WMGB01, WNGB03, WN03, WH04c, WLO09, WWYZ23, Wes93a, Wes93b, WvdVM07, Wil93, WJK15, WHB21, XLC23, YGM99, YW08, dV94, lS06, NW23]. analytic [AM92a, IMBH98, ODM23]. analytical [AM92b, SS22]. analytics [BCPP24, CZL^{+}24, GLÖS23, JPBL^{+}23, PMT23]. analyze [CS06, DGP10]. Analyzing [AW12, Lan09, LH14, BC99, JY15, KGS^{+}24, PG11]. anatomy [Win87]. ancestry [OGOF22]. anchors [BCM12]. ancient [Guy96b]. André [Ano13h]. Andrea [Edw04]. Andrew [Dod88, Edw04, Gra03, Sod08, Tra00]. Andy [Sap10]. angel [ES05, LCRL22, Ano94-27]. Angelo [Gra01]. angels [HSLS24]. Anglo [TI02]. Anglo-Japanese [TI02]. Anmol [Kaz03]. Anne [Fra03b, Mur91]. Annie [Coo97]. Announcement [Ano83a, Ano91a, Ano97a, Ano08a, Ano85e, Ano87i]. Announcements [Ano85a, Ano90b]. Anomaly [CZL^{+}24]. answer [Bal10]. Antecedents [FGTB23, SVV15, TWBC21, WWY15, MWC22, RB09, SO20, VB05, BC12, EH14]. anthropomorphism [QMP24]. anti [LL22, TW02]. anti-dumping [LL22]. anti-hypertensive [TW02]. anticipate [SB02]. Anticipating [KKL19, BTP^{+}23, FHBDO16]. Anticipation [QMP24]. antique [CP11]. Antonelli [Eva02]. Antonio [Gra01]. Anwendung [Ano94j]. anyway [BM16a]. app [CP23, Kan22]. apparel [Moo03]. appeal [BNV05b]. appendices [Kaz03]. Apple [Kan22, dHOvdKS16]. applicable [YH11]. applicaciones [Ano94t]. Application [LD01, Rue88, Sto99, Ano94v, Ano94j, Arc88, AFL^{+}24, CCF23, CL24, FJ12, GTW02, HPB03, JWS92a, JWS92b, KGS^{+}24, LLI06, LL04a, ML08a, MV06, MFH03, PWF07, PP04, PFP06, RUNS23, SCH93a, SCH93b, SCFD18, SK92, WKN03, vRGJ^{+}12, AW12]. Applications [SCM24, Ano94t, Cha94, CLC13, CGS^{+}24, FB93, Fun06, GKK23, Gup90, IAA94b, IAA94a, LYY^{+}22, hL05, OM18, TFS^{+}22, Fra02d]. Applied [HG92, AO10, BCV21, Cha08, KS10a, Nob04, SG22, Smi85, Ved81, VM97]. Applying [KKP17, NVMD22, Kuc07]. appraisal [HKJ07, OC85, Tho88b, Tho88c]. Approach [Fra01b, AAM24, ANI^{+}22, Bar95, Bar92b, BCM^{+}22, BCPP23, BSMC23, BMRG00, CBF24, Car04, CHT99, CJH^{+}07, CG11, CLZ^{+}22, Chi21, Clo91, Dow93a, Dow93b, EEY22, EM85, FSA23, FRWT06, FMV00, GLYP13, GHHH22, Gra24, Gre06, HKW^{+}22, HW01, HW02, II92a, II92b, KS21, KB09, KA06, LB21, LYP09, LJAK21, LKKH22, LLL23, LW21, Lin86, MG21, MAMRRR24, MAMD22, Meh88, MAKKM22, MG08, MCT06, MBA^{+}12, NPG17, NSYL18, Nwo02, OR06, OBMOM24, PGKF19, PPMC24, Pra86, SMvFM24, Sgo95, SSS17, Sma01, Sou83b, SCS06, TCC^{+}22, Ven96, Wes93a, Wes93b, YJLL20, YJLL21, YLH22, ZZL22, Szw07, Par03b]. approaches [Gil96, Lin11e, TAM11, Wec15]. Appropriability [GÁNA07, DL10, HKJ07, SCY17, YHL22]. appropriating [RHL09]. Appropriation [Sha86, BBZ20, Cab19, FMF21, HLH12, MPV13, Neu12, ZLH17]. April [Ano21a, Ano22a, Ano23a, Ano24a]. Arabia [ASGC09]. Aragon [ST95, SP98a, SP98b]. Aragonese [PS00a]. arbuscular [SO02]. Arch [Gof07]. Archer [Pot93]. archetype [MPO23]. archetypes [MRGTDG21]. Archibugi [Hay01a, Hay01c, Pap04]. Architecting [CvZ05]. architectural [HKRR20]. architecture [JSP22, LGY05, LCLC03, PPM24]. architectures [BWF18, ZDWG14]. area [BS01, CT12, Die00]. areas [BS08, CP11, ETM96, LS07, RS03]. arena [Car98a]. Argentina [AL11]. Argentine [MPV13]. arising [Whi89a, Whi90a, Whi90b, Whi90c]. Armistead [Del97]. Arnold [Hay97a, Lim06]. arrangements [BK89, CH01, HB89]. Art [Hay86c, BB95, CM94, Hit89b, Hui11, Lin02, NF24]. Arthur [Ano94b]. Article [Ano12a, Hit93]. artifacts [DAR07]. Artificial [CM96, MMMD23, PK23, ZWD09, ABvZ23, CDED21, Cha94, CL24, DDFS23, ZAI23]. ASEAN [Ano94-30, JL98, SS94, WC07]. Asia [Won08, Hob95, HCK01, Hos99, Hui17, Min95, Won13, Kap01]. Asia-Innovation [Won08]. Asia-The [Won13]. Asian [BWF18, Kim87]. ASIC [CD92a, CD92b]. Asking [ABGS13]. aspect [CL98, LSB07]. aspects [Bar96, Gup90, dAOCM02, SL98b, Tho88b]. aspirant [ZYZZ24]. assembled [Lag17, LS23]. assertiveness [vWvRN17]. assess [CCF23, SPLB19]. Assessing [AC08, BRV23, BDNP06, BC92, CL05a, JH18, LO10, Mow88, Soh96, BAWK12, McG98]. Assessment [DB05, FOL^{+}03, FJ12, KET22, OTA01, VLVH23, AM01, AKKS12, Bes02d, BLQS00b, CMA07, CN90, CN92, FAH08, GG05, Hay85b, IMIY87, Jol08, Lan98b, MII90, MBA^{+}12, OIN96, PR96, PR97, PMP24, PPP98a, PPP98b, PdW00, RMGDLC08, Sap89, SH08, SD23, TOS^{+}00, Van10, ZWD09, dVdWNF^{+}10]. assessments [NVMD22]. asset [FN21, YNS18]. asset-creation [YNS18]. assets [ASA^{+}22, AW11, GBM98a, GBM98b, KMID11, LT09, SSF12, SM95, SS07]. Assimakopoulos [Giu09]. assimilation [WA03, WAN03]. assist [OC85]. assistance [SBMP13]. assistants [Hay97c]. Assisted [BP01]. assisting [SC10]. associated [BWS11]. Associates [Hay88c]. associating [Sna98]. Association [Hay93c, Hay87d, Hay88e, Hay89a, Smi85]. associations [Ros89b]. assumptions [CWBG03]. assurance [Red14]. Aston [Hay93c, Guy96a]. asymmetric [AMP22, HF18]. Athens [FLKF^{+}06]. Atipol [Ano94-30]. attachés [Van85]. attempt [RSP10]. attempting [Wit85]. attendant [II98]. attention [Lin12d]. attitudes [dABKS14, CDED21, CHD18, CLC13, MGGOS21, Cur89]. attitudinal [TDA05]. ATTRACT [WPR^{+}22]. attraction [SvG15]. attractiveness [MP09a]. attributed [MCMA09]. attributes [CW22, GBV23, HMM14, MK96]. Atul [Ano94-37]. auction [Cha08]. audit [LL04a, LTT23]. Auditing [CO04, Mer94a]. Audretsch [Dib00]. Augsdorfer [Har97c]. August [Ano21b, Ano21c, Ano22b, Ano23b]. Austin [Hay97c]. Australasian [Orr97, ZS04a, ZS04b]. Australia [RAS15, Soh97b, SPP98, STZ03, TNSG96, TSH02]. Australian [Phi99, BS97, Bea98, Bea05, BS06, MB04, NC23, PS00b, Rya90, Rya96, Sha03, SGFS99, SPS01, TS00, Wan89, Wan91, Wit85]. Austria [KT02]. Austrian [WSF13]. authentic [LCWC09]. Author [Ano89a, Ano90a, Ano91b, Ano02a, Ano03a, Ano04a, Ano95n, Ano01a, Ano06a]. Authoritarianism [SRLR23]. Authority [PR97]. authors [Ano89r, Ano92e, Ano93g, Ano94o, Ano96n, Ano97p, Lin14]. auto [KK04]. automakers [AS18]. Automation [Ano94c, Hit94, Orr97, Alc99, Gol86, GPR89, Hay88f, KKRM05, San95, Tan95, vDvM89]. automobile [Kha97, Kha98d, Kha98c, KS10b, LVHR10, MY02]. automobiles [Kha98d, Kha98c]. Automotive [Ame95, AAZ22, BWF18, Hel94a, Hel94b, HW18, IKL01, LPS03, LG22, Nob04, PS00a, PDHW23, Thu18]. autonomous [MWC22, TSPJ23]. autonomously [RWP19]. Available [Rus94, Kra15, LOCSBGMH16]. Avatar [KMF09]. Avatar-based [KMF09]. Avebury [Har97c, Rad97]. Avenue [Per97b]. Avgerou [De 04]. aviation [Gee06]. avoid [Lin14]. awake [Sch09]. Award [AHAAEA03, Ugh88, vAdMSC94b, vAdMSC94a]. aware [IPP14]. awareness [Ber96, SY14, dBJ02]. AY [Rus94]. Aydalot [Hay91a].

B [Ada07, Ano81a, Bes94c, Bes02a, Dib00, Hay87d, Hay88e, Pap04, LLEB05]. B-42 [Ada07]. B-to-B [LLEB05]. B. [Wya07]. B.-A [Pap04]. B.-K. [Wya07]. B2B [RKFR24, TAM11]. B2C [DSS05, SH06, WPF06b]. B7 [Hay93c]. babelizing [Lin09c]. Baby [MWC22]. back [Hel94a, Hel94b, RGMW18]. backfires [QMP24]. background [JL96, MCA^{+}24]. backgrounds [ZWDZ21]. backward [CD00]. backwards [O’K19]. bad [CL05a, Lin20]. Badaracco [Dav92]. Baffins [Gri98, Per95, Pov97]. bag [Ano94b]. Bahrain [Pal00]. Bainbridge [Gri98]. Baker [Hay89d, Hay97b, Her92c]. Bakkers [Pap02a]. balance [Sir91, SvG15]. Balanced [WFW02, AAT14, FRW06]. Balancing [KL08a, SDT^{+}22, BHS05, Sun96, YXH21]. Baldwin [Cre04]. Ball [Ano93f]. Ballmoor [Sha95]. Bancroft [Bes94b]. Band [Gil97]. Bangalore [Sub05]. Banji [Coo07]. Bank [Kuc06, Won07]. banking [ASGC09, AFL^{+}24, EN07, GC99, Knu86, May86, UOIA00, YPF03]. banks [GCS01]. Banta [KS04a]. Baranson [Hay88f]. Barbara [Aug06a]. Barcelona [Die00, MP09a]. Bargaining [BA10, Rad99]. barrier [Bad11, BG19, GL10, Lic11, Lin10a, NHV05, dB10, vZS14, vdVdM11, vH10, vK11]. Barriers [Ayr88, Had99, TNSG96, UEGE^{+}14, GBFvdS11, JPBL^{+}23, KH06, KRA^{+}22c, PSdLC02, SDN23]. Barry [Ano94f]. Bartlett [Ver90b]. Base [AAAA23, AS22, CMtW10, Paw94, Sik94, FGOX23, Hal14, HMM14, OKvB23, Se14]. Based [Fra01b, AAT14, AMS^{+}24, And21, Ano94x, ABUS05, AVA22, Aut94, ALCM14, BG10, BKM10, BPU21, BS06, Bha05, BS99, BLLEM23, Bul85, CS08a, Can16, CK22, CBF24, CKR00, CF01a, CLZ18, CLZ21, xCZZ^{+}22, CX24, CPXC02, CTC02, Clo91, CHO98, CFC12, DGP10, DVdCAS16, DVSL19, DLL06, Dre97, FSA23, FMV00, FC95, Fre86, FGOX23, Fun06, Gal05, Gar96, GLYP13, GBFvdS11, GARM24, Gre06, GKB12, GG03, HV96, HSLS24, HESPSL11, HY04, HLC05, IAO^{+}02, IMIY87, II92a, II92b, JBL23, JLY23, JY15, JWY23, Kes03, KJE96, Kod05, Kod06, Kod07a, KMF09, KK87, KB06, LVS10, LVVH23, Lag17, LS05, Lee00, LS07, LYP09, LCLC03, LL08b, LL05a, LL05b, LOCSBGMH16, LOM14, MLG12, MRGT11, MRGTDG21, MYR^{+}21, MII90, NK22, NJ04, NP10, OSG04, OR06, PGLS21, Pap03, Pfi99]. based [PPMC24, Pur91, RS06, RMGDLC08, RZ21, Sae05, SC17, SCS22, SPÁG22, Sch94, SKTM08, DGMR21, Sma01, SJ16, STZ03, SKL17, Ste89, Sun96, TCMS22, TLL07, Tsu02, VM12, WCT10, WO08, XW21, Yas05, YJLL20, YJLL21, YHDR08, ZIB23, ZL01, ZWZ^{+}23, ZZL22, dV94, dlHBR22, lS06, VP14, Mer02]. Basic [Ano93b, CW93, BBA15, CLP23, CL22, Ham96, HV96, NB09]. Basil [Bro90, Cou90, Cou91, Dod88, Her91, Sha91a, Ver90a, Ver91a]. Basin [dSSD09]. Basingstoke [Coo07]. basis [Lin17a]. Bass [Fra01a, Fra01c, Hay88a, Hay88b, Hay88c, Mer01]. Bastow [Tan08]. batteries [McG98]. battle [Gal12]. battles [Gal12, dHOvdKS16]. BAYER [Car98a]. be [DR99, LVS10, Lin11c, Rei97]. Beach [Ano07k]. beacon [PS14]. beam [WMGB01]. bean [AT99]. Beast [Lin17c]. Beauty [Lin17c]. Beckers [Bec89]. become [HdWN21, OWCvdB15]. Becoming [HKY22]. Bedford [BB95, Bes94a, Bes95c, Mer94a]. Beer [Per95, llo91]. before [BEE24, FD21, KR12, Szá94a, Szá94b]. begins [KFL11]. Behavior [HW17, WNGB03, BG10, CDD^{+}11, CDD^{+}14, KWvdS21, LC07, LS16, wtcoLSDD91, SGWL16, Tsu02, WKOW03]. behavioral [Ano93n, CDED21, Pau93]. behaviors [CH04, HM11, Tsa13b]. behaviour [BCG01, MII90, San07, TP24]. behavioural [WSF13]. behaviours [AL11]. behind [BMW14, XX21, ZIAO^{+}21]. Being [JBZ23, Whi94]. Belarus [AHP16]. Belgian [Van85]. Belgium [Ano81a, Hay87b, LK20, TB24]. beliefs [BKA^{+}23]. Bell [Mow88]. Ben [Lae04, RA86]. benchmarking [ET98, LL98, Pov97]. Benedikt [Hol93]. beneficial [SP22b]. benefit [CZZM21, Gui11, MCH00, Sha86]. benefiting [VH06]. Benefits [DSS05, KA06, PSdLC02, BAWK12, BWS11, GKK23, IPP13, JMPR04, Joh08, KH06, ZC19, ZS04a, FCH01]. Bentley [For85]. Beobachtungen [Ano94g]. Berger [Gra03]. Berlin [Hop04]. Bessant [Hay97a, Why02]. Best [Bea05, Bes03b, RGJ13, SK23, TSH02, BN08, CvZ05, CL97b, CL97a, CO04, Guy95, Hal96, Hay97a, RMGBA08, SW04, vAdMSC94a, vAdMSC94b, Fit00]. bets [LTVA22]. Better [Ugh88, BZIZ21, GdWNPC03, Kra22b, Rus94, ZLJ^{+}21, dGSJ^{+}21]. Bettley [Szw07]. Between [RM09, AAGL23, Ahn95, AO10, AMP22, ALCLdF^{+}15, AC04, AEP21, BT22, BWF18, BF06, BOBT22, Cai23, Car99a, Che12, CL97b, CL97a, CK12b, Cor87a, CD00, Cur89, DTD13, DLL06, FMFMA16, FD21, GBFvdS11, GF21, GFvT22, GMWK23, Ham96, HWC06, HW01, JL96, JW06, KOX21, Koc04, KLP24, KBRR24, KS10a, KRM^{+}24, LW04, LKKH22, LD95, hLFL08, Lin10b, LCT04, LCT05, LR09, MTB01, MP09a, MM08, MLC24, NMN11, Nar89, ND15a, Neu12, O’K19, OC91, OM89, OW00b, Ott98b, OIN96, RV21, Ran24, RF12, Rie89, Sae05, SPÁG22, SR23, SH02, SM12b, TNSG96, TCT05, TKT13, Ved97, VTRM12, VNP^{+}22, VH06, WWM00, WA03, WAN03, WH04b, WA04b, XHM^{+}16, XLC23, YGM99, YI05, YMMYK24, ZXT10, ZC19, ZC22, ZXDG19, dlHBR22]. Betz [Dod87b]. beverages [IIA97b, IIA97a]. Beverley [Szw07]. Beyond [Boa08, JC01, KLL24, Phi99, Wat90, BBLW24, BLNP05, Bou15, DJ13, HCMMH21, JMT03, MSR10, NWGB09, NGK08, Rea21, Del97, OH04, Per95]. Bhanich [Ano94-30]. bias [CMB23, ET98]. biased [NVMD22]. bibliography [Kap03, Mer94b]. bibliometric [ADDS09, BJ05, CYJ07, DGR23, FHA22, Ho09, PK23, RPASLM23]. bibliometrics [HS14, KHKL23]. Biesbosch [Lan09]. Big [PDHW23, RFB24, ACBO23, BCPP24, BM05, CBGvdV22, GUMZI14, KLL24, Lin08b, MMCD23, PMT23, RvdSdE22, VH06, WPR^{+}22, JPBL^{+}23, LYST22, SP22a, SPPG20, UBCF19]. big-science [VH06]. bike [RHW14]. bio [BCC^{+}11, HMM15]. bio-pharmaceutical [BCC^{+}11, HMM15]. Biochemicals [Lid90]. biodiesel [MBA^{+}12]. biogas [AIOA00, RG10]. Biology [Fra01c, HMFS24, RL06]. biomass [HO14]. biomedical [JFCFMC^{+}12, OBvE97, Rep08, Zaf81]. biopharmaceutical [PEER23, VJ23a, XR08]. biosensor [HS14]. BioTech [PvdMB^{+}06, GEG12, MS19, TA10]. Biotechnologien [Ano94j]. biotechnologies [Ano94v, Ano94j, IAA94b, IAA94a, Ano94t]. Biotechnology [AA05, AAN06, AR99, BC12, BS06, BNV05a, Boa08, BA10, CAdNFP04, CC06, FMF21, Fon05, HBS02, HBS07, HN92a, HN92b, HST05, IH07, Jai23, KP12, Mar03, MT05, Meh01, MS99, NPV05, Pfi99, QGBV04, SO02, Sap89, Sen96, Sha87, TM06, WNM95, WvdVM07, WH96, vdVMM09]. biotecnologías [Ano94t]. Bird [Cou90, Cou91]. Birds [HW17]. Birmingham [Hay93c]. Biswas [Bas04]. black [BPU21, BM14, CS01b, KJ97, LM14]. Blackburn [Aug06b, Per97a]. blacksmith [BO81]. Blackwell [Bro90, Cou90, Cou91, Dod88, Her91, Sha91a, Ver90a, Ver91a]. Blazing [URG^{+}24]. Blockchain [SMI23, ABSZ23, CX24, KRA^{+}22c, Mas23, RIMP22, RPASLM23, ZXX23, CCR^{+}23]. blockchain-enabled [CX24, ZXX23]. blue [Gal12]. Bo [Gra01]. Board [Ano11g, Ano18b, Ano18c, Hay89c, AABY^{+}23, Ano88j, WW14, Ano81c, Ano83b, Ano85c, Ano86b, Ano86c, Ano87c, Ano87d, Ano88f, Ano89n, Ano90j, Ano91f, Ano92d, Ano93d, Ano94h, Ano95m, Ano96m, Ano97o, Ano02d, Ano02c, Ano02e, Ano02f, Ano03d, Ano03m, Ano03n, Ano03o, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano04c, Ano04l, Ano04m, Ano04n, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano05a, Ano05j, Ano05k, Ano05l, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano06b, Ano06j, Ano06k, Ano06l, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i, Ano07a, Ano07h, Ano07i, Ano07j, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano08b, Ano08j, Ano08k, Ano08l]. Board [Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano09a, Ano09i, Ano09j, Ano09k, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano10a, Ano10h, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano11b, Ano11i, Ano11j, Ano11c, Ano11d, Ano11e, Ano11f, Ano11h, Ano12c, Ano12j, Ano12k, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano13a, Ano13f, Ano13g, Ano13b, Ano13c, Ano13d, Ano13e, Ano14a, Ano14i, Ano14j, Ano14k, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano18a, Ano18d, Ano18e, Ano19a, Ano19b, Ano19c, Ano19d]. Board [Ano19e, Ano19f, Ano19g, Ano20b, Ano20c, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano21p, Ano21q, Ano21r, Ano22m, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g]. boards [Bla16]. Bodies [Wya00, Cor87c]. Body [YKW10]. Body-of-Knowledge [YKW10]. Boeing [SS16]. Boekema [Pap02a]. BoK [YKW10]. Bond [FD21]. Bone [Joh92e]. Book [Ada07, Ano81a, Ano81b, Ano91d, Ano91c, Ano94d, Ano95a, Ano96a, Ano97b, Ano98l, Ano98k, Ano98f, Ano98g, Ano98h, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano98i, Ano98j, Ano99a, Ano99b, Ano99c, Ano02g, Ano03b, Ano04b, Arc03, Arr04, Aug09, Aug06a, Aug06b, Bai84, Ban04, Bas04, BB95, Bes88, Bes93, Bes94a, Bes94b, Bes94c, Bes94d, Bes95a, Bes95b, Bes95c, Bes96, Bes97, Bes01a, Bes01b, Bes02a, Bes02b, Bes02c, Bes02d, Bes03a, Bes03b, Bon02, Box01, BH03, Bro90, Bro07, Buc91, Bul97, But96, Col97, CF07, Col08, Coo07, Coo97, Cou90, Cou91, Cre04, Dav92, De 04, Del97, Dib00, Dib01, Dib02a, Dib02b, Dib04, Dod87a, Dod87b, Dod88, Dre04, Dre06, Dut07, Edw04, Eva02, Fit00, For90, For91, For85, Fra99]. Book [Fra00b, Fra00a, Fra01a, Fra01c, Fra01b, Fra02b, Fra02c, Fra02d, Fra02a, Fra03a, Fra03b, Gar85, Gar87, Gar94, Gar95, Gar96, Gil96, Gil97, Giu09, Gof07, Gra01, Gra03, Gra04, Gri98, Hal97a, Har93a, Har93b, Har97c, Hay83, Hay84, Hay85a, Hay85d, Hay85b, Hay85c, Hay86c, Hay86d, Hay86e, Hay86a, Hay86f, Hay87a, Hay87d, Hay87b, Hay87c, Hay88a, Hay88b, Hay88c, Hay88d, Hay88g, Hay88f, Hay88e, Hay89a, Hay89d, Hay90a, Hay90b, Hay90c, Hay90d, Hay91a, Hay92b, Hay92a, Hay92c, Hay92d, Hay92e, Hay93d, Hay93a, Hay93b, Hay93c, Hay95, Hay96a, Hay97b, Hay97c, Hay97d, Hay97a, Hay01a, Hay01b, Hay01c, Her91, Her92a, Her92b, Her92c, Her04, Hob87, Hob97, Hob03, Hol93, Hop04, Jel07, Joh92a, Joh92b, Joh92f, Joh92c, Joh92d, Joh92e, Joh92g, Kap01, Kap03, Kaz03]. Book [Koc09, Kra15, Kuc06, Kuc07, Lae03, Lae04, Lal00, Leh07, Lim06, Lin90, Lin91, Lop06, Mer94a, Mer94b, Mer01, Mer02, Mil85, Mil88, Min94, Min96, Mor91, Mor04, Mur91, New04, Nuv07, O’H99, Oak97, OH04, Pap02a, Pap02b, Pap03, Pap04, Pap06, Par03a, Par06, Par07, Par03b, Per04, Per95, Per96, Per97b, Per97a, Pot93, Pov97, Poy97, Pra10, Pre01, Que06, Rad97, Ram86, Rea06, Ric95, RL06, Ros89c, Ros89a, Ros89b, Ros90, Rot87a, Rot88, Rus94, Saa98, SW04, Sap01, Sap05, Sap10, Sco89, Sha95, Sha91a, Sha91b, Sod08, Szw07, Tan08, Tat04, Tra00, Ver90a, Ver90b, Ver91a, Ver91b, Wan04, Whi94, Why02, Why05, Wil90, Won08, Wya00, Wya04, Wya07, dC07, Ano92f, Ano93h, Ano01b]. Books [For90, For91, Gar87, Joh92d, Lin90, Lin91, Ver90b]. bookstores [Kra15]. boost [Bes03a]. boosting [SBOB23]. bootlegging [Har97c]. BoP [AS22, RM14]. border [CM95]. borders [CYHC19, Ver90b]. Borja [Why05]. born [ARR23]. born-international [ARR23]. Bosch [LDK22]. Boston [Bes95a, Bes96, Bes01b, Dav92, Hay97d, Hay97a, Joh92a, Per04, SW04, Ano94-28, Ano94-29, Ano94w, Bes15, dCLSNL08, Töd94b, Töd94a]. Boston/London/Dordrecht [Hay97d]. Both [Bes94a, ABGS13, FCCE21, GGJ22]. bottlenecks [Sch85]. bottom [BK04, Bøl12, PRA22]. bottom-up [BK04, Bøl12]. Boucke [Ano94-38]. Boulder [For85, Hob87]. Boundaries [BK15, FLYTB24, PCQ24, SPMRS23, JBL23]. Boundary [CPT23, LPN22, BC10, CLP23, CE19, Kod07b, SJCT24, ZMJM23]. boundary-spanning [BC10, ZMJM23]. Boundaryless [Ano98l]. bounded [MAKKM22]. Bower [Bes93]. box [BPU21, BM14, CS01b, LM14]. boxes [SM05]. Bozeman [Ano94f]. BPR [Del97]. Braa [Gra01]. Brabantia [dV94]. Bradbury [Mor91, Sco89]. Brand [Why05, MGdG11, WF10]. Brazil [KdOL08, PP09, dSSD09, SFBK23, WAdN19, vAdMSL95, Que06]. Brazilian [Ano93n, Cam93, FJ12, JJS^{+}18, Pau93, vAdMSC94a, vAdMSC94b, dMvAdMSPB93]. Breadth [BC10, MAMRRR24]. Breadth-of-impact [BC10]. break [Pal06]. break-even [Pal06]. breakthrough [FGTB23, RM14, Aug06a]. breakthroughs [Ayr88]. breast [CAdNFP04]. breed [HSLS24]. Breeding [Liy95]. brewing [llo91]. Brian [Per97b, Rus94]. bricks [YWL22]. bricolage [FB06]. bridge [SCY22, TNY^{+}22]. Bridges [GM23]. Bridging [CJ98, CL24, Hal97a, HBO14, LZLL21, SOKE15, MH05, Cai23, PPM24]. Brief [Hit89a]. briefs [HBH21]. bright [DSPS23]. Brighton [Gar87, Hal97a, Kuc06]. Brigitte [Why05]. Brilliant [Kra15]. Bringing [RGMW18, SR22, Sha95]. Britain [AH84, Seg86, Tet00, WNM95]. British [Ano81b, Hay89a, Hay93b, Bac98, BYCS00, BBHW93, Har96, Mas86, RG85, Sur92, VOTH05, Wal85, Whe86, Woo87]. Britton [Hay88b]. broad [Woo87, vdWFvdB23]. broadband [CA01, CCY05, NLLC04]. broadcasting [WKN03]. broader [WHB21]. Brockhoff [Ros89a]. Brooking [Coo97]. Brown [Ano88h, Hay01b]. Bruce [Par06, Why02, Ano11k, HKW^{+}13, New13]. Bruxelles [Ano81a]. Brynjolfsson [Kra15]. BTP [RGJ13]. bubbles [Kap03]. Buckingham [Sha95]. Budapest [Pál04]. Buddy [Li05]. Buddy-finding [Li05]. budget [CG05, YWC09]. Bueno [Ano93l]. Build [Why05, Mol98]. build-up [Mol98]. Building [Chi21, Chu02, GHU21, HdWN21, SLA^{+}23a, Sch94, SCK10, SCK11, Bas01, DDG98, Gra03, HO14, LHI13, LWM24, NKL02, WCZZ14, Bes03a, Bul97, Fra00a]. Buildings [Hay93b]. built [dVV16]. bullet [BF23]. Burroughs [BN88]. buses [SM12b]. Business [Aug09, BCPP23, BB95, Bes94a, Bes95a, Bes95c, Bes96, Bes97, Bes01a, Bes01b, BNV05a, CWY14, Coo07, Dav92, Edw04, GG05, Hay97a, HS21a, IKL01, Kuc06, KS04b, LDF^{+}21, LNB21, Lin12b, MLF16, Mrk17, OS99, OW93a, OW93b, RAS15, Saa98, SW04, Sod08, Vel15, Ver90b, WZNG17, vDZ21, ACBO23, AMV07, AMS^{+}24, ABSZ23, Ano93n, AHP16, AC04, ASOB23, ABAHP^{+}23, BBT19, BN06, BH16, BS97, Bes94c, BMSJ24, Bøl12, Bro92, Bro94, BRCG12, CGM07, CK22, CvZ05, Car01, CW22, CP07, Dea81, DBL07, Del97, DHT22, Des96, DSS05, EKK^{+}21, EFC23, FB06, FMOPSVA21, FJ12, Fra02a, FGOX23, Gal06b, GMNL07, GGS08, GUMZI14, Gre06, Gri98, Guy97, HBS02, HSLL22, Han06, Han95, HPB03, HJM08, HLL12, HKY22, II98, INO^{+}03, IH07, JWL24, JL13, Jon05, KBO22]. business [KAV17, Kod03, Kod05, KLP24, KGK07, KZ03, LO10, LACH13, Lee23, LCRL22, LDK22, Li09, Li21, LL08a, Lin13b, Lin13e, LW91, LL03, LLZ19, LP06, MCH00, ML08b, Mas23, MKGL05, MRDA24, Moo03, NPV05, Obe88, Öbe23, OKvB23, OB21, Pal00, Pau93, POGBG06, PPP98a, PPP98b, PG11, Ran24, RH10, RAC84, Ros89b, RI04, SM09, SVLYT24, SB02, SH08, SPPG20, Sha91b, Shi24, DGMR21, Soh99, SL16, STH06, SH02, SG24, SSP04, Swi92a, Swi92b, TSPJ23, TCMS22, TCT05, TE14, TPBM23, VM12, VvWM20, VNJ18, WCZZ14, Why02, WvdVM07, YI05, YCL22, YT09, YXH21, ZQW^{+}19, ZMH22, ZXDG19, ZMJM23, dAASA21, CAP23, Kuc07, VNP^{+}22, Fra01c, Gil97]. business-level [YT09]. business-planning [LP06]. Business-to-business [IKL01]. business-to-government [JWL24]. businesses [GHHH22, Kar86, Lic05, MCT06, SAM^{+}23, SBMP13]. butter [Wat90]. Butterworth [Bes02c, Hay92d, Sap01]. Butterworth-Heinemann [Sap01]. Butterworths [Hay85c, Hay88d, Rot88]. buying [WPF06a].

C [Ano81b, Ano91d, Ano97b, Bes02d, De 04, Fra03a, Hay85d, Hay86d, Hay88e, Hay97d, Koc09, Mur91, Par07, Pov97, Ros90, Rus94, Sap10]. CA [Bes03a, Box01]. Cabibbo [Cab89]. cacao [Ano94v, Ano94t]. CAD [DC00]. cadre [Ano94-36]. Caird [Ano94n]. Cal [Fra01c]. calculation [Sto99]. Calendar [Ano92a, Ano88a, Ano89b, Ano89c, Ano90c, Ano90d, Ano90e, Ano90f, Ano90g, Ano90h, Ano90i, Ano91e, Ano95b, Ano95k, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95i, Ano95j, Ano96b, Ano96k, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano96i, Ano96j, Ano97c, Ano97l, Ano97m, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano98m]. Calibration [Sna90d]. Call [Ano84a, Ano12b, Ano85e, Lin13c]. Calling [RH23]. Cambridge [Cre04, Dut07, Hay88g, Hay01a, Hay01c, Hol93, Kap01, Nuv07, RL06, Ros90, Sod08, Ano94b, Lid90, Seg86, Sod08]. came [Lin11c]. Campania [Ano92c, CT92]. Campos [dSSD09]. Campus [KM97, Per97b, Bes93, Guy96b]. Can [BSG^{+}24, MP09a, OWCvdB15, RM14, WW21, YZY10, ACBO23, Bes01a, DR99, Dre01, HS02, JP96, Lin20, MCD06, MWC22, PMP24, SAWT06, SCY22, SK11, VNP^{+}24]. Canada [Bai84, Mer94b, AL05, Ano94-30, Ano07m, AR99, BS01, BBA15, Ras08, SDFN01, SS94, WNM95]. Canadian [Cre04, Ano94p, HBS02, Han94, Hay81, KMK03, LB03, LT05, NB92, wtcoLSDD91, TM03, TDC11, WH96]. Canberra [Hay85a]. cancer [CAdNFP04]. candidate [MAKKM22, SBF15]. candidates [YH11]. CANDU [Lee04]. CANDU-type [Lee04]. cannibalize [NHV05]. Canon [MOWGB02, WLO09]. can’t [vWvRN17]. Cantner [Ano94-38]. Cantor [Ano91d]. capabilities [ABSZ23, AAZ22, Arc88, AVA22, Bac98, BF06, CKT04, CCC^{+}12, CCMW23, CNN23, Chi21, DDN24, ES21, FN21, FSA23, FCM21, GMNL07, GHU21, GMWK23, HBvdB09, HKY22, HDE21, JRS06, KUH21, LL98, LT21, PKH15, Pet98, RBBF23, RGH19, SPO08, SR23, SY05, ST21, SSRO24, SD89a, TBSM21, XLG^{+}23, YXH21, dAASA21]. Capability [JBZ23, LHI13, Se14, SCM11, TC97, AKBA07, AHS^{+}23, AC08, AS18, CPP09, CCGK23, EB23, EWJ09, EEY22, Fan06, FCM21, Fig02, FB05, FdlCU22, GSP^{+}22, GHHH22, GM03, HO14, JvZ08, JL98, Kha99, KLL87, LSB07, LCT04, LLY^{+}18, LH19, MKK04, MRGT11, PR96, PR97, POGBG06, QGBV04, Ran24, RRB^{+}22, SGFS99, Tik23, VLIC12, hWyLbC08, WCZZ14, WH21, XR08, ZIB23, wK98]. capability-accumulation [Fig02]. capability-based [MRGT11]. capable [Wal85]. capacities [ALMW21, OIOA01]. Capacity [CS98, MS19, BBS22, Cab19, CJL22, DLTQ23, EH14, EHHW17, FMFMA16, dCHDAALMM07, HLO14, JBGMM11, KC07, LS05, LS04, LWC^{+}12, LP07, MJ24, MP09b, NQ05, PS21, PMT23, PA15, RV21, SC17, SBP21, SS23, SCK10, SCK11, WA03, WAN03, WSH23, XR08, Yam21, YMMYK24]. capital [AAAF22, Ano93j, CKR00, Cor05, CP06, DVdCAS16, Dib02b, FOW97, HHV05, Hon21, Kaa09, Kap03, Kha87, LH11, LDB13, MZG17, PJ96, Pan98, SCC17, SH95, TPFD13, Tom90, TS88, TB84, Wan89, Wan91, Wan23a, Wil93, Won07, YK08, ZSMB16, Hob03, Coo97]. capitalism [TSA13a]. capitalist [GZXX23]. capitalists [FOW97]. capitalize [OM88]. captive [CD92a, CD92b]. capture [CWB98b, CWB98a, Far23, ML05, PGB^{+}23]. Capturing [MZG17, TTC06, UBCF19]. car [BK99, MWC22, MK00, MSR10, Pal93a, Pal93b, RR11a]. Carayannis [Bon02, Giu09]. carbon [AMS^{+}24, CX24, HS14, Tit90b, ZHY22, ZZL22]. Card [Kuc06, M’C05, Won04]. cardamom [Meh01]. cardiac [BEH99, Goe10]. cards [WMGB01]. care [CJH^{+}07, EFC23, KRK^{+}22]. career [DBV97, Fox86b]. Carlo [Arc03, BH03]. Carlos [Ban04]. Carlota [Kap03]. Carol [Fra02c]. Carroll [Box01]. carrying [Sza90]. Cars [BDW06, YNS18]. carve [Pow10]. carve-outs [Pow10]. Cary [Dib02a]. cas [Ano94w]. Case [HB84, PR97, WZNG17, AY05, AM23, AM10, ASI99, AT93, AA91, AW11, Ano93s, Ano94-28, Ano94-29, Ano94w, AHP16, Arc88, ALCM14, AM97, BCM12, BBLW24, Bas01, Bha05, BSMC23, BDNP06, Boh18, BLLEM23, BM07, BHPB06, BWS11, BP01, BLQS00a, BHPW00, CLZ23, CD92a, CD92b, Cam93, CCP23, CG02, CC05, CT06, CS06, CM89, CEY11, CM94, CHT99, CJH^{+}07, CF01b, CGS^{+}24, CK12b, Day02, DVMS16, DSVM05, ETOA00, Eri00, EN07, ET98, EF11, FAH08, FRA05, FJV96, FLKF^{+}06, FPP24, Gar05, GC99, GSFS21, Goe10, GFK12, Gra82, GARM24, GPR89, HKRR20, HS91, HV96, HS14, HO14, HS02, Hen01, Hob88, HMFS24, Hos01, HLC05, HH08, II21, IS97, IMBH98, Jar86, JKF19, Joh08, JRS06, KdOL08, KC08, KI03, KJ97, KS21]. case [Kih03, Kin00, KJE96, Kod03, Kod05, Kod07b, KH89, KS04a, KB06, KM06b, KFL11, KKP17, LS05, LY96, LPMV22, LK20, Lee04, LS07, LKPP07, LVHR10, LDK22, Let03, LM24, hLwChLbL04, LfSS19, Liu93, LH19, LH11, Mac97a, Maj22, MGdG11, MCD06, MP09a, MFH03, McM94a, MG98c, Meh01, Meh88, MW09, Møl91, MPCRB^{+}03, MSvdK^{+}21, MBA^{+}12, NW03a, ND15a, NLC98, ND15b, NI05, dAOCM02, OKvB23, OW00a, OS98, Ove15, PLG24, Par05, PPB89, PGKF19, PV00, Pow10, PS00b, Poz86, PGB^{+}23, Pur91, RM14, RAC84, RG06, RFMC21, Ros82, RD05, RG85, RHW14, RI04, RNJ^{+}23, STA24, SO02, SG93, ST95, SP98a, SP98b, SU11, Sch94, SSS17, Sir91, Sna90c, Soh97a, Soh97b, Soh99, STZ03, Sto03, SF98, TKSM08, TPB12, Tho99, Tit90a, Töd94b, Töd94a, Tõn15, THHS22]. case [UBCF19, Uzu01, Vav06, VBA06, VK03, VOTH05, VM92, Wal97, Wal07, WC06a, WWM00, WA03, WKN04a, WM88, YNS18, YH09, YI05, YA11, YLB06, ZMPRP14, ZXX23, dMvAdMSPB93, dVV16, dHOvdKS16, vBDK21, vGS09, vdVMM09, CPSS06, Haa14, IH07, Kod07a, TB24, Tre98, VX97]. casebook [Per96]. casebound [Rus94]. cases [FJ12, Hay93d, KS14, SM12b]. cash [Ott97]. caso [Ano94-28]. cassava [OOJ^{+}00]. Cassell [Her92a]. Cassiolato [Que06]. casting [II98]. Catalan [LVS10]. catalysing [FHBDO16]. Catalysis [Fah05]. catalyst [HV96]. catalysts [CvZ05, HHV05, SS23]. Catch [Lal00, BG19, WCW^{+}14]. Catch-Up [Lal00, BG19, WCW^{+}14]. Catching [Fan06, Win87, ZIAO^{+}21]. catching-up [Win87]. category [MH04]. causal [DLKM24, LNB21]. causes [Fab96]. cautionary [Eek84]. Cave [Gil96]. CCI [RGFLRP21]. CCI-USOs [RGFLRP21]. CD [San07, But96]. CD-ROM [But96]. CE [dAASA21]. cell [Pil04]. cells [KH89]. cellular [HH08]. Celtic [Sha95]. census [SD10]. center [GC10a, SH08]. center-affiliation [GC10a]. centered [DBR24, SL14a]. centers [ÁRS17, Boa08, Col81, Fuk09, LW04, LC09, Old97, SD23, YLB06, Ahn95]. Central [CP05, Cha99, SH08]. centrality [CK22]. centre [CBGvdV22, Guy97, LGZ22, MPCRB^{+}03]. centred [FC06, FM07, Tit90a]. Centres [Lin07b, MGS99, Hay87c, MKGL05, Rot87a, Ste89, YHDR08, OKvB23]. Centric [Rea06, SR22]. century [Ano88m, Hit96, JMK05]. CEO [Ene22, ECB22, LLZ23]. CEOs [Mer01, ASOB23]. ceramic [Che12, CA07, LD01]. ceramics [GSFS21, OOMW03]. CERN [BHPW00]. Certification [Ano07l, DLC21, EFV02, GGJ22, LL04a, Ano07l]. chain [AAZ22, AZ90, Bar14, BSG^{+}24, Bes02c, BS05, Boy14, BVHJ24, CCGK23, CLZ18, CLZ21, CH01, Eng06, FC06, FdlCU22, Gra03, Kee14, KI14, KK05, KM06b, KK06, KRA^{+}22c, KRM^{+}24, KY06, LBA14, Mus09, NUSR21, PMT23, Pri17, Red14, RY14, SMC14, Sto03, Ven14, WH21, Wil14, XL23, Hay90b]. Chains [Lin17b, BCP22, Gra03, NJ04]. chairman [Ber89]. challenge [Ano89p, Arc03, ABKL08, BLS85, Bes88, CMP^{+}06, JCT23, JR91, LBA14, Moo03, MPWG22, Rea21, Shi24, SS16, Won07]. Challenger [Min96]. Challenges [DFFM22, LD02, Lee00, Lop06, WNLY06, AGFdMCM24, ALCM14, BN06, BAR16, CM95, CWBG03, DMRS23, DBR24, GKK23, Hea83, HBG12, Hos99, Hos01, HKR^{+}22, HC06, IPP13, JD09, JFCFMC^{+}12, KS21, KK05, KY06, LL02, PLG24, RS14, RB21, SPP06, DGMR21, SC05, SSKB24, Szá05, VSC16, vdVdJVdR09, Eng06]. challenging [CWBG03]. Chalmers [LWM24, MW82, Wal97]. championing [CDD^{+}11, CDD^{+}14]. champions [KA13, Mer01]. chance [FM07]. Chang [Ano94l]. Change [Gri98, Hay98, Mer01, Acs88, AW11, AP96, Bam94, BC12, BFO88a, CR08, DDOR08, DO99, FS84, FCH01, For85, Gil96, Gol82, Gre06, Gru93, HKRR20, Hal97b, Har96, HS02, Hay88a, Hob92, Ken13, KJ97, KL87, LO10, LVHR10, LM24, Lov85, Mac86b, MCD06, MVS^{+}22, May86, MC00, Mon06, Mow89, NW23, NBA22, Oak97, Pur91, RAC84, Sci83, SJSP23, SJP23, Shi24, SHS^{+}23, Sur92, SM12b, Tho88c, Tre98, Wal85, WD06, Whe86, Whi89a, Whi90a, Whi90b, Whi90c, Woo85, YLH22, Zeg87, ZZL22, Mer01, Tra00, Bes94b, Dod88, Eva02, Hay87d, Hay92a, Hay01c]. Changes [NWGB09, CCP23, Fra02a, KHL15, MTB01, dVV16, Obe86]. Changing [CBK03, LDK22, AB99, BF04, BFM86, BFM88, For99, Rot93, Sha95, Tho12, Whe88]. channel [CSL06]. Channels [Har97a, Har97b]. Chantal [Ano93a]. Chapman [Col97, Gar96, Hay93d, Hay93a, Hob97, Per97a]. Characterisation [MdCÁ06]. characterise [ØD22, RPASLM23]. Characteristics [Lüt04, NPV05, Rob84, AF08, AEP21, BG10, Cha08, CbLjJL12, Col81, Dan19, DB05, GR23, JLL10, LLS00, LQ09, NPG14, PV01, PG11, RAC84, SCY17, SM12b, WSF13]. characteristics-cases [SM12b]. characterized [TTC06]. charged [LC07]. Charles [Dib01, Hay89a]. Charlie [Dre06]. charts [FGS07]. Chase [Bes94a]. chasm [PS14]. ChatGPT [CGS^{+}24, TKYM24]. checklist [QvdR21]. Chee [Arr04, Ano93e]. Cheltenham [Arc03, Aug09, Ban04, BH03, CF07, De 04, Dib00, Dib01, Dib02a, Dib04, Dre06, Gra04, Her04, Lae03, Lal00, Mor04, Pap02a, Par07, Que06, Sap05, Won08, Wya07, Fra01b]. chemical [NB09]. chemistry [MFH03]. Chen [CF07, Saa98]. Chicago [Min96]. Chichester [Del97, Fra99, Gil97, Gri98, Hay85d, Hay85b, Hay86d, Her92c, Joh92g, Mil85, Mor91, Per95, Pov97, Pre01, Sco89, Ver91b]. Chien [CF07]. Chien-Hsun [CF07]. Children [TDA05]. Chile [PGKF19]. Chilton [Joh92a]. China [CLZ21, Col08, YHcHW21, YGW08, AMS^{+}24, BCO14, BSN21, Boh18, Cai23, CS04, CS05, CS00, CQ03, CR06, CLZ18, CYHC19, CZM23, DC22, Fan06, GC10b, GG11, GGJ18, GMWK23, HFWW16, HL22, JWS92a, JWS92b, JvZ08, KL08b, LY96, LT16, LWL^{+}22, LJ01, LH19, MLC24, ND15a, PCG22, RHW14, SD10, SJCT24, WWY15, WTJ23, WCW^{+}14, WZNC15, WSN^{+}18, WL21, WLTH23, WW14, XW21, Xie04, XX21, YC05, YHcHW17, YGW06, ZI01, ZQW^{+}19, ZYZZ24, ZMH22, Zho92a, Zho92b, CF07, Arr04, Wan04]. Chinese [CL98, CS00, CF01b, CCV11, CC14, CZM23, GM03, HMWT23, HW18, Hit03, Jol08, KS10a, LW97, PCQ24, RY14, SCH23, Thu18, YNS18, ZLJ^{+}21, ZL24]. Chintay [Ano94l]. Chiung [Ano94l]. Chiung-Wen [Ano94l]. Cho [Kap01]. choice [AdA08, CY00, HT08, MAB^{+}23, PWL20, WZY23]. choices [TWBC21]. choose [JLN22, LVS10, MAB^{+}23]. Choosing [JLN22]. chosen [LP03]. Christensen [Sap05]. Christian [Ano94-38]. Christiansen [Fra00a]. Christie [Ano91d]. Christopher [Bes01a, Har93a, Har93b, Hay86c, Joh92d, Que06, Ver90b]. Chu [OH04]. Ciborra [Gra01]. CIIPM [JW06]. CIM [Gup96]. Circular [dVGCOF^{+}23, RES^{+}24, dAASA21]. Cisco [Li09]. citation [SKTM08]. citations [Yam21]. cities [PC18, Tsa05b]. City [SSS17, BSSF24, MGAP23, SI98]. claims [SAWT06]. Clarendon [Tan08]. clarification [RSP10]. Clarifying [YK08]. Clark [Wil90]. Clarke [Mer94b]. clash [Lin85]. class [Tom90, VM92]. classic [CGBV21]. classical [MHKZ24]. Classics [Dib02a]. Classification [LOCSBGMH16, Par00, SS21]. classifying [Mou15, Mou17]. classroom [LB82]. Claudia [Cou90, Cou91]. Claudio [Gra01]. client [WFW02]. client-supplier [WFW02]. clients [RS03]. Cliffs [Dod87b]. climate [Ken13, MVS^{+}22, Moh95, NBA22, WHP85]. clinical [BCPP23, BN88, MWJ23]. Clippinger [Fra01c]. clockspeed [DM11]. closed [HT12, KM06b]. closed-loop [KM06b]. closer [NHV05]. cloth [Gar87, Hay97c]. clothback [Ric95]. cluster [Ano12a, CY12, FGPP22, LS07, LTH06, Se14, TKSM08]. clustering [SY05]. Clusters [Giu09, AM10, BCM12, Car04, Car05, CC06, Day02, GG11, HLH12, KTSM10, Liy95]. CMOS [TB24]. CNC [VX97]. CO [Hob87, BBK23, Gee06, Hay84, Hay92d, Hay97b, HKR^{+}22, JSP22, LB06, Men11, WH04b, Wya07, BGR91, CA01, CNN23, Cul00, EFC23, Goe10, Hag90, JLY23, LVR23, LP06, DGMR21, SDK24, TWGB02b, VvWM20]. Co-alignment [HKR^{+}22]. Co-creating [BBK23]. co-creation [CNN23, DGMR21, SDK24]. co-development [VvWM20]. Co-evolution [JSP22, WH04b, Wya07, EFC23, JLY23, TWGB02b]. Co-evolutionary [Gee06]. co-innovation [LVR23]. Co-locating [LB06]. Co-operation [Men11, Hag90, LP06]. co-operations [BGR91]. co-operative [Cul00]. co-opetition [Men11]. co-opetitive [CA01]. co-patent [Goe10]. Co. [BN88]. coach [MKL19]. coal [Sur92, XW21]. coastal [DM08]. coat [HPT10]. Cockburn [Sha95]. cocoa [Ano94t, Ano94j, IAA94b, IAA94a]. code [MSR10]. codification [AGS01]. Coevolution [Nuv07, YIKC21, HS21b]. coexistence [ZAI23]. cognition [Car99a]. Cognitive [CS06, FB91, FB93, Fox95, Car04, CNS06, DDN24, GMWK23, KFEC15, Lin85, RG10, BF23, MMC^{+}23]. Cohabitation [Bul85]. cohesion [MMM05]. cohesiveness [Hua09]. Coimbra [MCD06]. coincidence [FR15]. colaboración [Ano94i]. Cold [Mer94b]. Colin [Del97, Gri98]. Collaborating [GBY22]. Collaboration [RV21, dMD05, ADDS09, AA92b, AA92a, Ano94u, Ano94g, Ano94i, BK15, Bje10, BHS05, BJ05, Cai23, CBGvdV22, CZZM21, CKF24b, DDN24, EH14, FORA17, HCMMH21, KGK07, LfSS19, ML08a, MZL15, MS19, Nar04, NI05, OGOF22, Old97, ØD22, RV17, SLA^{+}23a, TS13, Tit94a, Tit94b]. collaborations [BH15, DJ05, HBO14, KP12, Pet11, PUFC23]. Collaborative [DDG98, FH87, MOM08, Mig19, PWA11, SBP21, AD09, BMW14, BLL95, CMA07, CYHC19, GLYP13, Hob88, IBPGML23, Joh08, KKG07, Liy95, NS07, SBF15, SMvFM24, SVV15, WH21, XLC23, ZC22, vRGJ^{+}12]. collective [AAM24, LBG14, RFB24, SA22]. Collectively [Ove15]. collectives [LBG14]. College [Dre04, Hay86c, Lae04]. Collier [Hay84]. Collins [Mur91]. Colombo [Col89]. colour [Xie04]. combinations [TA10]. combined [Cha08]. combining [LB06]. combustion [WTAW97, Sch94]. come [SBF15]. comedy [McM96a]. coming [CCR^{+}23]. Comment [Lin10a, Tho88a, dB10, vH10, HH89, McM94b, Suc87]. commentary [vK11]. Comments [Ho09]. Commerce [Kuc07, IKL01, KY06, LL02, LLEB05, MB04, WPF06b]. commercial [Ali86, Ano94n, Cai94, MTU23, Mow90, Oye01, TDC11, Whi89a, Whi90a, Whi90b, Whi90c, Won07]. Commercialisation [Mac97a, CFC12, HY04, KK23, KL08b, Pal06, Pil04, Sik97a, Sik97b]. Commercialization [KJ03, WWH15, Ano94-30, BPU21, BB22, cLhWCH12, NC23, Pow10, RMG06, Ras08, SS94, Wal12, Wec15, vRGJ^{+}12]. commercialize [ZT21]. commercializers [HP09]. Commercializing [PG11, SSKB24, Bro85, HGB23, JFCFMC^{+}12]. Commission [MPCRB^{+}03]. commitment [xCZZ^{+}22, Huc85]. common [Bro07, CE89, EBG24, GKK23]. communication [ADK08, Ano93f, Bad11, Car05, GL10, JCT23, LD05, Lic11, Lin10a, LL08b, PB93, RMGDLC08, Sou01, dB10, vdVdM11, vH10, vK11, Ada07]. communications [KZ03, LM24]. communities [CB22, CX24, Far23, GZXX23, HLC05, Hu13, MGdG11, MW14, Rei97, TPB12, TAM11, TDA05]. Community [OKvB23, BBK23, CS21, Cor87c, CL88, Kas94a, Kas94b, Kod05, Kod07a, MGdG11, SA22, WL21, Wit85, WF10, XWZ23]. community-led [BBK23]. companie [ACA24, SLSCS24]. Companies [Bes01a, Fra01a, Fra01b, Mer01, Ano94-29, Bac98, tBL86, Bar92b, BT22, BP09a, BP09b, BBHA01, Bul85, CW22, DB05, FdlCU22, FMvT00, GBM98a, GBM98b, dCHDAALMM07, IOS00, JBZ23, KM01, KAV17, KBC22, KM97, KC07, KBRR24, Lin86, MW82, MM11, Neu12, OFR24, OM00, OE87, PLC13, Pro21, Rie89, RRB^{+}22, Rog86, Rue86, Sal03, SW04, SCZ14, SCY22, Tit94c, Tit94d, UBCF19, VJ12, VP14, VM92, WR95, XW21, dGSJ^{+}21, dBJ02, Mat22, Lin91]. Companion [Ano91d]. Company [JC01, Kra15, ASI99, BLLEM23, CF01b, Ern95, Gar05, GC98a, GC98b, GTW02, INO^{+}03, Kod03, Kod07a, MT05, Nob04, PSdLC02, RI04, Ruo05, SPO08, Ste14b, YHDR08, Lun88, Hay97b, Bes93]. company-based [YHDR08]. Comparative [DTI^{+}23, RG10, CJ98, CEH03, CS04, CbLjJL12, Cor87a, Cor87c, DLL06, DPH^{+}22, FR15, GMWK23, IMIY87, JL98, Ket09, Kum06, MKK04, Nwo02, Sub05, iSKB02, TM03, WMGB01, Yas05]. compare [SS22]. compared [Gru93, HW18, LDB13, Rag93a, Rag93b, TM03]. Comparing [BRCG12, CS05, JB19, JJS^{+}18, Ano94f, BP94, SU11, CZL^{+}24, vAdMSL95]. comparison [AC04, Bac98, CL97b, CL97a, GEG12, HG92, Kha98b, Kha98a, Kha99, LS05, LLS00, LL93a, LL93b, Lin11f, OG05, PC89, Sou01, Tre98, VJ12, WNM95, XR08, YHDR08]. comparisons [DBL07, IH07].

Compatibility [KR12]. compete [BH03, Hay88a, De 23]. Competence [Bas01, DR99, ACV08, Ban03a, CA02, CL21, DVSL19, Dib04, FdJ09, Pet98, STH06]. competence-oriented [Pet98]. competence-seeking [FdJ09]. competencies [Ban03b, Dre01, HPMI^{+}23, LOH91, Lee23, MRGMBR13, WVG^{+}98, WGV^{+}98]. Competency [KC08, MJ24, ZL01]. competency-based [ZL01]. Competing [Dod87b, Per96, HW18]. Competition [Kuc06, MAKKM22, Sna01, Whe86, BVT21, BG19, CT02, CYK19, DC00, Knu86, LHY04, MFH16, May86, MPW03, Obe86, Old97, QGBV04, RkhR97, Sci83, SS07, TLL23, Wer03, Gra04]. competitiva [Ano94q]. Competitive [Arr04, LL94a, LL94b, Nuv07, AAAF22, Ano90o, Ano94-33, BS01, Bes03b, Bro94, Bro90, BA89, CW94, FCM21, GL22, GCS01, IDMD04, Kar93a, Kar93b, LB03, LH07, NGK08, Sch92, SGFS99, TWGB02b, TCT05, WKN03, WAN03, YGM99, ZL01, AD09, Ano94q, Ano94-45, Bes94a, Bes95a]. Competitiveness [SC05, BBDM01, BF99, BS05, Cam93, CR00, CS01a, CbLjJL12, Coc12, D’I14, Day02, Dib01, EL93, Gol82, KK04, LV23, LCT04, Pap95, SL14c, dMvAdMSPB93, Hay89d, Ver90a, Ver91a]. complement [WTN^{+}02]. complementarities [CM04, KA13, LR09]. complementarity [AMP22, OM88, SH22, ZSMB16]. complementary [BR16, CUF13, LKCK22, MW09, SSF12, SM95, vdEJG08]. complementing [KTE04]. Complementor [JWY23]. complementors [MPO23]. complements [CKT04]. completeness [KR12]. completeness-Identifying [KR12]. Complex [CJS00, Kuc06, Sod08, ALCM14, BBC06, Bas01, CC05, CG04a, DLKM24, HR98a, HR98b, KT04, PÅ06, Ree99, Rep08, RAATM05, Wil14]. Complexities [BLL95]. Complexity [CH04, Jac13, Car98b, DZX22, DGS23, ILR^{+}12, Mus09, PLC13, SP03, Tag97, Won05a]. compliance [KRM^{+}24, LTT23]. component [Ame95, LY96]. components [FPP24, KK04, SH95, Web96]. Composing [TKS18]. Composite [MOM08]. composition [Jan21, VP14]. compositional [SHT23]. Compound [CEY11]. comprehensive [SDT^{+}22, VM11, Sæt24]. computational [DHT22]. Computer [Mil85, AK00, BF05, BLS85, Cam93, CN05, CHT99, FB91, JWY22, Kha99, KLL87, Lay93, Mil88, OM00, Pur92, Sch92, Sci83, Tan87, Wan04, dMvAdMSPB93, dHOvdKS16]. computer-aided [Pur92]. computer-integrated [BLS85]. Computerised [Soh97a]. computerization [LT05]. computerized [LW91]. Computers [KKRM05, EM02, LC22a, RA86]. computing [BC99, FB93, GMB^{+}23, GMWK23, Par03a, Pyk07, RL06]. Concentration [AES95]. concept [Ano94c, AB89, HBvdB09, Hit89a, Hit94, KR12, Kli06, Kob97, MGS99, OGOF22, OW00b, PFP06, BPU21, CL24]. ‘Concept-Product’ [CL24]. concepts [Ano90o, Ano94-33, CEH03, CWY14, CW94, Hit93, Møl91, OM18, PT06]. Conceptual [KW05, Ano93o, BBHA01, FSA23, Fox86a, GH21, HLL93, Lag17, LGY05, LO94, Mal02, MMMD23, OW93a, OW93b, RS03, SDT^{+}22, DGMR21, WSH23, ZXX23]. conceptualization [NJ04]. Conceptualizing [KLP24, PD10, TSPJ23, vdVMM09, JBZ23]. Concerns [Bha05]. Concorde [Ano81a]. concurrent [Dow93a, Dow93b, HPB03, Kli06, WA04a, Kle13]. concurrentielle [Ano94s]. conditional [Coz06]. conditioning [CL22]. conditions [AW11, BSWC23, BFM86, BFM88, CLP23, JSR19, MMR24, WZNC15]. conducted [Cor87c, Fon05, NS90]. Conference [Ano88a, Ano89b, Ano89c, Ano89d, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90g, Ano90h, Ano90i, Ano91e, Ano95b, Ano95k, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95i, Ano95j, Ano96b, Ano96k, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano96i, Ano96j, Ano97c, Ano97l, Ano97m, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano98m, Ano07k, Ano07m, Coo16, Lop06, Ano90o, HdWN21, CR19, Pis89, TKS18]. conferences [Ano88g]. configuration [Pri17]. configurational [CW22, EEY22, TP24]. configurations [LVR23, LM14, MMC^{+}23, WSH23, Pro21]. configure [CW22]. Configuring [FLC17, KSHP21, PA15, BOBT22]. conflicting [LMMPVO09]. conflicts [KR12, KA13]. confront [SZSR23]. conglomerate [CLT06]. congruency [Mac86b, Mac88]. conjoint [LB21]. connectedness [NT16, SSGN23]. Connecting [CKF^{+}24a, LWL^{+}22, FHA22]. Connection [MPO23]. connections [WSN^{+}18]. connectivity [FPP24, MPO23, XR08]. connectivity-Absorptive [XR08]. consciousness [Sar95]. consequences [Ano94-35, ACA24, BC12, CCMW23, Fab96, GGS08, HO11, HCL09, Kan94, Knu86, LW21, SG24, Tei22, NG09]. conservation [AEP21, CK04, ZWZ^{+}23]. considerations [AGM89, Dre97, ETAO98, Ott98b, OB04, TOA01]. considered [GD08, LO96]. considering [Tsa13b, ZHY22]. console [CUF13]. consortia [AW12, LFFT09, WKN04a]. constant [BFO88a]. constituencies [Mol98]. Constituents [SRHL12]. constrained [BM16a, Cha22, HKL24, YWL22]. constraints [SY14]. construct [CS08b]. Constructed [Lak08, Har93a]. Constructing [WA04a, RZ21]. construction [AM00, BN06, DQ06, FRWT06, Hun00, PPM24, PC89, RD05]. constructive [HS21b]. constructivist [BMRG00]. constructs [IBL04]. consultancies [HBvdB09]. consultant [KI03]. consultant/researcher/innovator/entrepreneur [KI03]. Consultants [Mer94b, Ano94b, Old97, WvT23]. Consulting [Hol90, dCHDAALMM07, Sna98, VM92]. Consumer [CSL06, PMN^{+}24, RWP19, WKF24, wK97, BF04, CZL^{+}24, DAOO02, Fox95, GFE95, Haa14, KK07, LA15, Lüt04, MWC22, OW00a, Tsa13b, WPF06b, WF10, ZC19, ZHY22, dRP95, BVT21]. consumer-to-consumer [WF10]. consumers [Lüt04]. consumption [CP23, OTA01]. contact [CH01]. contacts [BBTW00]. contemporary [Cha88, KLL24, Ola98a, Ola98b]. Content [Ano95l, Ano93i, Dav93, SW91]. Contents [Ano82a, Ano84b, Ano85b, Ano86a, Ano87a, Ano87b, Ano88b, Ano88c, Ano89e, Ano89f, Ano92b, Ano93c, Ano94e, Ano96l, Ano97n, Ano02b, Ano06a, Ano01c, Ano03c, Ano04q]. Contest [TP24, HBH21]. contests [FDN^{+}23]. Context [HL24, MSvdK^{+}21, AAM24, ASOB23, BCPP24, BKM10, DM11, Kra22a, Lic10, LZD24, MKL19, PW99, RGFLRP21, RC23, RB21, SM09, SC05, Sou01, TC97, TPB12, WPF06a, YMMYK24, dV94]. contexts [BBLW24, HOGGP11, JB19, PUFC23, YK08]. contextual [DTD13, YMK^{+}21]. contingencies [DTD13]. contingency [Dre02, GSP^{+}22, KOX21, KB09, LD05, Pri17, VvWM20, WLY14]. contingent [DLTQ23, LLY^{+}18, SP22b, ZXDG19]. continued [EN07, Mas86]. continuing [BBTW00, Gol86]. Continuous [BBHW93, BCG^{+}94, BCG01, Bes02b, IS97, JW06, MC00, Pov97, SMvFM24, TS00, WC06b]. contract [Rad99]. Contracting [Cul00, CH01, BHS05, GAS96, HN99]. contractual [BH15, NH99, SK23]. Contrasting [HBG12]. contribute [EHHW17, HDE21, SFVI21, WTJ23]. Contributing [VJ12]. contribution [CA11, Lic11, MFH16]. Contributions [Can16, Per97a, TFS^{+}22, EKMS12, LW97, SSC23, SW91, WPF06c]. Contributors [Ano88d, Ano88e, Ano89h, Ano89i, Ano89j, Ano89k, Ano89l, Ano89m, Ano89g, Ano86i, Ano88i, SJCT24]. Control [Ros89a, Bes02b, Boy14, CM96, FLKF^{+}06, Gra01, Kuh88, KK05, LVHR10, LW21, OBvE97, Rue88, dHOvdKS16]. controls [DDN24, RGB23]. converge [CWAB09]. Convergence [Bat12, AK00, Bje10, BST03, CP05, Cav16, CMH14, HMFS24, KKL19, Kod14, LS24, RD92, SPLB19, ZZL^{+}23]. Convergent [Jol08]. Converging [Lee23, RH10]. conversion [Vek99]. convivial [RPGP23]. convolutional [HKW^{+}22]. Cook [Fit00, PF98]. Cooke [Hay93c]. Coombs [Lae03]. Cooper [Dib02a]. cooperate [GF21]. Cooperation [Ali90, BG10, QGBV04, SH10a, Sta98, Ano89p, Ano93m, AdA08, Bøl12, Cai23, CBK03, Cor87a, Cur89, DH06, Fuj98, Gei97, GF21, JKF19, KJE96, OM89, PWL20, SPFG12, SCY17, Sko87, Tei93, ZXT10, dBJ02]. Cooperation-based [BG10]. Cooperative [Ano94f, BP94, YLB06, Ahn95, BM07, Gra24, LB87, Miy96, NL11, SYCH23, WH04c]. cooperators [BG10]. Coopers [Gil96]. coopete [FCCE21]. Coopetition [CKV^{+}23, LRH22, MNL22, RHL09, RFB24, TLL23, XLG^{+}23]. coopetitive [JKF19]. Coordinating [DR07]. Coordination [ZDWG14, CG04b, EW06]. cope [SVLYT24]. Copenhagen [Pis89]. Coping [HLO14, Shu22, SFF21, ZWZ^{+}23]. CoPS [ZI01]. Copyright [DC00]. Cordella [Gra01]. core [Ban03a, CA02, Car01, CTC02, CL21, Oll08]. core-technology [CL21]. coreana [Ano94q]. coréenne [Ano94s]. corkscrew [dV94]. corn [SK92]. corner [McM96b]. Corning [Bes02a]. Corporate [ALvKL06, DMP05, DM91, HSLS24, Lic05, Why05, AGFE22, ABvZ23, AP08, Bac98, Ber96, BV09, Bul97, CS87, Dod87b, ES21, FF00, Gal12, GvBR10, Gra01, Gui11, Har97c, Her91, Hol94a, Hol94b, IS97, IBL04, KL24, Kod06, LZLL21, MRGMBR13, MPBC21, MPW03, Rya96, SR22, SXS24, VOTH05, Wan23a, WR95, XW21, Dib04, Hay85a]. Corporation [Bai84, Sou83b, LG22, Sch94]. corporations [Tag97, TRR^{+}24, ZL24].

Corral [Ban04]. Correa [Ano93l]. correlates [ZMPRP14]. correlation [LCT04, LCT05, TCT05]. corridor [RCC04]. Corrigendum [Mou17, SJSP23, YJLL21]. Corsten [Ano93b, Ano94-33]. cosmetic [Kum05]. Cost [AO03, Fru92, Ram23, Ano93i, AR99, Dav93, Yas05]. Cost-effective [Fru92]. costing [BB95, GWM05, GG03]. costing/management [GG03]. costs [AES95, BWS11, DDBM23, IEG97a, IEG97b, MY02, SH95]. Council [CR02, Tuz05]. Counselling [Lin85]. count [AT93]. countries [AHSS19, Alc99, Bas01, BBHA01, CP05, CWAB09, CS89, Der14, DGS23, ETOA00, Els95, Gui11, IH07, JL98, KA15, Kim87, LPMV22, LWP14, McQ05, MAKKM22, MPWG22, OOL06, RH10, Sar08, Sir91, Sto99, Sto03, SD04, WC07]. country [Chi90, FAH08, Had99, KL87, KP23, Lee11, LV09, LT16, LfSS19, ND15a, Sou01, SH99, Tre98, VGQK^{+}10, WW21, Yam21, YA11]. country-Taiwan [Chi90]. coupled [AABY^{+}23]. Course [AAD14, LP06]. Courses [GHU21, CK22, Jac22]. Court [Sha95]. courtroom [Whi94]. cover [Lae03, Lin11c]. COVID [SJSP23, YVG24, ABAHP^{+}23, BKA^{+}23, BEE24, CCBP22, DGR23, FGB^{+}24, GP22, KRK^{+}22, Kra22b, MWJ23, SJP23, TFS^{+}22, EHSB24, GMB^{+}23, Rea21, SSKB24]. COVID-19 [SJSP23, ABAHP^{+}23, BKA^{+}23, BEE24, DGR23, FGB^{+}24, GP22, KRK^{+}22, Kra22b, MWJ23, SJP23, TFS^{+}22, EHSB24, GMB^{+}23, Rea21, SSKB24]. cowpea [Tai98]. cradle [LY00]. Craft [CS21, Bes02a]. Crang [Wya00]. CRC [Bon02]. Create [Mor04, Haa14, PE22, RM14]. created [ACA24, Ove15]. Creating [EF11, KS07, KB09, MZL15, RWP19, UBCF19, WH04a, YH11, BBK23, FBBBM22, Fra01a, Hyv06, Joh92c, VRS^{+}22, CP11, RHL09]. Creation [Bas04, Kar86, AAM24, Ano93o, BBA15, BGF23, CMA07, CE19, CNN23, CC06, Cre04, DJ05, FMF21, Far23, FPP24, GH15, GL22, GD08, GvBR10, GG05, HLL93, HLH12, JB19, Kap01, Kih03, Kod05, Kod07a, MLG12, MSD15, dCLSNL08, McA04, NPS02, NG97, PWL20, RUNS23, DGMR21, SMI23, SDK24, Tet00, WWM00, WH04b, YNS18, ZIB23, ZZXW23]. Creative [CC05, SBP21, CSL23, KLV06, LDF^{+}21, Li21, Mal87, OGOF22, Rob84, RPGP23, RGFLRP21, TWGB02b, ZLJ^{+}21]. Creativity [Fit00, Bes01b, BD05, FCM21, FHA22, Hal96, Hay90a, Leo21, LD95, MCA^{+}24, MM02, MAU07, NKL02, Ric95, Rob84, TWGB02b, Ved81, Fit00, Sap10]. credibility [JH18]. credit [CS13, Paf04]. credits [DPH^{+}22]. CREST [LD95]. crises [VSE24]. Crisis [SZLR19, ABAHP^{+}23, JLWS23, KRK^{+}22, Mac05, Szá94a, Szá94b, Yam21]. Cristiano [Eva02]. criteria [ANI^{+}22, Col81, CNS06, Kak03, Lan98b, MNW14]. Criterion [WHB21]. Critical [AMV07, CS84, Dib00, PC89, Sch86, SW05, Bes97, Boy14, CD92a, CD92b, CPXC02, DSS05, KMK03, LL98, McA04, Moh95, OPPL16, THC^{+}23, CA06]. Criticality [Red14]. critique [AC09, BM14]. critiques [Ros89c]. CRM [DB09]. CRO [Pia02]. Croft [Tat04]. Croom [Hay85a, Hay86a, Hay88e]. crop [EKT^{+}22, SYK94]. crop-production [EKT^{+}22]. crops [IIA97b, IIA97a, Ree99]. Cross [CLGFVCM23, CM95, LC07, MP09b, CEH03, EH14, Fra03a, GGGR22, GC10b, KP23, LLS00, LCRL22, LR09, RGJ13, Sna90d, SSGN23, WW21, Yam21, ZWDZ21, ZYSL21, AAT14]. Cross-border [CM95]. cross-country [KP23, Yam21]. Cross-cultural [MP09b, CEH03, Fra03a, LLS00]. Cross-Entropy [AAT14]. cross-functional [LR09]. Cross-functionality [LC07]. Cross-industry [CLGFVCM23, EH14, ZWDZ21]. cross-institutional [ZYSL21]. cross-instrumental [GGGR22]. cross-quantile [SSGN23]. cross-region [GC10b]. cross-sectional [Sna90d]. Crossing [EWE22, PS14]. crossroad [SL16]. crossvergent [FLC17]. Crowd [BVS24, LS16]. Crowdfunding [ZC19, EHSB24, JWY23, JLN22, NK22]. crowding [CGM22]. crowds [MW14]. Crowdsourcing [LS16, CMB23, GBV23, MM24, MAB^{+}23, NPG17, RB21, TP24]. CSR [RM14]. CTV [Xie04]. cuisine [Leo21]. Cultura [CS21]. Cultural [CL98, Sap10, CEH03, FLC17, Fra03a, HBO14, LDF^{+}21, LLS00, MP09b, RGJ13, RGFLRP21, ZLJ^{+}21, SBP21]. culturally [Mil02]. culture [CS01b, Des96, Efr14, Hol94a, Hol94b, IS97, IBL04, LS04, Min96, RW86]. Cultures [Tra00, DQ06]. cumulative [LLR95]. curiosity [CSL23, LLZ23]. Current [Lin08a, Sun00, Bro89, CvZ05, KH06, Lic11, MGAP23]. curve [NLC98]. curves [AH85]. custodial [FWL06]. Custom [DFD^{+}91]. Custom-made [DFD^{+}91]. Customer [EKMS12, Rea06, TTC06, Fox86a, FM07, LS23, LZD24, MH98, RI04, Ruo05, Ski84, SCS06, YK08, YJLL20, YJLL21, Lin90, Lin91]. customer-active [Fox86a]. Customer-capturing [TTC06]. customer-centred [FM07]. Customer-driven [Lin90, Lin91]. customer-nomothetic [Ruo05]. customer-perceived [YJLL20, YJLL21]. customers [HWC06, XWZ23]. Customisation [SDK24]. Customization [SSP04, HJ02, JMT03, WCZZ14, WZSZ16]. Cyber [Bar14, Boy14, Kee14, KI14, LBA14, RY14, SMC14, Ven14, Wil14, GPCNS^{+}23]. Cyberdyne [II21]. cyberinfrastructure [YIKC21]. cybermastery [Gra03]. Cyberspace [Hol93]. cycle [Ayr88, AM92a, AM92b, DMP05, EM02, GC93a, GC93b, Kim03, RG06, TM06, WWM00, WAN03, WA04a, Wer03, Won05b, XL23, vdBS02]. cycles [AT93, CA14, Hay86c, Pyk00, San95, SH23, SB91]. cyclic [TB24]. Cynthia [Sha95]. Cypriot [ETAO98, ETOA00]. Cyprus [Had99].

D [Ano93f, Ano94b, Ano94m, Bes93, Bes02d, Dib00, Fra99, Fra01a, Fra02d, Giu09, Hay88c, Hay88g, Hay88e, Hay89d, Hay01a, Her91, Jel07, Joh92b, Joh92e, Kaz03, Mil85, Pap04, Sha91b, ALFM23, CB19, CNN23, DTI^{+}23, FMF21, SA22, UF22, WAdN19, ZYSL21]. D’Adderio [Mor04]. Dahlbom [Gra01]. Daniel [Hay90d, Hay97d, Wya04, Ano94-42]. Daniele [Hay01c]. Danish [Knu86, Obe88, PC05, TS88]. Dankbaar [Lae04]. dans [Ano94v, Ano94u, Ano94s]. dark [FEP15, McM96a, PEER23]. DARPA [Par03a]. Data [BBLW24, JLY23, MMCD23, ACBO23, BCPP23, BCPP24, BLLEM23, Bro92, Cav16, CGS^{+}24, JPBL^{+}23, KZ18, KHKL23, KLL24, Lee11, LG23, MW14, NSYL18, PK23, PMT23, PDHW23, QGBV04, RAG11, RKFR24, Sma01, SCS06, THBW22, JPBL^{+}23, SPPG20, UBCF19]. Data-driven [JLY23, BCPP23, CGS^{+}24, NSYL18, RKFR24]. data-mining [SCS06]. Data-push [BBLW24]. database [LJAK21]. David [Dib02b, Fra03a, Hay90c, Hay91a, Hob03, Mer01, O’H99, Ros90, Szw07]. Davids [CR00]. Davies [Ano93i, Edw04, Sod08]. Davis [Hay88c]. Day [Joh92c, KS00]. DBM [KLP24]. De-babelizing [Lin09c]. dead [MHKZ24]. Dealing [Ott04a]. Dearing [Ada07]. death [EWE22, DZX22, LCRL22, RO22, SCY22, TNY^{+}22]. debate [Hit89a, Lic11]. debates [EFC23]. decade [BCP22]. decade-long [BCP22]. decades [LWM24]. December [Ano20a, Ano21d, Ano22c, Ano23c, Tit90b]. decentralized [Ver91b]. deciphering [WTGB01]. Decision [DO99, ANI^{+}22, BCPP23, CDED21, Clo91, DLR07, Gim06, HJM08, II97, IEG97a, IEG97b, JGM98b, JGM98a, Kap99, KR12, MSD15, Mal87, ML05, Min96, PLG24, SI98, SW91, TLL23, WCW^{+}14, Lan98b]. decision-based [Clo91]. decision-maker [HJM08]. decision-making [ANI^{+}22, CDED21, IEG97a, MSD15, SI98, SW91, Lan98b]. Decisions [Hüb96, BC10, Bea98, BCP11, CLIH05, CFC12, Jar86, LVS10, LCRL22, PMT23, SCS22, SRLR23, VGAL24, ZHY22, ZCR22, Mil85, Rus94]. decline [BS01]. Decoding [TRR^{+}24]. decommissioning [Hay92c]. decompartmentalization [JCT23]. decoupling [MV06]. decreasing [BZIZ21]. dedicated [QGBV04]. deep [CZL^{+}24]. deeper [FMFMA16]. deepfake [WML^{+}23]. defence [Gui05, KPK94a]. Defending [BCP22]. defensa [Ano94r]. Defense [IMBH98, Vek99, Ano94-36, Ano94r, Ano94-41, CM95, KPK94b, RA17, Ano94-36]. deficit [Tom03]. define [Dre01]. Defining [CBF24, MKGL05, YKW10]. definition [Bar92a, BK89, Ehr95, GR21, GH21, vDvM89]. degree [PCG22, dOK12]. degrees [OB20]. déjà [Mow90]. del [Ano94t]. Delay [Llo07]. Delhi [Ada07, Bas04, Kaz03, Kuc07]. delicate [Pap95]. delivery [SWW12, SDT^{+}22, VNJ18]. delocalisation [AC07]. Demand [Rie89, AOO02, BRCG12, CA14, DLC21, ETM96, Lae03, TCC^{+}22]. demand-pull [DLC21]. demand-side [CA14]. democracies [CU13]. Democracy [Wal85, WFWC21]. Democratic [For85]. demographic [KHL15]. demonstration [MTU23]. demonstrators [Mou15, Mou17]. demystified [Hob97, Hay93a]. Demystifying [xCZZ^{+}22]. Deniozos [Ano94-34]. Denmark [Møl91, RG10]. Department [Hay86c, LTT23]. Departmental [Sna90c]. Departure [BiKM95]. dependence [Ban03a, DLTQ23, Fab96, MMR24, SSGN23]. dependent [MSvdK^{+}21]. deployment [MH98]. depots [SI98]. depth [CCV11, GKB12, LW10, TRR^{+}24, vBDK21]. deregulation [BO86a, BO86b, CN86, Hüb96, Kuh88, Rue86]. Derek [Whi94]. derivados [Ano94t]. derivatives [Fab96]. derived [Ano94t]. description [WHP85, vDvM89]. descriptions [Vos05]. Design [BCM^{+}22, Hay86c, PDHW23, RkhR97, Why05, dV94, AIOA00, Ano16h, AS18, BBS22, BNS16, BM94a, BM94b, CLZ23, CD92a, CD92b, CS08a, Can16, CBF24, CC05, CCF23, CT02, CG04a, CA07, D’I14, DC22, DQ06, DC00, Dow93a, Dow93b, Fis89, GC99, GKSW19, Hay89d, HR07, HBG12, II92a, II92b, JWY23, Kuh88, Lag17, LS23, LOM14, MG21, MTY02, Mac88, MMCD23, Mou15, Mou17, OZ06, Pur92, RH23, RNJ^{+}23, SY14, SP96, SL14a, SSS17, SCZ14, Sna90a, SL14c, TP24, TCT05, Ugh88, WLG05, YXH21, Why02, Hay86c, Ram86, Why02, Why05]. design-driven [SCZ14]. design-manufacturing [AS18]. designers [GFvT22]. Designing [FDN^{+}23, Huc85, Mer22, BCM^{+}22, II01, CCR^{+}23, Gri98]. designs [NBA22]. desk [SL14b]. destruction [RPGP23]. Detecting [SKTM08]. detection [CZL^{+}24]. determinant [Car98a, RSP10]. Determinants [BSN21, CG05, EHSB24, EN07, Fuk09, LC22a, LL08a, LL03, PCWG04, RH21, SSW22, SD10, Vit90, WOL05, WPF06a, WPF06c, YVG24, AL05, BPU21, BMLL07, BM07, Coc12, FB91, Gar87, JLL10, MP09b, NB09, Pal06, Sou01, VLVH23, VM11, WPF06b, YMK^{+}21, HCL09, WPF06b]. determine [Jai23, dOK12]. determined [RV21]. Determining [CYJ07, Ho09, Lin10b, MS19, GL22, PS04]. Deutscher [Hay96a]. Devanna [Mur91]. develop [ES21, Guy95, SVLYT24]. Developed [Lop06, AHSS19, CWAB09, Had99, Hos03, LH14, VGQK^{+}10]. developers [BBTW00, Dan19, BBTW00]. Developing [BBDM01, DO03, GR21, IDMD04, IAO^{+}02, Lop06, MG98c, MR00, SPP98, Tom90, Vos05, Alc99, Ano94-34, Bas01, Bie91, BBHA01, Bro94, CS02, CvZ05, CWAB09, CS89, Da 01, Den94, ETOA00, Fan06, FAH08, Hos03, KL87, Liu93, MH05, ND15a, OOL06, Pan98, Sar08, Sto99, Sto03, SH99, SD04, TC97, TPB12, VGQK^{+}10, YH09, YA11, Ano94j, NHS24, Joh92b]. Development [Aug09, BAWK12, CTYH02, Hob87, JY15, KBS01, LL04a, LH11, Pap06, Que06, Ros89a, SH99, WMGB01, Woo85, AY05, AM23, AAGL23, AOO02, AA90, ABLK07, Ali86, AW12, AAN06, AD09, Ano86d, Ano93n, Ano94y, Ano94-31, Ano94-32, Ara08, AN16, Bad11, tBL86, BYCS00, BBTW00, BM23, BSN21, BKM10, BK01, BF97, BAR16, Big13, BDL92, BH03, BBHA01, BM94a, BM94b, BLL95, BC82, BH15, BRP94, CDD^{+}11, CDD^{+}14, CGM07, CAdNFP04, CJ98, CvZ05, CPSS06, CS06, CBGvdV22, CE89, CL98, CHT99, Chi81, Chi05, CBK03, CN90, CN92, CL24, Dan19, DNT98, DBV97, DNP01, Dow93a, Dow93b, DH06, DR99, Dre01, DG02, EKMS12, EW06, Fah05, FMOPSVA21, For90, For91, FB05, GAS96, GBFvdS11, GFK12, GdWNPC03, GFvT22, Gri06, GL10, HK02, HN92a, HN92b]. development [HJ02, HBvdB09, HG92, Hob95, Hol94a, Hol94b, HKW^{+}22, Hos99, Hos00, How94, HT12, HLC05, HKR^{+}22, HKL24, II01, Ilo94, IOS00, INO^{+}03, Ish89a, Ish89b, IMIY87, JLL10, JvZ08, JKF19, JR91, KdOL08, KI03, Ket09, KWE05, KvdEV01, KM11, Kod05, KLH05, KS04b, KS04a, KT04, KFL11, KS04c, LB02, Lar09, LOH91, LLI06, LM14, LY00, LLS00, dLP06, LC07, LVHR10, LW11, LD95, LG14, Lic11, Lin10a, LCT05, LTHS12, LZD24, LCWC09, MCGLA02, MOWGB02, MH98, MMMT12, ML05, MR05, MCT06, MR00, MSR10, MII90, Nob04, OW93a, OW93b, OOKBD14, OH04, OOMW03, Ola98a, Ola98b, Ott03a, Ott04b, PLG24, Pau93, PP09, PS03, PV00, PMN^{+}24, Pfi99, PFP06, PW99, Pri17, Puj06, Pur91, RG10, RMGDLC08, RR11b, RG85, Sae05, Sæt24, SW21, SAWT06, SZLR19, SP03, SBOB23]. development [SS23, SM97, Sch85, SCH93a, SCH93b, SM95, Shu22, Sik91, Sik96, Sik97a, Sik97b, SJ21, SZSR23, Sou83b, SCS06, STH06, SW05, SM12b, SN92, TC97, TWGB02b, TW02, TWGB02a, TRW23, TWBC21, Tho88b, Tho88c, Tik23, Tsu01, Udd06, VvWM20, VB05, Wal07, WTJ23, WC07, WWM00, WN03, WAN03, WA04b, WH04c, WLO09, Web96, WM88, Whi89b, WFW02, YIKC21, ZI01, ZCHC23, ZXDG19, dAASA21, dGSJ^{+}21, dOK12, dV94, dB10, vdVdM11, vdEJG08, vH10, vK11, Ame95, ETA02, FN21, MOM08, Coo07, Lal00, Leh07, Dod87a, Hay86c, Hay88b, Hay90c, Hay92d, Hob87, Sha91a]. development-based [GBFvdS11]. development-lessons [KT04]. Developmental [TAM11, Ano93p, Lev93]. Developments [MS99, Mas23, MBA^{+}12, Pis89, Pro21, SD95, vdVMM09]. deviance [Min96]. Device [WZNG17, FC06]. devices [BBZ20, Cha22, JFCFMC^{+}12]. devoid [PKH15]. Dfl [Bes88, Rot87a]. Dfl.195.00/US$105.00/UK£ [Ano97b, Hay97d]. diagnostic [SSKB24]. diamond [Tit90b]. Diane [Min96]. Diego [Leh07]. dietary [II91]. difference [DDG98, LV09, NP10]. Differences [GBFvdS11, LLS00, LSB07, Neu12, CS08a, CF01b, Fig02, LT16, OÁVH14, PSdLC02, PSdLCJ05, SSF12, Sap92, WW21, ALCLdF^{+}15]. Different [DTI^{+}23, LT05, LO96, NL11, TS13, BRCG12, Car01, CGLSNLDV15, DS23, HSLL22, JJS^{+}18, Kaa09, McQ05, NSB23, RF12, Sna98, TLK09, WJK15, Wür17, ZIB23, ZLJ^{+}21, vGS09, GWZ^{+}23, SLJ13]. differential [CL22, GdWNPC03, Lee11, SM12a]. differentiated [dVdWNF^{+}10]. differentiating [FWL06]. differentiation [And21, SP96, VM12]. differing [HW18]. difficulties [Lin08b, Sal03]. difficulty [BAWK12]. diffusing [Hay88a]. Diffusion [Ros82, SFC14, AH85, Alc99, Ame95, Ano92c, AR99, BBA15, Bas04, CT92, CLP23, Car98c, CT06, CS05, Che12, Chi91, CDK^{+}23, Da 01, FGTB23, HY99, Hu13, Hüb96, LL08a, MW09, Mol97b, Mol97a, STA24, SI98, Sap89, Sgo95, Se14, Tsa13b, TS05, Udd06, WMGB01, WKN03, WA04b, Hay84]. diffusion-oriented [Chi91]. Digital [BKA^{+}23, BKP23, CCP23, Fel22, FGB^{+}24, GHHH22, KS04a, LVR23, LDMS23, MM24, Mas23, MRDA24, RNJ^{+}23, SSCH23, Sap05, BTP^{+}23, BK24, BBLW24, BF21, BCP22, Bre23, CBGvdV22, CW22, CDK^{+}23, CCR^{+}24, DMRS23, DP23, EFC23, FBBBM22, FPP24, GPCNS^{+}23, GR21, GHU21, HPMI^{+}23, HMFS24, How24, JWY23, JSP22, Kan22, KRK^{+}22, KM22, KLP24, LFJA23, Li21, LP03, LZD24, NUSR21, NVMD22, NLC98, ODM23, Oll08, PGB^{+}23, RDAV23, RNFG24, Ran24, RUNS23, STA24, SVLYT24, SLA^{+}23a, SHS^{+}23, SOA^{+}24, TFS^{+}22, TPBM23, URG^{+}24, WTJ23, WLTH23, WKN03, YVG24, ZLS23, Tan08]. digitalising [BN06]. Digitalization [NSB23, AS21, KLP24, PPM24, SPPG20, SXS24]. digitalize [CCR^{+}23].

Digitally [CKS22, RES^{+}24, Chi21, KSHP21]. Digitally-enabled [CKS22, Chi21]. digitization [KK13]. dilemma [AFL^{+}24, PSS23, TE14]. dilemmas [SCM24]. Dilić [Sha95]. diligence [Box01]. dimension [AAA02, EIMS01, SAWT06]. dimensional [KK23]. dimensions [BH16, Bje10, EHHW17, Kaa09, NL11, TRR^{+}24, Ulh05, XHM^{+}16, Dib02b]. Dimitris [Ano94-34]. Ding [Arr04]. d’innovation [Ano94-36]. Direct [BLQS00b, DLC21, Efr14, LYY^{+}22, SHT23]. direction [WT19]. directions [BDNP06, DBR24, Lin02, MMMD23, RS14, SdVGdFB24, XL23, XLD23]. directiveness [VvWM20]. directors [WW14]. directory [Hay93c]. disabilities [BV24]. Disassembly [dRP95]. disasters [McB13, Min94]. discarded [SL98a, SL98b]. disciplinary [LP06]. discipline [Dre97, Sza90]. disciplines [Bar14]. disclosure [XW21]. disconnects [CJ98]. discontinuance [AH84, AFL^{+}24]. discontinuities [Ehr95, McG98]. Discovering [DHT22, SKL17, LYP09]. discovery [CBF24, DF98a, DF98b, Fel23, JB19, RZ21]. discussion [ALCM14, Ger04, Haa14, Nwo02]. disease [RNJ^{+}23]. Diseconomies [MC00]. disembodied [KL04]. Disentangling [LK20, SA22]. dishonesty [Lin14]. disintegration [FL15, Lin18]. disintermediation [Lin18]. disintermediation/disintegration [Lin18]. disk [KM03]. Disparities [Pap06]. disparity [ZHY22]. dispersed [CMP^{+}06, ZDWG14]. displacement [ND15b]. dispositions [YMMYK24]. dispute [Per95]. disruption [HGR15, Kra22b, LVR23]. disruptions [GP22]. Disruptive [BTP^{+}23, KKWM02, RHW14, VNP^{+}24, WWY15, YH11, ZZXW23]. disruptor [PSS23]. Dissecting [CS01b]. distance [BG19, DDN24, Pet11, ZHZL21]. Distant [BBA15, EH14, Wan24, WG23]. distributed [CCR^{+}23]. distribution [Ano92c, Arc88, AB89, BRA^{+}23, CT92, CSL06, SJSP23, SJP23]. distributors [HESPSL11]. district [DGP10, GP04]. districts [AGS98, Arc03]. DIVA [BBZ20]. diverge [CWAB09]. divergent [Jol08]. diverse [Bar14, NS07]. Diversification [ADD19, BF03, BF05, CL21, CL22, Lic05, ZL24, Dib04]. diversified [Lic05]. Diversity [CGM22, Hob95, CL97b, CL97a, HC10, JWY22, MZG17, OKP13, PWA11, WvdWG24, WQZ21, vRGJ^{+}12]. divestiture [Mow88]. divide [HBO14]. divided [Mas86]. Diving [MDT^{+}23]. Division [Hay84]. DNA [Bar14, Lin19a, SO06]. Do [Bal10, CZZM21, FBBBM22, GF21, HFWW16, HLL12, Jel07, KLRR22, LLZ19, MMC^{+}23, ODM23, SCH19, TLK09, Aha88, BMW14, Ban04, BSWC23, CCB20, Car01, CCC^{+}12, CH01, EHHW17, FCCE21, HSLL22, JP96, JLN22, De 23, Lin10c, Lin12b, MM11, NK22, SVLYT24, Ser13, TNY^{+}22, WSF13, XLG^{+}23, ZWDZ21, vBDK21, PLC13, SDK24]. Do-It-Together [FBBBM22]. Do-It-Yourself [FBBBM22]. doctor [ZWZ^{+}23]. doctoral [SBF15]. Doctorate [KI03]. Dodgson [Hay95, Jel07, Won08]. Does [BCO14, CWAB09, CJ23, DS23, Mar03, Paf04, SBF15, SAM^{+}23, SJSP23, SJP23, THC^{+}23, WTJ23, Wan24, Won05a, ZYZZ24, ZL24, BBT19, BF23, Ene22, Gar95, HDE21, JL24, LWP14, LKCK22, LW10, Lin11d, LLZ23, MR19, TKYM24, WLY14, WHWS22, NP10, ZCR22]. Dollar [DMH22]. Domain [KHKL23, NJ04, Wya04]. domain-based [NJ04]. Domain-specific [KHKL23]. domains [IM10]. domestic [AAZ22, CD00, LW97]. dominance [dHOvdKS16]. dominant [Ano16h, BNS16, NBA22]. Dominic [Har93a]. Donald [Hay88b, Hay88c, Per04]. Dong [Kap01]. Dong-Sung [Kap01]. Don’t [vRE21]. doom [De 85]. door [FGPP22]. doping [FLKF^{+}06]. Dordrecht [Hay97d]. Dorothy [Bes01b]. Dotlich [Fra01a]. Double [HWC06, ZSX24, WGW22, XWZ23]. Double-edged [ZSX24, WGW22, XWZ23]. down [vZS14]. down-and-out [vZS14]. DPD [Ott04b]. Dr [Hay86a]. Dr. [Hit89a]. draft [HAH84]. Dreamliner [SS16]. drift [Gra01]. drive [KM03, MWC22]. driven [BCO14, BCPP23, Bes03a, CGS^{+}24, CGW05, FPP24, GC99, JLY23, Joh92c, Kim16, KW05, KL05, LG14, NSYL18, PGB^{+}23, RS14, RKFR24, SCZ14, Lin90, Lin91]. driver [NWGB09, NHT23, OW00b]. Drivers [AL11, Hop04, RPL10, Cav16, DGS23, HKL24, KMID11, MNL22, OGG06, PUFC23, RDAV23, HO11, JPBL^{+}23]. drives [SPFG12]. Driving [MW06, PRA22, AAZ22, CDK^{+}23, HF18, Ved97]. drone [KKP17]. Drucker [Hay86e]. drugs [CAdNFP04]. DSP [NLC98]. DTA [BLQS00a]. Dual [MGS99, CJL23, LFU23, Wat90]. Dual-Use [MGS99, Wat90]. duality [ECB22]. Ducati [MGdG11]. due [Box01, NWGB09]. DUI [DS23, WSH23]. Dumbleton [Dod87a]. dumping [LL22]. Dunkel [Hay96a]. Dunleavy [Tan08]. durable [Bar89b]. durables [BF04]. duration [SYCH23]. Durbin [PGLS21]. During [Pap03, Ano94a, ABAHP^{+}23, BKA^{+}23, BEE24, JSR19, KRK^{+}22, MWJ23, RGSK94, RG06, Tei22, TFS^{+}22, Vit90, dHOvdKS16]. Dussauge [Bes94c]. Dutch [Bes95b, Ano82c, EG05, Pra86, SOKE15]. DVDs [Gal12]. Dyer [SD89b]. Dynamic [FCM21, HW02, Kha99, Ott03a, Ott04b, WA03, AS18, ALMW21, Ban03b, Bar92a, BK01, BR16, CPP09, CCMW23, Chi21, DJ05, EB23, EWJ09, ES21, EEY22, FN21, GHHH22, HDE21, LT09, LT21, LH21, OB04, Pet98, SPO08, SR23, SGF02, SSRO24, TSA13a, TKT13, Won06, YXH21, BBZ20, VSE24]. dynamic-capability [EWJ09]. dynamically [CCF11]. Dynamics [CT12, NMN11, OOA04, Sap05, WvdVM07, Won08, BC10, Bas04, BKP23, Bes96, CR00, DGP10, DV23, DDOR08, DM08, ECB22, FORA17, Gee06, Gra01, Hal14, Har93a, HCMMH21, Kap03, Koc04, LKKH22, LTH06, MTU23, MBA^{+}12, OW00a, PS03, RES^{+}24, Ros82, SGB12, WL21, WBM22, XPC21, ZC22, dV94, vdVMM09, GUMZI14]. dynamism [CCGK23, CGLSNLDV15, Ran24, SCM24, TLL23, WWM00, WA04b, YXH21, ZMJM23]. dysfunctional [GSFS21]. dysfunctions [ML05].

E-Business [Coo07, LO10, LL08a, Moo03, DSS05]. e-clusters [AM10]. E-Commerce [Kuc07, LL02, LLEB05, WPF06b]. e-government [NDP07]. E-health [BSMC23]. e-Healthcare [CWY14]. E-innovation [LD02]. e-procurement [CP05, GGS08]. E-product [HJ02]. E-Quality [Hal97a]. E-readiness [FAH08]. E-tivities [New04]. e-trade [OCA09]. e-trust [YPF03]. e.g [CE89]. each [Wür17]. Early [MPCRB^{+}03, CKR00, HKW^{+}22, LXODT15, MTB01, MNL22, PS03, SOKE15, VP14, WKF24, dHOvdKS16]. early-stage [CKR00, MNL22]. Earth [BBLW24, Kha98d, Kha98c]. East [HCK01, Hos99, Kap01, BWF18, DNS^{+}03, HG92, Hob95, HCK01]. Eastern [CP05, Cha99, Mar01, Szá92]. eat [AT99]. EC [Mer94a]. EC4P [Gar95]. Eccles [Hay93d]. Eco [LB02, Puj06, YWL22, BRA^{+}23, CS12, DGS23, GGGR22, HMWT23, KM06b, LC22a, RA17, Mer94a]. eco-friendly [LC22a]. Eco-industrial [LB02]. Eco-innovation [Puj06, YWL22, BRA^{+}23, CS12, GGGR22, HMWT23]. eco-innovations [DGS23]. Eco-management [Mer94a]. eco-non-profit [KM06b]. ecodesign [Ano94k, Roy94]. ECOLEAD [CMA07]. ecology [FFJ92, XPC21]. Economic [AT99, BK00, DGS23, GUMZI14, Kuc06, Lal00, Yam21, AIOA00, AA90, AZ90, Bar96, BF03, BF04, BKM10, CP05, CG11, Coc12, Coz06, DLR07, Fon05, FGS07, GAS96, GGS08, Gol87, HN92a, HN92b, Hay86c, Hay88g, Hit93, Hos99, HPT10, JRS06, Kra22a, KS10b, LB87, Lin17a, LOM14, MY02, MRC14, NWGB09, OW93a, OW93b, dAOCM02, Pap02a, Per97a, Ros90, Rot87a, SD10, Wal85, WC07, Wil83, Won06, Aug09]. economic-probabilistic [MRC14]. Economics [Dib00, Que06, CH01, CK04, De 04, Gal06b, Gar87, KM03, Per04, SHS^{+}23, Smi85, Pál04]. economie [Win87]. economies [Ano93l, Ano94-34, AHP16, BJB^{+}24, CvZ05, Da 01, Den94, Gru93, GG11, HO14, HW18, Hos03, LL02, Man07, MH05, PGKF19, PBC93, Sto21, Rad97]. Economy [CMH14, Lop06, AAA02, BCS06, BD23, BG19, CS01a, CS02, CPSS06, Coc12, CHO98, Cre04, De 85, IIA97b, IIA97a, IAO^{+}02, Kod03, KL08b, LSB07, LT21, Öbe23, Oye01, RB09, RM09, Rag93a, Rag93b, RES^{+}24, RR11a, She09, Shi24, TPB12, Tho99, TSA13a, VRS^{+}22, XPC21, Zho92a, Zho92b, dAASA21, Lae04, Par06, Sap05, Sap10, Hay01a, Pap04]. economy-Evidence [KL08b]. Ecopetrol [GAS96]. Ecosystem [HS21b, BSG^{+}24, Bes15, CCR^{+}23, CKF^{+}24a, GSFS21, dVGCOF^{+}23, GKSW19, KSHP21, KBSB22, Li09, MTU23, PE22, RUNS23, RA17, RNJ^{+}23, SFdO^{+}23, SdVGdFB24, SFBK23, TAG22, YCL22]. ecosystem-as-coevolution [HS21b]. Ecosystem-as-structure [HS21b]. ecosystem-level [RUNS23, SFdO^{+}23]. Ecosystems [Bai18, AG24, ABSZ23, ABG23, BCV21, Bre23, CS21, DLTQ23, FN21, GUMZI14, GH21, Gra24, HCMMH21, LYST22, LG14, MPO23, MTU23, NBA22, OPPL16, Ove15, Pro21, PGB^{+}23, Ree99, SLA^{+}23a, SOA^{+}24, SSRO24, VLVH23, ZLS23]. Ed [Aug09, Aug06b, Col08, Dib02a, Fra01c, Fra02b, Gof07, Kap03, Lae04, Pap06, Par06, Sap01, Bes94a, Dod88, Gar96]. edge [Bro94, JPBL^{+}23]. edged [WGW22, XWZ23, ZSX24]. EDI [JMPR01, JMPR04]. Edited [Ano91c, Fra03b, Lin90, Lin91, Ram86, Sha91b, Ano91d, Arc03, Har93a, Har93b, Hay86c, Hay87d, Hay90a, Hay91a, Hay92c, Hay93d, Hay93c, Hay95, Hay97d, Hob87, Hol93, Joh92a, Joh92g, Que06, Rad97, Rot87a, Sha95]. edition [Bai84, Bes94a, But96, Hay84, Hay85c, Hay88d, Hay93c, Per97b]. editor [Bes97, KV24, Ric95]. Editorial [Ano89n, Ano99d, Ano00a, Ano03p, Ano11g, Ano18b, Ano18c, BB81, Bro83, Hay85e, Hay85f, Hay86g, Hay86h, Hay86i, Hay86b, Hay88h, Hay88i, Hay89b, Hay90e, Hay91b, Hay94a, Hay94b, Hay96b, Hay05, LVVH23, Lin09a, PL21, Ros85, Ros87, Ros88, Ros92, Ano88j, Ano90l, Hit92a, Jen91, Ano81c, Ano83b, Ano85c, Ano86b, Ano86c, Ano87c, Ano87d, Ano88f, Ano90j, Ano91f, Ano92d, Ano93d, Ano94h, Ano95m, Ano96m, Ano97o, Ano02d, Ano02c, Ano02e, Ano02f, Ano03d, Ano03m, Ano03n, Ano03o, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano04c, Ano04l, Ano04m, Ano04n, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano05a, Ano05j, Ano05k, Ano05l, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano06b]. Editorial [Ano06j, Ano06k, Ano06l, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i, Ano07a, Ano07h, Ano07i, Ano07j, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano08b, Ano08j, Ano08k, Ano08l, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano09a, Ano09i, Ano09j, Ano09k, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano10a, Ano10h, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano11b, Ano11i, Ano11j, Ano11c, Ano11d, Ano11e, Ano11f, Ano11h, Ano12c, Ano12j, Ano12k, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano13a, Ano13f, Ano13g, Ano13b, Ano13c, Ano13d, Ano13e, Ano14a, Ano14i, Ano14j, Ano14k, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano15a, Ano15b, Ano15c]. Editorial [Ano15d, Ano15e, Ano15f, Ano15g, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano18a, Ano18d, Ano18e, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano20b, Ano20c, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano21p, Ano21q, Ano21r, Ano22m, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g, Hay89c]. Editors [Lop06, Hay97c]. edn [Ros89c]. Edquist [Dib01]. Eds [Ada07, Ano97b, Arr04, Bas04, De 04, Dib00, Dib01, Dre04, Dre06, Edw04, Hay01a, Hay01c, Lae03, Leh07, Pap02a, Par07, Que06, Rea06, Sap05, Sap10, Szw07, Wya00, Hay96a, Pap03, Pap04]. Eduardo [Que06]. Educating [CR81, Ano81a, Rad97]. Education [Ano81a, LC82, Why02, YKW10, AA91, BBTW00, CEH03, Kod00, LJ01, OR06, RS06]. Educational [Her92a]. Edward [Arc03, Arr04, Aug09, Ban04, BH03, CF07, De 04, Dib00, Dib01, Dib02a, Dib04, Dre06, Fra01b, Gra04, Hay95, Her04, Lae03, Lal00, Mor04, Pap02a, Par07, Que06, Sap05, Won08, Wya07]. Eesti [Tõn15]. Effect [ANI^{+}22, MPM09, WKN04a, ALMW21, BSWC23, CHD18, CYHC19, xCZZ^{+}22, CU13, FDVB10, FGPP22, GMNL07, HN08, HR07, HHH09, ILR^{+}12, JCT23, LLR95, LWL^{+}22, LTH06, MLC24, MB04, NW03a, OOKBD14, OBMOM24, PGLS21, PH08, SHT23, ST21, SYCH23, SH10b, VTRM12, WCW^{+}14, WF10, XWZ23, YK08]. Effective [CB19, Bar92b, CO04, DO03, Fru92, Han95, JBL23, KT02, Mer22, PC21, RES^{+}24, WH04c, Whi94, Bes95b]. Effectiveness [VLIC12, Ano94-36, Ano94-41, BC92, BH15, DPH^{+}22, DB97a, DB97b, Gup96, LD05, LbKL91, OC85, SW21, WBM22]. effectiveness- [WBM22]. effectivness [Sha03]. Effects [Kaa09, WZSZ16, WJK15, ZQW^{+}19, ZHZL21, Bar15, BF21, BR16, BO86a, CZM23, CL21, CL22, DLC21, DVdCAS16, DVSL19, Ell81, GAS96, GdWNPC03, GGJ18, GGJ22, HC13, KM11, Kra22a, LD05, LT05, Lee11, LLB89, LV09, LT10, LZLL21, Mow88, Mus09, Pat24, PSG22, RATW15, RMGBA08, SLSCS24, SY05, TTC06, Tsa13b, VRS^{+}22, WCZZ14, YXH21, ZCF19, ZCR22, ZLJ^{+}21, dMD05, GGGR22, WH21]. effectual [KBO22]. Effektivität [Ano94-41]. efficacité [Ano94-36]. efficacy [CJL22, GBV23, LCWC09, MAS13]. efficiencies [KK06]. Efficiency [Gru93, Hit04, ADL07, Ano94r, Bar96, DB97a, DB97b, HFWW16, Kaf06, Kli06, KPK94a, KPK94b, LH11, LvLB06, MLC24, TWGB02b, WBM22]. efficiency-concurrent [Kli06]. efficiency-oriented [WBM22]. efficient [CH01, HR07]. effort [JLL10, JRS06, NQ05]. efforts [KDH02, Wit85, XHM^{+}16, ZWZ^{+}23]. eficacia [Ano94r]. eggheads [LH89]. Egypt [Fah05]. eighties [Vit90]. EIS [Sal03]. elaboración [Ano94t]. Elaborating [Bar89a]. elaboration [Bar89b, Kli06]. elasticity [WK03]. electric [CK04, Kod07b, McG98, RHW14]. Electrical [Bes02d, KDH02, MW06, WA03, WN03]. Electricité [PR97]. electricity [Kap99, PR96, PR97]. Electron [Pow10]. electronic [CCR^{+}23, HT08, IKL01, LCLC03, Mil85, MB04, Oye00, YPF03, dRP95]. electronics [CTYH02, DNT98, Hay90d, Hob92, HCK01, Kha98b, Kha98a, LC22a, NB92, OW00a, Tsu02, Yeh13]. elements [CTC02]. Eleventh [Lop06]. Elgar [Arc03, Arr04, Aug09, Ban04, BH03, CF07, De 04, Dib00, Dib01, Dib02a, Dib04, Dre06, Fra01b, Gra04, Hay95, Her04, Lae03, Lal00, Mor04, Pap02a, Par07, Que06, Sap05, Won08, Wya07]. Eli [Fra02c]. Elias [Bon02]. elitism [Tit94c, Tit94d]. Eliyahu [Fra02c]. Eliza [Mur91]. Élizabeth [Ano93o]. Elsevier [Bes88, Bro07, Dod87a, Gof07, Leh07, Lop06, Rot87a]. elsewhere [Bai84]. Elucidating [LTH06]. emanating [RAC84]. embedded [MO03, PDHW23]. Embeddedness [Day02, GD08, LC09, LFFT09, LG22, PGB^{+}23, ZMH22]. Embedding [KL08a]. Embodied [KL04]. embracing [JJS^{+}18, LO10]. embryonic [VNP^{+}22]. emergence [FGPV24, HLC05, MJ94, PE22, RG06, RHW14, Sch92, TAG22]. emergent [Lee00]. emergent-technology [Lee00]. Emerging [Col08, IAA94b, IAA94a, Lin11e, Lin17b, WZNG17, ZAI23, AABY^{+}23, AAAA23, AM97, BJB^{+}24, CCF11, CCR^{+}24, FHBDO16, Gol86, GG11, HGB23, HN99, HR98a, HR98b, HO14, HW18, HC06, Jus24, KM97, KS14, KKP17, Li21, LWL^{+}22, MJ24, OM88, Ove15, PGKF19, RB09, RM09, RR11a, RvdSdE22, RB21, SKTM08, Sma01, THHS22, VRS^{+}22, YHcHW17, YHcHW21, YWL22, ZZXW23, vdVMM09, BVHJ24, CL24, NPG14, RRSS07, Shi24, vDZ21]. emission [AMS^{+}24, Boh18, ZZL22]. emissions [BAWK12, LVHR10]. Emotional [AKBA07, GP22, HPMI^{+}23]. emotions [Fel23]. empathy [NK22]. empathy-altruism [NK22]. emphasis [CS08a, Can16, Hit93, JWL24, KBRR24, Sha03]. Empirical [AAK^{+}24, MRGTDG21, NHT23, AC08, AF08, Ano94-36, Ano94r, Ano94x, Ano94z, AKKS12, AdA08, Aut94, BG10, BLA06, BSG^{+}24, Bro01, BF86, CLW10, CA07, CY00, Cor87b, CL88, CNPV04, Dag04, Die00, EFV02, FCS21, Fox86b, Goe10, GÁNA07, GPR89, GM00, GC10b, Har97a, Har97b, Hua09, HT08, HKJ07, IM10, Jac13, JWL24, JWY22, KS07, KLV06, KPK94a, KPK94b, KJ03, KMK03, LPS03, LPXO14, LRH22, LLB89, LH07, LH21, LCT04, LCT05, cLhWCH12, LL05a, Lüt04, LP06, Mac97b, MCGLA02, MOWGB02, MZL15, Mil94, Miy00, MMM05, NB92, NPV05, PMN^{+}24, PS04, PD10, RDAV23, RSP10, ROMH15, Sha03, SF05, Soh96, TCT05, TPH15, Tuz05, Ven96, WZSZ16, WMGB01, WKOW03, WH04c, WLO09, WWYZ23, WvdVM07, WH96, WJK15, XW21, XLG^{+}23, YGM99, YW08, AAZ22, AEP21, BGA09]. Empirical [BM16b, CKM13, HMWT23, KLRR22, KMID11, LXODT15, dCLSNL08, SH10a, WCW^{+}14]. empírico [Ano94r]. empirique [Ano94-36]. employability [AG24]. employee [ALM22, CLW10, Gil96, JCT23, LO10, PRID23, ST21]. employees [LO10, NKL02]. Employer [PG09]. employment [FV91, SG24, Tei22, Tet00, Dib01]. Empowerment [vAdMSL95, Sun96]. empresas [Ano94-28]. EMS [ZS04b]. enable [TWGB02a, Car98a]. enabled [AM10, AMS^{+}24, Bøl12, CKS22, CX24, Chi21, HMWT23, HMFS24, JFCFMC^{+}12, KSHP21, PRA22, PC18, RES^{+}24, ZXX23, YH09]. enablers [SPPG20]. Enabling [CW06, NBA22, AVA22, CFP^{+}23, HW18, YW08]. enactment [Jon05]. encapsulation [Gar05]. Enclave [Ada07, Ano94-42, Fel94]. Encountering [GP22]. encyclopaedia [Ano91c]. end [BK00, BK01, BV09, GSP^{+}22, KFL11, LHI13, MHKZ24, RGMW18, SFF21, Ste14b, TPV16]. end-of-life [BK00, BK01]. endeavor [WT19]. Endogenizing [BK04]. endogenous [BC12, LC09, MC00]. Endowing [TBSM21]. endowments [SXS24]. Energia [Tõn15]. energitical [SL98b]. Energy [CS89, CK04, AEP21, BK04, CU13, FGS07, HR07, JMK05, Lin17a, Pot93, Tõn15, TS05, Wal12, dVV16, LvLB06]. energy-systems [BK04]. enforced [CYK19]. engage [JLN22, TLL23]. engaged [FGTB23]. engagement [ABGS13, BAR16, Kan22, SG22, SDK24, TP24]. Engaging [MCH00]. engine [Ame95, WTAW97, WTN98b, WTN98a, Won06]. engineer [Ano81a, DFD^{+}91, Bes88]. Engineering [AAA02, AIOA00, Ano81a, Ano03q, Hay82, IIA97b, IIA97a, KM03, AAAF22, CF01a, CW06, Dow93a, Dow93b, Ern95, FD21, HPB03, dCHDAALMM07, Hay92d, Hen01, HMFS24, KDH02, MZL15, PPM24, Pur92, RAC84, ZZXW23, KI03, Sch94]. Engineers [Bes02d, CSL23, Hit93, Lin17a, OZ06, Rus94, MAU07]. England [O’H99, Sub05]. Englewood [Dod87b]. enhance [LLZ19, SAM^{+}23, ZYZZ24]. enhanced [DDN24, HPMI^{+}23, Nar89]. enhances [Gaz86, JCT23].

Enhancing [BBS22, IBPGML23, RRB^{+}22, Tei22]. Enlightened [Hay01b]. enquiry [Goe10]. Enrico [Oak97]. Ensign [Koc09]. Enststehende [Ano94j]. Entangled [JBL23]. entered [FL15]. Entering [Arr04, EFC23]. Enterprise [GK06, Ros89b, ZHY22, Bes03a, Cor87a, Der14, DG14, GZXX23, GWM05, KS10a, Pal93a, Pal93b, PRID23, PP02, PdW00, Tom03, Whi90b, YGW06, YGW08, Zaf81, Zho92a, Zho92b, Rea06]. Enterprises [HRHW16, AABY^{+}23, Ano94m, Ano94-28, Arc03, BT22, Bar97, CS00, Cor87a, Cor87b, CL88, Cul00, EDM05, ED87, FDN^{+}23, GC96, Hay90c, KdOL08, KV14, KMID11, KL08b, LL02, LQ94, LOCSBGMH16, MB04, Nwo02, Ola98a, Ola98b, OOJ^{+}00, OIN96, OOL06, PCG22, RATW15, RMGDLC08, Rie89, SPKK08, SPÁG22, Shi24, Sub05, Tõn15, Udd06, Wal12, YHcHW17, YHcHW21, wK98]. entrant [ZWDZ21]. entrants [PSS23]. ENTRE [TMG11]. ENTRE-U [TMG11]. entreprenenurship [Oak03]. entrepreneur [BEE24, HP09, JL96, KI03, MKL19, Ros89c, ZGSG23, vRE21, vWvRN17, DVSL19]. entrepreneur-coach [MKL19]. Entrepreneurial [CYM^{+}16, Fel22, FMOPSVA21, FGPV24, GP22, HL24, HW17, SFBK23, WW04, Ano81a, Ano93a, ABG23, BCO14, BDB21, BKP23, BCV21, BSWC23, Bøl12, Box01, CPSS06, CKF^{+}24a, CK12b, DH06, Far23, FCM21, FWL06, FL15, GT93, GMWK23, Hal14, HCMMH21, II01, KA15, KHL15, Lin91, LL03, LLZ19, LWM24, MSD15, McM96a, MR19, OR06, Pap02b, Par05, PDOL11, PD10, Pro21, SFdO^{+}23, Sch86, SP22b, TSA13a, TMG11, Tsa05b, Wür17, De 23, TBSM21]. entrepreneurial-driven [BCO14]. entrepreneurial-enabled [Bøl12]. entrepreneurs [Ano93s, CR81, DMNR12, FGOX23, HS21a, JLN22, Kan22, KM22, LVS10, Lan98b, Lan98a, De 23, LH89, MT08, MG98c, Min95, OIOA01, Rad97, Rei97, SAY09, SG93, Sch86, TNY^{+}22, TBSM21]. Entrepreneurship [Ano11a, Aug09, GK91, Lin06c, Pra10, SP98a, SP98b, SMGPP24, iSKB02, Wil83, Won08, AA90, AAD14, AP08, Ber96, BM14, CEH03, CvZ05, CS21, CL12, CJL22, CHT22, DLKM24, EFV00, Fel22, Hay81, Hay86e, HY04, LFJA23, LPXO14, LAC24, Lun14, MRGMBR13, MG98c, MPW03, Mos16, MR19, NGK08, Nwo97, Per04, RS06, Rei97, Seg86, SHS^{+}23, SSS17, SFBK23, SL16, SCH19, TEA00, TDA05, Ulh05, VM97, Wit85, Won13, YVG24, ZLS23, LWM24, Fra02b]. Entreprenuership [Ano11k]. entreprises [Ano94w]. Entropy [AAT14]. entry [CGM22, KLL87, ZWDZ21]. Entwicklung [Ros89a]. envelopment [KHKL23]. Environment [CLR96, BC10, Bar15, Bøl12, Car01, Chi05, DCD22, Dow93a, Dow93b, DM91, FC95, GC93a, GC93b, Kuh88, KS04c, Li05, MK00, SM97, Shi24, Ste14a, STH06, vW84, Kuc06]. environment-friendly [Chi05, MK00]. Environmental [BO86a, Her04, Kar93a, Kar93b, LL05a, Mac86b, Par07, PBOR23, ZS04a, AR99, BFM86, BFM88, CCGK23, CGLSNLDV15, HS21a, HC13, LH07, NG09, NBA22, OBMOM24, RVW17, Ran24, Ree99, Rot92a, Rot92b, Sar95, SCM24, SL98b, Tei22, TLL23, TCT05, VTRM12, XW21, YXH21, ZYZZ24, ZCR22, ZMJM23, ZS04b, dBJ02, dBJ03, Ban04]. Environmental-Organizational [BO86a]. Environmentally [BL97, Sto03]. environments [BBC06, BM16a, CA11, Car98b, HPB03, Hay91a, HKL24, KLV06, LW11, LH21, Mrk17, PS97a, PS97b, PS00b, PH08, SY14, SPLB19, VK03]. envisioning [KKP17]. ePD [HJ02]. epidemic [WKOW03]. epidemics [WWYZ23]. Epigenetic [GUMZI14]. Episodic [YLH22]. equal [ACA24]. equations [WKN03]. Equipment [Aka98b, Aka98a, WZNG17, ALRSD04, Bie91, Fan06, HO14, KSHP21, MW06, OTA01, Oye00, Sha86, Sha93a, Sha93b, TOA01]. Equity [Pow10, SSGN23, Won13]. equivocality [RGMW18]. Era [Tan08, CCBP22, GHU21, GP22, Pyk07]. eras [Lin11a]. Eric [Gra01]. Ericsson [RG06]. Erik [Hay97a]. ERIS [BK04]. Erno [Ano93n]. Ernst [Hay92d]. ERP [KMK03, LGY05, WH04a]. Errata [Ano90k]. Erratum [Ano84c, Ano87e, Ano89o, BP09a, YGW08]. error [TRW23]. ESG [KRM^{+}24]. ESPRIT [Bar95]. Essentials [Gun97a, Gun97b, Hay96a, Fra03a]. Essex [Lun88]. establish [SH06]. established [ALBO08, ALCLdF^{+}15, CCC^{+}12, Kar86, KAV17, MPO23, TCMS22]. Establishing [CCY05, HAA09, KS10a, Sza90, Tan95, Fuj98, HMM14]. establishment [CE89, Col81]. estate [DSVM05]. estimated [FNJV14]. estimating [KW05]. estratégicos [Ano94q]. estructura [Ano94q]. Ethernet [CCY05]. ethical [GKK23, JLWS23, SCM24, TKYM24]. ethics [Min94]. ethnography [Des96]. Ettlie [Pre01]. EU [Giu09, Gra04, Cai23, CP05, Gui11, HL24, MAKKM22, PWA11]. EU-China [Cai23]. EUREKA [Ano89p, Bar95, Pis89]. Europe [Ver90a, Ver91a, Ano89p, AC04, BDL92, Cha99, Kaa09, KV14, OM89, PBOR23, Rot91, Szá92, Tet00, XR08, ZIAO^{+}21, Sha95]. European [Ano81a, Hay91a, Per97b, AMV07, Ano86d, ABG23, BMW14, BKP23, BK99, CP05, Cor87c, CL88, CNPV04, FMF21, Hob92, Jol08, KA15, Kas94a, Kas94b, Lin86, LTVA22, Mar01, MRGMBR13, MPCRB^{+}03, PGB^{+}23, QGBV04, RBBB98, RD92, Sto21, TM03, lS06, Hay93d, Per96]. EV [Pot93, Ada07]. evaluate [Bar90a, JK03]. evaluated [MPBC21]. Evaluating [Hob88, MK87, OBMOM24, Swi92a, Swi92b, UB82, hWyLbC08, WHB21, Bar89b, ETM96, GLYP13, Kas94a, Kas94b, RZ21]. Evaluation [LPS03, AIOA00, Ano93e, Bar89a, CS87, CLZ18, CLZ21, CCST06, EM85, Gol87, INO^{+}03, MG08, NSYL18, OO02, PC18, Phi99, RCC04, Rya90, SPS01, WL91, YS93]. even [Pal06]. event [BCV21, BGM88]. event-study [BGM88]. events [EKT^{+}22, Sch86]. Everett [Hay84]. every [KS00, OWCvdB15]. Evidence [AABY^{+}23, ABS24, BCP22, CYHC19, CZM23, Der14, GGJ18, Kan22, KL24, Lee23, LBG14, MLC24, NC23, Que06, RIMP22, RDS17, SJ21, VRS^{+}22, WL21, YHcHW17, YHcHW21, ZQW^{+}19, ZMH22, AL05, Ano94-27, AAZ22, AES95, AEP21, BGA09, BKP23, BM16b, BMRG00, BA10, Bro01, CKM13, Die00, EFV00, Fan06, GUMZI14, GPR89, HMWT23, HST21, Kaf06, KLRR22, KMID11, KMK03, LXODT15, Lee11, Mac97b, dCLSNL08, Mon01, NHT23, PK23, PC21, Pat24, PS00b, Sán94, SH10a, SPFG12, SY05, SDFN01, Tuz05, Ven96, WCW^{+}14, WFWC21, WKOW03, AMS^{+}24, BRV23, BH15, CUF13, EHSB24, GST08, GG11, Hon21, Kaa09, KV14, KBC22, KL08b, LWP14, LLZ19, MPV13, PWA11, PWL20, PG09, PEER23, RH10, RAS15, SNM12, SPMRS23, SCH23, SFBK23, SCH19, SCY22, SD10, SG24, UF22, WWY15, WLTH23, WSF13, Yam21]. Evidence [ZYZZ24, ZL24, Oak97]. Evidences [dSSD09, ADNP01]. evolution [Ano94k, Ano94-35, BCKC24, BF04, BC99, BRCG12, Che94, DGP10, Dib04, EFC23, FJ91, GUMZI14, GP04, JLY23, JSP22, Kan94, KZ03, LM24, MdCÁ06, OÁVH14, PV01, RG10, RD92, Roy94, SH23, TWGB02b, WH04b, Wya07]. Evolutional [OW00a]. Evolutionary [SP10, BCV21, BCG01, FF92, Gee06, LH23, LG22, MZL15, RAATM05, SB91]. evolve [JL24]. evolved [SMD23]. Evolving [YHL22, CBGvdV22, SJCT24]. examination [ADDS09, HS21b, JWY22, LZLL21, MMM05, MSR10, NK22, O’K19, OPPL16, OWC03, Pia02, PS97a, PS97b, PS04, SSC23, Tsa05a, WAdN19, ZDWG14]. examined [LY96]. Examining [AM00, BJB^{+}24, KRM^{+}24, KL05, ST21, LT05, DC22, HPMI^{+}23, PDOL11]. example [CMA07, Che12, Lak08, MJ92, TS05, WC07, ZI01]. examples [Ger04, VS06]. exaptation [CAP23]. Excel [KK06]. Excellence [AHAAEA03, Can16, DMV^{+}21, Hit96, IBL04, Ahn95, Bro07]. excellent [Lin09d]. exchange [CE89, SLA^{+}23a, SR23, TI02, ZR93a, ZR93b]. execution [DP23]. executives [O’K19, O’K19]. exemplary [BP01]. exemptions [NKL02]. Exhausting [Pot93]. Existing [SD89a, DJ13, MPO23, NKL02, TTC06]. exit [VJ23a, VJ23b]. Exogenous [VSE24]. Expanding [JSP22, Ayr88, BS87]. expectation [SH23, WH04b]. expectations [Tet00]. expected [Meh88]. expedients [HB84]. Experience [AM92a, AM92b, AA90, BK04, CP05, Col81, CN86, DMNR12, DVSL19, FM07, HAA09, Hay81, HO05, JWY23, KM22, KM97, LP07, LCWC09, MK87, OZ06, Pia02, SOKE15, Ste89, SH99, TRW23, TI02, Ugh88, Van85, Wan23b, Won13, dV94, vRE21, Hay88g, Hay91a]. experience-based [JWY23, dV94]. experienced [RB21]. experiences [AKL05, BDW06, Cur89, Els95, Nar89, Nob04, BBC06]. experiencing [WKN04b]. Experiential [MJ24, TSPJ23]. experiment [Hay93b, Wit85]. Experimental [EKK^{+}21, CCF23, HST21, LC22a, MMR24, OGOF22]. expert [Gup90, Gup00, Whi94]. expertise [Ene22, Poy97]. expired [YSKL21]. explain [WW21]. Explaining [Aab09, CN05, GMNL07, KDH02, Ano94-44, Fig02, Ham94, KBSW08, KK04, LLY^{+}18, CS08a]. exploit [HV96, WQZ21]. Exploitation [HESPSL11, AO10, CM04, CK18, HF18, Kas94a, Kas94b, Lic10, NS90, OB20, RC23, ZCR22, Sco89, Mor91]. Exploitation- [HESPSL11]. exploitative [EHHW17, GSP^{+}22, MMR24, SCM24]. Exploiting [CS02, DDN24]. exploration [AO10, ÁRS17, BT22, GD08, GdWNPC03, HF18, HESPSL11, JCT23, Mou15, Mou17, PDHW23, RC23, SDN23, ZCR22, Ste14b]. exploration-based [HESPSL11]. explorative [MMR24, MSR10, RGH19, SCZ14, SCM24]. explorativeness [LWP14]. Exploratory [EHHW17, GSP^{+}22, Kum05, BCC^{+}11, BCP11, xCZZ^{+}22, DB09, Gal05, Gar05, Gui11, KDH02, KH06, Pow10, SJ21, SM01, XR08, Zak23]. explore [BJ05, LAYS19, WQZ21]. explored [LS24]. Exploring [AMA21, AMP22, BM23, Bar15, BWF18, BV24, BOBT22, CYHC19, CLW10, Che12, GKK23, HL24, KS21, KV24, LLEB05, LH11, MNW14, PC05, SPMRS23, SSRO24, WSN^{+}18, KW05, SRHL12, ALFM23, BSSF24, DVdCAS16, FEP15, LNB21, McM96a, NGK08, PCQ24, PC21, PSS23, SH23, VM12]. export [FNJV14, GM03, KM01, Klo87]. exporter [KB06]. exporters [CS00, LL22]. exsitu [EKMS12]. Extended [TB24, xCZZ^{+}22, Cul00]. Extending [HJM08, MGdG11]. extension [Fox86a, Old97, FFJ92]. extensive [DB19]. extent [EHHW17]. External [ES21, Rot91, Sou01, AMP22, BBS22, BM94a, BM94b, CKT04, CCV11, CJL23, DVHL22, FORA17, FDN^{+}23, GW10, GGJ22, KOX21, LW11, LPN22, LT10, LW10, Mac97b, MAMRRR24, PN11, Rie89, RDS17, RB21, SCH19, VM12, WZNC15, WH04b]. externality [DR07]. extinguishing [Poz86]. extraction [FOL^{+}03]. extranets [Car98c]. Extreme [EKT^{+}22].

F [Ano93f, Hay86e, Hay86a, Hay93a, Joh92e, Pap02a, Ros89c, SKL17]. F-term [SKL17]. F.Y [Pap02b]. Fabrik [KC08]. face [HKJ07, Nar04]. facilitate [Cur89, Fel23, HKY22, LBG14, WLY14]. facilitates [CYK19]. facilitator [NSB23]. facilities [DLL06]. facing [Hos99, JFCFMC^{+}12]. facteur [Ano94u]. facto [Ano16h, BNS16, Dan19]. factor [Ano94u, Ano94g, Ano94i, FRWT06, OÁVH14, SAY09, Tit94a, Tit94b]. Factors [Bar90b, DAOO02, HT08, KC07, Leh07, Lin11f, NDP07, OM00, OWC03, AM23, Alc99, BS05, CS84, CPXC02, CY00, CA06, DSS05, FORA17, GW10, HB03, IPP14, KWvdS21, LD05, LO96, LLY^{+}18, LLB96, MS19, Moh95, NG09, OTA01, ØD22, PC89, RFMC21, RI04, SSKB24, Ste14a, SW05, WNGB03, WHP85, WWH15, BL97, dLP06]. factory [MG21, NJ04]. facts [Gei97, SB91]. faculty [AC07, GC10a]. fads [KLL24]. Fai [Dib04]. fail [JP96, SC20]. failing [NF24, BH03]. Failla [Gra01]. fails [FM24]. Failure [VNP^{+}24, AKL05, ACA24, BM23, BSSF24, BJB^{+}24, Car98a, CT12, CCMW23, CU13, CTSJ24, FM24, GS10, dLP06, NF24, PPMC24, PUFC23, SMGPP24, SAM^{+}23, TRW23, TB24, Wan23b, Win87, WLG05, WC04]. failures [ACA24, AAK^{+}24, CMS23, DDBM23]. fairness [CKF24b, OGOF22]. Faktor [Ano94g]. Fall [Ano94-29, Gar94]. fallacies [GKB12]. fallacy [Tom03]. falling [ZIAO^{+}21]. Family [DVHL22, AIOA00, BM07, MAMRRR24, NSB23]. family-sized [AIOA00]. fantasy [BMV02]. far [WKF24]. far-future [WKF24]. Faraday [Rus94]. farmers [SO02]. Farnborough [Mil88]. Farrukh [Bes02d]. fashion [SH06]. Fast [OGS06, SC20]. fastText [LLL23]. favourable [FC95]. FDI [NB11, VRS^{+}22]. FDIs [AAZ22]. Feasibility [FGB^{+}24, DA01, WL91]. feather [HW17]. features [Fig02, RPASLM23, WKOW03]. February [Ano21t, Ano21s, Ano22n, Ano23n, Ano24h]. fecklessness [LZLL21]. federal [Ali86, Ano93l, PBC93, Ste89]. feedback [BGF23, Ott97, SCS22, SGWL16, WG23]. Feldman [Per04]. Felicia [Dib04]. Fels [Fel89]. Felsenstein [Ano94-42]. Female [AABY^{+}23, FWL06]. Fetter [Gar95]. fi [PMN^{+}24]. fiber [Web96]. fiction [BV24, FD21, MCA^{+}24, Sæt24, WKF24]. Field [Ano09l, Bes95a, Bad11, BJ05, CBF24, DBV97, GL10, Hu13, LS07, Lic11, Lin10a, Lüt04, Pat24, RvdSdE22, Sgo95, Soh96, Tho12, dB10, vdVdM11, vH10, vK11]. field-based [CBF24]. field-study [Pat24]. fields [HS19, Kod05]. filed [FMF21]. filing [LYY^{+}22, Llo07, VJ12]. Filling [BS24, PKH15, Bar14]. film [Tit90b]. films [FD21]. filter [Ano93e, YS93]. filtering [CJL23, JP99]. finance [ABS24, Oak97]. financed [Cor87c]. Financial [Hay01b, MT05, Bad08, BK89, CHL21, Gup00, HB89, HL22, Kap03, LC22b, Mac86b, PWF07, PS21, Poy97, SO20, XW21, Lin10c, Lin11f]. Financing [Wan89, Fre99, GGJ22, Let03, LLB89, PJ96, Sik97a, Sik97b, TPFD13, Wan23a, Won06, Won13]. Fincham [Poy97]. Finding [CT02, Li05, Lin08b, Lic10]. Findings [Giu09, Hol85, vAdMSL95]. fine [OOMW03]. Finland [Wal07]. Finnish [HGL98]. fintech [KP23, CL24, SO20]. FinTechs [AMA21]. firefighters [FOW97]. firing [McM94b]. Firm [AL11, CLZ^{+}22, FC95, Kog03, Lim06, SGF02, Aab09, AA88, AKBA07, AGS98, Ano94-27, AP08, AES95, ARR23, ABAHP^{+}23, AVA22, AC07, BGA09, BC10, Bar95, Bar15, BWF18, BMLL07, Boh18, BGR91, BF86, CK22, CLP23, CR00, CA02, xCZZ^{+}22, CJ23, CK18, CL22, CS13, CGLSNLDV15, D’I14, DTD13, DID21, DR88, DLTQ23, EB23, EFV02, Fal05, FCM21, Fig02, FCS21, Fre99, Gal05, GMNL07, GW10, GPR89, GGJ22, Hag90, HS91, HBS07, HF18, HESPSL11, HN08, HC13, HDE21, Jan21, JJS^{+}18, KP12, KJ97, KOX21, Kra22a, KKG07, LNB21, Lee11, LWP14, Lee23, LLB89, hLwChLbL04, Lic05, LLZ23, LC22b, Maj22, Mal87, Mal02, ML08b, MRDA24, MC00, MK96, NGK08, ODM23, Paf04, PR96, PSG22, PSdLCJ05, RF12, RB21, RGH19, Sán94, SRDS06, SLA^{+}23a]. firm [SF05, SY05, SJ21, Sla23b, Sna90a, Sna90b, SFVI21, TKYM24, Tei22, TLL23, TDC11, Tsa05a, Tsu02, WNM95, WOL05, hWyLbC08, WCW^{+}14, WLY14, WLTH23, WSF13, WH04b, WH04a, WQZ21, YK08, YI05, YW08, ZZL^{+}23, ZCR22, ZLJ^{+}21, ZXDG19, dVdWNF^{+}10, dlHBR22, KBRR24, LfSS19, Ver91b]. firm-level [BC10, BMLL07, FCS21, GW10, HBS07, ML08b, NGK08, SRDS06, TDC11, dVdWNF^{+}10]. Firmen [Ano94-29]. Firms [Coo16, Nuv07, Pra10, ALvKL06, AL05, And21, Ano94x, Ano94y, ALCLdF^{+}15, AP08, AdA08, ABUS05, AVA22, Aut94, BMW14, BYCS00, Ban03a, Ban04, BCS06, BSG^{+}24, BS06, Ber96, BF23, Bro92, BM05, BF06, BM07, BRP94, BOBT22, CS08a, CKR00, Car01, Cha22, CCC^{+}12, CHL21, CR06, CCV11, CCF11, CLR96, CY00, CBK03, CK12b, CKM13, DVdCAS16, DVSL19, DMW13, DLL06, DTI^{+}23, DH06, DVHL22, Dur86, Egb15, EW97, FKT95, FC95, FJ87a, FOW97, Fre05, FL15, GMLMVJ07, GST08, Gar96, GWZ^{+}23, GÁNA07, GGGR22, GM00, GM03, Har97a, Har97b, HdWN21, HN92a, HN92b, HW18, HY04, HPBP98, HST05, HHH09, HT08, HGL98, II01, IH07, Jar86, JMPR01, JR91, JJS^{+}18, KKWM02, KZ18, Kha99, Kim87, KJE96, Kra22b, KGK07, KK87, KFL11, LWC09, LDF^{+}21, LWP14, LAYS19]. firms [LDMS23, LL03, LL05a, LL05b, Mac97b, MKK04, MRGMBR13, dCLSNL08, MAMRRR24, MD03, MAKKM22, Moh95, MS99, NSB23, OSG04, OG05, OS08, OC91, Oak03, OCA09, Pap03, PA15, Per97a, Pfi99, PH08, QGBV04, Ram23, RWP19, RGFLRP21, Rot91, Rya96, Sae05, SC17, SCS22, SP98a, SP98b, SNM12, SVÁG07, Sar22, SU11, Sch09, SH10a, Sen85, SG22, SGWL16, Sna90a, Sna90b, Sna90c, Sna90d, Sna98, STS88, Ste89, SCM11, SH10b, Sun96, TWGB02a, TS00, Tet00, TM03, Töd94b, Töd94a, TA10, Tsa05b, VB05, VGAL24, Vel15, VTRM12, VX97, VRS^{+}22, WCT10, WSN^{+}18, WKN04b, WSH23, WvdVM07, WH96, WVG^{+}98, WGV^{+}98, XR08, YWL22, YZY10, ZYSL21, dAASA21, vdEJG08, JH18, Koc09, Aug06b, Cou90, Cou91, Gar85, Hay88e]. First [But96, Hol93, Coz06, Hay90c, Hay97b, IKL01, Lay93, RI04, Ruo05]. first- [IKL01]. fish [BM05, SI98]. fit [BBT19, BBLW24, LWC09, Lin10b, Lin11d, ZXDG19]. fittest [SBMP13]. Fitting [BN88]. five [Car99b, HL24, IH07, EWE22]. flash [WKF24]. flat [BM05]. Fleck [Poy97]. Flensburg [Dre06]. Fletcher [Her92b]. Flexibility [CGW05, KvdEV01, SP03, CCL^{+}06, KZ18, MAB^{+}23, YZY10, Hay97d]. Flexibility-driven [CGW05]. Flexible [GPR89, vDvM89, AGM89, Ano89p, Ano94m, Gup88, GC96, LQ94, San95]. floor [Gaz86, LB82]. Florida [Ano07k]. FLOSS [YA11]. flow [CYK19, CD00, FLH^{+}02, FGS07, Ott97, RAS15]. flows [Ano94p, Han94, VJ23b]. fluid [BBC06]. fly [Bes01b]. FN [Hob97]. foam [Sch09]. focal [DJ13]. focus [Hol94a, Hol94b, LB06, Lin11d, Nwo02, RUNS23, SR22, WC04, LGZ22]. focused [HGB23, Lin09b, TS05, Wec15]. focusing [KTE04, MOWGB02, McM94b]. folks [DTI^{+}23, LT05]. following [VJ23a, VJ23b]. food [ARvT04, IOS00, KB06, KRA^{+}22c, OTA01, RES^{+}24, SDT^{+}22, SK92, Tai98]. foods [MH04]. foot [SK23]. footwear [AS08]. Foray [Har93a]. Forbes [RL06]. Forbidden [Har97c]. force [Ved97, IMBH98]. forces [BCV21, HF18, PE22]. Forcing [LVHR10]. Ford [Sha91b]. forecast [SB02, Bai84]. Forecasting [LA15, WC06a, ADK08, HS14, Szá92, TCC^{+}22, Rus94]. forecasts [CYJ07, De 85, Ho09]. Foreign [EL93, CD00, EW97, FLH^{+}02, LYY^{+}22, LW97, MG08, NT16, YNS18]. foreign-market [NT16]. foreign-owned [EW97]. forerunner [POGBG06]. foresight [AJLR04, CAdNFP04, Lin11a, MCH00, Web99]. foresight-technology [AJLR04]. Foresighting [BS08, YA11]. Foreword [Ano89q, Rub89a]. forgotten [HOGGP11]. Formal [Sha93a, Sha93b, ZLH17, CH01, LS23, MNW14, SSF12, TLL23, XLC23]. formalisation [CKF24b]. formalised [Hos03]. formalization [WHWS22]. formalize [BMSJ24]. format [Dan19]. Formation [Hsu05, WKOW03, Yas05, Ano94l, CSH94, HBO14, HST05, KK87, TBSM21, WNM95]. formative [FGTB23, KV14]. formats [Dan19]. forming [CAI00]. Forms [CC06, LZD24, SF98]. formulation [Par00, PCWG04]. Forschung [Ros89a]. Fortescue [Hay92e]. Forthcoming [Ano88g]. Forum [Bai18]. forums [Haa14]. Forward [JSR19, CD00, DDFS23, Hel94a, Hel94b, LFJA23, SK23, Ser13]. Forward-looking [JSR19]. foster [Bal10, CB22, DVMS16, Per04]. Fostering [ALCLdF^{+}15, Bai18, Car99a, Hos99, Rei97, EL19, NHS24]. Foundation [Ano81b, CE19, LH19, Col81]. foundations [YMK^{+}21]. founder [Hon21, HLL12, SCC17, BDB21]. founders [FMF21, SG24]. Founding [VP14, SG24]. foundry [KGK07, KKG07]. Four [RKFR24, Ano93n, KA15, LWP14, Pau93, Wit85, MPO23]. fourth [PK06]. Foxall [Hay85a]. frame [RNFG24]. Framework [BMW14, PWA11, Tõn15, AT93, AHS^{+}23, AW12, Ano93f, Ano94-36, Ano94r, AVA22, BD23, BBZ20, BN08, BC99, Bre23, BD05, CPP09, CMS23, DMRS23, EKY24, FHBDO16, FJ12, GSP^{+}22, GFK12, Hos01, KSHP21, Kas94a, Kas94b, KGS^{+}24, KKG07, KPK94a, KPK94b, Lag17, LM14, LGY05, LO94, Li21, MAS13, MdVB16, MMMD23, MZL15, PB93, PS01, RS03, RBBB98, RGH19, SFdO^{+}23, SDT^{+}22, SPP06, SS22, SPLB19, DGMR21, SdVGdFB24, SSKB24, STH06, VMWV16, Ven96, WLH10, WSH23, WLG05, WO08, XLD23]. frameworks [vW88]. framing [MGAP23, RNFG24, Shu22, KLP24]. France [MP09a, CJ98, CAI00, CEH03, PR97, Vav06, WNM95, Hay87a]. Frances [Gar85, Har93a]. Francis [Bes02c, Hay83, Ram86]. Francisco [Fra01a, Fra01c, Mer01]. Frank [Dre04, Rea06]. Frankel [Hay92d]. Franklin [RA86]. Frederic [Ano94-30]. Frederick [Ano97b, Dod87b, Hay97c]. Free [Hay84, Joh92c, YA11]. free/libre/open [YA11]. Freeman [Har93a, Hay86c, Que06]. French [Ano94-36, Ano94u, Ano94w, Ano98o, Cho96, Gui05, KS21, PG09, VGAL24]. Frequency [SSGN23, ÖBA09]. frequent [Ott97]. friendly [BL97, Chi05, LC22a, MK00]. friends [SMW21]. front [BV09, GSP^{+}22, KFL11, RGMW18, SFF21, Ste14b, TPV16, Woo87]. front-end [Ste14b, TPV16]. Frontier [LPMV22, BC10, KBO22, Mat22]. frontiers [Ayr88]. fronts [SKTM08]. Frøslev [Sap05]. frozen [SI98]. Frugal [Cha22, OKvB23, WZNG17, BBK23, PRA22, RBBF23]. Fruit [Har97c]. fsQCA [DID21]. fuel [Lee04, Pil04]. fuelled [Mas86]. fulcrums [HS19]. Function [MYR^{+}21, BHPB06, CSL06, CA07, GK06, HW02, MH98, WH04b, WH04a]. functional [FDVB10, GdWNPC03, LR09, MAS13, MH04, MLC24]. Functionality [OOMW03, LC07, SAWT06, WN03, WAN03, WA04b, WLO09]. functions [LL04b, SW22]. funded [Ano86d, BRV23, SVV15, Wal07, YHDR08, ZLJ^{+}21]. funding [Ano93j, GGJ22, HSLS24, SR23, SvG15, Tit90a, Wil93, ZSMB16, ZLJ^{+}21]. funds [GKB12, LDB13]. Funnel [VSE24]. Fürst [Sha95]. further [Bro01]. Fusing [WLO09]. fusion [Ano12a, Cav16, CY12, FMvT00, II92a, II92b, Kod14, MII90, Tan87]. Future [CGS^{+}24, Hay01b, Lin06a, AGFdMCM24, AHS^{+}23, BTP^{+}23, BK24, BSMC23, Big13, CCY05, DMV^{+}21, DMRS23, DBR24, Fun06, GLÖS23, Gol87, Hui11, KA06, Lin02, MW12, MMMD23, MMMT12, MGAP23, NLLC04, RCC04, Ser13, SdVGdFB24, SCM24, Sun00, TSA13a, WML^{+}23, WKF24, XL23, XLD23, ZKW^{+}23, vDZ21, Gra03]. futures [BV24, DR24]. fuzziness [GSP^{+}22]. Fuzzy [Ste14b, ANI^{+}22, CCY05, FSA23, RGMW18, TPV16].

G [Ano91d, Bon02, CvZ05, Fra00b, Gar87, Hay92b, Hay92d, Her91, Joh92b, Joh92c, Mur91, Rad97, dC07]. G-RVINs [CvZ05]. Gagnon [Ano93a]. gain [Bro90, Dan19]. gained [Ste89]. Gains [BEH99, Fal05]. galore [LB82]. gamble [MAMRRR24]. Game [WG23, BHS05, Gra24, LRH22, LP03, Par03b, RBBB98, SGWL16, SCM11]. Games [FLKF^{+}06]. gamification [AGFdMCM24, BSG^{+}24, BJB^{+}24]. gaming [LDF^{+}21]. Gann [Jel07]. gap [CL24, MH05, Hal97a]. gaps [BS24, LZLL21]. gas [KI03, RFMC21, dSSD09]. Gassmann [Hop04]. gate [SY17]. GATT [Paw94, Sik94]. Gattorna [Gra03]. gazelles [SO06]. Gbegiri [AT99]. Gear [Ruo05]. Gear-up [Ruo05]. Geely [YNS18]. Gender [Sha95, GC10a, Mat22]. general [DP23]. generalism [CGM22]. generality [GFE95]. generality/specificity [GFE95]. generalizability [Lin20]. Generally [YH11]. generate [BC10, Wan24]. generated [Ano94a, Bøl12, RGSK94]. Generating [PR97]. generation [BDW06, Bro95, Car98c, CDK^{+}23, CKF24b, CZL^{+}24, HY99, HHV05, IKL01, Kap99, MRGTDG21, MOWGB02, McA04, MGS99, PK06, PCWV16, SY14, DGMR21, Tsa13b, WWM00, WF10]. generation-specific [Tsa13b]. generational [SFC14, Tsa13b]. generations [BRCG12, MV06]. generative [CGS^{+}24, SCM24]. generic [Ano94-33, CW94, HS91, MTY02, YGM99]. genetic [Thu05]. genetically [CYK19]. Geoff [Hay93a, Hob97]. geographic [DS23, RC23]. Geographical [dSSD09, Car04, Car05, FGPP22, GR23, Pet11, TKSM08, Ved97]. Geographies [Wya00]. geotechnics [HB84]. German [For85, Har93b, SH08, Ano94j, Ano94g, Ano94-29, Ano94-41, Ano94-45, Ano98o, AS18, Bac98, BSWC23, BV09, CR06, CL97b, CL97a, Ern95, GGGR22, Har96, HW18, HO05, KH06, KMID11, MK87, Sar22]. Germany [Hay87c, BM16b, BGR91, CJ98, CAI00, HG92, Lay93, LfSS19, MS19, MS99, SH10a, SPFG12, Ste89]. gerontechnology [CC14]. Gerrit [Hop04]. Geschka [Hay85b]. Get [DMH22, DR07, Lin12c, vWvRN17]. Getting [SvG15, Lin07a, Lin12d]. Ghana [AAK^{+}24, EF11]. Ghoshal [Hay93d, Per97b, Ver90b]. Giantommaso [Mug89]. Giaoutzi [Hay90c]. Gibb [Rot87a]. gift [Ano97b]. Gigabit [CCY05]. Gillian [Hal97a]. Gilly [New04]. Gina [Bes01a]. Giovanni [Hay85b]. GIS [Lan09]. GitHub [Far23]. give [KHL15, MCH00]. Global [LL98, Lin07b, OS98, Par06, WH13, WZGBA01, Arc03, Bar89b, Bar90a, Bes94c, CS01a, CvZ05, DR24, Day02, Hay01a, Hu13, IMBH98, Joh92a, KK04, Kum05, McM94a, NW03b, NW23, PK23, SS07, SC05, Ver90a, Ver91a, WNLY06, ZDWG14, ZL24, Min94]. Globalisation [YC05, AK00, Nar04]. Globalised [Pap06]. globalising [Pap04]. globalism [Mon01, Szá01]. globalization [Gui11, RK04, Dut07]. globalizing [MC00, SIOO01]. globally [GBL03]. gloCal [CvZ05]. goal [Ano94-38, BCH94, CDMV24]. goals [MT08, SDT^{+}22, SFVI21]. Godet [Rot88]. godfather [Smi07]. Going [BEE24, McM94a, Hay86d]. Gold [BCO14]. golden [Kap03, KHL15]. Goldratt [Fra02c]. Goliaths [CR00]. good [CL05a, GKK23, Kha97, Lin20, Tuz05, YK08]. goods [Bar89b, Lüt04, MOWGB02, MW09, SH95, WNGB03]. Gordon [Hay85a]. gorilla [BT22]. govern [HSLL22]. Governance [EW06, GGJ18, HCK01, ALvKL06, AN16, BS24, BA10, BH15, CBGvdV22, ECB22, HMFS24, JSP22, LH21, MGAP23, SK23, SR22, SS04, SS05, VOTH05, Won07, Wya04, Tan08]. governing [Sta98, WNGB03]. Government [Aug09, BO81, DS13, FJV96, GGJ22, HL22, Lay93, McN92, Ras08, RHW14, WLTH23, WM88, Won16, Ano82c, Ano94f, Ano94-31, Arn87, BC91, BC92, BP94, CDD^{+}11, CDD^{+}14, CJ98, CRKA98, CAI00, CLR96, CU13, CS89, CN90, CN92, GWZ^{+}23, Gru93, HFWW16, HC01, HHH09, Ilo94, JWL24, JJS^{+}18, KP12, LvLB06, LDB13, MCD06, NDP07, PCQ24, Rot92a, Rot92b, Ste14a, SJCT24, Tom03, Wal07, WKN04a, ZLJ^{+}21, Liu93, Won06]. government-funded [Wal07, ZLJ^{+}21]. government-sponsored [HHH09]. governmental [Cur89, Lee00, LLB89]. governments [Aha88, Ano94-34, Den94, dBJ02]. governs [GW10]. Gower [Fit00, Gra03, Mer02, O’H99, Tat04]. gradate [Hay97c]. grade [Ano94a, RGSK94]. Graduate [Ano07l, MG98c, NW03b, YKW10, vWGdG21, AG24]. Graham [Bes02a]. Grain [OO02]. grains [SK92]. grand [FFJ92, FJ91]. Granting [Rei97]. grants [HFWW16, WHB21]. Grasping [CKF^{+}24a]. Grayson [Hay93b]. Great [Tan08]. greater [Ano94-28, Ano94-29, Ano94w, WLO09, Ano94-28, Ano94-29, Ano94w, Bes15, Töd94b, Töd94a]. Greece [BMV02, Sou01]. Green [Box01, CCGK23, Lae03, And21, BSG^{+}24, BVHJ24, FJ12, JWL24, KRM^{+}24, LC22a, OBMOM24, SXS24, SZSR23, Sto21, TPBM23, THHS22, URG^{+}24, VGAL24, XL23, ZL24, dlHBR22]. Greenfield [Bes02b]. Greening [LV23]. grey [LM14]. Grinyer [Her91]. Group [Ano97b, Hay86f, JWS92a, JWS92b, ABGS13, Hua09, Nwo02, Woo87, ZMH22, KS04a]. groups [Hol90, TAC96, TA98a, TA98b, Bes01b]. grow [Mac05]. Growing [Kim87, CR19, Lin09e, LL03, Sae05]. Growth [BA89, Hop04, KFEC15, MO03, McQ05, Tsa05b, AAAF22, AVA22, AZ90, BCO14, BSSF24, CP05, CTSJ24, DMW13, Dib01, FNJV14, Hay86d, Hay88e, Hay01c, IH07, KK13, LAYS19, Lin17a, LLZ19, LOM14, ML08b, Møl91, NW23, OGG06, Pap02a, Rag93a, Rag93b, RPGP23, Sae05, SBOB23, SOKE15, SCC17, SBMP13, Ste89, Sur92, VNP^{+}22, WD06, WWM00, WLO09, Wil83, ZC22, vGS09, Col08, Per04, Ros90]. growth-boosting [SBOB23]. growth-enabling [AVA22]. GU11 [Tat04]. Guangdong [CYHC19]. Guerrieri [Arc03]. Guest [Ano90l, Hit92a, Jen91, KV24]. GUI [CAI00]. Guide [Ano89r, Bes02b, Bes02d, Box01, Gri98, Hay97c, Mer94a, Rus94, Bul97, Col97]. Guidelines [Ano07l, Ano81b]. Guimaraes [Ano93k, Leh07]. guns [Wat90]. Gupta [Ano93k, Ano94-37]. Gurdon [Ano93s].

H [Bes94b, Bes95b, Dod87a, For90, For91, Fra01c, Her91, Mil85, OH04, Pov97, Poy97, Ros89b, Sap01, Ver91b]. Haarder [Haa89]. habitual [HP09]. hacking [BSSF24]. Haiyang [Col08]. HAL [II21]. half [Joh92b]. Hall [Bes95b, Col97, Dod87b, Gar94, Hay93d, Hay93a, Hay01b, Her92b, Hob97, Min94, Per96, Per97b]. Halm [Hay90b]. Hampshire [Coo07, O’H99, Tat04]. handbook [Col97, Hay95, Hay97c]. Hanel [Cre04]. Hans [Ano93b, Ano94-33]. Hanseth [Gra01]. Hanusch [Ano94-38]. happen [Ric85]. Harald [Bes95b]. hard [KM03, Lae03, Pfi99, Gei97]. Hardback [Ano91d, Ano91c, Col08, Dut07, Hay84, Hay85a, Hay85b, Hay85c, Hay86d, Hay86e, Hay89a, Hay89d, Hay90a, Hay90b, Hay90c, Hay90d, Hay91a, Hay01b, Kap03, Kuc07, Pra10, Ram86, Ros90, Sha91b, Ver90a, Ver90b, Wil90, Arr04, Bes93, Bes96, Bes97, De 04, Del97, Dod88, Gil96, Giu09, Har93b, Hay83, Hay85d, Hay86a, Hay88a, Hay88b, Hay88c, Hay88d, Hay88g, Hay88f, Hay88e, Hay92c, Hay92d, Hay92e, Hay95, Hay97b, Hay97a, Hay01a, Hay01c, Hol93, O’H99, Oak97, Per95, Per97a, Rad97, Ros89b, Sap10, Sha95, Tra00, Wya07, Gri98, Mer02]. Hardbound [Ano97b, Hay97d, Her04, Bas04]. Hardcover [Kra15, BH03, Giu09, Hop04, Lae04, Rot88]. hardware [RH21]. Harlow [Why02]. Harmonizing [IH07]. harmony [ZGSG23]. harnessing [GSP^{+}22]. Harper [Hay87d]. Harrison [Bes02c]. Harry [Bec89]. Harryson [Fra01b]. Hart [Bes94c]. Harvard [Bes95a, Bes96, Bes01a, Bes01b, Dav92, SW04, Ver90b]. Harvey [Ano93o]. Harwell [Guy97]. haute [Leo21]. HCWS [xCZZ^{+}22]. head [JWY23]. Headhunting [Mac86a]. health [BSMC23, CCR^{+}23, CJH^{+}07, EFC23, GKK23, SSGN23, WH13]. health-tech [SSGN23]. Healthcare [CP23, CNN23, DDFS23, ABSZ23, BTP^{+}23, BCPP23, BCPP24, BKA^{+}23, BGF23, CCP23, CCR^{+}23, DMRS23, DP23, GKK23, GPCNS^{+}23, GMB^{+}23, II21, KS21, LDT^{+}23, Mas23, PGB^{+}23, RDAV23, RNJ^{+}23, STA24, SSCH23, SD23, SMI23, TFS^{+}22, YH09, ZSX24, CWY14, TFS^{+}22, VS06]. heart [Fel22, Haa14]. heart-rate [Haa14]. Heather [Ano93j]. heavy [SM12b]. Hébert [Ros89c]. hedonic [TLL07]. Heidelberg [Lim06, Pap02b]. Heinemann [Bes02c, Hay86e, Sap01]. Heinrich [Lim06]. Heinz [Hay96a, Rad97]. Helen [Tan08]. Helena [Que06]. Helix [BCM^{+}22, JLWS23, Joh08, RGH19, BBK23, CE19, LAYS19, Lin19a, LH19, ZCF19]. hell [BSSF24]. Helm [Hay85a, Hay86a, Hay88e]. help [BSG^{+}24, BF99, SMW21, YZY10]. Hemel [Bes95b, Her92b, Per96, Per97b].

Hempstead [Bes95b, Her92b, Per96, Per97b]. Hendry [Hay93d]. Henry [For90, For91, Gar94, Per97b]. Hepsø [Gra01]. her [Hay87a]. herding [ABvZ23]. here [AGFE22, dABKS14, IBPGML23]. heroes [Lin09d]. Herts [Rus94]. heterogeneity [CDMV24, HBM14]. heterogeneous [SBP21, WSH23, ZHZL21]. heuristics [FP08]. hi [Par05]. hi-tech [Par05]. Hicks [Joh92e]. hidden [JLY23]. hiding [HEC22]. hierarchical [IMBH98, LFU23, Par00, SS21]. High [CRKA98, Car01, JR91, LS05, MM08, ALvKL06, Ano93s, Ano94y, Ano94-28, Ano94-29, Ara08, BMV02, BDB21, Ber96, BG19, BM14, BRP94, Car98b, CLZ18, CLZ21, xCZZ^{+}22, CA06, CJL22, Cou90, Cou91, Dod87a, EH02, Gra03, Gra04, GC10b, HAA09, HdWN21, Hay88e, Hay93b, HFWW16, Huc85, Jar86, JMT03, JWY22, Kak03, Kod07a, KLH05, LWC09, LD05, LY00, Lin86, Liu93, Mac86a, MZG17, MP09a, OG05, OGG06, OS08, OC91, Oak03, OB20, QvdR21, RIMP22, Rog86, Rya96, SG93, SP98a, SP98b, Sen85, SFC14, SPLB19, SCC17, Ste14b, SMW21, SL14b, Töd94b, Töd94a, VTRM12, VK03, Wan89, WTJ23, WCT10, WTN^{+}02, WAN03, WKN04b, Won16, Yas05, Coo16, HdWN21, CF07, Gar85, Hay91a, Ver90a, Ver91a, Won08]. High- [Car01]. high-barrier [BG19]. high-level [WTN^{+}02]. high-potential [QvdR21]. High-Tech [LS05, ALvKL06, Ano93s, Ano94y, Ano94-28, Ano94-29, BDB21, Ber96, BM14, BRP94, CLZ18, CLZ21, CA06, CJL22, GC10b, HAA09, HFWW16, Kak03, Kod07a, KLH05, LWC09, LD05, Lin86, Liu93, RIMP22, SG93, SCC17, Ste14b, VTRM12, WTJ23, WCT10, Won16, CF07, Won08]. High-technology [CRKA98, Dod87a, Jar86, Rog86, SFC14, Töd94a, WAN03, WKN04b, Yas05]. high-value-added [Rya96]. high-velocity [VK03]. Higher [Car98a, Fel23, LJ01]. highest [Ver01]. Highway [BSSF24]. Hilary [Hay87a, Hay87b, Hay87c]. Hill [Mer02]. Hills [Rus94, Szw07]. Hilpert [Pap06]. hinder [ZL24]. Hiroko [Hob87]. Historical [Hit02, Hit03, Hit05, BCV21, CYJ07, Hit93, Ho09, SL98b]. historically [Ver01]. history [Ano91d, Ano91c, Ano94c, Hay97b, Hit94]. Hitomi [Ano94c]. hitting [CKF^{+}24a]. Hobday [Edw04, Hay97a, Sod08]. Hochstrasser [Hoc89]. Hodge [Ano91d]. holding [SCY22]. holes [HHV05, WQZ21]. Holistic [AKKS12, DG14, Li21, MNW14]. Holography [Bro89]. holonic [SV01]. Holt [Bes88, Hay85b, Hay85c, Hay88d]. home [CKF^{+}24a, FB93, GC99, Sha95]. Honda [Joh92f]. Hong [Kim87, Che94, CC14, CTC02, PCWG04, SW05]. Honoring [New13]. horizon [Hui17, MCMA09]. Horizons [Que06, MDT^{+}23]. Horizontal [CE89, LRH22]. Horst [Ano94-38, Hay85b]. Hörte [Dre06]. Horváth [BB95]. Hosni [Leh07, Lop06]. hospital [CCR^{+}23, ZWZ^{+}23]. hospitality [Lin09b, ÖBA09, SSM09]. hospitals [RDAV23]. host [MPWG22]. hostility [LL05a]. hot [Lin12d]. hotel [MROS09]. hotels [DB09]. Hotspots [HR98a, HR98b]. Houndmills [Coo07]. House [Rus94, Tat04, Ano81b, Rie89]. household [DAOO02, OTA01, TOA01]. houses [dVV16]. hovercraft [RG85]. Howard [Hay97a, Mer02, Ano93i]. Howells [Hay01a]. Høyang [Hol88a]. HP2 [Per97b]. HRM [Jon96]. Hsinchu [LS05, LY00]. HSIP [HLC05]. HSP [LY00]. Hsu [Ano94l]. Hsun [CF07]. Huawei [LfSS19]. Huber [Her04]. Hübner [Hay96a]. hubs [AGFE22, Bat12]. Huff [Fra03b]. Hugo [Ano97b]. Hui [CF07]. Hui-Tzu [CF07]. Huisman [Par03b]. Hull [Dre04]. humain [Ano94u]. Human [AAAF22, Kra22a, LLB96, PS00b, Ano94u, Ano94g, AN16, CE89, DBR24, Hon21, Jus24, LG23, MZG17, PSdLC02, PS97a, PS97b, SCC17, SV01, Tai98]. human-centered [DBR24]. humane [DR24]. humanitarian [De 23]. humankind [Ser13]. Hungarian [CN86]. Hungary [DNS^{+}03, Pál04, Szá94a, Szá94b]. Hurt [PF98]. Hurt-Joseph-Cook [PF98]. Hussain [Kaz03]. hybrid [ANI^{+}22, CE19, GFvT22, SM12b]. hybridity [SPÁG22]. hybridity-based [SPÁG22]. Hydrogen [JMK05, MW09]. hype [SH23]. hyperlearning [Car99b]. Hyperlinked [Hu13]. hyperlinks [KKL19]. hypertensive [TW02]. hypotheses [IKL01]. hypothesis [NK22].

I-Space [BC99]. Iammarino [Arc03]. IAMOT [Ano07k, Ano09l, YKW10]. Ian [Ano91c]. Ibadan [SI98]. IBM [dHOvdKS16]. IC [Ano94l, CSH94, CT02, KGK07, KKG07]. ICAM [Tan87]. ice [Ver05]. ICT [BMLL07, CS02, CZM23, Fal05, FAH08, HAA09, LD05, MCT06, SPKK08, VM11]. ICT-linked [Fal05]. ICTs [OOA04]. Idea [Sco89, BDW06, CZL^{+}24, MNW14, McA04, Mor91, SY14, WF10]. ideas [CMB23, Gil97, HKW^{+}22, MM02, NPG14, OE87, PKH15, UB82, PKH15]. ideation [OSSC21]. Identification [OSSC21, BK89, MPCRB^{+}03, ÖBA09, SC17, TB24, Ven96, WGW22, ZGSG23, Cav16]. identify [QvdR21, SS22, WTN98b, WTN98a, WVG^{+}98, WGV^{+}98]. Identifying [BH16, GLYP13, LW11, Lin13b, MWJ23, OS08, SS21, HS14, Lic05, YLH22, KR12, RS14, Wal12]. identity [TKYM24]. idle [OIOA01]. IEE [Rus94]. IFS [BB95, Bes95c, Mer94a]. Igniting [Bes01b]. ihre [Ano94j]. II [AHAAEA03, Box01, FFJ92, Ish89b, Sna90b, Tho88c, WPF06b, Whi90a]. III [Fra01c, Sna90c, WPF06c, Whi90b]. illustration [ET98, MZL15]. Ilori [Ano94-31]. image [TB24]. imbalance [HF18]. imitation [AAAA23, RL06]. Immediacy [WHB21]. immediate [KM11]. imogolite [Coo90]. Impact [Bam94, HMWT23, LFJA23, Lin11f, SAY09, SY14, VLIC12, dVV16, ASI99, AKBA07, AC08, AKKS12, AHP16, ALM22, AB23, BGA09, BC10, BS05, Boa08, BC92, dABKS14, CCB20, CMA07, CCF23, CLW10, DCD22, Der14, Efr14, EEY22, FCM21, FOW97, GZXX23, HPMI^{+}23, HBM14, HS14, dCHDAALMM07, Hon21, HBH21, HC10, HC13, ILR^{+}12, JWL24, JWY22, Kaf06, Kas94a, Kas94b, KL24, KZ03, KRM^{+}24, Lay93, LbKL91, LQ09, Lin09e, Lin13a, Lin16, LDMY22, LC22b, MFH16, MCA^{+}24, MG08, NHS24, NT16, OSG04, OTA01, PCG22, PCQ24, PRA22, Pet11, PG11, SGB12, Sar22, SP22a, SH10a, SMGPP24, SJ16, SMD23, SMW21, SL14b, SD04, UOIA00, UF22, WZGBA01, XW21, YWC09, ZWDZ21, dVdWNF^{+}10, dlHBR22, vdWFvdB23, KLL24, Lin12a, Rot87a]. impacting [KC07]. Impacts [Rue86, WA04b, WWYZ23, GC10a, GHU21, JSR19, KKP17, SVV15, TPH15, WFWC21, WHB21, YVG24]. impede [CH01]. impediments [DSS05, ZS04a]. imperative [Ree99]. Imperatives [Dre04]. imperfect [AAGL23]. Imperial [Dre04, Lae04]. implantable [JFCFMC^{+}12]. Implanting [OE87]. implement [CLZ23, CCF11, Kap11, MM11]. Implementation [GC98a, GC98b, Lin02, YGW06, YGW08, BAWK12, BM07, CCP23, CHT99, CS12, CPXC02, DFFM22, ETOA00, ETA02, EB23, Eng06, Gup96, HB89, KS21, KMK03, KKRM05, LPS03, Mol97b, Mol97a, PC89, Soh97a, SM01, TCT05, VM92, Per96]. implementations [Ano90o, DB09, Moh95, Soh96]. Implementing [BBC06, BB89, BF97, CWBG03, Har96, Pra99, BJB^{+}24, FRW06, GH91, Hay88c, Hos03, HK86, LCLC03, Suc87]. Implications [BO86b, DLL06, GH91, Gup00, AOO02, AM92a, AM92b, BE09, Bor01, CK22, CS84, DHT22, FKT95, Fox86b, FB93, FB05, GC96, GG03, GK06, Hyv06, II98, Jon96, LL04b, LHY04, LJ01, Mik01, Mil88, Mud95, Oak03, Oye01, RK04, Seg86, Sen85, SBMP13, SH99, SSP04, WWY15, ZLC18, dBJ03, Zeg87, Cre04]. Import [MKK04]. Import-led [MKK04]. importance [CR19, D’I14, Hus10, LL05a, Lun14, MCMA09, NGK08, NS07, SK23, SM95, Ste89, VLIC12, ZZXW23, Lin13d]. important [Ano94n, Cai94, Tre98]. imports [Win87]. Impressively [Lin10c]. imprint [Hay93a]. improve [BF99, CCC^{+}12, GBV23, Lee23, NSYL18, SRT18, WHP85]. improved [MAMD22, Ott03b, Sæt24]. improvement [Ano94-41, BBHW93, BCG^{+}94, BCG01, GC98a, GC98b, IS97, JW06, Kli06, LB03, MC00, Ram23, TS00, WCZZ14, WC04, WC06b, RMGBA08, Bes02b, Pov97]. Improving [Gol90, Lin16, MGGOS21, OC85, SO02, Ano94-36, BCPP24, KWE05, KPK94a, KPK94b, MKGL05, WA04a, Ano94r]. Improvisation [LLY^{+}18, AAGL23]. in-depth [TRR^{+}24, vBDK21, GKB12]. in-house [Rie89]. In-licensing [Sla23b]. in-vivo [JFCFMC^{+}12]. inbound [EB23, SP22b, SCK10, SCK11]. Inc. [Fra01c]. incentive [MAB^{+}23]. Incentives [Ano03q, DC22, Kog03, Sto21]. Incentivizing [SZSR23]. incl [Arr04, BH03, De 04, Lae03, Pre01, Wya00, Wya04]. include [HJM08]. including [Ada07, Ano91d, Ano91c, Bas04, Cre04, Kap03, Lae04, Mor04, Pra10, Wya07]. inclusion [SO20, WHB21]. Inclusion-Immediacy [WHB21]. inclusive [AN16, WAdN19]. incongruences [RNFG24]. incorporated [SG24]. incorporating [TKYM24]. increase [Wit85]. increased [LS04, SK92, SYK94]. Increasing [SFdO^{+}23, Cam93, OGOF22, dMvAdMSPB93]. incremental [OKP13, Paf04, Udd06]. incubated [Lun14]. incubating [GG05]. Incubation [LH23, MLF16, AMA21, BH16, BRCG12, CvZ05, CKS22, HSLL22, LC09, MMM16, Mrk17, PCWV16, SJ16, Swi92a, Swi92b, vRE21]. Incubator [BN08, BMSJ24, Bøl12, BRCG12, CvZ05, CL05a, CK12b, FJ12, MM08, NHS24, SH10a, SC10, SSC23, vWvRN17]. incubators [Aab09, AMV07, Ano94-43, BRCG12, GG05, HSLL22, LLZ19, MM08, Mia94, RH10, RD05, SH08, SMW21, VM12, VvWM20, GM23, RAS15]. incumbent [DDN24, SCM11]. incumbent-startup [DDN24]. incumbents [EKK^{+}21]. independent [Hit06, LDB13, Vos84]. Index [Ano89s, Ano90m, Ano91c, Ano92e, Ano92f, Ano93g, Ano93h, Ano94o, Ano95n, Ano96n, Ano97p, Ano98n, Ano99e, Ano00b, Ano02g, Arr04, Cre04, De 04, GLYP13, Lae03, Mor04, Pre01, Ada07, Ano86j, Ano88l, Ano89a, Ano89t, Ano89u, Ano89v, Ano89w, Ano89x, Ano90a, Ano91b, Ano91g, Ano01a, Ano03a, Ano03b, Ano04a, Ano04b, Bar89a, Bas04, BH03, Kap03, KK04, Lae04, Pra10, RMGBA08, Wya00, Wya04, Wya07, Ano02a, Ano06a]. Index-based [GLYP13]. Indexes [Ano91d]. India [Ada07, Sub05, Bas01, BAR16, GMWK23, Kha98b, Kha98a, KJ03, MD03, NB09, NB11, Pan98, Paw94, RIMP22, Sik91, Sik94, Sik97a, Sik97b, Sik99, Sou83b, ZQW^{+}19]. Indian [Ban03a, Ban03b, Bou15, Day02, For99, Kha99, Kha97, Kha98d, Kha98c, KS10b, MKK04, Meh01, RM14, Sik96, ZYSL21]. indicator [Pil04, Sir91]. indicators [Ban03b, Bar96, BJ05, DB19, FNJV14, HG92, Hob87, MCMA09, Rag93a, Rag93b]. indices [Bar89b]. Indigenisation [KJ97]. Indigenous [II98, Sik96, Tsa05b, WLO09]. indirect [DLC21, Efr14, GAS96, SHT23]. Individual [YMK^{+}21, YMMYK24, EHHW17, NVMD22, OWC03, wtcoLSDD91, ST21, VJ12, Wür17, KWvdS21, WWH15]. Individual-level [YMMYK24]. individuals [MCA^{+}24, SFVI21]. Indo [Ano94-37, PG94]. Indo-US [Ano94-37, PG94]. Indonesia [IDMD04]. Indonesian [MKK04]. induced [BS24]. inducement [OW00b]. inductive [Leo21, SC20]. industria [Ano94q]. Industrial [Ano89t, Ano89u, Ano89v, Ano89w, Ano89x, GP04, LS05, Pap06, Rot92a, Rot92b, Sap05, Ski84, SD95, WWM00, AH84, AH85, AA05, AA92b, AA92a, AGS98, Ano94-31, Arc03, BS01, Bes15, Boa08, Bro95, CHT99, CF01b, Chi91, CC06, Dur86, DNS^{+}03, Fah05, FH87, GC10a, GHHH22, Gol90, GG11, Hay95, Hea83, HG92, Hit06, HC01, Hsu05, Hu13, Hun00, II21, Ilo94, JSP22, KBS01, LB02, LPMV22, LTH06, LvLB06, MJ92, OE87, Pap95, PV01, Pyk00, RUNS23, Rep08, Sch85, SCH93a, SCH93b, Sen96, dSSD09, Sko87, SD10, TKSM08, VLVH23, VM92, WD06, YLH22, YW08, Zeg87, ZZL22, dOK12, dBJ02, KAV17, VBA06, Bai84, Hay83, Hay86a]. industrialised [Els95, Sou01]. Industrialization [Sch85]. industrialized [Chi90, CU13, Tre98]. industrialized-industrializing [Tre98]. industrializing [AT93, Ano93l, PBC93, Tre98]. industrially [Tit90a]. industrie [Ano94s]. Industries [CF07, AS21, ALCM14, Bha05, Car99b, CYK19, CHL21, CF01b, CZZM21, CPXC02, CTC02, Chi21, Cur89, FCS21, Fre86, GBFvdS11, HCK01, HFWW16, HC06, KSHP21, KL04, Klo87, LB21, LDF^{+}21, LD05, LC07, Li21, LW97, LCT04, LCT05, MJ94, MZG17, Nar89, NGK08, OB20, RGFLRP21, San95, SCK10, SCK11, TCT05, Tsu01, WJK15, YHL22, Yas05, ZCF19, ZLJ^{+}21, vDZ21, CS21, SBP21, Hay86d]. Industry [Bje10, LP07, LCWC09, OIOA01, OIN96, WZNG17, Won08, ADDS09, AS08, Acs88, Ahn95, AW11, AF08, ALRSD04, ABGS13, Ano93i, Ano94q, Ano94l, Ano94-35, Ano94-41, Ano94-45, Arc88, AK00, Ayr88, AC07, Bad08, tBL86, BS08, BF03, BF05, BBDM01, BWF18, BCPP23, BCH98, BCC^{+}11, BNV05a, BF21, BA10, Bou15, BV09, BEH99, BO81, BLQS00a, BJ05, BA89, Cai23, Cam93, CJ98, CR00, CAI00, CS01a, CA01, CS02, CA02, CLGFVCM23, CCF23, CSH94, CHT99, CTYH02, CT02, CYK19, Che94, CYHC19, CZM23, CM95, CY00, CJS00, Col81, Cre04, CN90, CN92, DJ13, Dav93, Day02, DNT98, DM11, DQ06, DR07, DM08, ETAO98, ETOA00, EH14, Eri00, Ern95, Fan06, FS84, FC06, For99, FPP24, Fuj98, FL15, GMB^{+}23, HBS02, HBS07, Ham96, HMWT23]. industry [HO14, HAA09, HK02, Hay91a, HBO14, Hit02, Hit03, Hit04, Hit05, Hob90, Hob92, Hoe85, HP09, HST07, Jai23, JWY22, JvZ08, JH18, KM06a, Kan94, Kap01, Kas94a, Kas94b, Kee14, KLRR22, Kha98b, Kha98a, Kha97, Kha98d, Kha98c, KLL87, KJE96, KBS01, Kum05, KB06, KS10b, LVVH23, LPS03, LRH22, LL94a, LL94b, LY00, LW04, LKPP07, LQ09, LD95, LV09, LWL^{+}22, LH07, Liu93, LJ01, LM94a, LM94b, LG22, Lov85, M’C05, Mac86b, MCD06, MROS09, Mas86, MM02, MW06, MI07, MR05, Mow90, NWGB09, NB09, NB11, NDP07, NG97, NLLC05, NP10, NB92, NPV05, Obe88, OVD21, OM89, OB21, ØD22, PPM24, PPB89, PS14, PS00a, PV00, Pet11, PDHW23, Poz86, Pra99, PSS00, PUFC23, QWM92, RVW17, Rad01, RDAV23, RV17, RG06, Rie89, RFMC21]. industry [RHW14, Rue88, RI04, Ruo05, SBF15, SH06, ST95, Sar22, SSM09, Sch92, Sci83, Sha03, Sha86, Sha93a, Sha93b, SY05, SPLB19, SGWL16, Sta98, SCM11, SW05, SH99, Sur92, SD89a, TWGB02b, TWGB02a, TM06, TPB12, TCC^{+}22, Uzu01, VJ23a, VJ23b, Ved81, Ved97, VOTH05, VNP^{+}22, VH06, WTAW97, WTN98b, WTN98a, Wal07, WTJ23, WC06a, WL21, WZGBA01, WTN^{+}02, WA03, WN03, WAN03, WH04c, WM88, WHP85, WJK15, Xie04, YI05, ZI01, ZWDZ21, ZCR22, dMvAdMSPB93, dHOvdKS16, lS06, BK24, CFP^{+}23, CKF24b, CCBP22, KRM^{+}24, Hop04, Har93b, Hay85c, Hay88d, Hay90d, Wan04]. industry-academia [BJ05]. Industry-academic [OIOA01, OIN96]. industry-Does [NP10]. industry-related [SD89a]. inertia [Coc12, hLFL08, TWGB02a, ZSX24]. inferences [RAC84]. Infertility [YLB06]. inflation [Bal85]. Influence [GMLMVJ07, KP12, KS10b, MRGMBR13, MMM05, RIMP22, Wya07, vRGJ^{+}12, AHSS19, And21, Ano16h, ABUS05, BDB21, BBS22, BL97, BNS16, CCV11, CKM13, Ene22, FORA17, FOW97, HLL12, JMPR04, KAV17, LS24, LWP14, Lin09e, MKL19, Mat22, Men11, MAB^{+}23, MP09b, MR19, NG09, SCS22, SvG15, SH02, THC^{+}23, WSN^{+}18, WQZ21, ZMPRP14, vdEJG08]. influences [BRA^{+}23, CN90, CN92, HT12, LW11, LAYS19, Puj06, TPBM23, WC07]. influencing [Bar90b, DAOO02, HB03, OTA01, OM00, RFMC21, IPP14]. Influential [CY00, Lin11f, PS21]. inform [MMCD23]. Informal [Lan98b, Lan98a, Pyk97, Pyk00, Tre98, ES05, MAKKM22, Neu12, SSF12, Sha93a, Sha93b, TLL23, Wet81, XLC23, ZLH17, wK97, wK98]. informality [ALM22]. informant [ET98]. informatics [GAS96]. Information [AK06, AC09, BS99, Car05, Des96, EM02, Hay93b, hLwChLbL04, OOJ^{+}00, Poy97, WL91, ZL01, AAGL23, ASI99, AL05, Ano93i, Ano93k, Ano94-33, Arn87, AZ90, BGA09, Bor01, BM07, Car99a, CE89, CW94, DB09, Dav93, DFD^{+}91, FB06, GS05, Gra01, Gup89, GG93, Hay90b, Her92b, HBvdB09, Hob88, Kum06, LC07, LL08b, MWJ23, MII90, OW93a, OW93b, Per96, PMN^{+}24, RMGDLC08, SS21, Sér96, Sna90c, UOIA00, Wan04, WO08, XW21, YWC09, BS08, VS06]. Informations [De 04]. informed [RH23]. InfoTech [Szá05]. Infrared [Hay82]. Infrastructure [Hos00, BN06, CDD^{+}11, CDD^{+}14, CLZ^{+}22, DO03, HN92a, HN92b, IEG97a, IEG97b, KL08a, Lar09, MW09, PGLS21, Tan95, VS06]. infrastructures [AM97, CE89, WPR^{+}22, Gra01]. infusion [TKYM24]. Initial [ABUS05, MMR24, ALM22, Bar97, Hon21, JH18, NPV05, RBBF23]. initiated [FJ87b]. initiation [GF21]. initiative [LD95, Tit90a]. Initiatives [RMG06, Per04, WLTH23]. initiators [Fox95]. Inner [Lin17c]. innerspring [Sch09]. INNOTECH [Pál04]. Innova8ion [CHP13]. innovación [Ano94r]. Innovate [Eek84, ACBO23, GBY22, SW04, Bro90, Hay89a]. Innovating [KRK^{+}22, NH99, CAP23, Lüt04, Pap02b, RWP19, WR95, Hay88a]. Innovation [AM01, AA88, Acs88, ARvT04, Ano87h, Ano09l, Ano11a, Ano16h, Bai18, BD23, Bes01a, BMRG00, BNS16, CGM07, CT06, Car04, Cha88, CCF11, CKF24b, CCR^{+}24, CL88, Da 01, DNT98, DP23, Dre04, DB19, FJ91, FFJ92, Fra01b, Gal06a, GH21, Hal14, HMM14, Hay98, Hob95, HT06, HMFS24, Hop04, How24, JV97, Jel07, JBZ23, KS00, Koc04, Kod07a, Kod07b, Kuc06, LB21, Leh07, Let03, LTT23, Lin11f, LBG14, LTHS12, ML08b, MH04, MROS09, MT08, MVS^{+}22, MPV13, Moh95, MG08, NPG17, NW23, NKL02, OM18, OPPL16, Pap06, PKH15, Par07, PPB89, Pea90, Que06, RGFLRP21, RAATM05, SdO^{+}15, San95, Sap10, SS23, Sco89, SdVGdFB24, SHT23, Sod08, Ste89, Szá94a, Szá94b, TDC11, Thu18, VMWV16, VSE24, Wer03, WZNG17, Won08, YKW10]. innovation [AM23, AH84, AGFdMCM24, AS08, AHSS19, AB99, AK06, AABY^{+}23, AA05, AMS^{+}24, AHS^{+}23, AC08, ABSZ23, AS22, AGFE22, AL05, ALBO08, Ame95, And21, AMP22, AAM24, Ano82c, Ano87g, Ano94-36, Ano94r, Ano94x, Ano94-42, Ano94-41, Ano12a, ALCLdF^{+}15, AHP16, ACA24, AES95, AP96, ABKL08, ASOB23, ARR23, ALM22, AB23, ABG23, Aug06b, Aut94, AFL^{+}24, Bad11, BS01, BBTW00, Ban04, BZIZ21, BBLW24, Bar89a, BSSF24, Bar90b, BDB21, BT22, Bar15, BWF18, BM16a, BLA06, BRV23, BJB^{+}24, BS24, BCL22, BS06, BLNP05, BG19, BAR16, BBK23, BRA^{+}23, BF21, BVS24, Bou15, BR16, Bre23, BV09, BD05, Bro98, BM05, BF06, BHPB06, BWS11, dABKS14, BOBT22, BVHJ24, CCB20, CPT23, CA11, Cai23, CLZ23, CHP13, Cal24, CPSM12, CvdEG13, Can16, CB19, CP11]. innovation [CCP23, CFP^{+}23, Can06, CLP23, CR00, CAI00, CLGFVCM23, CBGvdV22, CS84, CS87, Cha22, CL12, CS04, CS05, CCC^{+}12, CY12, CH04, CHD18, CCMW23, CKF^{+}24a, CCGK23, CNN23, CHL21, CL05b, CG11, CLZ18, CLZ21, xCZZ^{+}22, CLZ^{+}22, CJ23, CZM23, CMB23, CS08b, CLW10, CK12a, CS12, CW22, Chi05, Chi21, CP07, Chu01, Chu02, CJL22, CHT22, CKM13, Col81, CR19, CJL23, CKV^{+}23, CGBV21, CO04, CTSJ24, Cou90, Cou91, CMS23, CZL^{+}24, DDBM23, DLC21, DC22, DZX22, DSPS23, DVdCAS16, DVSL19, DQ06, DG14, Dib01, DVMS16, Dod87b, DS13, DM08, DS23, DLTQ23, DO03, DLR07, DLKM24, DVHL22, Dur86, EDM05, Efr14, Egb15, EB23, Ell81, EWJ09, EWE22, EKK^{+}21, EH14, EHHW17, EEY22, EF11, EL19, EIMS01, FGTB23, Fan06, Far23, FSA23, FHBDO16, Fel22]. innovation [Fel94, FCCE21, FB06, FCM21, FMOPSVA21, FLYTB24, FCS21, Fis89, FF00, Fox86b, FJ87a, FJ87b, FF92, FB05, FDN^{+}23, Fre99, Fre05, FdJ09, FdlCU22, FM24, FD21, Gal06b, GSP^{+}22, GR23, GMLMVJ07, GST08, GLÖS23, Ger04, GSFS21, GCH96, GF21, GK91, dVGCOF^{+}23, GWZ^{+}23, GÁNA07, Gra82, GARM24, Gra24, GGGR22, GKB12, GL10, GM00, GZXX23, GC10b, Gui05, GMB^{+}23, GMWK23, Haa14, HKRR20, Had99, HBS02, HBS07, HBM14, HMWT23, HR98a, HR98b, HB84, Hay81, Hay83, Hay85a, Hay85c, Hay86c, Hay86d, Hay86a, Hay87c, Hay88c, Hay88d, Hay96a, HN08, HR07, HW18, HOGGP11, HEC22, HBG12, HPBP98, Hol85, Hol87, Hol88a, Hol88b, Hol90, HFWW16, Hsu05, Hu13, HMM15, HBH21, HL22, HST07, HC10, HLL12, Hüb96, HKY22, HCT11, Hui11, Hui17, HAH84, HHA84, HC13, HLO14]. innovation [HLNP21, HDE21, HKL24, Hyv06, II21, IS97, IBPGML23, JLWS23, JLL10, JL24, JWL24, JSR19, Joh99, JC01, JJS^{+}18, Jus24, KBSW08, KSHP21, KP12, Kap99, Kap11, KBO22, KZ18, KC08, KT02, KS21, Kih03, KL08a, KA13, KET22, KOX21, KHL15, KMF09, KS14, KH06, KB09, KLL24, KBS01, Kos99, KMID11, Kra22a, KLH05, KPK94a, KPK94b, KS10b, LVR23, LVVH23, LS23, LS05, LD02, LDF^{+}21, Lar09, LOH91, LNB21, LRH22, LO94, dLP06, Lee23, LSB07, LYST22, LB82, LYY^{+}22, LWL^{+}22, hLFL08, Lic10, Lic11, LW21, LHI13, LWC^{+}12, LZLL21, Lin10a, Lin06b, Lin09c, Lin09e, Lin10c, Lin12d, Lin17a, Lin18, LW97, LLY^{+}18, LDMS23, LLZ23, LLB96, Liy95, LL05b, LC82, LMMPVO09, LR09, LC22b, Mac97b, MJ94, MTB01, MDT^{+}23, MAS13, MLG12, MdVB16, MMMD23, MCD06, Mar03, MRGT11]. innovation [MAMD22, MMC^{+}23, Mas86, MKGL05, McN92, Meh01, MHKZ24, Men11, MRDA24, MK87, Mik01, MAKKM22, MRC14, MI07, MTU23, MAB^{+}23, Mon06, MMM05, MGAP23, Mor91, MM11, Mud95, MP09b, NG09, NF24, NHT23, NS07, NSB23, NHV05, NT16, NBA22, OGS06, OC91, OVD21, ON88, OKP13, OGRA23, OBL14, OBvE97, OWC03, OKvB23, OB21, OBMOM24, OW00a, OW00b, ØD22, Ott04a, Pal93a, Pal93b, Pál04, PGLS21, PCG22, PCQ24, PWA11, Pap02a, Par05, PMP24, PBOR23, PC18, Pat24, PG09, Pen08, POGBG06, Per04, PGKF19, Phi99, PPMC24, Poy97, Pra99, PS01, PS04, Puj06, PUFC23, PEER23, PE22, QWM92, RVW17, RM09, RV17, Ram23, Ran24, RGB23, RR11a, RSP10, Rea21, RS14, RMGBA08, Ric85, RHL09, ROMH15, RA17, RRB^{+}22, RPGP23, RDS17, RH23, RBBF23]. innovation [RD05, RG85, Rot87a, Rot91, Rot92a, Rot92b, Rot93, RFB24, RkhR97, RHW14, Rya90, RGH19, RK04, STA24, SRLR23, Sæt24, SAWT06, SRHL12, SVLYT24, SSF12, SK23, SGB12, SNM12, SSCH23, SBP21, SVÁG07, SPÁG22, Sar22, SM12a, SBOB23, SPMRS23, SMvFM24, Sch92, SPFG12, SMGPP24, SCFD18, Sen96, SCY17, SP22b, Sgo95, SAM^{+}23, SR22, SXS24, SSW22, Sha86, SF05, SS22, SH23, SHS^{+}23, dSSD09, Se14, SFF21, SCM24, Ski84, SJ21, Sla23b, SS16, Smi07, SY17, SvG15, SJ16, SC05, SL16, SSKB24, SZSR23, Sou83a, Sou01, SVV15, SCK10, SCK11, SK11, SA22, Sta98, SW22, SGF02, Sto21, SDK24, SMD23, SYCH23, SFVI21, SD10, SJCT24, Sun96, SDN23, SD95, Swi92a, Swi92b, SF98, TKBC22, TW02, TKYM24, Tan87, Tan98b, Tan98a, TB24, TS00, TSH02]. innovation [TM06, Tet00, TPV16, TM03, Tho12, TPH15, Tik23, TLK09, TEA00, Tõn15, THC^{+}23, TS05, TKS18, URG^{+}24, UEGE^{+}14, Uzu01, VLVH23, VGAL24, Vav06, Vek99, Vel15, VTRM12, VSC16, VOTH05, VX97, Vos84, VRS^{+}22, VH06, WvdWG24, WOL05, WWY15, WTJ23, Wan04, WC06a, hWyLbC08, WCT10, WZSZ16, WFWC21, WH21, WL21, WGW22, WLTH23, Wan23a, Wan23b, WTGB01, WH04a, WHWS22, WSH23, WQZ21, WWYZ23, WHP85, WML^{+}23, Why02, Wil83, WvT23, WBM22, Wit85, WKF24, Won04, Won05a, WJK15, WAdN19, WLG05, WW14, XW21, XHM^{+}16, XWZ23, XLC23, XL23, XLD23, XX21, YNS18, Yam21, YK08, YWC09, YHcHW17, YHcHW21, YCL22, YMK^{+}21, YWL22, YW08, YT09, YH11, YZY10, ZIAO^{+}21, Zaf81, ZKW^{+}23, ZIB23, ZXT10, ZWD09, ZQW^{+}19, ZWDZ21, ZWZ^{+}23, ZZXW23, ZSX24, ZZL22, ZZL^{+}23, ZMH22, ZLC18, ZLJ^{+}21, ZL24]. innovation [ZXDG19, ZHZL21, ZMJM23, dGSJ^{+}21, dMD05, dVdWNF^{+}10, dVV16, dBJ03, dB10, lS06, vWGdG21, vdVdJVdR09, vdVdM11, vH10, vK11, Ano90o, Ano03q, AM97, BE09, BCC^{+}11, CAP23, Gof07, LTVA22, VNP^{+}22, Wal12, ZCHC23, Dib00, Her04, Pre01, Why05, Ano81b, Bes96, Bes02a, De 04, Hay85d, Hay85b, Hay95, Joh92g, Lae03, Ram86, Bes88, Cre04, Giu09, Hay86e, Hay90b, Hay01a, Lae04, Wil90]. innovation-enabling [CFP^{+}23]. innovation-engaged [FGTB23]. innovation-exploring [SRHL12]. innovation-focused [TS05]. innovation-investment [ZIB23]. innovation-oriented [KLH05, WH04a]. innovation-related [RHL09]. innovation/integration [Lin18]. Innovational [BLQS00a]. Innovations [Ada07, Hja09, Ver05, AH85, AO10, BTP^{+}23, BBC06, Bie91, BAWK12, CWY14, Coz06, DJ13, DAR07, DGS23, EM85, FGTB23, FGS07, GC10b, HPMI^{+}23, HV96, HESPSL11, HW97, HM11, HKJ07, KET22, LV09, LG23, De 23, MM24, MGGOS21, OCA09, OBMOM24, Pal06, PRA22, Pyk07, RM14, Sap92, SLSCS24, SO20, Sik99, Smi07, Sub05, THHS22, Udd06, VNP^{+}24, WKN03, WBM22, Won06, ZC19, Hay84]. Innovationsprozesses [Ano94-41]. Innovative [BDW06, Die00, Fra00a, GM03, Hea83, Mat22, SK92, SYK94, Ano81a, Ano94n, Ano16h, Bar95, BT22, BNS16, Cai94, CKT04, CCF23, CJH^{+}07, CCV11, CBK03, DDN24, DCD22, EKT^{+}22, FLC17, FGB^{+}24, GC99, Hay91a, HH08, HS19, JK03, JBZ23, JJS^{+}18, Kaa09, KDH02, KC07, KWvdS21, LS04, LDT^{+}23, LT10, LL22, LC22a, LLZ19, MRGT11, Mas86, MOWGB02, MZL15, MRDA24, Miy00, MG08, MPCRB^{+}03, MCT06, NQ05, PLC13, PC05, QGBV04, RPL10, SRT18, SPO08, SY05, SCS06, TWGB02a, TM03, TCC^{+}22, WLH10, WW21, WMGB01, WNGB03, YH09, dlHBR22, wK97]. Innovativeness [Bar97, SVÁG07, WPF06a, WPF06b, WPF06c, AKBA07, DLTQ23, FB93, GFE95, JSR19, JB19, LA15, Mud90, OB20, PF98, RB09, RPL10, Tan98c, ZLH17]. Innovator [CMtW10, Ano94n, Cai94, KI03, LG23]. innovators [BG10, CR81, Hay90a, HOGGP11, Ric95]. input [Ano94p, DB97a, Han94, PR97]. input-output [Ano94p, Han94]. inputs [DB97b, LW97, TAC96]. Inquiries [Min94]. insertion [SC05]. insight [Cho96, RRB^{+}22, SL14a]. insight-value-perception [SL14a]. Insights [Giu09, HKL24, Mol97b, Mol97a, Sar22, AAK^{+}24, Ban03a, CS01b, CS01a, CL24, LP06, MR19, SSKB24, BV09, LDF^{+}21, Mas23]. insitu [EKMS12]. inspiration [GFvT22]. inspired [Fis89]. inspiring [RL06]. Institute [Ano81b, Hay89a, Pot93, Ros89b, Why05, Hsu05, BH16]. institutes [CZZM21, Hay97a, LO96, LfSS19, MR05, ZCF19]. Institution [Ano07l, Bes02d]. Institutional [AAM24, Hew99, WK03, Won08, AW12, ABG23, Bar15, BK89, BSWC23, Bje10, CGBV21, EKY24, EEY22, FLC17, HB89, HS21a, Kra22a, LD05, OW00b, Sap89, SSW22, Tho88b, WMGB01, WKOW03, WWH15, ZYSL21]. institutionalisation [Cor05]. Institutionalization [YLB06]. institutionally [Mrk17]. institutions [BBTW00, Cur89, Gup00, KJ03, LJ01, XLC23, YHcHW17, YHcHW21, Nuv07]. Instructor [Pre01]. instrument [CS12, JBGMM11]. instrumental [GGGR22]. instrumentation [KI03]. Instruments [Bro92, BMW14, Mac97b, Ras08]. insurance [Kee14, UOIA00]. insured [Sap92]. Intangible [Jan21, KMID11, Car01, FMV00]. integral [Sto03]. integrate [FB06]. Integrated [KRA^{+}22c, WW14, BS06, BLS85, BB89, DR99, ETA02, HY04, KKRM05, Lay93, LGY05, LJAK21, Sgo95, VM11, WSH23, WC06b]. Integrating [BCKC24, Bes94c, FBE23, Hay98, IS97, IEG97a, IEG97b, JGM98b, JGM98a, XL23, Gol90, MH98, RMGBA08, SP96, TEA00]. Integration [BV09, BVHJ24, GBL03, AH85, AAN06, ÁRS17, CPT23, EKMS12, FdlCU22, GST08, GdWNPC03, HY99, HKR^{+}22, LB06, LT10, Lin18, O’K19, PDHW23, Whi89a, Whi90a, Whi90b, Whi90c, Edw04, Hay92b]. integrative [EKY24, Mud90, SdVGdFB24, Tan98b, Tan98a, XLD23]. Intellectual [CPSM12, DVdCAS16, Han06, Hob03, NHT23, Sar08, YHDR08, Ano94-30, ALCM14, Bad08, Bor01, Bou15, Dib02b, LPXO14, LH11, MFH16, SLJ13, SH10b, SDN23, SS94, Tam16, TLL23, Wes93a, Wes93b, WJK15, YC05, BE09, Aug06b, Coo97]. Intelligence [ZAI23, ABvZ23, BCPP23, CDED21, Cha94, CL24, DDFS23, IDMD04, MMMD23, Par03a, PK23]. Intelligent [Cha08, Car98c, NF24, Tan87]. intensity [GGS08, LWP14, LKCK22, TLL23, WSN^{+}18, WTN^{+}02, WA03, WN03]. intensive [Cha22, dCLSNL08, NGK08, PA15, SSF12, SFBK23]. intent [KOX21]. intention [CP23, xCZZ^{+}22]. intentions [BSWC23, CDED21, KFEC15, PD10]. Inter [BGR91, HS91, KGK07, Maj22, AGS98, Arc88, FGPP22, Fig02, Hag90, HESPSL11, KP12, KKG07, LV09, SLA^{+}23a, WQZ21]. Inter-firm [BGR91, HS91, Maj22, AGS98, Fig02, Hag90, HESPSL11, KP12, KKG07, SLA^{+}23a, WQZ21]. Inter-firms [KGK07]. inter-industry [Arc88, LV09]. inter-organizational [FGPP22]. Interaction [ZC19, AC07, Dea81, KA13, Nar89, OW00b, Sae05, Ved97, WKOW03, ZC22]. Interactions [BBZ20, GKB12, LS24, LM94a, LM94b, MCD06, O’K19, SC10, TLK09, TKT13, WA03, WF10, ZCF19]. interactive [HH08, SCH93a, SCH93b, TCHS95]. interacts [AC07, SB02]. intercompany [Hol90]. interconnected [Wil14]. interdependencies [AAT14]. interdisciplinarity [HS19]. Interdisciplinary [LDT^{+}23, AKKS12, ÁRS17, DMRS23]. interest [Kap99, WW04]. Interestingly [AS22]. interests [AD09, KL08a, SSS17]. Interface [CL05b, Hyv06, Ano93f, AS18, Dan19, ET98, KS10a, LM94a, LM94b, PWL20, PB93]. interfacing [RvdSdE22]. interference [LXODT15]. interfirm [KTSM10]. interfunctional [GST08]. interim [LD95]. Interindustry [Ano94p, Han94]. intermediarie [LZLL21, SVLYT24]. Intermediaries [HKL24, BS24, CCR^{+}24, CJL23, GKB12, HMFS24, How24, STA24]. intermediary [HCMMH21, Jus24]. intermediate [Joh08]. intermediation [HCMMH21, KL08a, Lar09, WvT23]. Internal [CKT04, GW10, LTT23, ZMH22, CMB23, DTD13, FORA17, GST08, HC13, JCT23, LW10, LG22, RUNS23, RB21, VM12, VLIC12, WH04b, CJL23, LSB07]. internalization [GBL03]. International [Ano07k, AC04, ARR23, BB95, Bes94a, Bes95b, Bes95c, Bes97, Coo97, Dib00, Hay88g, Hay97a, HG92, LY96, Lop06, Mer94a, ND15a, PSU97, Per96, Per97b, Rad99, Red90, Ano90o, Ano93m, Cai23, Cre04, DSPS23, Fra03a, Ger04, Gol82, Gun97a, Gun97b, Hay89d, JBL23, JW06, KL04, KL24, LL93a, LL93b, LL04b, NHS24, Nar89, NH99, PWL20, RC23, Saa98, SS21, Sér96, Shi24, TNSG96, Tei93, TI02, TA10, Ven96, Vit90, ZIB23, ZR93a, ZR93b, Ano87g, Ano87h, Lal00]. Internationalisation [RAG11, AEP21]. Internationalization [Pra10, AHSS19, KBSW08, RIMP22, SPKK08, ZCR22]. Internationalizing [Wya07]. Internet [Bul97, KAV17, Lop06, Par06, Wya07, ANI^{+}22, BVT21, BBZ20, CSL06, CHT22, DFFM22, EN07, GUMZI14, Kaf06, Kin00, KK07, KRA^{+}22c, LDK22, MSvdK^{+}21, Oye01, Pal00, PRA22, PC18, RMGDLC08, SRDS06, SPPG20, TLL07, WPF06a, ZWZ^{+}23, Wya04]. Internet-based [RMGDLC08]. Internet-enabled [PC18]. Internet-of-Things [PRA22]. interorganizational [SP22b]. Interplay [VNJ18, ASA^{+}22, AEP21, BWF18, BOBT22, KLP24, Ott98b, Ran24, YMMYK24]. interpretation [Ric85, SB91]. interpretive [CLZ^{+}22]. interregional [Ano93j, MS19, Wil93]. interrelationships [SOA^{+}24]. Intersector [Gei97]. intersectoral [CS05]. intervenes [MTB01]. intervention [CS13, Gru93, LvLB06, PW99]. intra [FGPP22, Mal02, LT16]. intra-cluster [FGPP22]. Intra-country [LT16]. intra-firm [Mal02]. Intraboard [ECB22]. Intranet [Fun06, Ott03a]. intranets [Car98c]. Intrapreneurial [AG24, HL24, FGPV24, GHU21, KUH21, vWGdG21]. intrapreneurs [EKY24, HS21a, MAU07]. Intrapreneurship [ALMW21, ABS24, DV23, EKY24, FGPV24, NHS24, PLG24, SSC23].

Intricacies [FCS21]. Intrinsic [Ish89a, Ish89b, KR12]. Introducing [JLWS23, OGOF22, Soh99, BLQS00a]. Introduction [BWF18, Ber89, Cam98, CY12, CMH14, GH15, Lin19a, OBL14, She09, Whi89a, Ano94-34, AEP21, BFM86, Den94, DR07, ETAO98, Hit89a, LBA14, MVS^{+}22, VSC16, SL14c, Joh92d]. Introductory [Lin06a]. intuitive [MNW14]. invented [AGFE22, dABKS14, IBPGML23]. inventing [LXODT15]. Invention [RG85, Gar87, JWY22, Per95, Vos84, Zak23]. Inventions [Wal97, BC10, BBA15, NS90, PG11, Thu05, Wan24, WWH15]. inventive [Gar87, JCT23, LP07, ML08a, Mas86, PNC15]. inventiveness [Gaz86]. Inventor [LLL23, Rob84, wtcoLSDD91, VJ23a, VJ23b]. Inventor-licensee [LLL23]. Inventors [DMNR12, AO10, VJ12]. inventory [Tan98c]. inverted [dlHBR22]. invest [NK22]. investee [TB84]. Investigating [CJL22, PRID23, PUFC23, GF21, YSKL21, Box01]. Investigation [Car00, ASGC09, BSG^{+}24, BR16, BF86, Clo91, Dag04, DTD13, GTW02, Har97a, Har97b, KS07, KMK03, LLB89, LH21, Pil04, RB21, SAM^{+}23, Sha03, WTN98b, WTN98a, WZSZ16, WH96, XLG^{+}23, lS06]. investigations [LL22]. investigators [O’K19]. Investing [JD09, Kim16, MPA05, Tom90]. Investment [Bea98, Gim06, CR06, CLIH05, GKB12, IEG97a, IEG97b, Lan98b, LCRL22, LYY^{+}22, MJ24, PGLS21, Par03b, Pen08, SY17, VLIC12, Wal07, WO08, ZIB23]. investments [ALM22, Kha87, LKCK22, PSG22, WL91, vZS14]. Investor [TB84, Wet81]. investors [ES05, HSLS24, Lan98b, Lan98a]. involved [BBTW00, OOJ^{+}00]. Involvement [WPF06a, WPF06c, vdEJG08, Bie91, Boa08, BB22, CCL^{+}06, FB91, Hal97a, HLNP21, HK86, LZD24, PWF07, Suc87, TDA05, WPF06b]. Involving [XWZ23, Ahn95, SM12b]. Inward [TCHS95, WLY14]. iOS [Kan22]. IoT [BVT21, BCV21, DHT22]. IP [EBG24]. IPS [Bes94a]. Iran [FAH08, Sar08, TNSG96, ZMPRP14]. Ireland [Hea83, MG98b, MG98a]. Irish [FC06, KM97]. Irreconcilable [Sap92]. irrigation [AO03]. ISBN [Ada07, Ano97b, Arr04, Aug09, Aug06a, Aug06b, Ban04, Bas04, BB95, Bes88, Bes93, Bes94a, Bes94b, Bes94c, Bes94d, Bes95a, Bes95b, Bes95c, Bes96, Bes97, Bes01a, Bes01b, Bes02a, Bes02b, Bes02c, Bes02d, Bes03a, Bon02, Box01, BH03, Buc91, Bul97, But96, Col97, CF07, Col08, Coo07, Coo97, Cou90, Cou91, Cre04, Dav92, De 04, Del97, Dre04, Dut07, Edw04, Fit00, For90, For91, Fra02b, Fra02c, Fra02d, Fra02a, Fra03a, Fra03b, Gar94, Gar95, Gar96, Gil96, Gil97, Giu09, Gri98, Hal97a, Har93a, Har93b, Har97c, Hay92c, Hay92d, Hay92e, Hay93d, Hay93a, Hay93b, Hay93c, Hay95, Hay96a, Hay97b, Hay97c, Hay97d, Hay97a, Hay01a, Hay01b, Hay01c, Her91, Her92b, Her92c, Her04, Hob97, Hob03, Hol93, Hop04, Joh92a, Joh92b, Joh92f, Joh92d, Joh92e]. ISBN [Joh92g, Kap01, Kap03, Koc09, Kra15, Kuc07, Lae03, Lae04, Lal00, Lim06, Lin90, Lin91, Lop06, Mer02, Min94, Min96, Mor91, Mor04, Mur91, New04, Nuv07, O’H99, OH04, Par03a, Per04, Per95, Per96, Per97b, Per97a, Pot93, Pov97, Poy97, Pra10, Pre01, Que06, Rad97, Ric95, Rot88, Rus94, Saa98, Sap05, Sap10, Sco89, Sha95, Szw07, Tan08, Tat04, Tra00, Ver90a, Ver90b, Ver91a, Ver91b, Wan04, Wil90, Won08, Wya00, Wya04, Wya07, dC07]. ISBN-10 [Kra15]. ISBN-13 [Kra15]. ISBN-3-7908-0051-1 [Lim06]. ISBN0 [Her92a]. ISBN0-304-31817-7 [Her92a]. Ishigai [Hit89a]. ISO [EFV02, MdVB16, MFH03]. isomorphic [UF22]. ISPIM [Bai18, CR19, Ano85d, Ano86e, Ano86f, Ano86g, Ano86h, Ano87f, Ano87g, Ano87h, Ano89y, Ano89z, Ano90n, HCT11, TKS18]. ISPIM-News [Ano89z, Ano89y]. ISPIM’91 [Ano90o]. Israel [Ano90o, Fre86, GM00, Har81, RAS15, Vek99, Won16]. Israeli [CA06, IMBH98, RD05]. Issue [Ano12a, Ano88d, Ano88e, Ano89h, Ano89i, Ano89j, Ano89k, Ano89l, Ano89m, Ano03q, CMH14, GH15, GFE95, HCT11, Jol03, Lin09b, Lin19a, MVS^{+}22, Ram81, She09, SL14c]. Issues [ADNP01, Ano93k, GG93, Kas94a, Kas94b, AB89, BK99, BK00, Col81, Gup88, HR98a, HR98b, HK02, Hit92b, KV24, KMK03, LL02, NK22, OS98, PÅ06, PS97a, PS97b, Ven14, Ver90a, Ver91a, Whi90c, KI14]. it-Advice [Lin14]. IT-enabled [YH09]. IT/IS [IEG97a, IEG97b]. Italian [Ano92c, Arc88, BCPP24, BNV05a, BDNP06, CT92, CCF23, CA07, CR02, DID21, EIMS01, FV91, GPR89, LLZ19, Mat22, NG97, NPV05, PV00, Rag93a, Rag93b, RDAV23, SCZ14, Tuz05, VGAL24, VOTH05]. Italy [Arc03, ADDS09, Bal10, GP04, NS90, PV01, Sir91]. item [LO10, MAMD22, SSP04]. items [Ano88k, Ano92g]. it(R) [O’H99]. IV [Sna90d, Whi90c]. ivory [HPT10]. iVP [SL14a]. ix [Gil97, Hay97d, Kuc07, Lal00].

J [Ada07, Ano91d, Ano97b, Bas04, Box01, Buc91, Dav92, Dib02b, Fra99, Fra00a, Fra01a, Fra01c, Fra01b, Gra03, Hay88e, Hay89d, Hay90a, Hay90b, Hay90c, Hay92c, Hay97b, Hay01a, Her92c, Mil85, Mor91, Nuv07, Par03b, Pot93, Pov97, Poy97, Ros89b, Rot87a, Sco89, Tat04, Wya04]. J. [Ano93l]. J.-C [Ano97b]. Jack [Hay88f]. James [Bes96, Box01, Hay97c, Per97b, FD21]. Jan [Gra01]. Jane [Bes93, Tan08]. Janice [Tra00]. January [Ano20d, Ano22o, Ano23o, Ano24i, Tit90b]. January-December [Tit90b]. Janus [HKJ07]. Japan [Ano94f, AKKS12, BiKM95, BCO14, BDL92, BP94, CJ98, Fuj98, Fuk09, Fun06, GW87, HBO14, Hit89a, Hit04, Hon21, HHA84, HT08, Kim87, Kod03, Kod07a, NWGB09, iSKB02, TNY^{+}22, TKSM08, WWM00, WA03, WK03, WKN03, WN03, WAN03, WH04c, Yam21, Giu09, Ros89b]. Japanese [AS18, CN90, FMvT00, Hit85, Hit92b, Mow90, OWC03, TWGB02b, TWGB02a, TI02, Tsu01, Tsu02, WKN04a, YW08]. Jarunee [Kuc06]. Java [JRS06]. Jean [Ano93a, Mer94b, Ano93o]. Jean-Marie [Ano93a]. Jeffcutt [Sap10]. Jeffrey [Bes95a]. Jelassi [Per96]. Jelinek [Cou90, Cou91]. Jenkins [Fra03b]. Jennings [Per97a]. Jens [Sap05]. Jenster [Hay93d]. Jerusalem [Ano90o]. JET [Ano03q]. JET-M [Ano03q]. jewellery [YSKL21]. Job [DL10, CLW10, CG04b, CGW05, Lic10]. job-shop [CGW05]. jockeys [HHV05]. Joe [Dre04]. Johan [Gra04, Hay90b]. John [Cre04, Del97, Dod87a, Fra00b, Fra02d, Gra03, Gri98, Hay85d, Hay85b, Hay86d, Hay92a, Hay93d, Hay97a, Her92c, Joh92g, Kap01, Mil85, Per95, Pov97, Pre01, Why02]. Johne [Hay86a, Sha91a]. join [MPO23]. Joining [PE22]. joint [CE89, FEP15, Gun97a, Gun97b, KS07, KGK07, Pet11, Pra99, RSA05, SD04, TI02, VH06, MPCRB^{+}03]. joint-patent [Pet11]. jointly [GD08, Jai23]. Jon [Wya00]. Jonathan [Hay01c]. Jonckheere [KHKL23]. Jones [dC07]. Jordan [AHAAEA03]. José [Que06]. Joseph [Her04, PF98, Mer94a]. Jossey [Fra01a, Fra01c, Hay88a, Hay88b, Hay88c, Mer01]. Jossey-Bass [Fra01c, Hay88a, Hay88b, Hay88c, Mer01]. Jøtul [Hol88a]. Journal [Ano11a, LE07, Lin10b, Lin10d, Lin11d, MdCÁ06, Ano03q, Lin11f]. journals [CbLjJL12, Lin06b, Lin09d, Lin10c, Lin11f, Lin13b, Lin13e, Tho12]. Journey [Ada07, dGSJ^{+}21, CCF11]. judgement [wK97]. judicial [DGS23]. Julian [Hay87c, Pav89]. July [Ano22p, Ano23p]. June [Ano07m, Coo16, Ano21u, Ano22q, Ano23q, Ano24j]. Just [AB89, McM96b, GARM24, GH91, PS97a, PS97b, ZT21, PS00b]. Just-in-time [AB89, GH91, PS97a, PS00b]. Justification [IMBH98].

K. [Wya07]. Kai [Gra01]. Kakao [Ano94j]. Kalman [JP99]. Kano [MH98]. Kapiti [AY05]. Karel [Ano93s]. Karlsson [Dre06]. Karolinska [BH16]. Kathleen [Pov97]. Katsundo [Ano94c]. Kauser [Pap03]. Kaushalesh [Coo07]. Keeble [Hay91a]. Keeping [ØD22, Sch09]. Kempston [BB95, Bes94a, Bes95c]. Ken [Bes97, Lae03]. Kenilworth [Bes02b]. Kent [Rot88]. Kenya [OOA04, wK98]. Kenyan [HKL24]. Key [JMPR01, Leh07, LG23, ACV08, BBTW00, BS08, CEH03, Hay85b, Ket09, Kum06, LS07, MLC24, New04, Ott03b, PA15, TRR^{+}24, WTN98b, WTN98a, vGS09]. keys [Sal03]. Keyword [LYP09]. Keyword-based [LYP09]. Khalil [Leh07, Lop06]. KIBS [BOBT22, RDS17, SLSCS24]. Kijne [Ano97b]. Kim [Wya07]. kinds [TLK09]. kinetic [HW01, Hal97b]. King [AHAAEA03]. Kingdom [Hay93c, Tit90b, WHP85]. Kirchhoff [New13, Ano11k, HKW^{+}13]. Kirton [Hay90a, Mud95, Ric95, Fox86b]. KISAs [LVS10]. kitchen [OTA01, TOA01]. kites [RA86]. Klandt [Rad97]. Klaus [Ros89a]. Kluwer [Ano97b, Hay97d, Par03b, Per04]. Kluwes [Joh92a]. KMS [LWC09]. knitting [PPB89]. Know [Fra01b]. Know-Who [Fra01b]. Knowing [BTP^{+}23, Mac05]. Knowledge [ALFM23, AGS98, BCL22, Bor01, Car99b, DQ06, Dre06, Fra01b, GR23, Giu09, Hen01, Hua09, JH09, KdOL08, Ken13, Kod06, Lae04, LS24, LH07, McA04, Mor04, PK06, ROMH15, RAS15, TI02, THHS22, WvdWG24, WH21, YKW10, ZIB23, AK06, AGS01, ASA^{+}22, AW11, AMP22, ABGS13, AAM24, AC04, ÁRS17, BKM10, BBS22, BCH98, BS99, BHPB06, CB22, CPT23, CKT04, CMtW10, Car98c, Car99a, CAI00, CPSS06, CYK19, CL21, CHO98, CJL23, Cre04, Dag04, DGP10, Dav92, Day02, DVMS16, DJ05, DF98a, DF98b, DVHL22, EBG24, FMF21, FLH^{+}02, FRA05, FMFMA16, Fon05, FDN^{+}23, FP08, FEP15, GGS02, GL22, GCH96, GBM98a, GBM98b, GG11, Han95, HHV05, HESPSL11, HEC22, Hu13, JBL23, Jai23, JWY22, KA15, Kes03, KL08a, KOX21, Kod05, Kod07a, LVS10]. knowledge [LWC09, LM14, LS04, LTT23, LT09, LT10, hLFL08, LW10, LCT04, LCT05, LOCSBGMH16, Maj22, dCLSNL08, MRGTDG21, MOM08, MLF16, Mig19, MJ24, MPWG22, NL11, NQ05, NI05, Ott03b, PWL20, PS21, PRID23, Pur91, RMG06, RZ21, RSA05, RR11b, RDS17, RvdSdE22, RC23, SPKK08, Sae05, SSF12, SBP21, SP22a, SR23, SP10, SFF21, SFBK23, SL16, SCS06, TLK09, VJ23b, VRS^{+}22, WCZZ14, WL21, Wan24, WJK15, XHM^{+}16, Zak23, ZZXW23, ZSX24, ZLJ^{+}21, CMP^{+}06, ZT21, Fra03b, Bas04, Dre06, Koc09, Nuv07, Pap02a, Sap01]. Knowledge-based [Kod06, BKM10, BS99, DGP10, Kes03, LVS10, Pur91, ZIB23]. knowledge-flow [FLH^{+}02]. knowledge-intensive [dCLSNL08, SSF12, SFBK23]. KNOWVATION [PK06]. Knudsen [Hay97d]. Knut [Hay85b, Hay85c, Hay88d]. Kogan [Gil96, Hay89a, Joh92f, Joh92e, New04]. Kogut [Par06]. Köhler [Har93b]. Kondo [Giu09]. Kondratieff [Won05b]. Kong [Arr04, CC14, Che94, CTC02, Kim87, PCWG04, SW05]. Kontrolle [Ros89a]. Korea [Ahn95, BH15, CLR96, GBL03, HBO14, HT08, Kha98b, Kha98a, Kha99, Kim87, KLL87, KI14, KL24, LLS00, Lee04, dLP06, LS07, LKKH22, RPL10, SCY22, SH10b]. Korean [AKL05, Ano94q, Ano94-45, tBL86, BA89, CY00, CL97b, CL97a, Chu01, CBK03, CN92, Hit02, HH08, KP12, LL94a, LL94b, Lee00, LKPP07, OCA09, PH08, SCH19, SH99]. koreanischen [Ano94-45]. Kristin [Gra01]. Kuno [Par03b]. Kuwait [AA91]. kwaliteit [Bes95b]. Kyoto [Lin13c].

L [Bec89, Bes94a, Dav92, De 04, Dib02a, Fra01a, Fra02d, Gra03, Hay97c, Her92a, Joh92a, Leh07]. lab [HM11, SSS17]. labor [ALMW21, KZ18, PSG22]. Laboratories [CG02, Ano94f, BC91, BC92, BP94, CRKA98, Mow88, NI05, Bat12]. Laboratory [WZNG17, CM04, FLKF^{+}06, RAC84]. labour [NW23]. labs [PC21, SSCH23, SS23]. laggards [AS21]. Lake [Mer94b]. Lal [Coo07]. Lalejin [GSFS21]. Lall [BH03]. landscape [HSLS24, LJAK21, SB02, Tho12, WML^{+}23, Fra02a]. landuse [SO02]. Lane [Gar95, Gri98, Hay93c, Per95, Pov97]. Langdon [Ram86]. language [CGS^{+}24, Jus24, Lin09c]. Lanham [Wya04]. laptop [LC22a]. large [AO03, ABKL08, Bac98, CGS^{+}24, DTI^{+}23, ED87, GMLMVJ07, Hus10, Kar86, KC07, Lin86, LLZ19, MS98, MM11, Rie89, RFB24, SU11, SPO08, VS06, WSN^{+}18]. large-scale [AO03, ABKL08, KC07, WSN^{+}18]. largest [GM00, Rue86]. laser [Gal12, WMGB01]. Lastres [Que06]. late [GBL03, OH04]. latecomer [FCS21, LfSS19]. latent [NMN11, WGW22]. launch [CCST06, Min96]. Launching [EH02, Won04]. law [Fuj98]. laws [Ish89a, Ish89b]. LCD [CLT06, Tsa13b]. LDCs [AHSS19]. lead [BCO14, BF23, DL10, GF21, HLNP21, MGdG11, Mas86, SAWT06, VNP^{+}24, WA04a]. lead-user [MGdG11]. leader [AGS98, KK05, SW22]. Leaders [Fra01a, AS21]. leadership [BC12, Far23, Fra02d, Gra04, Kod05, Kod07a, MMM05, OR06]. leadership-based [Kod05, Kod07a]. Leading [Far23, FWL06, LDF^{+}21, Sch09, Hop04]. leads [HWC06]. leakage [EBG24, FEP15]. leaking [ROMH15]. Lean [MG21, PS00a, AM97, Bes03a, JBZ23, KK04, MKL19, Oll08, CWBG03, JBZ23]. leap [RCC04, Wan04]. learn [Bes97, KLRR22, TRW23]. learned [CEH03, CvZ05, CPSS06, KT04, KS04c, RTY01]. Learning [AKL05, ALBO08, Bad08, CYK19, Chi05, Coo07, Fig02, Fra01a, HO14, Mac05, OOL06, SD04, UR18, Wec15, AKBA07, AGS01, AC08, Ano93p, BM23, BSSF24, BF99, CD92a, CD92b, CJ98, Car98a, Car98b, Car98c, Car00, CAI00, CF01a, CA02, CEH03, CP05, CPSS06, CH04, CQ03, CW06, CTSJ24, CZL^{+}24, Dag04, DR99, GMLMVJ07, GBL03, GG11, HF18, Har97c, ILR^{+}12, Kin02, Lee04, LLR95, Lev93, hLFL08, MRGMBR13, MD03, MMR24, MMM05, Mus09, Pap02a, Pap04, PSS00, RSA05, RPL10, Sen96, SP22b, SAM^{+}23, SPP98, Ste14b, SMW21, TSPJ23, TI02, WZNC15, WA04b, WLO09, Woo85, Xie04, YCL22, YXH21, HB97, Bon02, Mer02, New04, Ano98k]. learning-by-licensing [WZNC15]. leasing [Aka98b, Aka98a]. least [AHSS19]. leather [Sto03]. led [BBK23, GMB^{+}23, HL22, MKK04, Pat24]. Lefebvre [Ano93o, Ano13h]. legacy [HKW^{+}13]. legitimacy [EFC23, HMM14, De 23, LWM24, LDMY22]. legitimacy-building [LWM24]. legitimate [CP07]. Leibfried [Pov97]. Leifer [Bes01a]. Lembke [Gra04]. Leonard [Bes01b, Ros89b]. Leong [Hay97c]. Leonie [Pot93]. Lesley [Hay93b]. less [FC95, Had99, RS03, RB21]. lesson [HOGGP11]. Lessons [BLA06, Cou90, Cou91, FRW06, Hay88a, Lin17c, RTY01, Ano93l, Ano94-30, Arn87, BCO14, BP01, CEH03, CvZ05, CPSS06, DR24, HB97, Hob90, Ket09, KT04, KS04c, Lin91, NLLC05, PBC93, Rot92a, Rot92b, SS94, VK03, WH04a, WT19, Wer03, wK97, BMW14, DBV97, EIMS01, MMMT12, Ras08, Rea21, Sar08, Se14, ZMPRP14]. Let [MAB^{+}23]. Letter [Lin13c, KI14]. leur [Ano94v]. Level [Ano07l, BC10, BH16, BMLL07, BLLEM23, CM04, EB23, EHHW17, FSA23, FCS21, Gar05, Gee06, GW10, HBS07, Kaa09, KJ97, KAV17, LY96, ML08b, MMMT12, MMM16, NGK08, RUNS23, STA24, SRDS06, SFdO^{+}23, SPFG12, SM12b, TDC11, VMWV16, WTN^{+}02, XX21, YMMYK24, YT09, dVdWNF^{+}10]. levels [Fel23, JJS^{+}18, LOCSBGMH16, SM09]. Leverage [Bøl12, ACBO23, RI04]. Leveraging [CB22, CAI00, HF18, Kes03, SXS24, Car98c, Red14]. levers [LW21]. Levin [Ano93p]. Levy [Hal97a]. Lexington [For90, For91, Lin90, Lin91]. Li [Col08]. liberalization [For99]. libraries [KS00]. Library [Dib00, Hay93b]. libre [YA11]. LIC [JBZ23]. licensee [LLL23]. Licensing [CCB20, BEH99, Cab19, CHO98, DMH22, Ene22, Fuk09, HMM15, HJSYJ12, KLRR22, LLL23, RR88, SHT23, Sla23b, WLY14, WZNC15, WWH15]. licensing-out [Ene22]. licensing-patent-portfolios [BEH99]. licensors [KLRR22]. lieu [LTT23]. Life [OKvB23, AT93, Ayr88, AM92a, AM92b, BK00, BK01, CK22, CA14, EM02, GC93a, GC93b, HM11, KS00, Kim03, New13, Pyk00, RG06, RL06, SY17, SB91, Wer03, XL23, ZMPRP14, CL12]. life-cycles [AT93]. lifecycle [MM08, VNP^{+}22]. Lifetime [HKW^{+}13, SH95]. light [CW06, DTD13, MNW14, PP09, Szá92, WMGB01]. like [Lin11d, SL14b]. likely [BNV05b]. liminality [KM22]. limit [WKN04a]. limitation [Thu05]. limitations [Gol86, Nar04, Roe87]. Limited [Hay83, Hay86f, Mer02, Mor04, Why02, Hay82]. Limiting [AM23]. limits [AB23, DDBM23, De 85, Hüb96, Lin20]. line [CE19, McM94b]. lineage [JL24]. linear [xCZZ^{+}22, DVSL19, DB97a, DB97b, MRGT11, Phi99]. liners [BM05]. Link [Ano94m, KKL19, Per04, Ros89c, BG19, GF21, LFU23, Ran24, Dav92]. Linkage [YI05, CK12a]. Linkages [LWL^{+}22, AL11, CD92a, CD92b, CM95, JH18, dLP06, OIN96, Aug09]. linked [Fal05]. Linking [Bar92b, CA07, EWJ09, EFV00, EFV02, HW97, LB03, Nob04, SPKK08, SS04, SCS06, Pap95]. links [BO81, Cai23, DVMS16, LM24, QWM92, SS05]. liquor [SK92]. List [Ano86i, Ano88i, Ano09l]. listed [XW21]. Listen [PLC13]. literature [AGFdMCM24, ACBO23, BLA06, BD23, CPT23, CFP^{+}23, CLGFVCM23, D’I14, DMRS23, DB19, FHA22, FLYTB24, GLÖS23, GBY22, Han06, HPBP98, KK23, LAC24, hL05, MdVB16, NPG14, OS99, OGRA23, PMP24, PS97a, PS97b, PS01, RPASLM23, RC23, SP22a, SS23, Sna01, SMD23, ZKW^{+}23, dBJ03]. litigation [EBG24, WZY23]. Little [Mil88, SMW21, Ano94b]. Little/Cambridge [Ano94b]. Littlefield [Wya04]. lived [KM22]. Liverpool [Coo16, Per97a]. living [PC21]. Ljungberg [Gra01]. loading [NJ04]. Local [Boh18, Ano94-31, CvZ05, CT12, Fuk09, Ilo94, McN92]. located [DLL06]. locating [LB06]. Location [Jar86, LVS10, BS05, DLL06, DS23, HN08, MTB01, Mal87]. Locational [SI98]. Locations [Pap06]. lock [Gos05]. lock-out [Gos05]. logic [SM09]. logical [Ven96]. logiciel [Ano94s]. logics [EKY24, LDMY22]. logistics [SS03]. Lois [Bes01a]. Lomond [Hay88f]. London [Ano81b, Ano91d, Ano91c, Aug06a, Aug06b, Bes97, Bes02d, Col97, Dre04, Eva02, For85, Gar85, Gar95, Gar96, Gil96, Hay83, Hay84, Hay85a, Hay85c, Hay86c, Hay86e, Hay90a, Hay90b, Hay90c, Hay90d, Hay91a, Hay92c, Hay92d, Hay92e, Hay93d, Hay93a, Hay93b, Hay97b, Hay97c, Hay97d, Hay97a, Her92a, Hob97, Joh92f, Joh92d, Joh92e, Kuc07, Lae04, Leh07, Min96, New04, Pap06, Par03a, Par06, Per97a, Pra10, Ram86, Rea06, Ric95, Sap10, Sha91b, Whi94, Wil90, Wya00]. Long [GS05, BCP22, Hay86c, SCC17, ZLC18, Ano94l]. long-run [SCC17]. long-term [ZLC18]. longitudinal [ABUS05, CLZ23, CA02, GS10, PCWG04, Fel23]. Longman [Hay86f]. Look [SH22, Ano93l, FMFMA16, Guy96a, Lin11d, NHV05, PBC93, TM03]. Looking [ABG23, DJ13, NGK08, JSR19, LFJA23]. loop [CYK19, KM06b]. loop-out [CYK19]. loss [Jai23]. Lost [THBW22]. Lothian [Kin00]. Lotka [WKN03]. Louis [Ano93o, Ano13h]. Louis-André [Ano13h]. Love [Hay93c]. Loveridge [Joh92g]. low [AAN06, Boh18, BMRG00, Car01, CX24, HOGGP11, MZG17, OG05, RV21, Sun96, ZHY22]. low-capacity [RV21]. low-carbon [CX24, ZHY22]. low-emission [Boh18]. low-performance [Car01]. low-tech [Sun96]. Lowe [Hay87c]. lower [LHI13]. Loyalty [WPF06a, WPF06c, WPF06b]. LP [ND15b]. LSA [KKP17]. Ltd [Ada07, BB95, Bes94a, Bes95b, Bes95c, Gar85, Gri98, Hay86a, Hay92d, Hay95, Hay97b, Hay97c, Her92a, Mer94a, Mer94b, Oak97, Per95, Per96, Per97b, Per97a, Pov97, Rus94, Won08]. Ltd. [Fra00a]. Lu [Arr04]. Lucia [Que06]. Luciana [Mor04]. luck [WG23]. Lundvall [Pap04]. Lüst [Lüs89]. Lybrand [Gil96]. Lynn [Ros89b].

M [Ano91d, Ano93j, Ano03q, BB95, Bes96, Dib04, Dre04, For90, For91, Fra02c, Giu09, Hay84, Hay89d, Her92c, Jel07, Joh92a, Joh92d, Joh92g, Leh07, Lim06, LOCSBGMH16, Lop06, Mil85, Par03b, Poy97, Rea06, Rot87a]. M&S [Ott98b]. M-SME [LOCSBGMH16]. M. [Ano94-31]. MA [Bes95a, Bes96, Dib02a, For90, For91, Hol93, Lal00, Lin90, Lin91, Que06, Won08, Wya07, Sap01]. Machine [Kra15, Ano93i, CZL^{+}24, Dav93, KM01, KM06a, Par03a, Vit90, VX97]. machinery [tBL86, WA03, WN03, WH04c]. machining [Ano94m, LQ94]. Maciel [Que06]. Mackenzie [Hay90b]. MacMillan [Coo07, Fra00a, Giu09, Hay84, Koc09, Oak97]. macroeconomic [FNJV14]. made [DFD^{+}91]. MADM [FSA23]. Madrid [Pis89]. Mainstream [Ros89c]. maintain [Ser13]. maintenance [Oye00, Sna90c, TOA01]. Maitra [Bul97]. major [Kra15, Kum05, Llo07, Sal03, Sap89]. Make [YWL22, FCCE21, HPT10, MH98, NP10, PLC13, Gal06a]. maker [HJM08]. makers [AB99, MT08]. makes [Fel22]. Making [Gri06, HOGGP11, LWM24, Ric85, ANI^{+}22, BBLW24, BCPP23, CDED21, GC93a, GC93b, II97, IEG97a, IEG97b, JGM98b, JGM98a, KR12, Lan98b, MSD15, RI04, SI98, SW91, SL14b, Bro07]. Mal [Bes02b]. Malaysia [HO14, RCC04, SAY09]. Malaysian [Moh95, SS03]. Mall [HST07]. man [FG06, CCY05]. manage [DF98a, DF98b, FHA22, HKR^{+}22, SMvFM24]. manageable [BiKM95]. managed [Mat22]. Management [Ano81b, Ano03q, Ano07k, Ano07m, Ano09l, Ano11a, Bon02, Chi90, CNPV04, Fra02b, Gup89, Hay88a, Hay88b, Hay88c, Hay89a, Hol99, Hos01, HGL98, IOS00, Kuc06, Lae04, Lim06, Lin11f, Lop06, NW03a, OBvE97, PT06, Rus94, Szw07, TM06, Why05, Won07, WO08, YKW10, vW88, AH85, AHAAEA03, AK06, ABLK07, AS22, Ame95, ACV08, AMP22, Ano87g, Ano93k, Ano94-34, Ano94-33, Ano94-43, Ano97b, ABG23, Aug06b, ÁRS17, BBDM01, Bar15, BCPP24, BBS22, BSG^{+}24, BE09, Bes94c, Bes02c, Bes02d, BS99, Boy14, BRH^{+}97, BV09, BD05, BBHA01, BEH99, BGM88, dABKS14, BFM86, BVHJ24, CPT23, CPSM12, Car98c, Car99a, CWAB09, CPP09, CCF11, CTC02, CG04a, CA07, CA14, CR19, CO04, CW94, CS98, DB09, DSPS23, DHT22, Den94, DLL06, Dib02a]. management [Dod88, DBR24, DO99, DO03, Dre97, Dur86, EKT^{+}22, Eng06, EFV00, FSA23, FCH01, FF92, Fra03a, FP08, Gal06b, GLÖS23, dVGCOF^{+}23, GG93, GG03, Han06, HR98a, HR98b, Hay85d, Hay85c, Hay88d, Hay92d, Hay93a, Hay96a, Hen01, HJM08, Her92b, HB03, Hit85, Hit92b, Hit93, Hob97, IBL04, JH09, JL13, Joh92g, Jol03, JC01, JW06, Kha99, Kod07b, KLL24, KMK03, KK05, KB06, KS04c, LB21, LO94, hLwChLbL04, hL05, Lin06b, Lin09e, Lin10c, Lin11e, Lin12d, LCT04, LCT05, LM94a, LM94b, LL05a, LMMPVO09, MKGL05, MHKZ24, MH05, Mer94b, Mia94, Mik01, Mil02, Min94, MGAP23, MPM09, NW03b, OSG04, Oak03, OOKBD14, OKP13, OC85, OBL14, Ott03b, Oye00, Oye01, PPM24, Pap02b, PK06, PMT23, POGBG06, PPP98a, PPP98b, PFP06, PS00b, RES^{+}24, RSA05, Rue86, SCS22, SAY09]. management [SZLR19, SD23, She09, SS04, SS05, Soh97b, SL16, Sto03, STH06, TC97, Tan87, TB24, Tho99, Tho12, TAC96, Tõn15, Tsu02, TKS18, Ved81, WvdWG24, WLH10, WvT23, WFW02, WW14, XL23, XLG^{+}23, YHDR08, ZIB23, ZS04a, ZS04b, dBJ03, vdVdJVdR09, Ano87h, Ano07l, CMP^{+}06, Dre02, Lop06, Dre06, Fra01b, Bes95b, Bes03a, Hay86a, Hay89a, Hay93d, Her92c, Joh92f, Rot88, Dod87a, Hay92a, Kaz03, Leh07, Mer94a]. management-Insights [BV09]. manager [Bul97, CDED21, CHD18, Col97, DP23, KWvdS21, Mal02]. Managerial [GC96, HB03, PÅ06, Whi89a, Whi90a, Whi90b, Whi90c, dBJ02, Car99a, CS87, CMS23, FWL06, LOH91, WPF06c, WSN^{+}18, BSMC23, Hay92b]. Managers [Fox86b, Min95, Ban03b, MPA05, MCMA09, PSdLC02, SWW12, SR22, SDFN01, Van10, WC85]. Managing [AW11, Bad08, BLNP05, Bro01, BM94a, BM94b, CMH14, CMP^{+}06, DNP01, DFBC98a, DFBC98b, GGS02, HS02, Hay98, Hob97, Hob03, IBPGML23, Jon05, Kim03, KV14, KUH21, KLV06, LYST22, LG22, Lov85, NUSR21, NI05, RB21, SM97, SR22, Sod08, SW22, TA98a, TA98b, VSC16, VS06, Whi89b, ZI01, ZR93a, ZR93b, Ano93f, Ano12a, Bes94b, CY12, GBM98a, GBM98b, Hay87d, Hay88a, LO10, LT09, Mow89, OM88, PB93, Ugh88, Wat90, Gof07, RFB24, Ano98k, Del97, Dib02b, Dod87b, Fra01b, Gil96, Hay88b, Hay88c, Her92a, Pre01, Saa98, Tat04, Ver90b, Ver91b, Joh92e]. mandate [ZDWG14]. manifest [EDM05]. manifold [Bar96]. mannah [CA01]. manpower [HRHW16]. Manual [Pre01]. manufactory [Bes15]. manufacture [Ano94a, RGSK94]. manufacturer [HKY22, KM03, KM06b]. manufacturers [BS97, GF21, HH08, MS98, TCMS22]. Manufacturing [Bro98, ETA02, Hit96, MH05, Old97, Par00, Sar95, Soh97b, SGFS99, AA88, AT93, AL05, ALBO08, AGM89, And21, Ano94m, AP08, AS18, ANI^{+}22, BBDM01, BS97, Bea98, Bea05, BLA06, BLS85, BBHW93, BF99, BS05, BBHA01, Bro94, Bro01, Bro07, CD92a, CD92b, Cha22, CCL^{+}06, CF01b, CPXC02, CTC02, DVdCAS16, Die00, ETAO98, ETOA00, FRW06, FdlCU22, Gol87, GÁNA07, GTW02, Gup88, Gup90, GC96, Hal96, HMWT23, HS02, Hay86d, Hay88c, Hit85, Hit92b, Hit02, Hit03, Hit04, Hit05, HWC06, HO05, IAO^{+}02, IEG97a, IEG97b, Joh92a, JRS06, KBSB22, Ket09, KAV17, KL04, KBC22, Klo87, Kob97, KBRR24, KH89, KBS01, KKRM05, LVVH23, LB03, LPS03, Lay93, Lee23, LLEB05, hLwChLbL04, LQ94, LW97, LDMS23, LGZ22, MKK04, MS98, MW06, NG97, NJ04, OG05, OGS06]. manufacturing [OGG06, OVD21, OS98, PGKF19, PC05, PPP98a, PPP98b, Pia02, PSS00, PdW00, PCWG04, RSP10, RR88, RAATM05, Rot91, Rya96, San95, SNM12, Sar22, Sgo95, Sha03, Sna96, Sna98, SM01, Sto03, SSRO24, Sun00, SV01, SH02, SD04, Tag97, Tan95, TS00, Tit94c, Tit94d, Uzu01, VBA06, VM92, WC06a, WA04a, WSH23, ZYZZ24, dGSJ^{+}21, BOBT22, Hay01b, Bes94a, Hay01b]. manufacturing-based [IAO^{+}02]. many [DMH22]. map [Fra02a, LWC09, LYP09, MTB01, Pen08]. Mapping [BWS11, CG11, WML^{+}23, vdWFvdB23, Bre23, Gre06, Fra03b]. maps [CS06]. marathon [Cou90, Cou91]. marble [AAA02]. March [Ano88h, Ano21s, Ano21v, Ano22r, Ano23r, Ano24k]. marco [Ano94r]. Marconi [Hay97b]. Margaret [Bes02a, Why02]. Margetts [Tan08]. marginal [RS03]. marginalised [PW99]. Maria [Hay90c, Que06]. Mariann [Cou90, Cou91]. Marie [Ano93a]. Marina [Won08]. maritime [DM08]. Mark [Aug06a, Bes01a, Bro90, Fra03b, Hay95, Won08, Car98a]. Market [Arr04, CK18, Dre04, FdJ09, GC99, Hay98, Hol88a, JFCFMC^{+}12, AABY^{+}23, AAAA23, AKL05, AJLR04, BK04, BCS06, Bes95b, BC92, BV09, BRP94, BA89, CCB20, Can16, Car98a, Car00, CA02, CUF13, CK12a, CU13, CK12b, CFC12, Coo90, CN90, CN92, Dan19, DAOO02, GST08, HBvdB09, Hol94a, Hol94b, II01, JLL10, Joh99, Kim16, KET22, KB09, Kum05, Lae03, LRH22, LT21, NLLC04, ND15b, NT16, Ott04a, PLC13, PSS23, Puj06, SRHL12, SVÁG07, Sap92, TW02, TWGB02a, Thu18, TS88, VNJ18, WTAW97, Wan89, Wan91, WWM00, WLO09, Wer03, YHcHW17, YHcHW21, dHOvdKS16, dlHBR22, Col08, Joh92a, Joh92c, Pap02b]. market-based [dlHBR22]. Market-driven [GC99, Kim16]. Market-oriented [Hol88a, Hol94a, Hol94b, KB09, Pap02b]. market-upstream [KET22]. Marketing [Hay85a, ASI99, Ano93f, AB89, Bor01, ET98, Her92a, Her92c, JR91, Lin91, Mas86, MPCRB^{+}03, PB93, VB05, Lin90, Her92b, Joh92d]. Markets [HJSYJ12, WZNG17, CS13, DMH22, Hyv06, KBO22, LHY04, LHI13, NPG14, PNC15, RKFR24, VJ12, YWL22, ZWD09, ZZXW23, ZYZZ24, Sha91b]. markets-implications [Hyv06]. Markov [JLY23]. Marsh [Bes94d, Bes95c]. Martin [Hay92c]. Martínez [Ano94-27]. Martins [Que06]. Marwick [Hay88e]. Mary [Mur91]. Maryann [Per04]. Maryland [Hay88f, Wya04]. Mascitelli [Bes03a]. Maskell [Sap05]. Mass [Bes01b, Dib01, BM05, HJ02, Hit89b, JMT03, ND15b, WCZZ14, WZSZ16]. Massachusetts [RL06, SW04]. Massive [AAD14, GHU21, Jac22]. master [Wya04]. Mastering [Mer01, Bes96]. mastery [GMLMVJ07]. Matching [Wür17, Ski84]. matchmaking [LLL23]. material [Che12, NWGB09, Lin17a]. Materials [Fis89, HS91, JV97, Kim16, LKPP07, MPA05, WR95]. materials-innovating [WR95]. Materials-inspired [Fis89]. mathematical [CLW10, Sto99]. Mathews [Kap01]. matrix [Ano94p, Han94]. Matsushita [Kod07b]. matter [Bla16, DMNR12, DS23, Gal05, Gar95, HLL12, MFH16, SDK24, XLG^{+}23, ZCR22, vRE21]. Matters [Koc09, RR11b, TLL23]. mattress [Sch09]. Mature [Bes01a, Kim87, OB20, OB21, PPB89, VOTH05, WC85]. maturity [Cho96, SD23, WMGB01, dOK12]. Maureen [Dib01]. Maximilian [Hop04]. May [Ano88h, Ano07k, HdWN21, Wya00, DVMS16, Ano21w, Ano22s, Ano23s, Ano24l]. Mayes [Her91]. Maylands [Per97b]. Mayle [Szw07]. Mayon [Hay87d]. Mayon-White [Hay87d]. MBA [Fox86b, Mur91]. McAfee [Kra15]. McDermott [Bes01a]. MCDM [CCY05]. McDonald [Joh92d]. McKelvey [Dib01]. McKeown [Ros89b]. McKiernan [Her91]. McLaughlin [Tra00]. McLintock [Hay88e]. McNair [Pov97]. McNeil [Ano91c]. me [MAB^{+}23, RHL09]. Meadows [Mil85]. meals [II91]. Mean [WO08]. Mean-Variance [WO08]. meaningfully [DR99]. means [DB97a, DB97b, KL08b, OGOF22, SP96]. measure [Bar92a, Jac13, JBGMM11, LO10, Sna90a, Sna90d]. Measurement [Ano03q, BBS22, Cal24, CS08b, DG14, Ehr95, Gar05, GWM05, MAMD22, PF98]. Measures [McB13, CE89, CLZ^{+}22, Ell81, IPP14, Møl91, Rue88, SKTM08, Whe88]. Measuring [ADL07, AP96, DM11, DB97a, DB97b, EIMS01, GC10b, KTE04, KK13, LC22b, Ram23, RMGBA08, RBBF23, AGS01, ABKL08, Bar96, BS05, CS12, GMNL07, MAS13, Sna90b]. meat [FS84]. mechanic [KdOL08]. mechanical [CF01b, Ern95, FPP24, Hit93, KDH02]. mechanism [Bro85, CLP23, MLC24, PPM24]. Mechanisms [OGRA23, XX21, YH09, DL10, EW06, GBFvdS11, Hsu05, LNB21, LBG14, MPV13, Neu12, SK23, SLA^{+}23a, SFF21, SMD23, Tit90a, YHL22, ZLH17]. Media [Ved81, GLÖS23, McG98, OSSC21, PRID23, SPMRS23, ZXX23]. mediate [FD21]. mediated [TKYM24]. Mediating [FHBDO16, ALM22, BCPP24, BRA^{+}23, DL10, FMOPSVA21, LZLL21, MZG17, MRDA24, SAY09, SLSCS24, SXS24, SCM24, WCZZ14]. mediation [AAGL23, DLKM24, SWW12, RPL10]. Medical [WZNG17, Bie91, Cha22, FC06, Kod00, Sha86, Sha93a, Sha93b, SY17, WvdVM07, lS06]. Medicare [SSGN23]. medicine [Hja09, SSGN23]. Medium [HRHW16, Arc03, tBL86, CY00, Cor87b, CL88, DTI^{+}23, EDM05, FRW06, FDN^{+}23, Hay90c, KdOL08, LOCSBGMH16, OCA09, Ola98a, Ola98b, OOL06, RATW15, RMGDLC08, Rie89, Rot91, SPKK08, SU11, Tit94c, Tit94d, Wal12]. medium-size [Tit94c, Tit94d]. medium-sized [Arc03, Cor87b, CL88, EDM05, FDN^{+}23, GC96, Rie89, Rot91, SPKK08]. meet [KHL15]. Meeting [Ano85e, Ano87i]. mega [LHI13]. megacompetition [WH04c, WKN04b]. megaprojects [CLZ^{+}22]. Meil [Har93b]. mejora [Ano94r]. Melvin [Col97]. member [AC07, Cor87c, CL88]. members [Ano88j]. membrane [BJ05]. memoriam [Ano88h]. Memory [Ano11k, ZMJM23]. Menasco [For90, For91]. Meng [Ano93e]. Mensch [Ano94g]. merger [JGM98b, JGM98a, KKG07, LQ09]. merger/acquisition [KKG07]. mergers [ALRSD04, CS87, dMD05]. merrier [Egb15]. mess [GGGR22]. Messages [Kle13]. Meta [BS24, RVW17, DPH^{+}22, ODM23, XLC23]. meta-analysis [XLC23]. meta-analytic [ODM23]. Meta-organizations [BS24]. Meta-organizing [RVW17]. meta-regression [DPH^{+}22]. metal [II98, KdOL08]. metal-mechanic [KdOL08]. metamorphosis [GP22]. metaphor [DF98a, DF98b]. Method [Ano94l, Bar90a, CSH94, DB97a, DB97b, EEY22, GFvT22, LCT05, LGZ22, MTU23, PP04, Tan87, lS06]. Methodological [DBR24, PPW06, dV94]. methodologies [CC05, hL05, RRSS07]. methodology [AAT14, BGM88, CP05, Kim16, KGS^{+}24, MKL19, ML05, Sér96, WKF24]. methods [CR19, EKMS12, ETM96, HRHW16, IMIY87, SF98, WC85]. Methuen [Hay97b]. metric [AGS01, Tam16]. metropolitan [Die00]. Mexico [Bai84, SC05]. mHealth [GKK23]. Miami [Ano07k]. Michael [Ano93s, Bes03b, Hay90a, Hay97a, Hol93, Rad97, Ric95, Rus94, SW04, Sap01, Sod08]. Michel [Rot88]. Michie [Hay01a, Hay01c]. micro [CE19, FGPP22, LH19, LOCSBGMH16, OOJ^{+}00, Sma01, YMK^{+}21, wK98, Bje10]. micro-enterprises [OOJ^{+}00, wK98]. micro-foundation [CE19, LH19]. micro-foundations [YMK^{+}21]. micro-geographical [FGPP22]. Micro-institutional [Bje10]. microchip [Mas86]. microdata [MPV13, SG24]. microdynamics [Eva02]. microelectronics [GEG12]. microentrepreneurs [wK97]. microfoundations [BT22, RGH19]. microsatellite [SM97]. microscope [VGQK^{+}10]. microscopic [AH85]. Microsystems [Car98a]. mid [Fox86b, LAYS19]. mid-career [Fox86b]. mid-size [LAYS19]. middle [BBDM01, Thu18]. migration [FC06, MV06, NLLC05]. Mike [Edw04, Wya00]. Milan [Pis89]. milestones [THC^{+}23]. milieu [KLH05, LL03]. military [Har81, LCLC03, Rep08]. military-industrial [Rep08]. Millar [Hay88b]. millennium [Pap03]. Miller [Ano94z]. Min [Saa98]. mind [FWL06, Pen08, Rob84]. mindsets [AB99]. Mining [PN11, WCT10, BVS24, KKP17, LAYS19, OSSC21, SPMRS23, Sma01, SKL17, SCS06, YJLL20, YJLL21]. Mintzberg [Gar94, Per97b]. Misaligned [MT08]. Misalignments [GFvT22]. Mismanaging [Pal93a, Pal93b]. mismatching [WN03]. missing [Cai23, GF21]. Mission [Rea21, Chi91]. Mission-oriented [Rea21, Chi91]. Mitchell [Rea06]. MITI [NW03a]. mitigate [EKT^{+}22]. Mitigating [AGFE22, NBA22, RO22]. Mito [Joh92f]. Mitter [Hal97a]. MIVES [LGZ22]. mix [AT99, CLT06, GGGR22, Sna96, Sna98]. mixed [HRHW16, LB02, MAMRRR24]. mixed-methods [HRHW16]. MK18 [Sha95]. MK42 [BB95, Bes94a, Bes95c]. MNC [RS14]. MNE [CJL23, MPWG22]. MNEs [ABAHP^{+}23, FCCE21]. Mobile [Eng06, KZ03, Won08, AN16, AFL^{+}24, CEY11, CJH^{+}07, EF11, Fun06, GKK23, JvZ08, Kod03, KY06, LDF^{+}21, Li05, LM24, Wal07, ZWDZ21]. Mobilising [SM12a]. mobility [ALMW21, BV24, DL10, LKKH22, MPO23, Mig19]. mode [CY00, HT08, WSH23]. Model [CAP23, Pyk07, VNP^{+}22, AH84, AHAAEA03, AMS^{+}24, AMA21, Ano93o, AP08, ASOB23, Ayr88, BK04, BS06, BCG01, BBHA01, Cai23, CGM07, CK22, Cha08, CLW10, CW22, CCST06, CLIH05, DGP10, Dre02, EKK^{+}21, FSA23, FMOPSVA21, FGPV24, FF92, FFJ92, FGOX23, GEG12, GBL03, HLL93, HJM08, HKW^{+}13, HY04, HKY22, II92a, II92b, JLWS23, JLY23, KBO22, KI03, KK06, LDF^{+}21, LNB21, LLI06, LH19, MTY02, Mal02, MRGT11, MH98, MRDA24, MRC14, MII90, MP09b, NPS02, NJ04, NLC98, NSYL18, Öbe23, OKvB23, PWF07, PGLS21, PCWV16, PdW00, QWM92, RES^{+}24, RMGDLC08, Sér96, SD23, SL14a, SO06, Sto99, SM05, Sun96, Tan98b, Tan98a, TB24, TSPJ23, TCMS22, TE14, TPBM23, Vel15, WPF06a, Wer03, WC06b, YCL22, YPF03, YXH21, ZXX23, ZXDG19, ZMJM23, dOK12]. Modeling [MW09, MPW03, Tsa13b, CG11, ZZL22]. Modelling [Kha87, PPW06, RRSS07, SH06, XPC21, CLZ^{+}22, HPB03, LC22b, Gim06]. Models [RPASLM23, WZNG17, ACBO23, ABAHP^{+}23, BBT19, BNV05a, Bro07, CGS^{+}24, CP07, CNPV04, DLKM24, EFC23, GG05, KAV17, KLP24, LACH13, LDK22, Li21, Mil85, Mrk17, Mud90, NPV05, PG11, SVLYT24, SCH93a, SCH93b, VNJ18, WvdVM07, Wür17, dAASA21, vDZ21]. moderate [LW10]. moderated [LZLL21, NW23]. Moderating [KRM^{+}24, AAAF22, CL21, CJL22, CGLSNLDV15, FDVB10, FCM21, GR23, HC13, ILR^{+}12, KBRR24, Kra22a, LD05, LW11, LPN22, LZD24, OKP13, SBP21, SVÁG07, SCM24, TKYM24, VTRM12, WvdWG24, WHWS22, WW14, YXH21, ZLJ^{+}21, ZMJM23, CCGK23]. moderators [SW21]. modern [Ano91d, Coc12, Sou83a]. modernising [Rad97]. modernist [Des96]. Modes [CKM13, AHSS19, AHP16, BCC^{+}11, DS23, ES21, GBL03, Hag90, JBL23, LT10, SBP21, Ven96]. Modifications [DAR07, HM11]. modified [CYK19]. modular [CL05b, HKRR20, ZDWG14]. modularising [PÅ06]. Modularity [WCZZ14, WA04a]. modularization [WHWS22]. modularized [TKSM08]. module [WNGB03]. Monck [Hay88e]. monetary [PP04]. monetizing [RKFR24]. Money [vRE21, CS13, Lin08b]. monitoring [AKKS12]. monitors [Haa14]. Monozukuri [AS18]. Montalvo [Ban04]. Monteiro [Gra01]. MOOC [AAD14]. MOOCs [GHU21]. Moore [Bro07]. Morel [Leh07]. Morel-Guimaraes [Leh07]. Morgan [Bes02b, Her92a, Pap02a]. MoritaLou [Hob87]. Moroccan [Gos05]. morphological [Ano94-38, BCH94]. Morris [Buc91]. Morrison [Fra99]. mortem [KT04]. Morten [Ano93p]. Moser [Fra99]. most [BNV05b, Gal05]. MOT [Ano07l, CbLjJL12, Kod14, Lin07b, LE07]. mother [She05]. motivated [HEC22]. Motivating [SFVI21]. motivation [RV21, WH04b]. Motivations [PLG24, RWP19, SR22, SA22, TLL07]. Motives [BM16b, ABGS13, GKB12, VJ12, ZIB23, vdVdJVdR09]. Motor [SPS01, BCH98, FJ91, FFJ92]. Moutinho [Her92a]. move [Ser13]. movement [GZXX23]. moves [LTVA22]. Moving [HCMMH21, Hui17, ZZXW23, NH99, DDFS23]. Mowery [Ros90]. Mozota [Why05]. MRO [MTU23]. MSEs [EF11]. Mt [Hay88f]. Mugnozza [Mug89]. Multi [KK23, XX21, ANI^{+}22, BH16, BBC06, BT22, BCV21, CA02, CEY11, FSA23, Gee06, Hit89b, KK06, LGY05, Lee11, LP06, MMM16, NJ04, STA24, SFC14, SM12b, Tit90a, Tsa13b, LL04b]. multi-agent [LGY05]. multi-centred [Tit90a]. multi-country [Lee11]. multi-criteria [ANI^{+}22]. Multi-dimensional [KK23]. multi-disciplinary [LP06]. multi-generational [SFC14, Tsa13b]. multi-industry [CA02]. Multi-level [XX21, BH16, FSA23, Gee06, MMM16, STA24, SM12b]. multi-national [BT22]. multi-phase [CEY11]. multi-product [Hit89b, KK06]. multi-scalar [BCV21]. multi-site [NJ04]. multi-stakeholder [BBC06]. multidimensionality [PS04]. multidisciplinarity [Bat12]. multidisciplinary [WML^{+}23]. multihoming [BVT21]. multilateral [RFB24]. Multilevel [ÁRS17, CYM^{+}16]. Multimedia [SH08, Car98a, GC99, Kod00, Kod03, Kod05, RCC04]. Multinational [Koc09, AABY^{+}23, ASI99, AT93, ARvT04, ED87, KMID11, LG22, MM11, Pra86, Tag97, TRR^{+}24, WJK15, ZL24]. multinationals [HW18]. Multinetworked [Fra01b]. Multiple [Vos84, BBLW24, BSMC23, DG02, EKY24, FPP24, Hob97, KL08a, KA13, KFL11, LO10, LK20, LDMY22, Maj22, NI05, PGB^{+}23, RI04, SM05, UBCF19, CCP23, DMH22, SU11]. multiple-case [PGB^{+}23, UBCF19]. multiple-item [LO10]. multiplier [KK13]. Multiscale [KTSM10]. Mundt [Fra99]. Municipal [CA14]. Murmann [Nuv07]. Murray [Hay97a]. music [PS14, San07]. Mutually [VH06]. my [MWC22, SMW21]. mycorrhiza [SO02]. mystery [Ran24]. Myth [TW02]. myths [Gei97].

N [Ano91d, Ano94m, Hay93a, Her92a, Per04, Ros89c]. N1 [Per97a]. nacionales [Ano94-28]. Nadler [Mer01]. naissantes [Ano94v]. Name [Ano91d]. Names [Ano91c]. Nancy [Bes94b, RL06]. nano [BCO14, JFCFMC^{+}12, KS14, LS07, Won05b]. nano-enabled [JFCFMC^{+}12]. nano-patent [BCO14]. nano-revolution [Won05b]. nano-technology [LS07]. nano-textiles [KS14]. nanoclusters [BCM12]. nanoproducts [cLhWCH12]. nanotech [MLG12]. nanotechnologies [MW12, vRGJ^{+}12]. Nanotechnology [GFK12, MI07, Yeh13, AW12, BBA15, Bat12, Bha05, CLC13, GEG12, IM10, NP10, PWA11, WHB21]. nanotube [HS14]. Naperville [Bai84]. Narasimha [Bas04]. narrative [FP08]. narratives [Hja09]. Narula [Dut07]. NASA [MY02, Sch09, Wes93a, Wes93b, Min96]. nascent [Hu13, THC^{+}23]. Nathan [Ros90]. National [CG02, CbLjJL12, CR02, EEY22, Lan09, Nuv07, Par00, Sen96, Sha87, Tuz05, Ver01, vAdMSC94a, BT22, BWF18, Chi90, Chu02, GL22, Gru93, HN08, Hoe85, Hos00, Hos03, Lee04, dLP06, MAS13, PSU97, Pap95, RAC84, SGB12, SDN23, ZIAO^{+}21, vAdMSC94b, Ano94-28, Col81, LS07, Sou83b, YLB06]. nationalism [Mon01]. Nations [Hob87, CS02, Coc12, Har93a]. Natural [Jus24, And21, ABS24]. natural-resource-based [And21]. nature [CBGvdV22, Fab96, LJ01, Mol97b, Mol97a, Rob84, Rot93]. navigate [EKK^{+}21]. Navigating [Kra22b, LJAK21, LDMY22, EKY24]. Near [Lin06a]. necessary [BMRG00, Hos03, Fra02c]. need [Ell81, LB06, Lar09, NHT23, Ott04a, SVLYT24, SJSP23, SJP23, Ski84, Smi07, Hay85b]. needs [FHBDO16, HW97, IMIY87, KHL15, MII90, SSS17, TW02, WGW22, Whe88, Whi89a, Whi90a, Whi90b, Whi90c, Wür17, vWvRN17]. negative [ALM22, CZM23, KM11, QMP24]. neglected [ASA^{+}22]. negotiation [ZR93a, ZR93b]. negotiations [BA10]. Neil [Wil90]. Neither [dABKS14, Pfi99]. neo [TP24]. neo-configurational [TP24]. Netherlands [Ano97b, KDH02, RG10, SL98a, SL98b, WvdVM07, vdVMM09, vRGJ^{+}12, Hay87b]. Network [LFFT09, TA10, ZC22, BN06, BSWC23, CS05, CLT06, CZZM21, DR07, GZXX23, HKW^{+}22, Jai23, LHY04, MTB01, MCH00, MS19, Sér96, SP10, SHT23, SvG15, Sta98, SMW21, SJCT24, TTC06, WQZ21, XPC21, Egb15]. network-structure [Jai23]. Networked [BD05]. Networking [MSR10, Bøl12, CW06, FdJ09, Koc04, PS03, Pyk97, Pyk00, Rot91, VLIC12]. Networks [Giu09, MJ92, Ano94-28, Ano94-29, Ano94w, BEE24, BF99, BMSJ24, CGM07, CMtW10, CvZ05, CZZM21, Chi05, CM96, Die00, DNS^{+}03, EKK^{+}21, FORA17, GD08, Goe10, Hit06, HLL12, HLNP21, KTSM10, Ken13, KA13, LFU23, Liy95, LL05b, Maj22, NS07, PvdMB^{+}06, RZ21, RGMW18, RD05, RFB24, RK04, SSF12, SP98a, SP98b, Sha93a, Sha93b, SKTM08, SPP98, TKBC22, TKSM08, Töd94b, Töd94a, Vek99, WC06a, WH21, Zak23, ZXT10]. networks-Structural [GD08]. Netzwerke [Ano94-29]. neural [CM96, HKW^{+}22, WC06a]. neuro [ANI^{+}22]. neuro-fuzzy [ANI^{+}22]. newly [AT93, Ano93l, Chi90, LH14, PBC93, Sou01, dVV16]. newly-built [dVV16]. News [Ano88k, Ano92g, Coo16, HdWN21, Hay89c, Ano85d, Ano86e, Ano86f, Ano86g, Ano86h, Ano87f, Ano87g, Ano87h, Ano90n, Ano89z, Ano89y]. next [CDK^{+}23, FGPP22]. next-generation [CDK^{+}23]. nexus [LW21, ODM23]. Niagara [BS01]. niche [CT02, RG10]. niches [Boh18, CGM22]. Niemand [BB95]. Nigeria [AAA02, AOO02, AO03, AIOA00, AA05, ASI99, Aka98b, Aka98a, Ano94v, Ano94t, Ano94j, Ano94-31, DA01, FOL^{+}03, II98, II91, IAA94b, IAA94a, Ilo94, IOS00, IAO^{+}02, INO^{+}03, Nwo97, Nwo02, OTA01, OIOA01, Ola98a, Ola98b, OO02, OOJ^{+}00, OIN96, OOA04, SI98, SO02, SIOO01, Tai98, TOS^{+}00, UOIA00, llo91]. Nigerian [AA90, Egb15, Oye01]. night [Sch09]. Nijkamp [Hay90c]. nine [OFR24]. nine-year [OFR24]. nineties [Obe88]. Nintendo [SCM11]. Nirvana [Wal12]. No [Box01, BF23, Pot93]. Noel [Fra01a]. Nokia [RG06]. nomothetic [Ruo05]. Non [DVSL19, Hit89b, SCC17, ABKL08, xCZZ^{+}22, DVHL22, EFC23, FAH08, HL22, IIA97b, IIA97a, Jus24, KM06b, Lag17, LS23, LTT23, LL05b, MRGT11, OZ06, SBF15, Sgo95, SFVI21, HOGGP11]. non-academic [LL05b]. non-alcoholic [IIA97b, IIA97a]. non-assembled [Lag17, LS23]. non-collaborative [SBF15]. non-engineers [OZ06]. non-family [DVHL22]. non-financial [HL22]. Non-founder [SCC17]. non-human [Jus24]. Non-linear [DVSL19, xCZZ^{+}22, MRGT11]. non-manufacturing [Sgo95]. Non-mass [Hit89b]. non-pecuniary [SFVI21]. non-platformized [EFC23]. non-profit [FAH08]. non-standardized [LTT23]. non-technical [ABKL08]. nonlinear [Tsa05a]. nonprofits [SPÁG22]. nor [dABKS14, Pfi99]. Norbert [Har93b]. norm [Ano94-44, Ham94]. Normal [Kra22b]. Normann [Fra02a]. Normative [TPFD13]. North [Fra02c, Chu01]. Northampton [Dib01, Dib02a, Lal00, Que06, Won08, Wya07]. Northeast [Sub05]. Northern [MG98b, MG98a]. Northridge [Bes03a]. Norton [Kra15]. Norway [CK12b, CKM13]. Norwegian [Ano93p, Lev93]. Nostrand [Bul97, Joh92b]. not-invented-here [AGFE22]. not-sold-here [AGFE22]. Note [Ano19h, Bai18, Ano82b, Ano84d, Ano03p, Ano04o, Ano04p, BE09, CR19, PP02, TKS18, Ano06m]. Notes [Smi85, Lin10c, Rob84, O’H99]. notice [CDD^{+}14]. novel [Dow93a, Dow93b, GFvT22]. novelty [AL05, ALBO08, BF23, FdJ09, GD08, HT12, Men11, NS07, RDS17, TDC11, WCW^{+}14, vdEJG08]. NPD [ET98, ILR^{+}12, LB06, MPM09, QvdR21, TAC96, WLY14, dVdWNF^{+}10]. NSF [WHB21]. NTBFs [OÁVH14]. nuclear [Big13, Hay88g, Lee04, SH99, Tho88b, Tho88c, Hay92c]. Numerical [KZ18]. nurse [HPMI^{+}23]. nursing [HPMI^{+}23]. Nurturing [NF24, TNY^{+}22, TPB12]. NY [Giu09, Kra15].

O [Ano94-31, Hay89d]. Oakey [Gar85, Gar96, Pap03]. Oaks [Box01, Hay88d, Kuc07]. object [QMP24]. objective [Sou01]. objectives [Sna98]. objects [CPT23]. observaciones [Ano94i]. observation [AKKS12, LS16, BBLW24]. observations [Ano94u, Ano94g, Ano94i, OW93a, OW93b, Oak03, Tit94b]. observed [Ric85]. Observing [Haa14]. Obstacles [Fra00b, KA06, vGS09]. O’Connor [Bes01a]. October [But96, Hay97c]. OECD [DGS23, KL04, LPMV22, Rad01, Sir91]. off [ADD19, CFC12, DB05, DLL06, KL08b, LKCK22, MW82, NPS02, ODM23, Rie89, Rog86, SvG15, SYCH23, VP14]. Offensive [WZY23]. offering [Ski84]. offers [BSWC23]. office [BBT19, Bes02b, NC23]. offices [BB22, RGJ13, Wec15, ZT21]. offs [Ano94-37, BKM10, BHPW00, CRKA98, Fon05, GvBR10, MSvdK^{+}21, OÁVH14, PG94, PS03, SS21, Sch86, SOKE15, SJ16, Ste14a, TBSM21, TKT13, WD06, vGS09]. offsets [AR99]. O’Hare [Bro90]. Ohio [EW97]. Oil [CS08a, FOL^{+}03, Gos05, HK02, OOJ^{+}00, RVW17, RFMC21, dSSD09, TOS^{+}00]. Oilseed [II91]. OKID [EEY22]. Olby [Ano91d]. old [CS01a, Kod03, Mus09, RWP19]. old-economy [Kod03]. Oldenbourg [Ros89a]. older [CC14, KHL15]. Ole [Gra01]. Oliver [Hop04]. Olympic [FLKF^{+}06]. omit [ABGS13]. omitted [Ano12a]. Ondo [OOJ^{+}00, TOS^{+}00]. One [BBT19, CL97b, CL97a, DMH22, HW17]. ones [RWP19]. Onida [Oni89]. Online [AAD14, CP07, GHU21, Jac22, ASGC09, BEE24, BNV05b, CS21, Cha08, Kan22, MGdG11, MW14, New04, OGOF22, PvdMB^{+}06, SDT^{+}22, XWZ23]. only [Lin10b, Wya07]. Open [AAD14, BDB21, CGLSNLDV15, FLYTB24, FL15, GHU21, Hay87d, HT12, Hui11, Jac22, Lin18, MAMD22, OVD21, PC18, Sha95, THBW22, THC^{+}23, VSE24, vdVdJVdR09, AM23, AABY^{+}23, AB23, Bad11, BT22, BAR16, dABKS14, CA11, CBGvdV22, CJL23, Cre04, CMS23, DDBM23, DCD22, EB23, FN21, Far23, FSA23, GL10, HMM15, HCT11, HC13, HDE21, IBPGML23, JJS^{+}18, Lic10, Lic11, Lin10a, LC22b, MDT^{+}23, MMC^{+}23, Men11, MM11, MSR10, NHT23, NSB23, OGRA23, PUFC23, PEER23, RVW17, RH21, SPMRS23, SP22b, SR22, SCK10, SCK11, SFVI21, VNJ18, YA11, ZXDG19, dB10, vdVdM11, vH10, vK11, BCC^{+}11, CCF11, LTVA22]. open-source [RH21]. Open2 [CHP13]. Open2-Innova8ion [CHP13]. Opening [BPU21, LVVH23]. openness [KM11, MSvdK^{+}21, TPV16, THC^{+}23, WCW^{+}14, ZLH17]. operating [LH11, OE87]. operation [Hag90, LP06, Men11]. operational [KK06, SM09]. operationalising [SPP06]. operationalizing [JBZ23]. Operations [Gup96, GCS01, BGR91, CG04b, FC06, Gup00, GG03, GK06, KZ03, PSdLC02, SDFN01, Szw07]. operative [Cul00]. operator [KY06]. operators [BLQS00b, Hos01]. opetition [Men11]. opetitive [CA01]. opinion [YJLL20, YJLL21]. opinions [OSSC21]. Opportunistic [CG04b]. Opportunities [BGF23, Dre04, HGR15, Lop06, Box01, BVHJ24, CS02, Car05, FL15, KBO22, KV24, KHL15, LYP09, MAMRRR24, MZL15, Mos16, Nar04, OZ06, Ove15, SC17, SKL17, Ver01, WKF24, YSKL21, ABSZ23, Obe88, Fra00b]. Opportunity [LC09, Par05, CK04, GH15, JB19, KK23, LJAK21, LFU23, MSD15, NQ05, NT16, RZ21, SB02, SMGPP24, THBW22, SAY09]. opposite [YXH21]. optic [Web96]. optical [FG06, WMGB01]. Optimal [MOWGB02, LVR23, SH95]. optimally [vWGdG21]. optimised [Sto03]. optimization [Coc12, SP10]. optimizer [LV23]. optimizing [DB97a, DB97b]. option [CEY11, VNP^{+}24, vZS14]. Options [Cam93, dMvAdMSPB93, CLIH05, Par03b, VTRM12, WKN03, WO08, Pot93]. opto [CTYH02]. opto-electronics [CTYH02]. or/and [CZZM21]. oracle [Mil85]. orbit [WKN03]. Orchestrating [HLNP21, KA13]. orchestration [ABAHP^{+}23, DVSL19, Far23, SFdO^{+}23]. order [Car98a, CGW05]. Ordinal [Rue88]. organic [Whi90b]. Organisation [Hay86a, Hos01, Tra00, AGS01, Gre06, Wan04]. Organisational [BCC^{+}11, BM07, Hal97b, UOIA00, ASA^{+}22, BF06, CHP13, HPB03, Hay92b, IBL04, MD03, Mil88, RRB^{+}22, SS23, SM12b]. organisations [AL11, Bes97, CS98, GARM24, II97, Kin02, LCLC03, SSCH23, Tan98a, ZC22, ZS04a, ZS04b]. organise [SCK10, SCK11]. Organization [FLKF^{+}06, Ano93b, Ano94-33, BN88, CCGK23, CA07, CHT22, CW93, CW94, Des96, FRW06, FJ91, Hol99, KM06b, KL05, LSB07, hLFL08, Llo07, Obe86, Ott97, Pet98, Pur91, Tom90, WHWS22, Fra00a, Jel07, Wil90, Ano98l]. Organizational [ABKL08, BO86a, CD92a, CD92b, Car98b, Dib02b, Dre04, GGS08, Gup88, Hag90, Hay88c, Hay98, LS04, SM01, Tsu02, VvWM20, Zak23, AC08, AAK^{+}24, Bar95, CCP23, CDED21, Car99a, CM89, CCC^{+}12, CNPV04, FGPP22, FHA22, GMLMVJ07, GKSW19, GC93a, GC93b, HBvdB09, HC10, KS04b, LB06, LXODT15, LLS00, LLR95, hLFL08, Mac86b, Mac88, MRGTDG21, MMM05, Mud95, NG09, dAOCM02, OBMOM24, RGB23, SVÁG07, Sen96, SCM24, Sta98, Tan98c, Tom90, VMWV16, WZY23, Whi89a, Whi90a, Whi90b, Whi90c, Woo85, YGM99, YCL22, BK89, HB89, VMWV16]. Organizations [Mor04, AM23, ACBO23, BS24, CE19, Cur89, DL10, FHA22, HCMMH21, Huc85, Joh08, Kar86, KRK^{+}22, KMK03, LACH13, MAU07, NG09, Nar89, OIOA01, OM89, Ott98a, PLG24, SCH19, SK11, SD89a, Tan98b, TFS^{+}22, TKT13, ZDWG14, LYST22]. Organizing [CvdEG13, HAH84, LR09, Sou83a, AW12, ALCM14, Mon01, RVW17, RK04, Sun96, Ros89b]. orientation [AAAA23, ABG23, BC92, CBF24, CCV11, CK12a, CK12b, FCM21, FMOPSVA21, JLL10, LSB07, LL05a, OWCvdB15, Ran24, RPL10, SCS22, SP22b, TRW23, TMG11, ZYSL21]. orientation-new [CK12a]. orientations [SRHL12, ZHZL21]. oriented [CMB23, Chi91, Hay85b, HWC06, Hol85, Hol87, Hol88a, Hol94a, Hol94b, KB09, KLH05, Pet98, Rea21, RRB^{+}22, SVÁG07, Sto03, WLH10, WH04a, WBM22, Pap02b]. origin [RGFLRP21]. origins [EW97, Lim06]. orthopaedic [ALFM23]. OSL [AFL^{+}24]. Osun [OTA01, OOJ^{+}00, TOS^{+}00]. other [Cur89, ES05, Lin10c, Nar89, OÁVH14, SD89a]. Ottawa [Ano07m, Smi85]. our [Lin08a, Lin11c, Lin11b, Pot93]. out-licensing [HMM15, KLRR22]. Outbound [HMM15]. outbreak [TFS^{+}22]. outcome [CLZ^{+}22, VNJ18]. outcomes [ABLK07, BPU21, CG11, EWJ09, ILR^{+}12, KZ18, Lin20, Pro21, SFdO^{+}23, SSW22, Sto21, WQZ21, XHM^{+}16, vRGJ^{+}12]. outline [CEH03]. outlook [CM95, Rad01, SMC14]. output [Ano94p, Han94, MP09b, Ram23, SPFG12, SH22]. outputs [AG24, ZT21]. outs [Pow10]. outset [OÁVH14]. Outsmart [Bes01a]. Outsourcing [KS04a, NC23, Pia02, HT06, LL04a, LTHS12, TS13, UR18]. outward [NB11]. overcome [AGFdMCM24, DFFM22, NVMD22, PSS23, TWGB02a]. Overcoming [HKR^{+}22, JCT23, RGMW18, WC04, ZSX24]. overconfidence [WSN^{+}18]. overlooked [AG24]. overruns [CG05]. overstepping [PCQ24]. overview [Ano94-43, BO86b, DMV^{+}21, DA01, KUH21, Mia94, MLF16, ON88, Per04, VM11, Whi89a, dRP95]. Owen [Bes02b, Hay90b]. own [HLL12, TPH15, Wür17]. owned [EW97, GWZ^{+}23, PCG22, Tõn15]. Ownership [ZLC18, AK00, Bar15, TPH15, Wan23a, YHcHW17, YHcHW21, FMF21]. Oxford [Bes02a, Bes02c, Bes03b, Bro90, But96, Cou90, Cou91, Dib02b, Dod88, Edw04, Fra02b, Gra01, Her91, Hob03, Jel07, Leh07, Pap03, Pap04, Pot93, Poy97, Sha91a, Tan08, Ver90a, Ver91a, Wan04, dC07, Dib02b, Hed89, But96]. Oyelaran [Coo07]. Oyelaran-Oyeyinka [Coo07]. Oyeyinka [Coo07].

P [Ano93k, Ano94-37, Bes94c, Box01, Dib00, Fra00b, Hay88e, Her91, Joh92b, Nuv07, Per04, Sha91a, Rad01]. P019 [Gri98]. P2P [San07]. PA [Buc91]. pacemaker [Goe10]. packaging [AO10, BRA^{+}23]. packing [SM05]. packing-multiple-boxes [SM05]. Page [Gil96, Hay89a, Joh92f, Joh92e, New04]. Pages [Kaz03, Ano91d, Ano91c, Arr04, Bai84, Bas04, Bes88, Bes93, Bes02b, Bes02c, Cou90, Cre04, Dod88, Dre04, Edw04, For90, Fra02b, Fra02d, Gar85, Hay85a, Hay85b, Hay85c, Hay86c, Hay86d, Hay86e, Hay86a, Hay86f, Hay87a, Hay87d, Hay87b, Hay87c, Hay88a, Hay88b, Hay88c, Hay88d, Hay88g, Hay88f, Hay88e, Hay93a, Hay97d, Her04, Hob03, Hop04, Lae04, Lin90, Lop06, Mil88, O’H99, OH04, Ram86, Ros89c, Ros89b, Rot87a, Rot88, Sco89, Tat04, Ver90a, Ver90b, Bro90, Buc91, Cou91, Dav92, Dod87a, For91, Fra03b, Hay83, Hay84, Hay85d, Hay90a, Hay90b, Hay90c, Hay90d, Hay91a, Hay92b, Hay92a, Hay92c, Hay92d, Hay92e, Her91, Her92a, Her92b, Her92c, Joh92a, Joh92f, Joh92c, Joh92d, Joh92e, Joh92g, Lin91, Mor91, Mur91, Ros89a, Ros90, Sha91a, Sha91b, Ver91a, Ver91b]. pages [Why02, Wil90]. pains [BEH99]. Pakistan [EKT^{+}22]. Palgrave [Coo07, Giu09, Koc09]. palm [Ano94t, Ano94j, IAA94b, IAA94a, OOJ^{+}00, TOS^{+}00]. Palmenerzeugnissen [Ano94j]. palmera [Ano94t]. palmier [Ano94v]. Pamela [Har93b]. Pan [Joh92d]. pandemic [BKA^{+}23, BEE24, EHSB24, FGB^{+}24, MWJ23]. Pandey [Ano94f]. panel [KZ18, QGBV04, RAG11]. Panoramic [vW84]. Panscheel [Ada07]. Pao [Ano94l]. Pao-Long [Ano94l]. Paolo [Arc03]. paper [Ano82c, Fra03a, Hay97c, JRS06, Sch92, SL14b]. Paperback [Ada07, Dut07, Fra03b, Kap03, Kuc07, Rus94, Szw07, BB95, Bes95b, Bes97, Bul97, Dod87b, Hal97a, Hay93a, Hay93b, Hay93c, Hay01a, Hay01c, Hol93, Lim06, Per96, Per97b, Pre01, Sha95, Wya00, Wya04, dC07, Col97, Pot93]. Papers [Lop06, Ano84a, Ano85e, Ano12b, Lin12c, Lin12d, Lin16, SL14b]. paradigm [CFP^{+}23, Chi91, CCF11, Fox86a, FG06, How24, KM97, LWL^{+}22, NWGB09, Szá01]. paradise [BSSF24]. paradox [Leo21, MLC24]. paradoxes [BVHJ24]. Paradoxical [Shu22, PCQ24, SOA^{+}24, XLG^{+}23]. parallel [SD95]. parametric [KW05]. Paré [Wya04]. parent [TKT13]. Paris [Rad01]. Park [BB95, Bes94a, Bes95c, Hay93c, LY00, BMV02, BDNP06, CL05a, DVMS16, Pál04, Phi99, QWM92, Guy96a, LS05, Lan09, MM08, Phi99]. Parker [Ano81b, Hay85d, Hay86d, Hay89a]. Parkinson [RNJ^{+}23]. Parks [Fre86, Ano94-42, BMV02, DSVM05, Fel94, HHV05, Hay87b, Hay88e, Hay93b, Kih03, LS05, LB02, LK20, RH10, Rot87a, Ved97, DLL06, LL03, LL05b, Hay87c]. part [SSP04, Ish89a, Ish89b, Tho88b, Tho88c, Whi89a, Whi90a, Whi90b, Whi90c, ZS04a]. participant [HL22]. participants [LS16]. Participation [Ott03b, FM07, PW99, RBBB98, SRLR23, Pap06]. participative [AH85]. particularly [WVG^{+}98, WGV^{+}98]. Partner [JH18, FGPP22, vRGJ^{+}12]. Partnering [BCH98, ADNP01, MR05, VK03]. partners [AdA08, EBG24, GLYP13, JH18]. partnership [AA05, CA11, LD95, MG98b, MG98a, RA86, CG02, Bes93]. Partnerships [SPÁG22, AC04, BT22, Bar15, BCL22, Bes94b, CAI00, DDN24, HS91, Jon05, NH99, SPO08, SPP98, VS06, WM88]. parts [LKPP07, Soh97a]. party [Bie91, SS03, Vel15]. Parvatiyar [Ano94-37]. Pasqualetti [Hay92c]. Passenger [SPS01]. passive [Kap99]. past [BTP^{+}23, CP11, FGTB23, Rot92a, Rot92b, Gol86]. patch [PDHW23]. Patent [FNJV14, GKB12, ML08a, WTGB01, AM01, Ano94p, BCO14, BLLEM23, Bro92, BEH99, Cav16, CN05, CK18, FR15, Goe10, Han94, JY15, LYP09, LJAK21, LYY^{+}22, Llo07, MFH16, NMN11, NS90, PK23, Pet11, SAWT06, SS21, SCH23, Tsu02, VJ12, WZY23, Yam21, lS06, vBDK21, HJSYJ12].

Patent-based [GKB12]. patent-product [LJAK21]. patent-selling [SCH23]. patented [WJK15]. Patenting [Ern95, ED87, Arc88, Bal10, EL93, KL24, LXODT15, Mar01, OM00, PSU97, Pil04]. Patents [Thu05, ZSMB16, AP96, BBA15, Bar15, FMF21, FNJV14, GLYP13, GUMZI14, HS19, JP96, KS07, MFH16, MO03, McQ05, Mig19, Neu12, SLJ13, SKL17, Wal97, Wan24, WWH15, YSKL21, vdWFvdB23]. path [CG11, MMR24, VNJ18]. Pathak [Kaz03]. paths [GSP^{+}22]. pathways [BM23, HMM14]. patient [BGF23, CJH^{+}07]. patients [RNJ^{+}23, BKA^{+}23]. Patrick [Tan08]. Pattern [Sub05, Ano93l, JL24, Pfi99, PBC93, SI98]. Patterns [BC82, Fre05, GG11, KL87, tBL86, CS05, CBK03, ED87, JY15, Kim87, Lan09, LL05b, McG98, Pal00, RAG11, RG06, SH10a, Sun00, vdVMM09, Fra99]. patterns-academic [LL05b]. Paul [Hal97a, Per97a, Sap10, Tra00]. Paulinyi [Ano93n]. Pavon [Pav89]. pay [Ano94-37, LKCK22, ODM23, PG94, SLJ13]. pay-offs [Ano94-37, PG94]. payment [SF98]. payments [CEY11, Sir91, Won08]. pays [ADD19]. Pearson [Ano93f, Why02]. Peat [Hay88e]. pecuniary [SFVI21]. Peer [BEE24]. penetration [LLR95, TWGB02a]. Pennies [DMH22]. Pennsylvania [Buc91, RA86]. people [Ano94i, BV24, KHL15, NK22, RR11a, Tit94a, Tit94b, Bes95a]. perceive [BMW14]. Perceived [GM00, Car01, HF18, JMPR04, OGOF22, YJLL20, YJLL21, vWvRN17]. perception [LA15, PSdLC02, SL14a, WC85, ZHY22]. perceptions [AL11, HBM14, HB03, JH18, KWvdS21, KJ03, FGTB23]. perceptual [Sou01]. Peregrinus [Rus94]. Perez [Kap03]. perfil [Ano94i]. Performance [BF21, CK22, CLZ18, CLZ21, GWM05, NLC98, OO02, Whe88, ALvKL06, AAGL23, AAAA23, AKBA07, AC08, AS22, Ano94-43, Ano16h, AP08, AHP16, ANI^{+}22, AB23, BC12, Bar89b, BS97, BBS22, BCL22, BRA^{+}23, BNS16, BM05, BF86, BOBT22, CHP13, Cal24, CKT04, Can16, Car00, CA02, Car01, CS87, CCC^{+}12, CSL23, CCGK23, CHL21, CCV11, CYHC19, CK12a, Cho96, CDMV24, Coz06, CGLSNLDV15, DTD13, DZX22, DCD22, Der14, DG14, DVHL22, Egb15, Ern95, FDVB10, FCM21, FMOPSVA21, FCS21, FJ12, GMLMVJ07, GST08, Gar05, Gol90, GdWNPC03, Gra03, GZXX23, GM03, HBS02, HG92, HL22, Hua09, HLL12, Huc85, HC13, INO^{+}03, JCT23, JL96, JKF19, JRS06, JJS^{+}18, KBSW08, KM01, KET22, Klo87, KM11, KLV06, Kuh88, KS10b, KRM^{+}24, LNB21, LRH22, Lee23, LL98, LQ09]. performance [LC09, LT10, LFFT09, LW10, LWC^{+}12, LH21, LCT05, LDMS23, cLhWCH12, LC22b, LZD24, Mac97b, MFH16, MRDA24, Mia94, MCMA09, MMM05, MPM09, NG09, OC91, ODM23, OOKBD14, OBMOM24, PNC15, PS21, PC05, PRID23, PS04, PH08, Puj06, Ran24, RGB23, RPL10, ROMH15, Rue86, Rue88, SRT18, SCS22, SBF15, SGB12, SAY09, SBP21, Sar22, SSM09, SCY17, SAM^{+}23, SY05, SHT23, SCM24, SGWL16, SH10b, SMW21, SH02, SD04, TKYM24, Tei22, TWBC21, TDC11, VP14, WG23, WTJ23, WC06a, WC07, WCT10, WLH10, WCW^{+}14, WLY14, WH21, Wer03, WvT23, WH96, XW21, XLC23, YGM99, YK08, YWC09, YHcHW17, YHcHW21, YMK^{+}21, ZXT10, ZQW^{+}19, ZC22, ZYSL21, ZZL^{+}23, ZMH22, ZCR22, ZHZL21, dVdWNF^{+}10, dVV16, dlHBR22, Lim06]. performance-oriented [WLH10]. performances [BDNP06, OSG04, ZWDZ21]. Pergamon [Pap03]. period [Bar97, INO^{+}03]. periods [ABAHP^{+}23]. peripheral [BG10, ST95]. perish [Eek84]. Perpignan [MP09a]. Perry [Ano94m]. persistent [ØD22]. personal [EM02, FB91, GMLMVJ07, SAY09, dHOvdKS16]. personalization [LWC09]. personas [Ano94i]. personnel [TS13]. perspective [Bad11, BSN21, BJB^{+}24, BK01, BF21, BS99, BST03, BRH^{+}97, Bro89, CP23, CL98, CCMW23, CSL06, CS12, DGP10, EB23, EG05, GKSW19, GC93a, GC93b, How24, JH09, Kee14, KAV17, KFEC15, KP23, KB06, KY06, LKKH22, LT21, LL04b, LTH06, LL08a, LWC^{+}12, LH14, LDMS23, LL08b, LG22, LH11, McA04, MMM16, Mol97b, Mol97a, MGAP23, NW23, NLLC04, Orr97, Par05, PH08, RNFG24, RPGP23, RY14, STA24, Sik96, SOA^{+}24, SS03, Sub05, SDN23, TP24, UEGE^{+}14, WLY14, WZY23, Whi90a, WO08, XL23, XPC21, YVG24, YCL22, YA11, YMK^{+}21, ZWZ^{+}23, ZR93a, ZR93b, ZXDG19, Pra10, Fra03a]. Perspectives [Bes03a, Gol86, Hay88f, BM16a, BSMC23, HBG12, Hui11, IKL01, LG23, MT08, PV01, Sik97b, SLJ13, SL16, TB84, Whi89b, JD09, Sik97a, Her92c]. pertinent [MDT^{+}23]. Peru [PGKF19]. Peruvian [SLSCS24]. Peter [BB95, Cre04, Fit00, Har97c, Hay86e, Hay90c, Hay93d, Per97a, Rad97, Rus94, Sap05, Wil90, SD89b]. Peterlongo [Hay85b]. Peters [Bes01a]. Petersburg [Kih03]. Pettigrew [Dod88]. Pfeffer [Bes95a]. Ph.D. [MR19, New13, SBF15]. Phaal [Bes02d]. pharma [RvdSdE22]. Pharmaceutical [Hop04, AW11, Bac98, BCC^{+}11, Bou15, CYK19, Gar05, HMM15, JK03, Jon96, KLRR22, Lov85, LTHS12, OBvE97, Pia02, RRB^{+}22, Sta98, TWGB02b, TW02, TWGB02a, Wan24, WTN^{+}02]. pharmaceuticals [Gra82]. phase [CEY11, FGTB23]. phase-specific [CEY11]. phenomenon [Lid90, Seg86]. Phil [Wya00]. Philip [Del97, Par03a]. Philippe [Hay91a, Ver90a, Ver91a]. Philips [OKvB23]. Phillips [Pap02b]. phone [Fun06, HH08, JvZ08, Wal07, ZWDZ21]. phone-based [Fun06]. phones [AN16]. photographs [Hay97b]. photonics [Yeh13]. photovoltaic [Kap99, WWM00]. photovoltaics [KGS^{+}24]. Physica [Lim06]. Physica-Verlag [Lim06]. physical [ETM96, GFvT22]. picture [KTE04, KJ97]. Pietrobelli [Arc03, BH03]. Piller [Rea06]. Pilot [Ott97]. piloting [TSPJ23]. Pimenta [Ano93l]. Pimenta-Bueno [Ano93l]. Pinter [Gar85, Har93a, Hay83, Ram86]. pioneering [LV09]. piracy [CP07]. Pitt [Joh92g]. PKI [CT06]. place [AMS^{+}24, JL13, LTVA22, Mud90]. place-based [AMS^{+}24]. placement [LCWC09]. plan [ETA02, Hos03, SPS01]. Planet [Kha98d, Kha98c]. Planetary [Ott98a]. planners [AOO02]. Planning [Ano94y, BRP94, MK00, MS98, Ros89a, Ano93k, BN88, CM96, ETA02, Gup89, GG93, GK06, HS21a, Hos00, Hos01, Kar93a, Kar93b, LKPP07, LP06, OG05, Ott97, PR97, Rya96, Sza90, YGW06, YGW08, Gar94, Hay87d]. Plant [Gar05, HS02]. plants [BS87, Bro01]. Planung [Ros89a]. plastics [BK99, Mus09]. Platform [FGOX23, TCMS22, dHOvdKS16, BBLW24, BVT21, CT12, CBGvdV22, CUF13, EFC23, FGB^{+}24, JSP22, KBSB22, Lag17, MSvdK^{+}21, Ott03a, RNJ^{+}23, ZXX23, ZWZ^{+}23]. Platform-based [FGOX23, TCMS22, Lag17, ZWZ^{+}23]. platform-to-platform [MSvdK^{+}21]. platformized [EFC23]. platforms [BGF23, JWY23, JSP22, Kan22, MR00, NK22, NPG17, Oll08, PGB^{+}23, RES^{+}24]. Play [Jel07, SW04, WZNC15]. player [PS14]. players [Kum05, Kum06, WTN98b, WTN98a]. Playing [BHS05, LP03]. Plaza [AM97]. PO19 [Per95, Pov97]. point [SSP04]. pointe [Ano94w]. pointers [SBOB23]. points [Ano93b, CW93]. Polaris [Hol88a]. Policies [Gui11, HBG12, Nar89, Ano94-43, Arn87, Bou15, BM05, CJ98, GWZ^{+}23, Kap11, Kra22a, Lar09, Lee00, LL93a, LL93b, MK87, Mia94, NG97, SGB12, Sha87, SH95, Web99, Won16, dBJ03, OH04]. Policy [Ano11k, GvBR10, GGGR22, HST05, LG14, Ros89b, Rya90, WKN03, AB99, AMS^{+}24, Ano94-38, BMW14, BCO14, Ban04, BCH94, BM14, CS84, Cha88, Chi91, CM95, CL97b, CL97a, CU13, CA14, CS89, Dib02b, DR88, DS13, Fab96, GW87, Gra82, Gra04, HY99, Hay83, Hay92e, Hay01a, HN08, Hoe85, HST21, HST07, II98, LHY04, LL03, LBG14, MJ92, MT08, MS19, Ola98a, Ola98b, Oye01, Par00, Pen08, Rad99, Rea21, RR88, Roe87, RW86, RD92, SIOO01, SBMP13, SJCT24, Szá96, Thu18, Tom03, Tõn15, TS05, Vav06, Vek99, WS14, Whi90c, WLG05, Zeg87, SL16, Par07]. Policy-driven [LG14]. policymakers [HOGGP11]. Polish [Bar97]. Political [CU13, Coc12, CYHC19, DVSL19, Koc04, Ram81, Wal85, WSN^{+}18, ZQW^{+}19]. politics [LPXO14, Smi07]. Polity [Dut07]. pollution [Sto99]. Polytechnic [AA92a]. Polytechnics [AA92b, OIOA01]. poor [PW99, RM14, SO02]. poor/marginalised [PW99]. popular [AB89]. population [FFJ92, XPC21]. portable [Mur91, PS14]. Portals [FRA05]. Porter [Hay88e]. Portfolio [Mik01, Egb15, dVGCOF^{+}23, Jol03, LQ09, MZG17, OKP13, Pil04, Van10, VGAL24]. portfolios [AAT14, BEH99, EWJ09, HST07, SR23]. Portland [Kuc06]. portrayal [FD21]. Portugal [FC95, Lar09, MCD06, RH10]. Portuguese [AS08]. posibles [Ano94t]. position [CT02, MTB01]. positioning [CX24, IMBH98]. positions [CZZM21, ZC22]. Positivity [Fel23]. possible [Ano94t]. Post [ACV08, LQ09, O’H99, AHP16, BFM86, HO11, HM11, INO^{+}03, KM06a, KT04, RPGP23, SH22, YVG24]. post-1990 [KM06a]. post-acquisition [SH22]. post-adoptive [HM11]. post-COVID-19 [YVG24]. post-growth [RPGP23]. post-introduction [BFM86]. Post-it(R) [O’H99]. Post-merger [LQ09]. post-privatisation [INO^{+}03]. Post-project [ACV08]. post-Soviet [AHP16]. post-use [HO11]. Potential [Cha94, KH06, Ano93m, Ano94j, Bre23, EH14, II91, IAA94b, IAA94a, KBS01, MW14, QvdR21, RV21, SAWT06, SS21, SD89a, Tai98, Tei93, THBW22, Her91]. potentials [Gol86]. potentielle [Ano94v, Ano94j]. pottery [GSFS21]. pound [Fel22]. powder [AT99]. powders [Che12]. power [BA10, HO14, HF18, Hay88g, KK07, MTB01, SH99, Tho88b, Tho88c, WWM00]. powered [AHS^{+}23, MWC22]. powerhouse [Won16]. pp [Ada07, Ano97b, Aug09, BB95, Bes94a, Bes94d, Bes95a, Bes95c, Bes97, Box01, BH03, Col97, Col08, Coo07, De 04, Del97, Dod87b, Dut07, Fra02c, Gar87, Gar94, Gar95, Gar96, Gil96, Giu09, Gri98, Hal97a, Hay93b, Hay93c, Hay95, Hay96a, Hay97b, Hay97a, Hay01a, Hay01b, Hay01c, Kap03, Koc09, Kra15, Kuc07, Lae03, Lim06, Lop06, Min94, Mor04, Oak97, Pap02b, Par03a, Per04, Per95, Per96, Per97b, Per97a, Pot93, Pov97, Poy97, Pra10, Pre01, Ric95, Sap10, Sha95, Szw07, Tan08, Won08, Wya00, Wya04, dC07, Bes94b, Bes94c, Coo97, Gil97, Hob97]. pp. [Ano81a, Ano81b, Bes96, Dib04, For85, Hay89a, Hay89d, Hay97c, Min96]. pp.176 [Mer02]. pp.197 [Bon02]. PPPs [TKBC22]. practical [Bes02b, Box01, Gri98, IPP14, Mer94a, PPP98a, PPP98b, Rus94]. Practice [Fit00, Hay86e, BN08, Bro07, CL97b, CL97a, CWB98b, CWB98a, CO04, CL88, DDBM23, DV23, HPMI^{+}23, Hal96, Hol87, Hos00, IPP13, LBA14, RRB^{+}22, SPP06, TSH02, TPB12, VM97, VM92, Bea05, Hay97a]. practice-oriented [RRB^{+}22]. practices [AMV07, BYCS00, BDB21, Bar15, Bar14, BS06, BS05, Bøl12, BM05, dABKS14, CJ98, CvZ05, CTC02, GBL03, Han06, Hen01, HKR^{+}22, KRM^{+}24, LC22b, MM02, Men11, MGAP23, OOKBD14, PS00a, PS00b, PS04, RMGBA08, SP03, SPP98, SGFS99, TM06, WZNC15, vAdMSC94a, vAdMSC94b, RGJ13]. Practicing [Fra02d]. practise [Pap02a]. practitioner [Whi89b]. practitioners [FGTB23]. Praeger [Ros89c]. pragmatic [CF01a, ML05]. Prashantham [Pra10]. Pratt [Sap10]. praxis [KK23]. pre [INO^{+}03, Mer22, TBSM21, VLVH23, YVG24, ZWDZ21]. pre- [INO^{+}03, YVG24]. pre-acceleration [Mer22]. pre-determinants [VLVH23]. pre-entry [ZWDZ21]. pre-formation [TBSM21]. Precarn [Joh08]. predevelopment [MK96]. Predicting [Sma01]. prediction [KKL19, LFU23]. Predictors [CC14, SM01]. Preface [Ano89-27, BO86c, BFO88b, Rot87b]. prefecture [TKSM08]. prehydrolysate [Ano94a, RGSK94]. Preliminary [WTAW97, Ano93l, HAH84, IKL01, PBC93]. Prencipe [Edw04]. Prentice [Bes95b, Dod87b, Gar94, Hay01b, Her92b, Min94, Per96, Per97b]. Prentice-Hall [Dod87b, Min94]. Preparing [EH14, Hay83, LP06, KS10a]. prerequisite [Hos03]. Prerequisites [KBC22]. Prescott [Koc09]. presence [Ell81]. Present [Hit92b, Hit02, Hit03, Hit05]. Press [Aug06a, Bes95a, Bes96, Bes97, Bes01a, Bes01b, Bes02a, Bes03b, Bon02, Buc91, But96, Cre04, Dav92, Dib02b, Dre04, Dut07, Edw04, Fra00a, Fra02b, Fra02c, Gra01, Hal97a, Hay84, Hay88g, Hay97a, Hay01a, Hay01c, Hob87, Hob03, Hol93, Jel07, Joh92c, Kap01, Kuc06, Lae04, Min96, Nuv07, Oak97, OH04, Pap04, Par03a, Par06, Poy97, RL06, Ros90, Saa98, SW04, Sha95, Sod08, Tan08, Wan04, Why05, dC07, Sha91b]. pressure [HS21a, KRK^{+}22, Sch09, XLG^{+}23]. pressure-relieving [Sch09]. pressures [RVW17, UF22]. Prestige [SCH23]. preventative [KS14]. Preventing [CMS23]. Prevention [KB06]. previously [QvdR21]. Price [Aug09, Bes02d, Fra02c, Fra02a, Hay01a, Hay01c, Mer01, Mer02, Bai84, Box01, CM94, Gar85, Giu09, Lal00, NMN11, Pap02b, Pra10, SBF15, Tsa13b, WWM00, Web96]. prices [SCH23, SSGN23, WNGB03]. pricing [SM09, SH95]. primer [Lin14]. principal [O’K19]. Principle [Kob97, HJSYJ12]. principles [AS08, Fra02d, GvBR10, Pur92, VM92, Hay86e]. printer [MOWGB02, Tsu01, WMGB01]. printers [WLO09]. printing [ALFM23, CB19, CNN23, DTI^{+}23, SA22, UF22, WAdN19, ZYSL21]. prior [Dag04, DVSL19, Pet11]. priorities [LB03]. prioritizing [YJLL20, YJLL21]. private [BBA15, BP09a, BP09b, DJ05, HB97, Hew99, Jon05, LO96, LBG14, Nwo97, Nwo02, Old97, Pen08, Sto21, SN92, TKBC22, WSN^{+}18]. privately [GWZ^{+}23]. privatisation [INO^{+}03]. privatization [PCG22, Whe86]. privatize [Aha88]. prix [FFJ92, FJ91]. Pro [HEC22, LT21]. pro-market [LT21]. Pro-socially [HEC22]. proactive [BLQS00b, Lar09]. probabilistic [MRC14]. probe [VGQK^{+}10]. Probert [Bes02d]. Problem [Fit00, O’H99, GBV23, Hen01, Pen08, PDHW23, Hay90a, Ric95]. Problem-Solving [Fit00, PDHW23, Hay90a, Ric95]. problematization [BJB^{+}24]. Problems [Cor87a, AGFE22, II98, MAB^{+}23, Moh95, NPG17, NJ04, OZ06, Pal00, RFB24, SFF21]. procedure [IAA94a, PP04, Bes02d]. procedures [Pat24]. proceedings [LXODT15]. proceso [Ano94r]. Process [BCKC24, HB97, MLG12, VX97, AS21, AGFdMCM24, ABLK07, Alc99, ABSZ23, Ano93p, Ano94-36, Ano94l, Ano94r, Ano94-41, Ano16h, BBTW00, BBZ20, BS06, BF97, Bes02b, BF21, BNS16, Bro01, BF06, BM94a, BM94b, BWS11, BN88, CLZ23, CB19, CM04, CSH94, CHL21, CG11, CLZ18, CLZ21, Chi21, CA06, DFBC98a, DFBC98b, DM91, DB19, EDM05, ETA02, EM85, FCS21, FGPV24, Fon05, FF92, FM24, FGS07, Gal06a, GCH96, Gim06, Gre06, GC10b, HPB03, Hos01, HLH12, HW02, IEG97a, IEG97b, II92a, II92b, IMBH98, KSHP21, Kap99, KBO22, KS21, KB09, KPK94a, KPK94b, KS04b, KFL11, LB21, LS23, LW11, Lev93, LH23, Lin07a, LLB96, LvLB06, MCGLA02, MD03, Mud90, MR00, OS99, OB20, Pan98, Par00, PA15, PGKF19, Per97b, PPP98a, PPP98b, Pia02, Poz86]. process [RNFG24, Ram23, Rot93, Sae05, SdO^{+}15, Shu22, SFF21, SY17, SW91, SH99, TSPJ23, TAC96, TCHS95, Tsu02, VB05, Vos84, VH06, WCZZ14, WBM22, ZS04b, dOK12, dV94, vDZ21, Moh95]. Process-based [MLG12, Gre06]. process-industry [BF21]. processability [LS23]. processed [KB06]. Processes [DV23, Joh92c, dCLSNL08, TAG22, AK06, AM00, Bar96, BBHA01, Can06, Car04, CS06, CKS22, Del97, DB97a, DLKM24, EWE22, FRWT06, Fig02, HKY22, HKR^{+}22, KdOL08, Kes03, KGS^{+}24, LL04a, Mat22, MKGL05, MS98, SdO^{+}15, SCFD18, SdVGdFB24, SS04, SS05, Tre98, WCT10, XHM^{+}16, dVdWNF^{+}10]. Processing [SL98a, SL98b, Ano94t, Ano94j, HBvdB09, IAA94b, IAA94a, Jus24, NLC98, OOJ^{+}00]. processus [Ano94-36]. Proctor [Poy97]. procurement [BRA^{+}23, CP05, CCF23, DLC21, GGS08, LCLC03, MGGOS21, OM18, Pat24, SRT18, UEGE^{+}14]. procurement-led [Pat24]. producción [Ano94t]. produce [IAA94b]. produced [Han95]. producers [BM05, CA07]. producing [OM00]. Product [Ano87h, BS01, BF03, BF05, CX24, DH06, Gof07, IMIY87, KWE05, MCGLA02, MOM08, RR11a, SP96, SCZ14, Tsu01, Web96, AM23, AAGL23, ASI99, AAAA23, AKBA07, ABLK07, AT93, AC08, AD09, And21, Ano81b, Ano87g, ALCLdF^{+}15, AS18, Ara08, AEP21, BYCS00, BM23, BSN21, BWF18, Bes94c, Bes95b, BF97, BF21, BM94a, BM94b, BLL95, BH15, BOBT22, BVHJ24, Cal24, CS06, CH04, CCGK23, CL05b, CLT06, CCST06, CG04a, CL97a, CKM13, Coo90, CO04, CL24, CZL^{+}24, DNT98, DVSL19, DR07, Dow93a, Dow93b, DG02, DM91, EW06, FC06, FDVB10, FMOPSVA21, For90, For91, FJ87b, FB93, Fre99, FdlCU22, FM07, GST08, GdWNPC03, GFvT22, Guy95, Haa14, HMWT23, HR98a, HR98b, Hay85b, Hay86a, HJ02, Hit89b, Hol85, Hol87, Hol88a, Hol88b, HKY22, HKR^{+}22, IOS00]. product [JL24, JR91, KdOL08, KI03, Ket09, KM11, Kod05, KBRR24, KB09, KS04a, KT04, KB06, KK06, LS23, LM14, LRH22, LLS00, LC07, LW11, LJAK21, LQ09, LCT05, LL05b, LTHS12, LZD24, Lüt04, MLG12, MdVB16, MH04, MH98, MM02, MW06, MAKKM22, MCMA09, MR00, MII90, MK96, NS07, Nob04, OC91, OOKBD14, OZ06, OE87, OB20, OW00a, OW00b, Ott03a, Ott04b, Par05, Pra99, Puj06, Pur91, RR11b, RkhR97, SZLR19, SP03, SM12a, SY14, Sha91a, Shu22, SJ21, SDK24, SCS06, SW05, SSP04, TWGB02b, TW02, TWGB02a, Tan87, TRW23, TSH02, TWBC21, TEA00, UB82, VLVH23, VB05, WGW22, WHWS22, Wer03, WBM22, WFW02, WW14, XWZ23, XL23, YZY10, ZDWG14, ZI01, ZQW^{+}19, ZYSL21, ZXDG19, dGSJ^{+}21, dOK12, CL24, JL24, LVVH23, Hay85c, Hay88d, Mor91, Sco89]. product-based [MLG12]. product-service [BOBT22, BVHJ24, CCGK23, VLVH23, XL23, LVVH23]. product-to-market [Bes95b]. production [AAA02, AT99, Ano93b, Ano94v, Ano94t, Ano94j, Ano94-33, AB89, Bar96, BHPB06, CM04, CG11, CL97b, CW93, CW94, CM96, CG04b, CGW05, CCBP22, DB97a, DB97b, EKT^{+}22, EL19, FC06, GC10b, Hit85, Hit89b, Hit96, Hit06, IAA94b, IAA94a, IIA97b, IIA97a, KK04, LB21, PS00a, SK92, SDFN01, Sto03, TOS^{+}00, vDvM89]. production/operations [SDFN01]. productive [BF04, SRDS06]. Productivity [BS87, Gol82, Sna90a, Sna90b, Sna90c, Sna90d, AAN06, BGA09, Cam93, CL21, CL22, EW06, Fal05, Gol90, GC98a, GC98b, HAH84, HHA84, Jai23, KTE04, ML08b, NW23, OÁVH14, PK23, RF12, SYK94, Sur92, Tsa05a, ZIAO^{+}21, dMvAdMSPB93]. productos [Ano94t]. Products [Sod08, Ano94t, Ano94j, Ano94n, Ara08, BL97, BFM86, Cai94, CUF13, CX24, EH02, FM07, HKRR20, JMT03, Joh92b, Lag17, LS23, LT16, MTY02, PS14, PÅ06, Red90, RWP19, SFC14, WMGB01, dRP95, vdEJG08, wK97]. produits [Ano94v]. Produktion [Ano94j]. Produktmarktbeleid [Bes95b]. professional [CS01b, CSL23, Mil88, WvT23, ZSX24]. professionals [BM94a, BM94b, DBV97, Hay88b, TC97, TNY^{+}22]. professions [CGBV21]. Professor [Ano13h, DR24]. proficiencies [KWE05]. proficiency [DM91]. profil [Ano94u]. profile [Ano94u, Ano94i, JP96, Tit94b, Ano94g]. Profiles [Bac98, Ano94g, Car01, SR23]. Profiling [CP05]. profit [FAH08, KM06b, Pap02b, YZY10, Fra99]. profitability [DMW13, KBRR24, PNC15]. profiting [Pyk07, SSKB24]. profits [Bes03a]. Program [Ano07l, YLB06, Ahn95, BCP22, Car98c, HRHW16, VS06, Hob90, VS06, Buc91]. Programme [OO02, DC22, Hob88, Hyv06, Kas94a, Kas94b, MG98c, SM97, PWA11, Hob97]. programmes [Ano93p, Bar95, CLR96, DDG98, Hob88, HHH09, Lev93, RBBB98, Won06]. programming [DB97a, DB97b]. programs [BPU21, CL05a, Chi90, LB03, MH05, Mer22, NW03b, OOKBD14, SW21, Ste14a, BMW14, LS07, LH11]. Progress [Kra15, Hja09, LPMV22, YIKC21, Zho92a, Zho92b]. progression [MM08]. Project [BK89, Dur86, HB89, KS04c, PWL20, SPFG12, STH06, AIOA00, ABLK07, ACV08, AM97, BM23, Bes03a, BDW06, CMA07, CCP23, Car98c, CLZ^{+}22, HT12, ILR^{+}12, JW06, KKG07, KL05, LB06, LLI06, LO96, Maj22, Mil02, MSR10, NW03a, OR06, PPM24, PC89, SdO^{+}15, SM05, TRW23, WPR^{+}22, vRGJ^{+}12, BPU21, Hay92d, Hay93a]. project-based [OR06]. project-driven [KL05]. Project-level [SPFG12]. project/program [Car98c]. Projects [PPM24, Sod08, AAT14, AO03, ABLK07, AKKS12, BCP11, CDD^{+}11, CDD^{+}14, CG05, CvdEG13, DFFM22, FEP15, Gun97a, Gun97b, Hob97, IBPGML23, JK03, JSR19, Kap11, KET22, KW05, LYY^{+}22, LW21, MH98, Meh88, ML05, Mik01, MRC14, MMCD23, NL11, PC89, Pra99, QvdR21, RGMW18, RRSS07, SBF15, SFF21, SVV15, SSRO24, SD89a, Sza90, TWBC21, WLH10, WH04a, WL91, YJLL20, YJLL21, ZLJ^{+}21, Tat04]. promote [HFWW16, KWvdS21, RMG06, SCH19, Thu05, WHWS22, ZL24]. promotes [HL22, ZLS23]. Promoting [NHS24, Fuj98, RR88, Tan95, RM09]. Promotion [Fuj98, Kih03, CL97b, CL97a, Els95, Hew99, KK87, Nar89, Nwo97]. proof [TWGB02b, BPU21]. Proof-of-Concept [BPU21]. propagating [WH04b, WH04a]. propeller [Gee06]. propensity [CN05]. Properties [JB19, PF98, SLJ13]. property [Ano94-30, Aug06b, ALCM14, Bad08, BE09, Bor01, Bou15, CPSM12, DGS23, Han06, LPXO14, MFH16, NHT23, Sar08, SH10b, SDN23, SS94, Tam16, TLL23, Wes93a, Wes93b, WJK15, YC05, YHDR08]. proponents [RvdSdE22]. proportionality [HJSYJ12]. proposal [PPW06, Tan87]. proposals [DB05, Lan98b]. proposed [Bro85, CS12, Kob97, SFdO^{+}23, YPF03]. proposition [RBBF23]. propositions [RKFR24]. propriety [OFR24]. prosocial [NK22]. prospective [Fah05]. prospects [II98, Swi92a, Swi92b, Hay88f]. Prosperity [Kra15]. prostheses [lS06]. protection [Bou15, DC00, HKJ07, MFH16, NHT23, TLL23]. protein [II91, SYK94]. Prototype [Mor04, BV24, LGY05]. prototyping [SY14]. provider [BVT21, CJH^{+}07, HKY22, LL04a, SSRO24, UBCF19, SOA^{+}24]. providers [AAK^{+}24]. province [CYHC19, MRGT11]. provision [Lin10d]. proximity [FGPP22, MS19, dSSD09, Ved97]. proxy [WTGB01]. pseudo [WTGB01]. pseudoinnovations [Bal85]. Psychological [RFMC21, Fel22, SMD23, MWC22]. psychology [Hay97c]. Psychometric [PF98]. PT [Ano81b]. Ptak [Fra02c]. Pub [Aug09]. Public [CLC13, CS13, Gra82, HY04, RR88, Ros89b, TKBC22, Ano86d, Ano94-34, AL11, BBA15, CT12, CCF23, CHD18, CR08, Cor87c, DLC21, Den94, DJ05, FLYTB24, Fuk09, Gre06, GGJ18, HB97, Hay83, HC06, INO^{+}03, Jon05, KS10a, KK87, LO96, Lee11, Llo07, MR05, MGGOS21, NG09, Nwo02, OM18, PC18, RATW15, RW86, SRT18, SBMP13, SCH19, UEGE^{+}14, WSF13, Zeg87, dBJ02, dBJ03, vZS14, vRGJ^{+}12]. Public-private [TKBC22, DJ05, Jon05]. publication [SL14b]. Publications [Ada07, Bas04, Box01, Hay83, Hay88f, Hay97c, Kaz03, Kuc07, Szw07, GLYP13, Lin09d, SKTM08]. Publicly [Han95, Ano86d, BRV23, Sap92, SVV15, YHDR08]. publicly-funded [Ano86d]. Published [Bai84, Bes03a, Hay01b, Ano81a, Ano81b, Bes94a, Hay88e, Hay90c, Hay97b, Hit93, Lin07a]. Publisher [Ano82b, Gar85, Ano04o, Ano04p, Ano06m, Ano19h]. Publishers [Ano84d, Ano97b, Bes88, Giu09, Hay84, Hay88a, Hay88b, Hay88c, Hay97d, Mer01, Per04, Ram86, Sap10, Par03b]. Publishing [Bes02b, Coo97, De 04, Fit00, Gar96, Hay84, Hay95, Mer02, Mor04, O’H99, Per97a, Tat04, Whi94, Won08, Dib04, Lin07a, Lin10b]. pull [BV09, DLC21, Lar09]. pulp [Ano94a, RGSK94]. Punjab [EKT^{+}22]. purchase [CP23, IMBH98, San07]. purchaser [MGGOS21]. purchasing [PMT23]. purpose [NF24, PGB^{+}23]. pursue [SRT18]. Pursuing [OB20]. pursuit [ASA^{+}22, Ros90]. push [BBLW24, BV09, CA11, Ott04a]. Pushing [PCQ24]. puzzle [Bar14]. PV [WWM00, WNGB03]. Pvt [Ada07]. pyramid [AAAA23, AS22, Thu18, FGOX23, Hal14, HMM14, OKvB23, PRA22, Se14].

QFD [HS14]. Qiwen [Wan04]. quadratic [DVdCAS16]. Quadruple [BBK23]. Qualcomm [Lak08]. qualitative [RGJ13]. qualities [SAY09]. Quality [Ano94z, CTC02, Mil94, NG97, vAdMSC94a, AHAAEA03, AM00, ABG23, Bar89a, BCPP24, Bes95c, Cam93, GR23, Gol90, GMWK23, HBH21, IBL04, JH09, JW06, LMMPVO09, MH98, McA04, MAB^{+}23, PG09, POGBG06, PS04, Ram23, SVÁG07, SSW22, SHT23, YWC09, ZZL^{+}23, vAdMSC94b, vAdMSL95, dMvAdMSPB93, wK97, Hal97a]. quantile [SSGN23]. quantitative [BNV05a, SAM^{+}23]. Quantity [ZZL^{+}23]. Quantum [GMB^{+}23]. Quasi [HST21, CCF23]. Quasi-experimental [HST21, CCF23]. Quebec [DM08]. Queen [Ugh88]. Queensland [Cha88]. Quella [Fra99]. quelques [Ano94u]. quest [LDMY22, Par03a]. question [BMSJ24, Car98a, Hoe85, McM94b, Tet00]. questioning [QWM92]. questions [NPG14, Whi90c]. Quick [Ano94m]. quickly [SC20]. Quinn [Per97b]. Quintas [Hay88e]. Quintuple [BCM^{+}22]. quo [Zho92a, Zho92b]. quoted [Box01]. Quotient [Fra00b].

R [Ano81b, Ano91d, Ano93l, Bes95b, Bes02d, Cre04, De 04, Dut07, Fra02b, Gar95, Gar96, Hay85a, Hay85d, Hay86d, Hay88e, Her92a, Joh92e, Joh92g, Kaz03, Lin90, Lin91, Pap02a, Pap03, Poy97, Ros89a, FMF21]. R&D [Koc09, LS07, SBOB23, YJLL21, CSL23, WQZ21, AAT14, AA05, Ali90, Ano93f, Ano93l, Ano93m, Ano94f, Ano94x, Ano94-27, Ano07m, AEP21, AdA08, ARR23, ALM22, Aut94, BiKM95, BMW14, BK15, BK04, BRV23, BCL22, Bje10, BC91, BC92, BP94, BGR91, CRKA98, CAI00, CLW10, CLR96, CL21, CL22, CKM13, Coc12, CLIH05, DC22, DBV97, DPH^{+}22, EBG24, EG05, ET98, FLH^{+}02, FJV96, FKT95, FEP15, Fru92, GLYP13, GW10, Gui05, Gui11, Gun97a, Gun97b, GGJ18, GGJ22, HBS02, Har97a, Har97b, HOGGP11, Hob88, HPBP98, HST21, HHH09, Hua09, HCL09, HC13, JK03, JP96, Jai23, JH09, Joh08, Jon96, Kaf06, KP12, KTE04, Kes03, KL04, Klo87, Kog03, KLV06, LPMV22, LO96, Lee04, LKPP07, Lee11, LWP14, LKCK22, LLB89, LL93a, LL93b, LYY^{+}22, LL04b, LfSS19, LFFT09, LWC^{+}12]. R&D [Lin11e, LGZ22, LG22, LL05b, Lov85, Maj22, Mal87, MZG17, Meh88, MLC24, Mik01, Miy96, MR05, NW03a, Nar04, NL11, OM89, Ott98b, Pap95, PK06, PB93, PBC93, PC89, PN11, PH08, RSP10, RF12, Rie89, SCS22, Sán94, SPFG12, SCY17, SH22, Sha03, SF05, SY05, SW91, SVV15, Sta98, SM05, SD89a, Sza90, TC97, Tag97, Tei93, TS13, Tei22, Tsa05a, Tsu02, UR18, Vek99, WCT10, WLH10, WCW^{+}14, WSN^{+}18, WSF13, WWM00, WZGBA01, WTN^{+}02, WA03, WN03, WAN03, WKN04a, WH04c, WT19, WR95, YJLL20, YHDR08, YH11, ZDWG14, ZCR22, vZS14, vRGJ^{+}12, Har97c, Paf04]. R&D-active [Tei22]. R&D-alliance [LL04b]. R&D-lessons [WT19]. R&D/marketing [Ano93f, PB93]. R. [Kuc07]. RACE [Bar95]. racing [FJ91, FFJ92, MFH03]. Radical [Mer01, SS16, AW11, Ano16h, BNS16, CCC^{+}12, DJ13, DZX22, DVdCAS16, NHV05, OKP13, PCQ24, Ros89c, SRHL12, SFF21, UF22, VMWV16, WGW22, Bes01a]. radicalism [JJS^{+}18]. radicalness [HKRR20]. Radio [ÖBA09]. Radiocarbon [Hed89]. Radosevic [Lal00]. rail [MP09a]. raised [AB89]. rallying [KC08]. Ramanantsoa [Bes94c]. Ranking [Lin06b, Lin13b, Lin13e, Tho12]. rankings [LE07, Lin12b]. Rapid [O’H99]. rapidly [BAWK12, BFM86, KL87]. Ratatouille [Leo21]. rate [Haa14]. rates [CT97]. rating [CHP13]. Rationale [SW91]. rationales [BMW14]. rationality [MAKKM22, TPFD13]. Ray [Gar85]. Raymond [Ano81a, For85]. rayon [Ano94a, RGSK94]. RDT&E [Har81]. Re [CS08b, Fra01a, ABAHP^{+}23, EH02, HKJ07, RNFG24, RG85, TS88, Bad08]. re-appraisal [HKJ07]. Re-creating [Fra01a]. Re-innovation [CS08b, RG85]. re-launching [EH02]. re-orchestration [ABAHP^{+}23]. re-regulation [TS88]. reaching [THC^{+}23]. Reacting [PPMC24]. reaction [PMN^{+}24]. reactions [AZ90]. reader [Bes97]. Readers [Fra02b]. readiness [FAH08, LO10, SFBK23, TRR^{+}24]. ready [AT99]. ready-to-eat [AT99]. Real [vZS14, CvZ05, CEY11, DSVM05, KMF09, Par03b, Sch92, VTRM12, VM97, WTGB01, WO08]. real-estate [DSVM05]. real-time [Sch92]. real-world [KMF09]. realist [KK23]. realistic [AH84]. realities [ÁRS17, Hay88g]. reality [FRWT06, FD21, MGS99, RM09]. really [Bal10, FM07, Lin12b, LLZ19]. realm [DV23]. reasons [AC07, SH06, ZS04a]. reassessment [Hit93]. Reavley [Mer94b]. rebirth [KM06a]. rebound [HR07]. recalls [KB06]. receiving [LYY^{+}22]. recently [LP03]. recipient [HHH09, LS24]. reciprocal [GKB12]. recognition [GH15, MSD15, Par05, SAY09]. Recognizing [Lin13d]. recombination [BBA15, THHS22, Wan24]. Reconceptualization [RV17]. Reconceptualizing [LS23, LAC24, Vav06]. reconfiguration [CGM07, PSS23, SCM11]. Reconfiguring [HST07, HKY22]. reconsidered [BiKM95]. record [CCR^{+}23, JL98, ND15b]. recovering [PPMC24]. recovery [AO03, CCBP22, KM06a]. recruitment [Kin00]. recruitment-case [Kin00]. recycling [Ano94a, BK99, BK00, BK01, RGSK94, SYK94, Tho99, dRP95, vdBS02]. Redes [Ano94-28]. redesign [KS04b]. Redesigning [Pur91]. Rediscovering [BCG^{+}94]. reduce [Sto99]. reducing [AT93, WNGB03, WA04a]. reduction [AMS^{+}24, Ram23, WWM00]. reengineering [HB97, OS99]. Reepmeyer [Hop04]. Reference [But96, Hay93b, Kha98d, Kha98c, NJ04, Rie89, RI04, Ruo05, dOK12]. references [Kaz03, Wya07]. refineries [Gos05]. Reflections [HKW^{+}13]. reform [Zho92a, Zho92b]. reforms [Kod07b, LT21]. Reframing [Fra02a]. refrigerators [SL98a, SL98b]. refueling [MW09]. regard [EL93]. regarding [GST08, Jol08]. regeneration [Per97a]. regime [HKJ07]. regimes [GP04, LPXO14, RF12, SCY17]. Region [Ano94-29, Ano92c, Ano94-28, BS01, CT92, GC10b, KT02, OWCvdB15, Töd94b, Ano94-28, Ano94w]. Regional [BHPB06, Ger04, Töd94b, Töd94a, BKM10, CMtW10, Chu02, EIMS01, HN92a, HN92b, Hay90c, Hit06, HLC05, HL22, Kaa09, KTSM10, Ken13, KMID11, LWL^{+}22, MMMT12, MMM16, OW93a, OW93b, RGFLRP21, RD92, Sen85, SB91, Ste14a, TKSM08, Ano94-29, Ano94w, HPT10, Ste14a]. regional-independent [Hit06]. Regionale [Ano94-29]. Regionalisation [Pap06]. régionaux [Ano94w]. Regions [Lop06, BKP23, Mal87, ST95, Pap02a]. regression [DPH^{+}22]. Regulation [FS84, II21, LKKH22, LBG14, McN92, Rot92a, Rot92b, TS88, ZYZZ24, Mer94a]. regulators [BLQS00b, Hos99, RO22]. regulatory [BS24, GFK12, Hay88g, dVV16, AMA21, Tho88b]. Reindustrialization [Hay86f]. Reinertsen [Joh92b]. reinforce [MP09a, TKYM24]. Reinhold [Bul97, Joh92b]. reinvestment [LCRL22]. Reiss [Hay93a, Hob97]. reject [Ban04, SL14b]. rejected [Lin12c, QvdR21]. Rejuvenating [BCM12]. related [Ano94-42, BN06, Dre97, Fel94, Lüt04, MO03, McQ05, RHL09, SD89a]. relatedness [Hus10, Pet11]. relating [PS97a, PS97b]. relation [OIOA01, RF12]. relational [And21, BH15, EM02, GD08]. relations [BO86a, Cul00, Hel94a, Hel94b, Kha97, OIN96, PS00a, Wec15, Gil96, Wya00]. relations-focused [Wec15]. Relationship [JW06, YGM99, ZXT10, AAGL23, CLP23, Che12, CJL22, CK12b, CD00, DTD13, EW06, Ern95, FLH^{+}02, FMFMA16, GMWK23, JWL24, JL96, KOX21, KBRR24, LKKH22, hLwChLbL04, MM08, OC91, SP03, SR23, SM12b, VTRM12, XHM^{+}16, XLC23, dlHBR22]. Relationships [FDVB10, JJS^{+}18, hLFL08, AGS98, BN06, BSWC23, Bor01, CH01, DMW13, ECB22, FGPP22, HSLL22, HESPSL11, LV09, MKL19, MMMT12, NT16, PT06, ST95, SCH19, WFW02, WF10]. relative [DPH^{+}22, LW97, ROMH15, TWBC21, UF22]. releases [CGW05]. relevance [BZIZ21, CDD^{+}11, CDD^{+}14, KK87, YHL22]. reliability [Oye00]. reliable [Ell81]. relieving [Sch09]. religions [WW21]. Relocation [EG05]. rely [TLK09]. Remaining [TWGB02a, Whi90c, HBG12]. remark [Rub89b]. remedies [WHP85]. renewable [TS05, Wal12]. Renewal [Lin86]. Renewing [GSFS21, dC07]. rent [DMP05, MPW03]. rent-seeking [MPW03]. reorganisation [Fal05, Soh97b]. repair [SH95]. Repeat [HP09]. Repercussions [CG04a]. replace [TTC06]. replacement [SH95]. Report [Pot93, Ano89g, Ano89f, Ano89d, Ano89q, Ano89-27, LD95, Wit85]. repositioning [Gui05]. repository [FRA05, LS04]. representation [AABY^{+}23, CF01a, Lin11d, MTY02]. representational [Red14]. Repressed [Bal85]. Reprint [CLZ21, YHcHW21]. reprinted [Hay90c, Hay97b]. Republic [Ste89, For85]. Reputation [Koc09, EBG24]. reputational [RM14]. require [RFB24]. required [WVG^{+}98, WGV^{+}98]. Requirement [CWB98a]. requirements [TS13, CWB98b]. requiring [Ano93b, CW93]. Research [BE09, CR02, Hay86c, HJ02, JKF19, KI14, dLP06, Lid90, Lin17c, MPCRB^{+}03, OWCvdB15, OB04, Ros89a, Ros89b, Sou83b, TS13, Tit90b, Tuz05, AM23, ADD19, Ahn95, AHS^{+}23, AAN06, Ano86d, Ano94z, AKKS12, AL11, ÁRS17, BCKC24, BBTW00, BPU21, BD23, BCP22, BB22, Bro95, BGM88, Buc91, Bul85, CMA07, CPSM12, CLP23, CRKA98, CWAB09, CYM^{+}16, CYHC19, CZZM21, CLW10, CbLjJL12, CL22, CR08, Col81, CFC12, CR19, CKV^{+}23, CP06, Cor87c, Cur89, Da 01, DMV^{+}21, DMRS23, DDG98, DBV97, DNP01, Dod87a, ET98, EFV02, FBE23, Fel22, FHA22, FH87, GC10a, GLÖS23, GFE95, Gri06, Ham96, HV96, Hay88b, Hay97c, HBO14, HY04, HG92, Hol85, Hsu05, HM11, IM10, JL13, JFCFMC^{+}12, KTE04, KS21, KET22, KFEC15, KS10a, LLI06, LPMV22]. research [LbKL91, LW04, LS07, LAC24, LDT^{+}23, LfSS19, LB87, Lin02, Lin10b, Lin12b, Lin13c, LBA14, Lin16, Liy95, Llo07, MdVB16, MCGLA02, MMMD23, Mas23, MZL15, Mil94, MS19, MI07, MGAP23, Mos16, Nar89, Nob04, NC23, NI05, OVD21, OIOA01, OBvE97, OM89, Ott03b, PLG24, PPM24, PA15, PV01, PS01, PP02, PG11, PUFC23, Ras08, Ree99, RS03, RS14, Rei97, RSA05, Rie89, Rog86, RH23, Ruo05, SBOB23, SOKE15, Sér96, SD23, SKTM08, SO06, SCH19, SD89a, Tit90a, VM97, VS06, WTN98b, WTN98a, Wal07, Wec15, WML^{+}23, XL23, XLD23, YHDR08, YLB06, ZKW^{+}23, ZWD09, ZCF19, ZC22, ZLJ^{+}21, ZT21, ZAI23, vAdMSL95, vDZ21, Lin07b, YLB06, Dod87b, Hay96a]. research- [Ott03b]. research-based [Bul85, CFC12]. research-industry [PUFC23]. researcher [KI03]. Researchers [Ano09l, BP09a, BP09b, Lin10b, LP07, RV21]. réseaux [Ano94w]. reservation [Sch92]. Reserve [OO02]. reshapes [CHT22]. residential [NLLC04]. residents [CL12]. Resilience [GPCNS^{+}23, WKN04b, HPMI^{+}23, TFS^{+}22, VSE24]. resistance [QMP24, SF98, ZSX24]. resolve [Lin85]. resolving [Mac05]. Resonance [Hay01b, BF06, Hal97b]. Resonant [WH04c, BM05]. Resort [Ano07k]. Resource [ABAHP^{+}23, Ban03a, Cha22, DVSL19, And21, ABUS05, BM16a, BVS24, xCZZ^{+}22, FMV00, Gal05, GK06, HKL24, Jan21, Kra22a, LLY^{+}18, PS00b, SO02, SY14, SXS24, Sun96, TA10, Yas05, YWL22, YGW06, YGW08, ZWZ^{+}23, vRE21, vWvRN17]. resource-based [xCZZ^{+}22, FMV00, Gal05, Yas05]. resource-constrained [BM16a, HKL24, YWL22]. resource-poor [SO02]. resources [Can16, Car01, CE89, CKM13, CS13, FMV00, Gal05, HLL12, HAH84, JKF19, Joh08, LC09, MM08, NGK08, PKH15, PSdLC02, PSdLCJ05, RS03, SAWT06, CNN23]. respond [Shi24]. Responding [Lin19b]. response [Ano89p, BK99, MAMD22, WG23, vdVdM11]. Responses [Dre04, Lim06, ACA24, AAK^{+}24, Boh18, FC95, GWZ^{+}23]. responsibilities [SDFN01]. Responsible [ZCHC23, Fel22]. responsive [Mil02]. responsiveness [STS88]. rest [Hal97a, ND15a]. restorative [CK04]. restriction [DL10]. restructuring [Cha99, MJ92]. result [MPCRB^{+}03]. resulting [BF04, PG11, TWGB02a]. Results [CMA07, Ano86d, BSG^{+}24, Cor87c, CNPV04, EFV02, GMNL07, GÁNA07, Lay93, NS90, NPV05, Sou01, WTAW97, Wit85]. retail [BJB^{+}24, HWC06, JMPR01, SF98]. Rethinking [VM11]. RETRACTED [BK01, CDD^{+}11]. Retraction [CDD^{+}14]. retraining [WC85]. retrieval [DF98a, DF98b]. retrospective [EKT^{+}22, LCRL22, WHB21, ZWDZ21]. return [Wal07]. returns [BZIZ21]. Revealing [Lin17c]. revenue [SM09]. reverse [AAAF22, Els95, HS14, VJ23b, ZZXW23]. Review [Ada07, AHS^{+}23, Ano81a, Ano81b, Ano91d, Ano91c, Ano97b, Arc03, Arr04, Aug09, Aug06a, Aug06b, Bai84, Ban04, Bas04, BB95, Bes88, Bes93, Bes94a, Bes94b, Bes94c, Bes94d, Bes95a, Bes95b, Bes95c, Bes96, Bes97, Bes01a, Bes01b, Bes02a, Bes02b, Bes02c, Bes02d, Bes03a, Bes03b, Bon02, Box01, BH03, Bro90, Bro07, Buc91, Bul97, But96, Col97, CF07, Col08, Coo07, Coo97, Cou90, Cou91, Cre04, Dav92, De 04, Del97, Dib00, Dib01, Dib02a, Dib02b, Dib04, Dod87a, Dod87b, Dod88, Dre04, Dre06, Dut07, Edw04, Eva02, Fit00, For90, For91, For85, Fra99, Fra00b, Fra00a, Fra01a, Fra01c, Fra01b, Fra02b, Fra02c, Fra02d, Fra02a, Fra03a, Fra03b, Gar85, Gar87, Gar94, Gar95, Gar96, Gil96, Gil97, Giu09, Gof07, Gra01, Gra03]. Review [Gra04, Gri98, Hal97a, Har93a, Har93b, Har97c, Hay83, Hay84, Hay85a, Hay85d, Hay85b, Hay85c, Hay86c, Hay86d, Hay86e, Hay86a, Hay86f, Hay87a, Hay87d, Hay87b, Hay87c, Hay88a, Hay88b, Hay88c, Hay88d, Hay88g, Hay88f, Hay88e, Hay89a, Hay89d, Hay90a, Hay90b, Hay90c, Hay90d, Hay91a, Hay92b, Hay92a, Hay92c, Hay92d, Hay92e, Hay93d, Hay93a, Hay93b, Hay93c, Hay95, Hay96a, Hay97b, Hay97c, Hay97d, Hay97a, Hay01a, Hay01b, Hay01c, Her91, Her92a, Her92b, Her92c, Her04, Hob87, Hob97, Hob03, Hol93, Hop04, Jel07, Joh92a, Joh92b, Joh92f, Joh92c, Joh92d, Joh92e, Joh92g, Kap01, Kap03, Kaz03, Koc09, Kra15, Kuc06, Kuc07, Lae03, Lae04, Lal00, Leh07, Lim06, Lin90, Lin91, Lop06, Mer94a, Mer94b, Mer01, Mer02, Mil85, Mil88, Min94, Min96, Mor91, Mor04, Mur91]. Review [New04, Nuv07, O’H99, Oak97, OH04, Pap02a, Pap02b, Pap03, Pap04, Pap06, Par03a, Par06, Par07, Par03b, Per04, Per95, Per96, Per97b, Per97a, Pot93, Pov97, Poy97, Pra10, Pre01, Que06, Rad97, Ram86, Rea06, Ric95, RL06, Ros89c, Ros89a, Ros89b, Ros90, Rot87a, Rot88, Rus94, Saa98, SW04, Sap01, Sap05, Sap10, Sco89, Sha95, Sha91a, Sha91b, Sod08, Szw07, Tan08, Tat04, Tra00, Ver90a, Ver90b, Ver91a, Ver91b, Wan04, Whi94, Why02, Why05, Wil90, Won08, Wya00, Wya04, Wya07, dC07, AGFdMCM24, ACBO23, Ano95a, Ano98l, Ano98k, Ano98f, Ano98g, Ano98h, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano98i, Ano98j, Ano99a, Ano99b, Ano99c, Ano03b, Ano04b, BLA06, BD23, CPT23, CPSM12, CFP^{+}23, CLGFVCM23, CYJ07, CKV^{+}23]. review [Cor87c, CZL^{+}24, D’I14, DMRS23, EKMS12, FLYTB24, GLÖS23, GBY22, GW87, GH21, HGB23, Hit89a, Ho09, HPBP98, IPP13, LAC24, hL05, MDT^{+}23, MdVB16, MMMD23, MVS^{+}22, MGAP23, Mud90, NPG14, OS99, OVD21, ODM23, OGRA23, OZ06, PMP24, PS01, RH21, RPASLM23, RC23, SP22a, SS23, SG22, SSKB24, Sou83a, SW22, SMD23, WML^{+}23, XL23, XLD23, ZKW^{+}23, ZWD09, dMD05, dBJ03, BB95, Hob03]. reviewed [RRSS07]. reviewers [Lin08a, Lin10d, Lin11b, Lin12a, Lin13a, Lin19b, Lin13d]. Reviewing [Lin09d, Lin10b, AS22, Jus24]. Reviews [Ano02g, ACV08, Ano92f, Ano93h, Ano94d, Ano96a, Ano01b]. revisited [Can06, SDK24, Wan91, WFWC21]. Revisiting [HKRR20, LC22b, TKBC22, WBM22]. revitalization [YW08]. revival [NWGB09]. revolution [II21, JV97, Won05b]. revolutionizing [Gil97, Boy14]. revolutions [Kap03]. reward [EHSB24]. rewards [ZC19]. RFID [Ho09, ÖBA09, BWS11, CYJ07, DR07, WNLY06]. rhythm [BEH99]. Rice [Bes01a]. rich [SPÁG22]. Richard [Bes01a, Box01, Fra02a, Hay88a, Ram86]. Richards [Lae03]. richness [LL08b]. Rickards [Fit00]. right [BHS05, Lic10, SvG15]. right-balancing [BHS05]. rights [Ano94-30, CPSM12, DGS23, Han06, LPXO14, MFH16, Sar08, SH10b, SDN23, SS94, WJK15]. rise [GKK23, Gar94, KHL15, KP23, TSA13a]. Risk [KS14, SGWL16, BCPP24, Boy14, Car98b, EKT^{+}22, FSA23, HBM14, IPP14, KV14, KS04c, KL05, MRC14, MPM09, OOKBD14, OBL14, WLH10, Web99, WW14]. risk-aware [IPP14]. Risk-taking [SGWL16]. risks [BiKM95, DDBM23, HT06, LTHS12, MDT^{+}23, SR22]. Risky [Min96]. rival [HLO14, PS21]. rivals [EBG24]. River [Fra02c]. Road [Bes95a, Per97a, Tat04, CDD^{+}11, CDD^{+}14]. roadmap [DMV^{+}21, JY15, Li09]. Roadmapping [CA11, IPP14, LKPP07, NVMD22]. Robert [Aug06b, Bes01a, Min94, Per97a, Ros89c]. Robin [Hay97a, Ano94k]. Robinson [Mil85]. robot [II21]. robotics [Mac97a, Sgo95]. robotisation [AAZ22]. robots [Bai84]. Rod [Lae03]. Roddam [Bas04]. Rodney [Tat04]. Roger [Ano94z]. Rogers [Hay84, Ada07]. Roland [Par03a]. Role [EL19, FdlCU22, SCY17, SSC23, AAAF22, AMV07, AA91, AGS98, ASA^{+}22, Ali86, AGFE22, Ano94i, Ano94-30, Ano94-31, Ano94-32, Ano94-34, AAZ22, AN16, ABAHP^{+}23, AFL^{+}24, AM97, BSSF24, BBDM01, BCPP24, BRA^{+}23, Car98c, Car99a, CR00, CUF13, CCGK23, CR06, CK12a, CL22, CJL22, CJL23, CGLSNLDV15, DC22, Dan19, DZX22, DDN24, DID21, DL10, DVSL19, DP23, Den94, DLTQ23, EBG24, Fah05, FLC17, Far23, FCM21, FMOPSVA21, GR23, GST08, GL22, GHU21, GMWK23, HKRR20, Hal96, Hay89d, HESPSL11, HO11, HW18, HBvdB09, Hol88b, How94, Ilo94, KBSW08, KKWM02, KBRR24, LPN22, Liu93, LLY^{+}18, LZD24, Mac97b, MSD15, MCA^{+}24, MZG17, MRDA24, Mol98, Mou15, Mou17, MK96, NUSR21, Nwo97, NBA22, OKP13, OB21, PS21, PMT23, RH10, RGB23, RO22, RGFLRP21, Rog86, RvdSdE22]. role [RG85, RHW14, SW21, SRHL12, SSCH23, SBP21, SVÁG07, SOKE15, SC10, SLSCS24, SP22b, SXS24, SSW22, SH23, SCM24, SGWL16, SS94, TKYM24, TLL23, TPV16, Tho88c, Tõn15, Udd06, URG^{+}24, VGAL24, WvdWG24, WGW22, WK03, WKN04a, Won13, WW14, XLC23, XLG^{+}23, YHcHW17, YHcHW21, ZDWG14, ZLJ^{+}21, ZT21, ZMJM23, ZLH17, ZGSG23, dAASA21, ZS04b]. Roles [Joh08, ALM22, BB22, Bre23, BLQS00b, Dag04, KY06, MV06, QMP24, SDFN01, SJCT24, TRW23, WA04a, WHWS22, ZZL^{+}23]. ROM [But96]. Romania [Tho99]. Ron [Bro07]. Ronald [Bes03a]. Roomba [HKRR20]. Rory [Bes94a]. Rosa [Rad97]. Rosen [Tra00]. Rosenberg [Ros90]. Rothwell [Hay83, Hay86f, Hay95, Ram86]. rough [WCT10]. round [Ano94b]. Route [dCLSNL08]. routine [Fis89, JC01, WZY23]. Routines [Gos05, HKR^{+}22]. Routledge [Ano91d, Ano91c, Aug06b, Bes93, Eva02, Gar95, Har93b, Hay89d, Hay90a, Hay90b, Hay90c, Hay90d, Hay91a, Hay92c, Hay92e, Hay97b, Pap06, Pra10, Ric95, Sap10, Tra00, Wil90, Wya00]. Rovere [De 04]. Row [Hay87d]. Rowland [Del97]. Rowman [Wya04]. Roy [Ano94k, Hay83, Hay86f, Hay95, Ram86]. Royal [Hay86c]. royalties [Bal10]. RTD [MPCRB^{+}03, RBBB98]. rubber [Buc91]. Rubenstein [Ver91b]. Rue [Ano81a]. rules [KH89, SS22]. run [CKF^{+}24a, SCC17]. rural [CP11, EKT^{+}22, LVS10, TPB12, TDA05, ZMPRP14]. Rush [Hay97a, BCO14]. Russia [GFK12, Kih03]. Russian [SMC14]. Rutten [Pap02a]. Ruza [Sha95]. RVINs [CvZ05].

S [Ano88h, Ano91d, Ano93l, Bas04, BB95, Bes94c, Fra03b, Hay88g, Hay88e, Hay90b, Joh92f, Pap02a, Pap03, Par07, NLC98]. S. [Aug09]. sacrificing [TWGB02a]. Safeguarding [SPO08]. safety [CJH^{+}07, Big13]. SAGE [Bas04, Szw07, Ada07, Box01, Fra03a, Fra03b, Hay97c, Kuc07]. Saharan [AHSS19, AN16]. sales [SSM09, SLSCS24, SSP04, WA03]. salesperson [LD05]. salients [HS14]. Sally [Ano94n]. Salmon [New04]. Salter [Jel07]. sample [KBC22, LLZ19]. Samsom [Ano93s]. Samsung [GBL03]. San [Fra01a, Fra01c, Leh07, Mer01]. Sanchéz [Ano94-27]. sandboxes [AMA21]. Sandia [CG02]. Sanjay [Ano94f]. Sanjaya [BH03]. Sankar [Hay92a]. Sanofi [BLLEM23]. Santarelli [Oak97]. São [KdOL08]. Sartorius [Par07].

SAS [Obe86]. SaskPower [CK04]. satellite [Bas01, CR00, CA01]. satisfaction [CLW10, LK20, MH98, OGOF22]. Saudi [ASGC09]. Scaffolding [KM22]. scalar [BCV21]. Scale [SC20, AO03, ABKL08, BZIZ21, Cal24, KC07, LO10, MPWG22, Ola98a, Ola98b, Ros82, RBBF23, TMG11, WSN^{+}18, ZCHC23, Tik23]. scales [PF98]. scaling [SS16]. scanning [BDW06, VGQK^{+}10, vW84]. Scarbrough [Poy97]. scarce [CS13]. scenario [MCD06, Sæt24]. scenarios [Sæt24, Rot88]. scheduling [SM05]. scheme [Par00, JC01]. schemes [CL97b, CL97a, RATW15]. scholarship [VM97]. scholarships [Rei97]. School [Bes95a, Bes96, Bes01a, Bes01b, Dav92, LWM24, SW04, Lin12b, LC82]. schools [Kin02, LD95, Lin12b]. Schoonhoven [Cou90, Cou91]. Schrage [SW04]. Schragenheim [Fra02c]. Schumpeter [Won05b]. sci [PMN^{+}24]. sci-fi [PMN^{+}24]. Science [BMV02, Bes88, Col81, Dib00, GC10a, Guy96a, Hay93b, Hay93c, Hob87, Kos99, LY00, LYST22, Lin13c, MM08, Rad01, Sæt24, Szá96, Ved97, Ano94-42, ABS24, BV24, BDNP06, CBGvdV22, CL05a, Cha99, DVMS16, DSVM05, Fel94, Fra02b, Fre86, GBFvdS11, HHV05, Hay87c, HPT10, IAO^{+}02, IH07, Kih03, LS05, LK20, LD95, LG23, Lin08b, LDMY22, MG21, MCA^{+}24, MT05, NP10, PWL20, Per04, Phi99, QWM92, RH10, RH23, DGMR21, Tom03, Tuz05, VH06, WPR^{+}22, ZZL^{+}23, De 85, DLL06, LL03, LL05b, SP22a, Ano91d, Hay87b, Hay88e, Hay92e, Hay93b, Hob87, Mer94b, Rot87a]. Science-based [LS05, Fre86, GBFvdS11, NP10, DGMR21]. sciences [SY17]. Scientific [ZCF19, ADD19, CM04, CR08, Col81, Fon05, Mac97b, Mou15, Mou17, SKTM08, Szá92, Wan24, YIKC21, Ano97b]. scientist [GS10, TBSM21, Wür17]. scientist-entrepreneurs [TBSM21]. scientist-started [GS10]. Scientists [CSL23, Koc09, AOO02, Ano93s, BSWC23, Boa08, HM11, KFEC15, LPN22, Lin17a, NL11, SG93]. SCM [DB09]. Scope [MPWG22, CCV11, CS89, SAWT06, SH22]. score [MYR^{+}21]. score-based [MYR^{+}21]. scorecard [FRW06]. Scotland [Kin00]. Scott [Rad97]. scrap [vdBS02]. Screening [HKW^{+}22, AMV07, MNW14]. SDH [CCY05]. SEA [Hos00]. Search [CHL21, AMP22, BBS22, Coo90, CGLSNLDV15, FMFMA16, FDN^{+}23, GBV23, JK03, JSR19, Jus24, Kli06, LB21, Mac86b, MAMRRR24, NPG17, OGG06, SO06, WG23, XHM^{+}16, Zak23, ZMJM23, vGS09]. Searching [BFM88, RZ21]. Second [Bes94a, HHV05, Hay85c, Kra15, OB20, Bai84, IKL01, CL12]. second-tier [IKL01]. secondment [HRHW16]. secrecy [DL10, FBE23]. secret [FBE23]. secrets [Her91]. section [Sna90a, Sna90b]. sectional [Sna90d]. sections [Sna90a, Sna90b, Sna90c, Sna90d]. sector [ARvT04, AEP21, Bad08, BBA15, BLA06, BE09, CCP23, Cha22, CBK03, CH01, FLYTB24, Gre06, HB97, Hew99, HC06, HKL24, JP99, JRS06, KdOL08, KDH02, KS21, KK04, KS10a, Lee00, LGZ22, Mas23, Men11, MPO23, Moo03, NG09, Nwo97, Nwo02, Oye01, PR96, PGKF19, PP02, PEER23, RDS17, SPKK08, SMI23, SN92, TDC11, Thu18, UOIA00, WSN^{+}18, WSF13, YW08, ZYZZ24, vZS14, wK97, wK98, Poy97]. sectoral [LM24, MG08, OS08, SCY17, SSF12]. sectors [BJB^{+}24, Bes15, CF01b, Der14, EFC23, HW01, Kod00, Neu12, SSF12, YW08]. Security [RY14, Ven14, Wil14, Bar14, GPCNS^{+}23, Kee14, KI14, LBA14, SMC14]. seed [CKR00, CYK19, FOL^{+}03]. seed-oil [FOL^{+}03]. seedbed [Ano94-42, Fel94]. seeds [TW02]. seekers [BVS24]. seeking [FdJ09, MPW03]. segmentation [Thu18]. seizing [KBO22]. Selected [Lop06, BDNP06, FOL^{+}03, Ger04, IOS00, Kim87, MPBC21, Hja09]. Selecting [AAT14, LS07, Meh88, Mil02]. Selection [CMB23, LL08b, Ano93e, BC10, CLT06, CLIH05, CNS06, JH18, Kim87, LLI06, LO96, LGZ22, SPP06, SM05, Sza90, Ven96, YS93]. Self [RK04, Bøl12, EN07, FWL06, LCWC09, OC85, Vos05, WH04b, WH04a, Joh92e]. self-descriptions [Vos05]. self-efficacy [LCWC09]. self-generated [Bøl12]. self-leading [FWL06]. Self-managing [Joh92e]. Self-organizing [RK04]. self-propagating [WH04b, WH04a]. self-service [EN07]. sell [SLJ13]. Selling [RKFR24, SCH23]. semantic [LFU23]. SEMATECH [CG02]. semiconductor [Ano94-35, BA89, FL15, Hob90, Kan94, Kap01, LH07, Liu93, SY05, VJ23b, YI05]. senior [CHD18]. seniority [Bro98]. sense [Bro07, Gal06a, HOGGP11]. Senseable [SSS17]. sensemaking [AAM24]. sensing [HBvdB09, WKF24]. Sensor [AJLR04, TB24]. sentiment [Fel23]. September [Ano21b, Dre04, Ano21x, Ano22t, Ano23t]. serendipity [BBLW24, WPR^{+}22]. Series [Hay88a, Hay88b, Hay88c, Hob87, Lae04, Rus94, MOWGB02, Que06]. Serious [SW04]. serve [RM14]. server [TCC^{+}22]. Service [Hay93b, SNM12, VM12, AHS^{+}23, AM00, Ano12a, BF21, BOBT22, BVHJ24, Can16, CY12, CMH14, CCGK23, CNN23, CWY14, CCY05, CK12a, CS98, EKMS12, EN07, HWC06, HT12, HLH12, HKR^{+}22, HGL98, II01, JLL10, Kod14, LVVH23, LL04a, LH14, MG21, MTY02, Men11, Ove15, Sar22, SMvFM24, SWW12, She09, Sun96, TPV16, TDC11, TPBM23, VLVH23, VvWM20, VNJ18, XL23, YH09, CMH14, JLL10, Dre04]. Service-based [VM12]. services [Ano93o, Ara08, Bad08, BCPP24, BE09, BRCG12, BFM86, CCR^{+}23, CJH^{+}07, GC99, HLL93, Hay92d, Her92a, JMT03, JSP22, Kod00, Mac86b, ML08b, McN92, PWF07, PC18, Poy97, RMGDLC08, SVLYT24, SS03, TKBC22, WvT23, YH09, LACH13]. Servitization [XLD23, KBSB22, KBRR24, KRM^{+}24, LVVH23, LV23, RNFG24, SLA^{+}23a, SOA^{+}24, TKBC22, TCMS22, URG^{+}24, WGW22, WHWS22, ZYZZ24]. servitization-based [LVVH23]. Servitized [BRA^{+}23, BSG^{+}24, FdlCU22]. Sessions [Kle13]. set [Bar14, DLKM24, WCT10]. set-theoretic [DLKM24]. setting [SPÁG22]. settings [DS23, SdVGdFB24]. Seven [Hay88d, CTC02]. Sevenoaks [Rot88]. Seville [MRGT11]. SFr [Hay96a]. SGI [Rus94]. SGRSP [OO02]. Shameen [Pra10]. Shane [Aug09]. Shanghai [CPXC02]. shape [BK24, HCMMH21]. shaped [BG19, CLP23]. shapes [ABvZ23, CS21]. Shapiro [Hal97a]. share [AD09, CK18, Vit90]. shared [dABKS14, LKKH22]. Sharing [Chi81, Koc09, ALFM23, BD23, FN21, Hua09, LM14, MOM08, Öbe23, RES^{+}24, ROMH15, WH21, vAdMSC94a, vAdMSC94b]. Sharpbenders [Her91]. Shedding [DTD13, PP09]. Shelf [But96, Gri06]. shift [FG06, Gee06, Vit90]. shift-share [Vit90]. Shifting [RUNS23, RS14]. shifts [NWGB09, SCZ14, TE14]. Shih [Ano94l, CF07]. Shiman [Par03a]. shipping [THHS22]. Shocks [Lim06]. Shocks-origins [Lim06]. Shonku [DR24]. shop [BNV05b, CG04b, CGW05, Gaz86, LB82]. shop-floor [Gaz86]. Shopping [TLL07]. short [Ano94c, GW87, Hit94]. shortage [II91]. shortfalls [ZCR22]. Should [SRT18]. shoulders [JWY23]. Shriberg [Fra02d]. Shuldiner [Bes02a]. side [CA14, DSPS23, FEP15, McM96a, PEER23, SW21, SV01]. sided [CT12]. Siegel [Per04]. Siemens [OS98]. sigma [KA06]. signal [NLC98]. signals [CCB20, PLC13]. Significance [Els95, BF04, Gal12]. significant [WK03]. Sigvald [Fra01b]. Silicon [BC82, De 23, LG22, OWCvdB15, iSKB02, Won06]. silver [BF23, KHL15]. Silverman [Col97]. similar [Mas86]. Simon [Gra01, Tan08]. Simona [Arc03]. simple [Ree99, SS22]. simulate [SW04]. simulation [Gre06, NPG17, RES^{+}24]. simultaneous [Ano93b, CW93, DGMR21, ZC22]. Simultaneously [Ano94-33, CW94]. Singapore [HRHW16, LW04, MJ92, OWC03, WOL05, WW04, Won13]. Singhal [Ada07]. single [ABGS13, ET98, IMIY87]. single-group [ABGS13]. SIP [HST07]. site [NJ04]. sites [Bea05]. Situated [MMM16]. situation [Cor87c, FB91]. Six [Rus94, KA06]. size [AA88, Ano94-27, AES95, BBT19, CF01b, CY00, Cor87a, Egb15, FRW06, GPR89, Hay90c, JJS^{+}18, Kog03, LAYS19, Man07, RF12, Sán94, SF05, SGF02, Tit94c, Tit94d, Tsa05a]. sized [AIOA00, Arc03, tBL86, Cor87b, CL88, EDM05, FDN^{+}23, GC96, Hit89b, HRHW16, Kim87, OCA09, Rie89, Rot91, SPKK08]. skill [Sch94, WG23]. skill-based [Sch94]. skilled [PSG22]. Skills [Mer01, Bas04, CA07, DMNR12, Fre05, MRGMBR13, SAY09, vWGdG21]. Skinner [Tra00]. slack [HC10, Tei22, YWC09]. slash [Bes03a]. slaughterhouse [Sto99]. Slavo [Lal00]. sleeping [BCM12]. slip [Ram81]. Slovene [KBS01]. Slovenia [DNS^{+}03]. Slow [LvLB06]. Slywotzky [Fra99]. Small [Arc03, AVA22, Coo16, DR88, HF18, HRHW16, HPBP98, Pra10, Sen85, Tet00, WNM95, ALvKL06, And21, Aug06b, tBL86, Ber96, CR00, CY00, CBK03, Cor87b, CL88, CS13, Dea81, DLL06, DTI^{+}23, Dur86, EDM05, FRW06, FDN^{+}23, Fre99, Fre05, Gar85, GC98a, GC98b, GC96, Had99, HdWN21, Hit89b, JR91, JH18, KdOL08, KKWM02, Kim87, KJE96, LAYS19, hLwChLbL04, Lin08b, LOCSBGMH16, MY02, MB04, OG05, OS08, OC91, Oak03, OCA09, Ola98a, Ola98b, OM00, OOL06, RATW15, RMGDLC08, RH21, Rie89, Rot91, SPKK08, SU11, SAM^{+}23, Soh99, SBMP13, SL16, Sub05, Tit94c, Tit94d, TA10, Udd06, VJ12, VD04, VX97, Wal12, WVG^{+}98, WGV^{+}98, Bul85, Hay90c, Per97a]. Small-and-Medium-sized [HRHW16]. Small-firm [AVA22, WNM95]. small-sized [Hit89b]. small/medium [GC96]. small/medium-sized [GC96]. smaller [FJ87a, VM92]. Smart [M’C05, MGAP23, HKY22, HKR^{+}22, KS14, Lee23, NUSR21, PC18, QMP24, SM95, SSRO24, WH04c, Won04, BOBT22]. SMBs [BMRG00]. SME [LOCSBGMH16, AHSS19, AHP16, ASOB23, FORA17, MCT06, RB09, RIMP22, SPO08, Sik99]. SMEs [KdOL08, ALBO08, Ano94n, BDB21, BRA^{+}23, BM07, BP01, Cai94, CW06, CAP23, CL24, EFV00, FSA23, FAH08, FB06, FMOPSVA21, GMLMVJ07, GKB12, Had99, HRHW16, Hus10, Hyv06, JPBL^{+}23, KP12, KT02, KDH02, KH06, KS10b, LB03, LNB21, LLI06, LLR95, LLEB05, LTVA22, MCH00, MCGLA02, MRGT11, MMC^{+}23, MT08, MRDA24, Nar04, NT16, OGS06, OGG06, PNC15, PCQ24, PBOR23, PvdMB^{+}06, RV21, RIMP22, RAG11, RSP10, RPL10, RD92, SRT18, SB02, SPO08, SDFN01, TS13, Tei22, VSE24, Vos05, ZXT10, dOK12, dBJ03, vdVdJVdR09, Coo07]. SMEs. [KDH02]. Smilor [Lin90, Lin91]. Smith [Joh92b, Whi94]. Snelson [Sha91a]. Social [BAR16, GLÖS23, Hun00, OSSC21, Sto21, AM23, AK06, BC99, Fra02b, HMM14, HHV05, Kaa09, KRK^{+}22, KKP17, De 23, MTB01, Mil85, NHS24, NK22, PRA22, PRID23, PGB^{+}23, QMP24, RVW17, RGMW18, RD05, Rot87a, SLA^{+}23a, SPÁG22, SPMRS23, SHS^{+}23, DGMR21, Ulh05, ZXX23, ZR93a, ZR93b, ZGSG23]. social-purpose-driven [PGB^{+}23]. Socialization [FF00]. socially [HEC22]. societal [Hea83, LDMY22, DGMR21, vdWFvdB23]. societies [FLC17, Zeg87]. Society [Ano81a, Ano87h, Lin13c, Ano87g, BCM^{+}22, GS05, HHV05, Web99, Hay92c]. Socio [RG10, AA90, JLY23, LKKH22, LAC24, PPW06, SS16]. Socio-cognitive [RG10]. socio-economic [AA90]. socio-tech [LAC24]. socio-technical [LKKH22, PPW06, SS16]. Sociometric [MTB01]. Sociotechnical [Pur92, Mol97b, Mol97a, Mol98, Pur91]. soft [Gei97, Pfi99]. Softback [Hay85a, Hay86c, Hay86f, Sha91b, Hay87a, Hay87b, Hay87c, Hay96a, Hay87d]. Software [Ara08, Ano94q, Ano94-45, Ban03a, Ban03b, BAR16, BV09, CHP13, CN05, Day02, DO03, FB93, HJM08, HKR^{+}22, II01, Ket09, KH06, KS04c, LL94a, LL94b, LQ09, MO03, McQ05, MD03, MSR10, OM00, PP02, Pyk07, RI04, Ruo05, SM09, SH10b, Tsa05b, YA11, Car98a, Mor04]. Softwareindustrie [Ano94-45]. solar [Kap99, KGS^{+}24, Ove15]. sold [AGFE22]. Soldiers [Bes95a]. solution [AAN06, GBV23, HKY22, HKR^{+}22, II91, Ver90b]. solutions [BCM^{+}22, HM11, HBH21, LDT^{+}23, MAB^{+}23, NPG17, OSSC21, TSPJ23]. solvers [HBH21, TP24]. Solving [Fit00, O’H99, Hen01, PDHW23, Hay90a, Ric95]. Some [Arn87, AB89, Ban03b, KM11, Seg86, Ano94u, Ano94-27, Die00, FOL^{+}03, GPR89, HB97, IIA97b, IIA97a, Jac22, LVS10, OW93a, OW93b, Oak03, OB04, OIN96, Rag93a, Rag93b, Sán94, Sen85, SB91, SL98b, TM03, WHP85, dGSJ^{+}21, vAdMSC94a, vAdMSC94b, AM92a, AM92b, AC07, BCO14, HB84, Hol85, Lin10b, Pat24, RAC84, Rob84, Rot92a, Rot92b]. SONET [CCY05].

SONET/SDH [CCY05]. Sons [Del97, Fra00b, Fra02d, Gil97, Gri98, Hay85d, Hay85b, Hay86d, Hay92a, Mil85, Per95, Pov97]. sophisticated [NC23]. sorghum [llo91]. Souder [Ano93e]. source [BAR16, CKR00, FN21, Far23, GFvT22, JMK05, LS24, MSR10, PMN^{+}24, QMP24, RH21, TW02, WH04b, WKN04b, YA11, YW08]. source-recipient [LS24]. sourced [LS16]. sourcer [ARR23]. Sources [AL05, Cab19, MM02, Pat24, SFF21, WN03, AOO02, CKT04, CCV11, DVHL22, LTT23, Pal06, SBP21]. sourcing [ARR23, BCP11, HCL09, JBL23, LT10, LW10, MG98b, MG98a, MPWG22, NB09, NB11, ND15a, PEER23, Rad99, RDS17, TWBC21]. South [HCK01, DR24, AOO02, BBDM01, Cha22, Chu01, GBL03, HBO14, HB03, Hos99, HH08, KK04, LS07, LKKH22, LP07, Moo03, MBA^{+}12, PH08, RPL10, SPO08, SH10b, Ven14]. South-East [HCK01]. Southampton [Hay93b]. Soviet [Hay92e, AHP16, Nay88, Win87, Hay88f]. Soviet-type [Win87]. soybean [DAOO02]. Spa [Ano07k]. Space [CNPV04, PV00, CCF23, CE19, EF11, Wya00, BC99]. spaces [CB22, KM22]. Spain [MRGT11, MP09a, Ano94-27, BGA09, BMLL07, dCLSNL08, NDP07, RDS17, Sán94, ST95, SNM12, VBA06]. Spanish [Ano94t, Ano94q, Ano94i, Ano94r, Ano94-28, Ano98o, AdA08, Bar95, BHPB06, CA07, EFV00, FJV96, GST08, GÁNA07, MCGLA02, MG08, dAOCM02, RAG11, SPÁG22, VB05, dBJ02]. spanners [SJCT24]. spanning [BC10, LPN22, ZMJM23]. spare [SSP04]. spare-part [SSP04]. sparks [Bes01b]. spatial [Ano92c, CT92, PGLS21]. SPC [ZI01]. Special [Ano03q, Ano12a, Ano93s, CY12, CMH14, GH15, HCT11, Kha98d, Kha98c, KV24, Lin09b, Lin19a, MVS^{+}22, SG93, She09, SL14c, Ano12a]. specialisation [Ver01]. specialists [JR91, GKK23]. Specialization [SH08, ADD19, PSU97, SH10a]. specialized [FL15]. specialty [CbLjJL12]. speciation [PPM24]. specific [AHSS19, Bar90a, CEY11, KHKL23, TLK09, Tsa13b]. specificity [GFE95]. spectrum [LS24]. speed [AAGL23, BSN21, CJ23, CK12b, MP09a, RAG11, THC^{+}23, Won05a, ZXDG19]. spells [ARvT04]. Spender [Ano97b]. spending [Ano93l, PBC93]. spill [Eri00]. spill-over [Eri00]. spillover [HW01, HW02, KGS^{+}24, NWGB09, SP22a, VRS^{+}22, WZGBA01, WTN^{+}02, WKN04a]. Spillovers [Dre06, BRV23, BCL22, CL21, Har97a, Har97b, KL04, NQ05, WA03]. Spin [BHPW00, KL08b, MW82, BKM10, CRKA98, CFC12, DB05, Fon05, GvBR10, NPS02, OÁVH14, PS03, Rog86, SS21, Sch86, SOKE15, SvG15, SJ16, Ste14a, TBSM21, TKT13, VP14, WD06, vGS09]. Spin-off [KL08b, MW82, CFC12, DB05, NPS02, Rog86, SvG15, VP14]. Spin-offs [BHPW00, BKM10, CRKA98, Fon05, GvBR10, OÁVH14, PS03, SS21, Sch86, SOKE15, SJ16, TBSM21, TKT13, WD06, vGS09]. spinoff [BCKC24, Pro21]. spinoffs [ABS24, CDMV24]. spiral [HWC06]. Spon [Hay93a, Hob97]. Sponsored [WD06, Ano94-43, HHH09, Mia94]. sponsorship [LbKL91, VvWM20]. sport [Lüt04]. sport-related [Lüt04]. spread [Car05]. spreadsheet [KK06]. Springer [Hay92b, Hop04, Pap02b, Rea06]. Springer-Verlag [Hay92b, Hop04, Pap02b]. spurring [WKN04a]. Srinivasan [Bas04]. St. [Kih03]. stability [Koc04, TWGB02a]. stack [Lin10c]. staff [ST21]. Stafford [Per95]. Stage [SY17, CKR00, CHT99, Cho96, Kim87, LZD24, MNL22, NPS02, WL91]. Stage-gate [SY17]. stages [AY05, CBK03, HKW^{+}22, VNP^{+}22]. stagnation [WN03]. Stakeholder [SG22, YCL22, BBC06, HBM14, MMMT12, MMM16, SR22]. stakeholder-centric [SR22]. stakeholders [SSS17, ZS04b]. stand [Lin12b]. standard [LP03]. standardisation [FHBDO16]. Standardization [XHM^{+}16, Ano16h, BNS16, HL22, JKF19, Red90, Tam16, VSC16, WZSZ16]. standardize [BM16b]. standardized [HJSYJ12, LTT23]. standards [Dan19, FL15, Gal12, Hyv06, dVV16]. Standing [JWY23, SL16]. standpoint [Cor87b]. star [LPN22]. Stargate [Mer94b]. Starkey [Bes97]. stars [Boa08, Goe10]. start [AF08, BC12, CA06, CTSJ24, DVSL19, GM23, HSLS24, HLL12, KZ18, LLZ19, MM08, MG98c, PKH15, Par05, RI04, Ruo05, SOKE15, vRE21]. start-up [DVSL19, GM23, HLL12, KZ18, MM08, MG98c, RI04, Ruo05, SOKE15, vRE21]. start-ups [AF08, BC12, CA06, CTSJ24, HSLS24, LLZ19, MM08, PKH15, Par05]. started [GS10]. starter [GBL03]. Startup [DCD22, BEE24, Bla16, DDN24, Lak08, MKL19, McM94a, OFR24]. startups [BC82, CNS06, DID21, MNL22, SMW21]. State [BB95, OTA01, Thu18, BLNP05, GWZ^{+}23, Hit89b, Hit02, Hit03, Hit05, Lin02, MLF16, PCG22, Tõn15, YHcHW17, YHcHW21, Hui11, Mac97b]. state-owned [GWZ^{+}23, PCG22]. stated [Lal00]. stategy [WKN04b]. states [Cor87c, CL88, FOL^{+}03, OOJ^{+}00, TOS^{+}00, BDL92, Miy96, Miy00, Roe87]. statistical [Llo07, NS90, Bes02b]. statistics [AL05, Lin08a, Lin10d, WTGB01]. Status [Zho92a, Zho92b, EFV02, Lin08a]. Staunton [Wil90]. Stay [GZXX23]. staying [LDK22]. steady [BLNP05]. steel [Acs88, Poz86]. steep [SK92]. steer [Joh99]. Steering [HPMI^{+}23]. Steersmanship [Kuh88]. Stefik [Aug06a]. STEM [MCA^{+}24, SG24]. Stephan [Dib00]. Stephen [Hay92e, Mil88]. Steps [Ano94-34, Den94, Hel94a, Hel94b, Hol93]. Steve [Hay01b]. Stevenage [Rus94]. STI [DS23]. still [PP02]. Stimulating [HC06, Ano94-30, BD05, LB02, SS94]. Stimulation [Gaz86, KJE96]. stochastic [Mat22]. stock [LFJA23, ZLJ^{+}21]. Stoltz [Fra00b]. stone [AAA02]. stones [YSKL21]. Storage [OO02, DF98a, DF98b]. Store [Kan22]. Storey [Hay88e, Hay90c]. stories [SZLR19]. story [Car98a]. Strait [LS05]. Straker [O’H99]. Strategic [BM05, BF06, CKR00, CR08, CNS06, DBV97, ETAO98, ETA02, Fra03b, Hay01b, Jon96, Kha97, KOX21, KM06b, LW91, LM94a, LM94b, NLLC05, OG05, OM88, Ott98b, Pra10, SB02, Szw07, Tsu02, Ver90a, VK03, AGM89, Ano93b, Ano94q, Ano94-32, Ano94-37, Ano94-45, Bar92a, Ber96, Bes94d, BVS24, BST03, Boy14, BEH99, BGM88, BN88, Car98b, Car98c, CAI00, xCZZ^{+}22, CA14, CW93, DB09, Dib02b, Dod88, EFV00, Far23, FMV00, FF92, Gar94, GLYP13, Har96, Hay93d, HW97, HT06, Hos01, How94, HKY22, Joh92g, Kar93a, Kar93b, Kod05, Kod07a, KS04a, LL94a, LL94b, LL98, LSB07, LGZ22, LL05a, MH05, NH99, Nay88, NLLC04, PR97, PG94, PMT23, PSS23, Rot88, RFB24, SRHL12, Sch92, STS88, SW91, SM01, Sun96, Ver91a, VD04, Vos05]. strategic [VLIC12, WZY23, Why02, YNS18, Yas05, YI05, YZY10, OO02, SM09, Sap92, Bes94c, Bon02, Par03a]. stratégie [Ano94s]. Strategies [FJ87a, FJ87b, GBM98a, GBM98b, Hay97a, Hob90, HHA84, II01, Par07, Sik91, Swi92a, Swi92b, WH96, AF08, Ano90o, Ano93b, Ano94-33, AR99, BWF18, Bes94c, BDL92, CL05b, CW93, CW94, CTSJ24, CGLSNLDV15, EKT^{+}22, Ern95, FMFMA16, GC98a, GC98b, GCS01, HBS07, HK02, Hay88c, Hay89a, HW18, HLC05, HCL09, HPT10, KSHP21, KvdEV01, KS14, LW10, Lin86, MV06, MPO23, NG97, OB21, PS00b, Poy97, Ric85, SPPG20, SS22, SCZ14, Ste14b, SDK24, SH02, TWGB02a, TTC06, TS00, TM06, VM12, Wal12, WZGBA01, WR95, Won04, YGM99, YH11, lS06, Kuc06]. strategische [Ano94-45]. Strategy [Ano94-35, GC93a, GC93b, HLH12, Kan94, MR05, ADD19, AO03, AAAF22, Aka98b, Aka98a, AT93, Ara08, BiKM95, BSN21, BK01, Bro98, Bul97, BA89, CEH03, CT02, CHL21, CCY05, CLW10, CCST06, CLR96, CN90, CN92, DP23, FCH01, FKT95, FF00, FJ91, FB93, Gal12, GAS96, Gar95, GBL03, HRHW16, HC01, Joh92c, KWE05, Kod06, LKCK22, LCT05, Mil02, MPM09, Obe86, OOMW03, Pea90, PSdLC02, Per96, Per97b, PV00, Pow10, PCWG04, RR11b, Rue88, Rya96, Sæt24, Sar95, SB02, SH08, SG22, Sha03, SJSP23, SJP23, SJ16, STZ03, TWGB02b, Tan95, Tan87, VNJ18, WAN03, WS14, Win87, YI05, YXH21, Sap01, Hay85a]. straw [YWL22]. straws [CKF^{+}24a]. Street [Tan08, Ano81b]. strength [Mas86]. strengthening [CE89]. stress [ZWZ^{+}23]. stressors [ABLK07]. stroke [WTAW97, WTN98b, WTN98a]. strokes [DTI^{+}23, LT05]. Structural [YW08, dVdWNF^{+}10, Ano94-35, BF05, CLZ^{+}22, HHV05, Hob92, Kan94, LLY^{+}18, MTB01, MP09b, NW23, WN03, WQZ21, WC06b, ZZL22, GD08]. structure [Ano94q, Ano94s, Ano94-45, Hon21, Hos03, HS21b, Jai23, JWY22, LL94a, LL94b, NWGB09, Wan23a, WH04c]. structured [BD05]. structures [EEY22, RS14, SS04, SS05, Whi89a, Whi90a, Whi90b, Whi90c, dVdWNF^{+}10, Roe87]. Structuring [SL14b]. Struktur [Ano94-45]. students [CEH03, Hay97c, LP06, WW04]. studies [AAA02, ABGS13, BLA06, BDNP06, CCP23, CPSS06, HB84, IIA97b, IIA97a, Kin00, Let03, NI05, PR97, Pot93, RM14, RAC84, SU11, Tre98, VBA06, VOTH05, VX97]. Study [MD03, AM10, ASI99, AA91, AF08, Ano94x, Ano94z, Ano94-38, Arc88, ABUS05, Aut94, Bad11, BC12, BDB21, BSMC23, BCH94, Bou15, BGM88, BM05, BLQS00a, BN88, CLZ23, CJ98, CG02, CA02, CEH03, CS06, CM89, CEY11, CBGvdV22, CM94, CL12, CGS^{+}24, CPXC02, CA07, Cor87a, CNPV04, DB09, Day02, Dea81, DH06, DSVM05, ET98, FRA05, FLKF^{+}06, FPP24, FOW97, Gal05, Gar87, Gar05, GÁNA07, GL10, GM00, GC10b, Gui11, GMWK23, GS10, HBS02, HV96, HS02, HJM08, HRHW16, HMFS24, Hol90, HLC05, HH08, HM11, Hua09, IS97, Jar86, JP99, JRS06, KdOL08, KI03, Ket09, KS21, KJE96, Kod03, Kod07a, Kod07b, KH06, KH89, KLV06, KJ03, KB06, KFL11, LY96, LPS03, LD05, LPXO14, LRH22, LK20, LS07, LCRL22, Leo21, hLwChLbL04, LH07]. study [LC22a, Lic11, Lin10a, LCT04, LCT05, Llo07, cLhWCH12, LH11, Lüt04, Maj22, MRGT11, MFH03, MOM08, McM94a, MG98c, Meh01, Meh88, Mil94, MMR24, Møl91, MPCRB^{+}03, NW03a, NW03b, NB09, NB11, ND15a, NPV05, OGOF22, OFR24, OOA04, PGLS21, Par05, Pat24, PPB89, PRID23, Pow10, PS00b, Poz86, PCWG04, PGB^{+}23, Pur91, RAG11, ROMH15, RFMC21, Ros82, RD05, RG85, RI04, Ruo05, RNJ^{+}23, SO02, SBF15, SC20, SCZ14, Sna90d, Soh96, Soh97a, Soh97b, Soh99, SGFS99, STZ03, SM01, Sto03, SSRO24, iSKB02, TKSM08, TB24, Tho99, TCT05, TPH15, Tit90a, TLL07, UBCF19, VGAL24, VK03, VM92, Wal07, WZNC15, WM88, XW21, YH09, Yas05, ZMPRP14, ZXX23, ZWDZ21, ZYSL21, ZS04b, dB10, vBDK21, vGS09, vdVdM11, vH10, vK11, vRGJ^{+}12]. Studying [DLKM24, KK06]. style [FB91, FB93, KFEC15, PWF07, WPF06a]. style/involvement [PWF07, WPF06a]. Styles [Hay90a, Ric95, Fox95, Lin85]. stylized [SB91]. Sub [AHSS19, AN16]. Sub-Saharan [AHSS19, AN16]. subcontracting [ADNP01, KS10b]. subcontractors [LL98]. Subject [Ano86j, Ano88l, Ano91d, Ano91g]. subsequent [PG11]. subsidiaries [CJL23, LG22, MG08, Tag97]. Subsidiary [RS14, MPWG22, OWC03, ZDWG14]. subsidies [DPH^{+}22, GGJ18, Klo87, PCQ24, WSF13]. subsidized [SPFG12]. substantial [GWZ^{+}23]. substantive [GARM24]. substitutes [CKT04]. substitution [JP99]. Success [CA06, Kak03, Leh07, Ste14a, Ano93m, Ano94n, BM23, Cai94, Car98a, CT12, DH06, DSS05, EHSB24, Gal05, GMNL07, GvBR10, GS10, Jan21, JWY23, LWC09, dLP06, MT05, MW06, Moh95, MK96, Pal06, RRSS07, Sal03, SSKB24, SM01, SDK24, SW05, SL14b, Tei93, VOTH05, ZMPRP14, ZLC18]. Successful [Hay89a, FGTB23, GBL03, GCH96, Gup96, KKRM05, MH98, Rob84, RkhR97, RI04, TWGB02a, WC06b, Sha91a]. successfully [Kap11, MPO23, Hob97]. successive [BS24]. successively [TWGB02b]. such [Cor87b]. Sufficient [Fra02c]. Suggested [Gui05, DS13]. Sumantra [Hay93d, Per97b, Ver90b]. Summary [Ano82c, Bai84, Lin12d]. Sun [Car98a]. Sung [Kap01]. Sunman [Hay87a, Hay87b, Hay87c]. Supapol [Ano94-30]. super [RCC04]. supercomputer [NW03a]. Supererogatory [JLWS23]. supplements [Pre01]. Supplier [CCL^{+}06, AD09, Hel94a, Hel94b, IKL01, Kha97, LM14, LV09, PS00a, SW21, UEGE^{+}14, WFW02]. suppliers [And21, CD00, HW18, KSHP21]. Supply [BS05, KK05, Lin17b, SMC14, Bar14, BSG^{+}24, Boy14, BCP22, BRCG12, BVHJ24, CCGK23, CH01, Eng06, FdlCU22, Gra03, Kee14, KI14, KH06, KM06b, KK06, KRA^{+}22c, KRM^{+}24, LBA14, NUSR21, NJ04, PMT23, Red14, RY14, SW21, Ven14, WH21, Wil14, XL23, Bes02c, Gra03]. supply-side [SW21]. support [AK06, ABLK07, BCPP23, BRV23, BSWC23, CS02, CHD18, Coc12, CH01, DO99, DO03, FJV96, FDN^{+}23, Gre06, GGJ22, JJS^{+}18, KP12, KT02, Lay93, Lee11, Mac97b, MCH00, MRGMBR13, MMM05, PLG24, PMP24, RATW15, Ras08, Rya90, Sik97a, Sik97b, SJ16, Ste14a, WKF24, Won13]. supported [Tit90a]. Supporting [MG98b, MG98a, AS08, Ano94-33, BMW14, CW94, CM96, EKY24, Joh08, Par05]. Supportive [Ano94-33, CW94]. surgery [ALFM23]. surrogate [Lun14]. survey [AL05, Cor87b, CL88, DLL06, EKT^{+}22, Han06, Har96, KDH02, KBS01, Mal87, NS90, PS00a, SP03, SU11, EIMS01]. surveys [AP96, ABKL08, TM03, PG09]. Survival [AF08, SBMP13, ABUS05, ABAHP^{+}23, Col97, DID21, DMW13, OOMW03, SCC17, Vel15, VOTH05, WK03, WKN04b]. sus [Ano94t]. Susan [Hay93c]. Sushil [Kaz03]. Sussex [Gri98, Hal97a, Kuc06, Per95]. sustain [Bro90]. Sustainability [ZYSL21, AS22, Boh18, CMB23, DQ06, EKK^{+}21, Ger04, HMFS24, KBRR24, LH14, MBA^{+}12, RGB23, Sæt24, Ser13, SDT^{+}22]. sustainability-oriented [CMB23]. sustainability-selected [Ger04]. Sustainable [ABG23, HK02, KBRR24, Leh07, AZ90, BSG^{+}24, BRA^{+}23, Cal24, De 85, Fah05, Gos05, GMB^{+}23, KvdEV01, LB02, LOM14, OSSC21, PMP24, SBOB23, Shu22, SFBK23, SZSR23, TS05, dGSJ^{+}21, Ano88m]. Sustained [FPP24]. Sustaining [LCRL22]. Sven [Dre06]. Sven-Åke [Dre06]. Swap [Bes01b]. Swasti [Hal97a]. Swedberg [Fra02b]. Sweden [ALvKL06, JP96]. Swedish [Lun14, PvdMB^{+}06, Sch86, SG24]. Swierczek [Ano94-30]. Swinburne [MG98c]. Swinnerton [SD89b]. Swinnerton-Dyer [SD89b]. Swiss [Bad08, Har97a, Har97b]. switch [GZXX23, ZI01]. switching [BVT21]. Switzerland [RBBB98]. sword [WGW22, XWZ23, ZSX24]. symbiosis [FF00]. Symbolic [GWZ^{+}23, GARM24]. symmetry [RV21]. symphony [TKS18]. symposium [Ano86d]. symptoms [ASOB23]. syndrome [IBPGML23]. synergetic [XW21]. synergies [Car99a]. Synergistic [Ano94-37, PG94]. synergy [RMGBA08, YK08]. syntegrity [Per95]. synthesis [KK23, KL87, OGRA23, PMP24, Sou83a]. Synthetic [Coo90, Buc91]. System [Kuh88, Ano93e, BH16, Bar97, BBS22, CC05, Chu02, DGP10, DG14, DDFS23, DGS23, DO99, FLH^{+}02, Gui05, GWM05, KKRM05, LGY05, dLP06, LKKH22, LWL^{+}22, LTH06, M’C05, MWJ23, Mar03, Meh01, MBA^{+}12, NS90, PWA11, PK06, RES^{+}24, RAATM05, SO02, SGB12, SBOB23, Sch92, Sik97a, Sik97b, SDN23, Vek99, Wal85, WLG05, YS93, ZMJM23, ZS04a, ZS04b, vdEJG08, MG08, VMWV16]. systematic [ACBO23, BLA06, BD23, CPT23, CFP^{+}23, CLGFVCM23, GLÖS23, GBY22, HS14, LAC24, MDT^{+}23, MMMD23, PMP24, RPASLM23, RC23, SS23, SW22, SMD23, WML^{+}23, XL23, XLD23, YJLL20, YJLL21, ZKW^{+}23]. systematically [Lic05, UB82]. Systematizing [WPR^{+}22]. systemic [BLLEM23, CWY14, DM11, DO03, VMWV16]. Systems [Edw04, AGM89, Ano93k, Ano94m, BF04, BK04, BWF18, Bas01, BK01, Boy14, BH03, BHPB06, CF01a, CC05, CS04, xCZZ^{+}22, CG04a, Chu01, Chu02, De 04, DO99, EEY22, Gal06a, GPCNS^{+}23, Gee06, Ger04, Gol87, Gra24, GG11, Gup88, Gup89, Gup90, GH91, GG93, Gup00, GK06, HR98a, HR98b, HGL98, IMBH98, KR12, KMID11, LG23, LM24, LQ94, LW91, MI07, Mon01, OB04, PLG24, Par00, PDHW23, Pur91, Pur92, SGB12, Sen96, SC05, Sto03, SV01, WKOW03, WL91, ZIAO^{+}21, ZL01, LVVH23, Sod08, Dib01, Que06]. systems-empirical [WKOW03]. Systems-Opening-up [LVVH23]. systems-some [OB04].

T [Ano93l, Bes02a, Buc91, Fra99, Hay97c, Leh07, Lop06, Rea06, HRHW16]. T-Up [HRHW16]. tackle [DGR23]. Taguspark [DSVM05]. Tailoring [BNV05b]. Taiwan [OH04, Ano94l, CSH94, CHT99, CTYH02, CT02, CS04, CS05, CJH^{+}07, CLC13, CSL06, Chi90, FLH^{+}02, GP04, HC01, HST05, HHH09, HLC05, HST07, Hua09, HT08, LS05, LY00, LC07, LC09, hLwChLbL04, LH07, LFFT09, Liu93, LCT04, LCT05, LH11, PJ96, PWL20, SY05, TC97, Tan95, TCT05, Arc03]. Taiwanese [FLH^{+}02, LH21, WC06a, YH09]. Takada [For90, For91]. takeoff [LT16]. Taking [HPT10, JL13, SGWL16, LFJA23]. tale [Eek84, GBM98a, GBM98b, PCG22, SZLR19]. TAM [Cha08]. Tanega [Mer94a]. Tangible [Jan21]. Tangible-Intangible [Jan21]. tango [BT22]. Tantoush [Szw07]. Tanzania [Hew99]. Tanzanian [vdBS02]. Tarek [Szw07]. target [SH22, BB95]. Targeting [CFC12, FB05]. targets [Hus10]. task [Cha08, Pap95]. tasks [Sna98]. Tassabehji [Kuc07]. Tawfik [Per96]. Tax [LL93a, LL93b, CJ23, DC22, DPH^{+}22, HST21, Kog03]. taxonomy [CGBV21, LT09, MG08, RSA05]. Taylor [Ano97b]. tea [Ano94b]. Teaching [AAD14, Mos16, OZ06, WC85, JC01]. Team [ABLK07, JWY22, RR11b, STZ03, Tik23, AAM24, BPU21, KLV06, Per95, SCS22, VP14, Mer02]. Team-based [STZ03, BPU21, Mer02]. teams [AAGL23, ALFM23, BM23, BBS22, HEC22, Hua09, KLV06, LC07, LR09, LP06, WvdWG24, Joh92e]. teamwork [MMM05]. Teare [Her92a]. Tech [Bai84, LS05, ALvKL06, Ano93s, Ano94y, Ano94-28, Ano94-29, Ara08, BMV02, BDB21, Ber96, BMRG00, BM14, BRP94, CLZ18, CLZ21, CA06, CJL22, GC10b, HAA09, HdWN21, HFWW16, Kak03, Kod07a, KLH05, LWC09, LD05, LAC24, Lin86, Liu93, MZG17, MM08, Par05, RIMP22, RGJ13, SG93, SCC17, SSGN23, Ste14b, SMW21, Sun96, VTRM12, WTJ23, WCT10, Won16, CF07, Won08]. tech-transfer [RGJ13]. Technical [AA90, Ber96, NB92, Oak03, Sur92, Acs88, Ano94-44, ABKL08, Bar96, CC05, CR81, Col97, DB97a, DB97b, Ene22, Gru93, Ham94, HB84, Hay01c, IDMD04, KC08, KR12, LKKH22, Mac97b, Mol98, PPW06, Red90, Rei97, Sae05, Sci83, SS16, SW22, Tho88c, Tre98, VB05, Whi89a, Whi90a, Whi90b, Whi90c, SBMP13]. technical/marketing [VB05]. Technics [HH89, Hit89a, Ish89a, Ish89b]. technique [KW05, KKP17, MTY02]. techniques [CP11, CM89, CZL^{+}24, DF98a, DF98b, Gup89, JK03, OSG04, OC85, Bes95c].

Techno [DR24, Mon01, BF03, JLY23, Kha99, KM97]. techno-academic [KM97]. techno-economic [BF03]. Techno-globalism [Mon01]. Techno-humane [DR24]. techno-management [Kha99]. techno-nationalism [Mon01]. techno-socio [JLY23]. technoinnovation [DGR23]. Technological [Ano93n, Ano03q, Ano11a, AK00, BCS06, Bon02, BST03, CGM22, CFP^{+}23, CAdNFP04, Can06, CPSS06, CL21, De 85, Dre04, Fab96, FMvT00, HN92a, HN92b, Hay86d, Hay98, HR07, JvZ08, JRS06, Kan22, LOH91, Lee04, LLR95, LM24, LHY04, Man07, May86, McG98, ND15b, PR96, PR97, Pau93, Pre01, Rag93a, Rag93b, RD05, SH22, Sik99, TOS^{+}00, TOA01, TEA00, Uzu01, Xie04, ZZL22, wK98, AH84, AH85, Ano89p, Ano94-31, Ano94-44, AP08, Arc88, AP96, AFL^{+}24, Bac98, Bam94, Ban03b, Ban04, Bar89a, Bat12, Bes94b, BLLEM23, BBHA01, BF06, CGM07, CvdEG13, CJ98, Car98a, Car98c, Car99b, CR00, Car00, CF01a, CA02, CS84, CL98, CQ03, CLT06, Chi81, CW06, CL22, CBK03, Col81, CD00, CGLSNLDV15, DVSL19, DDOR08, Dib04, DS13, DLKM24, Ehr95]. technological [EFV02, Eva02, FS84, FDVB10, Fig02, Fon05, FF92, Gal06b, GMNL07, GCH96, Gol82, Gra04, GP22, GP04, GG11, HKRR20, HMM14, HBM14, Ham94, HS14, HO14, dCHDAALMM07, Hay81, Hay88c, Hay92a, Hea83, HW97, HLC05, HC10, HS19, Hus10, ILR^{+}12, Ilo94, JP96, JP99, Joh92g, JL98, Kar93a, Kar93b, KJ97, KLL87, KL87, KKL19, LPMV22, LO94, LVHR10, Li09, LC82, M’C05, ML08a, MKK04, MRGMBR13, MG98c, MRC14, MC00, Mon01, Mow89, MYR^{+}21, NQ05, Oak97, ON88, OBL14, OBMOM24, PSU97, PGLS21, PP09, Pet11, PMN^{+}24, PSS00, RAG11, RAC84, RF12, RH23, SP03, SS21, SB02, Seg86, Sgo95, SP10, SCZ14, Sir91, Ski84, Sta98, SGF02, Sub05, SM12b, Szá92, Thu05, THHS22, URG^{+}24, Van85, VTRM12, Ver01, Vos84, VLIC12, Wal85, WvdWG24]. technological [hWyLbC08, WHP85, ZC19, ZWDZ21, ZHZL21, lS06, vW84, RAS15, For85, Kap03]. technologically [BS24, NGK08]. technologically-induced [BS24]. technologie [Ano94u, Ano94w]. Technologies [Coo07, ETA02, Her04, Web99, Ame95, Ano92c, Ano93a, BGA09, Ban04, Bar96, CT92, CDK^{+}23, CL97b, CL97a, CCR^{+}24, CFC12, DJ13, ETAO98, FBBBM22, FHBDO16, GT93, GGS02, GS05, Gri06, GCS01, HGB23, HS91, HN99, HMFS24, HC06, HJSYJ12, KKWM02, KRK^{+}22, KHKL23, KBC22, KS14, KJ03, KK05, KKP17, LFJA23, Lee23, LD01, LGZ22, LL08b, LvLB06, MRGTDG21, MPO23, MPCRB^{+}03, MJ24, NUSR21, NVMD22, NLC98, NP10, NBA22, ODM23, OM88, OOL06, Poz86, STA24, Sma01, SSKB24, Sun00, Tõn15, TS05, Wat90, ZYSL21, vdVMM09, Kra15, Hay88c]. technologies-on [GS05]. technologies/firm [ODM23]. Technologietransfer [Ano94g]. technologist [Whi94]. technologists [Rus94]. Technology [AMS^{+}24, ADK08, Ano93o, Ano93p, Ano03q, Ano07k, Ano07l, Ano09l, Ano11k, Ano11a, AZ90, tBL86, BS97, Bes03a, BCP11, BDL92, BC91, CDD^{+}14, Cav16, CS87, Cha99, CS00, CL97b, CL97a, CR02, Col08, Coo16, Cor87b, CN90, CN92, EW97, Eri00, FCH01, For99, FD21, GW87, Ham96, HdWN21, HLL93, HRHW16, Hoe85, HLC05, II97, Jel07, Lae04, Lal00, LACH13, LW04, LKPP07, Leh07, Lev93, hL05, Lic10, Lin09e, Lin11f, Lin13c, Lin17b, LJ01, LFU23, Lop06, LOM14, MMMT12, MLF16, Møl91, NB09, NB11, Nuv07, O’K19, Ola98a, Ola98b, Pál04, Par00, PPP98a, PPP98b, PFP06, Phi99, Pil04, PW99, Pra86, PEER23, RZ21, Roe87, RW86, Rus94, Rya96, Sap89, SWW12, SD23, SHS^{+}23, Sik97a, Sik97b, SB91, SN92, Tra00, Van10, Van85, VJ23a, VGQK^{+}10]. Technology [VS06, WZNC15, WTN^{+}02, WS14, Wes93a, Wes93b, YKW10, Zeg87, AK06, ASI99, Aka98b, Aka98a, AA91, AC09, Ali86, ABvZ23, AAN06, AJLR04, ADL07, Ano93e, Ano93j, Ano93i, Ano93s, Ano94u, Ano94g, Ano94i, Ano94l, Ano94p, Ano94x, Ano94w, Ano94-32, Ano94-34, Ano94-33, Ano94-43, Ano12a, AAZ22, AEP21, Arn87, ABUS05, ABAHP^{+}23, AVA22, Aut94, AAK^{+}24, Ayr88, AM97, BBT19, BBTW00, Bar90a, Bar92a, Bar92b, BCM^{+}22, Bas04, BV24, BE09, BB89, Bes94c, Bes02d, BG19, Bha05, BF21, BS99, BLLEM23, BC92, BRH^{+}97, BB22, BF23, BV09, BH03, Bro85, Bro01, BM14, BM07, BP01, CA11, CS08a, Can16, CK22, CRKA98, Car99a, CKR00, CT06, Car05, CWAB09, CPP09, Cha22, CL05a, CS21, CSH94, CHT99, Cha08, CY12, CMH14, CYJ07, CCGK23]. technology [Che94, CR06, CCV11, Che12, Chi90, Chi91, CLR96, CY00, CL21, CU13, CJS00, CA14, CHO98, CS89, CW94, Cou90, Cou91, CD00, Dag04, DB09, DMH22, DGP10, Dav93, DB05, De 85, Dea81, DBL07, DVdCAS16, DVSL19, Den94, DAR07, Des96, DLL06, Dod87a, Dod87b, Dre97, DR99, DSVM05, DFBC98a, DFBC98b, EH02, ETOA00, Els95, Ene22, EN07, EM02, FOL^{+}03, Fel22, FORA17, FB06, FGB^{+}24, FC95, FF00, FJ91, FV91, FG06, Fuj98, Fuk09, GKK23, GR23, Gar85, Gar96, Gei97, GEG12, GBL03, GBFvdS11, Gim06, Goe10, GKSW19, GARM24, HGB23, Hag90, HY99, HV96, HMWT23, Han94, Hay87b, Hay88e, Hay89a, Hay91a, Hay93b, HSLS24, HO11, HJM08, HOGGP11, HY04, HB03, Ho09, Hob87, Hob88, Hob92, HCK01, HPBP98, HO05, HKW^{+}22, How94, HP09]. technology [HC01, Hua09, HS19, HT08, HW01, HW02, HK86, Hyv06, II98, IAO^{+}02, INO^{+}03, JBL23, JGM98b, JGM98a, Jar86, JY15, JWS92a, JWS92b, JR91, Jol03, Jol08, JD09, KK23, KA15, Kap01, Kim87, Kim03, KS07, KJE96, KvdEV01, Kod14, KLL24, KGS^{+}24, Kos99, KL08b, KK87, KJ03, Kum05, Kum06, KW05, KKP17, LS24, LY96, Lan09, LPS03, LPXO14, Lar09, Lee00, LY00, LC07, LS07, LYP09, LVHR10, LJAK21, Lee23, LLL23, LPN22, LD95, Let03, LC09, LT21, hLwChLbL04, LL08a, LFFT09, LH21, Lin91, Lin11a, Lin06b, Lin10c, Lin12d, LW97, Llo07, LL05a, LL05b, LCWC09, Lun14, MV06, Mac97a, Mac86a, MPA05, Mal02, MRGMBR13, Mas23, Mat22, MKGL05, MG98b, MG98a, Mer94b, MW09, Mia94, Mil88, MS98, Min96, MR05, MGS99, Mol97b, Mol97a, Mos16, MPWG22, MBA^{+}12, NWGB09, NW03b]. technology [NVMD22, ND15a, NLC98, NSYL18, NC23, OSG04, OG05, OS08, OC91, Oak03, OKP13, Oye01, PNC15, PGLS21, PJ96, Pap02b, Pap03, PP04, PMP24, Paw94, Per04, PS03, Per96, PV00, Pfi99, PT06, Pow10, PdW00, PG11, Rad99, Rad01, RIMP22, Ram81, RFMC21, Rog86, RTY01, RD92, Saa98, Sae05, SC17, SCS22, SG93, SP98a, SP98b, SIOO01, SPO08, SP22a, Sch09, Sch85, SCH93a, SCH93b, Sen85, Sér96, SG22, SR22, SXS24, Sha95, SM95, SPP06, SCH23, She09, SFC14, SPLB19, Sik91, Sik94, Sik96, Se14, Sir91, SS04, SS05, SGWL16, Sko87, Sla23b, SOA^{+}24, SS07, ST21, Soh99, SCH19, SCY22, SKL17, SZSR23, Sou83a, Sou83b, Ste89, SM01, SSRO24, Suc87, Szá96, TNY^{+}22, TW02, Tan95, TB24, TTC06, TPB12, Tho99, Tho12, Tit90a]. technology [Tit90b, Tit94a, Tit94b, Töd94b, Töd94a, TE14, TCHS95, UOIA00, UF22, Ved81, VJ12, Vek99, Ven96, Ver90a, Ver91a, Ver91b, VD04, VK03, VLIC12, Wan89, WTJ23, WC07, WLY14, WZGBA01, WNGB03, WA03, WAN03, WKN04a, WA04b, WH04c, WKN04b, WLO09, Wec15, WM88, Whi89b, Wil90, Wil93, Win87, Woo87, WVG^{+}98, WGV^{+}98, WO08, YS93, Yas05, YT09, YH11, YSKL21, ZLS23, ZHY22, ZR93a, ZR93b, ZZL^{+}23, vW88, BS08, CDD^{+}11, Dre02, ES05, Lop06, LH11, MW82, SL16, Wal97, WvT23, Wya07, Dut07, Ano91c, Gra04, Hay86f, Her92b, Kaz03, Har93a, Har93b, Hay97a, Lae03, Lim06, Lin11f, Par03b, Ros90]. Technology-based [HLC05, LOM14, Ano94x, ABUS05, AVA22, Aut94, CS08a, Can16, CK22, CKR00, DVdCAS16, DVSL19, DLL06, FC95, Gar96, GARM24, HV96, HSLS24, JBL23, KJE96, KK87, Lee00, LL05a, LL05b, OSG04, Pap03, Sae05, SC17, SCS22, Ste89]. technology-driven [KW05]. Technology-enabled [AMS^{+}24, HMWT23]. technology-focused [HGB23]. Technology-Internet [Lop06]. technology-service [Ano12a, CY12, CMH14, Kod14]. technology-transaction [SCH23]. technoparks [RM09]. technopoles [Cho96, Hay87a]. Technopolises [Ano94-38, BCH94]. technopreneuress [FWL06]. technopreneurism [FF00]. Technovation [Mou17, SJSP23, YJLL21, Ano93a, Ano93b, Ano93f, Ano93e, Ano93j, Ano93i, Ano93k, Ano93l, Ano93m, Ano93o, Ano93n, Ano93p, Ano93s, Ano94c, Ano94f, Ano94k, Ano94l, Ano94m, Ano94n, Ano94z, Ano94-27, Ano94-30, Ano94-31, Ano94-34, Ano94-37, Ano94-33, Ano94-38, Ano94-42, Ano12a, BP09a, BM94b, CDD^{+}14, Hol94b, Paw94, Szá94b, YGW08, Hay89c, Ano11a, DMV^{+}21, GH15, II92b, JWS92b, KV24, Lin06a, Lin08a, MdCÁ06, Swi92b, Wet81, Zho92b]. tecnologia [Ano94i, Ano94-28]. Teece [Dib02b, Hob03]. Teichert [Ano93m]. telecare [BBC06]. telecenters [ZMPRP14]. telecentre [AY05, AY05]. telecentres [dAOCM02]. telecom [Fan06, Hos01, Whe86]. telecom-equipment [Fan06]. Telecommunication [Arr04, CT97, Ano92c, CT92, CY00, CBK03, KY06]. telecommunications [AKL05, ALRSD04, Cha94, FV91, Gru93, Hos99, Hos00, Hos01, Hos03, JP99, NDP07, Whe88]. teleconsultation [BKA^{+}23]. Telemedicine [DGR23, CCP23, FGB^{+}24, FMvT00, KS21]. telephone [McN92, Mow88, Hos01]. telephony [EF11]. telework [PSdLC02]. teleworking [PSdLCJ05]. tell [JP96]. temporary [ALFM23, KM06a]. Ten [Gil97, CbLjJL12, Gru93]. tenants [LK20]. Tendering [HN99]. tensions [Hew99, PDOL11, SOA^{+}24, TA98a, TA98b]. term [SKL17, ZLC18]. Terminology [OM18]. Terpstra [KHKL23]. territorial [CT97]. territories [VLVH23]. Tesla [HKRR20]. test [AC08, Fox86b, LL08a, PD10, WPF06a]. testbed [WK03]. Testing [AP08, For90, For91]. tests [KHKL23]. Text [SKL17, KKP17, Sma01]. text-based [Sma01]. Text-mining [SKL17]. Textile [DC00, BN06, MM02]. textile-related [BN06]. textiles [KS14]. TFT [CLT06]. TFT-LCD [CLT06]. Thai [Hos99, LPS03, PSS00, RI04, Ruo05]. Thailand [Hos00, Hos01, NI05, PR97, Swi92a, Swi92b, SN92, Won13]. Thales [ALCM14]. Thames [Whi94]. Thank [Lin11b, Lin08a]. Thanks [Lin10d, Lin12a, Lin13a]. Thayer [LC82]. theater [GARM24]. Their [Fra01a, Rot87a, AG24, ACBO23, AKBA07, ABGS13, Ano94t, Ano94j, BBTW00, BG10, BL97, Bre23, CS84, CZZM21, CKM13, CR08, Col81, Dre01, DNS^{+}03, Ern95, FCM21, Goe10, IAA94b, IAA94a, JH18, LS24, Lan98b, LTVA22, LOCSBGMH16, LC22b, Mer01, MG08, SVLYT24, SR23, SOA^{+}24, Ste89, TKT13, ZMPRP14, ZWDZ21, ZSMB16, vBDK21]. them [DFFM22]. thematic [DHT22]. themes [PT06, ZAI23]. Themselves [Fra01a]. theoretic [DLKM24, Par03b]. Theoretical [HKJ07, MGAP23, AC09, How24, LLB89, WPF06c]. theories [RH23]. Theory [Lee11, Oak97, Bad11, CWB98b, CWB98a, DDBM23, DV23, Fox86b, Gal05, GK91, Gra24, GL10, Hal97b, Hol87, JLWS23, Kaf06, Lic11, LH14, Lin10a, Lin20, LL08b, MGdG11, MAMD22, Mud95, Pap02a, SG22, SPP06, WCT10, WBM22, WO08, Yas05, ZWZ^{+}23, dB10, vdVdM11, vH10, vK11]. therapeutic [SJ21]. There [FG06]. Thermo [Pow10]. thick [MMCD23]. thickets [FR15]. thin [Ver05]. thing [VM97]. Things [BVT21, Ric85, ANI^{+}22, BBZ20, CHT22, DFFM22, KAV17, KRA^{+}22c, LDK22, MSvdK^{+}21, PRA22, SPPG20]. Think [Jel07]. thinkering [Tre98]. thinking [AB99, CLZ23, Gal06a, MMCD23, BCM^{+}22]. third [Bie91, Hay88d, SS03, Vel15]. third-party [Bie91, Vel15]. Thomas [Ano93b, Ano94-33, Fra03a, Hay88g, Hay92b, Mer94b]. Thomson [Bes97, Coo97, Hay97a, Saa98]. Thorsten [Ano93m]. those [BNV05b]. thought [Dib02a]. Thoughts [Lin17a]. Thousand [Box01, Kuc07]. threats [Obe88]. Three [Hel94a, Hel94b, Lin11a, SF98, KL08b, LDF^{+}21, OFR24, SBOB23]. threshold [RMGBA08]. throughout [DB19]. Tidd [Dre04]. tier [IKL01]. Tierney [Poy97]. ties [Pet11, SSF12, SOKE15, SHT23, SMW21, YCL22, ZQW^{+}19]. Tiger [Kap01]. tile [CA07]. TIM [Lin07b, LE07, YKW10]. Time [Kra15, MJ94, Tit94c, Tit94d, AB89, BJ05, DL10, FORA17, GH91, HH89, KK23, MY02, Mus09, PS21, PS97a, PS97b, Sch92, WA04a, MCMA09, PS00b, Joh92b, Par07]. timeliness [MJ94]. Times [Hay01b, Pal06, Lin10c, Lin11f, Pap02b]. timing [MOWGB02, SY14, SCFD18, Wan23b]. Timothy [Ros89b]. Tinkler [Tan08]. tipos [Ano94q]. titles [Ano89s, Ano90m]. tivities [New04]. TL4ED [CPSS06]. TMAP [Bes02d]. today [Lin06c]. Todd [Hay90d]. Together [AD09, FBBBM22, DR07, ØD22]. tolerance [KL05]. Tomkins [Ano81a]. Tony [Hay93d]. tool [Ano93i, BAWK12, CHP13, Dav93, JP99, KM01, KM06a, KS04b, KW05, MPA05, MW14, Pov97, RGJ13, SB02, TB24, Thu05, CHP13, FRA05]. tool-VIZCon [FRA05]. tooling [Ano94m, LQ94]. toolkit [Bes95c, Bes94d]. Tools [BRH^{+}97, Bes95c, CA14, HST05, Hu13, MM24, PMP24, PFP06, RH23, Vit90, VX97]. Top [Ano09l, FBE23, Fra01a, BBS22, Lin13b, Lin13e, SCS22, WvdWG24]. topic [JLY23, LC22b]. topics [Lin12d, Ano91c]. topological [SKTM08]. Topologies [SSF12]. Tor [Ano93k]. Torsten [Hay96a]. Total [AHAAEA03, IBL04, POGBG06, Hal97a, JW06, SVÁG07, vAdMSL95, OÁVH14]. Totemics [DF98a, DF98b]. Touchstones [Gil97]. Toulouse [Ano93a]. tourism [AC09, Hja09, JD09, Lin09b, MP09a]. tourism-Selected [Hja09]. tower [HPT10]. town [FG06]. Towne [Hit93]. toxicity [BLQS00b]. toy [SW05]. TQM [JW06, AS08, CPXC02, DB09, FKT95, PS01, PS04, PH08]. traceability [Soh97a]. Tracking [Ell81, OGS06]. Trade [Ros89b, CZM23, EL93, MSvdK^{+}21, OCA09, Rie89, TNSG96, Vit90, ZZL22, Hay01c]. trade-offs [MSvdK^{+}21]. trademark [SCFD18]. trademarks [ZSMB16]. Tradition [dC07]. traditional [Bro94, SO02, SCK10, SCK11]. Traditiovations [CP11]. tragi [McM96a]. tragi-comedy [McM96a]. trail [URG^{+}24]. training [AA91, ALM22, BBTW00, CJ98, HPMI^{+}23, OC85]. trajectories [LLEB05]. trajectory [BLLEM23, HWC06, WMGB01, WLO09]. Tran [Bai84]. transaction [SCH23, SS04, SS05, Yas05]. transaction-cost [Yas05]. transactions [Ano93i, Dav93, Wec15]. transactions-focused [Wec15]. Transactive [ZMJM23]. Transfer [Lal00, RGJ13, Sko87, ZT21, AA91, AGS98, ADL07, ABGS13, Ano93p, Ano93s, Ano94u, Ano94g, Ano94l, AM97, BBT19, BBTW00, BCH98, BC91, BC92, BB22, BP01, Car99b, CSH94, Cha99, Che94, CS00, CR02, Cor87b, Dag04, Day02, Dea81, ETOA00, ETA02, Els95, FRA05, FEP15, Fuj98, GGS02, GEG12, GCH96, GBFvdS11, GKSW19, Hag90, Ham96, HP09, HW02, KA15, KdOL08, KKWM02, LY96, LACH13, LW04, Lev93, LT21, LH07, LFFT09, LJ01, Llo07, Maj22, Mal02, Mat22, MKGL05, MGS99, Nob04, NC23, O’K19, PS03, PV00, PRID23, Pra86, Rad99, RAC84, RTY01, RD92, SG93, Sap89, SP22a, SCH93a, SCH93b, Sér96, Sha86, Sir91, ST21, SCH19, Sou83b, TNY^{+}22, Tan95, Tho99, Tit90a, Tit94a, Tit94b, TCHS95, Van85, Ven96, WL21, Wec15, Wes93a]. transfer [Wes93b, Woo87, ZR93a, ZR93b, GBFvdS11, MMMT12, DBL07, Hay93b, Saa98]. transfer-UK [DBL07]. transferencias [Ano94i]. Transferring [NP10, TPB12, ZZL22]. transfers [Ano93i, Ano94i, Dav93]. transfert [Ano94u]. Transformation [OB21, dC07, AO10, Ano94v, Bar97, BKA^{+}23, Ber96, Car98b, DMRS23, DP23, Fon05, GPCNS^{+}23, Gee06, GHHH22, GR21, Li21, Mas23, MLC24, MRDA24, Öbe23, RDAV23, SSCH23, Gri98]. transformations [PPW06]. Transformative [WT19]. Transforming [ABSZ23, Kod03, KL08b]. transition [AHP16, Fox86b, GS05, Hay97d, Kim03, Kod14, KK04, LL02, LSB07, LT21, PP02, RB09, Tho99, WKN03, Web99, Wya04, YLH22, ZGSG23]. transitioning [CvZ05]. Transitions [AB99, Boh18, EKK^{+}21, Gee06, LKKH22, PPM24, SS16]. translate [CTSJ24, SSS17]. Translated [Ano92h, Bes95b]. Translation [Ano98o]. Translations [Ano93q, Ano93r, Ano94-39, Ano94-40, Ano95r, Ano95x, Ano95o, Ano95s, Ano95p, Ano95q, Ano95t, Ano95u, Ano95v, Ano95w, Ano96o, Ano96u, Ano96p, Ano96q, Ano96r, Ano96s, Ano96t, Ano97r, Ano97q, Ano97s, Ano97t, Ano97u, Ano97v, Ano97w, Ano97x, Ano97y]. transnational [Ver90b]. transparency [TPBM23]. transport [WSF13, WH04c]. transportation [BV24, Rue86, Sna01]. trap [MJ24]. trash [QvdR21]. travel [Hja09]. traveler [MY02]. treasure [QvdR21]. treatment [BJ05, PPW06]. treble [OBMOM24]. tree [Ano94t]. trend [Chi91, HS14, LD01, MYR^{+}21]. Trends [Hop04, BS87, BJ05, CTYH02, CYJ07, Hit92b, Hit02, Hit03, Hit05, Ho09, KM01, Kum05, Li21, SB02, VSC16, Web96, vDZ21, vdVdJVdR09]. Triangle [Hay93c]. trigger [PDHW23]. triggering [SYCH23]. Triggers [Tei22, Mon06]. TripAdvisor [BGF23]. triple [JLWS23, Joh08, RGH19, CE19, LAYS19, Lin19a, LH19, ZCF19]. tripod [BSN21]. TRIZ [IPP13]. tropical [IIA97b, IIA97a]. tropics [Oye00]. trouble [Lin08b]. troubled [GM23]. true [LDK22]. trust [BHS05, DDG98, GST08, HO11, LCRL22, LL08b, YPF03, HBO14]. trusting [BKA^{+}23]. try [VM97]. Tseng [Rea06]. TTF [Cha08]. TTO [ST21]. tube [KJ97]. Tudor [Fit00]. Tunisian [HAA09]. TuoviWDM [BHPW00]. turbine [KI03]. turbines [RH21]. turbojet [Gee06]. turbulence [CvdEG13, FDVB10, HC13, SVÁG07, Tei22]. turbulent [STH06]. Turkey [Uzu01, YA11]. Turner [Tat04]. Turning [Fra00b]. TV [KJ97, WKN03, Xie04]. TVs [Tsa13b]. twelve [RAC84]. Twente [HdWN21]. twin [BK24]. twins [RUNS23]. Twiss [Rus94]. Twitter [Fel23]. Two [LZD24, Mas86, CT12, GBM98a, GBM98b, Hel94a, Hel94b, HW01, Kaz03, Let03, Lin10b, PG09, SZLR19, SM12b, WTAW97, WTN98b, YK08]. two-sided [CT12]. two-stroke [WTAW97, WTN98b]. two-way [SM12b]. type [AC07, GST08, GBV23, Lee04, Win87]. Typen [Ano94-45]. Types [EBG24, Ano93n, Ano94q, Ano94-45, AAK^{+}24, DMNR12, LL94a, LL94b, Pau93, SLA^{+}23a, SB91, SDK24, TLK09, BBTW00]. typologies [Ano94s, MDT^{+}23]. Typology [FMV00, ACA24, BKM10, DS13, Öbe23, SWW12, VB05]. Tzu [CF07].

U [Gra01, Pap06, CLP23, SH22, TMG11, dlHBR22]. U-shaped [CLP23]. U.K [Gar87]. U.K. [BCH98, DR88, Hob88, Hob90, PPB89, Sha86, Whi89a, Whi90a, Whi90b, Whi90c]. U.S. [Acs88, BC91, CJ98, Chi91, FL15, FH87, Rue88, Sap89, Sch92]. UAE [DB09]. Ughanwa [Hay89d]. ugly [CL05a]. UK [Ano81b, Ban04, BB95, Bes95b, Bes95c, Bes02b, Box01, CF07, Cre04, Dib02a, Gar95, Gil97, Gra04, Gri98, Hal97a, Hay93c, Hay95, Jel07, Lae03, Lal00, Mer94a, Mor04, Per95, Per96, Per97b, Per97a, Pov97, Que06, Rot88, Rus94, Sha95, Sha91b, Sub05, Tat04, Won08, Wya07, dC07, AA92b, AA92a, Ano93j, AK00, ABS24, BRV23, Bes94a, BLQS00a, CJS00, CS13, DB05, DBL07, Fre99, HPBP98, KMID11, MM02, PPM24, Sha93a, Sha93b, Tag97, Tom03, WR95, Wil93, Hay93c]. Ulm [Ano94-38, BCH94]. unanswered [NPG14]. uncertainties [MDT^{+}23]. uncertainty [ABvZ23, BS05, CLT06, CJ23, DNP01, Har97c, HS21a, Kar93a, Kar93b, LH07, OBL14, RO22, SRHL12, VTRM12, hWyLbC08]. unconventional [Szá05]. uncover [MDT^{+}23]. Uncovering [SDN23]. under-investment [Pen08]. undergoing [BC12]. underlying [EWE22, PDOL11]. underperformance [SYCH23]. understand [Dre01]. Understanding [AH84, BB22, CDED21, CPP09, CS21, EDM05, GCH96, GD08, MV06, Mol98, Mou15, Mou17, PCWV16, SS07, VJ23b, WvT23, YT09, Ban03b, CWBG03, EKY24, Hos01, LS23, Puj06, Wan23b, GARM24, LS24, Sha91b]. Unearthing [SO20]. unexplored [MDT^{+}23]. Unification [Chu01]. unified [GR21]. uniform [LT16]. Unilever [dC07]. unintended [SCY17]. Unintentional [HW01]. Union [Nay88, Hay92e]. unique [WKOW03]. Unit [Hed89]. United [Hay93c, Hob87, BDL92, Miy96, Miy00, Roe87, Tit90b, WHP85]. units [CR08, FLH^{+}02, FDVB10]. unity [CU13]. Universitäts [Hay96a]. universitie [GHU21, GP22, SR23]. Universities [Seg86, Bul85, CRKA98, CYM^{+}16, CZZM21, Cor87a, Cor87b, Cur89, DVMS16, EKY24, FGPV24, GC10a, KL08b, LfSS19, MLC24, Miy00, Nar89, NC23, OIOA01, OM89, OOA04, SCH23, SD89a, TMG11, VJ12, Wal97, ZCF19]. University [Ano93s, Ano94-42, Aug09, BBT19, Bes02a, Bes03b, BP09a, BP09b, BO81, Buc91, But96, Coo16, Cre04, Dib02b, Edw04, Fel94, Fra02b, Gra01, HdWN21, Hay87d, Hay88g, Hay01a, Hay01c, HL24, Hob03, Jel07, Kap01, LH19, MW82, Min96, Nuv07, Pál04, Pap04, Pat24, Poy97, RV17, Ros90, SG93, Sha95, SJ21, Sod08, Tan08, Wal97, Wan04, WL21, dC07, ADL07, ABGS13, Ano94-43, AC07, BCKC24, Bal10, BKM10, BPU21, BSWC23, Boa08, BC91, BB22, Bro85, Cai23, CLP23, CJ98, CAI00, CKS22, Che94, CYHC19, CKF24b, Col81, DB05, Dea81, FGPV24, Fuj98, GC10a, Hay81, HBO14, HP09, JH18, KK23, KA15, KHKL23, KJE96, LW04, LLL23, LT21, LWM24, Mac97a, MCD06, MKGL05, MMM16, Mia94, MR19, NL11, NC23, NI05, O’K19, ØD22, OIN96]. university [PS03, Pet11, PDOL11, Pra99, PG11, RMG06, Ras08, Rog86, SM97, Sla23b, SJ16, SCY22, Ste14a, TPB12, TBSM21, Tit90a, Ved97, VGQK^{+}10, VP14, WW04, Wec15, WM88, WWH15, YHDR08, YLB06, ZT21, vBDK21, ADDS09, CHO98, DBL07, MCD06, MM08, MMMT12, MG98c, OÁVH14, Pro21, ST95, SvG15]. university-based [SJ16, YHDR08, VP14, CHO98]. University-industry [BO81, RV17, CYHC19, JH18, ØD22, ADDS09, ST95]. University-related [Ano94-42, Fel94]. university-spin-offs [Ste14a]. university-sponsored [Ano94-43, Mia94]. university/small [Dea81]. unleashing [Her91]. unlocking [LDT^{+}23]. Unpacking [ALM22, LG23]. Unraveling [BCV21, NHV05, Ran24, SOA^{+}24]. unrealized [THBW22]. unserved [LHI13]. Unsolicited [Lin07a]. unsung [Lin09d]. unsustainable [MHKZ24].

Untangling [DMW13]. unused [Gri06]. Unveiling [MAMRRR24, DV23, YMMYK24]. updated [Tho12]. Updating [Lin13e]. upgrading [HRHW16, HC01, OH04, YNS18]. upon [LbKL91, NKL02]. Upper [KT02]. ups [AF08, BC12, CA06, CTSJ24, HSLS24, LLZ19, MM08, PKH15, Par05]. Upstart [Lak08]. Upstarts [Bes01a]. upstream [DLTQ23, KET22, LPN22, Thu05]. urgency [Rea21]. Urs [Hoc89]. US$22.95 [Hay01c]. US$24.95 [Bes96, Hay01a, Min96]. US$26.95 [Kra15]. US$27 [Mer01]. US$32.50 [Hay88f]. US$34.95 [Min94]. US$50.00 [Bai84]. US$64.95 [Hay01c]. US$65.00 [Bai84]. US$65/£ [Dre04]. US$69.95 [Hay01a]. US$90 [Mor04]. US-Europe [XR08]. USA [Ano07k, Bai84, Bes95a, Bes03a, Giu09, Hay88f, Hay93a, Hay95, Hay97a, Lal00, Mer01, Que06, Won08, Wya07, Ano94f, Arc88, AC04, DBL07, Dib02a, Hal97a, Mac86b, Mar01, vAdMSL95]. usage [HM11, OOA04]. Use [MGS99, ALFM23, ABvZ23, Ano88m, BBLW24, Bar15, BJ05, CB19, CAdNFP04, EN07, FB91, FB93, Gri06, HO11, Kin00, KK04, LPXO14, MM08, MPO23, NDP07, Neu12, ND15b, OC85, Sæt24, SS03, SK11, Wat90, llo91]. use-innovativeness [FB93]. used [Lan98b, RPASLM23, vWGdG21]. Useful [Tuz05, HPT10]. User [Bie91, CL12, Hol85, Hol87, TAC96, FJ87b, Hay85b, Hol88b, Hol94a, Hol94b, HLNP21, IMIY87, MGdG11, RG85, SL14a, SA22, SDK24, TPH15, TA98a, TA98b, ZLC18, BAR16]. user-centered [SL14a]. user-initiated [FJ87b]. User-oriented [Hol85, Hol87, Hay85b]. userness [GF21]. users [GBY22, GF21, GFvT22, Lae03, Lüt04, San07]. Using [AGFdMCM24, Ano94-44, BF99, Cor87c, FB06, Gre06, Ham94, HV96, Joh99, LAYS19, Why05, Aab09, AAD14, Bar14, BSMC23, BGM88, CDED21, Cav16, Gal06a, GLYP13, GCS01, HS14, IMBH98, KKL19, KHKL23, KK06, Lan09, Lar09, LS07, LJAK21, LFU23, MM24, MRGT11, MAMD22, NLC98, Ott97, SS21, SO06, Sna90d, Sun00, VRS^{+}22, YJLL20, YJLL21, KMF09]. USOs [RGFLRP21]. utilisation [TOA01]. utilitarian [TLL07]. utilities [CK04]. Utility [DLL06, Meh88]. Utilization [DA01, Ano86d, BCPP23, DAOO02, KH06, OIOA01, RUNS23, WCZZ14]. Utterback [Bes96]. Uwe [Ano94-38].

vaccine [BLLEM23, LG14, SJSP23, SJP23]. Valais [SSM09]. validated [Cal24]. Validation [CS12, JBGMM11, Car00, CS08b, MP09b, Tik23, ZCHC23]. validity [OFR24]. valley [EWE22, BCV21, BC82, DZX22, LCRL22, De 23, LG22, OWCvdB15, RO22, SCY22, iSKB02, TNY^{+}22, Won06]. valorisation [MPCRB^{+}03]. valuable [Wan24]. valuation [BCS06, CEY11, KHKL23, PP04, VGQK^{+}10, vRE21, vZS14]. Value [BBZ20, Bro95, FC06, PGB^{+}23, SL14c, Why05, Ano93o, AAZ22, BGF23, CCB20, CNN23, CLZ18, CLZ21, FMF21, Far23, Fon05, FPP24, FNJV14, Gra24, Haa14, HLL93, HLH12, JMT03, JSP22, KY06, LDB13, MWJ23, MZG17, MY02, Mus09, NG97, NLLC05, PP04, PSG22, PE22, RM14, RUNS23, RHL09, RKFR24, Rya96, SL14a, Soh97b, SMI23, SDN23, TLL07, UBCF19, WJK15, YJLL20, YJLL21, CS13, MLG12, Joh92c]. value-added [JMT03, Soh97b]. value-chain-complexity [Mus09]. Value-for-money [CS13]. values [CP23, PvdMB^{+}06, DGMR21]. Valuing [SLJ13, Tra00]. VAM [SO02]. variable [PS97a, PS97b]. variables [JMPR01, NMN11]. Varian [Soh97b, STZ03, TSH02]. variance [BCKC24, WNGB03, WO08]. Variation [YHL22, JB19]. varied [AAK^{+}24]. varieties [FLC17]. Variety [BF86, RDS17, AS18, BMW14, KWE05, Man07]. Various [Gup90]. VAT [But96]. Vaughan [Min96]. vegetable [SYK94]. Vehicle [SPS01, McG98]. vehicles [BF03, BK00, BK01, Boh18, MWC22, MW09, Sap89]. veiled [PvdMB^{+}06]. velocity [VK03]. vendors [ZWDZ21]. Venezuela [CM89]. Venture [ECB22, PJ96, TB84, AAAF22, Ano93j, ABUS05, Clo91, Cor05, CP06, FOW97, GG05, GZXX23, JL96, Jon05, Kha87, KGK07, LP06, LDB13, Pan98, SAY09, TPFD13, VM97, Wan89, Wan91, Wan23a, Wil93, Won07, ZSMB16]. Ventures [Col08, ALCLdF^{+}15, DB05, DSPS23, Dod87b, Ene22, GS10, Guy96b, HSLL22, Kak03, Lee00, LC09, Lun14, MLG12, MPBC21, OFR24, Ove15, RSA05, SC10, SP22b, SCC17, SD04, TI02, THC^{+}23]. Venturing [Mer22, Ano93s, DMP05, ES21, LCWC09, McM96a, SG93, TA10]. venue [ZWD09]. Verarbeitung [Ano94j]. Verbesserung [Ano94-41]. Verifying [SS05]. Verlag [Hay92b, Hay96a, Hop04, Lim06, Pap02b, Ros89a]. Vermont [Hay95]. versus [ADD19, BVT21, Car00, DVHL22, FMF21, HT12, Hus10, Let03, Lin18, MY02, MT05, MJ24, Ram23, SRLR23, Ste14a, VM97, WTGB01, dHOvdKS16]. Verteidigungsindustrie [Ano94-41]. Vertical [LRH22, LT10, ZKW^{+}23, EW06, FL15]. vertically [FL15]. vertically-specialized [FL15]. Veryzer [Bes01a]. vesicular [SO02]. vi [Hay93b]. via [WC06a, ZZL22]. viability [BFM88]. Vidar [Gra01]. video [CUF13, LRH22, LP03, OC85, SGWL16, SCM11]. Vienna [TKS18]. Vietnam [NHT23]. View [XX21, And21, CCC^{+}12, xCZZ^{+}22, CW22, FM24, Fun06, KBSB22, Kes03, Kod06, Min95, Phi99, SR23, SG22, SSP04, SN92, TKT13, VvWM20, ZIB23, ZL01, Fra02b]. view-implications [SSP04]. viewpoint [Rie89, SD89a]. views [Jol08, Ros89c]. viii [Kaz03, Ros90]. Virtual [CA01, DSVM05, FM07, KKG07, Mor04, CvZ05, CS21, CL12, FRWT06, KMF09, KLV06, LL02, NHS24, TAM11, WF10, ZLC18, Wya00]. Virtuelle [KC08]. Virtuous [WAN03, WWM00, WA04a]. visibility [PRID23]. vision [Joh99, RM09, RR11b]. visions [SM12a]. visual [Lin11d]. Visualising [MTU23]. viticulture [Chi05]. Vivien [Lae03]. vivo [JFCFMC^{+}12]. VIZCon [FRA05]. void [Mrk17, PKH15]. vol [Ano92e, Ano92f, Ano93g, Ano93h, Ano94d, Ano94o, Ano95a, Ano95n, Ano96n, Ano97p, Lae04]. volatility [CEY11]. Volterra [WKN03]. Volume [Ano04q, Ano82a, Ano84b, Ano85b, Ano86a, Ano87a, Ano87b, Ano88b, Ano88c, Ano92b, Ano93c, Ano94e, Ano95l, Ano96l, Ano97n, Ano01c, Ano03c, FNJV14, Her92c, Ano02b, Ano06a]. volumes [Dib00, Dib01]. Volvo [BDW06, Eri00, YNS18]. Vorstman [Bes95b]. vs [CR00, CS01a, CSL23, Jol08, LRH22, MLG12, OÁVH14, SBF15, Szá96, UR18, WvT23]. vu [Mow90]. vulnerable [LL22].

W [Ada07, Ano93f, Bes02a, For90, For91, Hay97b, Joh92a, Kra15, Pap03]. wake [AT93]. Walsh [Lae03]. Walter [Bes01b, Hay83, Hay86f]. Walton [Hay88a]. Wan [OH04]. Wan-wen [OH04]. want [MAB^{+}23, vWvRN17]. war [CZM23]. warmer [Gri06]. warp [PPB89]. Warren [Joh92a]. Waste [Tho99, Bes03a, CS13, RES^{+}24, SDT^{+}22, SYK94]. Wasteland [Wal12]. wastes [DA01]. wastewater [BLQS00a, PPW06]. watching [MCA^{+}24]. Water [AOO02, Ano88m, Lun88, Sto03, BLQS00a, BJ05, CS08a, CJS00, GM23, HKL24, Sto99]. watershed [Szá94a, Szá94b]. waves [GS05, WK03]. Way [Rus94, CAP23, EKK^{+}21, ES21, Guy95, Haa14, HS14, Hit85, HWC06, MAU07, SM12b, TTC06]. Wayne [Ano88h]. ways [Ugh88]. WC2A [Hay93b]. WC2B [Ano81b]. weak [Bar15]. wealth [Har93a]. weapon [Sch92]. weapons [BEH99]. weather [EKT^{+}22]. weaving [CJL23]. Web [Tat04, LAYS19, LCLC03, AM10]. web-based [LCLC03]. Website [WPF06a, WPF06c, WPF06b, Pre01, SSM09]. websites [BNV05b]. Webster [Tra00]. Weidenbaum [Joh92a]. weightings [Jol03]. Welcoming [Rub89b]. welfare [KK13, Kod00]. wellbeing [AG24]. Wellcome [BN88]. Wen [Ano94l, OH04]. West [Gri98, Per95, DNS^{+}03, JRS06, Kin00, Hay87c]. Western [AOO02, CS00, Phi99]. Westview [For85, Hob87]. Wettbewerbsfähige [Ano94-45]. Wheatsheaf [Gar87]. Where [JL24, JL13, KFL11, Lin12b]. Which [ADD19, Gal05, GEG12, Men11, ØD22, JSR19, SdO^{+}15, Kra22b, SdO^{+}15]. while [LDK22, Tsa13b]. Whiston [Hay92b]. White [Hay87d, Ano82c, KJ97, LM14]. Whittington [Gar95]. Who [Fra01b, Wya04, CP06, HGB23, CP06]. whole [LWM24]. whose [Hew99]. wider [SAWT06]. Wikipedia [KKL19, LS16]. wikis [SK11]. Wiley [Bes94b, Bes94c, Del97, Fra99, Fra00b, Fra02d, Gil97, Gri98, Hay85d, Hay85b, Hay86d, Hay92a, Her92c, Joh92g, Mil85, Mor91, Mur91, Per95, Pov97, Pre01, Sco89, Ver91b]. Will [Ano93b, Ano94-33, BK24]. William [Ano94-30, Gil97]. Williams [Poy97]. Williamson [Hay93d]. willingness [NHV05, SLJ13]. Wilson [Ano93j]. win [RBBB98]. win-win [RBBB98]. wind [OB20, RH21]. winner [Ugh88, vAdMSC94a, vAdMSC94b]. Winslow [Ano97b]. wired [She05]. wireless [ADK08, CA01, CT12]. wisdom [MW14]. Wissenschaftsstadt [Ano94-38, BCH94]. within [AM00, BBDM01, BDB21, BF06, Car04, DNT98, EKMS12, FRWT06, Fre99, Gui05, GG11, Hal96, LP03, LBG14, MAMRRR24, MFH03, OWC03, Pis89, RS14, RO22, RGH19, SM97, Sna90a, Sna90b, Sna90c, Sna90d, TKYM24, TPB12, Whi90b, Wit85, ZS04b]. without [TWGB02a, YWL22]. witness [Whi94]. Wm [Ano93e]. Woburn [Sap01]. Wolbold [BB95]. Wolseley [BB95, Bes94a, Bes95c]. women [Kan22, KM22]. Wong [Arr04]. Wonglimpiyarat [Kuc06]. wood [DA01]. Woodside [Gof07]. Wool [Ver91a]. Woot [Ver90a]. word2vec [HKW^{+}22, LJAK21]. Words [Lin06a, Lin10b]. words-Some [Lin10b]. Work [CSL23, Fit00, KH89, Kra15, Ano94-33, xCZZ^{+}22, CW94, ETM96, Har93b, KWvdS21, LS23, LP07, New13, YMMYK24, Mil88]. workbook [Hay85c, Hay88d, Bes02c]. Worker [HK86, Suc87, ZAI23]. workers [CMP^{+}06, Har96, LT05, Mig19]. Working [RNFG24, BSWC23, BP09a, BP09b, Moh95, SBF15]. workplace [Gil96, Orr97, ZAI23, Hay88a]. workstation [JWY22]. World [Fra01a, Lop06, SJSP23, Ano97b, BFO88a, BA89, CL12, Coz06, Fis89, Hal97a, Hay90d, HPT10, KMF09, ND15a, Poz86, SW04, SIOO01, SJP23, Szá92, Tom90, VM92, Wil14, ZLC18]. world-first [Coz06]. worlds [KMF09, Mas86, VD04]. worldwide [KBC22]. would [Rei97]. would-be [Rei97]. Writings [Dib00]. Wyatt [Gar87]. Wynarczyk [Hay88e].

X [Bes97, For91, Fra02b, Fra03a, Hay97c, Hol93, Pre01, Hal97a, Hay87d, Oak97, Del97, For90, HKRR20]. XI [Hay01b, Hay86e, Lop06]. xii [Hay95]. xiii [Hay97a]. xiv [Bes88, Del97]. xv [Hay88e, Hay90a]. xvi [Hay88f, Hay97c, Hay01a, Hay01c, Per04]. xvii [Hay88a, Hay90c, Mil88, dC07]. xviii [Ros89c, Ros89b]. xx [Ano97b, For85]. xxi [Hay86d, Hay88c, Hay96a, Ram86]. xxiv [Hay88b].

Y. [Lop06]. yachting [Lan09]. Yamagata [TKSM08]. Yap [Ano93e]. Yash [Ano94-37, Ano93k]. Yassin [Hay92a]. year [NWGB09, OFR24]. Years [MdCÁ06, DMV^{+}21, FBE23, Hit93, Hol99, KV14, TB24, Wit85]. York [Ano91d, Ano91c, Col97, Fra00b, Fra02d, Gar94, Gil97, Giu09, Hay84, Hay90a, Hay90b, Hay90c, Hay90d, Hay91a, Hay92c, Jel07, Joh92b, Joh92c, Koc09, Kra15, Lim06, Mac97b, Mur91, Nuv07, Pap06, Pre01, Ros90, Why05, dC07]. Young [Mat22, AVA22, BT22, KV14, Lin10b, Nwo97, Sae05]. Yourself [FBBBM22, LDK22]. Youth [BM16a]. Yue [Won08]. Yugo [Pal93a, Pal93b]. Yves [Ano93a]. Yves-Chantal [Ano93a].

Z [Bes95c]. Zack [Sap01]. Zafar [Kaz03]. Zealand [AF08, FKT95, Mar03]. Zedtwitz [Hop04]. Zegveld [Hay83, Hay86f]. Zhang [Won08]. Zhangjiang [LS05]. Zimbabwe [CW06]. Zundel [Par07]. Zusammenarbeit [Ano94g].

References

Acs:1988:IFS

[AA88] Zoltan J. Acs and David B. Audretsch. Innovation and firm size in manufacturing. Technovation, 7(3):197–210, July 1988. CODEN ???? ISSN 0166-4972 (print), 1879-

                 2383 (electronic). URL http://www.sciencedirect.com/
                 science/article/pii/016649728890020X.

Adjebeng-Asem:1990:TES

[AA90] Selina Adjebeng-Asem. Technical entrepreneurship and socio-economic development: the Nigerian experience. Technovation, 10(4):213–234, June 1990. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http:

//www.sciencedirect.com/science/article/pii/016649729090054N.

Al-Ali:1991:RTE

[AA91] Salahaldeen Al-Ali. The role of training and education in technology transfer: a case study of Kuwait. Technovation, 11(3):133–142, April 1991. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497291900308.

Al-Ali:1992:PICb

[AA92a] Salahaldeen Al-Ali. Polytechnic and industrial collaboration in the UK. Technovation, 12(7):490–491, November 1992. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729290056N.

Al-Ali:1992:PICa

[AA92b] Salahaldeen Al-Ali. Polytechnics and industrial collaboration in the UK. Technovation, 12(7):459–464, November 1992. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729290051I.

Adeoti:2005:BRP

[AA05] John Adeoti and Adetola Adeoti. Biotechnology R&D partnership for industrial innovation in Nigeria. Technovation, 25(4):349–365, April 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497203001512.

Adebimpe:2002:EES

[AAA02] R. A. Adebimpe, B. A. Arogundade, and O. Adeoti. Engineering economy studies on the production of dimension stone from marble in Nigeria. Technovation, 22(4):259–265, April 2002. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497200000900.

Adomako:2023:BPO

[AAAA23] Samuel Adomako, Joseph Amankwah-Amoah, and Mujtaba Ahsan. Base of the pyramid orientation, imitation orientation and new product performance in an emerging market. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222001614.

Adomako:2022:HCR

[AAAF22] Samuel Adomako, Joseph Amankwah-Amoah, and Kwabena Frimpong. Human capital, reverse engineering and new venture growth: the moderating role of competitive strategy. Technovation, 114(??):??, June 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497222000670.

Aaboen:2009:EIU

[Aab09] Lise Aaboen. Explaining incubators using firm analogy. Technovation, 29(10):657–670, October 2009. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209000613.

Adams:2023:FBR

[AABY^{+}23] Kweku Adams, Rexford Attah-Boakye, Honglan Yu, Jeaneth Johansson, and Eric Tchouamou Njoya. Female board representation and coupled open innovation: Evidence from emerging market multinational enterprises. Technovation, 124(??):??, June 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000603.

Al-Atabi:2014:TEU

[AAD14]  Mushtak Al-Atabi and Jennifer DeBoer.  Teaching

entrepreneurship using Massive Open Online Course (MOOC). Technovation, 34(4):261–264, April 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214000078.

Acikgoz:2023:RBI

[AAGL23] Atif Açikgöz, Fulya Acikgoz, Ayse Günsel, and Gary P. Latham. The relationship between imperfect information and performance speed: the mediation of improvisation in new product development teams. Technovation, 121(??): ??, March 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497222001869.

Ayamga:2024:AOR

[AAK^{+}24] Matthew Ayamga, Maria Carmela Annosi, Ayalew Kassahun, Wilfred Dolfsma, and Bedir Tekinerdogan. Adaptive organizational responses to varied types of failures: Empirical insights from technology providers in Ghana. Technovation, 129(??):??, January 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001980.

Annosi:2024:ICA

[AAM24] Maria Carmela Annosi, Francesco Paolo Appio, and Antonella Martini. Institutional context and agile team innovation: a sensemaking approach to collective knowledge creation. Technovation, 129(??):??, January 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223002055.

Amir-Aslani:2006:TIS

[AAN06] Arsia Amir-Aslani and Syoum Negassi. Is technology integration the solution to biotechnology’s low research and development productivity? Technovation, 26(5–6):573–582, May/June 2006. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649720500088X.

Abbassi:2014:SBP

[AAT14] Mohammad Abbassi, Maryam Ashrafi, and Ebrahim Sharifi Tashnizi. Selecting balanced portfolios of R&D projects with interdependencies: a Cross-Entropy based methodology. Technovation, 34(1):54–63, January 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213001041.

Anzolin:2022:WDR

[AAZ22] Guendalina Anzolin, Antonio Andreoni, and Antonello Zanfei. What is driving robotisation in the automotive value chain? Empirical evidence on the role of FDIs and domestic capabilities in technology adoption. Technovation, 115(??): ??, July 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000232.

Arnold:1989:JTS

[AB89] Ulli Arnold and Kenneth N. Bernard. Just-in-time: Some marketing issues raised by a popular concept in production and distribution. Technovation, 9(5):401–430, August 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900138.

Acha:1999:TTC

[AB99] Virginia Acha and Katalin Balazs. Transitions in thinking: changing the mindsets of policy makers about innovation. Technovation, 19(6–7):345–353, June 1999. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299000322.

Audretsch:2023:LOI

[AB23] B. David Audretsch and Maksim Belitski. The limits to open innovation and its impact on innovation performance. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497222000669.

Attah-Boakye:2023:RRO

[ABAHP^{+}23] Rexford Attah-Boakye, Kweku Adams, Elvis Hernandez-Perdomo, Honglan Yu, and Jeaneth Johansson. Resource re-orchestration and firm survival in crisis periods: the role of business models of technology MNEs during COVID-19. Technovation, 125(??):??, July 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000809.

Audretsch:2023:SOM

[ABG23] David Bruce Audretsch, Maksim Belitski, and Maribel Guerrero. Sustainable orientation management and institutional quality: Looking into European entrepreneurial innovation ecosystems. Technovation, 124(??):??, June 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000536.

Ankrah:2013:ABU

[ABGS13] S. N. Ankrah, T. F. Burgess, P. Grimshaw, and N. E. Shaw. Asking both university and industry actors about their engagement in knowledge transfer: What single-group studies of motives omit. Technovation, 33(2–3):50–65, February/ March 2013. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497212001228.

Armbruster:2008:OIC

[ABKL08] Heidi Armbruster, Andrea Bikfalvi, Steffen Kinkel, and Gunter Lay. Organizational innovation: the challenge of measuring non-technical innovation in large-scale surveys. Technovation, 28(10):644–657, October 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208000394.

Akgun:2007:TSM

[ABLK07] Ali E. Akgün, John C. Byrne, Gary S. Lynn, and Halit Keskin. Team stressors, management support, and project and process outcomes in new product development projects. Technovation, 27(10):628–639, October 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000272.

Audretsch:2024:IAA

[ABS24] David B. Audretsch, Maksim Belitski, and Deepa Scarra. Intrapreneurship activity and access to finance in natural science: Evidence from the UK academic spinoffs. Technovation, 129(??):??, January 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497223001992.

Aloini:2023:THE

[ABSZ23] Davide Aloini, Elisabetta Benevento, Alessandro Stefanini, and Pierluigi Zerbino. Transforming healthcare ecosystems through blockchain: Opportunities and capabilities for business process innovation. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222001043.

Aspelund:2005:IRI

[ABUS05] Arild Aspelund, Terje Berg-Utby, and Rune Skjevdal. Initial resources’ influence on new venture survival: a longitudinal study of new technology-based firms. Technovation, 25(11):1337–1347, November 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204000963.

Ameye:2023:HUS

[ABvZ23] Nicolas Ameye, Jacques Bughin, and Nicolas van Zeebroeck. How uncertainty shapes herding in the corporate use of artificial intelligence technology. Technovation, 127(??):??, September 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497223001578.

Archibugi:2004:IPK

[AC04] Daniele Archibugi and Alberto Coco. International partnerships for knowledge in business and academia: a comparison between Europe and the USA. Technovation, 24(7): 517–528, July 2004. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649720300141X.

Azagra-Caro:2007:WTF

[AC07] Joaquín M. Azagra-Caro. What type of faculty member interacts with what type of firm? Some reasons for the delocalisation of university–industry interaction.

Technovation, 27(11):704–715, November 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000636.

Alegre:2008:AIO

[AC08] Joaquín Alegre and Ricardo Chiva. Assessing the impact of organizational learning capability on product innovation performance: an empirical test. Technovation, 28(6):315– 326, June 2008. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497207001320.

Alford:2009:ITT

[AC09] Philip Alford and Steve Clarke. Information technology and tourism a theoretical critique. Technovation, 29(9):580–587, September 2009. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497209000753.

Appio:2024:AIF

[ACA24] Francesco Paolo Appio, Francesca Capo, and Maria Carmela Annosi. Not all (innovation) failures are created equal: a typology of companies’ responses to the consequences of innovation failure. Technovation, 130(??):??, February 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223002481.

Acciarini:2023:HCO

[ACBO23] Chiara Acciarini, Francesco Cappa, Paolo Boccardelli, and Raffaele Oriani. How can organizations leverage big data to innovate their business models? A systematic literature review. Technovation, 123(??):??, May 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722300024X.

Acs:1988:ITC

[Acs88] Zoltan J. Acs. Innovation and technical change in the U.S. steel industry. Technovation, 7(3):181–195, July 1988. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900193.

Anbari:2008:PPR

[ACV08] Frank T. Anbari, Elias G. Carayannis, and Robert James Voetsch. Post-project reviews as a key project management competence. Technovation, 28(10):633–643, October 2008. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207001514.

Andersen:2009:TWS

[AD09] Poul Houman Andersen and Ina Drejer. Together we share? Competitive and collaborative supplier interests in product development. Technovation, 29(10):690–703, October 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208001545.

Adams:2007:BRC

[Ada07] R. Adams. Book review: Communication of Innovations: a Journey with Ev Rogers, Singhal, A., Dearing, J. W. (Eds.), Sage Publications India Pvt Ltd, B-42, Panscheel Enclave, New Delhi 110 017, India (2006). Paperback, 260 pp. (including index) £14.99, ISBN 0-7619-3477-4. Technovation, 27(12):797–798, December 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720700017X.

Arranz:2008:CPR

[AdA08] Nieves Arranz and J. Carlos Fdez. de Arroyabe. The choice of partners in R&D cooperation: an empirical analysis of Spanish firms. Technovation, 28(1–2):88–100, January/ February 2008. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497207000958.

Abramo:2019:DVS

[ADD19] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. Diversification versus specialization in scientific research: Which strategy pays off? Technovation, 82–83 (??):51–57, April/May 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497217303462.

Abramo:2009:UIC

[ADDS09] Giovanni Abramo, Ciriaco Andrea D’Angelo, Flavia Di Costa, and Marco Solazzi. University-industry collaboration in Italy: a bibliometric examination. Technovation, 29 (6–7):498–507, June/July 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497208001405.

Anderson:2008:TFW

[ADK08] Timothy R. Anderson, Tugrul U. Daim, and Jisun Kim. Technology forecasting for wireless communication. Technovation, 28(9):602–614, September 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208000035.

Anderson:2007:MEU

[ADL07] Timothy R. Anderson, Tugrul U. Daim, and François F. Lavoie. Measuring the efficiency of university technology transfer. Technovation, 27(5):306–318, May 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497206001076.

Amesse:2001:IPE

[ADNP01] Fernand Amesse, Liliana Dragoste, Jean Nollet, and Silvia Ponce. Issues on partnering: evidences from subcontracting in aeronautics. Technovation, 21(9):559–569, September 2001. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497200000754.

Ardito:2021:IBT

[AEP21] Lorenzo Ardito, Holger Ernst, and Antonio Messeni Petruzzelli. The interplay between technology characteristics, R&D internationalisation, and new product introduction: Empirical evidence from the energy conservation sector. Technovation, ??(??):??, ???? 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497217307654.

Archibugi:1995:CFS

[AES95] Daniele Archibugi, Rinaldo Evangelista, and Roberto Simonetti. Concentration, firm size and innovation: evidence from innovation costs. Technovation, 15(3):153–163, April 1995. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497295966162.

Almeida:2008:SSC

[AF08] Shamika Almeida and Mario Fernando. Survival strategies and characteristics of start-ups: an empirical study from the New Zealand IT industry. Technovation, 28(3):161– 169, March 2008. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497207000600.

Avornyo:2024:ROM

[AFL^{+}24] Philip Avornyo, Ziyun Feng, Li Liu, Evans Asante Boadi, John Coffie Azamela, and Christian Narh Opata. The role of OSL in mobile banking application discontinuance: a technological innovation dilemma. Technovation, 131(??): ??, March 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223002572.

Abreu:2024:IEA

[AG24] Maria Abreu and Vadim Grinevich. Intrapreneurial ecosystems in academia and their overlooked outputs: Graduate employability and wellbeing. Technovation, 133(??):??, May 2024. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497224000464.

Abril:2024:UGO

[AGFdMCM24] Carmen Abril, Elena M. Gimenez-Fernandez, and María del Mar Camacho-Miñano. Using gamification to overcome innovation process challenges: a literature review and future agenda. Technovation, 133(??):??, May 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000701.

Amann:2022:MIH

[AGFE22] Marie Amann, Gabriel Granström, Johan Frishammar, and Jenny Elfsberg. Mitigating not-invented-here and not-sold-here problems: the role of corporate innovation hubs. Technovation, 111(??):??, March 2022. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001589.

Amoako-Gyampah:1989:AFM

[AGM89] Kwasi Amoako-Gyampah and Mary Jo Maffei. The adoption of flexible manufacturing systems: strategic considerations. Technovation, 9(6):479–491, October 1989. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649728990028X.

Agnelli:1989:UA

[Agn89] Umberto Agnelli. Umberto Agnelli. Technovation, 9(2–3): 251–254, June 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497289900643.

Albino:1998:KTI

[AGS98] Vito Albino, A. Claudio Garavelli, and Giovanni Schiuma. Knowledge transfer and inter-firm relationships in industrial districts: the role of the leader firm. Technovation, 19(1):53–63, November 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497298000789.

Albino:2001:MMK

[AGS01] V. Albino, A. C. Garavelli, and G. Schiuma. A metric for measuring knowledge codification in organisation learning. Technovation, 21(7):413–422, July 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000584.

Abraham:1984:UDT

[AH84] S. C. S. Abraham and G. Hayward. Understanding discontinuance: Towards a more realistic model of technological innovation and industrial adoption in Britain. Technovation, 2(3):209–231, June 1984. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649728490004X.

Abraham:1985:TMA

[AH85] Charles Abraham and George Hayward. Towards a microscopic analysis of industrial innovations: From diffusion curves to technological integration through participative management. Technovation, 3(1):3–17, February 1985. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649728590032X.

Aharoni:1988:WDG

[Aha88] Yair Aharoni. Why do governments privatize? Technovation, 8(1–3):7–23, ???? 1988. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649728890051X.

Abu-Hamatteh:2003:TQM

[AHAAEA03] Z. S. H. Abu-Hamatteh, T. A. Al-Azab, and M. El-Amyan. Total quality management achievement: King Abdullah II Award for Excellence of Jordan as a model. Technovation, 23(7):649–652, July 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720200127X.

Ahn:1995:NPC

[Ahn95] Soon Il Ahn. A new program in cooperative research between academia and industry in Korea, involving Centers of Excellence. Technovation, 15(4):241–257, May 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729596606T.

Apanasovich:2016:IBI

[AHP16] Natalja Apanasovich, Henar Alcalde Heras, and Mario Davide Parrilli. The impact of business innovation modes on SME innovation performance in post-Soviet transition economies: the case of Belarus. Technovation, 57–58(??):30–40, November/December 2016. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649721630030X.

Akter:2023:FAP

[AHS^{+}23] Shahriar Akter, Md Afnan Hossain, Shahriar Sajib, Saida Sultana, Mahfuzur Rahman, Demetris Vrontis, and Grace McCarthy. A framework for AI-powered service innovation capability: Review and agenda for future research. Technovation, 125(??):??, July 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000792.

Abubakar:2019:WSM

[AHSS19] Yazid Abdullahi Abubakar, Chris Hand, David Smallbone, and George Saridakis. What specific modes of internationalization influence SME innovation in Sub-Saharan least developed countries (LDCs)? Technovation, 79(??):56–70, January 2019. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497218303547.

Adeoti:2000:EDE

[AIOA00] O. Adeoti, M. O. Ilori, T. O. Oyebisi, and L. O. Adekoya. Engineering design and economic evaluation of a familysized biogas project in Nigeria. Technovation, 20(2):103– 108, February 2000. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497299001054.

Andersen:2004:SFT

[AJLR04] Per Dannemand Andersen, Birte Holst Jørgensen, Lars Lading, and Birgitte Rasmussen. Sensor foresighttechnology and market. Technovation, 24(4):311–320, April 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649720200072X.

Athreye:2000:TCG

[AK00] Suma Athreye and David Keeble. Technological convergence, globalisation and ownership in the UK computer industry. Technovation, 20(5):227–245, May 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001352.

Adamides:2006:ITS

[AK06] Emmanuel D. Adamides and Nikos Karacapilidis. Information technology support for the knowledge and social processes of innovation management. Technovation, 26(1):50– 59, January 2006. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497204001336.

Akarakiri:1998:ELS

[Aka98a] J. B. Akarakiri. Equipment leasing: a strategy for technology acquisition in Nigeria. Technovation, 18(5):347–352, ???? 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297001156.

Akarakiri:1998:ELA

[Aka98b] J. B. Akarakiri. Equipment leasing: a strategy for technology acquisition in Nigeria. Technovation, 18(5):365–366, ???? 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649729880031X.

Akgun:2007:ELC

[AKBA07] Ali E. Akgün, Halit Keskin, John C. Byrne, and Selim Aren. Emotional and learning capability and their impact on product innovativeness and firm performance. Technovation, 27(9):501–513, September 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720700034X.

Anzai:2012:HOM

[AKKS12] Tomohiro Anzai, Ryoichi Kusama, Hiroyuki Kodama, and Shintaro Sengoku. Holistic observation and monitoring of the impact of interdisciplinary academic research projects: an empirical assessment in Japan. Technovation, 32(6):345– 357, June 2012. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497211001842.

Ahn:2005:LFE

[AKL05] Jae-Hyeon Ahn, Myoung-Soo Kim, and Dong-Joo Lee. Learning from the failure: experiences in the Korean telecommunications market. Technovation, 25(1):69–82, January 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203000658.

Amara:2005:SID

[AL05] Nabil Amara and Réjean Landry. Sources of information as determinants of novelty of innovation in manufacturing firms: evidence from the 1999 statistics Canada innovation survey. Technovation, 25(3):245–259, March 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203001135.

Arza:2011:FLP

[AL11] Valeria Arza and Andrés López. Firms’ linkages with public research organisations in Argentina: Drivers, perceptions and behaviours. Technovation, 31(8):384–400, August 2011. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211000587.

Amara:2008:LNI

[ALBO08] Nabil Amara, Réjean Landry, Nizar Becheikh, and Mathieu Ouimet. Learning and novelty of innovation in established manufacturing SMEs. Technovation, 28(7):450–463, July 2008. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497208000151.

Alcorta:1999:DAA

[Alc99] Ludovico Alcorta. The diffusion of advanced automation in developing countries: factors and adoption process. Technovation, 19(3):163–175, January 1999. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000972.

Antolin-Lopez:2015:FPI

[ALCLdF^{+}15] Raquel Antolín-López, José Céspedes-Lorente, Nieves García de Frutos, Javier Martínez del Río, and Miguel Pérez-Valls. Fostering product innovation: Differences between new ventures and established firms. Technovation, 41–42(??):25–37, July/August 2015. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497215000127.

Ayerbe:2014:NCO

[ALCM14] C. Ayerbe, N. Lazaric, M. Callois, and L. Mitkova. The new challenges of organizing intellectual property in complex industries: a discussion based on the case of Thales. Technovation, 34(4):232–241, April 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214000029.

Ahmadpour:2023:KST

[ALFM23] Leila Ahmadpour, Alberto Leardini, Riccardo Fini, and Daniele Mascia. Knowledge sharing in temporary teams: Exploring the use of 3D printing in orthopaedic surgery. Technovation, 123(??):??, May 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000342.

Alic:1986:FRC

[Ali86] J. A. Alic. The federal role in commercial technology development. Technovation, 4(4):253–267, August 1986. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497286900295.

Alic:1990:CRI

[Ali90] John A. Alic. Cooperation in R&D. Technovation, 10(5): 319–332, July 1990. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729090016D.

Assefa:2022:UNI

[ALM22] Dawit Z. Assefa, Ching T. Liao, and Bisrat A. Misganaw. Unpacking the negative impact of initial informality on innovation: the mediating roles of investments in R&D and employee training. Technovation, 114(??):??, June 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000025.

Audretsch:2021:IAC

[ALMW21] David B. Audretsch, Erik E. Lehmann, Matthias Menter, and Katharine Wirsching. Intrapreneurship and absorptive capacities: the dynamic effect of labor mobility. Technovation, ??(??):??, ???? 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497218307922.

Amesse:2004:TEI

[ALRSD04] Fernand Amesse, Robert Latour, Claudia Rebolledo, and Louise Séguin-Dulude. The telecommunications equipment industry in the 1990s: from alliances to mergers and acquisitions. Technovation, 24(11):885–897, November 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000312.

Aaboen:2006:CGP

[ALvKL06] Lise Aaboen, Peter Lindelöf, Christopher von Koch, and Hans Löfsten. Corporate governance and performance of small high-tech firms in Sweden. Technovation, 26(8):955– 968, August 2006. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497205001069.

Ayres:1992:ELCa

[AM92a] Robert U. Ayres and Katalin Martinàs. Experience and the life cycle: Some analytic implications. Technovation, 12(7):465–486, November 1992. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729290052J.

Ayres:1992:ELCb

[AM92b] Robert U. Ayres and Katalin Martinás. Experience and the life cycle: Some analytical implications. Technovation, 12 (7):491–492, November 1992. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729290057O.

Azzone:1997:ERL

[AM97] Giovanni Azzone and Paolo Maccarrone. The emerging role of lean infrastructures in technology transfer: the case of the Innovation Plaza project. Technovation, 17(7):391–402, July 1997. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497296001198.

Al-Momani:2000:ESQ

[AM00] Ayman H. Al-Momani. Examining service quality within construction processes. Technovation, 20(11):643–651, November 2000. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649720000002X.

Abraham:2001:IAT

[AM01] Biju Paul Abraham and Soumyo D. Moitra. Innovation assessment through patent analysis. Technovation, 21(4): 245–252, April 2001. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497200000407.

Adebanjo:2010:AWE

[AM10] Dotun Adebanjo and Roula Michaelides. Analysis of Web 2.0 enabled e-clusters: a case study. Technovation, 30(4): 238–248, April 2010. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497209001370.

Abhari:2023:LFO

[AM23] Kaveh Abhari and Summer McGuckin. Limiting factors of open innovation organizations: a case of social product development and research agenda. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000736.

Alaassar:2021:ENI

[AMA21] Ahmad Alaassar, Anne-Laure Mention, and Tor Helge Aas. Exploring a new incubation model for FinTechs: Regulatory sandboxes. Technovation, 103(??):??, May 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000183.

Amey:1995:ACI

[Ame95] Robert George Amey. Automotive component innovation: Development and diffusion of engine management technologies. Technovation, 15(4):211–223, May 1995. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729596604R.

Angelidou:2022:EAC

[AMP22] Sofia Angelidou, Matthew Mount, and Krsto Pandza. Exploring the asymmetric complementarity between external knowledge search and management innovation. Technovation, 115(??):??, July 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222000190.

Ai:2024:TEB

[AMS^{+}24] Hongshan Ai, Sachin Kumar Mangla, Malin Song, Xiaoqing Tan, and Shangfeng Zhang. Technology-enabled business model innovation and carbon emission reduction: Evidence from a place-based policy in China. Technovation, 134(??): ??, June 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497224001020.

Aerts:2007:CRS

[AMV07] Kris Aerts, Paul Matthyssens, and Koen Vandenbempt. Critical role and screening practices of European business incubators. Technovation, 27(5):254–267, May 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497206001209.

Asongu:2016:RGM

[AN16] Simplice A. Asongu and Jacinta C. Nwachukwu. The role of governance in mobile phones for inclusive human development in Sub-Saharan Africa. Technovation, 55–56(??):1–13, September/October 2016. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497216300281.

Andersen:2021:RNR

[And21] Jim Andersén. A relational natural-resource-based view on product innovation: the influence of green product innovation and green suppliers on differentiation advantage in small manufacturing firms. Technovation, 104(??):??, June 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221000353.

Asadi:2022:EIT

[ANI^{+}22] Shahla Asadi, Mehrbakhsh Nilashi, Mohammad Iranmanesh, Sunghyup Sean Hyun, and Azadeh Rezvani. Effect of Internet of Things on manufacturing performance: a hybrid multi-criteria decision-making and neurofuzzy approach. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497221002078.

Anonymous:1981:BRE

[Ano81a] Anonymous. Book review: Educating the engineer for innovative and entrepreneurial activity: by Raymond Tomkins, published by the European Society for Engineering Education, Rue de la Concorde, 51, B 1050 Bruxelles, Belgium. 146 pp. Technovation, 1(1):81, February 1981. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497281900092.

Anonymous:1981:BRG

[Ano81b] Anonymous. Book review: Guidelines for product innovation: by R. C. Parker, published by the British Institute of Management Foundation, management house, Parker street, London WC2B 5 PT, UK 55 pp. Technovation, 1(1): 81, February 1981. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497281900109.

Anonymous:1981:EB

[Ano81c] Anonymous. Editorial Board. Technovation, 1(1):v, February 1981. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497281900018.

Anonymous:1982:CV

[Ano82a] Anonymous. Contents volume 1 (1981/1982). Technovation, 1(4):327–328, August 1982. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497282900128.

Anonymous:1982:PN

[Ano82b] Anonymous. Publisher’s note. Technovation, 1(4):253, August 1982. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497282900062.

Anonymous:1982:SDG

[Ano82c] Anonymous. Summary of the Dutch government white paper on innovation. Technovation, 1(3):225–252, February 1982. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497282900050.

Anonymous:1983:A

[Ano83a] Anonymous. Announcement. Technovation, 2(1):62, February 1983. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497283900093.

Anonymous:1983:EB

[Ano83b] Anonymous. Editorial Board. Technovation, 2(1):v, February 1983. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497283900019.

Anonymous:1984:CP

[Ano84a] Anonymous. Call for papers. Technovation, 2(4):299, July 1984. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497284900105.

Anonymous:1984:CV

[Ano84b] Anonymous. Contents volume 2 (1983/1984). Technovation, 2(4):301–302, July 1984. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497284900117.

Anonymous:1984:E

[Ano84c] Anonymous. Erratum. Technovation, 2(4):298, July 1984. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497284900099.

Anonymous:1984:PN

[Ano84d] Anonymous. Publishers note. Technovation, 2(2):67–68, May 1984. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497284900129.

Anonymous:1985:A

[Ano85a] Anonymous. Announcements. Technovation, 3(2):155–158, June 1985. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497285900094.

Anonymous:1985:CV

[Ano85b] Anonymous. Contents volume 3 (1985). Technovation, 3 (4):319–320, October 1985. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649728590029X.

Anonymous:1985:EB

[Ano85c] Anonymous. Editorial Board. Technovation, 3(1):v, February 1985. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497285900306.

Anonymous:1985:IN

[Ano85d] Anonymous. ISPIM news. Technovation, 3(3):231–232, August 1985. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497285900203.

Anonymous:1985:MAC

[Ano85e] Anonymous. Meeting announcement and call for papers. Technovation, 3(1):74, February 1985. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497285900409.

Anonymous:1986:CV

[Ano86a] Anonymous. Contents volume 4 (1986). Technovation, 4(4):321–322, August 1986. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497286900374.

Anonymous:1986:EBa

[Ano86b] Anonymous. Editorial Board. Technovation, 4(1):v, January 1986. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497286900179.

Anonymous:1986:EBb

[Ano86c] Anonymous. Editorial Board. Technovation, 5(1–3):iii, October 1986. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497286900386.

Anonymous:1986:ESU

[Ano86d] Anonymous. European symposium on the utilization of the results of public or publicly-funded research and development. Technovation, 4(2):163–164, April 1986. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497286900076.

Anonymous:1986:INa

[Ano86e] Anonymous. ISPIM news. Technovation, 4(1):87–88, January 1986. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497286900271.

Anonymous:1986:INb

[Ano86f] Anonymous. ISPIM news. Technovation, 4(2):165–166, April 1986. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497286900088.

Anonymous:1986:INc

[Ano86g] Anonymous. ISPIM news. Technovation, 4(3):247–249, June 1986. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497286900167.

Anonymous:1986:INd

[Ano86h] Anonymous. ISPIM news. Technovation, 4(4):320–321, August 1986. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497286900362.

Anonymous:1986:LC

[Ano86i] Anonymous. List of contributors. Technovation, 5(1– 3):201–203, October 1986. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497286900532.

Anonymous:1986:SI

[Ano86j] Anonymous. Subject index. Technovation, 5(1–3):205–207, October 1986. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497286900544.

Anonymous:1987:CVa

[Ano87a] Anonymous. Contents volume 5. Technovation, 5(4):331– 332, February 1987. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649728790071X.

Anonymous:1987:CVb

[Ano87b] Anonymous. Contents volume 6 (1987). Technovation, 6 (4):311–312, September 1987. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497287900848.

Anonymous:1987:EBa

[Ano87c] Anonymous. Editorial Board. Technovation, 6(1):v, April 1987. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497287900332.

Anonymous:1987:EBb

[Ano87d] Anonymous. Editorial Board. Technovation, 7(1):v, December 1987. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497287900423.

Anonymous:1987:E

[Ano87e] Anonymous. Erratum. Technovation, 7(1):79, December 1987. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497287900496.

Anonymous:1987:IN

[Ano87f] Anonymous. ISPIM news. Technovation, 6(2):150, June 1987. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497287900186.

Anonymous:1987:INIa

[Ano87g] Anonymous. ISPIM news: International society for product innovation management. Technovation, 6(1):73–74, April 1987. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497287900411.

Anonymous:1987:INIb

[Ano87h] Anonymous. ISPIM news: International Society for Product Innovation Management. Technovation, 7(1):83–84, December 1987. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497287900514.

Anonymous:1987:MA

[Ano87i] Anonymous. Meeting announcement. Technovation, 6(4): 309–310, September 1987. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497287900836.

Anonymous:1988:CC

[Ano88a] Anonymous. Conference calendar. Technovation, 8(4):327– 330, December 1988. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497288900387.

Anonymous:1988:CVa

[Ano88b] Anonymous. Contents volume 7 (1988). Technovation, 7 (4):427–429, October 1988. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649728890017X.

Anonymous:1988:CVb

[Ano88c] Anonymous. Contents volume 8. Technovation, 8(4):331– 332, December 1988. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497288900399.

Anonymous:1988:CIa

[Ano88d] Anonymous. Contributors to this issue. Technovation, 7 (4):415–417, October 1988. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497288900107.

Anonymous:1988:CIb

[Ano88e] Anonymous. Contributors to this issue. Technovation, 8 (4):313–314, December 1988. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497288900326.

Anonymous:1988:EB

[Ano88f] Anonymous. Editorial Board. Technovation, 8(1–3):ifc, ???? 1988. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649728890048X.

Anonymous:1988:FC

[Ano88g] Anonymous. Forthcoming conferences. Technovation, 7 (4):425–426, October 1988. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497288900168.

Anonymous:1988:MWB

[Ano88h] Anonymous. In memoriam Wayne S. Brown: 19 March 1928–17 May 1988. Technovation, 7(3):179–180, July 1988. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900181.

Anonymous:1988:LC

[Ano88i] Anonymous. List of contributors. Technovation, 8(1–3): 239–241, ???? 1988. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497288900636.

Anonymous:1988:NME

[Ano88j] Anonymous. New members of the editorial board. Technovation, 7(4):276–277, October 1988. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0166497288900028.

Anonymous:1988:NI

[Ano88k] Anonymous. News items. Technovation, 8(4):315–322, December 1988. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497288900338.

Anonymous:1988:SI

[Ano88l] Anonymous. Subject index. Technovation, 8(1–3):ix–xiii, ???? 1988. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497288900648.

Anonymous:1988:WSU

[Ano88m] Anonymous. Water 2020: Sustainable use for water in the 21st century. Technovation, 8(4):244–245, December 1988.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900260.

Anonymous:1989:AI

[Ano89a] Anonymous. Author index. Technovation, 9(8):699, December 1989. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497289900096.

Anonymous:1989:CCa

[Ano89b] Anonymous. Conference calendar. Technovation, 9(4):377– 379, July 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497287900095.

Anonymous:1989:CCb

[Ano89c] Anonymous. Conference calendar. Technovation, 9(7):613– 617, November 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497289900266.

Anonymous:1989:CRc

[Ano89d] Anonymous. Conference report. Technovation, 9(8):697– 698, December 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497289900084.

Anonymous:1989:C

[Ano89e] Anonymous. Contents. Technovation, 9(2–3):129–130, June 1989. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497289900503.

Anonymous:1989:CRb

[Ano89f] Anonymous. Contents of the report. Technovation, 9(2–3): 135, June 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497289900540.

Anonymous:1989:CRa

[Ano89g] Anonymous. Contributors to the report. Technovation, 9(2– 3):133, June 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497289900527.

Anonymous:1989:CIa

[Ano89h] Anonymous. Contributors to this issue. Technovation, 9(1): 113–114, May 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649728990045X.

Anonymous:1989:CIb

[Ano89i] Anonymous. Contributors to this issue. Technovation, 9(4): 371–372, July 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497287900071.

Anonymous:1989:CIc

[Ano89j] Anonymous. Contributors to this issue. Technovation, 9 (5):453–454, August 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497289900151.

Anonymous:1989:CId

[Ano89k] Anonymous. Contributors to this issue. Technovation, 9 (6):531–533, October 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497289900321.

Anonymous:1989:CIe

[Ano89l] Anonymous. Contributors to this issue. Technovation, 9 (7):607–608, November 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497289900230.

Anonymous:1989:CIf

[Ano89m] Anonymous. Contributors to this issue. Technovation, 9 (8):695–696, December 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497289900072.

Anonymous:1989:EB

[Ano89n] Anonymous. Editorial Board. Technovation, 9(1):v, May 1989. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497289900333.

Anonymous:1989:E

[Ano89o] Anonymous. Erratum. Technovation, 9(4):380, July 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497287900101.

Anonymous:1989:EFR

[Ano89p] Anonymous. EUREKA: a flexible response to the challenge of technological cooperation in Europe. Technovation, 9 (2–3):125–126, June 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497289900485.

Anonymous:1989:FR

[Ano89q] Anonymous. Foreword to the report. Technovation, 9(2–3): 134, June 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497289900539.

Anonymous:1989:GA

[Ano89r] Anonymous. Guide for authors. Technovation, 9(4):383– 386, July 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497287900125.

Anonymous:1989:IT

[Ano89s] Anonymous. Index of titles. Technovation, 9(8):701–702, December 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497289900102.

Anonymous:1989:IIa

[Ano89t] Anonymous. Industrial index. Technovation, 9(1):1, May 1989. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497289900345.

Anonymous:1989:IIb

[Ano89u] Anonymous. Industrial index. Technovation, 9(4):307, July 1989. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497287900010.

Anonymous:1989:IIc

[Ano89v] Anonymous. Industrial index. Technovation, 9(5):387, August 1989. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497289900114.

Anonymous:1989:IId

[Ano89w] Anonymous. Industrial index. Technovation, 9(7):535, November 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497289900187.

Anonymous:1989:IIe

[Ano89x] Anonymous. Industrial index. Technovation, 9(8):619, December 1989. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497289900011.

Anonymous:1989:INa

[Ano89y] Anonymous. ISPIM-news. Technovation, 9(1):117–119, May 1989. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497289900473.

Anonymous:1989:INb

[Ano89z] Anonymous. ISPIM-News. Technovation, 9(4):381–382, July 1989. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497287900113.

Anonymous:1989:PR

[Ano89-27] Anonymous. Preface to the report. Technovation, 9(2–3): 132, June 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497289900515.

Anonymous:1990:AI

[Ano90a] Anonymous. Author index. Technovation, 10(8):559, December 1990. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729090051K.

Anonymous:1990:CA

[Ano90b] Anonymous. Conference announcements. Technovation, 10(4):285–286, June 1990. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729090061N.

Anonymous:1990:CCa

[Ano90c] Anonymous. Conference calendar. Technovation, 10(1):61– 63, February 1990. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729090045L.

Anonymous:1990:CCb

[Ano90d] Anonymous. Conference calendar. Technovation, 10 (2):137–140, March 1990. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729090033G.

Anonymous:1990:CCc

[Ano90e] Anonymous. Conference calendar. Technovation, 10(4): 287–289, June 1990. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729090062O.

Anonymous:1990:CCd

[Ano90f] Anonymous. Conference calendar. Technovation, 10(5): 359–361, July 1990. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729090020K.

Anonymous:1990:CCe

[Ano90g] Anonymous. Conference calendar. Technovation, 10(6): 443–447, September 1990. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497290900118.

Anonymous:1990:CCf

[Ano90h] Anonymous. Conference calendar. Technovation, 10(7): 505–506, October 1990. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729090028I.

Anonymous:1990:CCg

[Ano90i] Anonymous. Conference calendar. Technovation, 10(8): 555–557, December 1990. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729090050T.

Anonymous:1990:EB

[Ano90j] Anonymous. Editorial Board. Technovation, 10(1):iii, February 1990. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729090036J.

Anonymous:1990:E

[Ano90k] Anonymous. Errata. Technovation, 10(6):451–452, September 1990. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729090013A.

Anonymous:1990:GE

[Ano90l] Anonymous. Guest editorial. Technovation, 10(1):3, February 1990. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729090037K.

Anonymous:1990:IT

[Ano90m] Anonymous. Index of titles. Technovation, 10(8):560, December 1990. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729090052L.

Anonymous:1990:IN

[Ano90n] Anonymous. ISPIM news. Technovation, 10(1):65–68, February 1990. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729090046M.

Anonymous:1990:IJI

[Ano90o] Anonymous. ISPIM’91: in Jerusalem, Israel 7th international conference on: Innovation strategies concepts, implementations, competitive advantage. Technovation, 10 (6):449–450, September 1990. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497290900129.

Anonymous:1991:A

[Ano91a] Anonymous. Announcement. Technovation, 11(8):505, December 1991. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729190012S.

Anonymous:1991:AI

[Ano91b] Anonymous. Author index. Technovation, 11(8):509, December 1991. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729190014U.

Anonymous:1991:BRE

[Ano91c] Anonymous. Book review: An encyclopaedia of the history of technology: Edited by Ian McNeil, Routledge, London and New York, 1990, Hardback £65.00, 1062 pages, including an Index of Names and an Index of Topics. Technovation, 11(2):121–122, March 1991. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497291900434.

Anonymous:1991:BRC

[Ano91d] Anonymous. Book review: Companion to the history of modern science: edited by R. C. Olby, G. N. Cantor, J. R. R. Christie and M. J. S. Hodge, Routledge, London and New York, 1990, Hardback £65.00, 1081 pages, including Name and Subject Indexes. Technovation, 11(2):121, March 1991. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497291900423.

Anonymous:1991:CC

[Ano91e] Anonymous. Conference calendar. Technovation, 11(8): 506–507, December 1991. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729190013T.

Anonymous:1991:EB

[Ano91f] Anonymous. Editorial Board. Technovation, 11(1):ifc, January 1991. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729190001K.

Anonymous:1991:SI

[Ano91g] Anonymous. Subject index. Technovation, 11(8):511–513, December 1991. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729190015V.

Anonymous:1992:C

[Ano92a] Anonymous. Calendar. Technovation, 12(6):421–422, September 1992. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729290015A.

Anonymous:1992:CV

[Ano92b] Anonymous. Contents of volume 12, 1992. Technovation, 12(8):541–545, December 1992. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729290081R.

Anonymous:1992:DSD

[Ano92c] Anonymous. The diffusion and spatial distribution of new telecommunication technologies in the Italian region of Campania. Technovation, 12(6):407–413, September 1992. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497292900064.

Anonymous:1992:EB

[Ano92d] Anonymous. Editorial Board. Technovation, 12(1):ifc, February 1992. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729290027F.

Anonymous:1992:IAV

[Ano92e] Anonymous. Index of authors, vol. 12, 1992. Technovation, 12(8):549, December 1992. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729290083T.

Anonymous:1992:IBR

[Ano92f] Anonymous. Index of book reviews, vol. 12, 1992. Technovation, 12(8):547, December 1992. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/016649729290082S.

Anonymous:1992:NI

[Ano92g] Anonymous. News items. Technovation, 12(4):279–280, May 1992. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729290047L.

Anonymous:1992:TA

[Ano92h] Anonymous. Translated abstracts. Technovation, 12(5): 343–349, July 1992. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729290042G.

Anonymous:1993:ANT

[Ano93a] Anonymous. Adopting new technologies: an entrepreneurial act: Yves-Chantal Gagnon and Jean-Marie Toulouse, Technovation, 13(7) (1993) 411–424. Technovation, 13(7):471– 472, November 1993. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729390029U. See [GT93].

Anonymous:1993:BPS

[Ano93b]  Anonymous.  Basic points for strategic production

organization requiring simultaneous strategies: Hans Corsten and Thomas Will, Technovation, 13:2 (1993) 000–000. Technovation, 13(2):110, March 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497293900609. See [CW93].

Anonymous:1993:CV

[Ano93c] Anonymous. Contents of volume 13, 1993. Technovation, 13(8):551–553, December 1993. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497293900724.

Anonymous:1993:EB

[Ano93d] Anonymous. Editorial Board. Technovation, 13(1):ifc, January 1993. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729390008J.

Anonymous:1993:FST

[Ano93e] Anonymous. A filter system for technology evaluation a selection: Chee Meng Yap and Wm. E. Souder, Technovation, 13(7) (1993) 449–469. Technovation, 13(7):475–476, November 1993. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729390032Q. See [YS93].

Anonymous:1993:FMC

[Ano93f] Anonymous. A framework for managing communication at the R&D/marketing interface: A. W. Pearson and D. F. Ball, Technovation, 13(7) (1993) 439–

447. Technovation, 13(7):474–475, November 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390031P. See [PB93].

Anonymous:1993:IAVb

[Ano93g] Anonymous. Index of authors, vol. 13, 1993. Technovation, 13(8):556, December 1993. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497293900746.

Anonymous:1993:IBR

[Ano93h] Anonymous. Index of book reviews. vol. 13, 1993. Technovation, 13(8):555, December 1993. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497293900735.

Anonymous:1993:ICT

[Ano93i] Anonymous. The information content of technology transfers: a transactions cost analysis of the machine tool industry: Howard Davies, Technovation, 13:2 (1993) 000–000. Technovation, 13(2):113, March 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390062Z. See [Dav93].

Anonymous:1993:IAVa

[Ano93j] Anonymous. An interregional analysis of venture capital and technology funding in the UK: Heather I. M. Wilson, Technovation, 13(7) (1993) 425–438. Technovation, 13(7):472–474, November 1993. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729390030Y.

See [Wil93].

Anonymous:1993:IMI

[Ano93k] Anonymous. Issues in management information systems planning: Yash P. Gupta and Tor Guimaraes, Technovation, 13(8), (1993) 533–544. Technovation, 13(8):549–550, December 1993. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497293900713. See [GG93].

Anonymous:1993:NLU

[Ano93l] Anonymous. A new look at the US pattern of federal R&D spending: preliminary lessons for newly industrializing economies: J. A. Pimenta-Bueno and R. T. S. Correa, Technovation, 13:2 (1993) 000–000. Technovation, 13(2):113, March 1993. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497293900632. See [PBC93].

Anonymous:1993:SPI

[Ano93m] Anonymous. The success potential of international R&D cooperation: Thorsten Teichert, Technovation, 13(8) (1993) 519–532. Technovation, 13(8):548–549, December 1993. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390070C. See [Tei93].

Anonymous:1993:TDB

[Ano93n] Anonymous. Technological development in Brazilian business: four behavioral types: Erno I. Paulinyi, Technovation, 13:2 (1993) 000–000. Technovation, 13(2):111, March 1993. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390061Y. See [Pau93].

Anonymous:1993:TCV

[Ano93o] Anonymous. Technology and the creation of value in services: a conceptual model: Jean Harvey, Élizabeth Lefebvre and Louis A. Lefebvre, Technovation, 13(8) (1993) 481–495. Technovation, 13(8):545–546, December 1993. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497293900687. See [HLL93].

Anonymous:1993:TTL

[Ano93p] Anonymous. Technology transfer as a learning and developmental process: an analysis of Norwegian programmes on technology transfer: Morten Levin, Technovation, 13(8) (1993) 497–518. Technovation, 13(8):546–548, December 1993. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497293900698. See [Lev93].

Anonymous:1993:TAa

[Ano93q] Anonymous. Translations of abstracts. Technovation, 13(4): 255–259, May 1993. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729390022N.

Anonymous:1993:TAb

[Ano93r] Anonymous. Translations of abstracts. Technovation, 13(5): 325–330, July 1993. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729390005G.

Anonymous:1993:USE

[Ano93s] Anonymous. University scientists as entrepreneurs: a special case of technology transfer and high-tech venturing: Karel A. Samsom and Michael A. Gurdon, Technovation, 13:2 (1993) 000–000. Technovation, 13(2):109, March 1993. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497293900595. See [SG93].

Anonymous:1994:ARP

[Ano94a] Anonymous. Anaerobic recycling of prehydrolysate generated during the manufacture of rayon grade pulp. Technovation, 14(8):564–565, October 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0166497294901589.

Anonymous:1994:ADL

[Ano94b] Anonymous. Arthur D. Little/Cambridge consultants and the round tea bag. Technovation, 14(9):571–572, November 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294900396.

Anonymous:1994:ACS

[Ano94c] Anonymous. Automation — its concept and a short history: Katsundo Hitomi, Technovation, 14(2) 121–

128. Technovation, 14(2):134–135, March 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294901074. See [Hit94].

Anonymous:1994:BRV

[Ano94d] Anonymous. Book reviews, vol. 14, 1994. Technovation, 14(10):705, December 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729490121X.

Anonymous:1994:CV

[Ano94e] Anonymous. Contents of volume 14, 1994. Technovation, 14(10):699–703, December 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497294901201.

Anonymous:1994:CRG

[Ano94f] Anonymous. Cooperative R&D in government laboratories: comparing the USA and Japan: Barry Bozeman and Sanjay Pandey, Technovation, 14(3) (1994) 145–

159. Technovation, 14(3):205–206, April 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900574. See [BP94].

Anonymous:1994:FMZ

[Ano94g] Anonymous. Der Faktor Mensch in Zusammenarbeit und Technologietransfer: Beobachtungen und Profile. (German)

[The human factor in collaboration and technology transfer: observations and profiles]. Technovation, 14(5):345, June 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294900787.

Anonymous:1994:EB

[Ano94h] Anonymous. Editorial Board. Technovation, 14(1):ifc, February 1994. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497294900639.

Anonymous:1994:FPC

[Ano94i] Anonymous. El factor de las personas en la colaboración y las transferencias de tecnologia: observaciones y perfil. (Spanish) [The role of people in collaboration and technology transfers: observations and profile]. Technovation, 14(5):345–346, June 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900795.

Anonymous:1994:EBI

[Ano94j] Anonymous. Enststehende Biotechnologien und ihre potentielle Anwendung in der Produktion und Verarbeitung von Kakao und Palmenerzeugnissen in Nigeria. (German) [Developing biotechnologies and their potential application in the production and processing of cocoa and palm products in Nigeria]. Technovation, 14(5):346, June 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900825.

Anonymous:1994:EER

[Ano94k] Anonymous. The evolution of ecodesign: Robin Roy, Technovation, 14(6) (1994) 363–380. Technovation, 14(6):419, August 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294900205. See [Roy94].

Anonymous:1994:FPT

[Ano94l] Anonymous. The formation process of Taiwan’s IC industry — method of technology transfer: Pao-Long Chang, Chintay Shih and Chiung-Wen Hsu, Technovation, 14(3) (1994) 161–171. Technovation, 14(3):206–207, April 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900582. See [CSH94].

Anonymous:1994:HFM

[Ano94m] Anonymous. How flexible manufacturing systems affect tooling and machining enterprises: Albert N. Link and Perry D. Quick, Technovation, 14(2) 85–92. Technovation, 14(2):131, March 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729490104X.

See [LQ94].

Anonymous:1994:HII

[Ano94n] Anonymous. How important is the innovator for the commercial success of innovative products in SMEs: Sally Caird, Technovation, 14(2) 71–83. Technovation, 14(2):129, March 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294901031. See [Cai94].

Anonymous:1994:IAV

[Ano94o] Anonymous. Index of authors, vol. 14, 1994. Technovation, 14(10):707, December 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497294901228.

Anonymous:1994:IFT

[Ano94p] Anonymous. Interindustry flows of technology: an analysis of the Canadian patent matrix and input-output matrix for 1978–1989. Technovation, 14(8):562–564, October 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294901570.

Anonymous:1994:ECT

[Ano94q] Anonymous. La estructura competitiva y los tipos estratégicos en la industria coreana de software. (Spanish) [Competitive structure and strategic types in the Korean software industry]. Technovation, 14(5):348, June 1994.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900876.

Anonymous:1994:MEP

[Ano94r] Anonymous. La mejora de la eficacia en el proceso de la innovación de la defensa: un marco empírico. (Spanish) [Improving efficiency in the defense innovation process: an empirical framework]. Technovation, 14(5):349, June 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900914.

Anonymous:1994:SCT

[Ano94s] Anonymous. La structure concurrentielle et les typologies de la stratégie dans l’industrie coréenne du logiciel. Technovation, 14(5):347, June 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729490085X.

Anonymous:1994:BSP

[Ano94t] Anonymous. Las biotecnologías y sus posibles applicaciones en la producción y elaboración del cacao y los productos derivados de la palmera en Nigeria. (Spanish) [Biotechnologies and their possible applications in the production and processing of cocoa and products derived from the palm tree in Nigeria]. Technovation, 14(5):346–347, June 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900833.

Anonymous:1994:FHC

[Ano94u] Anonymous. Le facteur humain dans la collaboration et le transfert de la technologie: quelques observations et un profil. (French) [The human factor in collaboration and transfer of technology: some observations and a profile]. Technovation, 14(5):345, June 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900779.

Anonymous:1994:BNL

[Ano94v] Anonymous. Les biotechnologies naissantes et leur application potentielle dans la production et la transformation du cacao et des produits du palmier au Nigeria. Technovation, 14(5):346, June 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900817.

Anonymous:1994:RRE

[Ano94w] Anonymous. Les réseaux régionaux des entreprises de technologie de pointe — le cas de la région de greater Boston. (French) [Regional networks of technology — the case of Greater Boston]. Technovation, 14(5):349–350, June 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900930.

Anonymous:1994:NTB

[Ano94x] Anonymous. New, technology-based firms as agents of R&D and innovation: an empirical study. Technovation, 14(4): 277, May 1994. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497294900132.

Anonymous:1994:PDH

[Ano94y] Anonymous. Planning development in high-tech firms. Technovation, 14(8):559, October 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0166497294901546.

Anonymous:1994:QRE

[Ano94z] Anonymous. Quality in research: an empirical study: Roger Miller, Technovation, 14(6) (1994) 381–394. Technovation, 14(6):420–421, August 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497294900213.

See [Mil94].

Anonymous:1994:RFS

[Ano94-27] Anonymous. R&D and firm size: some evidence from Spain: Angel Martínez Sanchéz, Technovation, 14(2)

63–69. Technovation, 14(2):129, March 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294901023. See [Sán94].

Anonymous:1994:RNE

[Ano94-28] Anonymous. Redes nacionales de empresas de alta tecnologia — el caso de la región de greater Boston. (Spanish) [National networks of high-tech enterprises — the case of the Greater Boston region]. Technovation, 14(5):350, June 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294900957.

Anonymous:1994:RNH

[Ano94-29] Anonymous. Regionale Netzwerke von high-tech Firmen — der Fall der Region greater Boston. (German) [Regional networks of high-tech companies — the case of the Greater Boston region]. Technovation, 14(5):350, June 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900949.

Anonymous:1994:RIP

[Ano94-30] Anonymous. The role of intellectual property rights in stimulating commercialization in ASEAN: lessons from Canada: Atipol Bhanich Supapol and Frederic William Swierczek, Technovation, 14(3) (1994) 181– 195. Technovation, 14(3):208–210, April 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900604. See [SS94].

Anonymous:1994:RLG

[Ano94-31] Anonymous. The role of local government in technological and industrial development in Nigeria: M. O. Ilori, Technovation, 14(3) (1994) 173–179. Technovation, 14(3):207–208, April 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294900590. See [Ilo94].

Anonymous:1994:RSA

[Ano94-32] Anonymous. The role of strategic alliances in the development of technology. Technovation, 14(4):276–277, May 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294900124.

Anonymous:1994:SSG

[Ano94-33] Anonymous. Simultaneously supporting generic competitive strategies by production management: Supportive concepts of information technology and work organization: Hans Corsten and Thomas Will, Technovation, 14(2) 111–120. Technovation, 14(2):133–134, March 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294901066. See [CW94].

Anonymous:1994:SIT

[Ano94-34] Anonymous. Steps for the introduction of technology management in developing economies: the role of public governments: Dimitris Deniozos, Technovation, 14(3) (1994) 197–203. Technovation, 14(3):210–211, April 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900612. See [Den94].

Anonymous:1994:SES

[Ano94-35] Anonymous. Strategy evolution and structural consequences in the US semiconductor industry. Technovation, 14(4):275–276, May 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497294900116.

Anonymous:1994:AED

[Ano94-36] Anonymous. Sur l’améliorisation de l’efficacité du processus d’innovation de la défense: un cadre empirique. (French) [On improving the effectiveness of the process of innovation in defense: an empirical framework]. Technovation, 14(5): 348, June 1994. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497294900892.

Anonymous:1994:SPO

[Ano94-37] Anonymous. Synergistic pay-offs in Indo-US strategic alliances: Atul Parvatiyar and Yash P. Gupta, Technovation, 14(6) (1994) 395–406. Technovation, 14(6):421, August 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294900221. See [PG94].

Anonymous:1994:TPG

[Ano94-38] Anonymous. ‘Technopolises’ as a policy goal: a morphological study of the Wissenschaftsstadt Ulm: Christian Boucke, Uwe Cantner and Horst Hanusch, Technovation, 14(6) (1994) 407–418. Technovation, 14(6):422, August 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729490023X. See [BCH94].

Anonymous:1994:TAa

[Ano94-39] Anonymous. Translations of abstracts. Technovation, 14 (1):55–59, February 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497294900701.

Anonymous:1994:TAb

[Ano94-40] Anonymous. Translations of abstracts. Technovation, 14 (7):485–491, September 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729490006X.

Anonymous:1994:VEI

[Ano94-41] Anonymous. Über die Verbesserung der Effektivität des Innovationsprozesses in der Verteidigungsindustrie. (German) [On the improvement of the effectiveness of the innovation process in the defense industry]. Technovation, 14(5):349, June 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294900906.

Anonymous:1994:URS

[Ano94-42] Anonymous. University-related science parks — ‘seedbeds’ or ‘enclaves’ of innovation?: Daniel Felsenstein, Technovation, 14(2) 93–110. Technovation, 14(2):131–133, March 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294901058. See [Fel94].

Anonymous:1994:UUS

[Ano94-43] Anonymous. US university-sponsored technology incubators: an overview of management, policies and performance. Technovation, 14(8):561–562, October 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294901562.

Anonymous:1994:UTN

[Ano94-44] Anonymous. Using technical norm in explaining technological action. Technovation, 14(8):560–561, October 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294901554.

Anonymous:1994:WSS

[Ano94-45] Anonymous. Wettbewerbsfähige Struktur und strategische Typen in der koreanischen Softwareindustrie. (German) [Competitive structure and strategic types in the Korean software industry]. Technovation, 14(5):347–348, June 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294900868.

Anonymous:1995:BRV

[Ano95a] Anonymous. Book review, vol. 15, 1995. Technovation, 15(10):637, December 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497295900535.

Anonymous:1995:CCa

[Ano95b] Anonymous. Conference calendar. Technovation, 15(1):53– 54, February 1995. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497295900098.

Anonymous:1995:CCc

[Ano95c] Anonymous. Conference calendar. Technovation, 15 (2):131–132, March 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497295900217.

Anonymous:1995:CCd

[Ano95d] Anonymous. Conference calendar. Technovation, 15(3): 193–194, April 1995. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497295900314.

Anonymous:1995:CCe

[Ano95e] Anonymous. Conference calendar. Technovation, 15(4): 267–268, May 1995. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497295900675.

Anonymous:1995:CCf

[Ano95f] Anonymous. Conference calendar. Technovation, 15(5): 337–338, June 1995. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497295900373.

Anonymous:1995:CCg

[Ano95g] Anonymous. Conference calendar. Technovation, 15(6): 405–406, August 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497295900616.

Anonymous:1995:CCh

[Ano95h] Anonymous. Conference calendar. Technovation, 15(7): 473–474, September 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497295900713.

Anonymous:1995:CCi

[Ano95i] Anonymous. Conference calendar. Technovation, 15(8): 533–534, October 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497295900136.

Anonymous:1995:CCj

[Ano95j] Anonymous. Conference calendar. Technovation, 15(9): 589–590, November 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497295900756.

Anonymous:1995:CCb

[Ano95k] Anonymous. Conference calendar. Technovation, 15(10): 641, December 1995. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497295900578.

Anonymous:1995:CV

[Ano95l] Anonymous. Content of volume 15, 1995. Technovation, 15 (10):631–635, December 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497295900519.

Anonymous:1995:EB

[Ano95m] Anonymous. Editorial Board. Technovation, 15(1):ifc, February 1995. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497295900012.

Anonymous:1995:IAV

[Ano95n] Anonymous. Index of author, vol. 15, 1995. Technovation, 15(10):639, December 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497295900551.

Anonymous:1995:TAc

[Ano95o] Anonymous. Translations of abstract. Technovation, 15 (2):121–127, March 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497295900152.

Anonymous:1995:TAe

[Ano95p] Anonymous. Translations of abstract. Technovation, 15(4): 259–264, May 1995. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497295900632.

Anonymous:1995:TAf

[Ano95q] Anonymous. Translations of abstract. Technovation, 15(5): 329–333, June 1995. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497295900330.

Anonymous:1995:TAa

[Ano95r] Anonymous. Translations of abstracts. Technovation, 15 (1):45–48, February 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497295900039.

Anonymous:1995:TAd

[Ano95s] Anonymous. Translations of abstracts. Technovation, 15(3): 183–187, April 1995. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497295900233.

Anonymous:1995:TAg

[Ano95t] Anonymous. Translations of abstracts. Technovation, 15 (6):399–404, August 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497295900594.

Anonymous:1995:TAh

[Ano95u] Anonymous. Translations of abstracts. Technovation, 15 (7):467–471, September 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497295900691.

Anonymous:1995:TAi

[Ano95v] Anonymous. Translations of abstracts. Technovation, 15 (8):527–531, October 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729590011X.

Anonymous:1995:TAj

[Ano95w] Anonymous. Translations of abstracts. Technovation, 15 (9):585–588, November 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729590073X.

Anonymous:1995:TAb

[Ano95x] Anonymous. Translations of abstracts. Technovation, 15 (10):627–629, December 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497295900497.

Anonymous:1996:BR

[Ano96a] Anonymous. Book reviews. Technovation, 16(10):617, October 1996. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497296901042.

Anonymous:1996:CCa

[Ano96b] Anonymous. Conference calendar. Technovation, 16(1):I– III, January 1996. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497296902400.

Anonymous:1996:CCc

[Ano96c] Anonymous. Conference calendar. Technovation, 16(2):I– III, February 1996. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497296900103.

Anonymous:1996:CCd

[Ano96d] Anonymous. Conference calendar. Technovation, 16(3):i– ii, March 1996. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497296901716.

Anonymous:1996:CCe

[Ano96e] Anonymous. Conference calendar. Technovation, 16(4):i–ii, April 1996. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497296901091.

Anonymous:1996:CCf

[Ano96f] Anonymous. Conference calendar. Technovation, 16(5):i–ii, May 1996. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497296900206.

Anonymous:1996:CCg

[Ano96g] Anonymous. Conference calendar. Technovation, 16(6):i–ii, June 1996. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497296901935.

Anonymous:1996:CCh

[Ano96h] Anonymous. Conference calendar. Technovation, 16(7):i–ii, July 1996. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649729690259X.

Anonymous:1996:CCi

[Ano96i] Anonymous. Conference calendar. Technovation, 16(8):i– iii, August 1996. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497296903466.

Anonymous:1996:CCj

[Ano96j] Anonymous. Conference calendar. Technovation, 16(9):i–ii, September 1996. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497296903041.

Anonymous:1996:CCb

[Ano96k] Anonymous. Conference calendar. Technovation, 16(10): I–II, October 1996. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649729690108X.

Anonymous:1996:CV

[Ano96l] Anonymous. Contents of volume 16, 1996. Technovation, 16(10):611–615, October 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497296901029.

Anonymous:1996:EB

[Ano96m] Anonymous. Editorial Board. Technovation, 16(1):ifc, January 1996. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497296902370.

Anonymous:1996:IAV

[Ano96n] Anonymous. Index of authors, vol. 16, 1996. Technovation, 16(10):619, October 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497296901066.

Anonymous:1996:TAa

[Ano96o] Anonymous. Translations of abstracts. Technovation, 16 (1):47–50, January 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497296902394.

Anonymous:1996:TAc

[Ano96p] Anonymous. Translations of abstracts. Technovation, 16 (3):153–157, March 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497296901777.

Anonymous:1996:TAd

[Ano96q] Anonymous. Translations of abstracts. Technovation, 16(4): 211–216, April 1996. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497296901212.

Anonymous:1996:TAe

[Ano96r] Anonymous. Translations of abstracts. Technovation, 16(6): 313–316, June 1996. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497296901959.

Anonymous:1996:TAf

[Ano96s] Anonymous. Translations of abstracts. Technovation, 16(7): 371–375, July 1996. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497296902606.

Anonymous:1996:TAg

[Ano96t] Anonymous. Translations of abstracts. Technovation, 16 (8):443–449, August 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497296903478.

Anonymous:1996:TAb

[Ano96u] Anonymous. Translations of abstracts. Technovation, 16 (10):601–607, October 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497296901005.

Anonymous:1997:A

[Ano97a] Anonymous. Announcement. Technovation, 17(3):iii, March 1997. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297902078.

Anonymous:1997:BRS

[Ano97b] Anonymous. Book review: Scientific management: Frederick Winslow Taylor’s gift to the world?: J.-C. Spender and Hugo Kijne (Eds.) Kluwer Academic Publishers Group, The Netherlands, 1996, Hardbound, xx + 191 pp., ISBN 0-7923-9758-4, Dfl.195.00/US$105.00/ UK£74.75. Technovation, 17(8):475, August 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297834657.

Anonymous:1997:CCa

[Ano97c] Anonymous. Conference calendar. Technovation, 17(1):I– II, January 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497297902959.

Anonymous:1997:CCd

[Ano97d] Anonymous. Conference calendar. Technovation, 17(2):I– II, February 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497297901735.

Anonymous:1997:CCe

[Ano97e] Anonymous. Conference calendar. Technovation, 17(3):I– II, March 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649729790211X.

Anonymous:1997:CCf

[Ano97f] Anonymous. Conference calendar. Technovation, 17(4):I–II, April 1997. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297901176.

Anonymous:1997:CCg

[Ano97g] Anonymous. Conference calendar. Technovation, 17(5):I– II, May 1997. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297903321.

Anonymous:1997:CCh

[Ano97h] Anonymous. Conference calendar. Technovation, 17(6):I–II, June 1997. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297903011.

Anonymous:1997:CCi

[Ano97i] Anonymous. Conference calendar. Technovation, 17(7):I– II, July 1997. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297900079.

Anonymous:1997:CCj

[Ano97j] Anonymous. Conference calendar. Technovation, 17(8):I– II, August 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497297903096.

Anonymous:1997:CCk

[Ano97k] Anonymous. Conference calendar. Technovation, 17(9):I– II, September 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497297903357.

Anonymous:1997:CCb

[Ano97l] Anonymous. Conference calendar. Technovation, 17(10): i, October 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497297900389.

Anonymous:1997:CCc

[Ano97m] Anonymous. Conference calendar. Technovation, 17(11– 12):I–II, November/December 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0166497297904673.

Anonymous:1997:CV

[Ano97n] Anonymous. Contents of volume 17, 1997. Technovation, 17(11–12):737–744, November/December 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297887743.

Anonymous:1997:EB

[Ano97o] Anonymous. Editorial Board. Technovation, 17(1):ifc, January 1997. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297902935.

Anonymous:1997:IAV

[Ano97p] Anonymous. Index of authors, vol. 17, 1997. Technovation, 17(11–12):745, November/December 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297887755.

Anonymous:1997:TAb

[Ano97q] Anonymous. Translations of abstract. Technovation, 17 (2):101–105, February 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497297901711.

Anonymous:1997:TAa

[Ano97r] Anonymous. Translations of abstracts. Technovation, 17 (1):47–50, January 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497297902947.

Anonymous:1997:TAc

[Ano97s] Anonymous. Translations of abstracts. Technovation, 17 (3):161–165, March 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497297902091.

Anonymous:1997:TAd

[Ano97t] Anonymous. Translations of abstracts. Technovation, 17(4): 221–224, April 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497297901152.

Anonymous:1997:TAe

[Ano97u] Anonymous. Translations of abstracts. Technovation, 17(5): 279–283, May 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649729790331X.

Anonymous:1997:TAf

[Ano97v] Anonymous. Translations of abstracts. Technovation, 17(6): 341–346, June 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649729790300X.

Anonymous:1997:TAg

[Ano97w] Anonymous. Translations of abstracts. Technovation, 17(7): 409–414, July 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497297900067.

Anonymous:1997:TAh

[Ano97x] Anonymous. Translations of abstracts. Technovation, 17 (8):471–474, August 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497297903084.

Anonymous:1997:TAi

[Ano97y] Anonymous. Translations of abstracts. Technovation, 17 (9):529–533, September 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497297903333.

Anonymous:1998:BRe

[Ano98a] Anonymous. Book review. Technovation, 18(3):221, March 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649729700120X.

Anonymous:1998:BRf

[Ano98b] Anonymous. Book review. Technovation, 18(3):221–222, March 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297001004.

Anonymous:1998:BRg

[Ano98c] Anonymous. Book review. Technovation, 18(4):294–295, April 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297000850.

Anonymous:1998:BRh

[Ano98d] Anonymous. Book review. Technovation, 18(4):295, April 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297000254.

Anonymous:1998:BRi

[Ano98e] Anonymous. Book review. Technovation, 18(6–7):491–493, January 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497298000352.

Anonymous:1998:BRa

[Ano98f] Anonymous. Book review. Technovation, 18(10):661–662, October 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497298000637.

Anonymous:1998:BRc

[Ano98g] Anonymous. Book review. Technovation, 18(11):719–720, November 1998. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497298000649.

Anonymous:1998:BRd

[Ano98h] Anonymous. Book review. Technovation, 18(12):775–786, October 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497298000844.

Anonymous:1998:BRj

[Ano98i] Anonymous. Book review. Technovation, 19(1):69, November 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497298001023.

Anonymous:1998:BRk

[Ano98j] Anonymous. Book review. Technovation, 19(2):133–134, December 1998. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497298001205.

Anonymous:1998:BRM

[Ano98k] Anonymous. Book review: Managing Learning. Technovation, 18(1):76–77, January 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497297000849.

Anonymous:1998:BRB

[Ano98l] Anonymous. Book review: the Boundaryless Organization. Technovation, 18(1):75–76, January 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297000953.

Anonymous:1998:CC

[Ano98m] Anonymous. Conference calendar. Technovation, 18(2):I– II, February 1998. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649729890075X.

Anonymous:1998:I

[Ano98n] Anonymous. Index. Technovation, 18(12):787, October 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497298001199.

Anonymous:1998:TAF

[Ano98o] Anonymous. Translation of Abstracts into French, German and Spanish. Technovation, 18(1):69–74, January 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649729700103X.

Anonymous:1999:BRa

[Ano99a] Anonymous. Book review. Technovation, 19(3):201, January 1999. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497298001291.

Anonymous:1999:BRb

[Ano99b] Anonymous. Book review. Technovation, 19(9):589–591, September 1999. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497299000693.

Anonymous:1999:BRc

[Ano99c] Anonymous. Book review. Technovation, 19(9):591–592, September 1999. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497299000681.

Anonymous:1999:E

[Ano99d] Anonymous. Editorial. Technovation, 19(6–7):339–344, June 1999. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497299000577.

Anonymous:1999:I

[Ano99e] Anonymous. Index. Technovation, 19(12):768, December 1999. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497299001509.

Anonymous:2000:E

[Ano00a] Anonymous. Editorial. Technovation, 20(1):1, January 2000. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497299001674.

Anonymous:2000:I

[Ano00b] Anonymous. Index. Technovation, 20(12):x, December 2000. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497200000596.

Anonymous:2001:AI

[Ano01a] Anonymous. author index. Technovation, 21(12):x, December 2001. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649720100092X.

Anonymous:2001:BR

[Ano01b] Anonymous. book reviews. Technovation, 21(12):ix, December 2001. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201000918.

Anonymous:2001:CV

[Ano01c] Anonymous. contents of volume 21, 2001. Technovation, 21(12):I–VIII, December 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497201000906.

Anonymous:2002:AI

[Ano02a] Anonymous. Author Index. Technovation, 22(12):x, December 2002. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497202001463.

Anonymous:2002:CV

[Ano02b] Anonymous. Contents of Volume 22. Technovation, 22 (12):I–VIII, December 2002. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016649720200144X.

Anonymous:2002:EBb

[Ano02c] Anonymous. Editorial Board. Technovation, 22(10):645– 653, October 2002. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497201000554.

Anonymous:2002:EBa

[Ano02d] Anonymous. Editorial Board. Technovation, 22(10):ifc, October 2002. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497202000834.

Anonymous:2002:EBc

[Ano02e] Anonymous. Editorial Board. Technovation, 22(11):ifc, November 2002. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497202001050.

Anonymous:2002:EBd

[Ano02f] Anonymous. Editorial Board. Technovation, 22(12):ifc, December 2002. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497202001438.

Anonymous:2002:IBR

[Ano02g] Anonymous. Index — book reviews. Technovation, 22(12): ix, December 2002. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497202001451.

Anonymous:2003:AI

[Ano03a] Anonymous. Author index. Technovation, 23(12):XI–XII, December 2003. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203001895.

Anonymous:2003:BRI

[Ano03b] Anonymous. Book review index. Technovation, 23(12):x, December 2003. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203001883.

Anonymous:2003:CV

[Ano03c] Anonymous. contents of volume. Technovation, 23(12):I– IX, December 2003. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203001871.

Anonymous:2003:EBa

[Ano03d] Anonymous. Editorial Board. Technovation, 23(1):ifc, January 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497202001591.

Anonymous:2003:EBe

[Ano03e] Anonymous. Editorial Board. Technovation, 23(2):ifc, February 2003. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497202001724.

Anonymous:2003:EBf

[Ano03f] Anonymous. Editorial Board. Technovation, 23(3):ifc, March 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203000063.

Anonymous:2003:EBg

[Ano03g] Anonymous. Editorial Board. Technovation, 23(4):ifc, April 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203000166.

Anonymous:2003:EBh

[Ano03h] Anonymous. Editorial Board. Technovation, 23(5):ifc, May 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203000373.

Anonymous:2003:EBi

[Ano03i] Anonymous. Editorial Board. Technovation, 23(6):ifc, June 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649720300052X.

Anonymous:2003:EBj

[Ano03j] Anonymous. Editorial Board. Technovation, 23(7):ifc, July 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203000853.

Anonymous:2003:EBk

[Ano03k] Anonymous. Editorial Board. Technovation, 23(8):ifc, August 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203001019.

Anonymous:2003:EBl

[Ano03l] Anonymous. Editorial Board. Technovation, 23(9):ifc, September 2003. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203001202.

Anonymous:2003:EBb

[Ano03m] Anonymous. Editorial Board. Technovation, 23(10):ifc, October 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203001342.

Anonymous:2003:EBc

[Ano03n] Anonymous. Editorial Board. Technovation, 23(11):ifc, November 2003. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203001603.

Anonymous:2003:EBd

[Ano03o] Anonymous. Editorial Board. Technovation, 23(12):ifc, December 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203001822.

Anonymous:2003:EN

[Ano03p] Anonymous. Editorial note. Technovation, 23(1):iii, January 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497202001645.

Anonymous:2003:SIJ

[Ano03q] Anonymous. Special issue, Journal of Engineering and Technology Management (JET-M): Accounting Measurement, Incentives, and Technological Innovation. Technovation, 23(11):903–904, November 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0166497203001275.

Anonymous:2004:AI

[Ano04a] Anonymous. Author index. Technovation, 24(12):XII–XIII, December 2004. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497204001786.

Anonymous:2004:BRI

[Ano04b] Anonymous. Book review index. Technovation, 24(12):X– XI, December 2004. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497204001774.

Anonymous:2004:EBa

[Ano04c] Anonymous. Editorial Board. Technovation, 24(1):ifc, January 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203002086.

Anonymous:2004:EBe

[Ano04d] Anonymous. Editorial Board. Technovation, 24(2):ifc, February 2004. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203002244.

Anonymous:2004:EBf

[Ano04e] Anonymous. Editorial Board. Technovation, 24(3):ifc, March 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204000173.

Anonymous:2004:EBg

[Ano04f] Anonymous. Editorial Board. Technovation, 24(4):ifc, April 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204000458.

Anonymous:2004:EBh

[Ano04g] Anonymous. Editorial Board. Technovation, 24(5):ifc, May 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204000574.

Anonymous:2004:EBi

[Ano04h] Anonymous. Editorial Board. Technovation, 24(6):ifc, June 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204000756.

Anonymous:2004:EBj

[Ano04i] Anonymous. Editorial Board. Technovation, 24(7):ifc, July 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204000859.

Anonymous:2004:EBk

[Ano04j] Anonymous. Editorial Board. Technovation, 24(8):ifc, August 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204001014.

Anonymous:2004:EBl

[Ano04k] Anonymous. Editorial Board. Technovation, 24(9):ifc, September 2004. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497204001191.

Anonymous:2004:EBb

[Ano04l] Anonymous. Editorial Board. Technovation, 24(10):ifc, October 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204001464.

Anonymous:2004:EBc

[Ano04m] Anonymous. Editorial Board. Technovation, 24(11):ifc, November 2004. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497204001580.

Anonymous:2004:EBd

[Ano04n] Anonymous. Editorial Board. Technovation, 24(12):ifc, December 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204001683.

Anonymous:2004:PNa

[Ano04o] Anonymous. Publisher’s note. Technovation, 24(1):i, January 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203002165.

Anonymous:2004:PNb

[Ano04p] Anonymous. Publisher’s note. Technovation, 24(8):597, August 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204000926.

Anonymous:2004:VC

[Ano04q] Anonymous. Volume contents. Technovation, 24(12):III– IX, December 2004. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497204001762.

Anonymous:2005:EBa

[Ano05a] Anonymous. Editorial Board. Technovation, 25(1):ifc, January 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204001981.

Anonymous:2005:EBe

[Ano05b] Anonymous. Editorial Board. Technovation, 25(2):ifc, February 2005. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497204001919.

Anonymous:2005:EBf

[Ano05c] Anonymous. Editorial Board. Technovation, 25(3):ifc, March 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649720400224X.

Anonymous:2005:EBg

[Ano05d] Anonymous. Editorial Board. Technovation, 25(4):ifc, April 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204002329.

Anonymous:2005:EBh

[Ano05e] Anonymous. Editorial Board. Technovation, 25(5):ifc, May 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649720400238X.

Anonymous:2005:EBi

[Ano05f] Anonymous. Editorial Board. Technovation, 25(6):ifc, June 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204002457.

Anonymous:2005:EBj

[Ano05g] Anonymous. Editorial Board. Technovation, 25(7):ifc, July 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205000088.

Anonymous:2005:EBk

[Ano05h] Anonymous. Editorial Board. Technovation, 25(8):ifc, August 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205000143.

Anonymous:2005:EBl

[Ano05i] Anonymous. Editorial Board. Technovation, 25(9):ifc, September 2005. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497205000313.

Anonymous:2005:EBb

[Ano05j] Anonymous. Editorial Board. Technovation, 25(10):ifc, October 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205000374.

Anonymous:2005:EBc

[Ano05k] Anonymous. Editorial Board. Technovation, 25(11):ifc, November 2005. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497205000544.

Anonymous:2005:EBd

[Ano05l] Anonymous. Editorial Board. Technovation, 25(12):CO2, December 2005. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497205001355.

Anonymous:2006:CVA

[Ano06a] Anonymous. Contents of Volume 26, 2006 and Author Index. Technovation, 26(12):I–XVIII, December 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497206000988.

Anonymous:2006:EBa

[Ano06b] Anonymous. Editorial Board. Technovation, 26(1):CO2, January 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205001586.

Anonymous:2006:EBe

[Ano06c] Anonymous. Editorial Board. Technovation, 26(2):CO2, February 2006. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497206000046.

Anonymous:2006:EBf

[Ano06d] Anonymous. Editorial Board. Technovation, 26(3):CO2, March 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497206000137.

Anonymous:2006:EBg

[Ano06e] Anonymous. Editorial Board. Technovation, 26(4):CO2, April 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497206000198.

Anonymous:2006:EBh

[Ano06f] Anonymous. Editorial Board. Technovation, 26(5–6):CO2, May/June 2006. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497206000307.

Anonymous:2006:EBi

[Ano06g] Anonymous. Editorial Board. Technovation, 26(7):CO2, July 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497206000423.

Anonymous:2006:EBj

[Ano06h] Anonymous. Editorial Board. Technovation, 26(8):CO2, August 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497206000563.

Anonymous:2006:EBk

[Ano06i] Anonymous. Editorial Board. Technovation, 26(9):CO2, September 2006. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497206000630.

Anonymous:2006:EBb

[Ano06j] Anonymous. Editorial Board. Technovation, 26(10):CO2, October 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497206000691.

Anonymous:2006:EBc

[Ano06k] Anonymous. Editorial Board. Technovation, 26(11):CO2, November 2006. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497206000848.

Anonymous:2006:EBd

[Ano06l] Anonymous. Editorial Board. Technovation, 26(12):CO2, December 2006. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497206000782.

Anonymous:2006:PN

[Ano06m] Anonymous. Publisher’s Note. Technovation, 26(1):1, January 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205001677.

Anonymous:2007:EBa

[Ano07a] Anonymous. Editorial Board. Technovation, 27(1–2): CO2, January/February 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497206001180.

Anonymous:2007:EBe

[Ano07b] Anonymous. Editorial Board. Technovation, 27(3):CO2, March 2007. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497207000077.

Anonymous:2007:EBf

[Ano07c] Anonymous. Editorial Board. Technovation, 27(4):ifc, April 2007. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497207000296.

Anonymous:2007:EBg

[Ano07d] Anonymous. Editorial Board. Technovation, 27(5):ifc, May 2007. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497207000405.

Anonymous:2007:EBh

[Ano07e] Anonymous. Editorial Board. Technovation, 27(6–7):ifc, June/July 2007. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649720700051X.

Anonymous:2007:EBi

[Ano07f] Anonymous. Editorial Board. Technovation, 27(8):ifc, August 2007. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497207000855.

Anonymous:2007:EBj

[Ano07g] Anonymous. Editorial Board. Technovation, 27(9):ifc, September 2007. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497207000995.

Anonymous:2007:EBb

[Ano07h] Anonymous. Editorial Board. Technovation, 27(10):ifc, October 2007. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497207001113.

Anonymous:2007:EBc

[Ano07i] Anonymous. Editorial Board. Technovation, 27(11):ifc, November 2007. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497207001174.

Anonymous:2007:EBd

[Ano07j] Anonymous. Editorial Board. Technovation, 27(12):ifc, December 2007. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649720700123X.

Anonymous:2007:IIC

[Ano07k] Anonymous. IAMOT 2007: 16th International Conference on Management of Technology Miami Beach Resort & Spa, Miami Beach, Florida, USA May 13– 17, 2007. Technovation, 27(4):239, April 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000211.

Anonymous:2007:PGM

[Ano07l] Anonymous. Program guidelines for Management of Technology (MOT) Graduate Level Academic Institutions’ Accreditation Certification Certification. Technovation, 27(4): 236–238, April 2007. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497207000156.

Anonymous:2007:RMC

[Ano07m] Anonymous. The R&D Management Conference 2008, June 18–20, Ottawa, Canada. Technovation, 27(11):718–719, November 2007. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497207001289.

Anonymous:2008:A

[Ano08a] Anonymous. Announcement. Technovation, 28(10):711, October 2008. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497208001016.

Anonymous:2008:EBa

[Ano08b] Anonymous. Editorial Board. Technovation, 28(1–2): CO2, January/February 2008. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497207001411.

Anonymous:2008:EBe

[Ano08c] Anonymous. Editorial Board. Technovation, 28(3):ifc, March 2008. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497208000060.

Anonymous:2008:EBf

[Ano08d] Anonymous. Editorial Board. Technovation, 28(4):ifc, April 2008. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497208000254.

Anonymous:2008:EBg

[Ano08e] Anonymous. Editorial Board. Technovation, 28(5):ifc, May 2008. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497208000345.

Anonymous:2008:EBh

[Ano08f] Anonymous. Editorial Board. Technovation, 28(6):ifc, June 2008. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497208000461.

Anonymous:2008:EBi

[Ano08g] Anonymous. Editorial Board. Technovation, 28(7):ifc, July 2008. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497208000606.

Anonymous:2008:EBj

[Ano08h] Anonymous. Editorial Board. Technovation, 28(8):ifc, August 2008. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497208000680.

Anonymous:2008:EBk

[Ano08i] Anonymous. Editorial Board. Technovation, 28(9):ifc, September 2008. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497208000771.

Anonymous:2008:EBb

[Ano08j] Anonymous. Editorial Board. Technovation, 28(10):ifc, October 2008. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497208000965.

Anonymous:2008:EBc

[Ano08k] Anonymous. Editorial Board. Technovation, 28(11):ifc, November 2008. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497208001120.

Anonymous:2008:EBd

[Ano08l] Anonymous. Editorial Board. Technovation, 28(12):ifc, December 2008. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497208001247.

Anonymous:2009:EBa

[Ano09a] Anonymous. Editorial Board. Technovation, 29(1):ifc, January 2009. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497208001429.

Anonymous:2009:EBe

[Ano09b] Anonymous. Editorial Board. Technovation, 29(2):ifc, February 2009. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649720900008X.

Anonymous:2009:EBf

[Ano09c] Anonymous. Editorial Board. Technovation, 29(3):ifc, March 2009. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497209000261.

Anonymous:2009:EBg

[Ano09d] Anonymous. Editorial Board. Technovation, 29(4):ifc, April 2009. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497209000194.

Anonymous:2009:EBh

[Ano09e] Anonymous. Editorial Board. Technovation, 29(5):ifc, May 2009. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497209000339.

Anonymous:2009:EBi

[Ano09f] Anonymous. Editorial Board. Technovation, 29(6–7):ifc, June/July 2009. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497209000510.

Anonymous:2009:EBj

[Ano09g] Anonymous. Editorial Board. Technovation, 29(8):ifc, August 2009. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497209000662.

Anonymous:2009:EBk

[Ano09h] Anonymous. Editorial Board. Technovation, 29(9):ifc, September 2009. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497209000935.

Anonymous:2009:EBb

[Ano09i] Anonymous. Editorial Board. Technovation, 29(10):ifc, October 2009. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497209001126.

Anonymous:2009:EBc

[Ano09j] Anonymous. Editorial Board. Technovation, 29(11):ifc, November 2009. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497209001266.

Anonymous:2009:EBd

[Ano09k] Anonymous. Editorial Board. Technovation, 29(12):ifc, December 2009. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497209001515.

Anonymous:2009:ITL

[Ano09l] Anonymous. IAMOT — top 50 list of researchers in the technology innovation management field. Technovation, 29(4):235–236, April 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497209000066.

Anonymous:2010:EBa

[Ano10a] Anonymous. Editorial Board. Technovation, 30(1):ifc, January 2010. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497209001643.

Anonymous:2010:EBc

[Ano10b] Anonymous. Editorial Board. Technovation, 30(2):ifc, February 2010. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497210000039.

Anonymous:2010:EBd

[Ano10c] Anonymous. Editorial Board. Technovation, 30(3):ifc, March 2010. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497210000143.

Anonymous:2010:EBe

[Ano10d] Anonymous. Editorial Board. Technovation, 30(4):ifc, April 2010. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497210000234.

Anonymous:2010:EBf

[Ano10e] Anonymous. Editorial Board. Technovation, 30(5–6):ifc, May/June 2010. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497210000350.

Anonymous:2010:EBg

[Ano10f] Anonymous. Editorial Board. Technovation, 30(7–8):ifc, July/August 2010. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497210000532.

Anonymous:2010:EBh

[Ano10g] Anonymous. Editorial Board. Technovation, 30(9–10): ifc, September/October 2010. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497210000738.

Anonymous:2010:EBb

[Ano10h] Anonymous. Editorial Board. Technovation, 30(11–12): ifc, November/December 2010. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497210000970.

Anonymous:2011:TJT

[Ano11a] Anonymous. Technovation: the journal of technological innovation, entrepreneurship and technology management. Technovation, 31(8):349, August 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211000782.

Anonymous:2011:EBa

[Ano11b] Anonymous. Editorial Board. Technovation, 31(1):ifc, January 2011. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497210001203.

Anonymous:2011:EBd

[Ano11c] Anonymous. Editorial Board. Technovation, 31(2–3): ifc, February/March 2011. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497211000034.

Anonymous:2011:EBe

[Ano11d] Anonymous. Editorial Board. Technovation, 31(4):ifc, April 2011. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497211000265.

Anonymous:2011:EBf

[Ano11e] Anonymous. Editorial Board. Technovation, 31(5–6):ifc, May/June 2011. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497211000435.

Anonymous:2011:EBg

[Ano11f] Anonymous. Editorial Board. Technovation, 31(7):ifc, July 2011. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497211000629.

Anonymous:2011:EBh

[Ano11g] Anonymous. Editorial board. Technovation, 31(8):ifc, August 2011. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497211000800.

Anonymous:2011:EBi

[Ano11h] Anonymous. Editorial Board. Technovation, 31(9):ifc, September 2011. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497211000976.

Anonymous:2011:EBb

[Ano11i] Anonymous. Editorial Board. Technovation, 31(10–11): ifc, October/November 2011. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497211001131.

Anonymous:2011:EBc

[Ano11j] Anonymous. Editorial Board. Technovation, 31(12):ifc, December 2011. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497211001362.

Anonymous:2011:TEP

[Ano11k] Anonymous. Technology entreprenuership and policy: In memory of Bruce A. Kirchhoff. Technovation, 31(8):347– 348, August 2011. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497211000770.

Anonymous:2012:AOS

[Ano12a] Anonymous. Article omitted from the Special Issue “Special cluster: managing technology-service fusion innovation, Technovation”. Technovation, 32(7–8):440, July/August 2012. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497212000521.

Anonymous:2012:CP

[Ano12b] Anonymous. Call for papers. Technovation, 32(9–10): i, September/October 2012. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497212000910.

Anonymous:2012:EBa

[Ano12c] Anonymous. Editorial Board. Technovation, 32(1):ifc, January 2012. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649721100157X.

Anonymous:2012:EBd

[Ano12d] Anonymous. Editorial Board. Technovation, 32(2):ifc, February 2012. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497211001775.

Anonymous:2012:EBe

[Ano12e] Anonymous. Editorial Board. Technovation, 32(3–4):ifc, March/April 2012. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497212000089.

Anonymous:2012:EBf

[Ano12f] Anonymous. Editorial Board. Technovation, 32(5):ifc, May 2012. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497212000375.

Anonymous:2012:EBg

[Ano12g] Anonymous. Editorial Board. Technovation, 32(6):ifc, June 2012. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497212000478.

Anonymous:2012:EBh

[Ano12h] Anonymous. Editorial Board. Technovation, 32(7–8):ifc, July/August 2012. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497212000582.

Anonymous:2012:EBi

[Ano12i] Anonymous. Editorial Board. Technovation, 32(9–10): ifc, September/October 2012. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016649721200082X.

Anonymous:2012:EBb

[Ano12j] Anonymous. Editorial Board. Technovation, 32(11):ifc, November 2012. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497212000995.

Anonymous:2012:EBc

[Ano12k] Anonymous. Editorial Board. Technovation, 32(12):ifc, December 2012. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497212001083.

Anonymous:2013:EBa

[Ano13a] Anonymous. Editorial Board. Technovation, 33(1):ifc, January 2013. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497212001265.

Anonymous:2013:EBd

[Ano13b] Anonymous. Editorial Board. Technovation, 33(2–3): ifc, February/March 2013. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497213000059.

Anonymous:2013:EBe

[Ano13c] Anonymous. Editorial Board. Technovation, 33(4–5):ifc, April/May 2013. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497213000400.

Anonymous:2013:EBf

[Ano13d] Anonymous. Editorial Board. Technovation, 33(6–7):ifc, June/July 2013. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497213000539.

Anonymous:2013:EBg

[Ano13e] Anonymous. Editorial Board. Technovation, 33(8–9):ifc, August/September 2013. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497213000837.

Anonymous:2013:EBb

[Ano13f] Anonymous. Editorial Board. Technovation, 33(10–11): ifc, October/November 2013. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497213000953.

Anonymous:2013:EBc

[Ano13g] Anonymous. Editorial Board. Technovation, 33(12):ifc, December 2013. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649721300117X.

Anonymous:2013:PLA

[Ano13h] Anonymous. Professor Louis-André Lefebvre. Technovation, 33(10–11):386, October/November 2013. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213000783.

Anonymous:2014:EBa

[Ano14a] Anonymous. Editorial Board. Technovation, 34(1):ifc, January 2014. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497213001302.

Anonymous:2014:EBe

[Ano14b] Anonymous. Editorial Board. Technovation, 34(2):ifc, February 2014. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649721300151X.

Anonymous:2014:EBf

[Ano14c] Anonymous. Editorial Board. Technovation, 34(3):ifc, March 2014. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497214000121.

Anonymous:2014:EBg

[Ano14d] Anonymous. Editorial Board. Technovation, 34(4):ifc, April 2014. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497214000273.

Anonymous:2014:EBh

[Ano14e] Anonymous. Editorial Board. Technovation, 34(5–6):ifc, May/June 2014. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497214000418.

Anonymous:2014:EBi

[Ano14f] Anonymous. Editorial Board. Technovation, 34(7):ifc, July 2014. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497214000571.

Anonymous:2014:EBj

[Ano14g] Anonymous. Editorial Board. Technovation, 34(8):ifc, August 2014. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649721400073X.

Anonymous:2014:EBk

[Ano14h] Anonymous. Editorial Board. Technovation, 34(9):ifc, September 2014. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497214000868.

Anonymous:2014:EBb

[Ano14i] Anonymous. Editorial Board. Technovation, 34(10):ifc, October 2014. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497214001102.

Anonymous:2014:EBc

[Ano14j] Anonymous. Editorial Board. Technovation, 34(11):ifc, November 2014. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497214001321.

Anonymous:2014:EBd

[Ano14k] Anonymous. Editorial Board. Technovation, 34(12):ifc, December 2014. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649721400145X.

Anonymous:2015:EBa

[Ano15a] Anonymous. Editorial Board. Technovation, 35(??):ifc, January 2015. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497214001606.

Anonymous:2015:EBb

[Ano15b] Anonymous. Editorial Board. Technovation, 36–37(??): ifc, February/March 2015. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497214001758.

Anonymous:2015:EBc

[Ano15c] Anonymous. Editorial Board. Technovation, 38(??):ifc, April 2015. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649721500005X.

Anonymous:2015:EBd

[Ano15d] Anonymous. Editorial Board. Technovation, 39–40(??):ifc, May/June 2015. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649721500022X.

Anonymous:2015:EBe

[Ano15e] Anonymous. Editorial Board. Technovation, 41–42(??):ifc, July/August 2015. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497215000334.

Anonymous:2015:EBf

[Ano15f] Anonymous. Editorial Board. Technovation, 43–44(??): ifc, September/October 2015. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497215000504.

Anonymous:2015:EBg

[Ano15g] Anonymous. Editorial Board. Technovation, 45–46(??): ifc, November/December 2015. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497215000668.

Anonymous:2016:EBa

[Ano16a] Anonymous. Editorial Board. Technovation, 47(??):ifc, January 2016. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497215000851.

Anonymous:2016:EBb

[Ano16b] Anonymous. Editorial Board. Technovation, 48–49(??): ifc, February/March 2016. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497216000110.

Anonymous:2016:EBc

[Ano16c] Anonymous. Editorial Board. Technovation, 50–51(??):ifc, April/May 2016. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497216000201.

Anonymous:2016:EBd

[Ano16d] Anonymous. Editorial Board. Technovation, 52–53(??):ifc, June/July 2016. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497216300219.

Anonymous:2016:EBe

[Ano16e] Anonymous. Editorial Board. Technovation, 54(??):ifc, August 2016. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497216300578.

Anonymous:2016:EBf

[Ano16f] Anonymous. Editorial Board. Technovation, 55–56(??): ifc, September/October 2016. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497216301262.

Anonymous:2016:EBg

[Ano16g] Anonymous. Editorial Board. Technovation, 57–58(??): ifc, November/December 2016. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497216303066.

Anonymous:2016:IFS

[Ano16h]  Anonymous.  Innovation and de facto standardiza-

tion: the influence of dominant design on innovative performance, radical innovation, and process innovation. Technovation, 50–51(??):iii, April/May 2016. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497216000250.

Anonymous:2017:EBa

[Ano17a] Anonymous. Editorial Board. Technovation, 59(??):ifc, January 2017. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497216304503.

Anonymous:2017:EBb

[Ano17b] Anonymous. Editorial Board. Technovation, 60–61(??):ifc, February 2017. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497217302699.

Anonymous:2017:EBc

[Ano17c] Anonymous. Editorial Board. Technovation, 62–63(??):ifc, April 2017. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497217304388.

Anonymous:2017:EBd

[Ano17d] Anonymous. Editorial Board. Technovation, 64–65(??):ifc, June 2017. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497217305515.

Anonymous:2017:EBe

[Ano17e] Anonymous. Editorial Board. Technovation, 66–67(??):ifc, August 2017. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497217307277.

Anonymous:2017:EBf

[Ano17f] Anonymous. Editorial Board. Technovation, 68(??):ifc, December 2017. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497217308313.

Anonymous:2018:EBa

[Ano18a] Anonymous. Editorial Board. Technovation, 69(??):ifc, January 2018. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497217308696.

Anonymous:2018:EBb

[Ano18b] Anonymous. Editorial board. Technovation, 70–71(??):ifc, ???? 2018. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497218302591.

Anonymous:2018:EBc

[Ano18c] Anonymous. Editorial board. Technovation, 72–73 (??):ifc, April/May 2018. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497218303018.

Anonymous:2018:EBd

[Ano18d] Anonymous. Editorial Board. Technovation, 74–75 (??):ifc, June/July 2018. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497218304140.

Anonymous:2018:EBe

[Ano18e] Anonymous. Editorial Board. Technovation, 78(??):ifc, December 2018. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497218306746.

Anonymous:2019:EBa

[Ano19a] Anonymous. Editorial Board. Technovation, 76–77(??): ifc, August/September 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497218305170.

Anonymous:2019:EBb

[Ano19b] Anonymous. Editorial Board. Technovation, 79(??):ifc, January 2019. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497218307600.

Anonymous:2019:EBc

[Ano19c] Anonymous. Editorial Board. Technovation, 80–81(??): ifc, February/March 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497219300057.

Anonymous:2019:EBd

[Ano19d] Anonymous. Editorial Board. Technovation, 82–83 (??):ifc, April/May 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497219300173.

Anonymous:2019:EBe

[Ano19e] Anonymous. Editorial Board. Technovation, 84–85 (??):ifc, June/July 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497219300343.

Anonymous:2019:EBf

[Ano19f] Anonymous. Editorial Board. Technovation, 86–87(??): ifc, August/September 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497219300446.

Anonymous:2019:EBg

[Ano19g] Anonymous. Editorial Board. Technovation, 88(??):Article 102086, December 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497219300513.

Anonymous:2019:PN

[Ano19h] Anonymous. Publisher’s note. Technovation, 88(??):Article 102092, December 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497219300574.

Anonymous:2020:D

[Ano20a] Anonymous. December 2020. Technovation, 98(??):??, December 2020. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2020:EBa

[Ano20b] Anonymous. Editorial Board. Technovation, 89(??):??, January 2020. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497219300653.

Anonymous:2020:EBb

[Ano20c] Anonymous. Editorial Board. Technovation, 98(??):??, December 2020. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497220300547.

Anonymous:2020:J

[Ano20d] Anonymous. January 2020. Technovation, 89(??):??, January 2020. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2021:Aa

[Ano21a] Anonymous. April 2021. Technovation, 102(??):??, April 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2021:AS

[Ano21b] Anonymous. August–September 2020. Technovation, ?? (??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2021:Ab

[Ano21c] Anonymous. August 2021. Technovation, 106(??):??, August 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2021:D

[Ano21d] Anonymous. December 2021. Technovation, 108(??):??, December 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2021:EBa

[Ano21e] Anonymous. Editorial Board. Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497220300055.

Anonymous:2021:EBb

[Ano21f] Anonymous. Editorial Board. Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497220300183.

Anonymous:2021:EBc

[Ano21g] Anonymous. Editorial Board. Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497220300286.

Anonymous:2021:EBd

[Ano21h] Anonymous. Editorial Board. Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497220300390.

Anonymous:2021:EBe

[Ano21i] Anonymous. Editorial Board. Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497220300730.

Anonymous:2021:EBf

[Ano21j] Anonymous. Editorial Board. Technovation, 100(??):??, February 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497221000080.

Anonymous:2021:EBg

[Ano21k] Anonymous. Editorial Board. Technovation, 101(??):??, March 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221000237.

Anonymous:2021:EBh

[Ano21l] Anonymous. Editorial Board. Technovation, 102(??):??, April 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221000432.

Anonymous:2021:EBi

[Ano21m] Anonymous. Editorial Board. Technovation, 103(??):??, May 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221000596.

Anonymous:2021:EBj

[Ano21n] Anonymous. Editorial Board. Technovation, 104(??):??, June 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221000778.

Anonymous:2021:EBk

[Ano21o] Anonymous. Editorial Board. Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221000985.

Anonymous:2021:EBl

[Ano21p] Anonymous. Editorial Board. Technovation, 106(??):??, August 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221001218.

Anonymous:2021:EBm

[Ano21q] Anonymous. Editorial Board. Technovation, 107(??):??, September 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497221001395.

Anonymous:2021:EBn

[Ano21r] Anonymous. Editorial Board. Technovation, 108(??):??, December 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497221001814.

Anonymous:2021:FM

[Ano21s] Anonymous. February–March 2020. Technovation, ??(??): ??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2021:F

[Ano21t] Anonymous. February 2021. Technovation, 100(??):??, February 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2021:J

[Ano21u] Anonymous. June 2021. Technovation, 104(??):??, June 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2021:Ma

[Ano21v] Anonymous. March 2021. Technovation, 101(??):??, March 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2021:Mb

[Ano21w] Anonymous. May 2021. Technovation, 103(??):??, May 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2021:S

[Ano21x] Anonymous. September 2021. Technovation, 107(??):??, September 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2022:Aa

[Ano22a] Anonymous. April 2022. Technovation, 112(??):??, April 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2022:Ab

[Ano22b] Anonymous. August 2022. Technovation, 116(??):??, August 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2022:D

[Ano22c] Anonymous. December 2022. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2022:EBa

[Ano22d] Anonymous. Editorial Board. Technovation, 109(??):??, January 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221002108.

Anonymous:2022:EBb

[Ano22e] Anonymous. Editorial Board. Technovation, 110(??):??, February 2022. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497222000116.

Anonymous:2022:EBc

[Ano22f] Anonymous. Editorial Board. Technovation, 111(??):??, March 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497222000360.

Anonymous:2022:EBd

[Ano22g] Anonymous. Editorial Board. Technovation, 112(??):??, April 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497222000505.

Anonymous:2022:EBe

[Ano22h] Anonymous. Editorial Board. Technovation, 113(??):??, May 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016649722200075X.

Anonymous:2022:EBf

[Ano22i] Anonymous. Editorial Board. Technovation, 114(??):??, June 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222001122.

Anonymous:2022:EBg

[Ano22j] Anonymous. Editorial Board. Technovation, 115(??):??, July 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222001183.

Anonymous:2022:EBh

[Ano22k] Anonymous. Editorial Board. Technovation, 116(??):??, August 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222001481.

Anonymous:2022:EBi

[Ano22l] Anonymous. Editorial Board. Technovation, 117(??): ??, September 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222001766.

Anonymous:2022:EB

[Ano22m] Anonymous. Editorial Board. Technovation, 118(??): ??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222002048.

Anonymous:2022:F

[Ano22n] Anonymous. February 2022. Technovation, 110(??):??, February 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2022:Ja

[Ano22o] Anonymous. January 2022. Technovation, 109(??):??, January 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2022:Jc

[Ano22p] Anonymous. July 2022. Technovation, 115(??):??, July 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2022:Jb

[Ano22q] Anonymous. June 2022. Technovation, 114(??):??, June 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2022:Ma

[Ano22r] Anonymous. March 2022. Technovation, 111(??):??, March 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2022:Mb

[Ano22s] Anonymous. May 2022. Technovation, 113(??):??, May 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2022:S

[Ano22t] Anonymous. September 2022. Technovation, 117(??):??, September 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2023:Aa

[Ano23a] Anonymous. April 2023. Technovation, 122(??):??, April 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2023:Ab

[Ano23b] Anonymous. August 2023. Technovation, 126(??):??, August 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2023:D

[Ano23c] Anonymous. December 2023. Technovation, 128(??):??, December 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2023:EBa

[Ano23d] Anonymous. Editorial Board. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222002206.

Anonymous:2023:EBb

[Ano23e] Anonymous. Editorial Board. Technovation, 120(??):??, February 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223000056.

Anonymous:2023:EBc

[Ano23f] Anonymous. Editorial Board. Technovation, 121(??):??, March 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223000378.

Anonymous:2023:EBd

[Ano23g] Anonymous. Editorial Board. Technovation, 122(??):??, April 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223000135.

Anonymous:2023:EBe

[Ano23h] Anonymous. Editorial Board. Technovation, 123(??):??, May 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649722300086X.

Anonymous:2023:EBf

[Ano23i] Anonymous. Editorial Board. Technovation, 124(??):??, June 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223001086.

Anonymous:2023:EBg

[Ano23j] Anonymous. Editorial Board. Technovation, 125(??):??, July 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223001165.

Anonymous:2023:EBh

[Ano23k] Anonymous. Editorial Board. Technovation, 126(??):??, August 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223001505.

Anonymous:2023:EBi

[Ano23l] Anonymous. Editorial Board. Technovation, 127(??):??, September 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223001736.

Anonymous:2023:EBj

[Ano23m] Anonymous. Editorial Board. Technovation, 128(??):??, December 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223002237.

Anonymous:2023:F

[Ano23n] Anonymous. February 2023. Technovation, 120(??):??, February 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2023:Ja

[Ano23o] Anonymous. January 2023. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2023:Jc

[Ano23p] Anonymous. July 2023. Technovation, 125(??):??, July 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2023:Jb

[Ano23q] Anonymous. June 2023. Technovation, 124(??):??, June 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2023:Ma

[Ano23r] Anonymous. March 2023. Technovation, 121(??):??, March 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2023:Mb

[Ano23s] Anonymous. May 2023. Technovation, 123(??):??, May 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2023:S

[Ano23t] Anonymous. September 2023. Technovation, 127(??):??, September 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2024:Aa

[Ano24a] Anonymous. April 2024. Technovation, 132(??):??, April 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2024:EBa

[Ano24b] Anonymous. Editorial Board. Technovation, 129(??):??, January 2024. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223002420.

Anonymous:2024:EBb

[Ano24c] Anonymous. Editorial Board. Technovation, 130(??):??, February 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497224000129.

Anonymous:2024:EBc

[Ano24d] Anonymous. Editorial Board. Technovation, 131(??):??, March 2024. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497224000300.

Anonymous:2024:EBd

[Ano24e] Anonymous. Editorial Board. Technovation, 132(??):??, April 2024. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497224000579.

Anonymous:2024:EBe

[Ano24f] Anonymous. Editorial Board. Technovation, 133(??):??, May 2024. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497224000907.

Anonymous:2024:EBf

[Ano24g] Anonymous. Editorial Board. Technovation, 134(??):??, June 2024. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497224001081.

Anonymous:2024:F

[Ano24h] Anonymous. February 2024. Technovation, 130(??):??, February 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2024:Ja

[Ano24i] Anonymous. January 2024. Technovation, 129(??):??, January 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2024:Jb

[Ano24j] Anonymous. June 2024. Technovation, 134(??):??, June 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2024:Ma

[Ano24k] Anonymous. March 2024. Technovation, 131(??):??, March 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Anonymous:2024:Mb

[Ano24l] Anonymous. May 2024. Technovation, 133(??):??, May 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic).

Adekalu:2003:CRS

[AO03] K. O. Adekalu and L. A. O. Ogunjimi. Cost recovery strategy for large-scale irrigation projects in Nigeria. Technovation, 23(1):77–83, January 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000578.

AlcidesGobbo:2010:TBE

[AO10] José Alcides Gobbo, Jr. and Annika Olsson. The transformation between exploration and exploitation applied to inventors of packaging innovations. Technovation, 30(5– 6):322–331, May/June 2010. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497210000106.

Adekalu:2002:WSD

[AOO02] K. O. Adekalu, J. A. Osunbitan, and O. E. Ojo. Water sources and demand in South Western Nigeria: implications for water development planners and scientists. Technovation, 22(12):799–805, December 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000566.

Archibugi:1996:MTC

[AP96] Daniele Archibugi and Mario Planta. Measuring technological change through patents and innovation surveys. Technovation, 16(9):451–468, 519, September 1996. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497296000314.

Antoncic:2008:ACT

[AP08] Bostjan Antoncic and Igor Prodan. Alliances, corporate technological entrepreneurship and firm performance: Testing a model on manufacturing firms. Technovation, 28(5): 257–265, May 2008. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497207000946.

Arundel:1999:DEB

[AR99] Anthony Arundel and Antoine Rose. The diffusion of environmental biotechnology in Canada: adoption strategies and cost offsets. Technovation, 19(9):551–560, September 1999. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497299000589.

Aramand:2008:SPS

[Ara08] Majid Aramand. Software products and services are high tech? New product development strategy for software products and services. Technovation, 28(3):154–160, March 2008. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497207001368.

Archibugi:1988:IID

[Arc88] Daniele Archibugi. The inter-industry distribution of technological capabilities. A case study in the application of the Italian patenting in the USA. Technovation, 7(3):259–274, July 1988. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497288900247.

Archibugi:2003:BRG

[Arc03] Daniele Archibugi. Book review: The global challenge to industrial districts. Small and medium-sized enterprises in Italy and Taiwan: edited by Paolo Guerrieri, Simona Iammarino and Carlo Pietrobelli; Edward Elgar, Cheltenham, 2002. Technovation, 23(11):901–902, November 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203000804.

Arnold:1987:SLG

[Arn87] Erik Arnold. Some lessons from government information technology policies. Technovation, 5(4):247–268, February 1987. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497287900666.

Arroio:2004:BRC

[Arr04] Ana Arroio. Book review: China’s Telecommunication Market — Entering a New Competitive Age: Ding Lu and Chee Kong Wong (Eds.); Edward Elgar, 2003, 177 pages (incl. Index), hardback, ISBN 1-84064-431-1. Technovation, 24(11):923–925, November 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001348.

Asimakopoulos:2023:IRS

[ARR23] Grigorios Asimakopoulos, Antonio Revilla, and Alicia Rodríguez. International R&D sourcing, innovation and firm age: the advantage of ‘born-international sourcers’. Technovation, 122(??):??, April 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222002139.

Avila-Robinson:2017:MER

[ÁRS17] Alfonso Ávila-Robinson and Shintaro Sengoku. Multilevel exploration of the realities of interdisciplinary research centers for the management of knowledge integration. Technovation, 62–63(??):22–41, April 2017. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497217300093.

Alfranca:2004:ISM

[ARvT04] Oscar Alfranca, Ruth Rama, and Nicholas von Tunzelmann. Innovation spells in the multinational agri-food sector. Technovation, 24(8):599–614, August 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202001293.

Abrunhosa:2008:TPS

[AS08] Ana Abrunhosa and Patrícia Moura E. Sá. Are TQM principles supporting innovation in the Portuguese footwear industry? Technovation, 28(4):208–221, April 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207001083.

Aoki:2018:MCD

[AS18] Katsuki Aoki and Thomas Staeblein. Monozukuri capability and dynamic product variety: An analysis of the design-manufacturing interface at Japanese and German automakers. Technovation, 70–71(??):33–45, ???? 2018. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497217307800.

Aaldering:2021:LLT

[AS21] Lukas Jan Aaldering and Chie Hoon Song. Of leaders and laggards — towards digitalization of the process industries. Technovation, ??(??):??, ???? 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497220300833.

Aman:2022:IIR

[AS22] Sadaf Aman and Stefan Seuring. Interestingly it’s innovation: Reviewing sustainability performance management in the base of the pyramid (BoP). Technovation, 112(??): ??, April 2022. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497221001759.

Ali:2022:NRK

[ASA^{+}22] Murad Ali, Muhammad Shujahat, Zulfiqar Ali, Aino Kianto, Minhong Wang, and Nick Bontis. The neglected role of knowledge assets interplay in the pursuit of organisational ambidexterity. Technovation, 114(??):??, June 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497221002339.

Al-Somali:2009:IAO

[ASGC09] Sabah Abdullah Al-Somali, Roya Gholami, and Ben Clegg. An investigation into the acceptance of online banking in Saudi Arabia. Technovation, 29(2):130–141, February 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208000850.

Adetayo:1999:IIT

[ASI99] J. O. Adetayo, S. A. Sanni, and M. O. Ilori. The impact of information technology on product marketing: a case study of a multinational company in Nigeria. Technovation, 19(11):691–699, November 1999. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299000516.

Artamoshina:2023:ASC

[ASOB23] Polina Artamoshina, Galina Shirokova, Oleksiy Osiyevskyy, and Virginia Bodolica. ADHD symptoms of CEOs and business model innovation in the SME context. Technovation, 128(??):??, December 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497223001566.

Akhtar:1993:FMS

[AT93] Khalid Akhtar and Mario T. Tabucanon. A framework for manufacturing strategy analysis in the wake of reducing product life-cycles: a case of a multinational in a newly industrializing count. Technovation, 13(5):265–281, July 1993. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729390001C.

Akanbi:1999:EAP

[AT99] C. T. Akanbi and K. A. Taiwo. Economic analysis of the production of ready-to-eat ‘Gbegiri’ bean mix powder. Technovation, 19(12):747–753, December 1999. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299000760.

Augsdorfer:2006:BRB

[Aug06a] Peter Augsdorfer. Book review: Breakthrough: Mark Stefik, Barbara Stefik. MIT Press, London (2004), ISBN 0-262-19514-3. Technovation, 26(3):418, March 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000787.

Augsdorfer:2006:BRI

[Aug06b] Peter Augsdörfer. Book review: Intellectual property and innovation management in small firms: Robert A. Blackburn (Ed.). Routledge, London (2003), ISBN 0-415-22884-0. Technovation, 26(1):135, January 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000246.

Auger:2009:BRS

[Aug09] Jean-François Auger. Book review: S. Shane (Ed.). Economic Development Through Entrepreneurship: Government, University, and Business Linkages. Edward Elgar Pub., Cheltenham (2006). 252 pp., Price: £59.95, ISBN 1-84376-855-0. Technovation, 29(1):74–75, January 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208000588.

Autio:1994:NTB

[Aut94] Erkko Autio. New, technology-based firms as agents of R&D and innovation: an empirical study. Technovation, 14(4):259–273, May 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497294900108.

Atzmon:2022:SFG

[AVA22] Morane Batya Atzmon, Johanna Vanderstraeten, and Sascha Albers. Small-firm growth-enabling capabilities: a framework for young technology-based firms. Technovation, 117(??):??, September 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016649722200089X.

Allarakhia:2011:MKA

[AW11] Minna Allarakhia and Steven Walsh. Managing knowledge assets under conditions of radical change: the case of the pharmaceutical industry. Technovation, 31(2–3):105– 117, February/March 2011. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497210001185.

Allarakhia:2012:AON

[AW12] Minna Allarakhia and Steven Walsh. Analyzing and organizing nanotechnology development: Application of the institutional analysis development framework to nanotechnology consortia. Technovation, 32(3–4):216–226, March/ April 2012. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497211001532.

Abbott:2005:STD

[AY05] Julie Abbott and Pak Yoong. The stages of telecentre development: the case of the Kapiti Telecentre. Technovation, 25(4):421–431, April 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203001469.

Ayres:1988:BBE

[Ayr88] Robert U. Ayres. Barriers and breakthroughs: an “expanding frontiers” model of the technology–industry life cycle. Technovation, 7(2):87–115, May 1988. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900417.

Ayres:1990:TIC

[AZ90] Robert Ayres and Ehud Zuscovitch. Technology and information: chain reactions and sustainable economic growth. Technovation, 10(3):163–183, May 1990. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729090022C.

Byun:1989:GKS

[BA89] Byung-Moon Byun and Byong-Hun Ahn. Growth of the Korean semiconductor industry and its competitive strategy in the world market. Technovation, 9(8):635–656, December 1989. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497289900047.

Bosse:2010:BPA

[BA10] Douglas A. Bosse and Sharon A. Alvarez. Bargaining power in alliance governance negotiations: evidence from the biotechnology industry. Technovation, 30(5–6):367–375, May/June 2010. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649721000012X.

Bachmann:1998:PCT

[Bac98] Astrid Bachmann. Profiles of corporate technological capabilities — a comparison of large British and German pharmaceutical companies. Technovation, 18(10):593–604, 655– 656, October 1998. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497298000534.

Bader:2008:MIP

[Bad08] Martin A. Bader. Managing intellectual property in the financial services industry sector: Learning from Swiss Re. Technovation, 28(4):196–207, April 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000685.

Badawy:2011:OIF

[Bad11] Michael K. Badawy. “Is open innovation a field of study or a communication barrier to theory development?”: a perspective. Technovation, 31(1):65–67, January 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210001082.

Bailey:1984:BRI

[Bai84] John E. Bailey. Book review: Industrial robots: a summary and forecast: Published by Tech Tran Corporation: Naperville, IL, second edition, 1983, 240 pages, price US$50.00 (in USA, Canada, Mexico), US$65.00 (elsewhere). Technovation, 2(2):148, May 1984. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497284900191.

Baiyere:2018:FIE

[Bai18] Abayomi Baiyere. Fostering innovation ecosystems — note on the 2017 ISPIM Innovation Forum. Technovation, 69(??):1, January 2018. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497217308817.

Balabkins:1985:RIP

[Bal85] Nicholas W. Balabkins. Repressed inflation and pseudoinnovations. Technovation, 3(3):187–197, August 1985. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497285900136.

Baldini:2010:DRR

[Bal10] Nicola Baldini. Do royalties really foster university patenting activity? An answer from Italy. Technovation, 30 (2):109–116, February 2010. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497209001448.

Bamane:1994:ITC

[Bam94] B. D. Bamane. Impact of technological change. Technovation, 14(1):3–5, February 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497294900655.

Banerjee:2003:RDC

[Ban03a] P. Banerjee. Resource dependence and core competence: insights from Indian software firms. Technovation, 23(3):251– 263, March 2003. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497201001201.

Banerjee:2003:SID

[Ban03b] P. Banerjee. Some indicators of dynamic technological competencies: understanding of Indian software managers. Technovation, 23(7):593–602, July 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720200007X.

Banerjee:2004:BRE

[Ban04] Bobby Banerjee. Book review: Environmental policy and technological innovation: Why do firms adopt or reject new technologies?: Carlos Montalvo Corral; Edward Elgar, Cheltenham UK, 2002, ISBN 1-84064-957-7. Technovation, 24(7):595, July 2004. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497204000641.

Barbiroli:1989:ETI

[Bar89a] Giancarlo Barbiroli. Elaborating a technological index for the evaluation of innovation quality. Technovation, 9(1): 101–107, May 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497289900436.

Barbiroli:1989:EGP

[Bar89b] Giancarlo Barbiroli. The elaboration of global performance indices for evaluating durable goods. Technovation, 9(1): 83–100, May 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497289900424.

Barbiroli:1990:NME

[Bar90a] Giancarlo Barbiroli. A new method to evaluate the specific and global advantage of a technology. Technovation, 10(2): 73–93, March 1990. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729090030N.

Baron:1990:FII

[Bar90b] P. J. Baron. Factors influencing innovation. Technovation, 10(6):379–387, September 1990. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497290900022.

Barbiroli:1992:TDD

[Bar92a] Giancarlo Barbiroli. Towards a definition and a dynamic measure of strategic technology. Technovation, 12(5):285– 296, July 1992. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729290037I.

Baron:1992:LCO

[Bar92b] Jonathan Baron. Linking companies with outside technology: an effective new approach. Technovation, 12(5):323– 332, July 1992. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729290040O.

Baranano:1995:SIF

[Bar95] Ana Maria Barañano. The Spanish innovative firm and the ESPRIT, RACE and EUREKA programmes: an organizational approach. Technovation, 15(6):339–350, August 1995. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729596595K.

Barbiroli:1996:NIM

[Bar96] Giancarlo Barbiroli. New indicators for measuring the manifold aspects of technical and economic efficiency of production processes and technologies. Technovation, 16(7):341– 356, July 1996. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497296000247.

Baruk:1997:IPE

[Bar97] Jerzy Baruk. Innovativeness of Polish enterprises in the initial period of system transformation. Technovation, 17(9):477–489, 529–530, September 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297000345.

Bartol:2014:CSC

[Bar14] Nadya Bartol. Cyber supply chain security practices DNA — Filling in the puzzle using a diverse set of disciplines. Technovation, 34(7):354–361, July 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214000066.

Barros:2015:EUP

[Bar15] Henrique M. Barros. Exploring the use of patents in a weak institutional environment: the effects of innovation partnerships, firm ownership, and new management practices. Technovation, 45–46(??):63–77, November/December 2015. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497215000449.

Bhatt:2016:SIO

[BAR16] Punita Bhatt, Ali J. Ahmad, and Muhammad Azam Roomi. Social innovation with open source software: User engagement and development challenges in India. Technovation, 52–53(??):28–39, June/July 2016. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497216000055.

Baskaran:2001:CBC

[Bas01] A. Baskaran. Competence building in complex systems in the developing countries: the case of satellite building in India. Technovation, 21(2):109–121, February 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000225.

Baskaran:2004:BRD

[Bas04]  Angathevar Baskaran.  Book review: The dynam-

ics of technology: Creation and diffusion of skills and knowledge: Roddam Narasimha, J. Srinivasan, S. K. Biswas (Eds.); SAGE Publications, New Delhi, 2003, 301 pages (including index), ISBN 0-7619-9670-2, hardbound. Technovation, 24(11):921–922, November 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001075.

Battard:2012:CMN

[Bat12] Nicolas Battard. Convergence and multidisciplinarity in nanotechnology: Laboratories as technological hubs. Technovation, 32(3–4):234–244, March/April 2012. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649721100126X.

Bocken:2012:DTR

[BAWK12] N. M. P. Bocken, J. M. Allwood, A. R. Willey, and J. M. H. King. Development of a tool for rapidly assessing the implementation difficulty and emissions benefits of innovations. Technovation, 32(1):19–31, January 2012. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211001301.

Brown:1981:E

[BB81] Wayne S. Brown and William R. Bowen. Editorial. Technovation, 1(1):1, February 1981. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649728190002X.

Bessant:1989:IIT

[BB89] John Bessant and Joanna Buckingham. Implementing integrated technology. Technovation, 9(4):321–336, July 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497287900046.

Berry:1995:BRT

[BB95] Aidan Berry and Guy Blomfield. Book review: Target costing — State of the art review: by Peter Horváth, S. Niemand and M. Wolbold. IFS International Ltd, Wolseley Business Park, Kempston, Bedford MK42 7PW, UK, 1995, 64 pp., ISBN 1-85907-004-3. £27.50 (paperback). Technovation, 15(3):189–190, April 1995. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497295900276.

Brantnell:2022:URI

[BB22] Anders Brantnell and Enrico Baraldi. Understanding the roles and involvement of technology transfer offices in the commercialization of university research. Technovation, 115(??):??, July 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222000724.

Barirani:2015:DRC

[BBA15] Ahmad Barirani, Catherine Beaudry, and Bruno Agard. Distant recombination and the creation of basic inventions: an analysis of the diffusion of public and private sector nanotechnology patents in Canada. Technovation, 36–37(??):39–52, February/March 2015. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214001370.

Barlow:2006:ICI

[BBC06] James Barlow, Steffen Bayer, and Richard Curry. Implementing complex innovations in fluid multi-stakeholder environments: Experiences of ‘telecare’. Technovation, 26 (3):396–406, March 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497205000970.

Barnes:2001:DMC

[BBDM01] Justin Barnes, John Bessant, Nikki Dunne, and Mike Morris. Developing manufacturing competitiveness within South African industry: the role of middle management. Technovation, 21(5):293–309, May 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720000047X.

BritoVinas:2001:CMD

[BBHA01] Beatriz Cristina Brito Viñas, John Bessant, Gilberto Hernández Pérez, and Arnaldo Alvarez González. A conceptual model for the development of technological management processes in manufacturing companies in developing countries. Technovation, 21(6):345–352, June 2001.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000523.

Bessant:1993:CIB

[BBHW93] John Bessant, Jane Burnell, Rebecca Harding, and Stephen Webb. Continuous improvement in British manufacturing. Technovation, 13(4):241–254, May 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390021M.

Bhattacharjya:2023:CCC

[BBK23] Bibhuti Ranjan Bhattacharjya, Saradindu Bhaduri, and Sashindra Kumar Kakoty. Co-creating community-led frugal innovation: an adapted Quadruple Helix? Technovation, 124(??):??, June 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497223000639.

Barbier:2024:DPI

[BBLW24] Raphaëlle Barbier, Skander Ben Yahia, Sylvain Lenfle, and Benoit Weil. Data-push innovation beyond serendipity: the case of a digital platform making Earth Observation data fit into multiple use contexts. Technovation, 132(??):??, April 2024. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497224000427.

Bedford:2022:EEK

[BBS22] David Bedford, Josep Bisbe, and Breda Sweeney. Enhancing external knowledge search: the influence of performance measurement system design on the absorptive capacity of top management teams. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016649722200133X.

Baglieri:2019:UTT

[BBT19] Daniela Baglieri, Francesco Baldi, and Christopher L. Tucci. University technology transfer office business models: One size does not fit all. Technovation, 76–77(??):

51–63, August/September 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497218303559.

Balthasar:2000:DKA

[BBTW00] Andreas Balthasar, Christoph Bättig, Alain Thierstein, and Beate Wilhelm. “Developers”: key actors of the innovation process. Types of developers and their contacts to institutions involved in research and development, continuing education and training, and the transfer of technology. Technovation, 20(10):523–538, October 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001807.

Benamar:2020:ITD

[BBZ20] Lamya Benamar, Christine Balagué, and Zeling Zhong. Internet of Things devices appropriation process: the Dynamic Interactions Value Appropriation (DIVA) framework. Technovation, 89(??):??, January 2020. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497218300221.

Bruno:1982:PDA

[BC82] Albert V. Bruno and Arnold C. Cooper. Patterns of development and acquisitions for Silicon Valley startups. Technovation, 1(4):275–290, August 1982. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497282900086.

Bozeman:1991:TTU

[BC91] Barry Bozeman and Michael Crow. Technology transfer from U.S. government and university R&D laboratories. Technovation, 11(4):231–246, May 1991. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497291900364.

Bozeman:1992:AET

[BC92] Barry Bozeman and Karen Coker. Assessing the effectiveness of technology transfer from US government R&D laboratories: the impact of market orientation. Technovation, 12(4):239–255, May 1992. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729290045J.

Boisot:1999:SFA

[BC99] Max Boisot and Benita Cox. The I-Space: a framework for analyzing the evolution of social computing. Technovation, 19(9):525–536, September 1999. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299000498.

Banerjee:2010:BIF

[BC10] Preeta M. Banerjee and Benjamin M. Cole. Breadthof-impact frontier: How firm-level decisions and selection environment dynamics generate boundary-spanning inventions. Technovation, 30(7–8):411–419, July/August 2010. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210000325.

Banerjee:2012:SBS

[BC12] Preeta M. Banerjee and Benjamin M. Cole. A study of biotechnology start-ups undergoing leadership change: Antecedents of change and endogenous performance consequences. Technovation, 32(9–10):568–578, September/ October 2012. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497212000557.

Bianchi:2011:OMO

[BCC^{+}11] Mattia Bianchi, Alberto Cavaliere, Davide Chiaroni, Federico Frattini, and Vittorio Chiesa. Organisational modes for Open Innovation in the bio-pharmaceutical industry: an exploratory analysis. Technovation, 31(1):22–33, January 2011. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497210000337.

Bessant:1994:RCI

[BCG^{+}94] J. Bessant, S. Caffyn, J. Gilbert, R. Harding, and S. Webb. Rediscovering continuous improvement. Technovation, 14 (1):17–29, February 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497294900671.

Bessant:2001:EMC

[BCG01] John Bessant, Sarah Caffyn, and Maeve Gallagher. An evolutionary model of continuous improvement behaviour. Technovation, 21(2):67–77, February 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000237.

Boucke:1994:TPG

[BCH94] Christian Boucke, Uwe Cantner, and Horst Hanusch. ‘Technopolises’ as a policy goal: a morphological study of the Wissenschaftsstadt Ulm. Technovation, 14(6):407–418, August 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294900191. See [Ano94-38].

Beecham:1998:PKT

[BCH98] M. A. Beecham and M. Cordey-Hayes. Partnering and knowledge transfer in the U.K. motor industry. Technovation, 18(3):191–205, March 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497297001132.

Bahuleyan:2024:PVR

[BCKC24] Athira Bahuleyan, Meena Chavan, Anna Krzeminska, and Francesco Chirico. Process and variance research: Integrating research on university spinoff evolution. Technovation, 130(??):??, February 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497223002316.

Bernal:2022:KSR

[BCL22] Pilar Bernal, Martin Carree, and Boris Lokshin. Knowledge spillovers, R&D partnerships and innovation performance. Technovation, 115(??):??, July 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://

www.sciencedirect.com/science/article/pii/S0166497222000037.

Baglieri:2012:RCS

[BCM12] Daniela Baglieri, Maria Cristina Cinici, and Vincent Mangematin. Rejuvenating clusters with ‘sleeping anchors’: the case of nanoclusters. Technovation, 32(3–4):245–256, March/April 2012. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497211001283.

Bartoloni:2022:TDS

[BCM^{+}22] Sara Bartoloni, Ernesto Calò, Luca Marinelli, Federica Pascucci, Luca Dezi, Elias Carayannis, Gian Marco Revel, and Gian Luca Gregori. Towards designing society 5.0 solutions: the new Quintuple Helix–Design Thinking approach to technology. Technovation, 113(??):??, May 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497221001942.

Baglieri:2014:DNP

[BCO14] Daniela Baglieri, Fabrizio Cesaroni, and Luigi Orsi. Does the nano-patent ‘gold rush’ lead to entrepreneurial-driven growth? Some policy lessons from China and Japan. Technovation, 34(12):746–761, December 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214001047.

Bonesso:2011:TSD

[BCP11] Sara Bonesso, Anna Comacchio, and Claudio Pizzi. Technology sourcing decisions in exploratory projects. Technovation, 31(10–11):573–585, October/November 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211000745.

Boyson:2022:DDS

[BCP22] Sandor Boyson, Thomas M. Corsi, and John-Patrick Paraskevas. Defending digital supply chains: Evidence from a decade-long research program. Technovation, 118 (??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497221001619.

Basile:2023:BIH

[BCPP23] Luigi Jesus Basile, Nunzia Carbonara, Roberta Pellegrino, and Umberto Panniello. Business intelligence in the healthcare industry: the utilization of a data-driven approach to support clinical decision making. Technovation, 120(??): ??, February 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497222000293.

Basile:2024:RBD

[BCPP24] L. J. Basile, N. Carbonara, U. Panniello, and R. Pellegrino. The role of big data analytics in improving the quality of healthcare services in the Italian context: the mediating role of risk management. Technovation, 133(??):??, May 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000609.

Bayona:2006:TAM

[BCS06] Cristina Bayona, Pilar Corredor, and Rafael Santamaría. Technological alliances and the market valuation of new economy firms. Technovation, 26(3):369–383, March 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001397.

Bessagnet:2021:UMS

[BCV21] Arnauld Bessagnet, Joan Crespo, and Jérôme Vicente. Unraveling the multi-scalar and evolutionary forces of entrepreneurial ecosystems: a historical event analysis applied to IoT Valley. Technovation, 108(??):??, December 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001103.

Brennan:2005:NCS

[BD05] A. Brennan and L. Dooley. Networked creativity: a structured management framework for stimulating innovation. Technovation, 25(12):1388–1399, December 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001373.

Belezas:2023:ISE

[BD23] Fernando Belezas and Ana Dias Daniel. Innovation in the sharing economy: a systematic literature review and research framework. Technovation, 122(??):??, April 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222000566.

Barrett:2021:OIW

[BDB21] Gillian Barrett, Lawrence Dooley, and Joe Bogue. Open innovation within high-tech SMEs: a study of the entrepreneurial founder’s influence on open innovation practices. Technovation, 103(??):??, May 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000134.

Boulton:1992:TDS

[BDL92] William R. Boulton, Michael J. Dowling, and Jurgen Lohmeyer. Technology development strategies in Japan, Europe and the United States. Technovation, 12(2):99–118, March 1992. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729290018D.

Bigliardi:2006:ASP

[BDNP06] Barbara Bigliardi, Alberto Ivo Dormio, Anna Nosella, and Giorgio Petroni. Assessing science parks’ performances: directions from selected Italian case studies. Technovation, 26(4):489–505, April 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000039.

Borjesson:2006:ISE

[BDW06] Sofia Börjesson, Fredrik Dahlsten, and Mats Williander. Innovative scanning experiences from an idea generation project at Volvo Cars. Technovation, 26(7):775–783, July 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205000064.

Berg:2009:RNI

[BE09] Daniel Berg and Norman G. Einspruch. Research note: Intellectual property in the services sector: Innovation and technology management implications. Technovation, 29(5): 387–393, May 2009. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497208001363.

Beaumont:1998:IDA

[Bea98] N. B. Beaumont. Investment decisions in Australian manufacturing. Technovation, 18(11):689–695, November 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000662.

Beaumont:2005:BPA

[Bea05] Nicholas Beaumont. Best Practice in Australian manufacturing sites. Technovation, 25(11):1291–1297, November 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204001543.

Beckers:1989:HB

[Bec89] Harry L. Beckers. Harry L. Beckers. Technovation, 9(2–3): 303–306, June 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497289900771.

Becker:2024:GOP

[BEE24] Kai Becker, Joris J. Ebbers, and Yuval Engel. Going online: Peer entrepreneur networks in a startup accelerator before and during the COVID-19 pandemic. Technovation, 130(??):??, February 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497223002286.

Brockhoff:1999:GPL

[BEH99] Klaus K. Brockhoff, Holger Ernst, and Eckhard Hundhausen. Gains and pains from licensing-patent-portfolios as strategic weapons in the cardiac rhythm management industry. Technovation, 19(10):605–614, October 1999.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299000656.

Bernardi:1989:IC

[Ber89] Luigi Rossi Bernardi. Introduction by the chairman. Technovation, 9(2–3):211–212, June 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/016649728990059X.

Berry:1996:TES

[Ber96] M. M. J. Berry. Technical entrepreneurship, strategic awareness and corporate transformation in small high-tech firms. Technovation, 16(9):487–498, 521–522, September 1996. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497296000569.

Bessant:1988:BRI

[Bes88] John Bessant. Book review: Innovation: a challenge to the engineer: By K. Holt, Elsevier Science Publishers, 1988, xiv + 452 pages, Dfl. 180.00, ISBN 0-444-42865-8. Technovation, 7(4):419–420, October 1988. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900119.

Bessant:1993:BRC

[Bes93] John Bessant. Book review: Company and campus partnership: by D. Jane Bower. Routledge, 1992, ISBN 0-415-07080-5, 197 pages, £35.00 (hardback). Technovation, 13 (5):331, July 1993. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729390006H.

Bessant:1994:BRC

[Bes94a] John Bessant. Book review: Competitive manufacturing: by Rory L. Chase (ed.). Second edition, 1994, 101 pp., ISBN 1-85907-005-1, £22.95. Both published by IPS International Ltd, Wolseley Business Park, Kempston, Bedford MK42 7PW, UK. Technovation, 14(9):625–626, November 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900493.

Bessant:1994:BRN

[Bes94b] John Bessant. Book review: New partnerships for managing technological change: by Nancy H. Bancroft. John Wiley, 1992, ISBN 0-471-54674-7, 268 pp. Technovation, 14(2): 137, March 1994. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497294901082.

Bessant:1994:BRSa

[Bes94c] John Bessant. Book review: Strategic technology management: Integrating product technology into global business strategies for the 1990s: by P. Dussauge, S. Hart and B. Ramanantsoa. John Wiley, 1992, ISBN 0-471-93418-6, 218 pp. Technovation, 14(2):137–138, March 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294901090.

Bessant:1994:BRSb

[Bes94d] John Bessant. Book review: The strategic toolkit: by John Marsh. 1993, 118 pp., ISBN 1-85907-002-7, £29.95. Technovation, 14(9):625, November 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900485.

Bessant:1995:BRC

[Bes95a] John Bessant. Book review: Competitive advantage through people: by Jeffrey Pfeffer. Harvard Business School Press, Soldiers Field Road, Boston, MA 02163, USA, 1994, 288 pp., ISBN 0-87584-413-8. $24.95. Technovation, 15(1):49– 50, February 1995. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497295900055.

Bessant:1995:BRE

[Bes95b] John Bessant. Book review: Effective product-to-market management: by H. R. Vorstman (translated from the Dutch Produktmarktbeleid en kwaliteit by R. Harald). Prentice Hall International (UK) Ltd, Hemel Hempstead, 1993, paperback, ISBN 0-13-0437875. £19.95. Technovation, 15 (1):50–51, February 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497295900071.

Bessant:1995:BRQ

[Bes95c] John Bessant. Book review: The quality toolkit: an A– Z of tools and techniques: by John Marsh. IFS International Ltd, Wolseley Business Park, Kempston, Bedford MK42 7PW, UK, 1993, 124 pp., ISBN 1-85423-106-

5. £29.95. Technovation, 15(2):129, March 1995. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497295900179.

Bessant:1996:BRM

[Bes96] John Bessant. Book review: Mastering the dynamics of innovation: by James M. Utterback, Harvard Business School Press, Boston, MA. 1994, ISBN 0-87584-342-5, US$24.95 (hardback), 253 pp. Technovation, 16(3):99, March 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649729690173X.

Bessant:1997:BRH

[Bes97] John Bessant. Book review: How organisations learn: a critical reader: by Ken Starkey (editor), International Thomson Business Press, London, 1996, 424 pp., ISBN 0-415-12609-6 hardback (£45.00), ISBN 0-415-12610-X paperback (£17.99). Technovation, 17(11–12):735–736, November/December 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497297887731.

Bessant:2001:BRR

[Bes01a] John Bessant. Book review: Radical Innovation: How Mature Companies can Outsmart Upstarts: Richard Leifer, Christopher McDermott, Gina O’Connor, Lois Peters, Mark Rice and Robert Veryzer; Harvard Business School Press, 2000, ISBN 1-57851-903-2. Technovation, 21(10): 706–707, October 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497201000438.

Bessant:2001:BRW

[Bes01b] John Bessant. Book review: When sparks fly: Igniting creativity in groups: Dorothy Leonard and Walter Swap; Harvard Business School Press, Boston, Mass, ISBN 0-87584-865-6, 1999. Technovation, 21(6):396–397, June 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000961.

Bessant:2002:BRC

[Bes02a] J. Bessant. Book review: Corning and the craft of innovation: Margaret B. W. Graham and Alec T. Shuldiner. Oxford University Press. £21.99. ISBN 0-19-514097-4. Technovation, 22(10):655, October 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000330.

Bessant:2002:BRSa

[Bes02b] John Bessant. Book review: Statistical process control in the office: a practical guide to continuous improvement: Mal Owen and John Morgan, Greenfield Publishing, Kenilworth, UK (2000), ISBN 0-9523328-4-1, 291 pages, £40.00. Technovation, 22(7):469, July 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201001006.

Bessant:2002:BRSb

[Bes02c] John Bessant. Book review: Supply chain management workbook: Francis Harrison, Butterworth Heinemann, Oxford, 2001, 181 pages, ISBN 0-7506-4999-2. Technovation, 22(9):593, September 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720100102X.

Bessant:2002:BRT

[Bes02d] John Bessant. Book review: TMAP: a guide to technology management assessment procedure: C. Farrukh, R. Phaal and D. Probert, Institution of Electrical Engineers, London, 2000, ISBN 0-85296-768-3, Price:

£95.00. Technovation, 22(7):469–470, July 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201001018.

Bessant:2003:BRB

[Bes03a] J. Bessant. Book review: Building a project driven enterprise: how to slash waste and boost profits through lean project management: Ronald Mascitelli, Published by Technology Perspectives, Northridge, CA, USA. 2002. ISBN 0-9662697-1-3. Technovation, 23(5):459, May 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000038.

Bessant:2003:BRN

[Bes03b] J. Bessant. Book review: The new competitive advantage: Michael Best, Oxford University Press, 2001. Technovation, 23(10):843, October 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497203000294.

Best:2015:GBI

[Bes15] Michael H. Best. Greater Boston’s industrial ecosystem: a manufactory of sectors. Technovation, 39–40(??):4–13, May/June 2015. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649721400039X.

Burton:1986:VFP

[BF86] Richard M. Burton and John D. Forsyth. Variety and the firm’s performance: an empirical investigation. Technovation, 5(1–3):9–21, October 1986. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497286900416.

Bessant:1997:INP

[BF97] John Bessant and David Francis. Implementing the new product development process. Technovation, 17(4):189–197, 222, April 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497297846901.

Bessant:1999:ULN

[BF99] John Bessant and David Francis. Using learning networks to help improve manufacturing competitiveness. Technovation, 19(6–7):373–381, June 1999. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497299000255.

Barbiroli:2003:PDV

[BF03]  G. Barbiroli and A. Focacci.  Product diversifica-

tion in the vehicles industry: a techno-economic analysis. Technovation, 23(6):461–513, June 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202001001.

Barbiroli:2004:CES

[BF04] Giancarlo Barbiroli and Antonio Focacci. The changing economic significance of consumer durables resulting from the evolution of productive systems. Technovation, 24(3):235– 263, March 2004. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203001147.

Barbiroli:2005:PDC

[BF05] Giancarlo Barbiroli and Antonio Focacci. Product diversification in the computer industry: a structural analysis. Technovation, 25(9):947–970, September 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204000379.

Brown:2006:SRB

[BF06] Steve Brown and Felicia Fai. Strategic resonance between technological and organisational capabilities in the innovation process within firms. Technovation, 26(1):60–75, January 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204001439.

Blichfeldt:2021:PED

[BF21] Henrik Blichfeldt and Rita Faullant. Performance effects of digital technology adoption and product and service innovation — a process-industry perspective. Technovation, ??

(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497221000560.

Brea:2023:NSB

[BF23] Edgar Brea and Jerad A. Ford. No silver bullet: Cognitive technology does not lead to novelty in all firms. Technovation, 122(??):??, April 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222001948.

Burton:1986:PIM

[BFM86] Richard M. Burton, John D. Forsyth, and Donald M. Melick. The post-introduction management of products and services under rapidly changing environmental conditions. Technovation, 5(1–3):125–143, October 1986. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497286900489.

Burton:1988:SVU

[BFM88] Richard M. Burton, John D. Forsyth, and Donald M. Melick. Searching for viability under changing environmental conditions. Technovation, 8(1–3):111–131, ???? 1988. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900569.

Burton:1988:NWC

[BFO88a] Richard M. Burton, John D. Forsyth, and Børge Obel. The new world of constant change. Technovation, 8(1–3):1–6, ???? 1988. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497288900508.

Burton:1988:P

[BFO88b] Richard M. Burton, John D. Forsyth, and Børge Obel. Preface. Technovation, 8(1–3):vii, ???? 1988. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900491.

Barge-Gil:2010:CBI

[BG10] Andrés Barge-Gil. Cooperation-based innovators and peripheral cooperators: an empirical analysis of their characteristics and behavior. Technovation, 30(3):195–206, March 2010. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497209001710.

Bessonova:2019:HIC

[BG19] Evguenia Bessonova and Ksenia Gonchar. How the innovation–competition link is shaped by technology distance in a high-barrier catch-up economy. Technovation, 86–87(??):15–32, August/September 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0166497219300021.

Badescu:2009:IIT

[BGA09] Mariela Badescu and Concepción Garcés-Ayerbe. The impact of information technologies on firm productivity: Empirical evidence from Spain. Technovation, 29(2):122–129, February 2009. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497208000862.

Bez:2023:THO

[BGF23] Sea Matilda Bez, Irène Georgescu, and Mohammad Saleh Farazi. TripAdvisor of healthcare: Opportunities for value creation through patient feedback platforms. Technovation, 121(??):??, March 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497222001729.

Bromiley:1988:UES

[BGM88] Philip Bromiley, Michele Govekar, and Alfred Marcus. On using event-study methodology in strategic management research. Technovation, 8(1–3):25–42, ???? 1988. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900521.

Brockhoff:1991:IFR

[BGR91] Klaus Brockhoff, Ashok K. Gupta, and Christian Rotering. Inter-firm R&D co-operations in Germany. Technovation, 11(4):219–229, May 1991. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497291900353.

Brew-Hammond:2003:BRF

[BH03] Abeeku Brew-Hammond. Book review: Failing to compete, technology development and technology systems in Africa: Sanjaya Lall and Carlo Pietrobelli, Edward Elgar, Cheltenham, 2002, pp. 268 (incl. index), ISBN 1-84064-640-3 (hardcover). Technovation, 23(12):963–964, December 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203000798.

Bstieler:2015:ERC

[BH15] Ludwig Bstieler and Martin Hemmert. The effectiveness of relational and contractual governance in new product development collaborations: Evidence from Korea. Technovation, 45–46(??):29–39, November/December 2015. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497215000486.

Baraldi:2016:IND

[BH16] Enrico Baraldi and Malena Ingemansson Havenvid. Identifying new dimensions of business incubation: a multilevel analysis of Karolinska Institute’s incubation system. Technovation, 50–51(??):53–68, April/May 2016. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497215000607.

Bhat:2005:CNT

[Bha05] Jyoti S. A. Bhat. Concerns of new technology based industries — the case of nanotechnology. Technovation, 25(5): 457–462, May 2005. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203001664.

Buesa:2006:RSI

[BHPB06] Mikel Buesa, Joost Heijs, Mónica Martínez Pellitero, and Thomas Baumert. Regional systems of innovation and the knowledge production function: the Spanish case. Technovation, 26(4):463–472, April 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204002214.

Byckling:2000:SOC

[BHPW00] Eero Byckling, Ari-Pekka Hameri, Thomas Pettersson, and Horst Wenninger. Spin-offs from CERN and the case of TuoviWDM. Technovation, 20(2):71–80, February 2000. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001133.

Blomqvist:2005:PCG

[BHS05] Kirsimarja Blomqvist, Pia Hurmelinna, and Risto Seppänen. Playing the collaboration game right-balancing trust and contracting. Technovation, 25(5):497–504, May 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001555.

Biemans:1991:UTP

[Bie91] Wim G. Biemans. User and third-party involvement in developing medical equipment innovations. Technovation, 11(3):163–182, April 1991. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729190032Y.

Bigot:2013:FNS

[Big13] Bernard Bigot. The future of nuclear: Safety and development. Technovation, 33(4–5):109, April/May 2013. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213000321.

Baba:1995:JRS

[BiKM95] Yasunori Baba, Jun ichi Kikuchi, and Shunsuke Mori. Japan’s R&D strategy reconsidered: Departure from the manageable risks. Technovation, 15(2):65–78, March 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729596611V.

Butcher:2005:UBI

[BJ05] Juliette Butcher and Paul Jeffrey. The use of bibliometric indicators to explore industry-academia collaboration trends over time in the field of membrane use for water treatment. Technovation, 25(11):1273–1280, November 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204000951.

Behl:2024:EFG

[BJB^{+}24] Abhishek Behl, Nirma Jayawardena, Shikha Bhardwaj, Vijay Pereira, Manlio del Giudice, and Justin Zhang. Examining the failure of gamification in implementing innovation from the perspective of problematization in the retail sectors of emerging economies. Technovation, 129(??): ??, January 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223002134.

Bjerregaard:2010:IAC

[Bje10] Toke Bjerregaard. Industry and academia in convergence: Micro-institutional dimensions of R&D collaboration. Technovation, 30(2):100–108, February 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209001692.

Beckers:1989:PID

[BK89] Harry L. Beckers and Alistair Keddie. Project identification and definition: Organizational, financial and institutional arrangements. Technovation, 9(2–3):265–270, June 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900667.

Bellmann:1999:ERI

[BK99] Klaus Bellmann and Anshuman Khare. European response to issues in recycling car plastics. Technovation, 19(12):721– 734, December 1999. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497299000814.

Bellmann:2000:EIR

[BK00] Klaus Bellmann and Anshuman Khare. Economic issues in recycling end-of-life vehicles. Technovation, 20(12):677– 690, December 2000. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497200000122.

Bellmann:2001:RSD

[BK01] Klaus Bellmann and Anshuman Khare. RETRACTED: a systems dynamic perspective on the development of recycling strategy for end-of-life vehicles. Technovation, 21 (8):489–499, August 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497200000638.

Barreto:2004:ERM

[BK04] Leonardo Barreto and Socrates Kypreos. Endogenizing R&D and market experience in the “bottom-up” energysystems ERIS model. Technovation, 24(8):615–629, August 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497202001244.

Back:2015:BRC

[BK15] Iivari Bäck and Marko Kohtamäki. Boundaries of R&D collaboration. Technovation, 45–46(??):15–28, November/ December 2015. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497215000474.

Barata:2024:HWD

[BK24] João Barata and Ina Kayser. How will the digital twin shape the future of Industry 5.0? Technovation, 134(??): ??, June 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497224000750.

Baudier:2023:DTH

[BKA^{+}23] Patricia Baudier, Galina Kondrateva, Chantal Ammi, Victor Chang, and Francesco Schiavone. Digital transformation of healthcare during the COVID-19 pandemic: Patients’ teleconsultation acceptance and trusting beliefs. Technovation, 120(??):??, February 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222000943.

Bathelt:2010:KBT

[BKM10] Harald Bathelt, Dieter F. Kogler, and Andrew K. Munro. A knowledge-based typology of university spin-offs in the context of regional economic development. Technovation, 30(9–10):519–532, September/October 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210000519.

Belitski:2023:DAE

[BKP23] Maksim Belitski, Julia Korosteleva, and Lucia Piscitello. Digital affordances and entrepreneurial dynamics: New evidence from European regions. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497221002236.

Bhate:1997:EFP

[BL97]  Seema Bhate and Kevin Lawler.  Environmentally

friendly products: Factors that influence their adoption. Technovation, 17(8):457–465, August 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297000060.

Becheikh:2006:LIE

[BLA06] Nizar Becheikh, Réjean Landry, and Nabil Amara. Lessons from innovation empirical studies in the manufacturing sector: a systematic review of the literature from 1993– 2003. Technovation, 26(5–6):644–664, May/June 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205001033.

Blagburn:2016:WSB

[Bla16] Natalia Blagburn. Why startup boards matter. Technovation, 57–58(??):45–46, November/December 2016. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497216302620.

Bruce:1995:CCP

[BLL95] Margaret Bruce, Fiona Leverick, and Dale Littler. Complexities of collaborative product development. Technovation, 15(9):535–552, November 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/016649729596584G.

Boutillier:2023:SAT

[BLLEM23] Sophie Boutillier, Blandine Laperche, Didier Lebert, and Sana Elouaer-Mrizak. A systemic analysis of the technological trajectory at company level based on patent data: the case of Sanofi’s vaccine technology. Technovation, 124 (??):??, June 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223000573.

Bessant:2005:MIB

[BLNP05] John Bessant, Richard Lamming, Hannah Noke, and Wendy Phillips. Managing innovation beyond the steady state. Technovation, 25(12):1366–1376, December 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000702.

Burgess:2000:IAU

[BLQS00a] J. E. Burgess, P. J. Longhurst, J. Quarmby, and T. Stephenson. Innovational adaptation in the UK water and wastewater industry: a case study of introducing DTA. Technovation, 20(1):37–45, January 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001091.

Burgess:2000:DTA

[BLQS00b] Joanna E. Burgess, P. Longhurst, J. Quarmby, and T. Stephenson. Direct toxicity assessment: adaptation and the proactive roles of regulators and operators. Technovation, 20(6):313–320, June 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001455.

Bessant:1985:CCI

[BLS85] John Bessant, Richard Lamming, and Peter Senker. The challenge of computer-integrated manufacturing. Technovation, 3(4):283–295, October 1985. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497285900252.

Bruce:1994:MEDa

[BM94a] Margaret Bruce and Barny Morris. Managing external design professionals in the product development process. Technovation, 14(9):585–599, November 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900418. See [BM94b].

Bruce:1994:MEDb

[BM94b] Margaret Bruce and Barny Morris. Managing external design professionals in the product development process: Technovation, 14(9) (1994) 585–599. Technovation, 14 (9):622–623, November 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497294900450.

See [BM94a].

Brown:2005:SRF

[BM05] Steve Brown and Harvey Maylor. Strategic resonant firms, mass producers, big fish & flat liners: a study of policies, practices & performance in innovation. Technovation, 25(4): 307–319, April 2005. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203001408.

Bruque:2007:ODI

[BM07] Sebastián Bruque and José Moyano. Organisational determinants of information technology adoption and implementation in SMEs: the case of family and cooperative firms. Technovation, 27(5):241–253, May 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497206001210.

Brown:2014:IHT

[BM14] Ross Brown and Colin Mason. Inside the high-tech black box: a critique of technology entrepreneurship policy. Technovation, 34(12):773–784, December 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214001084.

Baskaran:2016:WIA

[BM16a] Shruthi Baskaran and Khanjan Mehta. What is innovation anyway? Youth perspectives from resource-constrained environments. Technovation, 52–53(??):4–17, June/July 2016. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497216000067.

Blind:2016:MSE

[BM16b] Knut Blind and Axel Mangelsdorf. Motives to standardize: Empirical evidence from Germany. Technovation, 48–49(??):13–24, February/March 2016. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649721600002X.

Balzano:2023:EPL

[BM23] Marco Balzano and Giacomo Marzi. Exploring the pathways of learning from project failure and success in new product development teams. Technovation, 128(??):??, December 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649722300189X.

Bayo-Moriones:2007:FLA

[BMLL07] Alberto Bayo-Moriones and Fernando Lera-López. A firmlevel analysis of determinants of ICT adoption in Spain. Technovation, 27(6–7):352–366, June/July 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000168.

Boly:2000:ILT

[BMRG00] Vincent Boly, Laure Morel, Jean Renaud, and Claudine Guidat. Innovation in low tech SMBs: evidence of a necessary constructivist approach. Technovation, 20(3):161–168, March 2000. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497299000991.

Bibeau:2024:FFB

[BMSJ24] Jean Bibeau, Roxane Meilleur, and Étienne St-Jean. To formalize, or not to formalize, business incubators’ networks: That is not the question. Technovation, 130(??):??, February 2024. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223002158.

Bakouros:2002:SPH

[BMV02] Yiannis L. Bakouros, Dimitri C. Mardas, and Nikos C. Varsakelis. Science park, a high tech fantasy?: an analysis of the science parks of Greece. Technovation, 22(2):123– 128, February 2002. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497200000870.

Bach:2014:HDF

[BMW14] Laurent Bach, Mireille Matt, and Sandrine Wolff. How do firms perceive policy rationales behind the variety of instruments supporting collaborative R&D? Lessons from the European Framework Programs. Technovation, 34(5–6):327– 337, May/June 2014. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497214000327.

Burton:1988:FSP

[BN88] Richard M. Burton and Donald H. Namm. Fitting the strategic planning process to organization: a clinical study of Burroughs Wellcome Co. Technovation, 8(1–3):143–154, ???? 1988. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497288900582.

Baraldi:2006:CDB

[BN06] Enrico Baraldi and Giancarlo Nadin. The challenges in digitalising business relationships. The construction of an IT infrastructure for a textile-related business network. Technovation, 26(10):1111–1126, October 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205001483.

Bergek:2008:IBP

[BN08] Anna Bergek and Charlotte Norrman. Incubator best practice: a framework. Technovation, 28(1–2):20–28, January/ February 2008. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497207001046.

Brem:2016:IFS

[BNS16] Alexander Brem, Petra A. Nylund, and Gerd Schuster. Innovation and de facto standardization: the influence of dominant design on innovative performance, radical innovation, and process innovation. Technovation, 50–51(??):79– 88, April/May 2016. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497215000760.

Bigliardi:2005:BMI

[BNV05a] Barbara Bigliardi, Anna Nosella, and Chiara Verbano. Business models in Italian biotechnology industry: a quantitative analysis. Technovation, 25(11):1299–1306, November 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204002056.

Blake:2005:TNW

[BNV05b] Brian F. Blake, Kimberly A. Neuendorf, and Colin M. Valdiserri. Tailoring new websites to appeal to those most likely to shop online. Technovation, 25(10):1205–1214, October 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204000653.

Brown:1981:UIL

[BO81] George E. Brown, Jr. and Thomas C. O’Brien. Universityindustry links: Government as blacksmith. Technovation, 1(2):85–95, August 1981. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497281900110.

Burton:1986:EOR

[BO86a] Richard M. Burton and Børge Obel. Environmentalorganizational relations: the effects of deregulation. Technovation, 5(1–3):23–34, October 1986. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0166497286900428.

Burton:1986:IDO

[BO86b] Richard M. Burton and Børge Obel. Implications of deregulation: an overview. Technovation, 5(1–3):1–7, October 1986. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497286900404.

Burton:1986:P

[BO86c] Richard M. Burton and Børge Obel. Preface. Technovation, 5(1–3):vii, October 1986. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497286900398.

Boardman:2008:BSI

[Boa08] P. Craig Boardman. Beyond the stars: the impact of affiliation with university biotechnology centers on the industrial involvement of university scientists. Technovation, 28(5): 291–297, May 2008. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497207000806.

Bustinza:2022:EIB

[BOBT22] Oscar F. Bustinza, Marco Opazo-Basaez, and Shlomo Tarba. Exploring the interplay between Smart Manufacturing and KIBS firms in configuring product-service innovation performance. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497221000390.

Bohnsack:2018:LNF

[Boh18] René Bohnsack. Local niches and firm responses in sustainability transitions: The case of low-emission vehicles in China. Technovation, 70–71(??):20–32, ???? 2018. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497218301159.

Bollingtoft:2012:BBI

[Bøl12] Anne Bøllingtoft. The bottom-up business incubator: Leverage to networking and cooperation practices in a self-generated, entrepreneurial-enabled environment. Technovation, 32(5):304–315, May 2012. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211001672.

Bonecchi:2002:BRS

[Bon02] M. B. Burkle Bonecchi. Book review: Strategic Management of Technological Learning: By Elias G. Carayannis; CRC Press, pp.197, ISBN 0-8493-3741-0. Technovation, 22(8):535–536, August 2002. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497202000184.

Borg:2001:KII

[Bor01]  Erik A. Borg.  Knowledge, information and intel-

lectual property: implications for marketing relationships. Technovation, 21(8):515–524, August 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000663.

Bouet:2015:SIP

[Bou15] Dominique Bouet. A study of intellectual property protection policies and innovation in the Indian pharmaceutical industry and beyond. Technovation, 38(??):31–41, April 2015. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497214001540.

Boxer:2001:BRI

[Box01] Rosie Boxer. Book review: Investigating entrepreneurial opportunities: a practical guide for due diligence: Richard P. Green II and James J. Carroll; Sage Publications, Thousand Oaks CA. No UK price quoted, pp. 180, ISBN 0-8039-5942-7. Technovation, 21(4):263–264, April 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000675.

Boyson:2014:CSC

[Boy14] Sandor Boyson. Cyber supply chain risk management: Revolutionizing the strategic control of critical IT systems. Technovation, 34(7):342–353, July 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214000194.

Bozeman:1994:CRG

[BP94] Barry Bozeman and Sanjay Pandey. Cooperative R&D in government laboratories: comparing the US and Japan. Technovation, 14(3):145–159, April 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900523. See [Ano94f].

Buratti:2001:ATT

[BP01] Nicoletta Buratti and Lara Penco. Assisted technology transfer to SMEs: lessons from an exemplary case. Technovation, 21(1):35–43, January 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000158.

Boardman:2009:EUR

[BP09a] P. Craig Boardman and Branco L. Ponomariov. Erratum to “University researchers working with private companies” [Technovation 29 (2009) 142–153]. Technovation, 29(8):574, August 2009. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497209000170. See [BP09b].

Boardman:2009:URW

[BP09b] P. Craig Boardman and Branco L. Ponomariov. University researchers working with private companies. Technovation, 29(2):142–153, February 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497208000527.

See erratum [BP09a].

Battaglia:2021:OBB

[BPU21] Daniele Battaglia, Emilio Paolucci, and Elisa Ughetto. Opening the black box of university Proof-of-Concept programs: Project and team-based determinants of research commercialization outcomes. Technovation, 108(??):??, December 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221001152.

Bourke:2016:AAI

[BR16] Jane Bourke and Stephen Roper. AMT adoption and innovation: an investigation of dynamic and complementary effects. Technovation, 55–56(??):42–55, September/ October 2016. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497216300463.

Bhatti:2023:SSP

[BRA^{+}23] Sabeen Hussain Bhatti, Majid Rashid, Ahmad Arslan, Shlomo Tarba, and Yipeng Liu. Servitized SMEs’ performance and the influences of sustainable procurement, packaging, and distribution: the mediating role of ecoinnovation. Technovation, 127(??):??, September 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001426.

Bruneel:2012:EBI

[BRCG12] Johan Bruneel, Tiago Ratinho, Bart Clarysse, and Aard Groen. The evolution of business incubators: Comparing demand and supply of business incubation services across different incubator generations. Technovation, 32(2):110– 121, February 2012. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497211001659.

Brea:2023:FMA

[Bre23] Edgar Brea. A framework for mapping actor roles and their innovation potential in digital ecosystems. Technovation, 125(??):??, July 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497223000949.

Brady:1997:TTM

[BRH^{+}97] T. Brady, H. Rush, M. Hobday, A. Davies, D. Probert, and S. Banerjee. Tools for technology management: an academic perspective. Technovation, 17(8):417–426, August 1997. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297000175.

Brown:1983:E

[Bro83] Wayne S. Brown. Editorial. Technovation, 2(1):1, February 1983. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497283900020.

Brown:1985:PMC

[Bro85] Wayne S. Brown. A proposed mechanism for commercializing university technology. Technovation, 3(1):19–25, February 1985. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497285900331.

Brown:1989:HCP

[Bro89] Simon J. S. Brown. Holography — a current perspective. Technovation, 9(6):505–517, October 1989. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900308.

Brown:1990:BRI

[Bro90] Damian Brown. Book review: Innovate! How to gain and sustain competitive advantage: by Mark O’Hare, Basil Blackwell, Oxford, 1988, £15.00, 256 pages. Technovation, 10(4):281–282, June 1990. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729090059S.

Brockhoff:1992:IPD

[Bro92] Klaus K. Brockhoff. Instruments for patent data analyses in business firms. Technovation, 12(1):41–59, February 1992. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729290031C.

Brooks:1994:CED

[Bro94] D. G. M. Brooks. The competitive edge in developing a traditional manufacturing business. Technovation, 14(6):357– 362, August 1994. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497294900159.

Brockhoff:1995:VGI

[Bro95] Klaus Brockhoff. Value generation by industrial research. Technovation, 15(10):591–599, December 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729599327C.

Brown:1998:MSS

[Bro98] Steve Brown. Manufacturing seniority, strategy and innovation. Technovation, 18(3):149–162, March 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297001235.

Brown:2001:MPT

[Bro01] Steve Brown. Managing process technology — further empirical evidence from manufacturing plants. Technovation, 21(8):467–478, August 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497200000742.

Brown:2007:BRM

[Bro07] Steve Brown. Book review: Making common sense common practice — models for manufacturing excellence, Ron Moore. Elsevier, Amsterdam (2004). Technovation, 27(1–2):89–90, January/February 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497206000289.

Buskirk:1994:PMD

[BRP94] Bruce D. Buskirk, Allan C. Reddy, and Edward T. Popper. Planning market development in high-tech firms. Technovation, 14(8):493–503, October 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729490149X.

Becker:2023:AIS

[BRV23] Bettina Becker, Stephen Roper, and Enrico Vanino. Assessing innovation spillovers from publicly funded R&D and innovation support: Evidence from the UK. Technovation, 128(??):??, December 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497223001712.

Brennan:1987:PTE

[BS87] D. J. Brennan and G. K. Stephens. Productivity trends in expanding plants and new plants. Technovation, 6(3):189– 204, August 1987. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497287900228.

Beaumont:1997:TMP

[BS97] N. B. Beaumont and R. M. Schroder. Technology, manufacturing performance and business performance amongst Australian manufacturers. Technovation, 17(6):297–307, June 1997. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297000114.

Bolisani:1999:ITM

[BS99] Ettore Bolisani and Enrico Scarso. Information technology management: a knowledge-based perspective. Technovation, 19(4):209–217, February 1999. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0166497298001096.

Bagchi-Sen:2001:PIC

[BS01] Sharmistha Bagchi-Sen. Product innovation and competitive advantage in an area of industrial decline: the Niagara region of Canada. Technovation, 21(1):45–54, January 2001.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720000016X.

Bhatnagar:2005:SCC

[BS05] Rohit Bhatnagar and Amrik S. Sohal. Supply chain competitiveness: measuring the impact of location factors, uncertainty and manufacturing practices. Technovation, 25(5): 443–456, May 2005. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649720300172X.

Bernstein:2006:IIP

[BS06] Boaz Bernstein and Prakash J. Singh. An integrated innovation process model based on practices of Australian biotechnology firms. Technovation, 26(5–6):561–572, May/ June 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204002202.

Banuls:2008:FKA

[BS08] Victor A. Bañuls and Jose L. Salmeron. Foresighting key areas in the Information Technology industry. Technovation, 28(3):103–111, March 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000661.

Berkowitz:2024:FST

[BS24] Heloise Berkowitz and Antoine Souchaud. Filling successive technologically-induced governance gaps: Metaorganizations as regulatory innovation intermediaries. Technovation, 129(??):??, January 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0166497223002018.

Behl:2024:CGH

[BSG^{+}24] Abhishek Behl, Brinda Sampat, Jighyasu Gaur, Vijay Pereira, Benjamin Laker, Amit Shankar, Peter Shi, and Mohammad Roohanifar. Can gamification help green supply chain management firms achieve sustainable results in servitized ecosystem? An empirical investigation. Technovation, 129(??):??, January 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223002262.

Biancone:2023:HFM

[BSMC23] Paolo Biancone, Silvana Secinaro, Roberto Marseglia, and Davide Calandra. E-health for the future. Managerial perspectives using a multiple case study approach. Technovation, 120(??):??, February 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001875.

Bao:2021:DNP

[BSN21] Yongchuan Bao, Zhongfeng Su, and Charles H. Noble. Determinants of new product development speed in China: a strategy tripod perspective. Technovation, 106(??):??, August 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221000729.

Bargoni:2024:HHP

[BSSF24] Augusto Bargoni, Lubos Smrcka, Gabriele Santoro, and Alberto Ferraris. Highway to hell or paradise city? Exploring the role of growth hacking in learning from innovation failure. Technovation, 131(??):??, March 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223002560.

Bores:2003:TCS

[BST03] C. Borés, C. Saurina, and R. Torres. Technological convergence: a strategic perspective. Technovation, 23(1):1–13, January 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201000943.

Bijedic:2023:HDW

[BSWC23] Teita Bijedić, Christian Schröder, Arndt Werner, and Xiangyu Chen. How do working conditions, network relationships, and institutional support offers effect entrepreneurial intentions of German university scientists? Technovation, 123(??):??, May 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000263.

Barrett:2022:TGE

[BT22] Gillian Barrett and George Tsekouras. A tango with a gorilla: an exploration of the microfoundations of open innovation partnerships between young innovative companies and multi-national enterprises. Technovation, 117 (??):??, September 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222001080.

Bamel:2023:DDI

[BTP^{+}23] Umesh Bamel, Shalini Talwar, Vijay Pereira, Laura Corazza, and Amandeep Dhir. Disruptive digital innovations in healthcare: Knowing the past and anticipating the future. Technovation, 125(??):??, July 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000962.

Buckright:1991:BRA

[Buc91] John Buckright. Book review: The American synthetic rubber research program: by J. T. Morris, University of Pennsylvania Press, PA, 1989, ISBN 0-8122-8205-1, $34.95, 191 pages. Technovation, 11(2):128–129, March 1991. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729190050E.

Bullock:1985:CSR

[Bul85] Matthew Bullock. Cohabitation: Small research-based companies and the universities. Technovation, 3(1):27–38, February 1985. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497285900343.

Bull:1997:BRB

[Bul97] Gordon Bull. Book review: Building a corporate Internet strategy: the IT manager’s guide: by Amit K. Maitra. Van Nostrand Reinhold, 1996, ISBN 0-442-02246-8, £22.50 (paperback). Technovation, 17(8):475–476, August 1997.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297834669.

Butler:1996:BRO

[But96] Nick Butler. Book review: Oxford Reference Shelf on CD-ROM: Oxford University Press (First edition: October 1994), ISBN 0-19-268255-5, £68.08 (+VAT). Technovation, 16(10):609–610, October 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497297828994.

Brem:2009:IMP

[BV09] Alexander Brem and Kai-Ingo Voigt. Integration of market pull and technology push in the corporate front end and innovation management-Insights from the German software industry. Technovation, 29(5):351–367, May 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208000898.

Bennett:2024:EMT

[BV24] Roger Bennett and Rohini Vijaygopal. Exploring mobility and transportation technology futures for people with ambulatory disabilities: a science fiction prototype. Technovation, 133(??):??, May 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000518.

Bustinza:2024:IPS

[BVHJ24] Oscar F. Bustinza, Ferran Vendrell-Herrero, and Charbel Jose Chiappetta Jabbour. Integration of product-service innovation into green supply chain management: Emerging opportunities and paradoxes. Technovation, 130(??): ??, February 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223002341.

Bonazzi:2024:CMS

[BVS24] Riccardo Bonazzi, Gianluigi Viscusi, and Adriano Solidoro. Crowd mining as a strategic resource for innovation seekers. Technovation, 132(??):??, April 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000191.

Basaure:2021:ITI

[BVT21] Arturo Basaure, Alexandr Vesselkov, and Juuso Töyli. Internet of Things (IoT) platform competition: Consumer switching versus provider multihoming. Technovation, ?? (??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497218305303.

Bartnik:2018:IIE

[BWF18] Roman Bartnik, Miriam Wilhelm, and Takahiro Fujimoto. Introduction to innovation in the East Asian automotive industry: Exploring the interplay between product architectures, firm strategies, and national innovation systems. Technovation, 70–71(??):1–6, ???? 2018. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497218302736.

Bunduchi:2011:MBC

[BWS11] Raluca Bunduchi, Clara Weisshaar, and Alison U. Smart. Mapping the benefits and costs associated with process innovation: the case of RFID adoption. Technovation, 31 (9):505–521, September 2011. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497211000551.

Balbontin:2000:NPD

[BYCS00] A. Balbontin, B. B. Yazdani, R. Cooper, and W. E. Souder. New product development practices in American and British firms. Technovation, 20(5):257–274, May 2000. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001364.

Barbero:2021:MAB

[BZIZ21] Javier Barbero, Jon Mikel Zabala-Iturriagagoitia, and José L. Zofío. Is more always better? On the relevance of decreasing returns to scale on innovation. Technovation, 107(??):??, September 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722100095X.

Carayannis:2001:VWM

[CA01] Elias G. Carayannis and Jeffrey Alexander. Virtual, wireless mannah: a co-opetitive analysis of the broadband satellite industry. Technovation, 21(12):759–766, December 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000833.

Carayannis:2002:TLF

[CA02] Elias G. Carayannis and Jeff Alexander. Is technological learning a firm core competence, when, how and why? A longitudinal, multi-industry study of firm technological learning and market performance. Technovation, 22 (10):625–643, October 2002. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497201000475.

Chorev:2006:SIH

[CA06] Schaul Chorev and Alistair R. Anderson. Success in Israeli high-tech start-ups; Critical factors and process. Technovation, 26(2):162–174, February 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205001008.

Chiva:2007:LDM

[CA07] Ricardo Chiva and Joaquín Alegre. Linking design management skills and design function organization: an empirical study of Spanish and Italian ceramic tile producers. Technovation, 27(10):616–627, October 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000764.

Caetano:2011:RTP

[CA11] Mauro Caetano and Daniel C. Amaral. Roadmapping for technology push and partnership: a contribution for open innovation environments. Technovation, 31(7):320– 335, July 2011. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497211000198.

Cohen:2014:MDS

[CA14] Boyd Cohen and Jose Ernesto Amorós. Municipal demandside policy tools and the strategic management of technology life cycles. Technovation, 34(12):797–806, December 2014. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497214000960.

Cabibbo:1989:NC

[Cab89] Nicola Cabibbo. Nicola Cabibbo. Technovation, 9(2–3): 297–298, June 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497289900746.

Cabaleiro:2019:SAC

[Cab19] Goretti Cabaleiro. Sources of appropriation capacity in licensing agreements. Technovation, 86–87(??):48–61, August/September 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497217304960.

Canongia:2004:TFU

[CAdNFP04] Claudia Canongia, Adelaide Antunes, and Maria de Nazaré Freitas Pereira. Technological foresight — the use of biotechnology in the development of new drugs against breast cancer. Technovation, 24(4):299–309, April 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203001901.

Caird:1994:HII

[Cai94] Sally Caird. How important is the innovator for the commercial success of innovative products in SMEs? Technovation, 14(2):71–83, March 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497294900973.

See [Ano94n].

Carayannis:2000:LKL

[CAI00] Elias G. Carayannis, Jeffrey Alexander, and Anthony Ioannidis. Leveraging knowledge, learning, and innovation in forming strategic government–university–industry (GUI) R&D partnerships in the US, Germany, and France. Technovation, 20(9):477–488, September 2000. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001625.

Cai:2023:TNM

[Cai23] Yuzhuo Cai. Towards a new model of EU-China innovation cooperation: Bridging missing links between international university collaboration and international industry collaboration. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222001006.

Calik:2024:VMS

[Cal24] Eyup Calik. A validated measurement scale for sustainable product innovation performance. Technovation, 129(??):??, January 2024. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223001931.

Campelo:1993:OIC

[Cam93] M. Angela Campelo. Options for increasing competitiveness, quality and productivity: the case of the Brazilian computer industry. Technovation, 13(6):403–404, September 1993. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497293900816.

Cammell:1998:I

[Cam98] John E. Cammell. Introduction. Technovation, 18(8–9): 495, August/September 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497298000340.

Cantisani:2006:TIP

[Can06] Antonio Cantisani. Technological innovation processes revisited. Technovation, 26(11):1294–1301, November 2006.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205001550.

Candi:2016:CDE

[Can16] Marina Candi. Contributions of design emphasis, design resources and design excellence to market performance in technology-based service innovation. Technovation, 55–56(??):33–41, September/October 2016. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497216300554.

Codini:2023:BMI

[CAP23] Anna Paola Codini, Tindara Abbate, and Antonio Messeni Petruzzelli. Business Model Innovation and exaptation: a new way of innovating in SMEs. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000955.

Carayannis:1998:HOT

[Car98a] Elias G. Carayannis. Higher order technological learning as determinant of market success in the multimedia arena; a success story, a failure, and a question mark: AGFA/BAYER AG, Enable Software, and Sun Microsystems. Technovation, 18(10):639–653, 658–660, October 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497298000546.

Carayannis:1998:OTS

[Car98b] Elias G. Carayannis. Organizational transformation and strategic learning in high risk, high complexity environments. Technovation, 19(2):87–103, December 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000820.

Carayannis:1998:SMT

[Car98c] Elias G. Carayannis. The strategic management of technological learning in project/program management: the role of extranets, intranets and intelligent agents in knowledge generation, diffusion, and leveraging. Technovation, 18(11):697–703, November 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497298000650.

Carayannis:1999:FSB

[Car99a] Elias G. Carayannis. Fostering synergies between information technology and managerial and organizational cognition: the role of knowledge management. Technovation, 19 (4):219–231, February 1999. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497298001011.

Carayannis:1999:KTT

[Car99b] Elias G. Carayannis. Knowledge transfer through technological hyperlearning in five industries. Technovation, 19 (3):141–161, January 1999. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497298000819.

Carayannis:2000:IVT

[Car00] Elias G. Carayannis. Investigation and validation of technological learning versus market performance. Technovation, 20(7):389–400, July 2000. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497299001601.

Carmeli:2001:HLP

[Car01] A. Carmeli. High- and low-performance firms: do they have different profiles of perceived core intangible resources and business environment? Technovation, 21(10):661–671, October 2001. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201000505.

Carbonara:2004:IPW

[Car04] Nunzia Carbonara. Innovation processes within geographical clusters: a cognitive approach. Technovation, 24(1):17– 28, January 2004. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497202000469.

Carbonara:2005:ICT

[Car05] Nunzia Carbonara. Information and communication technology and geographical clusters: opportunities and spread. Technovation, 25(3):213–222, March 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000956.

Caviggioli:2016:TFI

[Cav16] Federico Caviggioli. Technology fusion: Identification and analysis of the drivers of technology convergence using patent data. Technovation, 55–56(??):22–32, September/ October 2016. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497216300293.

Candi:2019:EUP

[CB19] Marina Candi and Ahmad Beltagui. Effective use of 3D printing in the innovation process. Technovation, 80–81(??): 63–73, February/March 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497217308519.

Caccamo:2022:LAS

[CB22] Marta Caccamo and Sara Beckman. Leveraging accelerator spaces to foster knowledge communities. Technovation, 113(??):??, May 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497221002029.

Cankurtaran:2024:DDO

[CBF24] Pinar Cankurtaran, Michael B. Beverland, and Francis J. Farrelly. Defining design orientation: a field-based discovery approach. Technovation, 131(??):??, March 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223002547.

Cavallo:2022:ENO

[CBGvdV22] Angelo Cavallo, Henri Burgers, Antonio Ghezzi, and Vareska van de Vrande. The evolving nature of open innovation governance: a study of a digital platform development in collaboration with a big science centre. Technovation, 116(??):??, August 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497221001516.

Chung:2003:CPT

[CBK03] Jin-Woo Chung, Zong-Tae Bae, and Ji Soo Kim. Changing patterns of technological cooperation activities of innovative small firms along technological development stages in the Korean telecommunication sector. Technovation, 23 (2):163–173, February 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497201001110.

Choi:2012:NCC

[CbLjJL12] Dong Geun Choi, Yi beck Lee, Mi jung Jung, and Heesang Lee. National characteristics and competitiveness in MOT research: a comparative analysis of ten specialty journals, 2000–2009. Technovation, 32(1):9–18, January 2012. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211001295.

Carayannis:2005:CSD

[CC05] Elias Carayannis and John Coleman. Creative system design methodologies: the case of complex technical systems. Technovation, 25(8):831–840, August 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720400029X.

Chiaroni:2006:FCI

[CC06] Davide Chiaroni and Vittorio Chiesa. Forms of creation of industrial clusters in biotechnology. Technovation, 26(9): 1064–1076, September 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497205001458.

Chen:2014:PGA

[CC14] Ke Chen and Alan H. S. Chan. Predictors of gerontechnology acceptance by older Hong Kong Chinese. Technovation, 34(2):126–135, February 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497213001156.

Cabaleiro-Cervino:2020:LAS

[CCB20] Goretti Cabaleiro-Cerviño and Ana Burcharth. Licensing agreements as signals of innovation: When do they impact market value? Technovation, 98(??):??, December 2020. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722030047X.

Cugno:2022:IPR

[CCBP22] Monica Cugno, Rebecca Castagnoli, Giacomo Büchi, and Marco Pini. Industry 4.0 and production recovery in the COVID era. Technovation, 114(??):??, June 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497221002248.

Chang:2012:HDE

[CCC^{+}12] Yuan-Chieh Chang, Huo-Tsan Chang, Hui-Ru Chi, Ming-Huei Chen, and Li-Ling Deng. How do established firms improve radical innovation performance? The organizational capabilities view. Technovation, 32(7–8):441–451, July/August 2012. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497212000302.

Chiaroni:2011:OIJ

[CCF11] Davide Chiaroni, Vittorio Chiesa, and Federico Frattini. The Open Innovation Journey: How firms dynamically implement the emerging innovation management paradigm. Technovation, 31(1):34–43, January 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209001400.

Castelnovo:2023:QED

[CCF23] Paolo Castelnovo, Stefano Clò, and Massimo Florio. A quasi-experimental design to assess the innovative impact of public procurement: an application to the Italian space industry. Technovation, 121(??):??, March 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222002346.

Chaudhuri:2023:GSC

[CCGK23] Ranjan Chaudhuri, Sheshadri Chatterjee, Shivam Gupta, and Sachin Kamble. Green supply chain technology and organization performance: Moderating role of environmental dynamism and product-service innovation capability. Technovation, 128(??):??, December 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001682.

Chang:2006:SIM

[CCL^{+}06] Shih-Chia Chang, Rong-Huei Chen, Ru-Jen Lin, Shiaw-Wen Tien, and Chwen Sheu. Supplier involvement and manufacturing flexibility. Technovation, 26(10):1136–1146, October 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205001227.

Chatterjee:2023:CIF

[CCMW23] Sheshadri Chatterjee, Ranjan Chaudhuri, Marcello Mariani, and Samuel Fosso Wamba. The consequences of innovation failure: an innovation capabilities and dynamic capabilities perspective. Technovation, 128(??):??, December 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223001694.

Cannavacciuolo:2023:DIO

[CCP23] Lorella Cannavacciuolo, Guido Capaldo, and Cristina Ponsiglione. Digital innovation and organizational changes in the healthcare sector: Multiple case studies of telemedicine project implementation. Technovation, 120(??):??, February 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497222000979.

Cerchione:2023:BCH

[CCR^{+}23] Roberto Cerchione, Piera Centobelli, Emanuela Riccio, Stefano Abbate, and Eugenio Oropallo. Blockchain’s coming to hospital to digitalize healthcare services: Designing a distributed electronic health record ecosystem. Technovation, 120(??):??, February 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722200027X.

Colovic:2024:IIE

[CCR^{+}24] Ana Colovic, Annalisa Caloffi, Federica Rossi, Stefania Paladini, and Mehdi Bagherzadeh. Innovation intermediaries and emerging digital technologies. Technovation, 133(??): ??, May 2024. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497224000725.

Chiu:2006:EMN

[CCST06] Yi-Chia Chiu, Benson Chen, Joseph Z. Shyu, and Gwo-Hshiung Tzeng. An evaluation model of new product launch strategy. Technovation, 26(11):1244–1252, November 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205001185.

Chen:2011:ISD

[CCV11] Jin Chen, Yufen Chen, and Wim Vanhaverbeke. The influence of scope, depth, and orientation of external technology sources on the innovative performance of Chinese firms. Technovation, 31(8):362–373, August 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211000526.

Cheng:2005:EMA

[CCY05] Joe Z. Cheng, Peng-Ting Chen, and Hsiao-Cheng D. Yu. Establishing a MAN access strategy for future broadband service: a fuzzy MCDM analysis of SONET/SDH and Gigabit Ethernet. Technovation, 25(5):557–567, May 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203001718.

Callahan:1992:OLAa

[CD92a] John Callahan and Peter Diedrich. Organizational learning across critical linkages: the case of captive ASIC design and manufacturing. Technovation, 12(7):433–446, November 1992. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729290049N.

Callahan:1992:OLAb

[CD92b] John Callahan and Peter Diedrich. Organizational learning across critical linkages: the case of captive ASIC design and manufacturing. Technovation, 12(7):488–489, November 1992. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729290054L.

Craig:2000:FBF

[CD00] Steven G. Craig and Thomas R. DeGregori. The forward and backward flow of technology: the relationship between foreign suppliers and domestic technological advance. Technovation, 20(8):403–412, August 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000080.

Caerteling:2011:RTD

[CDD^{+}11] Jasper S. Caerteling, C. Anthony Di Benedetto, André G. Dorée, Johannes I. M. Halman, and Michael Song. RE-TRACTED: Technology development projects in road infrastructure: the relevance of government championing behavior. Technovation, 31(5–6):270–283, May/June 2011. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649721100023X. See retraction notice

[CDD^{+}14].

Caerteling:2014:RNT

[CDD^{+}14] Jasper S. Caerteling, C. Anthony Di Benedetto, André G. Dorée, Johannes I. M. Halman, and Michael Song. Retraction notice to “Technology development projects in road infrastructure: the relevance of government championing behavior” [Technovation 31 (5–6) (2011) 270– 283]. Technovation, 34(12):817, December 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214001308. See [CDD^{+}11].

Cao:2021:UMA

[CDED21] Guangming Cao, Yanqing Duan, John S. Edwards, and Yogesh K. Dwivedi. Understanding managers’ attitudes and behavioral intentions towards using artificial intelligence for organizational decision-making. Technovation, 106(??):??, August 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221000936.

Cho:2023:WDD

[CDK^{+}23] Jaehan Cho, Timothy DeStefano, Hanhin Kim, Inchul Kim, and Jin Hyun Paik. What’s driving the diffusion of nextgeneration digital technologies? Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000244.

Civera:2024:GPH

[CDMV24] Alice Civera, Alfredo De Massis, Michele Meoli, and Silvio Vismara. The goal and performance heterogeneity of academic spinoffs. Technovation, 131(??):??, March 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000221.

Chabbal:1989:HMG

[CE89] Robert Chabbal and Kerstin Eliasson. Horizontal measures, e.g., information exchange, joint development of human resources and establishment or strengthening of common infrastructures. Technovation, 9(2–3):275–279, June 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900680.

Champenois:2019:BLB

[CE19] Claire Champenois and Henry Etzkowitz. From boundary line to boundary space: the creation of hybrid organizations as a Triple Helix micro-foundation. Technovation, 76–77 (??):28–39, August/September 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://

www.sciencedirect.com/science/article/pii/S0166497217307976.

Carayannis:2003:CCL

[CEH03] Elias G. Carayannis, Dan Evans, and Mike Hanson. A cross-cultural learning strategy for entrepreneurship education: outline of key concepts and lessons learned from a comparative study of entrepreneurship students in France and the US. Technovation, 23(9):757–771, September 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000305.

Cassimon:2011:CRO

[CEY11] D. Cassimon, P. J. Engelen, and V. Yordanov. Compound real option valuation with phase-specific volatility: a multiphase mobile payments case study. Technovation, 31(5– 6):240–255, May/June 2011. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497211000149.

Carayannis:2001:PRS

[CF01a] Elias Carayannis and James Forbes. A pragmatic representation of systems engineering based on technological learning. Technovation, 21(4):197–207, April 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000286.

Chen:2001:IAM

[CF01b] Xiang Dong Chen and Lan Sheng Fu. IT adoption in manufacturing industries: differences by company size and industrial sectors — the case of Chinese mechanical industries. Technovation, 21(10):649–660, October 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720000078X.

Collinson:2007:BRH

[CF07] Simon Collinson and Lead Ghoshal Fellow. Book review: High-Tech Industries in China: Chien-Hsun Chen, Hui-Tzu Shih. Edward Elgar, Cheltenham, UK (2005), ISBN 1-84542-199X. Technovation, 27(5):322–323, May 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497206000496.

Conceicao:2012:CDR

[CFC12] Oscarina Conceição, Margarida Fontes, and Teresa Calapez. The commercialisation decisions of research-based spin-off: Targeting the market for technologies. Technovation, 32 (1):43–56, January 2012. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497211001118.

Cannavacciuolo:2023:TIE

[CFP^{+}23] Lorella Cannavacciuolo, Giovanna Ferraro, Cristina Ponsiglione, Simonetta Primario, and Ivana Quinto. Technological innovation-enabling Industry 4.0 paradigm: a systematic literature review. Technovation, 124(??):??, June 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000445.

Carayannis:2002:SSN

[CG02] Elias G. Carayannis and James Gover. The SEMATECH– Sandia National Laboratories Partnership: a case study. Technovation, 22(9):585–591, September 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000451.

Chiva-Gomez:2004:RCA

[CG04a] R. Chiva-Gomez. Repercussions of complex adaptive systems on product design management. Technovation, 24 (9):707–711, September 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497202001554.

Corsten:2004:OCO

[CG04b] Hans Corsten and Ralf Gössinger. Opportunistic coordination of operations in job shop production. Technovation, 24(3):219–234, March 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497202000743.

Calisir:2005:DBO

[CG05] F. Calisir and C. A. Gumussoy. Determinants of budget overruns on IT projects. Technovation, 25(6):631–636, June 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203002013.

Chen:2011:MIP

[CG11] Kaihua Chen and Jiancheng Guan. Mapping the innovation production process from accumulative advantage to economic outcomes: a path modeling approach. Technovation, 31(7):336–346, July 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649721100054X.

Cordasco:2021:ITA

[CGBV21] Carlo Cordasco, Cristian Gherhes, Chay Brooks, and Tim Vorley. An institutional taxonomy of adoption of innovation in the classic professions. Technovation, 107(??):??, September 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497221000535.

Cruz-Gonzalez:2015:OSS

[CGLSNLDV15] Jorge Cruz-González, Pedro López-Sáez, Jose Emilio Navas-López, and Miriam Delgado-Verde. Open search strategies and firm performance: the different moderating role of technological environmental dynamism. Technovation, 35(??):32–45, January 2015. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497214001230.

Calia:2007:INT

[CGM07] Rogerio C. Calia, Fabio M. Guerrini, and Gilnei L. Moura. Innovation networks: From technological development to business model reconfiguration. Technovation, 27(8):426– 432, August 2007. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497206000939.

Candiani:2022:TEN

[CGM22] Juan Antonio Candiani, Victor Gilsing, and Mariano Mastrogiorgio. Technological entry in new niches: Diversity, crowding and generalism. Technovation, 116(??):??, August 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000256.

Chiarello:2024:FAG

[CGS^{+}24] Filippo Chiarello, Vito Giordano, Irene Spada, Simone Barandoni, and Gualtiero Fantoni. Future applications of generative large language models: a data-driven case study on ChatGPT. Technovation, 133(??):??, May 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722400052X.

Corsten:2005:FDO

[CGW05] Hans Corsten, Ralf Gössinger, and Norbert Wolf. Flexibilitydriven order releases in job-shop production. Technovation, 25(8):815–830, August 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497204000112.

Cullen:2001:CEA

[CH01] Penny-Anne Cullen and Richard Hickman. Contracting and economics alliances in the aerospace sector: do formal contact arrangements support or impede efficient supply chain relationships? Technovation, 21(8):525–533, August 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000730.

Chapman:2004:CLB

[CH04] Ross Chapman and Paul Hyland. Complexity and learning behaviors in product innovation. Technovation, 24(7):553– 561, July 2004. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497202001219.

Chataway:1988:IPC

[Cha88] Graham Chataway. Innovation policy in contemporary Queensland. Technovation, 7(3):211–230, July 1988. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900211.

Chablo:1994:PAA

[Cha94] Alexander Chablo. Potential applications of artificial intelligence in telecommunications. Technovation, 14(7):431–435, September 1994. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497294900019.

Chataway:1999:TTR

[Cha99] Joanna Chataway. Technology transfer and the restructuring of science and technology in Central and Eastern Europe. Technovation, 19(6–7):355–364, June 1999. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299000292.

Chang:2008:IAT

[Cha08] Hsin Hsin Chang. Intelligent agent’s technology characteristics applied to online auctions’ task: a combined model of TTF and TAM. Technovation, 28(9):564–577, September 2008. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497208000400.

Chakravarty:2022:RCI

[Cha22] Sanghamitra Chakravarty. Resource constrained innovation in a technology intensive sector: Frugal medical devices from manufacturing firms in South Africa. Technovation, 112(??):??, April 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497221001784.

Chapman:2018:EPS

[CHD18] Gary Chapman and Nola Hewitt-Dundas. The effect of public support on senior manager attitudes to innovation. Technovation, 69(??):28–39, January 2018. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497216301201.

Chen:1994:EUI

[Che94] E. Yegin Chen. The evolution of university–industry technology transfer in Hong Kong. Technovation, 14(7): 449–459, September 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497294900035.

Cheng:2012:ERB

[Che12] An-Chin Cheng. Exploring the relationship between technology diffusion and new material diffusion: the example of advanced ceramic powders. Technovation, 32(3–4):163–167, March/April 2012. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497211001520.

Chico:1981:STD

[Chi81]  Leon V. Chico.  Sharing in technological develop-

ment. Technovation, 1(1):57–68, February 1981. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497281900079.

Chiang:1990:MNT

[Chi90] Jong-Tsong Chiang. Management of national technology programs in a newly industrialized country-Taiwan. Technovation, 10(8):531–554, December 1990. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729090049P.

Chiang:1991:MOD

[Chi91] Jong-Tsong Chiang. From ‘mission-oriented’ to ‘diffusionoriented’ paradigm: the new trend of U.S. industrial technology policy. Technovation, 11(6):339–356, September 1991. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729190017X.

Chiffoleau:2005:LAI

[Chi05] Y. Chiffoleau. Learning about innovation through networks: the development of environment-friendly viticulture. Technovation, 25(10):1193–1204, October 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204000719.

Chirumalla:2021:BDE

[Chi21] Koteshwar Chirumalla. Building digitally-enabled process innovation in the process industries: a dynamic capabilities approach. Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000377.

Cheah:2021:SSI

[CHL21] Sarah Lai-Yin Cheah, Yuen-Ping Ho, and Shiyu Li. Search strategy, innovation and financial performance of firms in process industries. Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000389.

Chorda:1996:TMS

[Cho96] Isidre March Chordá. Towards the maturity stage: an insight into the performance of French technopoles. Technovation, 16(3):143–152, March 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497295000429.

Conceicao:1998:UBT

[CHO98] P. Conceição, M. V. Heitor, and P. Oliveira. Universitybased technology licensing in the knowledge based economy. Technovation, 18(10):615–625, 657–658, October 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000510.

Caird:2013:OIT

[CHP13] Sally Caird, Stephen Hallett, and Stephen Potter. The Open2-Innova8ion Tool — a software tool for rating organisational innovation performance. Technovation, 33(10–11):381–385, October/November 2013. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213000655.

Chang:1999:SAI

[CHT99] Pao-Long Chang, Chiung-Wen Hsu, and Chien-Tzu Tsai. A stage approach for industrial technology development and implementation — the case of Taiwan’s computer industry. Technovation, 19(4):233–241, February 1999. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298001114.

Clarysse:2022:HIT

[CHT22] Bart Clarysse, Vivianna Fang He, and Christopher L. Tucci. How the Internet of Things reshapes the organization of innovation and entrepreneurship. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016649722200195X.

Chung:2001:USN

[Chu01] Sunyang Chung. Unification of South and North Korean innovation systems. Technovation, 21(2):99–107, February 2001. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497200000250.

Chung:2002:BNI

[Chu02] S. Chung. Building a national innovation system through regional innovation systems. Technovation, 22(8):485–491, August 2002. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201000359.

Carayannis:1998:BGU

[CJ98] Elias G. Carayannis and Jennifer Jorge. Bridging government–university–industry technological learning disconnects: a comparative study of training and development policies and practices in the U.S., Japan, Germany, and France. Technovation, 18(6–7):383–407, January 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000194.

Chen:2023:DTU

[CJ23] Wanyi Chen and Rong Jin. Does tax uncertainty affect firm innovation speed? Technovation, 125(??):??, July 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000822.

Chao:2007:IMA

[CJH^{+}07] Chia Chen Chao, Wen Yuan Jen, Ming Chien Hung, Yu Chuan Li, and Y. P. Chi. An innovative mobile approach for patient safety services: the case of a Taiwan health care provider. Technovation, 27(6–7):342–351, June/July 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720700003X.

Chung:2022:IRH

[CJL22] Doohee Chung, Haejun Jung, and Yunjeong Lee. Investigating the relationship of high-tech entrepreneurship and innovation efficacy: the moderating role of absorptive capacity. Technovation, 111(??):??, March 2022. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001747.

Conroy:2023:DKR

[CJL23] Kieran M. Conroy, Simon Jacobs, and Yang Liu. The dual knowledge role of open innovation intermediaries: Internal weaving and external filtering for MNE subsidiaries. Technovation, 123(??):??, May 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000329.

Clark:2000:CAN

[CJS00] T. Clark, P. Jeffrey, and T. Stephenson. Complex agendas for new technology adoption in the UK water industry. Technovation, 20(5):247–256, May 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001315.

Curtis:2004:ECE

[CK04] Monica Curtis and Anshuman Khare. Energy conservation in electric utilities: an opportunity for restorative economics at SaskPower. Technovation, 24(5):395–402, May 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202001165.

Cheng:2012:RSI

[CK12a] Colin C. Cheng and Dennis Krumwiede. The role of service innovation in the market orientation-new service performance linkage. Technovation, 32(7–8):487–497, July/ August 2012. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497212000351.

Clausen:2012:RBE

[CK12b] Tommy Clausen and Tor Korneliussen. The relationship between entrepreneurial orientation and speed to the market: the case of incubator firms in Norway. Technovation, 32(9–10):560–567, September/October 2012. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497212000569.

Choi:2018:MSF

[CK18] Donghyuk Choi and Yeonbae Kim. Market share and firms’ patent exploitation. Technovation, 72–73(??):13–23, April/ May 2018. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497217308982.

Candi:2022:PIB

[CK22] Marina Candi and Fumi Kitagawa. Performance implications of business model centrality over technology-based firms’ life courses. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222001730.

Chaudhary:2024:CEE

[CKF^{+}24a] Sanjay Chaudhary, Puneet Kaur, Alberto Ferraris, Stefano Bresciani, and Amandeep Dhir. Connecting entrepreneurial ecosystem and innovation. Grasping at straws or hitting a home run? Technovation, 130(??):??, February 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223002535.

Clauss:2024:IGT

[CKF24b] Thomas Clauss, Tobias Kesting, and Matheus Franco. Innovation generation through formalisation and fairness in university — Industry collaboration. Technovation, 134(??):??, June 2024. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497224000993.

Clausen:2013:MIR

[CKM13] Tommy Høyvarde Clausen, Tor Korneliussen, and Einar Lier Madsen. Modes of innovation, resources and their influence on product innovation: Empirical evidence from R&D active firms in Norway. Technovation, 33(6–7):225–233, June/ July 2013. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497213000187.

Carayannis:2000:SAS

[CKR00] Elias G. Carayannis, Suleiman K. Kassicieh, and Raymond Radosevich. Strategic alliances as a source of early-stage seed capital in new technology-based firms. Technovation, 20(11):603–615, November 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001613.

Chan:2022:DEU

[CKS22] Yolande E. Chan, Rashmi Krishnamurthy, and Arman Sadreddin. Digitally-enabled university incubation processes. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497222001079.

Caloghirou:2004:ICE

[CKT04] Yannis Caloghirou, Ioanna Kastelli, and Aggelos Tsakanikas. Internal capabilities and external knowledge sources: complements or substitutes for innovative performance? Technovation, 24(1):29–39, January 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0166497202000512.

Corbo:2023:CIR

[CKV^{+}23] Leonardo Corbo, Sascha Kraus, Božidar Vlačić, Marina Dabić, Andrea Caputo, and Massimiliano M. Pellegrini. Coopetition and innovation: a review and research agenda. Technovation, 122(??):??, April 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222001717.

Corsten:1988:IPS

[CL88] Hans Corsten and Otmar Lang. Innovation practice in small and medium-sized enterprises: an empirical survey of the member states of the European community. Technovation, 7(2):143–154, May 1988. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900442.

Chung:1997:TPBb

[CL97a] Sunyang Chung and Gunter Lay. Technology policy between “diversity” and “one best practice” — a comparison of Korean and German promotion schemes for new product technologies. Technovation, 17(11–12):728–729, November/ December 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497297904636.

Chung:1997:TPBa

[CL97b] Sunyang Chung and Gunter Lay. Technology policy between ‘diversity’ and ‘one best practice’ — a comparison of Korean and German promotion schemes for new production technologies. Technovation, 17(11–12):675–693, November/ December 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497297000564.

Chang:1998:CAT

[CL98] Zeph Yun Chang and Kee Beng Lee. Cultural aspect and technological development: a Chinese perspective. Technovation, 18(12):765–770, October 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000790.

Chan:2005:ATI

[CL05a] K. F. Chan and Theresa Lau. Assessing technology incubator programs in the science park: the good, the bad and the ugly. Technovation, 25(10):1215–1228, October 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204000677.

Chen:2005:ISM

[CL05b] Kuo-Min Chen and Ren-Jye Liu. Interface strategies in modular product innovation. Technovation, 25(7):771–782, July 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204000070.

Chandra:2012:UIE

[CL12] Yanto Chandra and Mark A. A. M. Leenders. User innovation and entrepreneurship in the virtual world: a study of Second Life residents. Technovation, 32(7–8):464–476, July/August 2012. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497212000284.

Choi:2021:TDR

[CL21] Mincheol Choi and Chang-Yang Lee. Technological diversification and R&D productivity: the moderating effects of knowledge spillovers and core-technology competence. Technovation, 104(??):??, June 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000304.

Choi:2022:DEB

[CL22] Jin-Uk Choi and Chang-Yang Lee. The differential effects of basic research on firm R&D productivity: the conditioning role of technological diversification. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222001067.

Cubric:2024:BCP

[CL24] Marija Cubric and Feng Li. Bridging the ‘Concept-Product’ gap in new product development: Emerging insights from the application of artificial intelligence in Fin-Tech SMEs. Technovation, 134(??):??, June 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000671.

Chen:2013:PAT

[CLC13] Mei-Fang Chen, Yi-Ping Lin, and Tsun-Jen Cheng. Public attitudes toward nanotechnology applications in Taiwan. Technovation, 33(2–3):88–96, February/March 2013. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497212001423.

Carmona-Lavado:2023:CII

[CLGFVCM23] Antonio Carmona-Lavado, Elena M. Gimenez-Fernandez, Vesna Vlaisavljevic, and Carmen Cabello-Medina. Crossindustry innovation: a systematic literature review. Technovation, 124(??):??, June 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497223000548.

Lo:2012:ESC

[cLhWCH12] Chih cheng Lo, Chun hsien Wang, Pei-Yu Chien, and Chien-Wei Hung. An empirical study of commercialization performance on nanoproducts. Technovation, 32(3–4):168–178, March/April 2012. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497211001246.

Coldrick:2005:ROS

[CLIH05] Simon Coldrick, Philip Longhurst, Paul Ivey, and John Hannis. An R&D options selection model for investment decisions. Technovation, 25(3):185–193, March 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000993.

Clouse:1991:DBA

[Clo91] Van G. H. Clouse. A decision-based approach to the investigation of new venture activity. Technovation, 11(4): 205–217, May 1991. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497291900342.

Cao:2023:UBR

[CLP23] Qinwei Cao, Yi Li, and Huatao Peng. From university basic research to firm innovation: diffusion mechanism and boundary conditions under a U-shaped relationship. Technovation, 123(??):??, May 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000299.

Cho:1996:ETS

[CLR96] Hyun-Dae Cho, Jae-Keun Lee, and Kony-Kyun Ro. Environment and technology strategy of firms in government R&D programmes in Korea. Technovation, 16(10):553–560, October 1996. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297828957.

Chen:2006:ANP

[CLT06] Hsing Hung Chen, Amy H. I. Lee, and Yunhuan Tong. Analysis of new product mix selection at TFT-LCD technological conglomerate network under uncertainty. Technovation, 26(11):1210–1221, November 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205001288.

Cheng:2010:EII

[CLW10] Cheng-Feng Cheng, Meng-Kuan Lai, and Wann-Yih Wu. Exploring the impact of innovation strategy on R&D employees’ job satisfaction: a mathematical model and empirical research. Technovation, 30(7–8):459–470, July/August 2010. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497210000477.

Chen:2018:PEC

[CLZ18] Xiafei Chen, Zhiying Liu, and Qingyuan Zhu. Performance evaluation of China’s high-tech innovation process: Analysis based on the innovation value chain. Technovation, 74– 75(??):42–53, June/July 2018. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497218301238.

Chen:2021:RPE

[CLZ21] Xiafei Chen, Zhiying Liu, and Qingyuan Zhu. Reprint of “Performance evaluation of China’s high-tech innovation process: Analysis based on the innovation value chain”. Technovation, ??(??):??, ???? 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497219300598.

Chen:2022:FPI

[CLZ^{+}22] Xiaoyan Chen, Giorgio Locatelli, Xinyue Zhang, Yunhao Gong, and Qinghua He. Firm and project innovation outcome measures in infrastructure megaprojects: an interpretive structural modelling approach. Technovation, 109(??): ??, January 2022. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497221001309.

Cai:2023:WHI

[CLZ23] Ying Cai, Jun Lin, and Ruxin Zhang. When and how to implement design thinking in the innovation process: a longitudinal case study. Technovation, 126(??):??, August 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722300127X.

Cartaya:1989:NOT

[CM89] Vanessa Cartaya and Ernesto Medina. New organizational techniques: a case study from Venezuela. Technovation, 9(6):519–530, October 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649728990031X.

Cave:1994:PAC

[CM94] Martin Cave and Jonathan Michie. The price of art: a case study. Technovation, 14(7):477–484, September 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900051.

Chinworth:1995:CBL

[CM95] Michael W. Chinworth and David C. Mowery. Cross-border linkages and the US defense industry: outlook and policy challenges. Technovation, 15(3):133–152, April 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729596615Z.

Corsten:1996:ANN

[CM96] Hans Corsten and Constantin May’. Artificial neural networks for supporting production planning and control. Technovation, 16(2):67–76, 92–93, February 1996. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497295000240.

Carayol:2004:ECS

[CM04] Nicolas Carayol and Mireille Matt. The exploitation of complementarities in scientific production process at the laboratory level. Technovation, 24(6):455–465, June 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720400063X.

Camarinha-Matos:2007:RAI

[CMA07] Luis M. Camarinha-Matos and Hamideh Afsarmanesh. Results assessment and impact creation in collaborative research — an example from the ECOLEAD project. Technovation, 27(1–2):65–77, January/February 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497206000940.

Chen:2023:SBI

[CMB23] Qian Chen, Mats Magnusson, and Jennie Björk. Selection bias of ideas for sustainability-oriented innovation in internal crowdsourcing. Technovation, 124(??):??, June 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722300072X.

Chang:2014:ISI

[CMH14] Yuan-Chieh Chang, Ian Miles, and Shih-Chang Hung. Introduction to special issue: Managing technology-service convergence in Service Economy 3.0. Technovation, 34(9): 499–504, September 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497214000790.

Corso:2006:MDW

[CMP^{+}06] Mariano Corso, Antonella Martini, Luisa Pellegrini, Silvia Massa, and Stefania Testa. Managing dispersed workers: the new challenge in Knowledge Management. Technovation, 26(5–6):583–594, May/June 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000490.

Cricelli:2023:POI

[CMS23] Livio Cricelli, Roberto Mauriello, and Serena Strazzullo. Preventing open innovation failures: a managerial framework. Technovation, 127(??):??, September 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722300144X.

Cantner:2010:INR

[CMtW10] Uwe Cantner, Andreas Meder, and Anne L. J. ter Wal. Innovator networks and regional knowledge base. Technovation, 30(9–10):496–507, September/October 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210000507.

Csath:1986:HED

[CN86] Magdolna Csath and Thomas H. Naylor. The Hungarian experience with deregulation. Technovation, 5(1–3):95–113, October 1986. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497286900465.

Crow:1990:TSD

[CN90] Michael M. Crow and Shrilata Nath. Technology strategy development in Japanese industry: an assessment of market and government influences. Technovation, 10(5):333–346, July 1990. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729090017E.

Crow:1992:TSD

[CN92] Michael M. Crow and Shrilata A. Nath. Technology strategy development in Korean industry: an assessment of market and government influences. Technovation, 12(2):119–136, March 1992. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729290019E.

Chabchoub:2005:EPP

[CN05] Norhène Chabchoub and Jorge Niosi. Explaining the propensity to patent computer software. Technovation, 25 (9):971–978, September 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497204000409.

Chaudhuri:2023:HPS

[CNN23] Atanu Chaudhuri, Hussein Naseraldin, and Gopalakrishnan Narayanamurthy. Healthcare 3D printing service innovation: Resources and capabilities for value cocreation. Technovation, 121(??):??, March 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222001432.

Cucit:2004:MOM

[CNPV04] L. Cucit, A. Nosella, G. Petroni, and C. Verbano. Management and organizational models of the European Space Agencies: the results of an empirical study. Technovation, 24(1):1–15, January 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497202000445.

Csaszar:2006:SCC

[CNS06] Felipe Csaszar, Miguel Nussbaum, and Marcos Sepulveda. Strategic and cognitive criteria for the selection of startups. Technovation, 26(2):151–161, February 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000453.

Cormican:2004:ABP

[CO04] Kathryn Cormican and David O’Sullivan. Auditing best practice for effective product innovation management. Technovation, 24(10):819–829, October 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000130.

Coccia:2012:PER

[Coc12] Mario Coccia. Political economy of R&D to support the modern competitiveness of nations and determinants of economic optimization and inertia. Technovation, 32(6):370– 379, June 2012. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649721200034X.

Colton:1981:NSF

[Col81] Robert M. Colton. National Science Foundation experience with university–industry centers for scientific research and technological innovation (an analysis of issues, characteristics and criteria for their establishment). Technovation, 1(2):97–108, August 1981. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497281900122.

Colombo:1989:UC

[Col89] Umberto Colombo. Umberto Colombo. Technovation, 9 (2–3):283–285, June 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497289900692.

Collins:1997:BRT

[Col97] G. C. S. Collins. Book review: The technical manager’s handbook — A survival guide: by Melvin Silverman, Chapman and Hall, London and New York, 1996. ISBN 0-412-991217, 505 pp., paperback. Technovation, 17(5):286, May 1997. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297882272.

Collinson:2008:BRH

[Col08] Simon Collinson. Book review: Haiyang Li (Ed.), Growth of New Technology Ventures in China’s Emerging Market (2006). 325 pp., Hardback, ISBN 1-84542-119-1. Technovation, 28(1–2):101, January/February 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000831.

Cooper:1990:SIP

[Coo90] Martin J. Cooper. Synthetic imogolite: a product in search of a market. Technovation, 10(2):69–71, March 1990. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729090029J.

Coombs:1997:BRI

[Coo97] Rod Coombs. Book review: Intellectual capital: by Annie Brooking, International Thomson Publishing, 1996, ISBN 1-861-52023-9, 204 pp. Technovation, 17(11–12):733–734, November/December 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497297887718.

Cooke:2007:BRS

[Coo07] Yolande Cooke. Book review: SMEs and New Technologies: Learning E-Business and Development: Banji Oyelaran-Oyeyinka, Kaushalesh Lal. Palgrave MacMillan, Houndmills, Basingstoke, Hampshire, 2006 (222 pp., £55), ISBN 0-230-00201-3. Technovation, 27(6–7):413, June/July 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000181.

Cook:2016:NHT

[Coo16] Gary Cook. News from the High Technology Small Firms Conference, University of Liverpool, 13–15 June 2016. Technovation, 57–58(??):41–42, November/December 2016. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497216302607.

Corsten:1987:PCB

[Cor87a] Hans Corsten. Problems with cooperation between universities and enterprises — a comparative study on size of enterprise. Technovation, 6(4):295–301, September 1987. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497287900769.

Corsten:1987:TTU

[Cor87b] Hans Corsten. Technology transfer from universities to small and medium-sized enterprises — an empirical survey from the standpoint of such enterprises. Technovation, 6(1): 57–68, April 1987. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497287900393.

Corsten:1987:URR

[Cor87c] Hans Corsten. Using the results of research financed and conducted by public bodies — a comparative review of the situation in the member states of the European community. Technovation, 6(3):155–167, August 1987. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497287900204.

Cornelius:2005:IVC

[Cor05] Barbara Cornelius. The institutionalisation of venture capital. Technovation, 25(6):599–608, June 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203002177.

Cousins:1990:BRI

[Cou90] Lesley Cousins. Book review: The innovation marathon: Lessons from high technology firms: By Mariann Jelinek and Claudia Bird Schoonhoven, Basil Blackwell, Oxford, 1990, £35.00, 469 pages, ISBN 0-631-15392-6. Technovation, 10(6):438–439, September 1990. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497290900077.

Cousins:1991:BRI

[Cou91] Lesley Cousins. Book review: The innovation marathon: Lessons from high technology firms: by Mariann Jelinek and Claudia Bird Schoonhoven, Basil Blackwell, Oxford, 1990, ISBN 0-631-15392-6, £35.00, 469 pages. Technovation, 11(2):127–128, March 1991. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729190049A.

Cozzarin:2006:WFI

[Coz06] Brian P. Cozzarin. Are world-first innovations conditional on economic performance? Technovation, 26(9):1017–1028, September 2006. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497205001665.

Carayannis:2005:PME

[CP05] Elias G. Carayannis and Denisa Popescu. Profiling a methodology for economic growth and convergence: learning from the EU e-procurement experience for Central and Eastern European countries. Technovation, 25(1):1–14, January 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203000713.

Cornelius:2006:WWV

[CP06] Barbara Cornelius and Olle Persson. Who’s who in venture capital research. Technovation, 26(2):142–150, February 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205000866.

Choi:2007:OPI

[CP07] David Y. Choi and Arturo Perez. Online piracy, innovation, and legitimate business models. Technovation, 27(4):168– 178, April 2007. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497206001040.

Cannarella:2011:TCI

[CP11] Carmelo Cannarella and Valeria Piccioni. Traditiovations: Creating innovation from the past and antique techniques for rural areas. Technovation, 31(12):689–699, December 2011. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497211001064.

Chakraborty:2023:HAP

[CP23] Debarun Chakraborty and Justin Paul. Healthcare apps’ purchase intention: a consumption values perspective. Technovation, 120(??):??, February 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222000281.

Cetindamar:2009:UTM

[CPP09] Dilek Cetindamar, Robert Phaal, and David Probert. Understanding technology management as a dynamic capability: a framework for technology management activities. Technovation, 29(4):237–246, April 2009. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208001338.

Candelin-Palmqvist:2012:IPR

[CPSM12] Hanni Candelin-Palmqvist, Birgitta Sandberg, and Ulla-Maija Mylly. Intellectual property rights in innovation management research: a review. Technovation, 32(9–10):502–512, September/October 2012. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497212000065.

Carayannis:2006:TLE

[CPSS06] Elias G. Carayannis, Denisa Popescu, Caroline Sipp, and McDonald Stewart. Technological learning for entrepreneurial development (TL4ED) in the knowledge economy (KE): Case studies and lessons learned. Technovation, 26(4):419–443, April 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497205000738.

Caccamo:2023:BOK

[CPT23] Marta Caccamo, Daniel Pittino, and Fredrik Tell. Boundary objects, knowledge integration, and innovation management: a systematic review of the literature. Technovation, 122(??):??, April 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497222001961.

Chin:2002:ABS

[CPXC02] K.-S. Chin, K.-F. Pun, Y. Xu, and J. S. F. Chan. An AHP based study of critical factors for TQM implementation in Shanghai manufacturing industries. Technovation, 22(11):707–715, November 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000657.

Chen:2003:NTL

[CQ03] J. Chen and W. G. Qu. A new technological learning in China. Technovation, 23(11):861–867, November 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720200038X.

Clarke:1981:ETE

[CR81] Thomas E. Clarke and Jean Reavley. Educating technical entrepreneurs and innovators for the 1980’s. Technovation, 1(2):125–134, August 1981. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497281900146.

Carayannis:2000:DVG

[CR00] Elias G. Carayannis and Robie I. Samanta Roy. Davids vs Goliaths in the small satellite industry:: the role of technological innovation dynamics in firm competitiveness. Technovation, 20(6):287–297, June 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001376.

Coccia:2002:TTA

[CR02] Mario Coccia and Secondo Rolfo. Technology transfer analysis in the Italian National Research Council. Technovation, 22(5):291–299, May 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000189.

Chen:2006:RTI

[CR06] Xiangdong Chen and Guido Reger. The role of technology in the investment of German firms in China. Technovation, 26(3):407–415, March 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497204002147.

Coccia:2008:SCP

[CR08] Mario Coccia and Secondo Rolfo. Strategic change of public research units in their scientific activity. Technovation, 28(8):485–494, August 2008. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497208000163.

Conn:2019:GIR

[CR19] Steffen Conn and Paavo Ritala. The growing importance of research methods in innovation management research: a note from the 2018 ISPIM Innovation Conference. Technovation, 80–81(??):1–2, February/March 2019. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497218306655.

Crespi:2004:BRI

[Cre04] Gustavo A. Crespi. Book review: Innovation and knowledge creation in an open economy. Canadian industry and international implications: John R. Baldwin and Peter Hanel; Cambridge University Press, Cambridge, UK, 2003, 512 pages (including Index), ISBN 0-521-81086-8. Technovation, 24(6):513–514, June 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204000100.

Carayannis:1998:HTS

[CRKA98] Elias G. Carayannis, Everett M. Rogers, Kazuo Kurihara, and Marcel M. Allbritton. High-technology spinoffs from government R&D laboratories and research universities. Technovation, 18(1):1–11, January 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297001016.

Chakrabarti:1984:CFT

[CS84] Alok K. Chakrabarti and William E. Souder. Critical factors in technological innovation and their policy implications. Technovation, 2(4):255–275, July 1984. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497284900063.

Chakrabarti:1987:TIP

[CS87] Alok K. Chakrabarti and William E. Souder. Technology, innovation and performance in corporate mergers: a managerial evaluation. Technovation, 6(2):103–114, June 1987. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497287900149.

Cook:1989:ETD

[CS89] P. Lesley Cook and John Surrey. Energy technology in developing countries: the scope for government policy. Technovation, 9(5):431–451, August 1989. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649728990014X.

Corsten:1998:CMS

[CS98] Hans Corsten and Stephan Stuhlmann. Capacity management in service organisations. Technovation, 18(3):163–178, March 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297901218.

Chen:2000:TTC

[CS00] X. Chen and C. Sun. Technology transfer to China: alliances of Chinese enterprises with Western technology exporters. Technovation, 20(7):353–362, July 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001480.

Carayannis:2001:NVO

[CS01a] Elias Carayannis and John Sagi. “New” vs. “old” economy: insights on competitiveness in the global IT industry. Technovation, 21(8):501–514, August 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000729.

Carayannis:2001:DPC

[CS01b] Elias G. Carayannis and John Sagi. Dissecting the professional culture: insights from inside the IT “black box”. Technovation, 21(2):91–98, February 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000274.

Carayannis:2002:EON

[CS02] Elias G. Carayannis and John Sagi. Exploiting opportunities of the new economy: developing nations in support of the ICT industry. Technovation, 22(8):517–524, August 2002. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201000463.

Chang:2004:IST

[CS04] Pao-Long Chang and Hsin-Yu Shih. The innovation systems of Taiwan and China: a comparative analysis. Technovation, 24(7):529–539, July 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202001177.

Chang:2005:CPI

[CS05] Pao-Long Chang and Hsin-Yu Shih. Comparing patterns of intersectoral innovation diffusion in Taiwan and China: a network analysis. Technovation, 25(2):155–169, February 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203000774.

Carbonara:2006:CMA

[CS06] Nunzia Carbonara and Barbara Scozzi. Cognitive maps to analyze new product development processes: a case study. Technovation, 26(11):1233–1243, November 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205001240.

Candi:2008:OWE

[CS08a] Marina Candi and Rognvaldur Saemundsson. Oil in water? Explaining differences in aesthetic design emphasis in new technology-based firms. Technovation, 28(7):464–471, July 2008. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497208000321.

Cheng:2008:RIC

[CS08b] Chi Jyun Cheng and Eric C. C. Shiu. Re-innovation: the construct, measurement, and validation. Technovation, 28 (10):658–666, October 2008. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497207001095.

Cheng:2012:VPI

[CS12] Colin C. Cheng and Eric C. Shiu. Validation of a proposed instrument for measuring eco-innovation: an implementation perspective. Technovation, 32(6):329–344, June 2012. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497212000272.

Cowling:2013:PIU

[CS13] Marc Cowling and Josh Siepel. Public intervention in UK small firm credit markets: Value-for-money or waste of scarce resources? Technovation, 33(8–9):265–275, August/ September 2013. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497212001241.

Chandna:2021:WTS

[CS21] Vallari Chandna and Manjula S. Salimath. When technology shapes community in the Cultura and Craft Industries: Understanding virtual entrepreneurship in online ecosystems. Technovation, ??(??):??, ???? 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497217300597.

Chang:1994:FPT

[CSH94] Pao-Long Chang, Chintay Shih, and Chiung-Wen Hsu. The formation process of Taiwan’s IC industry — method of technology transfer. Technovation, 14(3):161–171, April 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294900531. See [Ano94l].

Cheng:2006:CAI

[CSL06] Julian Ming-Sung Cheng, Gwo-Ji Sheen, and Guan-Cheng Lou. Consumer acceptance of the Internet as a channel of distribution in Taiwan — a channel function perspective. Technovation, 26(7):856–864, July 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000027.

Chang:2023:WCR

[CSL23] Yu-Yu Chang, Hui-Yu Shih, and Bou-Wen Lin. Work curiosity and RD professionals’ creative performance: Scientists vs. engineers. Technovation, 124(??):??, June 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000500.

Campisi:1992:DSD

[CT92] Domenico Campisi and Carlo Tesauro. The diffusion and spatial distribution of new telecommunication technologies in the Italian region of Campania. Technovation, 12(6):355–368, September 1992. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729290001X.

Campisi:1997:TRT

[CT97] Domenico Campisi and Carlo Tesauro. Telecommunication rates and territorial aggregations. Technovation, 17(5):267– 277, May 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497296001137.

Chang:2002:FNP

[CT02] Pao-Long Chang and Chien-Tzu Tsai. Finding the niche position — competition strategy of Taiwan’s IC design industry. Technovation, 22(2):101–111, February 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000936.

Carayannis:2006:IDT

[CT06] Elias G. Carayannis and Eric Turner. Innovation diffusion and technology acceptance: the case of PKI technology. Technovation, 26(7):847–855, July 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205001021.

Casey:2012:DTS

[CT12] Thomas R. Casey and Juuso Töyli. Dynamics of two-sided platform success and failure: an analysis of public wireless local area access. Technovation, 32(12):703–716, December 2012. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497212000958.

Chin:2002:QMP

[CTC02] K. S. Chin, V. M. Rao Tummala, and K. M. Chan. Quality management practices based on seven core elements in Hong Kong manufacturing industries. Technovation, 22(4):213– 230, April 2002. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497201000207.

Corvello:2024:HSU

[CTSJ24] Vincenzo Corvello, Ciro Troise, Giovanni Schiuma, and Paul Jones. How start-ups translate learning from innovation failure into strategies for growth. Technovation, 134 (??):??, June 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497224001019.

Chang:2002:DTT

[CTYH02] Pao-Cheng Chang, Nien-Tao Tsou, Benjamin J. C. Yuan, and Chien-Chun Huang. Development trends in Taiwan’s opto-electronics industry. Technovation, 22(3):161–173, March 2002. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201000086.

Cirone:2013:PMF

[CU13] Alexandra E. Cirone and Johannes Urpelainen. Political market failure? The effect of government unity on energy technology policy in industrialized democracies. Technovation, 33(10–11):333–344, October/November 2013. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213000631.

Cenamor:2013:RCP

[CUF13] Javier Cenamor, Belén Usero, and Zulima Fernández. The role of complementary products on platform adoption: Evidence from the video console market. Technovation, 33 (12):405–416, December 2013. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497213000692.

Cullen:2000:CCO

[Cul00] Penny-Anne Cullen. Contracting, co-operative relations and extended enterprises. Technovation, 20(7):363–372, July 2000. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497299001728.

Curien:1989:AFC

[Cur89] Hubert Curien. Actions to facilitate cooperation between industries, universities and other research organizations: Attitudes and experiences of governmental institutions. Technovation, 9(2–3):235–239, June 1989. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649728990062X.

Candi:2013:OIP

[CvdEG13] Marina Candi, Jan van den Ende, and Gerda Gemser. Organizing innovation projects under technological turbulence. Technovation, 33(4–5):133–141, April/May 2013.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213000163.

Carayannis:2005:AGG

[CvZ05] Elias G. Carayannis and Maximilian von Zedtwitz. Architecting gloCal (global–local), real–virtual incubator networks (G-RVINs) as catalysts and accelerators of entrepreneurship in transitioning and developing economies: lessons learned and best practices from current development and business incubation practices. Technovation, 25(2):95– 110, February 2005. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203000725.

Corsten:1993:BPS

[CW93] Hans Corsten and Thomas Will. Basic points for a strategic production organization requiring simultaneous strategies. Technovation, 13(2):73–81, March 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390055Z. See [Ano93b].

Corsten:1994:SSG

[CW94] Hans Corsten and Thomas Will. Simultaneously supporting generic competitive strategies by production management: supportive concepts of information technology and work organization. Technovation, 14(2):111–120, March 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294901007. See [Ano94-33].

Chipika:2006:ETL

[CW06] Stephen Chipika and Gordon Wilson. Enabling technological learning among light engineering SMEs in Zimbabwe through networking. Technovation, 26(8):969–979, August 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205001446.

Cheng:2022:HCC

[CW22] Cong Cheng and Limin Wang. How companies configure digital innovation attributes for business model innovation? A configurational view. Technovation, 112(??):??, April 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221001796.

Cetindamar:2009:DTM

[CWAB09] D. Cetindamar, S. N. Wasti, H. Ansal, and B. Beyhan. Does technology management research diverge or converge in developing and developed countries? Technovation, 29(1):45– 58, January 2009. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497208000552.

Cooper:1998:RCTb

[CWB98a] Rachel Cooper, Andrew Wootton, and Margaret Bruce. ‘Requirement capture’: theory and practice. Technovation, 18(8–9):585, August/September 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298901213.

Cooper:1998:RCTa

[CWB98b] Rachel Cooper, Andrew B. Wootton, and Margaret Bruce. “Requirements capture”: theory and practice. Technovation, 18(8–9):497–511, 585, August/September 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000339.

Crute:2003:ILA

[CWBG03] V. Crute, Y. Ward, S. Brown, and A. Graves. Implementing Lean in aerospace-challenging the assumptions and understanding the challenges. Technovation, 23(12):917–928, December 2003. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203000816.

Chen:2014:BCS

[CWY14] Shin-Horng Chen, Pei-Chang Wen, and Chih-Kai Yang. Business concepts of systemic service innovations in e-Healthcare. Technovation, 34(9):513–524, September 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214000340.

Chen:2024:PPL

[CX24] Ting Chen and Yu Xiong. Product positioning of lowcarbon products based on blockchain-enabled product communities. Technovation, 132(??):??, April 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000415.

Cho:2000:IFC

[CY00] Dae-Hyun Cho and Pyung-Il Yu. Influential factors in the choice of technology acquisition mode: an empirical analysis of small and medium size firms in the Korean telecommunication industry. Technovation, 20(12):691–704, December 2000. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497299001820.

Chang:2012:ISC

[CY12] Yuan-Chieh Chang and HsiuJu Rebecca Yen. Introduction to the special cluster on managing technology-service fusion innovation. Technovation, 32(7–8):415–418, July/ August 2012. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497212000454.

Chen:2019:EEP

[CYHC19] Guanghua Chen, Guoliang Yang, Feng He, and Kaihua Chen. Exploring the effect of political borders on universityindustry collaborative research performance: Evidence from China’s Guangdong province. Technovation, 82–83(??): 58–69, April/May 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497216301699.

Chao:2007:DTT

[CYJ07] Chia-Chen Chao, Jiann-Min Yang, and Wen-Yuan Jen. Determining technology trends and forecasts of RFID by a historical review and bibliometric analysis from 1991 to 2005. Technovation, 27(5):268–279, May 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497206001003. See comments [Ho09].

Chang:2019:ELK

[CYK19] Ching-Wen Chang, Takayuki Yamanaka, and Shingo Kano. An enforced loop-out knowledge flow facilitates industry competition: Learning from the pharmaceutical and genetically modified seed industries. Technovation, 79(??):11–24, January 2019. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497216300700.

Chang:2016:EUR

[CYM^{+}16] Yuan-Chieh Chang, Phil Yihsing Yang, Ben R. Martin, Hui-Ru Chi, and Tung-Fei Tsai-Lin. Entrepreneurial universities and research ambidexterity: a multilevel analysis. Technovation, 54(??):7–21, August 2016. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497216300074.

Cui:2024:ADC

[CZL^{+}24] Xiling Cui, Zhongshan Zhu, Libo Liu, Qiang Zhou, and Qiang Liu. Anomaly detection in consumer review analytics for idea generation in product innovation: Comparing machine learning and deep learning techniques. Technovation, 134(??):??, June 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000786.

Chen:2023:NEU

[CZM23] Yufeng Chen, Shun Zhang, and Jiafeng Miao. The negative effects of the US–China trade war on innovation: Evidence from the Chinese ICT industry. Technovation, 123(??):??, May 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223000457.

Chen:2021:DRI

[CZZM21] Kaihua Chen, Yi Zhang, Guilong Zhu, and Rongping Mu. Do research institutes benefit from their network positions in research collaboration networks with industries or/ and universities? Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497217307836.

DaSilveira:2001:IDR

[Da 01] Giovani Da Silveira. Innovation diffusion: research agenda for developing economies. Technovation, 21(12):767–773, December 2001. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497201000074.

Dionco-Adetayo:2001:UWW

[DA01] Erlinda A. Dionco-Adetayo. Utilization of wood wastes in Nigeria: a feasibility overview. Technovation, 21(1):55–60, January 2001. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497200000031.

deArroyabe:2021:DCB

[dAASA21] J. C. Fernandez de Arroyabe, N. Arranz, M. Schumann, and M. F. Arroyabe. The development of CE business models in firms: the role of circular economy capabilities. Technovation, 106(??):??, August 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000730.

Burcharth:2014:NIN

[dABKS14] Ana Luiza de Araújo Burcharth, Mette Praest Knudsen, and Helle Alsted Søndergaard. Neither invented nor shared here: the impact and management of attitudes for the adoption of open innovation practices. Technovation, 34(3):149– 161, March 2014. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497213001454.

Daghfous:2004:EIR

[Dag04] Abdelkader Daghfous. An empirical investigation of the roles of prior knowledge and learning activities in technology transfer. Technovation, 24(12):939–953, December 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203000592.

Dan:2019:HIF

[Dan19] Sujan M. Dan. How interface formats gain market acceptance: the role of developers and format characteristics in the development of de facto standards. Technovation, 88(??):Article 102054, December 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497217301943.

Obra:2002:EOA

[dAOCM02] Ana Rosa del Aguila Obra, Sebastián Bruque Cámara, and Antonio Padilla Meléndez. The economic and organizational aspects of telecentres: the Spanish case. Technovation, 22(12):785–798, December 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000329.

Dionco-Adetayo:2002:FIS

[DAOO02] E. A. Dionco-Adetayo, O. F. Olaniyan, and B. O. Ogunba. Factors influencing soybean utilization in the household consumer market. Technovation, 22(2):129–133, February 2002. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497200000766.

Desouza:2007:MIT

[DAR07] Kevin C. Desouza, Yukika Awazu, and Arkalgud Ramaprasad. Modifications and innovations to technology artifacts. Technovation, 27(4):204–220, April 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0166497206001015.

Davies:1992:BRK

[Dav92] Robert Davies. Book review: The knowledge link: by J. L. Badaracco, Harvard Business School Press, Boston, 1991, ISBN 0-87584-226-7, £22.95, 189 pages. Technovation, 12 (2):137–138, March 1992. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497292900219.

Davies:1993:ICT

[Dav93] Howard Davies. The information content of technology transfers: a transactions cost analysis of the machine tool industry. Technovation, 13(2):93–100, March 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497293900573. See [Ano93i].

Dayasindhu:2002:EKT

[Day02] N. Dayasindhu. Embeddedness, knowledge transfer, industry clusters and global competitiveness: a case study of the Indian software industry. Technovation, 22(9):551–560, September 2002. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497201000980.

Donini:1997:MOIa

[DB97a] Benito Romeo Donini and Giancarlo Barbiroli. Measuring and optimizing the input technical effectiveness and efficiency of production processes by means of the linear programming method. Technovation, 17(11–12):667–674, November/December 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297000588.

Donini:1997:MOIb

[DB97b] Benito Romeo Donini and Giancarlo Barbiroli. Measuring and optimizing the inputs technical effectiveness and efficiency of production by means of the linear programming method. Technovation, 17(11–12):728, November/ December 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497297904624.

DeCoster:2005:APN

[DB05] Rebecca De Coster and Clive Butler. Assessment of proposals for new technology ventures in the UK: characteristics of university spin-off companies. Technovation, 25(5):535– 543, May 2005. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203001913.

Daghfous:2009:SMI

[DB09] Abdelkader Daghfous and Reza Barkhi. The strategic management of information technology in UAE hotels: an exploratory study of TQM, SCM, and CRM implementations. Technovation, 29(9):588–595, September 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209000765.

diBenedetto:2010:COI

[dB10] Anthony di Benedetto. Comment on ‘is open innovation a field of study or a communication barrier to theory development?’. Technovation, 30(11–12):557, November/ December 2010. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497210000957.

Dziallas:2019:IIT

[DB19] Marisa Dziallas and Knut Blind. Innovation indicators throughout the innovation process: an extensive literature analysis. Technovation, 80–81(??):3–29, February/ March 2019. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497217301402.

delBrio:2002:MEA

[dBJ02] J. Á. del Brío and B. Junquera. Managerial environmental awareness and cooperation with public governments in Spanish industrial companies. Technovation, 22(7):445–452, July 2002. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201000335.

delBrio:2003:RLE

[dBJ03] Jesús Ángel del Brío and Beatriz Junquera. A review of the literature on environmental innovation management in SMEs: implications for public policies. Technovation, 23(12):939–948, December 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000366.

Decter:2007:UBT

[DBL07] Moira Decter, David Bennett, and Michel Leseure. University to business technology transfer-UK and USA comparisons. Technovation, 27(3):145–155, March 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497206000265.

Dong:2024:THC

[DBR24] Mengchen Dong, Jean-François Bonnefon, and Iyad Rahwan. Toward human-centered AI management: Methodological challenges and future directions. Technovation, 131 (??):??, March 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497224000038.

Debackere:1997:SCD

[DBV97] Koen Debackere, Dirk Buyens, and Tine Vandenbossche. Strategic career development for R&D professionals: Lessons from field research. Technovation, 17(2):53–62, 101, February 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497297841910.

Dickson:2000:TDP

[DC00] Keith Dickson and Anne-Marie Coles. Textile design protection:: Copyright, CAD and competition. Technovation, 20(1):47–53, January 2000. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497299000966.

deCampos:2007:BRR

[dC07] André Luiz de Campos. Book review: Renewing Unilever: Transformation and Tradition: G. Jones. Oxford University Press, Oxford, UK, New York (2005), pp. xvii + 447, £22.50 (paperback) and £60.00; ISBN 0-19-926944-0 (paperback) and 978-019-926944-0 (paperback), 0-19-926943-2 and 978-019-926943-3. Technovation, 27(11):716–717, November 2007. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649720600099X.

Dai:2022:RTI

[DC22] Xiaoyong Dai and Gary Chapman. R&D tax incentives and innovation: Examining the role of programme design in China. Technovation, 113(??):??, May 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497221002005.

DelSarto:2022:SAO

[DCD22] Nicola Del Sarto, Claudio Cruz Cazares, and Alberto Di Minin. Startup accelerators as an open environment: the impact on startups’ innovative performance. Technovation, 113(??):??, May 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497221002066.

Haro-Dominguez:2007:IAC

[dCHDAALMM07] M. del Carmen Haro-Domínguez, Daniel Arias-Aranda, Francisco Javier Lloréns-Montes, and Antonia Ruíz Moreno. The impact of absorptive capacity on technological acquisitions engineering consulting companies. Technovation, 27(8):417–425, August 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497207000582.

Martin-de-Castro:2008:PKC

[dCLSNL08] Gregorio Martín de Castro, Pedro López-Sáez, and José E. Navas-López. Processes of knowledge creation in knowledge-intensive firms: Empirical evidence from Boston’s Route 128 and Spain. Technovation, 28(4):222– 230, April 2008. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497207001344.

Dabic:2023:LOI

[DDBM23] Marina Dabić, Tugrul Daim, Marcel L. A. M. Bogers, and Anne-Laure Mention. The limits of open innovation: Failures, risks, and costs in open innovation practice and theory. Technovation, 126(??):??, August 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000974.

Dicuonzo:2023:HSM

[DDFS23] Grazia Dicuonzo, Francesca Donofrio, Antonio Fusco, and Matilda Shini. Healthcare system: Moving forward with artificial intelligence. Technovation, 120(??):??, February 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497222000578.

Davenport:1998:CRP

[DDG98] Sally Davenport, John Davies, and Charlotte Grimes. Collaborative research programmes: building trust from difference. Technovation, 19(1):31–40, November 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000832.

Dekker:2024:ECD

[DDN24] Henri C. Dekker, Carole Donada, and Gwenaëlle Nogatchewsky. Exploiting cognitive distance for enhanced innovative capabilities: the role of collaboration controls in incumbent-startup partnerships. Technovation, 134(??):??, June 2024. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649722400097X.

Dercole:2008:ADT

[DDOR08] Fabio Dercole, Ulf Dieckmann, Michael Obersteiner, and Sergio Rinaldi. Adaptive dynamics and technological change. Technovation, 28(6):335–348, June 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207001496.

DeGregori:1985:TLF

[De 85] Thomas R. De Gregori. Technological limits to forecasts of doom: Science, technology, and the sustainable economy. Technovation, 3(3):209–220, August 1985. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649728590015X.

DeSaulles:2004:BRI

[De 04] Martin De Saulles. Book review: Informations systems and the economics of innovation: C. Avgerou and R. L. La Rovere (Eds). Cheltenham: Edward Elgar Publishing. 2003. pp. 219 (incl. Index), £49.95, ISBN 1-84376-018-5 (hardback). Technovation, 24(1):82–83, January 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203001123.

Lima:2023:HDS

[De 23] Wenderson De Lima. How do social entrepreneurs compete? Entrepreneurial legitimacy in the Silicon Valley of humanitarian innovations. Technovation, 125(??):??, July 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000780.

Dean:1981:SUS

[Dea81] Charles W. Dean. A study of university/small business interaction for technology transfer. Technovation, 1(2):109– 123, August 1981. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497281900134.

Delany:1997:BRM

[Del97] Mike Delany. Book review: Managing business processes; BPR and beyond: by Colin Armistead and Philip Rowland, John Wiley and Sons, Chichester, 1996, xiv + 362 pp., £24.95 (hardback), ISBN 0-471-95490-X. Technovation, 17(10):599, October 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297890888.

Deniozos:1994:SIT

[Den94] Dimitris Deniozos. Steps for the introduction of technology management in developing economies: the role of public governments. Technovation, 14(3):197–203, April 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900566. See [Ano94-34].

Derbyshire:2014:IAE

[Der14] James Derbyshire. The impact of ambidexterity on enterprise performance: Evidence from 15 countries and 14 sectors. Technovation, 34(10):574–581, October 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214000716.

Despres:1996:ITC

[Des96] Charles Jean-Noel Despres. Information, technology and culture. An ethnography of information technology and modernist business organization. Technovation, 16(1):1–20, January 1996. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497295000283.

Dumas:1998:TNMa

[DF98a] Angela Dumas and Andrew Fentem. Totemics: new metaphor techniques to manage knowledge from discovery to storage and retrieval. Technovation, 18(8– 9):513–521, 585–586, August/September 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000364.

Dumas:1998:TNMb

[DF98b] Angela Dumas and Andrew Fentem. Totemics: new metaphor techniques to manage knowledge from discovery to storage and retrieval. Technovation, 18(8–9):585– 586, August/September 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497298901225.

Durrani:1998:MTAa

[DFBC98a] T. S. Durrani, S. M. Forbes, C. Broadfoot, and A. S. Carrie. Managing the technology acquisition process. Technovation, 18(8–9):523–528, 587, August/September 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000479.

Durrani:1998:MTAb

[DFBC98b] T. S. Durrani, S. M. Forbes, C. Broadfoot, and A. S. Carrie. Managing the technology acquisition process. Technovation, 18(8–9):587, August/September 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298901237.

Duret:1991:CMI

[DFD^{+}91] D. Duret, T. Ferrari, U. Debrun, E. Werner, and P. Degoul. Custom-made information and the information engineer. Technovation, 11(2):113–119, March 1991. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497291900412.

DaiPraMartens:2022:CII

[DFFM22] Cristina Dai Prá Martens, Luciano Ferreira da Silva, Deivison Feitosa Silva, and Mauro Luiz Martens. Challenges in the implementation of Internet of Things projects and actions to overcome them. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016649722100208X.

Drejer:2002:TMP

[DG02] Anders Drejer and Agnar Gudmundsson. Towards multiple product development. Technovation, 22(12):733–745, December 2002. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649720200024X.

Dewangan:2014:THE

[DG14] Vikas Dewangan and Manish Godse. Towards a holistic enterprise innovation performance measurement system. Technovation, 34(9):536–545, September 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214000376.

Silva:2021:ASC

[DGMR21] Muthu De Silva, Leonid Gokhberg, Dirk Meissner, and Margherita Russo. Addressing societal challenges through the simultaneous generation of social and business values: a conceptual framework for science-based co-creation. Technovation, 104(??):??, June 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000493.

Dangelico:2010:SDM

[DGP10] Rosa Maria Dangelico, Achille Claudio Garavelli, and Antonio Messeni Petruzzelli. A system dynamics model to analyze technology districts’ evolution in a knowledge-based perspective. Technovation, 30(2):142–153, February 2010. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209001436.

Drago:2023:TTT

[DGR23] Carlo Drago, Andrea Gatto, and Matteo Ruggeri. Telemedicine as technoinnovation to tackle COVID-19: a bibliometric analysis. Technovation, 120(??):??, February 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722100198X.

Donis:2023:ECP

[DGS23] Silvia Donis, Jaime Gómez, and Idana Salazar. Economic complexity, property rights and the judicial system as drivers of eco-innovations: an analysis of OECD countries. Technovation, 128(??):??, December 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001797.

deGuimaraes:2021:JTS

[dGSJ^{+}21] Julio Cesar Ferro de Guimarães, Eliana Andréa Severo, Charbel Jose Chiappetta Jabbour, Ana Beatriz Lopes de Sousa Jabbour, and Ariane Ferreira Porto Rosa. The journey towards sustainable product development: why are some manufacturing companies better than others at product innovation? Technovation, 103(??):??, May 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000201.

Dowling:2006:PDS

[DH06] Michael Dowling and Roland Helm. Product development success through cooperation: a study of entrepreneurial firms. Technovation, 26(4):483–488, April 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720500101X.

denHartigh:2016:PCD

[dHOvdKS16] Erik den Hartigh, J. Roland Ortt, Geerten van de Kaa, and Claire C. M. Stolwijk. Platform control during battles for market dominance: the case of Apple versus IBM in the early personal computer industry. Technovation, 48–49(??):4–12, February/March 2016. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497215000917.

Delgosha:2022:DII

[DHT22] Mohammad Soltani Delgosha, Nastaran Hajiheydari, and Mojtaba Talafidaryani. Discovering IoT implications in business and management: a computational thematic analysis. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497221000171.

DIppolito:2014:IDF

[D’I14] Beatrice D’Ippolito. The importance of design for firms’ competitiveness: a review of the literature. Technovation, 34(11):716–730, November 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016649721400008X.

Dibiaggio:2000:BRE

[Dib00] L. Dibiaggio. Book review: The Economics of Science and Innovation: P. E. Stephan and D. B. Audretsch (Eds); The International Library of Critical Writings in Economics, Edward Elgar, Cheltenham, 2000, 2 volumes. Technovation, 20(8):459–460, August 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000213.

Dibiaggio:2001:BRS

[Dib01] Ludo Dibiaggio. Book review: Systems of innovation: growth, competitiveness and employment: Charles Edquist and Maureen McKelvey (Eds); Edward Elgar, Cheltenham, and Northampton, Mass., 2000, 2 volumes. Technovation, 21(8):535–537, August 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000141.

Dibiaggio:2002:BRC

[Dib02a] Ludovic Dibiaggio. Book review: Classics in management thought: Cary L. Cooper (Ed.); Edward Elgar, Cheltenham, UK. Northampton, MA, USA. Technovation, 22(1):61–62, January 2002. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201000712.

Dibiaggio:2002:BRM

[Dib02b] Ludovic Dibiaggio. Book review: Managing intellectual capital. Organizational, strategic, and policy dimensions: David J. Teece; Oxford University Press, Oxford. Technovation, 22(1):62–64, January 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000724.

Dibiaggio:2004:BRC

[Dib04] Ludovie Dibiaggio. Book review: Corporate technological competence and the evolution of technological diversification: Felicia M. Fai, Cheltenham. Edward Elgar publishing. 2003. 194 pp. Technovation, 24(9):759–760, September 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204000665.

DelSarto:2021:RAF

[DID21] Nicola Del Sarto, Diane A. Isabelle, and Alberto Di Minin. The role of accelerators in firm survival: an fsQCA analysis of Italian startups. Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497218306424.

Diez:2000:INM

[Die00] Javier Revilla Diez. Innovative networks in manufacturing: some empirical evidence from the metropolitan area of Barcelona. Technovation, 20(3):139–150, March 2000. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001121.

Drejer:2005:DCK

[DJ05] Ina Drejer and Birte Holst Jørgensen. The dynamic creation of knowledge: Analysing public-private collaborations. Technovation, 25(2):83–94, February 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000750.

Datta:2013:LBF

[DJ13] Avimanyu Datta and Leonard M. Jessup. Looking beyond the focal industry and existing technologies for radical innovations. Technovation, 33(10–11):355–367, October/ November 2013. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497213000618.

Delerue:2010:JMR

[DL10] Hélène Delerue and Albert Lejeune. Job mobility restriction mechanisms and appropriability in organizations: the mediating role of secrecy and lead time. Technovation, 30(5–6):359–366, May/June 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210000118.

Dai:2021:DDP

[DLC21] Xiaoyong Dai, Yanchao Li, and Kaihua Chen. Direct demand-pull and indirect certification effects of public procurement for innovation. Technovation, 101(??):??, March 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497220300705.

laHiz:2022:IUR

[dlHBR22] Dante I. Leyva de la Hiz and María Teresa Bolívar-Ramos. The inverted U relationship between green innovative activities and firms’ market-based performance: the impact of firm age. Technovation, 110(??):??, February 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722100153X.

Du:2024:SCC

[DLKM24] Yunzhou Du, Qiuchen Liu, Phillip H. Kim, and Johannes Meuer. Studying complex causal processes in technological innovation and entrepreneurship with set-theoretic mediation models. Technovation, 134(??):??, June 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000658.

Dettwiler:2006:ULC

[DLL06] Paul Dettwiler, Peter Lindelöf, and Hans Löfsten. Utility of location: a comparative survey between small new technology-based firms located on and off Science Parks — implications for facilities management. Technovation, 26(4): 506–517, April 2006. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497205000878.

Lee:2006:RDL

[dLP06] Jeong dong Lee and Chansoo Park. Research and development linkages in a national innovation system: Factors affecting success and failure in Korea. Technovation, 26(9): 1045–1054, September 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497205001203.

Du:2007:IDE

[DLR07] Jun Du, James H. Love, and Stephen Roper. The innovation decision: an economic analysis. Technovation, 27(12):766– 773, December 2007. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497207000703.

Dong:2023:HUI

[DLTQ23] Caiting Dong, Xielin Liu, Fangcheng Tang, and Shumin Qiu. How upstream innovativeness of ecosystems affects firms’ innovation: the contingent role of absorptive capacity and upstream dependence. Technovation, 124(??):??, June 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223000469.

Dwyer:1991:CEP

[DM91] Larry Dwyer and Robert Mellor. Corporate environment and the proficiency of new product process activities. Technovation, 11(2):63–78, March 1991. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497291900386.

Doloreux:2008:DIQ

[DM08] David Doloreux and Yannik Melançon. On the dynamics of innovation in Quebec’s coastal maritime industry. Technovation, 28(4):231–243, April 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207001381.

Dedehayir:2011:MIC

[DM11] Ozgur Dedehayir and Saku J. Mäkinen. Measuring industry clockspeed in the systemic industry context. Technovation, 31(12):627–637, December 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649721100109X.

deMan:2005:CIR

[dMD05] Ard-Pieter de Man and Geert Duysters. Collaboration and innovation: a review of the effects of mergers, acquisitions and alliances on innovation. Technovation, 25(12): 1377–1387, December 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016649720400135X.

Dang:2022:GPM

[DMH22] Jianwei Dang, Kazuyuki Motohashi, and Dong Huo. Get pennies from many or get a dollar from one? Multiple licensing in markets for technology. Technovation, 113(??): ??, May 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497221002042.

DEste:2012:IEA

[DMNR12] Pablo D’Este, Surya Mahdi, Andy Neely, and Francesco Rentocchini. Inventors and entrepreneurs in academia: What types of skills and experience matter? Technovation, 32(5):293–303, May 2012. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497211001866.

Darroch:2005:CVR

[DMP05] Jenny Darroch, Morgan P. Miles, and Chris W. Paul. Corporate venturing and the rent cycle. Technovation, 25(12): 1437–1442, December 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497204001853.

DalMas:2023:CDT

[DMRS23] Francesca Dal Mas, Maurizio Massaro, Pierluigi Rippa, and Giustina Secundo. The challenges of digital transformation in healthcare: an interdisciplinary literature review, framework, and future research agenda. Technovation, 123(??): ??, May 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223000275.

Dabic:2021:YEO

[DMV^{+}21] Marina Dabić, Giacomo Marzi, Bozidar Vlacić, Tugrul U. Daim, and Wim Vanhaverbeke. 40 years of excellence: an overview of Technovation and a roadmap for future research. Technovation, 106(??):??, August 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000845.

deMelo:1993:OIC

[dMvAdMSPB93] M. Angela Campelo de Melo, T. Diana L. v. A. de Macedo-Soares, and Paulo Jacinto Pastor-Braga. Options for increasing competitiveness, quality and productivity: the case of the Brazilian computer industry. Technovation, 13(6):367–382, September 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497293900779.

Delmar:2013:URA

[DMW13] Frédéric Delmar, Alexander McKelvie, and Karl Wennberg. Untangling the relationships among growth, profitability and survival in new firms. Technovation, 33(8–9):276– 291, August/September 2013. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497213000199.

Doctor:2001:MUR

[DNP01] R. N. Doctor, D. P. Newton, and A. Pearson. Managing uncertainty in research and development. Technovation, 21(2):79–90, February 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497200000262.

Dyker:2003:EWN

[DNS^{+}03] D. Dyker, A. Nagy, P. Stanovnik, J. Turk, H. Usenik, and P. Vince. ‘East’–‘West’ networks and their alignment: industrial networks in Hungary and Slovenia. Technovation, 23(7):603–616, July 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497202000093.

DeToni:1998:IPD

[DNT98] Alberto De Toni, Guido Nassimbeni, and Stefano Tonchia. Innovation in product development within the electronics industry. Technovation, 19(2):71–80, December 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000984.

Dooley:1999:DSS

[DO99] Lawrence Dooley and David O’Sullivan. Decision support system for the management of systems change. Technovation, 19(8):483–493, August 1999. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0166497299000231.

Dooley:2003:DSI

[DO03] L. Dooley and D. O’Sullivan. Developing a software infrastructure to support systemic innovation through effective management. Technovation, 23(8):689–704, August 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000202.

Dodgson:1987:BRMa

[Dod87a] M. J. Dodgson. Book review: Management of hightechnology research and development: by John H. Dumbleton, Elsevier, Amsterdam, 1986, 399 pages. Technovation, 6(4):306–307, September 1987. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497287900812.

Dodgson:1987:BRMc

[Dod87b] Mark Dodgson. Book review: Managing technology: Competing through new ventures, innovation, and corporate research: Frederick Betz, Prentice-Hall, Englewood Cliffs, 1987, 268 pp., £20.35 (paperback). Technovation, 7(1):81, December 1987. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497287900502.

Dodgson:1988:BRM

[Dod88] Mark Dodgson. Book review: The management of strategic change: By Andrew Pettigrew (ed.), Basil Blackwell, Oxford, 1987. 370 pages, hardback £29.95. Technovation, 7(4):420–421, October 1988. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497288900120.

deOliveira:2012:RMD

[dOK12] Antonio Carlos de Oliveira and Paulo Carlos Kaminski. A reference model to determine the degree of maturity in the product development process of industrial SMEs. Technovation, 32(12):671–680, December 2012. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497212000934.

Dowlatshahi:1993:NAPa

[Dow93a] Shad Dowlatshahi. A novel approach to product design and development in a concurrent engineering environment. Technovation, 13(3):161–176, April 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390038W.

Dowlatshahi:1993:NAPb

[Dow93b] Shad Dowlatshahi. A novel approach to product design and development in a concurrent engineering environment. Technovation, 13(3):180–181, April 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390042T.

Denicolai:2023:ISD

[DP23] Stefano Denicolai and Pietro Previtali. Innovation strategy and digital transformation execution in healthcare: the role of the general manager. Technovation, 121(??):??, March 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649722200102X.

Dimos:2022:RER

[DPH^{+}22] Christos Dimos, Geoff Pugh, Mehtap Hisarciklilar, Ema Talam, and Ian Jackson. The relative effectiveness of R&D tax credits and R&D subsidies: a comparative meta-regression analysis. Technovation, 115(??):??, July 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497221002315.

Demaid:2006:KAC

[DQ06] Adrian Demaid and Paul Quintas. Knowledge across cultures in the construction industry: sustainability, innovation and design. Technovation, 26(5–6):603–610, May/ June 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649720500091X.

Dodgson:1988:SFP

[DR88] Mark Dodgson and Roy Rothwell. Small firm policy in the U.K. Technovation, 7(3):231–247, July 1988. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900223.

Drejer:1999:CDT

[DR99] Anders Drejer and Jens Ove Riis. Competence development and technology: How learning and technology can be meaningfully integrated. Technovation, 19(10):631–644, October 1999. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497299000644.

Dew:2007:MWG

[DR07] Nicholas Dew and Stuart Read. The more we get together: Coordinating network externality product introduction in the RFID industry. Technovation, 27(10):569–581, October 2007. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497207000028.

Das:2024:THF

[DR24] Arindam Das and Subhasis Ray. Techno-humane futures in the global south: lessons from Professor Shonku. Technovation, 132(??):??, April 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497224000373.

Drejer:1997:DMT

[Dre97] Anders Drejer. The discipline of management of technology, based on considerations related to technology. Technovation, 17(5):253–265, May 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497296001071.

Drejer:2001:HCW

[Dre01] Anders Drejer. How can we define and understand competencies and their development? Technovation, 21(3):135– 146, March 2001. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497200000316.

Drejer:2002:TMC

[Dre02] A. Drejer. Towards a model for contingency of Management of Technology. Technovation, 22(6):363–370, June 2002. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000293.

Drejer:2004:BRS

[Dre04] Ina Drejer. Book review: Service Innovation. Organizational Responses to Technological Opportunities & Market Imperatives: Joe Tidd and Frank M. Hull (Eds.); Imperial College Press, London, September 2003; 437 pages, ISBN 1-86094-367-5, US$65/£48. Technovation, 24(11):922–923, November 2004. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497204001257.

Drejer:2006:BRK

[Dre06] Ina Drejer. Book review: Knowledge Spillovers and Knowledge Management: Charlie Karlsson, Per Flensburg, Sven-Åke Hörte (Eds.). Edward Elgar, Cheltenham (2004). Technovation, 26(5–6):621–622, May/June 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000957.

deRon:1995:DRE

[dRP95] Ad de Ron and Kiril Penev. Disassembly and recycling of electronic consumer products: an overview. Technovation, 15(6):363–374, August 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729596597M.

Dolfsma:2013:GPT

[DS13] Wilfred Dolfsma and DongBack Seo. Government policy and technological innovation — a suggested typology. Technovation, 33(6–7):173–179, June/July 2013. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213000473.

Doloreux:2023:DLM

[DS23] David Doloreux and Richard Shearmur. Does location matter? STI and DUI innovation modes in different geographic settings. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222001560.

DelGiudice:2023:BSI

[DSPS23] Manlio Del Giudice, Veronica Scuotto, Armando Papa, and Sanjay Kumar Singh. The ‘bright’ side of innovation management for international new ventures. Technovation, 125 (??):??, July 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223001001.

Dubelaar:2005:BIC

[DSS05] Chris Dubelaar, Amrik Sohal, and Vedrana Savic. Benefits, impediments and critical success factors in B2C E-business adoption. Technovation, 25(11):1251–1262, November 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001403.

Silvestre:2009:GPI

[dSSD09] Bruno dos Santos Silvestre and Paulo Roberto Tavares Dalcol. Geographical proximity and innovation: Evidences from the Campos Basin oil & gas industrial agglomeration — Brazil. Technovation, 29(8):546–561, August 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209000030.

Durao:2005:VRE

[DSVM05] D. Durão, M. Sarmento, V. Varela, and L. Maltez. Virtual and real-estate science and technology parks: a case study of Taguspark. Technovation, 25(3):237–244, March 2005.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720300110X.

DeClercq:2013:SNL

[DTD13] Dirk De Clercq, Narongsak (Tek) Thongpapanl, and Dimo Dimov. Shedding new light on the relationship between contextual ambidexterity and firm performance: an investigation of internal contingencies. Technovation, 33(4–5):119–132, April/May 2013. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497212001447.

Dhir:2023:DSD

[DTI^{+}23] Amandeep Dhir, Shalini Talwar, Nazrul Islam, Rasha Alghafes, and Saeed Badghish. Different strokes for different folks: Comparative analysis of 3D printing in large, medium and small firms. Technovation, 125(??):??, July 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001037.

During:1986:PMM

[Dur86] W. E. During. Project management and management of innovation in small industrial firms. Technovation, 4(4):269– 278, August 1986. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497286900301.

Dutrenit:2007:BRG

[Dut07] Gabriela Dutrénit. Book review: Globalization and Technology. Narula, R. Polity Press Cambridge. 243 pp., Hardback £55/$77.95, Paperback £16.99/$29.95, ISBN 0-7456-2456-1. Technovation, 27(10):640–641, October 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000363.

deVries:1994:DPD

[dV94] Marc J. de Vries. Design process dynamics in an experiencebased context: a design methodological analysis of the Brabantia corkscrew development. Technovation, 14(7):

437–448, September 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497294900027.

DeKeyser:2023:PPR

[DV23] Bart De Keyser and Koen Vandenbempt. Processes of practice in the realm of theory: Unveiling the dynamics of academic intrapreneurship. Technovation, 126(??):??, August 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223001220.

Delgado-Verde:2016:ICR

[DVdCAS16] Miriam Delgado-Verde, Gregorio Martín de Castro, and Javier Amores-Salvadó. Intellectual capital and radical innovation: Exploring the quadratic effects in technologybased manufacturing firms. Technovation, 54(??):35–47, August 2016. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497216300050.

deVisser:2010:SAN

[dVdWNF^{+}10] Matthias de Visser, Petra de Weerd-Nederhof, Dries Faems, Michael Song, Bart van Looy, and Klaasjan Visscher. Structural ambidexterity in NPD processes: a firm-level assessment of the impact of differentiated structures on innovation performance. Technovation, 30(5–6):291–299, May/ June 2010. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649720900145X.

Gomes:2023:CEI

[dVGCOF^{+}23] Leonardo Augusto de Vasconcelos Gomes, Danika A. Castillo-Ospina, Ana Lucia Figueiredo Facin, Camila dos Santos Ferreira, and Aldo Roberto Ometto. Circular ecosystem innovation portfolio management. Technovation, 124 (??):??, June 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223000561.

Duong:2022:EKS

[DVHL22] Phuong-Anh Nguyen Duong, Wim Voordeckers, Jolien Huybrechts, and Frank Lambrechts. On external knowledge sources and innovation performance: Family versus non-family firms. Technovation, 114(??):??, June 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497221002297.

Diez-Vial:2016:HKL

[DVMS16] Isabel Díez-Vial and Ángeles Montoro-Sánchez. How knowledge links with universities may foster innovation: the case of a science park. Technovation, 50–51(??):41–52, April/ May 2016. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497215000620.

Deligianni:2019:NLE

[DVSL19] Ioanna Deligianni, Irini Voudouris, Yiannis Spanos, and Spyros Lioukas. Non-linear effects of technological competence on product innovation in new technology-based firms: Resource orchestration and the role of the entrepreneur’s political competence and prior start-up experience. Technovation, 88(??):Article 102076, December 2019. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497217306545.

deVries:2016:ICR

[dVV16] Henk J. de Vries and W. Pieter Verhagen. Impact of changes in regulatory performance standards on innovation: a case of energy performance standards for newlybuilt houses. Technovation, 48–49(??):56–68, February/ March 2016. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497216000092.

Dean:2022:RCV

[DZX22] Tereza Dean, Haisu Zhang, and Yazhen Xiao. The role of complexity in the Valley of Death and radical innovation performance. Technovation, 109(??):??, January 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497218306187.

ElMaalouf:2023:IIO

[EB23] Nicole El Maalouf and Hanna Bahemia. The implementation of inbound open innovation at the firm level: a dynamic capability perspective. Technovation, 122(??):??, April 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497222002103.

Edris:2024:TCR

[EBG24] Sarah Edris, René Belderbos, and Victor Gilsing. Types of common R&D partners and knowledge leakage to rivals: the role of IP litigation reputation. Technovation, 131(??): ??, March 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497224000051.

Erikson:2022:VGD

[ECB22] Truls Erikson, Chris Coleridge, and Ekaterina Bjornali. Venture governance and its dynamics: Intraboard relationships and CEO duality. Technovation, 115(??):??, July 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000876.

Etemad:1987:PPL

[ED87] Hamid Etemad and Louise Séguin Dulude. Patenting patterns in 25 large multinational enterprises. Technovation, 7(1):1–15, December 1987. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497287900435.

Edwards:2005:UIS

[EDM05] Tim Edwards, Rick Delbridge, and Max Munday. Understanding innovation in small and medium-sized enterprises: a process manifest. Technovation, 25(10):1119–1127, October 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204000732.

Edwards:2004:BRB

[Edw04] Rod Edwards. Book review: The Business of Systems Integration: Andrea Prencipe, Andrew Davies and Mike Hobday (Eds.); Oxford University Press, 374 pages, ISBN 0-19-926322-1, £50. Technovation, 24(6):514–515, June 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204000124.

Eekels:1984:IPC

[Eek84] J. Eekels. Innovate or perish: a cautionary tale. Technovation, 2(3):149–167, June 1984. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497284900014.

Erzurumlu:2022:NIS

[EEY22] S. Sinan Erzurumlu, Yaman O. Erzurumlu, and YongKi Yoon. National innovation systems and dynamic impact of institutional structures on national innovation capability: a configurational approach with the OKID method. Technovation, 114(??):??, June 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222000992.

Essegbey:2011:CSI

[EF11] George Owusu Essegbey and Godfred Kwasi Frempong. Creating space for innovation — the case of mobile telephony in MSEs in Ghana. Technovation, 31(12):679–688, December 2011. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497211001222.

Essen:2023:ENP

[EFC23] Anna Essen, Johan Frishammar, and Javier Cenamor. Entering non-platformized sectors: the co-evolution of legitimacy debates and platform business models in digital health care. Technovation, 121(??):??, March 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222001444.

Efrat:2014:DII

[Efr14] Kalanit Efrat. The direct and indirect impact of culture on innovation. Technovation, 34(1):12–20, January 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213001028.

Entrialgo:2000:LES

[EFV00] Montserrat Entrialgo, Esteban Fernández, and Camilo J. Vázquez. Linking entrepreneurship and strategic management: evidence from Spanish SMEs. Technovation, 20 (8):427–436, August 2000. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497299001686.

Escanciano:2002:LFT

[EFV02] Carmen Escanciano, Esteban Fernández, and Camilo Vázquez. Linking the firm’s technological status and ISO 9000 certification: results of an empirical research. Technovation, 22(8):509–515, August 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000529.

Erken:2005:RRD

[EG05] Hugo Erken and Victor Gilsing. Relocation of R&D — a Dutch perspective. Technovation, 25(10):1079–1092, October 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204000549.

Egbetokun:2015:MMN

[Egb15] Abiodun A. Egbetokun. The more the merrier? Network portfolio size and innovation performance in Nigerian firms. Technovation, 43–44(??):17–28, September/October 2015. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497215000450.

Easingwood:2002:LRL

[EH02] C. Easingwood and S. Harrington. Launching and relaunching high technology products. Technovation, 22(11): 657–666, November 2002. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497202000974.

Enkel:2014:PDC

[EH14] Ellen Enkel and Sebastian Heil. Preparing for distant collaboration: Antecedents to potential absorptive capacity in cross-industry innovation. Technovation, 34(4):242–260, April 2014. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497214000170.

Enkel:2017:EEI

[EHHW17] Ellen Enkel, Sebastian Heil, Monika Hengstler, and Henning Wirth. Exploratory and exploitative innovation: To what extent do the dimensions of individual level absorptive capacity contribute? Technovation, 60–61(??):29–38, February 2017. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497216302346.

Ehrnberg:1995:DMT

[Ehr95] Ellinor Ehrnberg. On the definition and measurement of technological discontinuities. Technovation, 15(7):437–452, September 1995. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729596593I.

Elrashidy:2024:DRC

[EHSB24] Zeinab Elrashidy, Roszaini Haniffa, Mohamed Sherif, and Sarra Baroudi. Determinants of reward crowdfunding success: Evidence from Covid-19 pandemic. Technovation, 132 (??):??, April 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649722400035X.

Evangelista:2001:MRD

[EIMS01] Rinaldo Evangelista, Simona Iammarino, Valeria Mastrostefano, and Alberto Silvani. Measuring the regional dimension of innovation. Lessons from the Italian Innovation Survey. Technovation, 21(11):733–745, November 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000845.

Engwall:2021:ENB

[EKK^{+}21] Mats Engwall, Matti Kaulio, Emrah Karakaya, Maxim Miterev, and Daniel Berlin. Experimental networks for business model innovation: a way for incumbents to navigate sustainability transitions? Technovation, 108(??):??, December 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497221001115.

Edvardsson:2012:CIW

[EKMS12] Bo Edvardsson, Per Kristensson, Peter Magnusson, and Erik Sundström. Customer integration within service development — a review of methods and an analysis of insitu and exsitu contributions. Technovation, 32(7–8):419–429, July/August 2012. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497211000605.

Elahi:2022:EWE

[EKT^{+}22] Ehsan Elahi, Zainab Khalid, Muhammad Zubair Tauni, Hongxia Zhang, and Xing Lirong. Extreme weather events risk to crop-production and the adaptation of innovative management strategies to mitigate the risk: a retrospective survey of rural Punjab, Pakistan. Technovation, 117 (??):??, September 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497221000365.

Engzell:2024:AIN

[EKY24] Jeanette Engzell, Solmaz Filiz Karabag, and Anna Yström. Academic intrapreneurs navigating multiple institutional logics: an integrative framework for understanding and supporting intrapreneurship in universities. Technovation, 129(??):??, January 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223002031.

Eto:1993:FPT

[EL93] Hajime Eto and Jung Hee Lee. Foreign patenting and trade with regard to competitiveness. Technovation, 13(4):221– 233, May 1993. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729390019R.

Eum:2019:RPF

[EL19] Wonsub Eum and Jeong-Dong Lee. Role of production in fostering innovation. Technovation, 84–85(??):1–10, June/ July 2019. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497217307708.

Ellin:1981:TIN

[Ell81] Everett Ellin. Tracking innovation: the need for reliable measures of its presence and effects. Technovation, 1(1):69– 80, February 1981. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497281900080.

Elshout:1995:SRT

[Els95] Jan Elshout. Significance of reverse transfer of technology and experiences with its promotion in industrialised countries. Technovation, 15(7):423–435, September 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729596592H.

Evans:1985:NAE

[EM85] James R. Evans and Samuel J. Mantel, Jr. A new approach to the evaluation of process innovations. Technovation, 3 (4):263–281, October 1985. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497285900240.

Esposito:2002:ITP

[EM02] Emilio Esposito and Michele Mastroianni. Information technology and personal computers: the relational life cycle. Technovation, 22(1):41–50, January 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000894.

Eriksson:2007:DCU

[EN07] Kent Eriksson and Daniel Nilsson. Determinants of the continued use of self-service technology: the case of Internet banking. Technovation, 27(4):159–167, April 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497206001088.

Ener:2022:HDC

[Ene22] Hakan Ener. How does CEO technical expertise influence licensing-out at technology ventures? Technovation, 114(??):??, June 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222000487.

Eng:2006:MSC

[Eng06] Teck-Yong Eng. Mobile supply chain management: Challenges for implementation. Technovation, 26(5–6):682–686, May/June 2006. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497205001070.

Eriksson:2000:TSA

[Eri00] Sören Eriksson. Technology spill-over from the aircraft industry: the case of Volvo Aero. Technovation, 20(12):653– 664, December 2000. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497200000092.

Ernst:1995:PSG

[Ern95] H. Ernst. Patenting strategies in the German mechanical engineering industry and their relationship to company performance. Technovation, 15(4):225–240, May 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729596605S.

Erikson:2005:TAO

[ES05] Truls Erikson and Roger Sørheim. ‘Technology angels’ and other informal investors. Technovation, 25(5):489–496, May 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203001731.

Enkel:2021:ECV

[ES21] Ellen Enkel and Veronika Sagmeister. External corporate venturing modes as new way to develop dynamic capabilities. Technovation, ??(??):??, ???? 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497218307983.

Ernst:1998:RMI

[ET98] Holger Ernst and Thorsten Teichert. The R&D marketing interface and single informant bias in NPD research: an illustration of a benchmarking case study. Technovation, 18 (12):721–739, October 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497298000698.

Efstathiades:2002:SPT

[ETA02] Andreas Efstathiades, Savvas Tassou, and Antonios Antoniou. Strategic planning, transfer and implementation of Advanced Manufacturing Technologies (AMT). Development of an integrated process plan. Technovation, 22(4): 201–212, April 2002. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497201000244.

Efstathiades:1998:SCI

[ETAO98] Andreas Efstathiades, Savvas A. Tassou, Antonios Antoniou, and George Oxinos. Strategic considerations in the introduction of advanced manufacturing technologies in the Cypriot industry. Technovation, 19(2):105–115, December 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497298000923.

Elola:1996:NME

[ETM96] Luis Navarro Elola, Ana Clara Pastor Tejedor, and Leonor González Menorca. New methods of evaluating physical demand at work areas. Technovation, 16(10):595– 599, October 1996. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497297828982.

Efstathiades:2000:AMT

[ETOA00] Andreas Efstathiades, Savvas A. Tassou, George Oxinos, and Antonios Antoniou. Advanced manufacturing technology transfer and implementation in developing countries: the case of the Cypriot manufacturing industry. Technovation, 20(2):93–102, February 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001005.

Evangelista:2002:BRM

[Eva02] R. Evangelista. Book review: The microdynamics of technological change: Cristiano Antonelli; Routledge, London, 1999. Technovation, 22(2):135–136, February 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000220.

Erdilek:1997:TOF

[EW97] Asim Erdilek and Milton A. Wolf. Technology origins of foreign-owned firms in Ohio. Technovation, 17(2):63– 72, 101–102, February 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497296000806.

Eng:2006:GMR

[EW06] Teck-Yong Eng and Veronica Wong. Governance mechanisms and relationship productivity in vertical coordination for new product development. Technovation, 26(7): 761–769, July 2006. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497204002093.

Ellwood:2022:CVD

[EWE22] Paul Ellwood, Ceri Williams, and John Egan. Crossing the valley of death: Five underlying innovation processes. Technovation, 109(??):??, January 2022. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497218306023.

Ellonen:2009:LDC

[EWJ09] Hanna-Kaisa Ellonen, Patrik Wikström, and Ari Jantunen. Linking dynamic-capability portfolios and innovation outcomes. Technovation, 29(11):753–762, November 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209000601.

Fabayo:1996:TDA

[Fab96] J. A. Fabayo. Technological dependence in Africa: its nature, causes, consequences and policy derivatives. Technovation, 16(7):357–370, July 1996. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497296000132.

Fahmy:2005:CRS

[Fah05] Yasser M. Fahmy. Catalysis role for sustainable industrial development in Egypt with prospective. Technovation, 25(6):645–655, June 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497203001962.

Fathian:2008:RAN

[FAH08] Mohammad Fathian, Peyman Akhavan, and Maryam Hoorali. E-readiness assessment of non-profit ICT SMEs in a developing country: the case of Iran. Technovation, 28(9):578–590, September 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208000175.

Falk:2005:ILF

[Fal05] Martin Falk. ICT-linked firm reorganisation and productivity gains. Technovation, 25(11):1229–1250, November 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001105.

Fan:2006:CTD

[Fan06] Peilei Fan. Catching up through developing innovation capability: evidence from China’s telecom-equipment industry. Technovation, 26(3):359–368, March 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720400183X.

Faridian:2023:LOI

[Far23] Parisa Haim Faridian. Leading open innovation: the role of strategic entrepreneurial leadership in orchestration of value creation and capture in GitHub open source communities. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000931.

Foxall:1991:CSP

[FB91] Gordon R. Foxall and Seema Bhate. Cognitive style, personal involvement and situation as determinants of computer use. Technovation, 11(3):183–199, April 1991. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729190033Z.

Foxall:1993:CSU

[FB93] Gordon R. Foxall and Seema Bhate. Cognitive style and use-innovativeness for applications software in home computing: implications for new product strategy. Technovation, 13(5):311–323, July 1993. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729390004F.

Francis:2005:TII

[FB05] Dave Francis and John Bessant. Targeting innovation and implications for capability development. Technovation, 25 (3):171–183, March 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497204000525.

Ferneley:2006:UBI

[FB06] Elaine Ferneley and Frances Bell. Using bricolage to integrate business and information technology innovation in SMEs. Technovation, 26(2):232–241, February 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000672.

Fauchart:2022:DIY

[FBBBM22] Emmanuelle Fauchart, Maya Bacache-Beauvallet, Marc Bourreau, and François Moreau. Do-it-yourself or do-it-together: How digital technologies affect creating alone or with others? Technovation, 112(??):??, April 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001930.

Fedorenko:2023:TSI

[FBE23] Ivan Fedorenko, Pierre Berthon, and Linda Edelman. Top secret: Integrating 20 years of research on secrecy. Technovation, 123(??):??, May 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000020.

Fontes:1995:NTB

[FC95] Margarida Fontes and Rod Coombs. New technologybased firms and technology acquisition in Portugal: Firms’ adaptive responses to a less favourable environment. Technovation, 15(8):497–510, October 1995. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729596588K.

Fennelly:2006:VCM

[FC06] Dermot Fennelly and Kathryn Cormican. Value chain migration from production to product centred operations: an analysis of the Irish medical device industry. Technovation, 26(1):86–94, January 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001117.

Fernandez:2021:WDM

[FCCE21] Anne-Sophie Fernandez, Paul Chiambaretto, Mathieu Chauvet, and Juliane Engsig. Why do MNEs both make and coopete for innovation? Technovation, 106(??):??, August 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221000948.

Fenwick:2001:TCM

[FCH01] P. Fenwick and M. Cordey-Hayes. Technology change and management strategy in the Benefits Agency. Technovation, 21(3):167–174, March 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497200000328.

Ferreira:2021:DCC

[FCM21] Jorge Ferreira, Arnaldo Coelho, and Luiz Moutinho. Dynamic capabilities, creativity and innovation capability and their impact on competitive advantage and firm performance: the moderating role of entrepreneurial orientation. Technovation, ??(??):??, ???? 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497217301335.

Figueiredo:2021:IFL

[FCS21] Paulo N. Figueiredo, Bernardo P. Cabral, and Felipe Q. Silva. Intricacies of firm-level innovation performance: an empirical analysis of latecomer process industries. Technovation, ??(??):??, ???? 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000833.

Fritzsche:2021:TBE

[FD21] Albrecht Fritzsche and Konrad Dürrbeck. Technology before engineering: How James Bond films mediate between fiction and reality in the portrayal of innovation. Technovation, ??(??):??, ???? 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497217301220.

Freel:2009:MNC

[FdJ09] Mark Freel and Jeroen P. J. de Jong. Market novelty, competence-seeking and innovation networking. Technovation, 29(12):873–884, December 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0166497209001072.

Freije:2022:RSC

[FdlCU22] Inmaculada Freije, Alberto de la Calle, and José V. Ugarte. Role of supply chain integration in the product innovation capability of servitized manufacturing companies. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497220300882.

Franzo:2023:DIC

[FDN^{+}23] Simone Franzò, Nicola Doppio, Angelo Natalicchio, Federico Frattini, and Luca Mion. Designing innovation contests to support external knowledge search in small and medium-sized enterprises. Technovation, 121(??):??, March 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497222002358.

Fernandez:2010:RAF

[FDVB10] Pilar Fernández, M. Luisa Del Río, José Varela, and Belén Bande. Relationships among functional units and new product performance: the moderating effect of technological turbulence. Technovation, 30(5–6):310–321, May/June 2010. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209001394.

Fels:1989:XF

[Fel89] Xavier Fels. Xavier Fels. Technovation, 9(2–3):293–294, June 1989. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497289900722.

Felsenstein:1994:URS

[Fel94] Daniel Felsenstein. University-related science parks — ‘seedbeds’ or ‘enclaves’ of innovation? Technovation, 14(2): 93–110, March 1994. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729490099X. See [LQ94].

Fellnhofer:2022:EAT

[Fel22]  Katharina Fellnhofer.  Entrepreneurial alertness to-

ward responsible research and innovation: Digital technology makes the psychological heart of entrepreneurship pound. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497221001656.

Fellnhofer:2023:PHA

[Fel23] Katharina Fellnhofer. Positivity and higher alertness levels facilitate discovery: Longitudinal sentiment analysis of emotions on Twitter. Technovation, 122(??):??, April 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222002176.

Frishammar:2015:DSK

[FEP15] Johan Frishammar, Kristian Ericsson, and Pankaj C. Patel. The dark side of knowledge transfer: Exploring knowledge leakage in joint R&D projects. Technovation, 41–42(??):75–88, July/August 2015. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497215000024.

Foxall:1992:EMT

[FF92] Gordon R. Foxall and John R. Fawn. An evolutionary model of technological innovation as a strategic management process. Technovation, 12(3):191–202, April 1992. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729290035G.

Foo:2000:STT

[FF00] Check-Teck Foo and Check-Tong Foo. Socialization of technopreneurism: towards symbiosis in corporate innovation and technology strategy. Technovation, 20(10):551– 562, October 2000. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497299001716.

Foxall:1992:IGP

[FFJ92] Gordon R. Foxall, John R. Fawn, and Brian R. Johnston. Innovation in grand prix motor racing. II. Extension of the population ecology model. Technovation, 12(1):1–14, February 1992. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729290029H.

Frietsch:2006:TNM

[FG06] Rainer Frietsch and Hariolf Grupp. There’s a new man in town: the paradigm shift in optical technology. Technovation, 26(1):13–29, January 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001130.

Foglia:2024:DTC

[FGB^{+}24] Emanuela Foglia, Elisabetta Garagiola, Daniele Bellavia, Federica Rossetto, and Francesca Baglio. Digital technology and COVID-19 pandemic: Feasibility and acceptance of an innovative telemedicine platform. Technovation, 130(??): ??, February 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223002523.

Fu:2023:PBB

[FGOX23] Xiaolan Fu, Pervez Ghauri, Nwamaka Ogbonna, and Xiaoqiang Xing. Platform-based business model and entrepreneurs from Base of the Pyramid. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497221002327.

Ferretti:2022:PND

[FGPP22] Marco Ferretti, Massimiliano Guerini, Eva Panetti, and Adele Parmentola. The partner next door? The effect of micro-geographical proximity on intra-cluster interorganizational relationships. Technovation, 111(??):??, March 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221001711.

Flores:2024:EUI

[FGPV24] Maria Cecilia Flores, Rosa Grimaldi, Silvia Poli, and Elisa Villani. Entrepreneurial universities and intrapreneurship: a process model on the emergence of an intrapreneurial university. Technovation, 129(??):??, January 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223002171.

Furnsinn:2007:AEF

[FGS07] Stefan Fürnsinn, Markus Günther, and Christian Stummer. Adopting energy flow charts for the economic analysis of process innovations. Technovation, 27(11):693–703, November 2007. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497207000478.

Fabra:2023:ASD

[FGTB23] Irene Talavera Fabra, Abby Ghobadian, Ciro Troise, and Stefano Bresciani. Antecedents of successful diffusion of breakthrough innovations past the formative phase: Perceptions of innovation-engaged practitioners. Technovation, 127(??):??, September 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001621.

Fusfeld:1987:CIR

[FH87] Herbert I. Fusfeld and Carmela S. Haklisch. Collaborative industrial research in the U.S. Technovation, 5(4):305–315, February 1987. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497287900691.

Fetrati:2022:HMC

[FHA22] Mahdieh A. Fetrati, David Hansen, and Payman Akhavan. How to manage creativity in organizations: Connecting the literature on organizational creativity through bibliometric research. Technovation, 115(??):??, July 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000207.

Featherston:2016:MCI

[FHBDO16] Charles R. Featherston, Jae-Yun Ho, Laure Brévignon-Dodin, and Eoin O’Sullivan. Mediating and catalysing innovation: a framework for anticipating the standardisation needs of emerging technologies. Technovation, 48–49(??):25–40, February/March 2016. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497215000772.

Figueiredo:2002:LPF

[Fig02] P. N. Figueiredo. Learning processes features and technological capability-accumulation: explaining inter-firm differences. Technovation, 22(11):685–698, November 2002. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000682.

Fischmeister:1989:MII

[Fis89] Hellmut Fischmeister. Materials-inspired innovation in a world of routine design. Technovation, 9(4):309–319, July 1989. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497287900034.

Fitzgerald:2000:BRB

[Fit00] P. Fitzgerald. Book review: Best Practice Creativity: by Peter Cook; Gower Publishing, Aldershot, 1998, ISBN 0-566-08027-3 Creativity and Problem-Solving at Work by Tudor Rickards; Gower Publishing, 1998, ISBN 0-566-07961-5. Technovation, 20(7):401–402, July 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001777.

Foxall:1987:SIS

[FJ87a] Gordon Foxall and Brian Johnston. Strategies of innovation in smaller firms. Technovation, 6(3):169–187, August 1987. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497287900216.

Foxall:1987:SUI

[FJ87b] Gordon Foxall and Brian Johnston. Strategies of userinitiated product innovation. Technovation, 6(2):77–102, June 1987. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497287900137.

Foxall:1991:IGP

[FJ91] Gordon R. Foxall and Brian R. Johnston. Innovation in Grand Prix motor racing: the evolution of technology, organization and strategy. Technovation, 11(7):387–402, November 1991. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497291900205.

Fonseca:2012:ABI

[FJ12] Sergio Azevedo Fonseca and Charbel José Chiappetta Jabbour. Assessment of business incubators’ green performance: a framework and its application to Brazilian cases. Technovation, 32(2):122–132, February 2012. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211001507.

Fernandez:1996:GSR

[FJV96] Esteban Fernández, Beatriz Junquera, and Camilo J. Vazquez. Government support for R&D: the Spanish case. Technovation, 16(2):59–65, 92, February 1996. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497295000208.

Fisher:1995:ITR

[FKT95] J. Fisher, C. Kirk, and D. Taylor. The implications of TQM for R&D strategy in New Zealand firms. Technovation, 15(1):1–9, February 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729596607U.

Funk:2015:OSV

[FL15] Jeffrey L. Funk and Jianxi Luo. Open standards, vertical disintegration and entrepreneurial opportunities: How vertically-specialized firms entered the U.S. semiconductor industry. Technovation, 45–46(??):52–62, November/ December 2015. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497215000462.

Fan:2017:CIS

[FLC17] Di Fan, Yi Li, and Liang Chen. Configuring innovative societies: the crossvergent role of cultural and institutional varieties. Technovation, 66–67(??):43–56, August 2017. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497216301936.

Fang:2002:RFR

[FLH^{+}02] Shih-Chieh Fang, Julia L. Lin, Luke Y. C. Hsiao, Chung-Ming Huang, and Shyh-Rong Fang. The relationship of foreign R&D units in Taiwan and the Taiwanese knowledgeflow system. Technovation, 22(6):371–383, June 2002. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720100027X.

Fragkaki:2006:ODC

[FLKF^{+}06] A. G. Fragkaki, I.-P. Leontiou, N. Kioukia-Fougia, M. Tsivou, M.-H. E. Spyridaki, and C. G. Georgakopoulos. Organization of the doping control laboratory in the Athens 2004 Olympic Games: a case study. Technovation, 26(10):1162–1169, October 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205001197.

Figenschou:2024:OIP

[FLYTB24] Therese Figenschou, Jason Li-Ying, Anne Tanner, and Marcel Bogers. Open innovation in the public sector: a literature review on actors and boundaries. Technovation, 131 (??):??, March 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223002511.

Fuller:2007:VPE

[FM07] Johann Füller and Kurt Matzler. Virtual product experience and customer participation — a chance for customercentred, really new products. Technovation, 27(6–7):378– 387, June/July 2007. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497206001039.

Freisinger:2024:WFW

[FM24] Elena Freisinger and Ian P. McCarthy. What fails and when? A process view of innovation failure. Technovation, 133(??):??, May 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497224000452.

Fabiano:2021:VDK

[FMF21] Gianluca Fabiano, Andrea Marcellusi, and Giampiero Favato. R versus D, from knowledge creation to value appropriation: Ownership of patents filed by European biotechnology founders. Technovation, 108(??):??, December 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001097.

Ferreras-Mendez:2016:RBK

[FMFMA16] José Luis Ferreras-Méndez, Anabel Fernández-Mesa, and Joaquín Alegre. The relationship between knowledge search strategies and absorptive capacity: a deeper look. Technovation, 54(??):48–61, August 2016. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497216300153.

Ferreras-Mendez:2021:EON

[FMOPSVA21] José Luis Ferreras-Méndez, Julia Olmos-Peñuela, Andrés Salas-Vallina, and Joaquín Alegre. Entrepreneurial orientation and new product development performance in SMEs: the mediating role of business model innovation. Technovation, 108(??):??, December 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001061.

Fernandez:2000:TSA

[FMV00] Esteban Fernández, José M. Montes, and Camilo J. Vázquez. Typology and strategic analysis of intangible resources: a resource-based approach. Technovation, 20 (2):81–92, February 2000. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497299001157.

Fujimoto:2000:TFT

[FMvT00] Masayo Fujimoto, Kumiko Miyazaki, and Nick von Tunzelmann. Technological fusion and telemedicine in Japanese companies. Technovation, 20(4):169–187, April 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001261.

Faridian:2021:AAA

[FN21] Parisa Haim Faridian and Donald O. Neubaum. Ambidexterity in the age of asset sharing: Development of dynamic capabilities in open source ecosystems. Technovation, ?? (??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497218308186.

Frietsch:2014:PIM

[FNJV14] Rainer Frietsch, Peter Neuhäusler, Taehyun Jung, and Bart Van Looy. Patent indicators for macroeconomic growth — the value of patents estimated by export volume. Technovation, 34(9):546–558, September 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214000662.

Faborode:2003:ASO

[FOL^{+}03] M. O. Faborode, O. K. Owolarafe, A. A. Lasisi, S. A. Kasali, and K. S. Oguntuase. Assessment of seed-oil extraction technology in some selected states in Nigeria. Technovation, 23(6):545–553, June 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000056.

Fontes:2005:PTS

[Fon05] Margarida Fontes. The process of transformation of scientific and technological knowledge into economic value conducted by biotechnology spin-offs. Technovation, 25(4):339– 347, April 2005. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203001561.

Fortescue:1985:BRT

[For85] Stephen Fortescue. Book review: Technological change in the German Democratic Republic: Raymond Bentley. Westview, Boulder and London, 1984, xx + 296 pp. Technovation, 3(4):315–316, October 1985. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497285900276.

Forman:1990:BRN

[For90] Andrew Forman. Book review: New product development and testing: By W. Henry, M. Menasco and H. Takada, Lexington Books, Lexington, MA, 1989, £25.00, 302 pages, ISBN 0-669-1277-X. Technovation, 10(6):437–438, September 1990. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497290900066.

Forman:1991:BRN

[For91] Andrew Forman. Book review: New product development and testing: W. Henry, M. Menasco and H. Takada, Lexington Books, Lexington, MA, 1989, ISBN 0-669-1277-X, £25.00, 302 pages. Technovation, 11(2):126–127, March 1991. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497291900489.

Forbes:1999:TII

[For99] Naushad Forbes. Technology and Indian industry: what is liberalization changing? Technovation, 19(6–7):403–412, June 1999. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649729900053X.

Fernandez-Olmos:2017:HIE

[FORA17] Marta Fernández-Olmos and Marisa Ramírez-Alesón. How internal and external factors influence the dynamics of SME technology collaboration networks over time. Technovation, 64–65(??):16–27, June 2017. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497216301705.

Fredriksen:1997:VCF

[FOW97] Öystein Fredriksen, Christer Olofsson, and Clas Wahlbin. Are venture capitalists firefighters? A study of the influence and impact of venture capital firms. Technovation, 17(9):503–511, 531–532, September 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297000308.

Foxall:1986:CEC

[Fox86a] Gordon Foxall. A conceptual extension of the customeractive paradigm. Technovation, 4(1):17–27, January 1986. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497286900209.

Foxall:1986:MTE

[Fox86b] Gordon R. Foxall. Managers in transition: an empirical test of Kirton’s adaption-innovation theory and its implications for the mid-career MBA. Technovation, 4(3):219–232, June 1986. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497286900143.

Foxall:1995:CSC

[Fox95] Gordon R. Foxall. Cognitive styles of consumer initiators. Technovation, 15(5):269–288, June 1995. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729596600X.

Friedman:2008:HNK

[FP08] Robert S. Friedman and Laurence Prusak. On heuristics, narrative and knowledge management. Technovation, 28 (12):812–817, December 2008. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497208000837.

Franze:2024:SVC

[FPP24] Claudia Franzè, Emilio Paolucci, and Elena Pessot. Sustained value creation driven by digital connectivity: a multiple case study in the mechanical components industry. Technovation, 129(??):??, January 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223002298.

Fischer:2015:CPT

[FR15] Timo Fischer and Philipp Ringler. The coincidence of patent thickets — a comparative analysis. Technovation, 38(??):42–49, April 2015. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497214001722.

Francis:1999:BRP

[Fra99] D. L. Francis. Book review: Profit Patterns: A. J. Slywotzky, D. Morrison, T. Moser, K. A. Mundt, J. A. Quella; Wiley, Chichester, 1999. Technovation, 19(10):649–650, October 1999. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649729900098X.

Francis:2000:BRB

[Fra00a] Dave Francis. Book review: Building the Innovative Organization: J. Christiansen; MacMillan Press Ltd. Technovation, 20(10):587, October 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000468.

Francis:2000:BRA

[Fra00b] David Francis. Book review: Adversity Quotient: Turning Obstacles into Opportunities: P. G. Stoltz; John Wiley and Sons, New York, 1997. Technovation, 20(7):402, July 2000. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000109.

Francis:2001:BRA

[Fra01a] D. L. Francis. Book review: Action Learning — How the World’s Top Companies are Re-creating Their Leaders and Themselves: D. L. Dotlich and J. L. Noel; Jossey Bass Inc., San Francisco, 1999. Technovation, 21(2):133–134, February 2001. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497200000560.

Francis:2001:BRM

[Fra01b] D. L. Francis. Book review: Managing Know-Who Based Companies: a Multinetworked Approach to Knowledge and Innovation Management: Sigvald J Harryson; Edward Elgar, Cheltenham. Technovation, 21(6):395–396, June 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720000095X.

Francis:2001:BRB

[Fra01c] D. L. Francis. Book review: The Biology of Business: Clippinger III, J. H., Ed.; (1999). San Francisco, Cal., Jossey-Bass Inc. Technovation, 21(7):464–465, July 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720100013X.

Francis:2002:BRR

[Fra02a] D. Francis. Book review: Reframing business: when the map changes the landscape: Richard Normann; 2001, Price £17-99. ISBN 0-471-48557-8. Technovation, 22(11):731– 732, November 2002. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497202000792.

Francis:2002:BRE

[Fra02b] D. L. Francis. Book review: Entrepreneurship: a social science view: R. Swedberg Ed. (2000). Oxford Management Readers. Oxford, Oxford University Press. ISBN 0-19-829462-X (403 pages). Technovation, 22(3):199, March 2002. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201000797.

Francis:2002:BRN

[Fra02c] D. L. Francis. Book review: Necessary but Not Sufficient: Eliyahu M. Goldratt, Eli Schragenheim and Carol A. Ptak; The North River Press, 2000, pp. 231, Price £14.99, ISBN 0-88427-170-6. Technovation, 22(4):267–268, April 2002. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000839.

Francis:2002:BRP

[Fra02d] D. L. Francis. Book review: Practicing leadership: principles and applications: A. Shriberg, D. L. Shriberg et al. (2002). New York, John Wiley & Sons Inc. ISBN 0-471-39283-9 (246 pages) £35.95. Technovation, 22(3):200, March 2002. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201000803.

Francis:2003:BRE

[Fra03a] D. Francis. Book review: Essentials of international management: a cross-cultural perspective: David C. Thomas; Sage. ISBN 0-7619-2468-X 0-7619-2181-8 (paper). Technovation, 23(1):85–86, January 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000809.

Francis:2003:BRM

[Fra03b] D. L. Francis. Book review: Mapping Strategic Knowledge: Edited by Anne S. Huff and Mark Jenkins, Sage, 2002. ISBN 0-7619-6948-9 (£60). ISBN 0-7619-6949-7 (£21.99) Paperback. 320 pages. Technovation, 23(6):555–556, June 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000129.

Fernandes:2005:PKR

[FRA05] Kiran Jude Fernandes, Vinesh Raja, and Simon Austin. Portals as a knowledge repository and transfer tool-VIZCon case study. Technovation, 25(11):1281–1289, November 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204000033.

Freier:1986:PSB

[Fre86] S. Freier. Parks of science-based industries in Israel. Technovation, 4(3):183–187, June 1986. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497286900118.

Freel:1999:FSF

[Fre99] Mark S. Freel. The financing of small firm product innovation within the UK. Technovation, 19(12):707–719, December 1999. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497299000826.

Freel:2005:PIS

[Fre05] Mark S. Freel. Patterns of innovation and skills in small firms. Technovation, 25(2):123–134, February 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000828.

Frumerman:1992:CER

[Fru92] Robert Frumerman. Cost-effective R&D. Technovation, 12(3):203–208, April 1992. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729290036H.

Fernandes:2006:LIB

[FRW06] Kiran Jude Fernandes, Vinesh Raja, and Andrew Whalley. Lessons from implementing the balanced scorecard in a small and medium size manufacturing organization. Technovation, 26(5–6):623–634, May/June 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000696.

Fernandes:2006:AVR

[FRWT06] Kiran Jude Fernandes, Vinesh Raja, Andrew White, and Christos-Dimitris Tsinopoulos. Adoption of virtual reality within construction processes: a factor analysis approach. Technovation, 26(1):111–120, January 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001270.

Feller:1984:RTC

[FS84] Irwin Feller and John D. Sink. Regulation and technological change in the American meat industry. Technovation, 2(2): 69–87, May 1984. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497284900130.

Farjam:2023:CMO

[FSA23] Foroogh Farjam, Payam Shojaei, and Kazem Askarifar. A conceptual model for open innovation risk management based on the capabilities of SMEs: a multi-level fuzzy MADM approach. Technovation, 127(??):??, September 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223001554.

Fujisue:1998:PAI

[Fuj98] Kenzo Fujisue. Promotion of academia-industry cooperation in Japan — establishing the “law of promoting technology transfer from university to industry” in Japan. Technovation, 18(6–7):371–381, January 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000558.

Fukugawa:2009:DLA

[Fuk09] Nobuya Fukugawa. Determinants of licensing activities of local public technology centers in Japan. Technovation, 29 (12):885–892, December 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497209000881.

Funk:2006:FMP

[Fun06] Jeffrey L. Funk. The future of mobile phone-based Intranet applications: a view from Japan. Technovation, 26(12): 1337–1346, December 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497205001100.

Frey:1991:NTE

[FV91] Marco Frey and Marco Vivarelli. New technology and employment in Italian telecommunications. Technovation, 11(5):303–314, July 1991. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497291900283.

Foo:2006:MTD

[FWL06] Check Teck Foo, Loh Sow Wai, and Tan Soo Lang. The mind of a technopreneuress: differentiating the selfleading, entrepreneurial from custodial, managerial female. Technovation, 26(2):175–184, February 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000234.

Galbreath:2005:WRM

[Gal05] Jeremy Galbreath. Which resources matter the most to firm success? An exploratory study of resource-based theory. Technovation, 25(9):979–987, September 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204000355.

Galanakis:2006:IPM

[Gal06a] Kostas Galanakis. Innovation process. Make sense using systems thinking. Technovation, 26(11):1222–1232, November 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205001057.

Galende:2006:ATI

[Gal06b] Jesús Galende. Analysis of technological innovation from business economics and management. Technovation, 26 (3):300–311, March 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497205000763.

Gallagher:2012:BBL

[Gal12] Scott R. Gallagher. The battle of the blue laser DVDs: the significance of corporate strategy in standards battles. Technovation, 32(2):90–98, February 2012. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211001489.

Gonzalez-Alvarez:2007:AIR

[GÁNA07] Nuria González-Álvarez and Mariano Nieto-Antolín. Appropriability of innovation results: an empirical study in Spanish manufacturing firms. Technovation, 27(5):280–295, May 2007. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497206001222.

Gardiner:1985:BRH

[Gar85] Paul Gardiner. Book review: High technology small firms: Ray Oakey, Frances Pinter (Publisher) Ltd., London, 1984, 179 pages, price £16.50. Technovation, 3(2):153–154, June 1985. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497285900082.

Gardiner:1987:BRE

[Gar87] Paul Gardiner. Book review: The economics of invention: a study of the determinants of inventive activity: G. Wyatt, Wheatsheaf Books, Brighton, U.K., 1986, 245 pp., £28.50 (cloth). Technovation, 6(2):147, June 1987. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497287900162.

Gardiner:1994:BRR

[Gar94] Paul Gardiner. Book review: The rise and fall of strategic planning: by Henry Mintzberg. Prentice Hall, New York, 1994, 458 pp., ISBN 0-13-781824-6, £19.95. Technovation, 14(9):626–628, November 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900507.

Gardiner:1995:BRW

[Gar95] Paul Gardiner. Book review: What is strategy — And does it matter?: by R. Whittington. Routledge, 11 New Fetter Lane, London EC4P 4EE, UK, 1993, 165 pp., ISBN 0-415-07385-5. £10.99. Technovation, 15(3):190–192, April 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497295900292.

Gardiner:1996:BRN

[Gar96] Paul Gardiner. Book review: New technology-based firms in the 1990s: by R. Oakey (ed.). Paul Chapman Publishing, London, 1994, 218 pp., ISBN 1-85396-274-0, £39.95. Technovation, 16(2):97, February 1996. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497296900085.

Gargeya:2005:PLP

[Gar05] Vidyaranya B. Gargeya. Plant level performance measurement: an exploratory case study of a pharmaceutical encapsulation company. Technovation, 25(12):1457–1467, December 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205000623.

Granados:2024:ISJ

[GARM24] Cristian Granados, Yarid Ayala, and Monica Ramos-Mejia. Is it substantive or just symbolic? Understanding innovation theater in organisations: the case of technologybased innovation. Technovation, 129(??):??, January 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001918.

Garcia:1996:IEE

[GAS96] Alberto Garcia, Fernand Amesse, and Mario Silva. The indirect economic effects of Ecopetrol’s contracting strategy for informatics development. Technovation, 16(9):469–485, 519–521, September 1996. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497296000417.

Gazdik:1986:SEI

[Gaz86] Igor Gazdik. Stimulation enhances inventiveness on the shop-floor. Technovation, 4(2):131–141, April 1986. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497286900040.

Gilsing:2011:DTT

[GBFvdS11] Victor Gilsing, Rudi Bekkers, Isabel Maria Bodas Freitas, and Marianne van der Steen. Differences in technology transfer between science-based and developmentbased industries: Transfer mechanisms and barriers. Technovation, 31(12):638–647, December 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211000927.

Gil:2003:IMT

[GBL03] Y. Gil, S. Bong, and J. Lee. Integration model of technology internalization modes and learning strategy: globally late starter Samsung’s successful practices in South Korea. Technovation, 23(4):333–347, April 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201001109.

Griffiths:1998:SMKa

[GBM98a] Dorothy Griffiths, Max Boisot, and Veronica Mole. Strategies for managing knowledge assets: a tale of two companies. Technovation, 18(8–9):529–539, 587–588, August/ September 1998. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649729800025X.

Griffiths:1998:SMKb

[GBM98b] Dorothy Griffiths, Max Boisot, and Veronica Mole. Strategies for managing knowledge assets: a tale of two companies. Technovation, 18(8–9):587–588, August/September 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497298901249.

Gurca:2023:ACT

[GBV23] Andrei Gurca, Mehdi Bagherzadeh, and Rezvan Velayati. Aligning the crowdsourcing type with the problem attributes to improve solution search efficacy. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222001602.

Ghasemzadeh:2022:CUI

[GBY22] Khatereh Ghasemzadeh, Guido Bortoluzzi, and Zornitsa Yordanova. Collaborating with users to innovate: a systematic literature review. Technovation, 116(??):??, August 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000347.

Gupta:1993:SMEa

[GC93a] Yash P. Gupta and David C. W. Chin. Strategy making and environment: an organizational life cycle perspective. Technovation, 13(1):27–44, January 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390012K.

Gupta:1993:SMEb

[GC93b] Yash P. Gupta and David C. W. Chin. Strategy making and environment: an organizational life cycle perspective. Technovation, 13(1):56–57, January 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390016O.

Gupta:1996:MIF

[GC96] Mahesh Gupta and Garret Cawthon. Managerial implications of flexible manufacturing for small/medium-sized enterprises. Technovation, 16(2):77–83, 93–94, February 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497295000232.

Gunasekaran:1998:IPIa

[GC98a] A. Gunasekaran and P. Cecille. Implementation of productivity improvement strategies in a small company. Technovation, 18(5):311–320, ???? 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000054.

Gunasekaran:1998:IPIb

[GC98b] A. Gunasekaran and P. Cecille. Implementation of productivity improvement strategies in a small company. Technovation, 18(5):362–363, ???? 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649729880028X.

Garrone:1999:MDD

[GC99] Paola Garrone and Massimo G. Colombo. Market-driven design of innovative services: the case of multimedia home banking. Technovation, 19(9):537–549, September 1999. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299000334.

Gaughan:2010:SFU

[GC10a] Monica Gaughan and Elizabeth A. Corley. Science faculty at US research universities: the impacts of university research center-affiliation and gender on industrial activities. Technovation, 30(3):215–222, March 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209001722.

Guan:2010:MIP

[GC10b] Jiancheng Guan and Kaihua Chen. Measuring the innovation production process: a cross-region empirical study of China’s high-tech innovations. Technovation, 30(5–6):348– 358, May/June 2010. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497210000301.

Gilbert:1996:UPK

[GCH96] Myrna Gilbert and Martyn Cordey-Hayes. Understanding the process of knowledge transfer to achieve successful technological innovation. Technovation, 16(6):301–312, June 1996. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497296000120.

Gupta:2001:OSB

[GCS01] Mahesh C. Gupta, Anthony Czernik, and Ramji D. Sharma. Operations strategies of banks — using new technologies for competitive advantage. Technovation, 21(12):775–782, December 2001. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497201000177.

Gilsing:2008:UNC

[GD08] Victor A. Gilsing and G. M. Duysters. Understanding novelty creation in exploration networks-Structural and relational embeddedness jointly considered. Technovation, 28 (10):693–708, October 2008. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497208000412.

Gomes:2003:MAB

[GdWNPC03] J. F. S. Gomes, P. C. de Weerd-Nederhof, A. W. Pearson, and M. P. Cunha. Is more always better? An exploration of the differential effects of functional integration on performance in new product development. Technovation, 23(3):185–191, March 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201001079.

Geels:2006:CEM

[Gee06] F. W. Geels. Co-evolutionary and multi-level dynamics in transitions: the transformation of aviation systems and the shift from propeller to turbojet (1930– 1970). Technovation, 26(9):999–1016, September 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205001276.

Genet:2012:WMT

[GEG12] Corine Genet, Khalid Errabi, and Caroline Gauthier. Which model of technology transfer for nanotechnology? A comparison with biotech and microelectronics. Technovation, 32(3–4):205–215, March/April 2012. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211001519.

Geisler:1997:ITC

[Gei97] Eliezer Geisler. Intersector technology cooperation: Hard myths, soft facts. Technovation, 17(6):309–320, June 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497296001174.

Gerstlberger:2004:RIS

[Ger04] Wolfgang Gerstlberger. Regional innovation systems and sustainability-selected examples of international discussion. Technovation, 24(9):749–758, September 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202001529.

Globocnik:2021:DLU

[GF21] Dietfried Globocnik and Rita Faullant. Do lead users cooperate with manufacturers in innovation? Investigating the missing link between lead userness and cooperation initiation with manufacturers. Technovation, 100(??):??, February 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497220300596.

Goldsmith:1995:GSI

[GFE95] Ronald E. Goldsmith, Jon B. Freiden, and Jacqueline K. Eastman. The generality/specificity issue in consumer innovativeness research. Technovation, 15(10):601–612, December 1995. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729599328D.

Gokhberg:2012:NDR

[GFK12] Leonid Gokhberg, Konstantin Fursov, and Oleg Karasev. Nanotechnology development and regulatory framework: the case of Russia. Technovation, 32(3–4):161–162, March/ April 2012. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649721200003X.

Granato:2022:MBU

[GFvT22] Giulia Granato, Arnout R. H. Fischer, and Hans C. M. van Trijp. “Misalignments between users and designers as source of inspiration: a novel hybrid method for physical new product development”. Technovation, 111(??):??, March 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221001723.

Gupta:1993:IMI

[GG93] Yash P. Gupta and Tor Guimaraes. Issues in management information systems planning. Technovation, 13(8):533– 544, December 1993. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497293900676. See [Ano93k].

Gupta:2003:ABC

[GG03] M. Gupta and K. Galloway. Activity-based costing/ management and its implications for operations management. Technovation, 23(2):131–138, February 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000931.

Grimaldi:2005:BIN

[GG05] Rosa Grimaldi and Alessandro Grandi. Business incubators and new venture creation: an assessment of incubating models. Technovation, 25(2):111–121, February 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000762.

Guo:2011:PTL

[GG11] Bin Guo and Jing-Jing Guo. Patterns of technological learning within the knowledge systems of industrial clusters in emerging economies: Evidence from China. Technovation, 31(2–3):87–104, February/March 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210001148.

Greco:2022:PMP

[GGGR22] Marco Greco, Francesca Germani, Michele Grimaldi, and Dragana Radicic. Policy mix or policy mess? Effects of cross-instrumental policy mix on eco-innovation in German firms. Technovation, 117(??):??, September 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497220300663.

Guo:2018:GEP

[GGJ18]  Di Guo, Yan Guo, and Kun Jiang.  Governance

and effects of public R&D subsidies: Evidence from China. Technovation, 74–75(??):18–31, June/July 2018. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497218302372.

Guo:2022:GRS

[GGJ22] Di Guo, Yan Guo, and Kun Jiang. Government R&D support and firms’ access to external financing: funding effects, certification effects, or both? Technovation, 115(??):??, July 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000165.

Garavelli:2002:MKT

[GGS02] A. Claudio Garavelli, Michele Gorgoglione, and Barbara Scozzi. Managing knowledge transfer by knowledge technologies. Technovation, 22(5):269–279, May 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000098.

Garrido:2008:OEC

[GGS08] M. José Garrido, Ana Gutiérrez, and Rebeca San José. Organizational and economic consequences of business e-procurement intensity. Technovation, 28(9):615–629, September 2008. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497207001538.

Gupta:1991:IIJ

[GH91] Yash P. Gupta and Sunderesh Heragu. Implications of implementing just-in-time systems. Technovation, 11(3):143– 162, April 1991. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729190031X.

Garnsey:2015:ISI

[GH15] Elizabeth Garnsey and Chang Chieh Hang. Introduction to special issue of Technovation on opportunity recognition and creation. Technovation, 39–40(??):1–3, May/June 2015. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497215000115.

Granstrand:2021:IEC

[GH21] Ove Granstrand and Marcus Holgersson. Innovation ecosystems: a conceptual review and a new definition. Technovation, ??(??):??, ???? 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497218303870.

Ghosh:2022:DTI

[GHHH22] Swapan Ghosh, Mat Hughes, Ian Hodgkinson, and Paul Hughes. Digital transformation of industrial businesses: a dynamic capability approach. Technovation, 113(??):??, May 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497221001954.

Guerrero:2021:BUI

[GHU21] Maribel Guerrero, Sohvi Heaton, and David Urbano. Building universities’ intrapreneurial capabilities in the digital era: the role and impacts of Massive Open Online Courses (MOOCs). Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497218307855.

Gilbert:1996:BRM

[Gil96] John A. Gilbert. Book review: Managing change in the workplace — New approaches to employee relations: by Alan Cave, Coopers & Lybrand, Kogan Page, London, 1994, ISBN 0-7494-1007-8, 209 pp. (hardback), £25.00. Technovation, 16(3):99–100, March 1996. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497296901753.

Gilbert:1997:BRT

[Gil97] John Gilbert. Book review: Touchstones: Ten new ideas revolutionizing business: by William A. Band. John Wiley & Sons, Chichester (UK) and New York, 1994. ISBN 0-471-31096-4, $24.95, ix + 306 pp. Technovation, 17(7):416, July 1997. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297868095.

Gimenez:2006:INT

[Gim06] Gregorio Gimenez. Investment in new technology: Modelling the decision process. Technovation, 26(3):345–350, March 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205000635.

Giuliani:2009:BRI

[Giu09] Elisa Giuliani. Book review: Innovation Networks and Knowledge Clusters. Findings and Insights from the US, EU and Japan: Carayannis, E., Assimakopoulos, D., Kondo, M. Palgrave MacMillan Publishers, New York, NY, USA (2008). 224 pp., price $80.00 (hardcover), hardback, ISBN 1-4039-4245-5. Technovation, 29(1):72–73, January 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720800093X.

Goldsmith:1991:EAI

[GK91] Ronald E. Goldsmith and John R. Kerr. Entrepreneurship and adaption-innovation theory. Technovation, 11(6): 373–382, September 1991. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729190019Z.

Gupta:2006:ERP

[GK06] Mahesh Gupta and Amarpreet Kohli. Enterprise resource planning systems and its implications for operations function. Technovation, 26(5–6):687–696, May/June 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001841.

Gredel:2012:PBI

[GKB12] Daniel Gredel, Matthias Kramer, and Boris Bend. Patentbased investment funds as innovation intermediaries for SMEs: In-depth analysis of reciprocal interactions, motives and fallacies. Technovation, 32(9–10):536–549, September/ October 2012. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497211001337.

Galetsi:2023:EBE

[GKK23] Panagiota Galetsi, Korina Katsaliaki, and Sameer Kumar. Exploring benefits and ethical challenges in the rise of mHealth (mobile healthcare) technology for the common good: An analysis of mobile applications for health specialists. Technovation, 121(??):??, March 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222001456.

Good:2019:TTE

[GKSW19] Matthew Good, Mirjam Knockaert, Birthe Soppe, and Mike Wright. The technology transfer ecosystem in academia. An organizational design perspective. Technovation, 82–83 (??):35–50, April/May 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497217301657.

Groen:2010:OIF

[GL10] Aard J. Groen and Jonathan D. Linton. Is open innovation a field of study or a communication barrier to theory development? Technovation, 30(11–12):554, November/ December 2010. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497210000921.

Ge:2022:RKC

[GL22] Shuang Ge and Xielin Liu. The role of knowledge creation, absorption and acquisition in determining national competitive advantage. Technovation, 112(??):??, April 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001772.

Geissinger:2023:SMA

[GLÖS23] Andrea Geissinger, Christofer Laurell, Christina Öberg, and Christian Sandström. Social media analytics for innovation management research: a systematic literature review and future research agenda. Technovation, 123(??):??, May 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223000238.

Geum:2013:IES

[GLYP13] Youngjung Geum, Sungjoo Lee, Byungun Yoon, and Yongtae Park. Identifying and evaluating strategic partners for collaborative R&D: Index-based approach using patents and publications. Technovation, 33(6–7):211–224, June/ July 2013. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497213000503.

Grupp:2000:PII

[GM00] Hariolf Grupp and Shlomo Maital. Perceived innovation of Israel’s largest firms:: an empirical study. Technovation, 20(3):129–138, March 2000. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497299001145.

Guan:2003:ICE

[GM03] J. Guan and N. Ma. Innovative capability and export performance of Chinese firms. Technovation, 23(9):737–747, September 2003. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497202000135.

Grilli:2023:BTW

[GM23] Luca Grilli and Riccardo Marzano. Bridges over troubled water: Incubators and start-ups’ alliances. Technovation, 121(??):??, March 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497222002401.

Gupta:2023:QCL

[GMB^{+}23] Shivam Gupta, Sachin Modgil, Priyanka C. Bhatt, Charbel Jose Chiappetta Jabbour, and Sachin Kamble. Quantum computing led innovation for achieving a more sustainable Covid-19 healthcare industry. Technovation, 120(??): ??, February 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497222000918.

Garcia-Morales:2007:IPM

[GMLMVJ07] Víctor J. García-Morales, Francisco Javier Lloréns-Montes, and Antonio J. Verdú-Jover. Influence of personal mastery on organizational performance through organizational learning and innovation in large firms and SMEs. Technovation, 27(9):547–568, September 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000375.

Garcia-Muina:2007:EMS

[GMNL07] Fernando E. García-Muiña and José E. Navas-López. Explaining and measuring success in new business: the effect of technological capabilities on firm results. Technovation, 27(1–2):30–46, January/February 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497206000526.

Gupta:2023:RIA

[GMWK23] Shivam Gupta, Sachin Modgil, Christina W. Y. Wong, and Arpan Kumar Kar. The role of innovation ambidexterity on the relationship between cognitive computing capabilities and entrepreneurial quality: a comparative study of India and China. Technovation, 127(??):??, September 2023.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001463.

Goetze:2010:EEC

[Goe10] Christiane Goetze. An empirical enquiry into co-patent networks and their stars: the case of cardiac pacemaker technology. Technovation, 30(7–8):436–446, July/August 2010. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210000428.

Goffin:2007:BRA

[Gof07] Keith Goffin. Book review: Arch Woodside (Ed.), Managing Product Innovation, Elsevier, Amsterdam, 2005. Technovation, 27(6–7):412, June/July 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497206000393.

Gold:1982:PTC

[Gol82] Bela Gold. Productivity, technological change and international competitiveness. Technovation, 1(3):203–213, February 1982. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497282900037.

Gold:1986:PCA

[Gol86] Bela Gold. Perspectives on continuing advances in automation: Past limitations and emerging potentials. Technovation, 4(2):153–162, April 1986. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497286900064.

Gold:1987:EEF

[Gol87] Bela Gold. On the economic evaluation of future manufacturing systems. Technovation, 6(4):227–245, September 1987. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497287900721.

Gold:1990:IIP

[Gol90] Bela Gold. Improving industrial performance through integrating quality and productivity advances. Technovation, 10(8):521–530, December 1990. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729090048O.

Gossart:2005:RSL

[Gos05] Cédric Gossart. Routines and the sustainable lock-out of Moroccan oil refineries. Technovation, 25(12):1468–1475, December 2005. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497205000921.

Guerrieri:2004:IDE

[GP04] Paolo Guerrieri and Carlo Pietrobelli. Industrial districts’ evolution and technological regimes: Italy and Taiwan. Technovation, 24(11):899–914, November 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000488.

Guerrero:2022:EUM

[GP22] Maribel Guerrero and Rhiannon Pugh. Entrepreneurial universities’ metamorphosis: Encountering technological and emotional disruptions in the COVID-19 era. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222001316.

Garcia-Perez:2023:RHS

[GPCNS^{+}23] Alexeis Garcia-Perez, Juan Gabriel Cegarra-Navarro, Mark Paul Sallos, Eva Martinez-Caro, and Anitha Chinnaswamy. Resilience in healthcare systems: Cyber security and digital transformation. Technovation, 121(??):??, March 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497222001304.

Gros-Pietro:1989:FAF

[GPR89] Gian Maria Gros-Pietro and Secondo Rolfo. Flexible automation and firm size: some empirical evidence on the Italian case. Technovation, 9(6):493–503, October 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900291.

Gong:2021:DUD

[GR21] Cheng Gong and Vincent Ribiere. Developing a unified definition of digital transformation. Technovation, 102(??): ??, April 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497220300894.

Gao:2023:KAI

[GR23] Xue Gao and Varun Rai. Knowledge acquisition and innovation quality: the moderating role of geographical characteristics of technology. Technovation, 125(??):??, July 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000779.

Grabowski:1982:PPI

[Gra82] Henry Grabowski. Public policy and innovation: the case of pharmaceuticals. Technovation, 1(3):157–189, February 1982. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497282900013.

Grantham:2001:BRC

[Gra01] Andrew Grantham. Book review: From control to drift: the dynamics of corporate information infrastructures: Claudio U. Ciborra, Kristin Braa, Antonio Cordella, Bo Dahlbom, Angelo Failla, Ole Hanseth, Vidar Hepsø, Jan Ljungberg, Eric Monteiro and Kai A. Simon; Oxford University Press, 2000. Technovation, 21(4):264–265, April 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000687.

Grantham:2003:BRS

[Gra03] A. Grantham. Book review: Supply chain cybermastery: building high performance supply chains of the future: Andrew J. Berger and John L. Gattorna, Gower, Aldershot, 2001. Technovation, 23(9):781, September 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000270.

Grantham:2004:BRC

[Gra04] Andrew Grantham. Book review: Competition for technological leadership: EU policy for high technology: Johan Lembke; Edward Elgar, Cheltenham, UK, 2002. Technovation, 24(4):369–370, April 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203001160.

Granstrand:2024:CGT

[Gra24] Ove Granstrand. A cooperative game theory systems approach to the value analysis of (innovation) ecosystems. Technovation, 130(??):??, February 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223002377.

Greasley:2006:UPM

[Gre06] A. Greasley. Using process mapping and business process simulation to support a process-based approach to change in a public sector organisation. Technovation, 26(1):95– 103, January 2006. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497204001142.

Grier:1998:BRD

[Gri98] David Grier. Book review: ‘Designing for Change’: a practical guide to business transformation: By Colin Bainbridge. John Wiley & Sons Ltd, Baffins Lane, Chichester, West Sussex, P019 1UD, UK, 1996, 250 pp. ISBN 0-471-96452-2, Hardback. Technovation, 18(2):148, February 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298800138.

Grimpe:2006:MUU

[Gri06] Christoph Grimpe. Making use of the unused: shelf warmer technologies in research and development. Technovation, 26(7):770–774, July 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497204001804.

Grupp:1993:EGI

[Gru93] Hariolf Grupp. Efficiency of government intervention in technical change in telecommunications: ten national economies compared. Technovation, 13(4):187–220, May 1993. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729390018Q.

Goransson:2005:LWI

[GS05] Bo Göransson and Johan Söderberg. Long waves and information technologies-on the transition towards the information society. Technovation, 25(3):203–211, March 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203001159.

Gurdon:2010:LSS

[GS10] Michael A. Gurdon and Karel J. Samsom. A longitudinal study of success and failure among scientist-started ventures. Technovation, 30(3):207–214, March 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720900159X.

Ghazinoory:2021:RDI

[GSFS21] Sepehr Ghazinoory, Alfred Sarkissian, Mojtaba Farhanchi, and Fatemeh Saghafi. Renewing a dysfunctional innovation ecosystem: the case of the Lalejin ceramics and pottery. Technovation, ??(??):??, ???? 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497220300122.

Gama:2022:EEC

[GSP^{+}22] Fábio Gama, David Sjödin, Vinit Parida, Johan Frishammar, and Joakim Wincent. Exploratory and exploitative capability paths for innovation: a contingency framework for harnessing fuzziness in the front end. Technovation, 113(??):??, May 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497221001978.

Garcia:2008:NPI

[GST08] Nuria García, M. José Sanzo, and Juan A. Trespalacios. New product internal performance and market performance: Evidence from Spanish firms regarding the role of trust, interfunctional integration, and innovation type. Technovation, 28(11):713–725, November 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208000047.

Gagnon:1993:ANT

[GT93] Yves-Chantal Gagnon and Jean-Marie Toulouse. Adopting new technologies: an entrepreneurial act. Technovation, 13 (7):411–424, November 1993. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729390025Q.

See [Ano93a].

Gunasekaran:2002:IAA

[GTW02] A. Gunasekaran, E. Tirtiroglu, and V. Wolstencroft. An investigation into the application of agile manufacturing in an aerospace company. Technovation, 22(7):405–415, July 2002. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201000396.

Guichard:2005:SRD

[Gui05] Renelle Guichard. Suggested repositioning of defence R&D within the French system of innovation. Technovation, 25(3):195–201, March 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497203001317.

Guimon:2011:PBG

[Gui11] José Guimón. Policies to benefit from the globalization of corporate R&D: an exploratory study for EU countries. Technovation, 31(2–3):77–86, February/March 2011. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210000891.

Gomez-Uranga:2014:EED

[GUMZI14] Mikel Gómez-Uranga, Juan Carlos Miguel, and Jon Mikel Zabala-Iturriagagoitia. Epigenetic Economic Dynamics: the evolution of big Internet business ecosystems, evidence for patents. Technovation, 34(3):177–189, March 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213001600.

Gunasekaran:1997:EIJa

[Gun97a] A. Gunasekaran. Essentials of international and joint R&D projects. Technovation, 17(11–12):637–647, November/ December 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497297000436.

Gunasekaran:1997:EIJb

[Gun97b] A. Gunasekaran. Essentials of international and joint R&D projects. Technovation, 17(11–12):725–727, November/ December 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497297904600.

Gupta:1988:OIF

[Gup88] Yash P. Gupta. Organizational issues of flexible manufacturing systems. Technovation, 8(4):255–269, December 1988. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497288900296.

Gupta:1989:MIS

[Gup89] Yash P. Gupta. Management information systems planning: analysis and techniques. Technovation, 9(1):63–81, May 1989. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497289900412.

Gupta:1990:VAE

[Gup90] Yash P. Gupta. Various aspects of expert systems: applications in manufacturing. Technovation, 10(7):487–504, October 1990. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729090027H.

Gupta:1996:OES

[Gup96] Mahesh Gupta. Operations effectiveness for a successful implementation of CIM. Technovation, 16(10):589–594, October 1996. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297828970.

Gupta:2000:IES

[Gup00] Mahesh Gupta. Implications of expert systems for the operations of financial institutions. Technovation, 20(9): 509–516, September 2000. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497299001650.

Guy:1995:WBW

[Guy95] Ian Guy. What’s the best way to develop your product? Technovation, 15(8):523–525, October 1995. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729596590P.

Guy:1996:LAS

[Guy96a] Ian Guy. A look at Aston Science Park. Technovation, 16(5):217–218, May 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497296000028.

Guy:1996:NVA

[Guy96b] Ian Guy. New ventures on an ancient campus. Technovation, 16(6):269–270, June 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497296000144.

Guy:1997:HNB

[Guy97] Ian Guy. Harwell’s new business centre. Technovation, 17(8):467–469, August 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497297000369.

Gilsing:2010:PPC

[GvBR10] Victor A. Gilsing, Elco van Burg, and A. Georges L. Romme. Policy principles for the creation and success of corporate and academic spin-offs. Technovation, 30(1):12– 23, January 2010. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497209001060.

Goto:1987:TPJ

[GW87] Akira Goto and Ryuhei Wakasugi. Technology policy in Japan: a short review. Technovation, 5(4):269–279, February 1987. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497287900678.

Griffiths:2010:WGF

[GW10] William Griffiths and Elizabeth Webster. What governs firm-level R&D: Internal or external factors? Technovation, 30(7–8):471–481, July/August 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210000465.

Gunasekaran:2005:PMC

[GWM05] A. Gunasekaran, H. James Williams, and Ronald E. Mc-Gaughey. Performance measurement and costing system in new enterprise. Technovation, 25(5):523–533, May 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203001767.

Gong:2023:SSD

[GWZ^{+}23] Tianyu Gong, Xiaolei Wang, Li Zhang, Xuehao Gao, and En Xie. Symbolic or substantial? Different responses of state-owned and privately owned firms to government innovation policies. Technovation, 127(??):??, September 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001384.

Gu:2023:SSI

[GZXX23] Jing Gu, Fujuan Zhang, Xun Xu, and Chaokai Xue. Stay or switch? The impact of venture capitalists’ movement across network communities on enterprises’ innovation performance. Technovation, 125(??):??, July 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000810.

Haarder:1989:BH

[Haa89] Bertel Haarder. Bertel Haarder. Technovation, 9(2–3):203– 204, June 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497289900576.

Harbi:2009:EHT

[HAA09] Sana Harbi, Mariam Amamou, and Alistair R. Anderson. Establishing high-tech industry: the Tunisian ICT experience. Technovation, 29(6–7):465–480, June/July 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208001399.

Haavisto:2014:ODF

[Haa14] Piia Haavisto. Observing discussion forums and product innovation — a way to create consumer value? Case heartrate monitors. Technovation, 34(4):215–222, April 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213001478.

Hadjimanolis:1999:BIS

[Had99] Athanasios Hadjimanolis. Barriers to innovation for SMEs in a small less developed country (Cyprus). Technovation, 19(9):561–570, September 1999. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299000346.

Hagedoorn:1990:OMI

[Hag90] John Hagedoorn. Organizational modes of inter-firm cooperation and technology transfer. Technovation, 10(1):17– 30, February 1990. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729090039M.

Hull:1984:ORI

[HAH84] Frank Hull, Koya Azumi, and Jerald Hage. Organizing resources for innovation and productivity: a preliminary abstract draft. Technovation, 2(4):293–297, July 1984. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497284900087.

Hall:1996:RCW

[Hal96] David J. Hall. The role of creativity within best practice manufacturing. Technovation, 16(3):115–121, March 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729500050X.

Hall:1997:BRQ

[Hal97a] Dave Hall. Book review: E-Quality: Bridging the total involvement gap: by Gillian Shapiro, Paul Levy and Swasti Mitter. Sussex Academic Press, Brighton, 1996, x + 118 pp., ISBN 1-89872312-5 (paperback), £12.95 in UK and rest of world, $19.95 in USA. Technovation, 17(6):347– 348, June 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497297863390.

Hall:1997:OCK

[Hal97b] David J. Hall. Organisational change: Kinetic theory and organisational resonance. Technovation, 17(1):11–24, January 1997. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497296000752.

Hall:2014:IED

[Hal14] Jeremy Hall. Innovation and entrepreneurial dynamics in the Base of the Pyramid. Technovation, 34(5–6):265–269, May/June 2014. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497214000364.

Hameri:1994:UTN

[Ham94] Ari-Pekka Hameri. Using technical norm in explaining technological action. Technovation, 14(8):505–514, October 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294901503.

Hameri:1996:TTB

[Ham96] Ari-Pekka Hameri. Technology transfer between basic research and industry. Technovation, 16(2):51–57, 91–92, February 1996. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497295000305.

Hanel:1994:IFT

[Han94] Petr Hanel. Interindustry flows of technology: an analysis of the Canadian patent matrix and input-output matrix for 1978–1989. Technovation, 14(8):529–548, October 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729490152X.

Hansen:1995:PPK

[Han95] Povl A. Hansen. Publicly produced knowledge for business: When is it effective? Technovation, 15(6):387–397, August 1995. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729596599O.

Hanel:2006:IPR

[Han06] Petr Hanel. Intellectual property rights business management practices: a survey of the literature. Technovation, 26(8):895–931, August 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497205001793.

Harman:1981:MRI

[Har81] Alvin Jay Harman. On military RDT&E in Israel. Technovation, 1(1):31–41, February 1981. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0166497281900055.

Harding:1993:BRTa

[Har93a] Rebecca Harding. Book review: Technology and the wealth of nations: the dynamics of constructed advantage: edited by Dominic Foray and Christopher Freeman. Frances Pinter, 1993, ISBN 1-85567-082-8. Technovation, 13(6):407, September 1993. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497293900838.

Harding:1993:BRTb

[Har93b] Rebecca Harding. Book review: Technology and work in German industry: edited by Norbert Altman, Christopher Köhler and Pamela Meil. Routledge, 1992, ISBN 0-415-07926-8, £45 (hardback). Technovation, 13(6):408, September 1993. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497293900849.

Harding:1996:ISC

[Har96] Rebecca Harding. Implementing strategic change: a survey of British and German workers. Technovation, 16(3):101– 113, March 1996. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497295000488.

Harabi:1997:CRSa

[Har97a] Najib Harabi. Channels of R&D spillovers: an empirical investigation of Swiss firms. Technovation, 17(11–12):627– 635, November/December 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497297000643.

Harabi:1997:CRSb

[Har97b] Najib Harabi. Channels of R&D spillovers: an empirical investigation of Swiss firms. Technovation, 17(11–12):724– 725, November/December 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497297904594.

Harding:1997:BRF

[Har97c] Rebecca Harding. Book review: Forbidden Fruit: an analysis of bootlegging, uncertainty and learning in corporate R&D: by Peter Augsdorfer, Avebury, Aldershot, 1996, ISBN 1-85972-330-0, £35. Technovation, 17(11–12):734–735, November/December 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649729788772X.

Hay:1981:CUE

[Hay81] D. Robert Hay. A Canadian university experience in technological innovation and entrepreneurship. Technovation, 1(1):43–55, February 1981. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497281900067.

Hayward:1982:IEL

[Hay82] George Hayward. Infrared Engineering Limited. Technovation, 1(4):317–325, August 1982. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497282900116.

Hayward:1983:BRI

[Hay83] George Hayward. Book review: Industrial innovation and public policy: Preparing for the 1980’s and 1990’s: by Roy Rothwell and Walter Zegveld, Francis Pinter Publications Limited, London, 1981, £13.50 hardback, 251 pages. Technovation, 2(1):61, February 1983. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497283900081.

Hayward:1984:BRD

[Hay84] George Hayward. Book review: Diffusion of innovations: Everett M. Rogers, The Free Press, New York (a Division of Macmillan Publishing Co. Inc. New York. Collier Macmillan Publishers, London), 3rd edition, 1983. Hardback £17.25, 453 pages. Technovation, 2(2):147, May 1984. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649728490018X.

Hayward:1985:BRC

[Hay85a] George Hayward. Book review: Corporate innovation: Marketing and strategy: by Gordon R. Foxall. Croom Helm, London and Canberra, 276 pages, Hardback £19.95, Softback £9.95. Technovation, 3(3):229, August 1985. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497285900197.

Hayward:1985:BRN

[Hay85b] George Hayward. Book review: Need assessment: a key to user-oriented product innovation: by Knut Holt, Horst Geschka and Giovanni Peterlongo. John Wiley & Sons, Chichester, 1984, 187 pages, Hardback £13.25. Technovation, 3(3):228, August 1985. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497285900185.

Hayward:1985:BRP

[Hay85c] George Hayward. Book review: Product innovation management: a workbook for management in industry: (Second edition), by Knut Holt. Butterworths, London, 1983, 295 pages, Hardback £15.00. Technovation, 3(3):227, August 1985. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497285900173.

Hayward:1985:BRM

[Hay85d] George Hayward. Book review: The management of innovation: by R. C. Parker, John Wiley & Sons, Chichester, 1983, £14.95 hardback, 221 pages. Technovation, 3(1):73, February 1985. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497285900392.

Hayward:1985:Ea

[Hay85e] George Hayward. Editorial. Technovation, 3(3):159–160, August 1985. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497285900100.

Hayward:1985:Eb

[Hay85f] George Hayward. Editorial. Technovation, 3(4):233, October 1985. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497285900215.

Hayward:1986:BRIb

[Hay86a] G. Hayward. Book review: Industrial product innovation: Organisation and management: By Dr. F. A. Johne. Croom Helm Ltd., 1985, 177 pages, hardback £15.95. Technovation, 4(4):318–319, August 1986. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497286900350.

Hayward:1986:Ed

[Hay86b] G. Hayward. Editorial. Technovation, 4(4):251, August 1986. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497286900283.

Hayward:1986:BRD

[Hay86c] George Hayward. Book review: Design, innovation and long cycles in economic development: edited by Christopher Freeman. Department of Design Research, Royal College of Art, London, 1984, 364 pages, Softback £32.00. Technovation, 4(1):85, January 1986. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497286900246.

Hayward:1986:BRG

[Hay86d] George Hayward. Book review: Going for growth: Technological innovation in manufacturing industries: by R. C. Parker. John Wiley & Sons, Chichester, 1985, xxi + 251 pages, Hardback £13.75. Technovation, 4(1):86, January 1986. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649728690026X.

Hayward:1986:BRIa

[Hay86e] George Hayward. Book review: Innovation and entrepreneurship: Practice and principles: by Peter F. Drucker. Heinemann, London, 1985. xi + 258 pages, Hardback £10.95. Technovation, 4(1):85–86, January 1986. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497286900258.

Hayward:1986:BRR

[Hay86f] George Hayward. Book review: Reindustrialization and technology: By Roy Rothwell and Walter Zegveld. Longman Group Limited, 1985, 282 pages, Softback £9.95. Technovation, 4(4):318, August 1986. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497286900349.

Hayward:1986:Ea

[Hay86g] George Hayward. Editorial. Technovation, 4(1):1, January 1986. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497286900180.

Hayward:1986:Eb

[Hay86h] George Hayward. Editorial. Technovation, 4(2):89, April 1986. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497286900015.

Hayward:1986:Ec

[Hay86i] George Hayward. Editorial. Technovation, 4(3):167, June 1986. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649728690009X.

Hayward:1987:BRF

[Hay87a] G. Hayward. Book review: France and her technopoles: by Hilary Sunman, 145 pages, softback £82.00. Technovation, 6(4):304–305, September 1987. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497287900782.

Hayward:1987:BRS

[Hay87b] G. Hayward. Book review: Science and technology parks in Belgium and the Netherlands: by Hilary Sunman, 87 pages, softback £66.00. Technovation, 6(4):304–305, September 1987. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497287900800.

Hayward:1987:BRW

[Hay87c] G. Hayward. Book review: West Germany — innovation centres and science parks: by Hilary Sunman and Julian Lowe, 129 pages, softback £82.00. Technovation, 6(4):304–305, September 1987. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497287900794.

Hayward:1987:BRP

[Hay87d] George Hayward. Book review: Planning and managing change: edited by B. Mayon-White, Harper and Row in association with The Open University, 1986, x + 271 pages, £8.50 softback. Technovation, 6(4):307–308, September 1987. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497287900824.

Hayward:1988:BRI

[Hay88a] G. Hayward. Book review: Innovating to compete: Lessons for diffusing and managing change in the workplace: By Richard E. Walton, Jossey-Bass Publishers (The Jossey-Bass Management Series), 1987, xvii + 361 pages, hardback £27.00. Technovation, 7(4):424, October 1988. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900156.

Hayward:1988:BRMa

[Hay88b] G. Hayward. Book review: Managing professionals in research and development: by Donald Britton Millar, Jossey-Bass Publishers (The Jossey-Bass Management Series), 1986, xxiv + 403 pages, hardback £32.00. Technovation, 7(2):178, May 1988. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497288900478.

Hayward:1988:BRMb

[Hay88c] G. Hayward. Book review: Managing technological innovation: Organizational strategies for implementing advanced manufacturing technologies: By Donald D. Davis and Associates, Jossey-Bass Publishers (The Jossey-Bass Management Series), 1986, xxi + 302 pages, hardback £29.00. Technovation, 7(4):423, October 1988. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900144.

Hayward:1988:BRP

[Hay88d] George Hayward. Book review: Product innovation management: a workbook for management in industry: third edition by Knut Holt, Butterworths, Seven Oaks, 1988, 350 pages, hardback £20.00. Technovation, 8(4):324, December 1988. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497288900351.

Hayward:1988:BRSb

[Hay88e] George Hayward. Book review: Science parks and the growth of high technology firms: C. S. P. Monck, P. R. Quintas, R. B. Porter, D. J. Storey and P. Wynarczyk, Croom Helm published in association with Peat Marwick McLintock, 1988, xv + 270 pages, hardback £35.00. Technovation, 8(4):326, December 1988. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900375.

Hayward:1988:BRSa

[Hay88f] George Hayward. Book review: Soviet automation: Perspectives and prospects: Jack Baranson, Lomond Publications Inc., Mt. Airy, Maryland, USA, 1987, xvi + 142 pages, hardback US$32.50. Technovation, 8(4):325, December 1988. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497288900363.

Hayward:1988:BRR

[Hay88g] George Hayward. Book review: The realities of nuclear power: International economic and regulatory experience: S. D. Thomas, Cambridge University Press, 1988, 289 pages, hardback £35.00. Technovation, 8(4):323, December 1988. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649728890034X.

Hayward:1988:Ea

[Hay88h] George Hayward. Editorial. Technovation, 7(2):85–86, May 1988. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497288900405.

Hayward:1988:Eb

[Hay88i] George Hayward. Editorial. Technovation, 8(4):243–244, December 1988. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497288900259.

Hayward:1989:BRI

[Hay89a] George Hayward. Book review: Innovate through technology: Successful strategies for management: Charles Parker, Kogan Page in association with the British Institute of Management, 1989, Hardback, £14.95, 128 pp. Technovation, 9(7):609–610, November 1989. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900242.

Hayward:1989:E

[Hay89b] George Hayward. Editorial. Technovation, 9(8):621–622, December 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497289900023.

Hayward:1989:NEB

[Hay89c] George Hayward. News of the Editorial Board of Technovation. Technovation, 9(4):308, July 1989. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497287900022.

Hayward:1989:BRR

[Hay89d] Gerorge Hayward. Book review: The role of design in international competitiveness: D. O. Ughanwa and M. J. Baker, Routledge, 1989, Hardback, £45.00, 375 pp. Technovation, 9(7):610–611, November 1989. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900254.

Hayward:1990:BRA

[Hay90a] George Hayward. Book review: Adaptors and innovators: Styles of creativity and problem-solving: edited by Michael J. Kirton, Routledge, London and New York, 1989, Hardback £35.00, xv + 234 pages. Technovation, 10(1):59–60, February 1990. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729090044K.

Hayward:1990:BRI

[Hay90b] George Hayward. Book review: Innovation in the information chain: by J. S. Mackenzie Owen and Johan van Halm, Routledge, London and New York, 1989, Hardback £25.00, 121 pages. Technovation, 10(1):59, February 1990. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729090043J.

Hayward:1990:BRS

[Hay90c] George Hayward. Book review: Small and medium size enterprises and regional development: by Maria Giaoutzi, Peter Nijkamp and David J. Storey, Routledge, London and New York, first published in 1988 and reprinted in 1989, Hardback £45.00, xvii + 314 pages. Technovation, 10(2): 141, March 1990. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729090034H.

Hayward:1990:BRW

[Hay90d] George Hayward. Book review: The world electronics industry: by Daniel Todd, Routledge, London and New York, 1990, Hardback £35.00, 335 pages. Technovation, 10(2):141–142, March 1990. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729090035I.

Hayward:1990:E

[Hay90e] George Hayward. Editorial. Technovation, 10(4):211–212, June 1990. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729090053M.

Hayward:1991:BRH

[Hay91a] George Hayward. Book review: High technology industry and innovative environments: the European experience: edited by Philippe Aydalot and David Keeble, Routledge, London and New York, 1988, Hardback £27.50, 241 pages. Technovation, 11(2):122–123, March 1991. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497291900445.

Hayward:1991:E

[Hay91b] George Hayward. Editorial. Technovation, 11(1):1, January 1991. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729190002L.

Hayward:1992:BRMb

[Hay92a] G. Hayward. Book review: Management of technological change: Yassin Sankar, John Wiley and Sons Inc., £39.80, 374 pages. Technovation, 12(6):419–420, September 1992. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497292900149.

Hayward:1992:BRMa

[Hay92b] G. Hayward. Book review: Managerial and organisational integration: Thomas G. Whiston, Springer-Verlag, 1992, 237 pages. Technovation, 12(6):418–419, September 1992. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497292900138.

Hayward:1992:BRN

[Hay92c] George Hayward. Book review: Nuclear decommissioning and society: edited by Martin J. Pasqualetti, Routledge, London and New York, 1990, ISBN 0-415-03480-9, £45.00 hardback, 237 pages. Technovation, 12(2):138–139, March 1992. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729290022A.

Hayward:1992:BRP

[Hay92d] George Hayward. Book review: Project management in engineering services and development: by Ernst G. Frankel, Butterworth & Co. Ltd, London, 1990, ISBN 0-408-03957-4, £49.50 hardback, 358 pages. Technovation, 12(2):140– 141, March 1992. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729290024C.

Hayward:1992:BRS

[Hay92e] George Hayward. Book review: Science policy in the Soviet Union: by Stephen Fortescue, Routledge, London, 1990, ISBN 0-415-02379-3, £35.00 hardback, 230 pages. Technovation, 12(2):139–140, March 1992. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729290023B.

Hayward:1993:BRP

[Hay93a] G. Hayward. Book review: Project management demystified: by Geoff Reiss. E. & F. N. Spon (an imprint of Chapman and Hall), London, 1992, ISBN 0-419-16920-2 and 0-442-31493-0 (USA), 213 pages, £14.95 (paperback). Technovation, 13(4):261, May 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390023O.

Hayward:1993:BRS

[Hay93b] G. Hayward. Book review: Science parks: an experiment in high technology transfer: Lesley Grayson. The British Library, Science Reference and Information Service, 25 Southampton Buildings, London WC2A 1AW, 1993, vi + 143 pp., ISBN 0-7123-0788-5,£50.00 (paperback). Technovation, 13(7):477, November 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390033R.

Hayward:1993:BRU

[Hay93c] G. Hayward. Book review: UK Science Park Association — Science park directory: 5th edition, edited by Susan Cooke. United Kingdom Science Park Association, Aston Science Park, Love Lane, Aston Triangle, Birmingham B7 4BJ, UK, 1993, 185 pp., ISBN 1-871786-04-5, £50.00 (paperback). Technovation, 13(7):477–478, November 1993. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390034S.

Hayward:1993:BRE

[Hay93d] George Hayward. Book review: European cases in strategic management: edited by John Hendry and Tony Eccles, with Sumantra Ghoshal, Per Jenster and Peter Williamson. Chapman and Hall, London, 1993, ISBN 0-412-48600-8, 431. Technovation, 13(5):331–332, July 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390007I.

Hayward:1994:Ea

[Hay94a] George Hayward. Editorial. Technovation, 14(1):1–2, February 1994. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497294900647.

Hayward:1994:Eb

[Hay94b] George Hayward. Editorial. Technovation, 14(6):353–355, August 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294900140.

Hayward:1995:BRH

[Hay95] George Hayward. Book review: The handbook of industrial innovation: edited by Mark Dodgson and Roy Rothwell. Edward Elgar Publishing Ltd, Aldershot (UK) and Vermont (USA), xii + 453 pp., 1994, ISBN 1-85278-655-8, hardback, £75.00. Technovation, 15(4):265, May 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497295900659.

Hayward:1996:BRR

[Hay96a] G. Hayward. Book review: Recent essentials in innovation management and research: Heinz Hübner and Torsten Dunkel (eds.), Deutscher Universitäts Verlag, 1995, xxi + 269 pp. (softback), ISBN 3-8244-6253-2, SFr 92. Technovation, 16(4):159, April 1996. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497296891231.

Hayward:1996:E

[Hay96b] G. Hayward. Editorial. Technovation, 16(1):iii, January 1996. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497296902382.

Hayward:1997:BRTb

[Hay97a] G. Hayward. Book review: Technology institutes: Strategies for best practice: By Howard Rush, Michael Hobday, John Bessant, Erik Arnold and Robin Murray, International Thomson Business Press, London and Boston, USA, 1996, xiii + 201 pp., £45.00 (hardback), ISBN 0-415-125030. Technovation, 17(11–12):733, November/December 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297887706.

Hayward:1997:BRH

[Hay97b] George Hayward. Book review: A history of the Marconi Company: by W. J. Baker. Routledge, London, reprinted 1996 (first published by Methuen & Co. Ltd, 1970). ISBN 0-415-14624-0, hardback, £50, 414 pp., 23 photographs. Technovation, 17(5):285–286, May 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297882260.

Hayward:1997:BRP

[Hay97c] George Hayward. Book review: The psychology research handbook: a guide for gradate students and research assistants: by Frederick T. L. Leong and James T. Austin, editors. Sage Publications Ltd, London, October 1996. ISBN 0-8039-7048-X, cloth, £49.00; ISBN 0-8039-7049-8, paper, £19.95; xvi + 388 pp. Technovation, 17(6):347, June 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297863389.

Hayward:1997:BRTa

[Hay97d] George Hayward. Book review: The transition to flexibility: edited by Daniel C. Knudsen. Kluwer Academic Publishers, Boston/London/Dordrecht, 1996, Hardbound, ix + 185 pages, ISBN 0-7923-9760-6, Dfl.195.00/US$105.00/ UK£74.75. Technovation, 17(9):535–536, September 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297903369.

Hayward:1998:MII

[Hay98] G. Hayward. Managing innovation: Integrating technological, market and organizational change. Technovation, 18(5): 369–370, ???? 1998. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497298800333.

Hayward:2001:BRI

[Hay01a] G. Hayward. Book review: Innovation policy in a global economy: D. Archibugi, J. Howells and J. Michie (Eds.); Cambridge University Press, 1999, pp 278+xvi, Price £15.95 (US$24.95) paperback, ISBN 0-521-63361-3; £45.00 (US$69.95) hardback, ISBN 0-521-63327-3. Technovation, 21(9):623, September 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000360.

Hayward:2001:BRM

[Hay01b] G. Hayward. Book review: Manufacturing the Future: Strategic Resonance for Enlightened Manufacturing: Steve Brown; Financial Times, Prentice Hall. Published 2000, 207 pp. XI. ISBN 0-273-64322-3 Hardback: £24.99. Technovation, 21(6):397–398, June 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000116.

Hayward:2001:BRT

[Hay01c] G. Hayward. Book review: Trade growth and technical change: Daniele Archibugi and Jonathan Michie (Eds.); Cambridge University Press, 1998, pp. 276 + xvi, Price £15.95 (US$22.95) paperback, ISBN 0-521-55641-4. £45.00 (US$64.95) hardback, ISBN 0-521-55393-8. Technovation, 21(9):623–624, September 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000372.

Hayward:2005:E

[Hay05] G. Hayward. Editorial. Technovation, 25(12):1365, December 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205001434.

Harvey:1984:CSG

[HB84] R. C. Harvey and E. Burley. Case studies in geotechnics: Some expedients for technical innovation. Technovation, 2(2):89–100, May 1984. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497284900142.

Hochstrasser:1989:PIO

[HB89] Urs Hochstrasser and Sergio Barabaschi. Project implementation: Organizational, financial and institutional arrangements. Technovation, 9(2–3):271–274, June 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900679.

Halachmi:1997:PRP

[HB97] Arie Halachmi and Tony Bovaird. Process reengineering in the public sector: Learning some private sector lessons. Technovation, 17(5):227–235, May 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649729600123X.

Hipkin:2003:MPF

[HB03] Ian Hipkin and David Bennett. Managerial perceptions of factors influencing technology management in South Africa. Technovation, 23(9):719–735, September 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000317.

Hobday:2012:PDP

[HBG12] Mike Hobday, Anne Boddington, and Andrew Grantham. Policies for design and policies for innovation: Contrasting perspectives and remaining challenges. Technovation, 32(5): 272–281, May 2012. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497211001696.

Hu:2021:IIC

[HBH21] Feng Hu, Tammo H. A. Bijmolt, and Eelko K. R. E. Huizingh. The impact of innovation contest briefs on the quality of solvers and solutions. Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497217308933.

Hall:2014:ISH

[HBM14] Jeremy Hall, Vernon Bachor, and Stelvia Matos. The impact of stakeholder heterogeneity on risk perceptions in technological innovation. Technovation, 34(8):410–419, August 2014. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649721300148X.

Hemmert:2014:BCD

[HBO14] Martin Hemmert, Ludwig Bstieler, and Hiroyuki Okamuro. Bridging the cultural divide: Trust formation in university– industry research collaborations in the US, Japan, and South Korea. Technovation, 34(10):605–616, October 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214000510.

Hall:2002:SRI

[HBS02] Linda A. Hall and Sharmistha Bagchi-Sen. A study of R&D, innovation, and business performance in the Canadian biotechnology industry. Technovation, 22(4):231–244, April 2002. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201000165.

Hall:2007:AFL

[HBS07] Linda A. Hall and Sharmistha Bagchi-Sen. An analysis of firm-level innovation strategies in the US biotechnology industry. Technovation, 27(1–2):4–14, January/ February 2007. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497206000757.

Heusinkveld:2009:MSN

[HBvdB09] Stefan Heusinkveld, Jos Benders, and Robert-Jan van den Berg. From market sensing to new concept development in consultancies: the role of information processing and organizational capabilities. Technovation, 29(8):509–516, August 2009. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497209000418.

Hsu:2001:GSU

[HC01] Chiung-Wen Hsu and Hsueh-Chiao Chiang. The government strategy for the upgrading of industrial technology in Taiwan. Technovation, 21(2):123–132, February 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000298.

Hung:2006:SNI

[HC06] Shih-Chang Hung and Yee-Yeen Chu. Stimulating new industries from emerging technologies: challenges for the public sector. Technovation, 26(1):104–110, January 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001324.

Huang:2010:ITD

[HC10] Yi-Fen Huang and Chung-Jen Chen. The impact of technological diversity and organizational slack on innovation. Technovation, 30(7–8):420–428, July/August 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210000210.

Hung:2013:IOI

[HC13] Kuang-Peng Hung and Christine Chou. The impact of open innovation on firm performance: the moderating effects of internal R&D and environmental turbulence. Technovation, 33(10–11):368–380, October/November 2013. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213000680.

Hobday:2001:GTE

[HCK01] Mike Hobday, Alan Cawson, and S. Ran Kim. Governance of technology in the electronics industries of East and South-East Asia. Technovation, 21(4):209–226, April 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000389.

Huang:2009:RSS

[HCL09] Yu-An Huang, Hsien-Jui Chung, and Chad Lin. R&D sourcing strategies: Determinants and consequences. Technovation, 29(3):155–169, March 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497208001089.

Hernandez-Chea:2021:MBI

[HCMMH21] Roberto Hernández-Chea, Maral Mahdad, Thai Thi Minh, and Carsten Nico Hjortsø. Moving beyond intermediation: How intermediary organizations shape collaboration dynamics in entrepreneurial ecosystems. Technovation, 108(??):??, December 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497221001139.

Huizingh:2011:ISI

[HCT11] Eelko Huizingh, Steffen Conn, and Marko Torkkeli. ISPIM special issue on open innovation. Technovation, 31(1):1, January 2011. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497210001124.

Hutton:2021:HDO

[HDE21] Steven Hutton, Robert Demir, and Stephen Eldridge. How does open innovation contribute to the firm’s dynamic capabilities? Technovation, 106(??):??, August 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000699.

Harms:2021:BHT

[HdWN21] Rainer Harms and Petra de Weerd-Nederhof. Building high tech small firms to become — news from the High Technology Small Firms conference, University Twente, 27–28 May 2019. Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649721930063X.

Healy:1983:IIT

[Hea83] Mel Healy. Innovative Ireland — technological, industrial and societal challenges. Technovation, 2(1):45–53, February 1983. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497283900068.

Hilliard:2022:PSM

[HEC22] Rachel Hilliard, James English, and Maébh Coleman. Prosocially motivated knowledge hiding in innovation teams. Technovation, 116(??):??, August 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://

www.sciencedirect.com/science/article/pii/S0166497222000608.

Hedges:1989:ORA

[Hed89] R. E. M. Hedges. The Oxford Radiocarbon Accelerator Unit. Technovation, 9(1):109–112, May 1989. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900448.

Helper:1994:TSFa

[Hel94a] Susan R. Helper. Three steps forward, two steps back in automotive supplier relations. Technovation, 14(10):633– 640, December 1994. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497294901104.

Helper:1994:TSFb

[Hel94b] Susan R. Helper. Three steps forward, two steps back in automotive supplier relations. Technovation, 14(10):693, December 1994. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497294901155.

Henriksen:2001:KME

[Hen01] Lars Bo Henriksen. Knowledge management and engineering practices: the case of knowledge management, problem solving and engineering practices. Technovation, 21 (9):595–603, September 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497201000232.

Heron:1991:BRS

[Her91] Thomas Heron. Book review: Sharpbenders: the secrets of unleashing corporate potential: by P. H. Grinyer, D. G. Mayes and P. McKiernan, Basil Blackwell, Oxford, 1988, ISBN 0-631-15304-7, £25.00, 287 pages. Technovation, 11(2):124–125, March 1991. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497291900467.

Heron:1992:BRMa

[Her92a] Thomas Heron. Book review: Managing and marketing services in the 1990s: by R. Teare, L. Moutinho and N. Morgan, Cassell Educational Ltd, London, 1990, ISBN0-304-31817-7, £25.00, 255 pages. Technovation, 12(2):142, March 1992. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729290026E.

Heron:1992:BRMb

[Her92b] Thomas Heron. Book review: Marketing management and information technology: K. Fletcher, Prentice Hall, Hemel Hempstead, 1990, ISBN 0-13-556507-3, £15.95, 354 pages. Technovation, 12(6):415, September 1992. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497292900075.

Heron:1992:BRP

[Her92c] Thomas Heron. Book review: Perspectives on marketing management, volume 1: M. J. Baker, John Wiley, Chichester, 1991, ISBN 0-471-92858-5, £50.00, 285 pages. Technovation, 12(6):417–418, September 1992. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497292900127.

Hertin:2004:BRN

[Her04] Julia Hertin. Book review: New Technologies and Environmental Innovation: Joseph Huber, Edward Elgar, Cheltenham, 2004, 384 pages, hardbound, ISBN 1-84376-799-6, £69.95. Technovation, 24(12):1011, December 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001361.

Hernandez-Espallardo:2011:EEB

[HESPSL11] Miguel Hernández-Espallardo, Manuel Sánchez-Pérez, and Cristina Segovia-López. Exploitation- and explorationbased innovations: the role of knowledge in inter-firm relationships with distributors. Technovation, 31(5–6):203–215, May/June 2011. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497211000216.

Hewitt:1999:ITP

[Hew99] Tom Hewitt. Institutional tensions and private sector promotion in Tanzania: whose agenda? Technovation, 19(6–7): 383–391, June 1999. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497299000267.

Hao:2018:LLF

[HF18] Bin Hao and Yanan Feng. Leveraging learning forces in asymmetric alliances: Small firms’ perceived power imbalance in driving exploration and exploitation. Technovation, 78(??):27–39, December 2018. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497216303844.

Hong:2016:DGG

[HFWW16] Jin Hong, Bing Feng, Yanrui Wu, and Liangbing Wang. Do government grants promote innovation efficiency in China’s high-tech industries? Technovation, 57–58(??):4– 13, November/December 2016. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497216301018.

Hinze:1992:ARI

[HG92] Sybille Hinze and Hariolf Grupp. Applied research and industrial development in East Germany: International comparison by performance indicators. Technovation, 12(4): 257–278, May 1992. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729290046K.

Haessler:2023:WHC

[HGB23] Philipp Haessler, Ferran Giones, and Alexander Brem. The who and how of commercializing emerging technologies: a technology-focused review. Technovation, 121(??):??, March 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497222001870.

Hussain:1998:MAS

[HGL98] Md. Mostaque Hussain, A. Gunasekaran, and Erkki K. Laitinen. Management accounting systems in Finnish service firms. Technovation, 18(1):57–67, January 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649729700062X.

Hang:2015:OD

[HGR15] Chang Chieh Hang, Elizabeth Garnsey, and Yi Ruan. Opportunities for disruption. Technovation, 39–40(??):83–93, May/June 2015. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497214001734.

Hayward:1989:TTC

[HH89] George Hayward and Joan Hayward. Technics over time: a comment. Technovation, 9(1):5–6, May 1989. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900369.

Hu:2008:MIM

[HH08] Jin-Li Hu and Yu-Hsueh Hsu. The more interactive, the more innovative? A case study of South Korean cellular phone manufacturers. Technovation, 28(1–2):75–87, January/February 2008. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649720700106X.

Hull:1984:SIP

[HHA84] Frank Hull, Jerald Hage, and Koya Azumi. Strategies for innovation and productivity in Japan and America. Technovation, 2(2):121–139, May 1984. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497284900166.

Hsu:2009:EGS

[HHH09] Fang-Ming Hsu, Der-Juinn Horng, and Chao-Chih Hsueh. The effect of government-sponsored R&D programmes on additionality in recipient firms in Taiwan. Technovation, 29(3):204–217, March 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497208000564.

Hansson:2005:SGS

[HHV05] Finn Hansson, Kenneth Husted, and Jakob Vestergaard. Second generation science parks: from structural holes jockeys to social capital catalysts of the knowledge society. Technovation, 25(9):1039–1049, September 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204000537.

Hitomi:1985:JWM

[Hit85] Katsundo Hitomi. The Japanese way of manufacturing and production management. Technovation, 3(1):49–55, February 1985. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497285900367.

Hitomi:1989:BRD

[Hit89a] Katsundo Hitomi. Brief review of the debate on technics in Japan and introduction to Dr. Ishigai’s concept. Technovation, 9(1):7–8, May 1989. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900370.

Hitomi:1989:NMM

[Hit89b] Katsundo Hitomi. Non-mass, multi-product, small-sized production: the state of the art. Technovation, 9(4):357– 369, July 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649728790006X.

Hitomi:1992:GE

[Hit92a] Katsundo Hitomi. Guest editorial. Technovation, 12(8): 497–498, December 1992. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729290071O.

Hitomi:1992:PTI

[Hit92b] Katsundo Hitomi. Present trends and issues in Japanese manufacturing and management. Technovation, 12(3):177– 189, April 1992. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729290034F.

Hitomi:1993:EME

[Hit93] Katsundo Hitomi. The emphasis of management and economic concepts for mechanical engineers: a historical reassessment of Towne’s article published 100 years ago. Technovation, 13(4):235–239, May 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390020V.

Hitomi:1994:ACS

[Hit94] Katsundo Hitomi. Automation — its concept and a short history. Technovation, 14(2):121–128, March 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294901015. See [Ano94c].

Hitomi:1996:MEC

[Hit96] Katsundo Hitomi. Manufacturing excellence for 21st century production. Technovation, 16(1):33–41, January 1996.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497295000186.

Hitomi:2002:HTP

[Hit02] Katsundo Hitomi. Historical trends and the present state of Korean industry and manufacturing. Technovation, 22(7): 453–462, July 2002. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497201000384.

Hitomi:2003:HTP

[Hit03] Katsundo Hitomi. Historical trends and the present state of Chinese industry and manufacturing. Technovation, 23(7): 633–641, July 2003. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497202000123.

Hitomi:2004:EAJ

[Hit04] K. Hitomi. Efficiency analysis of Japan’s industry and manufacturing. Technovation, 24(9):741–748, September 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202001517.

Hitomi:2005:HTP

[Hit05] Katsundo Hitomi. Historical trends and the present state of the US industry and manufacturing. Technovation, 25(6): 673–681, June 2005. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203002050.

Hitomi:2006:AIN

[Hit06] Katsundo Hitomi. An analysis of industrial networks for regional-independent production. Technovation, 26(9): 1077–1080, September 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497205001720.

Helander:2002:RPD

[HJ02] Martin G. Helander and Jianxin Jiao. Research on E-product development (ePD) for mass customization.

Technovation, 22(11):717–724, November 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000748.

Hjalager:2009:ITM

[Hja09] Anne-Mette Hjalager. Innovations in travel medicine and the progress of tourism-Selected narratives. Technovation, 29(9):596–601, September 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497209000777.

Hernandez:2008:ETA

[HJM08] Blanca Hernández, Julio Jiménez, and M. José Martín. Extending the technology acceptance model to include the IT decision-maker: a study of business management software. Technovation, 28(3):112–121, March 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207001484.

Hytonen:2012:MST

[HJSYJ12] Henri Hytönen, Toni Jarimo, Ahti Salo, and Erkki Yli-Juuti. Markets for standardized technologies: Patent licensing with principle of proportionality. Technovation, 32(9–10):523–535, September/October 2012. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211001210.

Hyclak:1986:WII

[HK86] Thomas J. Hyclak and Michael G. Kolchin. Worker involvement in implementing new technology. Technovation, 4(2): 143–151, April 1986. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497286900052.

Harris:2002:SDI

[HK02] Roger Harris and Anshuman Khare. Sustainable development issues and strategies for Alberta’s oil industry. Technovation, 22(9):571–583, September 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720100058X.

Hurmelinna:2007:JFA

[HKJ07] Pia Hurmelinna, Kalevi Kyläheiko, and Tiina Jauhiainen. The Janus face of the appropriability regime in the protection of innovations: Theoretical re-appraisal and empirical analysis. Technovation, 27(3):133–144, March 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720500129X.

Hyvarinen:2024:IDI

[HKL24] Anne M. J. Hyvärinen, Marko Keskinen, and Jarkko Levänen. Intermediaries as drivers of innovation development in resource-constrained environments: Insights from the Kenyan water sector. Technovation, 131(??):??, March 2024. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497224000063.

Huikkola:2022:OCS

[HKR^{+}22] Tuomas Huikkola, Marko Kohtamäki, Rodrigo Rabetino, Hannu Makkonen, and Philipp Holtkamp. Overcoming the challenges of smart solution development: Co-alignment of processes, routines, and practices to manage product, service, and software integration. Technovation, 118 (??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497221001632.

Habib:2020:RRM

[HKRR20] Tufail Habib, Jimmi Normann Kristiansen, Mohammad Bakhtiar Rana, and Paavo Ritala. Revisiting the role of modular innovation in technological radicalness and architectural change of products: the case of Tesla X and Roomba. Technovation, 98(??):??, December 2020. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497218308162.

Hills:2013:LMA

[HKW^{+}13] Gerald E. Hills, Judith J. Kirchhoff, Steven T. Walsh, Zoltan J. Acs, Robert H. Brockhaus, Thomas C. Dandridge, William J. Dennis, Jr., Patricia G. Greene, Bruce D. Phillips, and George Solomon. Lifetime model of achievement in academia: Reflections on Bruce A. Kirchhoff’s legacy. Technovation, 33(8–9):249–254, August/September 2013. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497213000710.

Hong:2022:SIE

[HKW^{+}22] Suckwon Hong, Juram Kim, Han-Gyun Woo, Young-Choon Kim, and Changyong Lee. Screening ideas in the early stages of technology development: a word2vec and convolutional neural network approach. Technovation, 112(??): ??, April 2022. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497221001887.

Huikkola:2022:BSS

[HKY22] Tuomas Huikkola, Marko Kohtamäki, and Juho Ylimäki. Becoming a smart solution provider: Reconfiguring a product manufacturer’s strategic capabilities and processes to facilitate business model innovation. Technovation, 118 (??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222000451.

Liao:2005:TMM

[hL05] Shu hsien Liao. Technology management methodologies and applications: a literature review from 1995 to 2003. Technovation, 25(4):381–393, April 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720300155X.

Hu:2022:GNF

[HL22] Yefei Hu and Dayong Liu. Government as a nonfinancial participant in innovation: How standardization led by government promotes regional innovation performance in China. Technovation, 114(??):??, June 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000712.

Henry:2024:EIA

[HL24] Colette Henry and Katja Lahikainen. Exploring intrapreneurial activities in the context of the entrepreneurial university: an analysis of five EU HEIs. Technovation, 129(??):??, January 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497223002043.

Hu:2005:TBR

[HLC05] Tai-Shan Hu, Chien-Yuan Lin, and Su-Li Chang. Technologybased regional development strategies and the emergence of technological communities: a case study of HSIP, Taiwan. Technovation, 25(4):367–380, April 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203001676.

Liao:2008:RBK

[hLFL08] Shu hsien Liao, Wu-Chen Fei, and Chih-Tang Liu. Relationships between knowledge inertia, organizational learning and organization innovation. Technovation, 28(4):183– 195, April 2008. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497207001502.

Hsieh:2012:SPV

[HLH12] Pi-Feng Hsieh, Chung-Shing Lee, and Jonathan C. Ho. Strategy and process of value creation and appropriation in service clusters. Technovation, 32(7–8):430–439, July/ August 2012. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497211000538.

Harvey:1993:TCV

[HLL93] Jean Harvey, Élisabeth Lefebvre, and Louis A. Lefebvre. Technology and the creation of value in services: a conceptual model. Technovation, 13(8):481–495, December 1993. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497293900643. See [Ano93o].

Huang:2012:DFO

[HLL12] Hao-Chen Huang, Mei-Chi Lai, and Kuo-Wei Lo. Do founders’ own resources matter? The influence of business networks on start-up innovation and performance. Technovation, 32(5):316–327, May 2012. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211001854.

Hurmelinna-Laukkanen:2021:OLU

[HLNP21] Pia Hurmelinna-Laukkanen, Satu Nätti, and Minna Pikkarainen. Orchestrating for lead user involvement in innovation networks. Technovation, 108(??):??, December 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221001073.

Hurmelinna-Laukkanen:2014:CRA

[HLO14] Pia Hurmelinna-Laukkanen and Heidi Olander. Coping with rivals’ absorptive capacity in innovation activities. Technovation, 34(1):3–11, January 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213000795.

Liao:2004:ITR

[hLwChLbL04] Shu hsien Liao, Yea wen Chem, Feng hao Liu, and Wei bang Liao. Information technology and relationship management: a case study of Taiwan’s small manufacturing firm. Technovation, 24(2):97–108, February 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000378.

Hu:2011:MIR

[HM11] Yansong Hu and Damien McLoughlin. Modifications and innovations to research solutions in the lab: a study on the post-adoptive usage behaviors of life scientists. Technovation, 31(10–11):539–553, October/November 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211000708.

Holland:2024:IIC

[HMFS24] Claire Holland, Adam McCarthy, Priscila Ferri, and Philip Shapira. Innovation intermediaries at the convergence of digital technologies, sustainability, and governance: a case study of AI-enabled engineering biology. Technovation, 129(??):??, January 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001864.

Hall:2014:IPB

[HMM14] Jeremy Hall, Stelvia V. Matos, and Michael J. C. Martin. Innovation pathways at the Base of the Pyramid: Establishing technological legitimacy through social attributes. Technovation, 34(5–6):284–294, May/June 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213001491.

Hu:2015:OOI

[HMM15] Yansong Hu, Peter McNamara, and Damien McLoughlin. Outbound open innovation in bio-pharmaceutical outlicensing. Technovation, 35(??):46–58, January 2015. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214000996.

Han:2023:ITE

[HMWT23] Myat Su Han, Shuang (Sara) Ma, Yonggui Wang, and Qinghong Tian. Impact of technology-enabled product ecoinnovation: Empirical evidence from the Chinese manufacturing industry. Technovation, 128(??):??, December 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001645.

Haug:1992:TIRa

[HN92a] Peter Haug and Philip Ness. Technological infrastructure and regional economic development of biotechnology firms. Technovation, 12(7):423–432, November 1992. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729290048M.

Haug:1992:TIRb

[HN92b] Peter Haug and Philip Ness. Technological infrastructure and regional economic development of biotechnology firms. Technovation, 12(7):487–488, November 1992. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729290053K.

Hameri:1999:TCN

[HN99] Ari-Pekka Hameri and Markus Nordberg. Tendering and contracting of new, emerging technologies. Technovation, 19(8):457–465, August 1999. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497299000358.

Herrera:2008:NIP

[HN08] Liliana Herrera and Mariano Nieto. The national innovation policy effect according to firm location. Technovation, 28(8):540–550, August 2008. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497208000230.

Hofmann:2005:AMT

[HO05] Christian Hofmann and Stuart Orr. Advanced manufacturing technology adoption — the German experience. Technovation, 25(7):711–724, July 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203002189.

Ho:2009:CDT

[Ho09] Yuh-Shan Ho. Comments on “Determining technology trends and forecasts of RFID by a historical review and bibliometric analysis from 1991 to 2005”. Technovation, 29 (10):725–727, October 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497209000029.

See [CYJ07].

Hernandez-Ortega:2011:RPU

[HO11] Blanca Hernández-Ortega. The role of post-use trust in the acceptance of a technology: Drivers and consequences. Technovation, 31(10–11):523–538, October/ November 2011. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497211000939.

Hansen:2014:LTC

[HO14] Ulrich Elmer Hansen and David Ockwell. Learning and technological capability building in emerging economies: the case of the biomass power equipment industry in Malaysia. Technovation, 34(10):617–630, October 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214000984.

Hobday:1987:BRS

[Hob87] Mike Hobday. Book review: Science and technology indicators for development: edited by Hiroko MoritaLou, United Nations Science and Technology for Development Series, Westview Press, Boulder, CO, 1985. Technovation, 6(2): 148–149, June 1987. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497287900174.

Hobday:1988:ECR

[Hob88] Michael Hobday. Evaluating collaborative R&D programmes in information technology: the case of the U.K. Alvey programme. Technovation, 8(4):271–298, December 1988. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497288900302.

Hobday:1990:SUK

[Hob90] Michael Hobday. Strategies for the U.K. semiconductor industry: lessons from the Alvey Program. Technovation, 10(3):193–210, May 1990. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729090024E.

Hobday:1992:EEI

[Hob92] Mike Hobday. The European electronics industry: technology and structural change. Technovation, 12(2):75–97, March 1992. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729290017C.

Hobday:1995:IEA

[Hob95] Mike Hobday. Innovation in East Asia: Diversity and development. Technovation, 15(2):55–63, March 1995. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497295966106.

Hobday:1997:BRP

[Hob97] Mike Hobday. Book review: Programme management demystified: Managing multiple projects successfully: by Geoff Reiss. E and FN Spon, Chapman and Hall, London, 1996. ISBN 0-419-21350-3, 261 pp. Technovation, 17(7):415–416, July 1997. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297868083.

Hobday:2003:BRR

[Hob03] Mike Hobday. Book review: Review of Managing Intellectual Capital: David Teece, Oxford University Press, 300 pages, ISBN 0-19-829542-1. Technovation, 23(2):183, February 2003. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497202001190.

Hochstrasser:1989:UH

[Hoc89] Urs Hochstrasser. Urs Hochstrasser. Technovation, 9(2–3): 301–302, June 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649728990076X.

Hoelscher:1985:TIQ

[Hoe85] H. E. Hoelscher. Technology, industry, and the question of national policy. Technovation, 3(3):161–166, August 1985. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497285900112.

Hervas-Oliver:2011:MSI

[HOGGP11] Jose-Luis Hervas-Oliver, Jose Albors Garrigos, and Ignácio Gil-Pechuan. Making sense of innovation by R&D and non–R&D innovators in low technology contexts: a forgotten lesson for policymakers. Technovation, 31(9):427–446, September 2011. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497211000897.

Holt:1985:UOP

[Hol85] Knut Holt. User-oriented product innovation — Some research findings. Technovation, 3(3):199–208, August 1985. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497285900148.

Holt:1987:UOP

[Hol87] Knut Holt. User-oriented product innovation in theory and practice. Technovation, 6(1):37–45, April 1987. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649728790037X.

Holt:1988:MOP

[Hol88a] Knut Holt. Market-oriented product innovation at Høyang Polaris and Jøtul. Technovation, 8(4):249–254, December 1988. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497288900284.

Holt:1988:RUP

[Hol88b] Knut Holt. The role of the user in product innovation. Technovation, 7(3):249–258, July 1988. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900235.

Holt:1990:CIT

[Hol90] Knut Holt. Consulting in innovation through intercompany study groups. Technovation, 10(5):347–354, July 1990. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729090018F.

Holtham:1993:BRC

[Hol93] Clive Holtham. Book review: Cyberspace: First steps: edited by Michael Benedikt. MIT Press, Cambridge, MA, 1992, ISBN 0-262-02327-X, £19.95 (hardback), £13.95 (paperback). Technovation, 13(4):261–262, May 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390024P.

Holt:1994:UFDa

[Hol94a] Knut Holt. The user in focus — the development of a market-oriented corporate culture. Technovation, 14(9): 573–584, November 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729490040X.

See [Hol94b].

Holt:1994:UFDb

[Hol94b] Knut Holt. The user in focus — the development of a market-oriented corporate culture: Technovation, 14(9) (1994) 573–584. Technovation, 14(9):621–622, November 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294900442. See [Hol94a].

Holt:1999:MOT

[Hol99] Knut Holt. Management and organization through 100 years. Technovation, 19(3):135–140, January 1999. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298001138.

Honjo:2021:IFH

[Hon21] Yuji Honjo. The impact of founders’ human capital on initial capital structure: Evidence from Japan. Technovation, 100(??):??, February 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497220300638.

Hopkins:2004:BRL

[Hop04] Michael Hopkins. Book review: Leading Pharmaceutical Innovation: Trends and Drivers for Growth in the Pharmaceutical Industry: Oliver Gassmann, Gerrit Reepmeyer, Maximilian von Zedtwitz; Springer-Verlag, Berlin, 2004, 178 pages, hardcover, ISBN 3-540-40717-

0. Technovation, 24(12):1012–1013, December 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001427.

Hossain:1999:FED

[Hos99] Liaquat Hossain. Fostering economic development through telecommunications in South East Asia: challenges facing the Thai telecommunications regulators. Technovation, 19 (11):681–689, November 1999. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016649729900067X.

Hossain:2000:IDP

[Hos00] Liaquat Hossain. Infrastructure development planning in SEA: an analysis of the national telecommunications planning practice in Thailand. Technovation, 20(5):275–285, May 2000. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497299001339.

Hossain:2001:MCT

[Hos01] Liaquat Hossain. Management challenges for the telecom operators: a framework for understanding the telecommunications strategic planning process for the case of the Telephone Organisation of Thailand (1954– 1996). Technovation, 21(7):449–461, July 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000572.

Hossain:2003:FSN

[Hos03] L. Hossain. Is a formalised structure a necessary prerequisite for implementing national telecommunications plan in developing and developed economies? Technovation, 23(1):39–49, January 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497201000840.

Howarth:1994:RSA

[How94] Cathy S. Howarth. The role of strategic alliances in the development of technology. Technovation, 14(4):243–257, May 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294900094.

Howells:2024:IID

[How24] Jeremy Howells. Innovation intermediaries in a digital paradigm: a theoretical perspective. Technovation, 129(??): ??, January 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223002006.

Hoye:2009:RCH

[HP09] Kate Hoye and Fred Pries. ‘Repeat commercializers,’ the ‘habitual entrepreneurs’ of university–industry technology transfer. Technovation, 29(10):682–689, October 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209000832.

Haque:2003:ABP

[HPB03] B. Haque, K. S. Pawar, and R. J. Barson. The application of business process modelling to organisational analysis of concurrent engineering environments. Technovation, 23(2):147–162, February 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201001031.

Hoffman:1998:SFR

[HPBP98] Kurt Hoffman, Milady Parejo, John Bessant, and Lew Perren. Small firms, R&D, technology and innovation in the UK: a literature review. Technovation, 18(1):39–55, January 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297001028.

Hack-Polay:2023:SRN

[HPMI^{+}23] Dieu Hack-Polay, Ali B. Mahmoud, Irene Ikafa, Mahfuzur Rahman, Maria Kordowicz, and Juan Manuel Verde. Steering resilience in nursing practice: Examining the impact of digital innovations and enhanced emotional training on nurse competencies. Technovation, 120(??):??, February 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497222000967.

Hussler:2010:TIT

[HPT10] Caroline Hussler, Fabienne Picard, and Ming Feng Tang. Taking the ivory from the tower to coat the economic world: Regional strategies to make science useful. Technovation, 30(9–10):508–518, September/October 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210000684.

Hansen:1998:HCPa

[HR98a] Karen Lee Hansen and Howard Rush. Hotspots in complex product systems: emerging issues in innovation management. Technovation, 18(8–9):555–561, 589–590, August/ September 1998. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497298000273.

Hansen:1998:HCPb

[HR98b] Karen Lee Hansen and Howard Rush. Hotspots in complex product systems: emerging issues in innovation management. Technovation, 18(8–9):589–590, August/September 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497298901262.

Herring:2007:TIE

[HR07] Horace Herring and Robin Roy. Technological innovation, energy efficient design and the rebound effect. Technovation, 27(4):194–203, April 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720600112X.

Ho:2016:TUS

[HRHW16] Yuen-Ping Ho, Yi Ruan, Chang-Chieh Hang, and Poh-Kam Wong. Technology upgrading of small-and-medium-sized enterprises (SMEs) through a manpower secondment strategy — a mixed-methods study of Singapore’s T-Up program. Technovation, 57–58(??):21–29, November/ December 2016. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497216301602.

Hagedoorn:1991:IFP

[HS91] John Hagedoorn and Jos Schakenraad. Inter-firm partnerships for generic technologies — the case of new materials. Technovation, 11(7):429–444, November 1991. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729190024X.

Harris:2002:MCA

[HS02] Andrew Harris and Amrik S. Sohal. Managing change in an aluminium can manufacturing plant: a case study. Technovation, 22(10):615–623, October 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000633.

Han:2014:SWI

[HS14] Kyunam Han and Juneseuk Shin. A systematic way of identifying and forecasting technological reverse salients using QFD, bibliometrics, and trend impact analysis: a carbon nanotube biosensor case. Technovation, 34(9):559–570, September 2014. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497214000704.

Huang:2019:IFT

[HS19] Hung-Chun Huang and Hsin-Ning Su. The innovative fulcrums of technological interdisciplinarity: an analysis of technology fields in patents. Technovation, 84–85 (??):59–70, June/July 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497217307149.

Honig:2021:BPI

[HS21a] Benson Honig and Mikael Samuelsson. Business planning by intrapreneurs and entrepreneurs under environmental uncertainty and institutional pressure. Technovation, ??(??):

??, ???? 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497218307892.

Hou:2021:ESE

[HS21b] Hong Hou and Yongjiang Shi. Ecosystem-as-structure and ecosystem-as-coevolution: a constructive examination. Technovation, 100(??):??, February 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497220300651.

Han:2022:HDB

[HSLL22] Shaojie Han, Jingqin Su, Yibo Lyu, and Qing Liu. How do business incubators govern incubation relationships with different new ventures? Technovation, 116(??):??, August 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000335.

Hegeman:2024:CAN

[HSLS24] Puck D. Hegeman, Roger Sørheim, Hans Landström, and Erik Andreas Saether. Corporate angels: a new breed of investors in the funding landscape for technology-based start-ups. Technovation, 134(??):??, June 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722400066X.

Hsu:2005:PTF

[HST05] Yeou-Geng Hsu, Joseph Z. Shyu, and Gwo-Hshing Tzeng. Policy tools on the formation of new biotechnology firms in Taiwan. Technovation, 25(3):281–292, March 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000786.

Huang:2007:RIP

[HST07] Chi-Yo Huang, Joseph Z. Shyu, and Gwo-Hshiung Tzeng. Reconfiguring the innovation policy portfolios for Taiwan’s SIP Mall industry. Technovation, 27(12):744–765, December 2007. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649720700048X.

Holt:2021:ART

[HST21] Jared Holt, Ahmed Skali, and Russell Thomson. The additionality of R&D tax policy: Quasi-experimental evidence. Technovation, 107(??):??, September 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000742.

Hsu:2005:FII

[Hsu05] Chiung-Wen Hsu. Formation of industrial innovation mechanisms through the research institute. Technovation, 25 (11):1317–1329, November 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016649720400094X.

Hoecht:2006:IRS

[HT06] A. Hoecht and P. Trott. Innovation risks of strategic outsourcing. Technovation, 26(5–6):672–681, May/June 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000507.

Hung:2008:FAC

[HT08] Shiu-Wan Hung and Ruei-Hung Tang. Factors affecting the choice of technology acquisition mode: an empirical analysis of the electronic firms of Japan, Korea and Taiwan. Technovation, 28(9):551–563, September 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720700137X.

Hsieh:2012:OVC

[HT12] Kuo-Nan Hsieh and Joe Tidd. Open versus closed new service development: the influences of project novelty. Technovation, 32(11):600–608, November 2012. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497212000788.

Hu:2013:HAG

[Hu13] Yansong Hu. Hyperlinked actors in the global knowledge communities and diffusion of innovation tools in nascent industrial field. Technovation, 33(2–3):38–49, February/ March 2013. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497212001046.

Huang:2009:KSG

[Hua09] Chi-Cheng Huang. Knowledge sharing and group cohesiveness on performance: an empirical study of technology R&D teams in Taiwan. Technovation, 29(11):786–797, November 2009. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497209000480.

Hubner:1996:DID

[Hüb96] Heinz Hübner. Decisions on innovation and diffusion and the limits of deregulation. Technovation, 16(7):327–339, July 1996. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497296000223.

Huczynski:1985:DHC

[Huc85] Andrzej A. Huczynski. Designing high commitment — high performance organizations. Technovation, 3(2):111– 118, June 1985. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497285900045.

Huizingh:2011:OIS

[Hui11] Eelko K. R. E. Huizingh. Open innovation: State of the art and future perspectives. Technovation, 31(1):2–9, January 2011. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497210001100.

Huizingh:2017:MIH

[Hui17] Eelko K. R. E. Huizingh. Moving the innovation horizon in Asia. Technovation, 60–61(??):43–44, February 2017. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497217300809.

Hung:2000:SCI

[Hun00] Shih-Chang Hung. Social construction of industrial advantage. Technovation, 20(4):197–203, April 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001273.

Hussinger:2010:ITR

[Hus10] Katrin Hussinger. On the importance of technological relatedness: SMEs versus large acquisition targets. Technovation, 30(1):57–64, January 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209001102.

Hameri:1996:UBR

[HV96] Ari-Pekka Hameri and Olli Vuola. Using basic research as a catalyst to exploit new technology-based innovations — a case study. Technovation, 16(10):531–539, October 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297828933.

Hill:1997:LTI

[HW97] Terry Hill and Roy Westbrook. Linking technological innovations to strategic needs. Technovation, 17(3):109–117, March 1997. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497296000922.

Hur:2001:UTS

[HW01] Kwang In Hur and Chihiro Watanabe. Unintentional technology spillover between two sectors: kinetic approach. Technovation, 21(4):227–235, April 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000377.

Hur:2002:DPT

[HW02] Kwang In Hur and Chihiro Watanabe. Dynamic process of technology spillover; a transfer function approach. Technovation, 22(7):437–444, July 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000268.

Holley:2017:AEB

[HW17] Alicia Castillo Holley and John Watson. Academic entrepreneurial behavior: Birds of more than one feather. Technovation, 64–65(??):50–57, June 2017. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http:

//www.sciencedirect.com/science/article/pii/S0166497217304819.

Hertenstein:2018:RSE

[HW18] Peter Hertenstein and Peter J. Williamson. The role of suppliers in enabling differing innovation strategies of competing multinationals from emerging and advanced economies: German and Chinese automotive firms compared. Technovation, 70–71(??):46–58, ???? 2018. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497218301251.

Hobo:2006:DST

[HWC06] Masayo Hobo, Chihiro Watanabe, and Chaojung Chen. Double spiral trajectory between retail, manufacturing and customers leads a way to service oriented manufacturing. Technovation, 26(7):873–890, July 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000891.

Wang:2008:EFT

[hWyLbC08] Chun hsien Wang, Iuan yuan Lu, and Chie bein Chen. Evaluating firm technological innovation capability under uncertainty. Technovation, 28(6):349–363, June 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207001393.

Hahn:1999:TNT

[HY99] Yoon-Hwan Hahn and Pyung-Il Yu. Towards a new technology policy: the integration of generation and diffusion. Technovation, 19(3):177–186, January 1999. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000960.

Hindle:2004:PRC

[HY04] Kevin Hindle and John Yencken. Public research commercialisation, entrepreneurship and new technology based firms: an integrated model. Technovation, 24(10):793–803, October 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203000233.

Hyvattinen:2006:ISC

[Hyv06] Heli Hyvättinen. Interface standards and creating innovation markets-implications on SMEs in a technology programme. Technovation, 26(2):262–273, February 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000611.

Ilori:1994:EBTb

[IAA94a] M. O. Ilori, O. R. Adeniyi, and S. O. O. Afilaka. Emerging biotechnologies and their potential applications in the production and processing of cocoa and palm procedure in Nigeria. Technovation, 14(5):346, June 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900809.

Ilori:1994:EBTa

[IAA94b] M. O. Ilori, O. R. Adeniyi, and S. O. O. Afilaka. Emerging biotechnologies and their potential applications in the production and processing of cocoa and palm produce in Nigeria. Technovation, 14(5):287–294, June 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900728.

Ilori:2002:DMB

[IAO^{+}02] M. O. Ilori, A. A. Adeniyi, A. A. Oyewale, S. A. Sanni, and I. A. Irefin. Developing a manufacturingbased economy in Nigeria through science and technology. Technovation, 22(1):51–60, January 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000547.

Irani:2004:TQM

[IBL04] Z. Irani, A. Beskese, and P. E. D. Love. Total quality management and corporate culture: constructs of organisational excellence. Technovation, 24(8):643–650, August 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497202001281.

Ismail:2023:EOI

[IBPGML23] Muhammad Ismail, Alejandro Bello-Pintado, Teresa García-Marco, and Valentina Lazzarotti. Enhancing open innovation: Managing not invented here syndrome in collaborative projects. Technovation, 128(??):??, December 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001906.

Ifan:2004:DCT

[IDMD04] Hadi-Kusuma Ifan, Jean-Marie Dou, Jr., Sri Manullang, and Henri Dou. Developing competitive technical intelligence in Indonesia. Technovation, 24(12):995–999, December 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203000695.

Irani:1997:ICMa

[IEG97a] Z. Irani, J.-N. Ezingeard, and R. J. Grieve. Integrating the costs of a manufacturing IT/IS infrastructure into the investment decision-making process. Technovation, 17(11–12):695–706, November/December 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297000606.

Irani:1997:ICMb

[IEG97b] Z. Irani, J.-N Ezingeard, and R. J. Grieve. Integrating the costs of a manufacturing IT/IS infrastructure into the investment decision making process. Technovation, 17(11–12):729–731, November/December 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297904648.

Ireland:2007:HSB

[IH07] David C. Ireland and Damian Hine. Harmonizing science and business agendas for growth in new biotechnology firms: Case comparisons from five countries. Technovation, 27(11):676–692, November 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000740.

Ilori:1991:OMP

[II91] M. O. Ilori and M. M. Ige. Oilseed meals as a potential solution to dietary protein shortage in Nigeria. Technovation, 11(5):315–322, July 1991. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497291900294.

Ishii:1992:PDAa

[II92a] Kazuyoshi Ishii and Takaya Ichimura. A process design approach based on the fusion model. Technovation, 12(8):499– 507, December 1992. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729290072P. See [II92b].

Ishii:1992:PDAb

[II92b] Kazuyoshi Ishii and Takaya Ichimura. A process design approach based on the fusion model: Technovation, 12:8 (1992) 499–507. Technovation, 12(8):535, December 1992. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729290076T. See [II92a].

Ilori:1997:TDM

[II97] M. O. Ilori and I. A. Irefin. Technology decision making in organisations. Technovation, 17(3):153–160, March 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497296000867.

Ibitoye:1998:IMC

[II98] S. A. Ibitoye and M. O. Ilori. Indigenous metal casting in Nigeria: its technology, attendant problems, business prospects and policy implications. Technovation, 18(11): 705–711, November 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497298000522.

Igel:2001:SSM

[II01] Barbara Igel and Nazrul Islam. Strategies for service and market development of entrepreneurial software designing firms. Technovation, 21(3):157–166, March 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000432.

Iizuka:2021:RIU

[II21] Michiko Iizuka and Yoko Ikeda. Regulation and innovation under the 4th industrial revolution: the case of a healthcare robot, HAL by Cyberdyne. Technovation, 108(??):??, December 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497221001164.

Ilori:1997:EESb

[IIA97a] M. O. Ilori, I. A. Irefin, and A. A. Adeniyi. Engineering economy studies on non-alcoholic beverages production from some tropical crops. Technovation, 17(11–12):731–732, November/December 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297904661.

Ilori:1997:EESa

[IIA97b] M. O. Ilori, I. A. Irefin, and A. A. Adeniyi. Engineering economy studies on the production of nonalcoholic beverages from some tropical crops. Technovation, 17(11–12):715–721, November/December 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649729700045X.

Iskandar:2001:BBE

[IKL01] Basuki Y. Iskandar, Susumu Kurokawa, and Larry J. LeBlanc. Business-to-business electronic commerce from first- and second-tier automotive suppliers’ perspectives: a preliminary analysis for hypotheses generation. Technovation, 21(11):719–731, November 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000530.

Ilori:1994:RLG

[Ilo94] M. O. Ilori. The role of local government in technological and industrial development in Nigeria. Technovation, 14(3): 173–179, April 1994. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729490054X. See [Ano94-31].

Ignatius:2012:ITL

[ILR^{+}12] Joshua Ignatius, Jasmine Yeap Ai Leen, T. Ramayah, Chai Kah Hin, and Muhamad Jantan. The impact of technological learning on NPD outcomes: the moderating effect of project complexity. Technovation, 32(7–8):452–463, July/August 2012. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497212000326.

Islam:2010:EAN

[IM10] Nazrul Islam and Kumiko Miyazaki. An empirical analysis of nanotechnology research domains. Technovation, 30(4): 229–237, April 2010. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497209001576.

Israeli:1998:JGP

[IMBH98] Aviad A. Israeli, Abraham Mehrez, Dov Bochen, and David Hakim. Justification of global positioning systems purchase using the analytic hierarchical process — the case of the Israeli Defense Force. Technovation, 18(6–7):409–424, January 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649729800011X.

Ishii:1987:PDB

[IMIY87] K. Ishii, R. Muramatsu, T. Ichimura, and K. Yamaguchi. Product development based on users’ needs assessment by single and comparative assessment methods. Technovation, 7(1):39–49, December 1987. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497287900460.

Ilori:2003:EBP

[INO^{+}03] M. O. Ilori, M. L. Nassar, J. O. Okolofo, J. B. Akarakiri, and T. O. Oyebisi. An evaluation of business performance and technology development in the pre- and post-privatisation period of a public company in Nigeria. Technovation, 23(2):175–182, February 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201001055.

Ilori:2000:MNP

[IOS00] M. O. Ilori, J. S. Oke, and S. A. Sanni. Management of new product development in selected food companies in Nigeria. Technovation, 20(6):333–342, June 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001479.

Ilevbare:2013:RTB

[IPP13] Imoh M. Ilevbare, David Probert, and Robert Phaal. A review of TRIZ, and its benefits and challenges in practice. Technovation, 33(2–3):30–37, February/March 2013. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497212001356.

Ilevbare:2014:TRA

[IPP14] Imoh M. Ilevbare, David Probert, and Robert Phaal. Towards risk-aware roadmapping: Influencing factors and practical measures. Technovation, 34(8):399–409, August 2014. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497214000650.

Irani:1997:ICI

[IS97] Zahir Irani and John M. Sharp. Integrating continuous improvement and innovation into a corporate culture: a case study. Technovation, 17(4):199–206, 223, April 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497296001034.

Ishigai:1989:ILDa

[Ish89a] Seikan Ishigai. Intrinsic laws of development of technics. Part I. Technovation, 9(1):9–21, May 1989. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900382.

Ishigai:1989:ILDb

[Ish89b] Seikan Ishigai. Intrinsic laws of development of technics. Part II. Technovation, 9(1):23–48, May 1989. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900394.

Suzuki:2002:EJS

[iSKB02] Kan ichiro Suzuki, Sang-Hoon Kim, and Zong-Tae Bae. Entrepreneurship in Japan and Silicon Valley: a comparative study. Technovation, 22(10):595–606, October 2002. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000992.

Jacobs:2013:CTE

[Jac13] Mark A. Jacobs. Complexity: Toward an empirical measure. Technovation, 33(4–5):111–118, April/May 2013. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213000035.

Jacqmin:2022:WSM

[Jac22] Julien Jacqmin. Why are some Massive Open Online Courses more open than others? Technovation, 112(??): ??, April 2022. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497221001760.

Jain:2023:HKL

[Jai23] Amit Jain. How knowledge loss and network-structure jointly determine R&D productivity in the biotechnology industry. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497222001547.

Jancenelle:2021:TIR

[Jan21] Vivien E. Jancenelle. Tangible-intangible resource composition and firm success. Technovation, 108(??):??, December 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221001188.

Jarboe:1986:LDH

[Jar86] Kenan Patrick Jarboe. Location decisions of hightechnology firms: a case study. Technovation, 4(2):117– 129, April 1986. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497286900039.

Jones:2019:POC

[JB19] Raymond J. Jones and Anat Barnir. Properties of opportunity creation and discovery: Comparing variation in contexts of innovativeness. Technovation, 79(??):1–10, January 2019. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497218301214.

Jimenez-Barrionuevo:2011:VIM

[JBGMM11] María Magdalena Jiménez-Barrionuevo, Víctor J. García-Morales, and Luis Miguel Molina. Validation of an instrument to measure absorptive capacity. Technovation, 31 (5–6):190–202, May/June 2011. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016649721000132X.

Jacob:2023:EMB

[JBL23] Jojo Jacob, René Belderbos, and Boris Lokshin. Entangled modes: Boundaries to effective international knowledge sourcing through technology alliances and technologybased acquisitions. Technovation, 122(??):??, April 2023.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222002127.

Johnson:2023:BLC

[JBZ23] William H. A. Johnson, Pelin Bicen, and Zhen Zhu. Being lean: Conceptualizing and operationalizing the Lean Innovation Capability (LIC) of innovative companies. Technovation, 126(??):??, August 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001050.

Jones:2001:BRI

[JC01] Oswald Jones and Martin Craven. Beyond the routine: innovation management and the Teaching Company Scheme. Technovation, 21(5):267–279, May 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000420.

Jiang:2023:OCE

[JCT23] Lin Jiang, Brent B. Clark, and Daniel B. Turban. Overcoming the challenge of exploration: How decompartmentalization of internal communication enhances the effect of exploration on employee inventive performance. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497222001584.

Jolly:2009:ITT

[JD09] Dominique Jolly and Frederic Dimanche. Investing in technology for tourism activities: Perspectives and challenges. Technovation, 29(9):576–579, September 2009. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209000741.

Jelinek:2007:BRT

[Jel07] Mariann Jelinek. Book review: Think, Play, Do: Technology, Innovation and Organization, M. Dodgson, D. Gann, A. Salter. Oxford University Press, Oxford, UK, New York (2005). Technovation, 27(4):234–235, April 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497206000927.

Jensen:1991:GE

[Jen91] John H. Jensen. Guest editorial. Technovation, 11(7):417, November 1991. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729190022V.

Juanola-Feliu:2012:MCF

[JFCFMC^{+}12] E. Juanola-Feliu, J. Colomer-Farrarons, P. Miribel-Català, J. Samitier, and J. Valls-Pasola. Market challenges facing academic research in commercializing nano-enabled implantable devices for in-vivo biomedical analysis. Technovation, 32(3–4):193–204, March/April 2012. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211001325.

James:1998:ITMb

[JGM98a] Andrew D. James, Luke Georghiou, and J. Standley Metcalfe. Integrating technology into merger and acquisition decision making. Technovation, 18(8–9):590–591, August/ September 1998. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497298901274.

James:1998:ITMa

[JGM98b] Andrew D. James, Luke Georghiou, and J. Stanley Metcalfe. Integrating technology into merger and acquisition decision making. Technovation, 18(8–9):563–573, 590–591, August/September 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497298000297.

Jayawarna:2009:KQM

[JH09] Dilani Jayawarna and Robin Holt. Knowledge and quality management: an R&D perspective. Technovation, 29 (11):775–785, November 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497209000492.

Johnston:2018:PSU

[JH18] Andrew Johnston and Robert Huggins. Partner selection and university-industry linkages: Assessing small firms’ initial perceptions of the credibility of their partners. Technovation, 78(??):15–26, December 2018. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497216300359.

Jugend:2018:RAO

[JJS^{+}18] Daniel Jugend, Charbel Jose Chiappeta Jabbour, Janaina A. Alves Scaliza, Robson Sø Rocha, Alcides Gobbo Junior, Hengky Latan, and Manoel Henrique Salgado. Relationships among open innovation, innovative performance, government support and firm size: Comparing Brazilian firms embracing different levels of radicalism in innovation. Technovation, 74–75(??):54–65, June/July 2018. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497218301147.

Jacob:2003:SIT

[JK03] William F. Jacob and Young Hoon Kwak. In search of innovative techniques to evaluate pharmaceutical R&D projects. Technovation, 23(4):291–296, April 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720100116X.

Johansson:2019:RDR

[JKF19] Magnus Johansson, Matts Kärreman, and Amalia Foukaki. Research and development resources, coopetitive performance and cooperation: the case of standardization in 3GPP, 2004–2013. Technovation, 88(??):Article 102074, December 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497217302626.

Jo:1996:RBE

[JL96] Hyungrae Jo and Jinjoo Lee. The relationship between an entrepreneur’s background and performance in a new venture. Technovation, 16(4):161–171, 211, April 1996.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497296891243.

Jones:1998:CRT

[JL98] Alan Jones and Ashish Lall. A comparative record of technological capability in ASEAN countries. Technovation, 18(4):263–274, April 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016649729700117X.

Jiang:2013:WRB

[JL13] Quanhong Jiang and Jonathan D. Linton. Where is research in business and management taking place? Technovation, 33(4–5):97–100, April/May 2013. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213000175.

Jeong:2024:WHD

[JL24] Dawoon Jeong and Jeong-Dong Lee. Where and how does a product evolve? Product innovation pattern in product lineage. Technovation, 131(??):??, March 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000087.

Jaw:2010:DNS

[JLL10] Chyi Jaw, Jyue-Yu Lo, and Yi-Hsing Lin. The determinants of new service development: Service characteristics, market orientation, and actualizing innovation effort. Technovation, 30(4):265–277, April 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209001709.

Junge:2022:CCW

[JLN22] Louise Bech Junge, Iben Cleveland Laursen, and Kristian Roed Nielsen. Choosing crowdfunding: Why do entrepreneurs choose to engage in crowdfunding? Technovation, 111(??):??, March 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497221001668.

James:2023:IET

[JLWS23] Steffan James, Zheng Liu, Gareth R. T. White, and Anthony Samuel. Introducing ethical theory to the triple helix model: Supererogatory acts in crisis innovation. Technovation, 126(??):??, August 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001438.

Jang:2023:DDT

[JLY23] Hyejin Jang, Suyeong Lee, and Byungun Yoon. Data-driven techno-socio co-evolution analysis based on a topic model and a hidden Markov model. Technovation, 126(??):??, August 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223001244.

Johnston:2005:HES

[JMK05] Brenda Johnston, Michael C. Mayo, and Anshuman Khare. Hydrogen: the energy source for the 21st century. Technovation, 25(6):569–585, June 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203002049.

Jimenez-Martinez:2001:KVE

[JMPR01] Julio Jiménez-Martinez and Yolanda Polo-Redondo. Key variables in the EDI adoption by retail firms. Technovation, 21(6):385–394, June 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497200000353.

Jimenez-Martinez:2004:IEA

[JMPR04] J. Jiménez-Martínez and Y. Polo-Redondo. The influence of EDI adoption over its perceived benefits. Technovation, 24(1):73–79, January 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497202000470.

Jiao:2003:THV

[JMT03] Jianxin Jiao, Qinhai Ma, and Mitchell M. Tseng. Towards high value-added products and services: mass customization and beyond. Technovation, 23(10):809–821, October 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497202000238.

Johne:1992:BRA

[Joh92a] Axel Johne. Book review: American manufacturing in a global market: edited by K. W. Chilton, M. E. Warren and M. L. Weidenbaum, Kluwes, Boston, 1990, ISBN 0-7923-9051-2, £28.35, 214 pages. Technovation, 12(2):141, March 1992. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729290025D.

Johne:1992:BRD

[Joh92b] Axel Johne. Book review: Developing products in half the time: by P. G. Smith and D. G. Reinertsen, Van Nostrand Reinhold, New York, 1991, ISBN 0-442-00243-2, £22. Technovation, 12(8):539, December 1992. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497292900802.

Johne:1992:BRMa

[Joh92c] Axel Johne. Book review: Market driven strategy: Processes for creating value: by G. Day, The Free Press, New York, 1990, £21.50, 405 pages. Technovation, 12(2):137, March 1992. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729290020I.

Johne:1992:BRMb

[Joh92d] Axel Johne. Book review: Marketing: an introduction: M. Christopher and M. McDonald, Pan Books, London, 1991, ISBN 0-330-32046-7, £6.99, 399 pages. Technovation, 12(6):417, September 1992. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497292900116.

Johne:1992:BRSa

[Joh92e] Axel Johne. Book review: Self-managing teams: R. F. Hicks and D. Bone, Kogan Page, London, 1991, ISBN 0-7494-0527-9, £5.99, 121 pages. Technovation, 12(6):415, September 1992. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497292900086.

Johne:1992:BRH

[Joh92f] Axel Johne. Book review: The Honda book of management: S. Mito, Kogan Page, London, 1990, ISBN 0-7494-0201-6, £7.99, 112 pages. Technovation, 12(6):415–416, September 1992. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497292900097.

Johne:1992:BRSb

[Joh92g] Axel Johne. Book review: The strategic management of technological innovation: edited by R. Loveridge and M. Pitt, John Wiley, Chichester, 1990, ISBN 0-471-92499-7, £39.95, 404 pages. Technovation, 12(6):416–417, September 1992. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729290010F.

Johne:1999:UMV

[Joh99] Axel Johne. Using market vision to steer innovation. Technovation, 19(4):203–207, February 1999. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298001126.

Johnson:2008:RRB

[Joh08] William H. A. Johnson. Roles, resources and benefits of intermediate organizations supporting triple helix collaborative R&D: the case of Precarn. Technovation, 28(8):495– 505, August 2008. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497208000217.

Jolly:2003:IWT

[Jol03] D. Jolly. The issue of weightings in technology portfolio management. Technovation, 23(5):383–391, May 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202001578.

Jolly:2008:CVE

[Jol08] Dominique R. Jolly. Chinese vs. European views regarding technology assessment: Convergent or divergent? Technovation, 28(12):818–830, December 2008. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208001193.

Jones:1996:SHI

[Jon96] Oswald Jones. Strategic HRM: the implications for pharmaceutical R&D. Technovation, 16(1):21–32, January 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497295000151.

Jones:2005:MPP

[Jon05] Oswald Jones. Managing public-private partnerships: the enactment of a new business venture. Technovation, 25(6): 587–597, June 2005. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497204001269.

Jacobsson:1996:STP

[JP96] Staffan Jacobsson and Joakim Philipson. Sweden’s technological profile: What can R&D and patents tell and what do they fail to tell us? Technovation, 16(5):245–253, 266– 267, May 1996. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497295000704.

Jimenez:1999:KFT

[JP99] Julio Jiménez and Yolanda Polo. Kalman filtering as a tool to study the technological substitution in the telecommunications sector. Technovation, 19(12):735–746, December 1999. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497299000796.

Justy:2023:EBD

[JPBL^{+}23] Théo Justy, Estelle Pellegrin-Boucher, Denis Lescop, Julien Granata, and Shivam Gupta. On the edge of Big Data: Drivers and barriers to data analytics adoption in SMEs. Technovation, 127(??):??, September 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722300161X.

Johne:1991:HTP

[JR91] Axel Johne and Stephen Rowntree. High technology product development in small firms: a challenge for marketing specialists. Technovation, 11(4):247–259, May 1991. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497291900375.

Jonker:2006:TET

[JRS06] Martijn Jonker, Henny Romijn, and Adam Szirmai. Technological effort, technological capabilities and economic performance: a case study of the paper manufacturing sector in West Java. Technovation, 26(1):121–134, January 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001816.

Jovanovic:2022:CEP

[JSP22] Marin Jovanovic, David Sjödin, and Vinit Parida. Coevolution of platform architecture, platform services, and platform governance: Expanding the platform value of industrial digital platforms. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497220300900.

Jissink:2019:FLS

[JSR19] Tymen Jissink, Fiona Schweitzer, and René Rohrbeck. Forward-looking search during innovation projects: Under which conditions it impacts innovativeness. Technovation, 84–85(??):71–85, June/July 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://

www.sciencedirect.com/science/article/pii/S0166497217305229.

Just:2024:NLP

[Jus24] Julian Just. Natural language processing for innovation search — Reviewing an emerging non-human innovation intermediary. Technovation, 129(??):??, January 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001943.

Janszen:1997:IMR

[JV97] Felix H. A. Janszen and Marc P. F. Vloemans. Innovation and the materials revolution. Technovation, 17(10): 549–556, 594, October 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497297000412.

Jin:2008:TCD

[JvZ08] Jun Jin and Maximilian von Zedtwitz. Technological capability development in China’s mobile phone industry. Technovation, 28(6):327–334, June 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720700079X.

Jung:2006:RBT

[JW06] Joo Y. Jung and Yong Jian Wang. Relationship between total quality management (TQM) and continuous improvement of international project management (CI-IPM). Technovation, 26(5–6):716–722, May/June 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497206000113.

Jia:2024:IBG

[JWL24] Xingping Jia, Jiangzhe Wang, and Tingting Liu. The impact of business-to-government relationship emphasis on green innovation: an empirical analysis. Technovation, 129(??):??, January 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497223002304.

Jiang:1992:GTAa

[JWS92a] Wen Bing Jiang, Zhi Bo Wang, and Hou Fang Sun. Group technology application in China. Technovation, 12(8):509– 514, December 1992. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729290073Q. See [JWS92b].

Jiang:1992:GTAb

[JWS92b] WenBing Jiang, ZhiBo Wang, and HouFang Sun. Group technology application in China: Technovation, 12:8 (1992) 509–514. Technovation, 12(8):536, December 1992. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729290077U. See [JWS92a].

Jiao:2022:TSI

[JWY22] Hao Jiao, Tang Wang, and Jifeng Yang. Team structure and invention impact under high knowledge diversity: an empirical examination of computer workstation industry. Technovation, 114(??):??, June 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497221002303.

Jiao:2023:SHS

[JWY23] Hao Jiao, Lindong Wang, and Jifeng Yang. Standing head and shoulders above others? Complementor experiencebased design and crowdfunding success on digital platforms. Technovation, 128(??):??, December 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001827.

Jeong:2015:DPR

[JY15] Yujin Jeong and Byungun Yoon. Development of patent roadmap based on technology roadmap by analyzing patterns of patent development. Technovation, 39–40(??):37– 52, May/June 2015. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497214000339.

Kwak:2006:BOF

[KA06] Young Hoon Kwak and Frank T. Anbari. Benefits, obstacles, and future of six sigma approach. Technovation, 26 (5–6):708–715, May/June 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497204001828.

Klerkx:2013:IMC

[KA13] Laurens Klerkx and Noelle Aarts. The interaction of multiple champions in orchestrating innovation networks: Conflicts and complementarities. Technovation, 33(6–7):193– 210, June/July 2013. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497213000242.

Kalar:2015:EUA

[KA15] Barbara Kalar and Bostjan Antoncic. The entrepreneurial university, academic activities and technology and knowledge transfer in four European countries. Technovation, 36–37(??):1–11, February/March 2015. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214001709.

Kaasa:2009:EDD

[Kaa09] Anneli Kaasa. Effects of different dimensions of social capital on innovative activity: Evidence from Europe at the regional level. Technovation, 29(3):218–233, March 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208000138.

Kafouros:2006:IIR

[Kaf06] Mario I. Kafouros. The impact of the Internet on R&D efficiency: theory and evidence. Technovation, 26(7):827– 835, July 2006. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497205000489.

Kakati:2003:SCH

[Kak03] M. Kakati. Success criteria in high-tech new ventures. Technovation, 23(5):447–457, May 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000147.

Kanz:1994:SES

[Kan94] John Kanz. Strategy evolution and structural consequences in the US semiconductor industry. Technovation, 14(4):

221–241, May 1994. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497294900086.

Kang:2022:TEW

[Kan22] Hye Young Kang. Technological engagement of women entrepreneurs on online digital platforms: Evidence from the Apple iOS App Store. Technovation, 114(??):??, June 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000694.

Kaplan:1999:PAA

[Kap99] Abram W. Kaplan. From passive to active about solar electricity: innovation decision process and photovoltaic interest generation. Technovation, 19(8):467–481, August 1999. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649729800128X.

Kaplinsky:2001:BRT

[Kap01] Raphael Kaplinsky. Book review: Tiger technology: the creation of a semiconductor industry in East Asia: John A. Mathews and Dong-Sung Cho; Cambridge University Press, Cambridge, 2000, ISBN 0-521-66269-9. Technovation, 21(5):333–334, May 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000699.

Kaplinsky:2003:BRT

[Kap03] R. Kaplinsky. Book review: Technological revolutions and financial capital: the dynamics of bubbles and golden ages: Carlota Perez (Ed.), 2002, pp 198 including index and bibliography, ISBN 1-84064-922-4 Hardback, £45. Paperback, £19.95. Technovation, 23(10):844–845, October 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000464.

Kapsali:2011:HII

[Kap11] Maria Kapsali. How to implement innovation policies through projects successfully. Technovation, 31(12):615– 626, December 2011. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497211001076.

Karlsson:1986:CNB

[Kar86] Christer Karlsson. Creation of new businesses in large established organizations. Technovation, 5(1–3):155–168, October 1986. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497286900507.

Karagozoglu:1993:EUSa

[Kar93a] Necmi Karagozoglu. Environmental uncertainty, strategic planning, and technological competitive advantage. Technovation, 13(6):335–347, September 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497293900757.

Karagozoglu:1993:EUSb

[Kar93b] Necmi Karagozoglu. Environmental uncertainty, strategic, planning, and technological competitive advantage. Technovation, 13(6):399–400, September 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390079B.

Kastrinos:1994:IEEa

[Kas94a] Nikolaos Kastrinos. Issues of exploitation: evaluating the impact of the European community framework programme on European industry. Technovation, 14(10):679–688, December 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294901139.

Kastrinos:1994:IEEb

[Kas94b] Nikos Kastrinos. Issues of exploitation: evaluating the impact of the European community framework programme on European industry. Technovation, 14(10):696–697, December 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729490118X.

Kiel:2017:III

[KAV17] Daniel Kiel, Christian Arnold, and Kai-Ingo Voigt. The influence of the Industrial Internet of Things on business models of established manufacturing companies — a business level perspective. Technovation, 68(??):4–19, December 2017. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497216303169.

Kazmi:2003:BRM

[Kaz03] Azhar Kazmi. Book review: Management of technology: by Zafar Hussain, Sushil & R. D. Pathak, Anmol Publications, New Delhi 2001; Pages 433 + viii; references, and two appendices. Technovation, 23(3):279–280, March 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202001670.

Kumar:2006:PMP

[KB06] Sameer Kumar and Erin M. Budin. Prevention and management of product recalls in the processed food industry: a case study based on an exporter’s perspective. Technovation, 26(5–6):739–750, May/June 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000817.

Kok:2009:CMO

[KB09] Robert A. W. Kok and Wim G. Biemans. Creating a market-oriented product innovation process: a contingency approach. Technovation, 29(8):517–526, August 2009. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208001557.

Kinkel:2022:PAA

[KBC22] Steffen Kinkel, Marco Baumgartner, and Enrica Cherubini. Prerequisites for the adoption of AI technologies in manufacturing — Evidence from a worldwide sample of manufacturing companies. Technovation, 110(??):??, February 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001565.

Karami:2022:EPB

[KBO22] Masoud Karami, William W. Baber, and Arto Ojala. The effectual process of business model innovation for seizing opportunities in frontier markets. Technovation, 117 (??):??, September 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222001420.

Kohtamaki:2024:SSP

[KBRR24] Marko Kohtamäki, Krishna Raj Bhandari, Rodrigo Rabetino, and Mikko Ranta. Sustainable servitization in product manufacturing companies: the relationship between firm’s sustainability emphasis and profitability and the moderating role of servitization. Technovation, 129(??): ??, January 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223002183.

Koschatzky:2001:DIP

[KBS01] Knut Koschatzky, Ulrike Bross, and Peter Stanovnik. Development and innovation potential in the Slovene manufacturing industry: analysis of an industrial innovation survey. Technovation, 21(5):311–324, May 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720000050X.

Kapoor:2022:PEV

[KBSB22] Kawaljeet Kapoor, Ali Ziaee Bigdeli, Andreas Schroeder, and Tim Baines. A platform ecosystem view of servitization in manufacturing. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497221000298.

Kafouros:2008:RIE

[KBSW08] Mario I. Kafouros, Peter J. Buckley, John A. Sharp, and Chengqi Wang. The role of internationalization in explaining innovation performance. Technovation, 28(1–2):63–74, January/February 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207001058.

Koc:2007:FII

[KC07] Tufan Koc and Cemil Ceylan. Factors impacting the innovative capacity in large-scale companies. Technovation, 27(3):105–114, March 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497205001549.

Katzy:2008:CRT

[KC08] Bernhard R. Katzy and Kevin Crowston. Competency rallying for technical innovation — the case of the Virtuelle Fabrik. Technovation, 28(10):679–692, October 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207001472.

Keizer:2002:EIE

[KDH02] Jimme A. Keizer, Lieuwe Dijkstra, and Johannes I. M. Halman. Explaining innovative efforts of SMEs.: an exploratory survey among SMEs in the mechanical and electrical engineering sector in The Netherlands. Technovation, 22(1):1–13, January 2002. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497200000912.

Kaminski:2008:KTP

[KdOL08] Paulo Carlos Kaminski, Antonio Carlos de Oliveira, and Tiago Marques Lopes. Knowledge transfer in product development processes: a case study in small and medium enterprises (SMEs) of the metal-mechanic sector from São Paulo, Brazil. Technovation, 28(1–2):29–36, January/February 2008. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497207000909.

Keegan:2014:CSS

[Kee14] Christopher Keegan. Cyber security in the supply chain: a perspective from the insurance industry. Technovation, 34(7):380–381, July 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497214000200.

Kennel:2013:KAN

[Ken13] Charles F. Kennel. Knowledge action networks and regional climate change adaptation. Technovation, 33(4–5):107, April/May 2013. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497213000308.

Kessler:2003:LRP

[Kes03] Eric H. Kessler. Leveraging e–R&D processes: a knowledgebased view. Technovation, 23(12):905–915, December 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203001081.

Kettunen:2009:AKL

[Ket09] Petri Kettunen. Adopting key lessons from agile manufacturing to agile software product development — a comparative study. Technovation, 29(6–7):408–422, June/July 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208001302.

Klessova:2022:AAM

[KET22] Svetlana Klessova, Sebastian Engell, and Catherine Thomas. Assessment of the advancement of marketupstream innovations and of the performance of research and innovation projects. Technovation, 116(??):??, August 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000426.

Knockaert:2015:GIA

[KFEC15] Mirjam Knockaert, Maw Der Foo, Truls Erikson, and Eva Cools. Growth intentions among research scientists: a cognitive style perspective. Technovation, 38(??):64–74, April 2015. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497214001801.

Kurkkio:2011:WPD

[KFL11] Monika Kurkkio, Johan Frishammar, and Ulrich Lichtenthaler. Where process development begins: a multiple case study of front end activities in process firms.

Technovation, 31(9):490–504, September 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211000733.

Ku:2007:IFC

[KGK07] Kuo-Cheng Ku, Charan K. Gurumurthy, and Hsin-Pei Kao. Inter-firms collaboration of joint venture in IC foundry business. Technovation, 27(5):296–305, May 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497206000952.

Kolesnikov:2024:FMA

[KGS^{+}24] Sergey Kolesnikov, Anna P. Goldstein, Bixuan Sun, Gabriel Chan, Venkatesh Narayanamurti, and Laura Diaz Anadon. A framework and methodology for analyzing technology spillover processes with an application in solar photovoltaics. Technovation, 134(??):??, June 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000981.

Kolchin:1989:WRM

[KH89] Michael G. Kolchin and Thomas J. Hyclak. Work rules and manufacturing cells: a case study. Technovation, 9(7):537– 550, November 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497289900199.

Kohn:2006:PBC

[KH06] Stefan Kohn and Stefan Hüsig. Potential benefits, current supply, utilization and barriers to adoption: an exploratory study on German SMEs and innovation software. Technovation, 26(8):988–998, August 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205001112.

Khan:1987:MVC

[Kha87] Arshad M. Khan. Modelling venture capital investments. Technovation, 6(1):25–35, April 1987. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497287900368.

Khare:1997:SAG

[Kha97] Anshuman Khare. Strategic advantages of good supplier relations in the Indian automobile industry. Technovation, 17(10):557–568, 594–595, October 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297000230.

Khan:1998:CEIb

[Kha98a] Mohsin U. Khan. A comparison of the electronics industry in India and Korea. Technovation, 18(2):144, February 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297000734.

Khan:1998:CEIa

[Kha98b] Mohsin U. Khan. A comparison of the electronics industry of India and Korea. Technovation, 18(2):111–123, February 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297000904.

Khare:1998:PEAb

[Kha98c] Anshuman Khare. Planet Earth and automobiles, with special reference to Indian automobile industry. Technovation, 18(5):366–367, ???? 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497298800321.

Khare:1998:PEAa

[Kha98d]  Anshuman Khare.  Planet Earth and automobiles,

with special reference to the Indian automobile industry. Technovation, 18(5):353–359, ???? 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000170.

Khan:1999:DTM

[Kha99] Mohsin U. Khan. Dynamic techno-management capability of Indian computer firms in comparison with Korea. Technovation, 19(4):243–259, February 1999. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000996.

Kim:2023:DSV

[KHKL23] Juram Kim, Suckwon Hong, Yubin Kang, and Changyong Lee. Domain-specific valuation of university technologies using bibliometrics, Jonckheere–Terpstra tests, and data envelopment analysis. Technovation, 122(??):??, April 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222002152.

Kohlbacher:2015:GOS

[KHL15] Florian Kohlbacher, Cornelius Herstatt, and Nils Levsen. Golden opportunities for silver innovation: How demographic changes give rise to entrepreneurial opportunities to meet the needs of older people. Technovation, 39–40(??):73– 82, May/June 2015. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497214000534.

Kerr:2003:EDM

[KI03] C. I. Kerr and P. C. Ivey. The Engineering Doctorate model of consultant/researcher/innovator/entrepreneur for new product development — a gas turbine instrumentation case study. Technovation, 23(2):95–102, February 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000251.

Kim:2014:RLI

[KI14] Ki-Chan Kim and Il Im. Research letter: Issues of cyber supply chain security in Korea. Technovation, 34(7):387– 388, July 2014. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497214000042.

Kihlgren:2003:PIA

[Kih03] Alessandro Kihlgren. Promotion of innovation activity in Russia through the creation of science parks: the case of St.

Petersburg (1992–1998). Technovation, 23(1):65–76, January 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201000773.

Kim:1987:GMS

[Kim87] Soung Hie Kim. Growing and mature stage technology selection patterns of small sized firms in selected Asian countries — Korea, Hong Kong, and Japan. Technovation, 7(1):63–78, December 1987. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497287900484.

Kim:2003:MTT

[Kim03] B. Kim. Managing the transition of technology life cycle. Technovation, 23(5):371–381, May 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202001682.

Kim:2016:IAM

[Kim16] Jin S. Kim. Investing in advanced materials: a marketdriven methodology. Technovation, 47(??):23–31, January 2016. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497215000802.

Kinder:2000:UIR

[Kin00] Tony Kinder. The use of the Internet in recruitment-case studies from West Lothian, Scotland. Technovation, 20 (9):461–475, September 2000. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497200000171.

Kinder:2002:SLO

[Kin02] Tony Kinder. Are schools learning organisations? Technovation, 22(6):385–404, June 2002. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497201000281.

Kharbanda:1997:ITC

[KJ97] V. P. Kharbanda and Ashok Jain. Indigenisation and technological change at the firm level — the case of the black and white TV picture tube. Technovation, 17(8):439–456, August 1997. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297000217.

Kumar:2003:CNT

[KJ03] V. Kumar and P. K. Jain. Commercialization of new technologies in India: an empirical study of perceptions of technology institutions. Technovation, 23(2):113–120, February 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201000955.

Klofsten:1996:STB

[KJE96] Magnus Klofsten and Dylan Jones-Evans. Stimulation of technology-based small firms — a case study of university– industry cooperation. Technovation, 16(4):187–193, 213, April 1996. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497295000526.

Kulicke:1987:FRP

[KK87] Marianne Kulicke and Helmar Krupp. The formation, relevance and public promotion of new technology-based firms. Technovation, 6(1):47–56, April 1987. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497287900381.

Kojima:2004:ULP

[KK04] Sakura Kojima and Raphael Kaplinsky. The use of a lean production index in explaining the transition to global competitiveness: the auto components sector in South Africa. Technovation, 24(3):199–206, March 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203001421.

Kumar:2005:SCM

[KK05] Sameer Kumar and William Krob. Supply chain management challenges for aerospace control technologies leader. Technovation, 25(1):53–58, January 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000634.

Kumar:2006:SOE

[KK06] Sameer Kumar and John Kropp. Studying the operational efficiencies of a multi-product supply chain using Excel spreadsheet model. Technovation, 26(10):1186–1200, October 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205001082.

Kucuk:2007:ACP

[KK07] S. Umit Kucuk and Sandeep Krishnamurthy. An analysis of consumer power on the Internet. Technovation, 27(1–2):47–56, January/February 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720600054X.

Katz:2013:MDG

[KK13] R. L. Katz and P. Koutroumpis. Measuring digitization: a growth and welfare multiplier. Technovation, 33(10–11):314–319, October/November 2013. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213000667.

Kalantaridis:2023:MDT

[KK23] Christos Kalantaridis and Merle Küttim. Multi-dimensional time and university technology commercialisation as opportunity praxis: a realist synthesis of the accumulated literature. Technovation, 122(??):??, April 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722200236X.

Ku:2007:VIF

[KKG07] Kuo-Cheng Ku, Hsing-Pei Kao, and Charan K. Gurumurthy. Virtual inter-firm collaborative framework — an IC foundry merger/acquisition project. Technovation, 27 (6–7):388–401, June/July 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497207000247.

Kim:2019:ATC

[KKL19] Juram Kim, Seungho Kim, and Changyong Lee. Anticipating technological convergence: Link prediction using Wikipedia hyperlinks. Technovation, 79(??):25–34, January 2019. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497216303881.

Kwon:2017:ALT

[KKP17] Heeyeul Kwon, Jieun Kim, and Yongtae Park. Applying LSA text mining technique in envisioning social impacts of emerging technologies: the case of drone technology. Technovation, 60–61(??):15–28, February 2017. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497217300044.

Kumar:2005:CMT

[KKRM05] K. Dhinesh Kumar, L. Karunamoorthy, Hubert Roth, and T. T. Mirnalinee. Computers in manufacturing: towards successful implementation of integrated automation system. Technovation, 25(5):477–488, May 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720300169X.

Kassicieh:2002:RSF

[KKWM02] Suleiman K. Kassicieh, Bruce A. Kirchhoff, Steven T. Walsh, and Paul J. McWhorter. The role of small firms in the transfer of disruptive technologies. Technovation, 22 (11):667–674, November 2002. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497201000645.

Kim:1987:PTC

[KL87] Linsu Kim and Hosun Lee. Patterns of technological change in a rapidly developing country: a synthesis. Technovation, 6(4):261–276, September 1987. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497287900745.

Kim:2004:EDI

[KL04] J. W. Kim and H. K. Lee. Embodied and disembodied international spillovers of R&D in OECD manufacturing industries. Technovation, 24(4):359–368, April 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000962.

Kwak:2005:ERT

[KL05] Young Hoon Kwak and Kenneth Scott LaPlace. Examining risk tolerance in project-driven organization. Technovation, 25(6):691–695, June 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497203001688.

Klerkx:2008:BMI

[KL08a] Laurens Klerkx and Cees Leeuwis. Balancing multiple interests: Embedding innovation intermediation in the agricultural knowledge infrastructure. Technovation, 28(6):364– 378, June 2008. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649720700065X.

Kroll:2008:SEM

[KL08b] Henning Kroll and Ingo Liefner. Spin-off enterprises as a means of technology commercialisation in a transforming economy-Evidence from three universities in China. Technovation, 28(5):298–313, May 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000624.

Kim:2024:IIP

[KL24] Sangdong Kim and Jihong Lee. The impact of international patenting on corporate patenting activities: Evidence from Korea. Technovation, 130(??):??, February 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223002493.

Kleiner:2013:MCS

[Kle13]  Matthias Kleiner.  Messages from Concurrent Ses-

sions. Technovation, 33(4–5):104, April/May 2013. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213000278.

Ku:2005:HTM

[KLH05] Yi Ling Ku, Shu-Jong Liau, and Woan-Chiau Hsing. The high-tech milieu and innovation-oriented development. Technovation, 25(2):145–153, February 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000749.

Kling:2006:SEC

[Kli06] Ragnar Kling. In search of efficiency-concurrent concept elaboration and improvement. Technovation, 26(7):753– 760, July 2006. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497205000933.

Kim:1987:KEC

[KLL87] Linsu Kim, Jangwoo Lee, and Jinjoo Lee. Korea’s entry into the computer industry and its acquisition of technological capability. Technovation, 6(4):277–293, September 1987. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497287900757.

Kokshagina:2024:BDF

[KLL24] Olga Kokshagina, Pascal Le Masson, and Jianxi Luo. Beyond the data fads: Impact of big data on contemporary innovation and technology management. Technovation, 134 (??):??, June 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497224000762.

Klodt:1987:RSE

[Klo87] Henning Klodt. R&D subsidies and export performance of manufacturing industries. Technovation, 7(1):51–61, December 1987. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497287900472.

Kohtamaki:2024:CDB

[KLP24] Marko Kohtamäki, Seppo Leminen, and Vinit Parida. Conceptualizing digital business models (DBM): Framing the interplay between digitalization and business models. Technovation, 133(??):??, May 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000634.

Kelchtermans:2022:DLL

[KLRR22] Stijn Kelchtermans, Bart Leten, Maarten Rabijns, and Massimo Riccaboni. Do licensors learn from outlicensing? Empirical evidence from the pharmaceutical industry. Technovation, 112(??):??, April 2022. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001863.

Kratzer:2006:MCT

[KLV06] Jan Kratzer, Roger Th. A. J. Leenders, and Jo M. L. Van Engelen. Managing creative team performance in virtual environments: an empirical study in 44 R&D teams. Technovation, 26(1):42–49, January 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001300.

Kinsella:1997:CCE

[KM97] Ray Kinsella and Vincent McBrierty. Campus companies and the emerging techno-academic paradigm: the Irish experience. Technovation, 17(5):245–251, May 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497296001149.

Kalafsky:2001:RTE

[KM01] Ronald V. Kalafsky and Alan D. MacPherson. Recent trends in the export performance of US machine tool companies. Technovation, 21(11):709–717, November 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000499.

Kumar:2003:EEH

[KM03]  Sameer Kumar and Thomas R. McCaffrey.  Engi-

neering economics at a hard disk drive manufacturer. Technovation, 23(9):749–755, September 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000408.

Kalafsky:2006:PRU

[KM06a] Ronald V. Kalafsky and Alan D. MacPherson. The post-1990 rebirth of the US machine tool industry: a temporary recovery? Technovation, 26(5–6):665–671, May/June 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000210.

Kumar:2006:SAC

[KM06b] Sameer Kumar and P. Malegeant. Strategic alliance in a closed-loop supply chain, a case of manufacturer and econon-profit organization. Technovation, 26(10):1127–1135, October 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205001094.

Knudsen:2011:SIN

[KM11] Mette Praest Knudsen and Thomas Bøtker Mortensen. Some immediate — but negative — effects of openness on product development performance. Technovation, 31(1):54– 64, January 2011. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497210000842.

Kelly:2022:SLL

[KM22] Gráinne Kelly and Maura McAdam. Scaffolding liminality: the lived experience of women entrepreneurs in digital spaces. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000840.

Kohler:2009:ABI

[KMF09] Thomas Kohler, Kurt Matzler, and Johann Füller. Avatarbased innovation: Using virtual worlds for real-world innovation. Technovation, 29(6–7):395–407, June/July 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208001491.

Kramer:2011:IAD

[KMID11] Jan-Philipp Kramer, Elisabetta Marinelli, Simona Iammarino, and Javier Revilla Diez. Intangible assets as drivers of innovation: Empirical evidence on multinational enterprises in German and UK regional systems of innovation. Technovation, 31(9):447–458, September 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211000885.

Kumar:2003:ICM

[KMK03] Vinod Kumar, Bharat Maheshwari, and Uma Kumar. An investigation of critical management issues in ERP implementation: empirical evidence from Canadian organizations. Technovation, 23(10):793–807, October 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000159.

Knudsen:1986:DBN

[Knu86] Niels Chr. Knudsen. Danish banking: the new competition and its consequences. Technovation, 5(1–3):61–71, October 1986. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497286900441.

Kobayashi:1997:PMP

[Kob97] Akira Kobayashi. Principle of manufacturing: a proposed new concept. Technovation, 17(4):181–187, 222, April 1997.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497296000934.

Koch:2004:INB

[Koc04] C. Koch. Innovation networking between stability and political dynamics. Technovation, 24(9):729–739, September 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497202001542.

Koch:2009:BRK

[Koc09] Anne Koch. Book review: Knowledge Sharing Among Scientists. Why Reputation Matters for R&D in Multinational Firms: Prescott C. Ensign. Palgrave Macmillan, New York (2009) pp. 1-220, £45. ISBN 0-230-61173-7. Technovation, 29(11):806, November 2009. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209000868.

Kodama:2000:NMS

[Kod00] Mitsuru Kodama. New multimedia services in the education, medical and welfare sectors. Technovation, 20(6):321– 331, June 2000. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497299001340.

Kodama:2003:TOE

[Kod03] Mitsuru Kodama. Transforming an old-economy company into a new economy — the case study of a mobile multimedia business in Japan. Technovation, 23(3):239–250, March 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201001043.

Kodama:2005:NKC

[Kod05] Mitsuru Kodama. New knowledge creation through leadership-based strategic community — a case of new product development in IT and multimedia business fields. Technovation, 25(8):895–908, August 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204000252.

Kodama:2006:KBV

[Kod06] Mitsuru Kodama. Knowledge-based view of corporate strategy. Technovation, 26(12):1390–1406, December 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205001148.

Kodama:2007:IKC

[Kod07a] Mitsuru Kodama. Innovation and knowledge creation through leadership-based strategic community: Case study on high-tech company in Japan. Technovation, 27(3):115– 132, March 2007. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649720500115X.

Kodama:2007:ITB

[Kod07b] Mitsuru Kodama. Innovation through boundary management — a case study in reforms at Matsushita electric. Technovation, 27(1–2):15–29, January/February 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205001239.

Kodama:2014:MTT

[Kod14] Fumio Kodama. MOT in transition: From technology fusion to technology-service convergence. Technovation, 34 (9):505–512, September 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497213000606.

Koga:2003:FSR

[Kog03] Tadahisa Koga. Firm size and R&D tax incentives. Technovation, 23(7):643–648, July 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016649720200010X.

Kostoff:1999:STI

[Kos99] Ronald N. Kostoff. Science and technology innovation. Technovation, 19(10):593–604, October 1999. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649729900084X.

Ko:2021:SIC

[KOX21] Young Jin Ko, Hugh O’Neill, and Xuanli Xie. Strategic intent as a contingency of the relationship between external knowledge and firm innovation. Technovation, 104(??):??, June 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221000419.

Kang:2012:IGR

[KP12] Kyung-Nam Kang and Hayoung Park. Influence of government R&D support and inter-firm collaborations on innovation in Korean biotechnology SMEs. Technovation, 32(1):68–78, January 2012. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497211001234.

Kowalewski:2023:RFC

[KP23] Oskar Kowalewski and Paweł Pisany. The rise of fintech: a cross-country perspective. Technovation, 122(??): ??, April 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497222001936.

Kumar:1994:IEDa

[KPK94a] Vinod Kumar, Aditha N. S. Persaud, and Uma Kumar. On improving the efficiency of the defence innovation process: an empirical framework. Technovation, 14(5):311–321, June 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294900744.

Kumar:1994:IEDb

[KPK94b] Vinod Kumar, Aditha N. S. Persaud, and Uma Kumar. On improving the efficiency of the defense innovation process: an empirical framework. Technovation, 14(5):348, June 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294900884.

Kihlander:2012:CBC

[KR12] Ingrid Kihlander and Sofia Ritzén. Compatibility before completeness-Identifying intrinsic conflicts in concept decision making for technical systems. Technovation, 32(2):79– 89, February 2012. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497211001490.

Kramer:2015:BRS

[Kra15] Brandy Kramer. Book review: The Second Machine Age: Work, Progress, and Prosperity in a Time of Brilliant Technologies: E. Brynjolfsson, A. McAfee. W. W. Norton & Company, New York, NY (2014). 320 pp., US$26.95, Hardcover, ISBN-10 1-4805-7747-2, ISBN-13 978-1-4805-7747-3, available in all major bookstores. Technovation, 35(??):71, January 2015. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497214001424.

Krammer:2022:HRP

[Kra22a] Sorin M. S. Krammer. Human resource policies and firm innovation: the moderating effects of economic and institutional context. Technovation, 110(??):??, February 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001474.

Krammer:2022:NNN

[Kra22b] Sorin M. S. Krammer. Navigating the New Normal: Which firms have adapted better to the COVID-19 disruption? Technovation, 110(??):??, February 2022. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001498.

Kumar:2022:IBI

[KRA^{+}22c] Shashank Kumar, Rakesh D. Raut, Nishant Agrawal, Naoufel Cheikhrouhou, Mahak Sharma, and Tugrul Daim. Integrated blockchain and Internet of Things in the food supply chain: Adoption barriers. Technovation, 118 (??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222001365.

Kateb:2022:IUP

[KRK^{+}22] Sanaz Kateb, Rebecca C. Ruehle, David P. Kroon, Elco van Burg, and Max Huber. Innovating under pressure: Adopting digital technologies in social care organizations during the COVID-19 crisis. Technovation, 115(??):??, July 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000839.

Kumar:2024:MEC

[KRM^{+}24] Mukesh Kumar, Rakesh D. Raut, Sachin Kumar Mangla, Soumyadeb Chowdhury, and Vikas Kumar Choubey. Moderating ESG compliance between Industry 4.0 and green practices with green servitization: Examining its impact on green supply chain performance. Technovation, 129(??): ??, January 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223002092.

Katsirikou:2000:IED

[KS00] Anthi Katsirikou and Elena Sefertzi. Innovation in the every day life of libraries. Technovation, 20(12):705–709, December 2000. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497200000043.

Kumar:2004:OSA

[KS04a] Sameer Kumar and Thomas Snavely. Outsourcing and strategic alliances for product development: a case of Banta Digital Group. Technovation, 24(12):1001–1010, December 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203000622.

Kumar:2004:BPR

[KS04b] Sameer Kumar and Russell Strehlow. Business process redesign as a tool for organizational development. Technovation, 24(11):853–861, November 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202001827.

Kwak:2004:PRM

[KS04c] Y. H. Kwak and J. Stoddard. Project risk management: lessons learned from software development environment. Technovation, 24(11):915–920, November 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000336.

Kim:2007:CNT

[KS07] Changsu Kim and Jaeyong Song. Creating new technology through alliances: an empirical investigation of joint patents. Technovation, 27(8):461–470, August 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000259.

Kroll:2010:EIB

[KS10a] Henning Kroll and Daniel Schiller. Establishing an interface between public sector applied research and the Chinese enterprise sector: Preparing for 2020. Technovation, 30(2):117–129, February 2010. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497209001588.

Kumar:2010:ISI

[KS10b] R. Sudhir Kumar and M. H. Bala Subrahmanya. Influence of subcontracting on innovation and economic performance of SMEs in Indian automobile industry. Technovation, 30(11–12):558–569, November/December 2010. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210000714.

Kohler:2014:RPI

[KS14] Andreas R. Köhler and Claudia Som. Risk preventative innovation strategies for emerging technologies the cases of nano-textiles and smart textiles. Technovation, 34(8):420– 430, August 2014. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649721300076X.

Khodadad-Saryazdi:2021:ETI

[KS21] Ali Khodadad-Saryazdi. Exploring the telemedicine implementation challenges through the process innovation approach: a case study research in the French healthcare sector. Technovation, 107(??):??, September 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000547.

Kamalaldin:2021:CES

[KSHP21] Anmar Kamalaldin, David Sjödin, Dusana Hullova, and Vinit Parida. Configuring ecosystem strategies for digitally enabled process innovation: a framework for equipment suppliers in the process industries. Technovation, ?? (??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497221000316.

Kaufmann:2002:HEI

[KT02] Alexander Kaufmann and Franz Tödtling. How effective is innovation support for SMEs? An analysis of the region of Upper Austria. Technovation, 22(3):147–159, March 2002. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720000081X.

Kumar:2004:PMC

[KT04] S. Kumar and D. Terpstra. The post mortem of a complex product development-lessons learned. Technovation, 24(10):805–818, October 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497202001839.

Karlsson:2004:MRP

[KTE04] Martin Karlsson, Lars Trygg, and Bengt-Olof Elfström. Measuring R&D productivity: complementing the picture by focusing on research activities. Technovation, 24(3):179– 186, March 2004. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203001111.

Kajikawa:2010:MAI

[KTSM10] Yuya Kajikawa, Yoshiyuki Takeda, Ichiro Sakata, and Katsumori Matsushima. Multiscale analysis of interfirm networks in regional clusters. Technovation, 30(3):168–180, March 2010. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497209001758.

Kuckertz:2006:BRS

[Kuc06] Andreas Kuckertz. Book review: Strategies of Competition in the Bank Card Business. Innovation Management in a Complex Economic Environment, Jarunee Wonglimpiyarat. Sussex Academic Press, Brighton, Portland (2004). Technovation, 26(7):820, July 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720500180X.

Kuckertz:2007:BRR

[Kuc07] Andreas Kuckertz. Book review: R. Tassabehji, Applying E-Commerce in Business, Sage Publications, London, Thousand Oaks, New Delhi, ISBN 0-7619-4875-9 (Paperback £26.99), ISBN 0-7619-4874-0 (Hardback £80.00), 2003 (pp. 325 + ix). Technovation, 27(3):156–157, March 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497206000502.

Kuhn:1988:SDP

[Kuh88] Arthur J. Kuhn. System design for performance control and deregulation: ‘steersmanship’ in the new environment. Technovation, 8(1–3):87–109, ???? 1988. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900557.

Klofsten:2021:MIC

[KUH21] Magnus Klofsten, David Urbano, and Sohvi Heaton. Managing intrapreneurial capabilities: an overview. Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497220300493.

Kumar:2005:EAG

[Kum05] Sameer Kumar. Exploratory analysis of global cosmetic industry: major players, technology and market trends. Technovation, 25(11):1263–1272, November 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001099.

Kumar:2006:CAK

[Kum06] Sameer Kumar. A comparative analysis of key information technology players. Technovation, 26(7):836–846, July 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000829.

Kim:2014:MRF

[KV14] YoungJun Kim and Nicholas S. Vonortas. Managing risk in the formative years: Evidence from young enterprises in Europe. Technovation, 34(8):454–465, August 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214000558.

Klofsten:2024:EOG

[KV24] Magnus Klofsten and Wim Vanhaverbeke. Exploring opportunities as a guest editor for special issues in Technovation. Technovation, 133(??):??, May 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000622.

Knot:2001:FSS

[KvdEV01] J. Marjolijn C. Knot, Jan C. M. van den Ende, and Philip J. Vergragt. Flexibility strategies for sustainable technology development. Technovation, 21(6):335–343, June 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000493.

Kwak:2005:CET

[KW05] Young Hoon Kwak and Rudy J. Watson. Conceptual estimating tool for technology-driven projects: exploring parametric estimating technique. Technovation, 25(12): 1430–1436, December 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497204001865.

Kim:2005:PVS

[KWE05] Jung-Yoon Kim, Veronica Wong, and Teck-Yong Eng. Product variety strategy for improving new product development proficiencies. Technovation, 25(9):1001–1015, September 2005. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497204000380.

Kor:2021:HPM

[KWvdS21] Burcu Kör, Ingrid Wakkee, and Peter van der Sijde. How to promote managers’ innovative behavior at work: Individual factors and perceptions. Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497218308010.

Kuo:2006:TOC

[KY06] Ying-Feng Kuo and Ching-Wen Yu. 3G telecommunication operators’ challenges and roles: a perspective of mobile commerce value chain. Technovation, 26(12):1347–1356, December 2006. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497205001124.

Kumar:2003:MCE

[KZ03] S. Kumar and C. Zahn. Mobile communications: evolution and impact on business operations. Technovation, 23(6): 515–520, June 2003. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497202001207.

Kato:2018:NLF

[KZ18] Masatoshi Kato and Haibo Zhou. Numerical labor flexibility and innovation outcomes of start-up firms: a panel data analysis. Technovation, 69(??):15–27, January 2018. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497216302085.

Lowe:2015:FCP

[LA15] Ben Lowe and Frank Alpert. Forecasting consumer perception of innovativeness. Technovation, 45–46(??):1–14, November/December 2015. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497215000103.

Leitao:2024:RST

[LAC24] Maria Eugénia Leitão, Miguel Amaral, and Ana Carvalho. Reconceptualizing socio-tech entrepreneurship: a systematic literature review and research agenda. Technovation, 134(??):??, June 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497224000683.

Landry:2013:TTO

[LACH13] Réjean Landry, Nabil Amara, Jean-Samuel Cloutier, and Norrin Halilem. Technology transfer organizations: Services and business models. Technovation, 33(12):431–449, December 2013. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497213001132.

Laestadius:2003:BRT

[Lae03] S. Laestadius. Book review: Technology and the market — demand, users and innovation: Rod Coombs, Ken Green, Albert Richards, Vivien Walsh (Eds); Edward Elgar, Cheltenham, UK, 2001, pp. 307 (incl. Index), ISBN 1-84064-469-9 (hard cover), £59.95. Technovation, 23(9):782–783, September 2003. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203000282.

Laestadius:2004:BRI

[Lae04] Staffan Laestadius. Book review: Innovation Management in the Knowledge Economy: Series on Technology Management, vol. 7; Ben Dankbaar (Ed); Imperial College Press, London, 2003, 371 pages (including index), hardcover, ISBN 1-86094-359-4. Technovation, 24(7):593–594, July 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204000240.

Lager:2017:CFP

[Lag17] Thomas Lager. A conceptual framework for platform-based design of non-assembled products. Technovation, 68(??):20– 34, December 2017. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649721630520X.

Lakoff:2008:USC

[Lak08] Sanford Lakoff. Upstart startup: “constructed advantage” and the example of Qualcomm. Technovation, 28(12):831– 837, December 2008. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497208000825.

Lall:2000:BRI

[Lal00] Sanjaya Lall. Book review: International Technology Transfer and Catch-Up in Economic Development: Slavo Radosevic; Edward Elgar, Cheltenham (UK) and Northampton, MA (USA), 1999, ix + 284, ISBN 1-85898-717-2 (price not stated). Technovation, 20(4):225–226, April 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001765.

Landstrom:1998:IIEb

[Lan98a] Hans Landström. Informal investors as entrepreneurs. Technovation, 18(5):363–365, ???? 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0166497298800291.

Landstrom:1998:IIEa

[Lan98b] Hans Landström. Informal investors as entrepreneurs: Decision-making criteria used by informal investors in their assessment of new investment proposals. Technovation, 18(5):321–333, ???? 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497298000017.

Landre:2009:AYP

[Lan09] Martin Landré. Analyzing yachting patterns in the Biesbosch National Park using GIS technology. Technovation, 29(9):602–610, September 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497209000789.

Laranja:2009:DTI

[Lar09] Manuel Laranja. The development of technology infrastructure in Portugal and the need to pull innovation using proactive intermediation policies. Technovation, 29(1):23– 34, January 2009. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497208000916.

Lay:1993:GSC

[Lay93] Gunter Lay. Government support of computer integrated manufacturing in Germany: first results of an impact analysis. Technovation, 13(5):283–297, July 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390002D.

Li:2019:UWM

[LAYS19] Yin Li, Sanjay Arora, Jan Youtie, and Philip Shapira. Using web mining to explore Triple Helix influences on growth in small and mid-size firms. Technovation, 76–77(??):3– 14, August/September 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497216000031.

Li:1982:IGC

[LB82] Yao Tzu Li and Roger A. Blais. The innovation galore, from classroom to the shop floor. Technovation, 1(4):255– 273, August 1982. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497282900074.

Link:1987:EAC

[LB87] Albert N. Link and Laura L. Bauer. An economic analysis of cooperative research. Technovation, 6(4):247–260, September 1987. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497287900733.

Lambert:2002:EIP

[LB02] A. J. D. Lambert and F. A. Boons. Eco-industrial parks: stimulating sustainable development in mixed industrial parks. Technovation, 22(8):471–484, August 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000402.

Lagace:2003:LMI

[LB03] Denis Lagacé and Mario Bourgault. Linking manufacturing improvement programs to the competitive priorities of Canadian SMEs. Technovation, 23(8):705–715, August 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497202000263.

Lakemond:2006:CLN

[LB06] Nicolette Lakemond and Christian Berggren. Co-locating NPD? The need for combining project focus and organizational integration. Technovation, 26(7):807–819, July 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720500074X.

Lager:2021:IPM

[LB21] Thomas Lager and Jessica Bruch. Innovation and production management in the process industries — in search of a conjoint approach. Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000882.

Linton:2014:CCS

[LBA14] Jonathan D. Linton, Sandor Boyson, and John Aje. The challenge of cyber supply chain security to research and practice — an introduction. Technovation, 34(7):339–341, July 2014. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497214000522.

Lopez-Berzosa:2014:IPW

[LBG14] David Lopez-Berzosa and Annabelle Gawer. Innovation policy within private collectives: Evidence on 3GPP’s regulation mechanisms to facilitate collective innovation. Technovation, 34(12):734–745, December 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649721400100X.

Lee:1991:IRS

[LbKL91] Dal Hwan Lee, Hong bumm Kim, and Jinjoo Lee. The impact of research sponsorship upon research effectiveness. Technovation, 11(1):39–57, January 1991. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729190005O.

Long:1982:ETI

[LC82] Carl F. Long and John P. Collier. Education for technological innovation at the Thayer school. Technovation, 1 (3):215–224, February 1982. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497282900049.

Lee:2007:CFC

[LC07] Chuan Lee and Wen-Jung Chen. Cross-functionality and charged behavior of the new product development teams in Taiwan’s information technology industries. Technovation, 27(10):605–615, October 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000351.

Li:2009:OEE

[LC09] Yan-Ru Li and Yiche Chen. Opportunity, embeddedness, endogenous resources, and performance of technology ventures in Taiwan’s incubation centers. Technovation, 29 (1):35–44, January 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497208000758.

Liao:2022:DIG

[LC22a] Chi-Shun Liao and Hui-Kai Chuang. Determinants of innovative green electronics: an experimental study of ecofriendly laptop computers. Technovation, 113(??):??, May 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497221002054.

Lu:2022:MOI

[LC22b] Qinli Lu and Henry Chesbrough. Measuring open innovation practices through topic modelling: Revisiting their impact on firm financial performance. Technovation, 114(??): ??, June 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497221002157.

Liao:2003:WBA

[LCLC03] Shu-Hsien Liao, Ching-Hsue Cheng, Wei-Bang Liao, and I-Liang Chen. A web-based architecture for implementing electronic procurement in military organisations. Technovation, 23(6):521–532, June 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000068.

Lefebvre:2022:STC

[LCRL22] Vincent Lefebvre, Gilles Certhoux, and Miruna Radu-Lefebvre. Sustaining trust to cross the Valley of Death: a retrospective study of business angels’ investment and reinvestment decisions. Technovation, 109(??):??, January 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497218306199.

Liu:2004:ESC

[LCT04] Pang-Lo Liu, Wen-Chin Chen, and Chih-Hung Tsai. An empirical study on the correlation between knowledge management capability and competitiveness in Taiwan’s industries. Technovation, 24(12):971–977, December 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000610.

Liu:2005:ESC

[LCT05] Pang-Lo Liu, Wen-Chin Chen, and Chih-Hung Tsai. An empirical study on the correlation between the knowledge management method and new product development strategy on product performance in Taiwan’s industries. Technovation, 25(6):637–644, June 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203002025.

Lucas:2009:IPA

[LCWC09] William A. Lucas, Sarah Y. Cooper, Tony Ward, and Frank Cave. Industry placement, authentic experience and the development of venturing and technology selfefficacy. Technovation, 29(11):738–752, November 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209000893.

Lethbridge:1995:IDC

[LD95] David Lethbridge and George Davies. An initiative for the development of creativity in science and technology (CREST): an interim report on a partnership between schools and industry. Technovation, 15(7):453–465, September 1995. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729596594J.

Liang:2001:ATA

[LD01] Yahong Liang and Sourin P. Dutta. Application trend in advanced ceramic technologies. Technovation, 21(1):61–65, January 2001. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497200000195.

Lan:2002:CAI

[LD02] Ping Lan and Howard H. Du. Challenges ahead Einnovation. Technovation, 22(12):761–767, December 2002. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000275.

Lapierre:2005:IAM

[LD05] Jozée Lapierre and Arnaud Denier. ICT adoption and moderating effects of institutional factors on salesperson’s communication effectiveness: a contingency study in hightech industries. Technovation, 25(8):909–927, August 2005.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204000136.

Luukkonen:2013:VAG

[LDB13] Terttu Luukkonen, Matthias Deschryvere, and Fabio Bertoni. The value added by government venture capital funds compared with independent venture capital funds. Technovation, 33(4–5):154–162, April/May 2013. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497212001411.

Landoni:2021:BMI

[LDF^{+}21] Paolo Landoni, Claudio Dell’era, Federico Frattini, Antonio Messeni Petruzzelli, Roberto Verganti, and Luca Manelli. Business model innovation in cultural and creative industries: Insights from three leading mobile gaming firms. Technovation, ??(??):??, ???? 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649721730086X.

Leiting:2022:ITC

[LDK22] Ann-Kathrin Leiting, Lien De Cuyper, and Christian Kauffmann. The Internet of Things and the case of Bosch: Changing business models while staying true to yourself. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S016649722200044X.

Liu:2023:DIP

[LDMS23] Yang Liu, Jiuyu Dong, Liang Mei, and Rui Shen. Digital innovation and performance of manufacturing firms: an affordance perspective. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000050.

Llopis:2022:NML

[LDMY22] Oscar Llopis, Pablo D’Este, Maureen McKelvey, and Alfredo Yegros. Navigating multiple logics: Legitimacy and the quest for societal impact in science. Technovation, 110(??):??, February 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497221001486.

Lepore:2023:IRU

[LDT^{+}23] Dominique Lepore, Koustabh Dolui, Oleksandr Tomashchuk, Heereen Shim, Chetanya Puri, Yuan Li, Nuoya Chen, and Francesca Spigarelli. Interdisciplinary research unlocking innovative solutions in healthcare. Technovation, 120(??): ??, February 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649722200058X.

Linton:2007:MTJ

[LE07] Jonathan D. Linton and Mark Embrechts. MOT TIM journal rankings 2006. Technovation, 27(3):91–94, March 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497206001192.

Lee:2000:CKT

[Lee00] Jangwoo Lee. Challenges of Korean technology-based ventures and governmental policies in the emergent-technology sector. Technovation, 20(9):489–495, September 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001704.

Lee:2004:TLN

[Lee04] Tae Joon Lee. Technological learning by national R&D: the case of Korea in CANDU-type nuclear fuel. Technovation, 24(4):287–297, April 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497202000524.

Lee:2011:DEP

[Lee11] Chang-Yang Lee. The differential effects of public R&D support on firm R&D: Theory and evidence from multicountry data. Technovation, 31(5–6):256–269, May/June 2011. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497211000204.

Lee:2023:CTI

[Lee23] Hyunmin Lee. Converging technology to improve firm innovation competencies and business performance: Evidence from smart manufacturing technologies. Technovation, 123 (??):??, May 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223000354.

Lehtonen:2007:BRM

[Leh07] Markku Lehtonen. Book review: Management of Technology: Key Success Factors for Innovation and Sustainable Development: L. Morel-Guimaraes, T. M. Khalil, Y. A. Hosni (Eds.). Elsevier, Amsterdam, San Diego, Oxford and London (2005). Technovation, 27(5):320–321, May 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497206001106.

Leone:2021:RPI

[Leo21] Ludovica Leone. The Ratatouille paradox. An inductive study of creativity in haute cuisine. Technovation, ??(??): ??, ???? 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497217301591.

Lethbridge:2003:ITV

[Let03] David Lethbridge. Innovation in technology versus innovation in financing: two case studies. Technovation, 23 (11):869–878, November 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497202000421.

Levin:1993:TTL

[Lev93] Morten Levin. Technology transfer as a learning and developmental process: an analysis of Norwegian programmes on technology transfer. Technovation, 13(8):497–518, December 1993. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497293900654. See [Ano93p].

Lin:2009:NET

[LFFT09] Julia L. Lin, Shih-Chieh Fang, Shyh-Rong Fang, and Fu-Sheng Tsai. Network embeddedness and technology transfer performance in R&D consortia in Taiwan. Technovation, 29(11):763–774, November 2009. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209000649.

Lamine:2023:IDT

[LFJA23] Wadid Lamine, Alain Fayolle, Sarah Jack, and David Audretsch. Impact of digital technologies on entrepreneurship: Taking stock and looking forward. Technovation, 126(??): ??, August 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223001347.

Liefner:2019:LFR

[LfSS19] Ingo Liefner, Yue fang Si, and Kerstin Schäfer. A latecomer firm’s R&D collaboration with advanced country universities and research institutes: the case of Huawei in Germany. Technovation, 86–87(??):3–14, August/September 2019. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497219300112.

Liu:2023:TOA

[LFU23] Zhenfeng Liu, Jian Feng, and Lorna Uden. Technology opportunity analysis using hierarchical semantic networks and dual link prediction. Technovation, 128(??):??, December 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223001839.

Li:2014:PDE

[LG14] Julia Fan Li and Elizabeth Garnsey. Policy-driven ecosystems for new vaccine development. Technovation, 34(12): 762–772, December 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497214000972.

Lo:2022:MIE

[LG22] Amadou Lô and Martha Geiger. Managing internal embeddedness in multinational corporations’ R&D subsidiaries: an evolutionary perspective on the automotive industry in Silicon Valley. Technovation, 113(??):??, May 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497221002030.

Li:2023:UHS

[LG23] Keyao Li and Mark A. Griffin. Unpacking human systems in data science innovations: Key innovator perspectives. Technovation, 128(??):??, December 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001803.

Lea:2005:PMA

[LGY05] Bih-Ru Lea, Mahesh C. Gupta, and Wen-Bin Yu. A prototype multi-agent ERP system: an integrated architecture and a conceptual framework. Technovation, 25(4):433–441, April 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203001536.

Lizarralde:2022:AMM

[LGZ22] Rafael Lizarralde, Jaione Ganzarain, and Mikel Zubizarreta. Adaptation of the MIVES method for the strategic selection of new technologies at an R&D centre. Focus on the manufacturing sector. Technovation, 115(??):??, July 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000098.

Link:1989:EE

[LH89] Albert N. Link and Robert F. Hébert. Of eggheads and entrepreneurs. Technovation, 9(1):3–4, May 1989. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900357.

Liao:2007:KTC

[LH07] Shu-Hsien Liao and Ta-Chien Hu. Knowledge transfer and competitive advantage on environmental uncertainty: an empirical study of the Taiwan semiconductor industry. Technovation, 27(6–7):402–411, June/July 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000223.

Lu:2011:EOE

[LH11] Wen-Min Lu and Shiu-Wan Hung. Exploring the operating efficiency of Technology Development Programs by an intellectual capital perspective — a case study of Taiwan. Technovation, 31(8):374–383, August 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210001161.

Lin:2014:ASN

[LH14] Fu-Ren Lin and Pei-Shan Hsieh. Analyzing the sustainability of a newly developed service: an activity theory perspective. Technovation, 34(2):113–125, February 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649721300103X.

Liu:2019:UCM

[LH19] Yipeng Liu and Qihai Huang. University capability as a micro-foundation for the Triple Helix model: the case of China. Technovation, 76–77(??):40–50, August/September 2019. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497218301263.

Lin:2021:AGA

[LH21] Liang-Hung Lin and Yu-Ling Ho. Ambidextrous governance and alliance performance under dynamic environments: an empirical investigation of Taiwanese technology alliances. Technovation, 103(??):??, May 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000213.

Lindelof:2023:IEP

[LH23] Peter Lindelöf and Roland Hellberg. Incubation — an evolutionary process. Technovation, 124(??):??, June 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000664.

Lim:2013:CBT

[LHI13] Chaisung Lim, Seokhee Han, and Hiroshi Ito. Capability building through innovation for unserved lower end mega markets. Technovation, 33(12):391–404, December 2013. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213000722.

Lim:2004:TCN

[LHY04] Kwang-Sun Lim, Yoon-Hwan Hahn, and Pyung-Il Yu. Technological competition in network markets with policy implications. Technovation, 24(9):721–728, September 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202001669.

Li:2005:BFM

[Li05] Xiaoqing Li. Buddy-finding in the mobile environment. Technovation, 25(9):1017–1023, September 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204000331.

Li:2009:TRC

[Li09] Yan-Ru Li. The technological roadmap of Cisco’s business ecosystem. Technovation, 29(5):379–386, May 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209000157.

Li:2021:DTB

[Li21] Feng Li. The digital transformation of business models in the creative industries: a holistic framework and emerging trends. Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497217300536.

Lichtenthaler:2005:CDI

[Lic05] Eckhard Lichtenthaler. Corporate diversification: identifying new businesses systematically in the diversified firm. Technovation, 25(7):697–709, July 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720300230X.

Lichtenthaler:2010:TEC

[Lic10] Ulrich Lichtenthaler. Technology exploitation in the context of open innovation: Finding the right ‘job’ for your technology. Technovation, 30(7–8):429–435, July/August 2010. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210000490.

Lichtenthaler:2011:OIF

[Lic11] Ulrich Lichtenthaler. ‘Is open innovation a field of study or a communication barrier to theory development?’ A contribution to the current debate. Technovation, 31(2–3):138–139, February/March 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210001318. See response [vdVdM11].

Liddle:1990:CRB

[Lid90] Ken Liddle. The Cambridge Research Biochemicals phenomenon. Technovation, 10(4):273–279, June 1990. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729090058R.

Lim:2006:BRT

[Lim06] Chaisung Lim. Book review: Technology Shocks-origins, Management Responses and Firm Performance: Heinrich M. Arnold; Physica-Verlag, Heidelberg and New York. 2003, pp. 264, ISBN-3-7908-0051-1 (paperback). Technovation, 26(1):137–138, January 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497205000684.

Lindsay:1985:CRC

[Lin85] Philip R. Lindsay. Counselling to resolve a clash of cognitive styles. Technovation, 3(1):57–67, February 1985. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497285900379.

Lindell:1986:RSL

[Lin86] Martin Lindell. Renewal strategies in large high-tech companies: the European approach. Technovation, 5(1–3):183– 199, October 1986. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497286900520.

Ling:1990:BRC

[Lin90] Gary Ling. Book review: Customer-driven marketing: Edited by R. Smilor, Lexington Books, Lexington, MA, 1989, $24.00, 190 pages, ISBN 0-669-21128-1. Technovation, 10(6):439–440, September 1990. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497290900088.

Ling:1991:BRC

[Lin91] Gary Ling. Book review: Customer-driven marketing: lessons from entrepreneurial technology companies: Edited by R. Smilor, Lexington Books, Lexington, MA, 1989, ISBN 0-669-21128-1, £24.00, 190 pages. Technovation, 11(2):123– 124, March 1991. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497291900456.

Linton:2002:IRS

[Lin02] Jonathan D. Linton. Implementation research: state of the art and future directions. Technovation, 22(2):65–79, February 2002. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649720100075X.

Linton:2006:IWN

[Lin06a] Jonathan Linton. Introductory words on the near and not so near future of Technovation. Technovation, 26(1):2, January 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205001689.

Linton:2006:RTI

[Lin06b] Jonathan Linton. Ranking of technology and innovation management journals. Technovation, 26(3):285–287, March 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497206000101.

Linton:2006:ET

[Lin06c] Jonathan D. Linton. Entrepreneurship today. Technovation, 26(2):139–141, February 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0166497206000022.

Linton:2007:UAG

[Lin07a] Jonathan Linton. Unsolicited Advice on getting published and the publishing process. Technovation, 27(1–2):1–3, January/February 2007. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649720600109X.

Linton:2007:MTC

[Lin07b] Jonathan D. Linton. MOT TIM Centres of Global Research 2006. Technovation, 27(9):491–500, September 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720700096X.

Linton:2008:TCS

[Lin08a] Jonathan Linton. Technovation: Current status, statistics, and a thank you to our reviewers. Technovation, 28(1–2): 1–5, January/February 2008. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497207001460.

Linton:2008:WBS

[Lin08b] Jonathan D. Linton. Why big science has trouble finding big money and small science has difficulties finding small money. Technovation, 28(12):799–801, December 2008. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208000849.

Linton:2009:E

[Lin09a] J. D. Linton. Editorial. Technovation, 29(12):807–809, December 2009. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497209001461.

Linton:2009:WSI

[Lin09b] Jonathan Linton. Why a special issue focused on tourism and hospitality? Technovation, 29(9):575, September 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720900073X.

Linton:2009:BLI

[Lin09c] Jonathan D. Linton. De-babelizing the language of innovation. Technovation, 29(11):729–737, November 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209000595.

Linton:2009:RUH

[Lin09d] Jonathan D. Linton. Reviewing: the unsung heroes of excellent journals and publications. Technovation, 29(1):1–4, January 2009. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497208001478.

Linton:2009:TIM

[Lin09e] Jonathan D. Linton. Technology innovation management’s growing influence and impact. Technovation, 29(10):643– 644, October 2009. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497209001035.

Linstone:2010:COI

[Lin10a] Harold A. Linstone. Comment on ‘is open innovation a field of study or a communication barrier to theory development?’. Technovation, 30(11–12):556, November/December 2010. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497210000945.

Linton:2010:DFB

[Lin10b] Jonathan D. Linton. Determining fit between research and journal with only two words-Some advice for young researchers on publishing and reviewing. Technovation, 30(11–12):551–553, November/December 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649721000091X.

Linton:2010:HDT

[Lin10c] Jonathan D. Linton. How do technology innovation management journals stack up against the Financial Times 45 — impressively — and other notes. Technovation, 30(9–10):483–484, September/October 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210000696.

Linton:2010:TRP

[Lin10d] Jonathan D. Linton. Thanks to reviewers and provision of recent journal statistics. Technovation, 30(1):1–2, January 2010. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497209001564.

Linstone:2011:TET

[Lin11a] Harold A. Linstone. Three eras of technology foresight. Technovation, 31(2–3):69–76, February/March 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210001094.

Linton:2011:TYO

[Lin11b] J. D. Linton. Thank you to our 2010 reviewers. Technovation, 31(5–6):187–189, May/June 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211000381.

Linton:2011:HON

[Lin11c] Jonathan Linton. How our new cover came to be. Technovation, 31(7):285, July 2011. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497211000575.

Linton:2011:WDJ

[Lin11d] Jonathan Linton. What does journal fit and focus look like: a visual representation. Technovation, 31(12):613–614, December 2011. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497211001349.

Linton:2011:ENA

[Lin11e] Jonathan D. Linton. Emerging and new approaches to R&D management. Technovation, 31(4):141, April 2011. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211000241.

Linton:2011:HIT

[Lin11f] Jonathan D. Linton. How influential are technology innovation management journals — Technology Innovation Management Journal 2010 impact factors in comparison with Financial Times 45. Technovation, 31(9):425–426, September 2011. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497211000952.

Linton:2012:ITR

[Lin12a] Jonathan D. Linton. 2011 Impact and thanks to reviewers. Technovation, 32(9–10):499–501, September/October 2012. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497212000776.

Linton:2012:BSR

[Lin12b] Jonathan D. Linton. Business school research and rankings: Where do the schools really stand? Technovation, 32(1): 1–5, January 2012. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497211001556.

Linton:2012:HGY

[Lin12c] Jonathan D. Linton. How to get your papers rejected (or not). Technovation, 32(1):6–8, January 2012. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211001313.

Linton:2012:WHW

[Lin12d] Jonathan D. Linton. What’s hot and what’s not: a summary of topics and papers in technology innovation management that are getting attention. Technovation, 32(12):653– 655, December 2012. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497212001058.

Linton:2013:ITRa

[Lin13a] Jonathan D. Linton. 2012 impact and thanks to reviewers. Technovation, 33(10–11):311–313, October/November 2013. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497213000801.

Linton:2013:ITRb

[Lin13b] Jonathan D. Linton. Identifying the top ranking business journals. Technovation, 33(2–3):25–29, February/March 2013. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497212001459.

Linton:2013:LKC

[Lin13c] Jonathan D. Linton. Letter from Kyoto — a call for research in science, technology, and society. Technovation, 33 (4–5):101–103, April/May 2013. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497213000266.

Linton:2013:RIR

[Lin13d] Jonathan D. Linton. Recognizing the Importance of Reviewers. Technovation, 33(1):1, January 2013. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497212001368.

Linton:2013:UTR

[Lin13e] Jonathan D. Linton. Updating the top ranking business journals. Technovation, 33(12):387–390, December 2013.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213001089.

Linton:2014:ADP

[Lin14] Jonathan D. Linton. Academic dishonesty primer: How to avoid it-Advice for authors. Technovation, 34(1):1–2, January 2014. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497213001387.

Linton:2016:IIR

[Lin16] Jonathan D. Linton. Improving impact of research papers. Technovation, 52–53(??):1–3, June/July 2016. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497216000171.

Linton:2017:EME

[Lin17a] Jonathan D. Linton. E = mc^{2}: Material and energy innovation as a basis for economic growth — thoughts for scientists and engineers. Technovation, 68(??):1–3, December 2017. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497217307940.

Linton:2017:ETS

[Lin17b]  Jonathan D. Linton.  Emerging technology supply

chains. Technovation, 62–63(??):1–3, April 2017. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649721730281X.

Linton:2017:RIB

[Lin17c] Jonathan D. Linton. Revealing the inner beauty of academic research: Lessons from the beauty and the beast. Technovation, 66–67(??):1–2, August 2017. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497217307241.

Linton:2018:OII

[Lin18] Jonathan D. Linton. Open innovation/integration versus disintermediation/disintegration. Technovation, 78 (??):1–3, December 2018. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016649721830405X.

Linton:2019:DTH

[Lin19a] Jonathan D. Linton. DNA of the Triple Helix: Introduction to the special issue. Technovation, 76–77(??):1– 2, August/September 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016649721830470X.

Linton:2019:RR

[Lin19b] Jonathan D. Linton. Responding to reviewers. Technovation, 86–87(??):1–2, August/September 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0166497219300422.

Linton:2020:LGW

[Lin20] Jonathan D. Linton. The limits of generalizability: Why good theory can have bad outcomes. Technovation, 89(??): ??, January 2020. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497219300628.

Liu:1993:GRD

[Liu93] Chung-Yuan Liu. Government’s role in developing a hightech industry: the case of Taiwan’s semiconductor industry. Technovation, 13(5):299–309, July 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390003E.

Liyanage:1995:BIC

[Liy95] Shantha Liyanage. Breeding innovation clusters through collaborative research networks. Technovation, 15(9):553– 567, November 1995. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729596585H.

Liu:2001:TTH

[LJ01] Hong Liu and Yunzhong Jiang. Technology transfer from higher education institutions to industry in China: nature and implications. Technovation, 21(3):175–188, March 2001. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497200000456.

Lee:2021:NPL

[LJAK21] Changyong Lee, Daeseong Jeon, Joon Mo Ahn, and Ohjin Kwon. Navigating a product landscape for technology opportunity analysis: a word2vec approach using an integrated patent-product database. Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497220300237.

Lecluyse:2020:DST

[LK20] Laura Lecluyse and Mirjam Knockaert. Disentangling satisfaction of tenants on science parks: a multiple case study in Belgium. Technovation, 98(??):??, December 2020. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497220300377.

Lee:2022:WDA

[LKCK22] Yong Suk Lee, Taekyun Kim, Sukwoong Choi, and Wonjoon Kim. When does AI pay off? AI-adoption intensity, complementary investments, and R&D strategy. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222001377.

Lee:2022:RBS

[LKKH22] Junmin Lee, Keungoui Kim, Jiyong Kim, and Junseok Hwang. The relationship between shared mobility and regulation in South Korea: a system dynamics approach from the socio-technical transitions perspective. Technovation, 109(??):??, January 2022. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001085.

Lee:2007:TRR

[LKPP07] Sungjoo Lee, Sungryong Kang, YeSeuk Park, and Yongtae Park. Technology roadmapping for R&D planning: the case of the Korean parts and materials industry. Technovation, 27(8):433–445, August 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000235.

Leyden:1993:TPAa

[LL93a] Dennis Patrick Leyden and Albert N. Link. Tax policies affecting R&D: an international comparison. Technovation, 13(1):17–25, January 1993. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729390011J.

Leyden:1993:TPAb

[LL93b] Dennis Patrick Leyden and Albert N. Link. Tax policies affecting R&D: an international comparison. Technovation, 13(1):56, January 1993. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729390015N.

Lee:1994:CSSa

[LL94a]  Jungwon Lee and Jinjoo Lee.  Competitive struc-

ture and strategic types in the Korean software industry. Technovation, 14(5):295–309, June 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900736.

Lee:1994:CSSb

[LL94b]  Jungwon Lee and Jinjoo Lee.  Competitive struc-

ture and strategic types in the Korean software industry. Technovation, 14(5):347, June 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900841.

Lefebvre:1998:GSB

[LL98] Élisabeth Lefebvre and Louis A. Lefebvre. Global strategic benchmarking, critical capabilities and performance of aerospace subcontractors. Technovation, 18(4):223–234, April 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497298000133.

Lefebvre:2002:CVE

[LL02] Louis A. Lefebvre and Élisabeth Lefebvre. E-commerce and virtual enterprises: issues and challenges for transition economies. Technovation, 22(5):313–323, May 2002. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000104.

Lofsten:2003:DEM

[LL03] Hans Löfsten and Peter Lindelöf. Determinants for an entrepreneurial milieu: Science Parks and business policy in growing firms. Technovation, 23(1):51–64, January 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000864.

Leem:2004:DCA

[LL04a] Choon Seong Leem and Hong Joo Lee. Development of certification and audit processes of application service provider for IT outsourcing. Technovation, 24(1):63–71, January 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203001457.

Lichtenthaler:2004:AFI

[LL04b] Ulrich Lichtenthaler and Eckhard Lichtenthaler. Alliance functions: implications of the international multi–R&Dalliance perspective. Technovation, 24(7):541–552, July 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203000932.

Lofsten:2005:EHS

[LL05a] Hans Löfsten and Peter Lindelöf. Environmental hostility, strategic orientation and the importance of management accounting — an empirical analysis of new technologybased firms. Technovation, 25(7):725–738, July 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204000057.

Lofsten:2005:RNP

[LL05b]  Hans Löfsten and Peter Lindelöf.  R&D networks

and product innovation patterns-academic and nonacademic new technology-based firms on Science Parks. Technovation, 25(9):1025–1037, September 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204000343.

Lin:2008:DBD

[LL08a] Hsiu-Fen Lin and Szu-Mei Lin. Determinants of e-business diffusion: a test of the technology diffusion perspective. Technovation, 28(3):135–145, March 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207001356.

Lo:2008:SCT

[LL08b] Shao-Kang Lo and Ting Lie. Selection of communication technologies — a perspective based on information richness theory and trust. Technovation, 28(3):146–153, March 2008. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000776.

Li:2022:IEV

[LL22] Yue Li and Wanli Li. Are innovative exporters vulnerable to anti-dumping investigations? Technovation, 112(??):??, April 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221001905.

Leyden:1989:EGF

[LLB89] Dennis Patrick Leyden, Albert N. Link, and Barry Bozeman. The effects of governmental financing on firms’ R&D activities: a theoretical and empirical investigation. Technovation, 9(7):561–575, November 1989. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900217.

Livesay:1996:HFI

[LLB96] Howard C. Livesay, David S. Lux, and Marilyn A. Brown. Human factors and the innovation process. Technovation, 16(4):173–186, 212, April 1996. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497295000461.

Lefebvre:2005:EBB

[LLEB05] Louis-A. Lefebvre, Élisabeth Lefebvre, Elie Elia, and Harold Boeck. Exploring B-to-B e-commerce adoption trajectories in manufacturing SMEs. Technovation, 25(12):1443–1456, December 2005. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497205000982.

Lawson:2006:ANR

[LLI06] C. P. Lawson, P. J. Longhurst, and P. C. Ivey. The application of a new research and development project selection model in SMEs. Technovation, 26(2):242–250, February 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204001312.

Lee:2023:ILM

[LLL23] Gyumin Lee, Sungjun Lee, and Changyong Lee. Inventorlicensee matchmaking for university technology licensing: a fastText approach. Technovation, 125(??):??, July 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000767.

llori:1991:USB

[llo91] M. O. llori. The use of sorghum for beer brewing in Nigeria. Technovation, 11(1):27–37, January 1991. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729190004N.

Llor:2007:DPF

[Llo07] Antoine Llor. Delay from patent filing to technology transfer: a statistical study at a major public research organization. Technovation, 27(8):446–460, August 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497206001052.

Lefebvre:1995:TPO

[LLR95] Élisabeth Lefebvre, Louis A. Lefebvre, and Marie-Josée Roy. Technological penetration and organizational learning in SMEs: the cumulative effect. Technovation, 15(8):511– 522, October 1995. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729596589L.

Lee:2000:DOC

[LLS00] Jaehee Lee, Jinjoo Lee, and William E. Souder. Differences of organizational characteristics in new product development: cross-cultural comparison of Korea and the US. Technovation, 20(9):497–508, September 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001698.

Liu:2018:IIC

[LLY^{+}18] Yang Liu, Diwei Lv, Ying Ying, Felix Arndt, and Jiang Wei. Improvisation for innovation: the contingent role of resource and structural factors in explaining innovation capability. Technovation, 74–75(??):32–41, June/July 2018. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497218301275.

Lukes:2019:DBI

[LLZ19] Martin Lukes, Maria Cristina Longo, and Jan Zouhar. Do business incubators really enhance entrepreneurial growth? Evidence from a large sample of innovative Italian start-ups. Technovation, 82–83(??):25–34, April/May 2019. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497218301627.

Liu:2023:HDC

[LLZ23] Zhidi Liu, Jia Li, and Xiao Zhang. How does CEO curiosity affect firm innovation? Technovation, 127(??):??, September 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223001591.

Liyanage:1994:SMIa

[LM94a] Shantha Liyanage and Helen Mitchell. Strategic management of interactions at the academic-industry interface. Technovation, 14(10):641–655, December 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294901112.

Liyanage:1994:SMIb

[LM94b] Shantha Liyanage and Helen Mitchell. Strategic management of interactions at the academic-industry interface. Technovation, 14(10):693–695, December 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294901163.

LeDain:2014:KSF

[LM14] Marie Anne Le Dain and Valéry Merminod. A knowledge sharing framework for black, grey and white box supplier configurations in new product development. Technovation, 34(11):688–701, November 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214001278.

Li:2024:TCE

[LM24] Daitian Li and Franco Malerba. Technological change and the evolution of the links across sectoral systems: the case of mobile communications. Technovation, 130(??):??, February 2024. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649722300247X.

Lopez-Mielgo:2009:QIM

[LMMPVO09] Nuria López-Mielgo, José M. Montes-Peón, and Camilo J. Vázquez-Ordás. Are quality and innovation management conflicting activities? Technovation, 29(8):537–545, August 2009. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649720900042X.

Latifi:2021:BMI

[LNB21] Mohammad-Ali Latifi, Shahrokh Nikou, and Harry Bouwman. Business model innovation and firm performance: Exploring causal mechanisms in SMEs. Technovation, 107(??): ??, September 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497221000559.

Lee:1994:CFT

[LO94] Mushin Lee and Kiyong Om. A conceptual framework of technological innovation management. Technovation, 14(1): 7–16, February 1994. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497294900663.

Lee:1996:DFC

[LO96] Mushin Lee and Kiyong Om. Different factors considered in project selection at public and private R&D institutes. Technovation, 16(6):271–275, June 1996. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497296000065.

Lai:2010:AME

[LO10] Jung-Yu Lai and Chorng-Shyong Ong. Assessing and managing employees for embracing change: a multiple-item scale to measure employee readiness for e-business. Technovation, 30(1):76–85, January 2010. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0166497209000716.

Lopez-Ortega:2016:CMS

[LOCSBGMH16] Eugenio López-Ortega, Damian Canales-Sanchez, Tomás Bautista-Godinez, and Santiago Macias-Herrera. Classification of micro, small and medium enterprises (M-SME) based on their available levels of knowledge. Technovation, 47(??):59–69, January 2016. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497215000747.

Larsen:1991:TID

[LOH91] Henrik H. Larsen, Michael P. O’Driscoll, and Maria Humphries. Technological innovation and the development of managerial competencies. Technovation, 11(7):419–428, November 1991. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729190023W.

Luo:2014:TBD

[LOM14] Jianxi Luo, Alison L. Olechowski, and Christopher L. Magee. Technology-based design and sustainable economic growth. Technovation, 34(11):663–677, November 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497212000752.

Lopez:2006:BRH

[Lop06] Vicente B. Ortiz Lopez. Book review: Y. A. Hosni and T. M. Khalil, Editors. Management of Technology-Internet Economy: Opportunities and Challenges for Developed and Developing Regions of the World. Selected Papers from the Eleventh International Conference on Management of Technology, Elsevier (2004). pp. xi. 556 pages, £74.50 ISBN 0-08-044214-5. Technovation, 26(7):891–892, July 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205001501.

Loveday:1985:MCP

[Lov85] D. E. E. Loveday. Managing change in pharmaceutical industry R&D. Technovation, 3(2):89–109, June 1985. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497285900033.

Lint:2003:RCD

[LP03] Onno Lint and Enrico Pennings. The recently chosen digital video standard: playing the game within the game. Technovation, 23(4):297–306, April 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201001183.

Luthje:2006:PBS

[LP06] Christian Lüthje and Reinhard Prügl. Preparing business students for co-operation in multi-disciplinary new venture teams: empirical insights from a business-planning course. Technovation, 26(2):211–219, February 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001890.

Lubango:2007:IWE

[LP07] Louis Mitondo Lubango and Anastassios Pouris. Industry work experience and inventive capacity of South African academic researchers. Technovation, 27(12):788–796, December 2007. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497207000739.

Le:2022:FAR

[LPMV22] Thanh Le, Hanh Pham, Sau Mai, and Ngoc Vu. Frontier academic research, industrial R&D and technological progress: the case of OECD countries. Technovation, 114(??):??, June 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497221002170.

Leone:2022:BST

[LPN22] Maria Isabella Leone, Antonio Messeni Petruzzelli, and Angelo Natalicchio. Boundary spanning through external technology acquisition: the moderating role of star scientists and upstream alliances. Technovation, 116(??):??, August 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000438.

Laosirihongthong:2003:ENM

[LPS03] T. Laosirihongthong, H. Paul, and M. W. Speece. Evaluation of new manufacturing technology implementation: an empirical study in the Thai automotive industry. Technovation, 23(4):321–331, April 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201001158.

Laplume:2014:PIP

[LPXO14] André O. Laplume, Saurav Pathak, and Emanuel Xavier-Oliveira. The politics of intellectual property rights regimes: an empirical study of new technology use in entrepreneurship. Technovation, 34(12):807–816, December 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214001011.

Link:1994:HFM

[LQ94] Albert N. Link and Perry D. Quick. How flexible manufacturing systems affect tooling and machining enterprises. Technovation, 14(2):85–92, March 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900981. See [Ano94m].

Leger:2009:PMP

[LQ09] Pierre-Majorique Léger and Louis Quach. Post-merger performance in the software industry: the impact of characteristics of the software product portfolio. Technovation, 29(10):704–713, October 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497209000856.

Love:2009:OIC

[LR09] James H. Love and Stephen Roper. Organizing innovation: Complementarities between cross-functional teams. Technovation, 29(3):192–203, March 2009. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208000928.

LeRoy:2022:VVH

[LRH22] Frédéric Le Roy, Frank Robert, and Rizlane Hamouti. Vertical vs horizontal coopetition and the market performance of product innovation: an empirical study of the video game industry. Technovation, 112(??):??, April 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001929.

Lemon:2004:OCK

[LS04] M. Lemon and P. S. Sahota. Organizational culture as a knowledge repository for increased innovative capacity. Technovation, 24(6):483–498, June 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202001025.

Lai:2005:CIC

[LS05] Hsien-Che Lai and Joseph Z. Shyu. A comparison of innovation capacity at science parks across the Taiwan Strait: the case of Zhangjiang High-Tech Park and Hsinchu Sciencebased Industrial Park. Technovation, 25(7):805–813, July 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203002062.

loStorto:2006:MBP

[lS06] Corrado lo Storto. A method based on patent analysis for the investigation of technological innovation strategies: the European medical prostheses industry. Technovation, 26(8):932–942, August 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205001641.

Lee:2007:SKR

[LS07] Yong-Gil Lee and Yong-Il Song. Selecting the key research areas in nano-technology field using technology cluster analysis: a case study based on National R&D Programs in South Korea. Technovation, 27(1–2):57–64, January/February 2007. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497206000514.

Lee:2016:CAS

[LS16] Jung Lee and DongBack Seo. Crowdsourcing not all sourced by the crowd: an observation on the behavior of Wikipedia participants. Technovation, 55–56(??):14–21, September/ October 2016. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497216300311.

Lager:2023:CUD

[LS23] Thomas Lager and Christopher Don Simms. From customer understanding to design for processability: Reconceptualizing the formal product innovation work process for nonassembled products. Technovation, 125(??):??, July 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000615.

Lai:2024:KSE

[LS24] I-Chun Lai and Hsin-Ning Su. Knowledge spectrum explored: Understanding source-recipient interactions and their influence on technology convergence. Technovation, 133(??):??, May 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497224000506.

Leskovar-Spacapan:2007:DOI

[LSB07] Gabrijela Leskovar-Spacapan and Majda Bastic. Differences in organizations’ innovation capability in transition economy: Internal aspect of the organizations’ strategic orientation. Technovation, 27(9):533–546, September 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000727.

Lavoie:2005:DSD

[LT05] Marie Lavoie and Pierre Therrien. Different strokes for different folks: examining the effects of computerization on Canadian workers. Technovation, 25(8):883–894, August 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204000264.

Li:2009:DTM

[LT09] Sheng-Tun Li and Ming-Hong Tsai. A dynamic taxonomy for managing knowledge assets. Technovation, 29(4):284– 298, April 2009. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497208001296.

Li:2010:VII

[LT10] Hsiu-Ling Li and Ming-Je Tang. Vertical integration and innovative performance: the effects of external knowledge sourcing modes. Technovation, 30(7–8):401–410, July/ August 2010. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649721000043X.

Li:2016:CUI

[LT16] Ying Li and Gerard J. Tellis. Is China uniform? Intra-country differences in the takeoff of new products. Technovation, 47(??):1–13, January 2016. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497215000590.

Li:2021:DCP

[LT21] Yang Li and Yujie Tang. A dynamic capabilities perspective on pro-market reforms and university technology transfer in a transition economy. Technovation, 103(??):??, May 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000055.

Lin:2006:EIC

[LTH06] Chin-Huang Lin, Chiu-Mei Tung, and Chih-Tai Huang. Elucidating the industrial cluster effect from a system dynamics perspective. Technovation, 26(4):473–482, April 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204002159.

Lowman:2012:IRO

[LTHS12] M. Lowman, P. Trott, A. Hoecht, and Z. Sellam. Innovation risks of outsourcing in pharmaceutical new product development. Technovation, 32(2):99–109, February 2012. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211001660.

Lhuillery:2023:ILC

[LTT23] Stéphane Lhuillery, Marion Tellechea, and Stéphanie Thiéry. Innovation in lieu of compliance: Internal audit departments’ standardized and non-standardized knowledge sources. Technovation, 123(??):??, May 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000317.

Livieratos:2022:OIM

[LTVA22] Antonios D. Livieratos, George Tsekouras, Wim Vanhaverbeke, and Antonios Angelakis. Open Innovation moves in SMEs: How European SMEs place their bets? Technovation, 117(??):??, September 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497222001389.

Lunn:1988:WEW

[Lun88] Martin Lunn. Water 1988 — Essex Water Company. Technovation, 8(4):245–248, December 1988. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900272.

Lundqvist:2014:ISE

[Lun14] Mats A. Lundqvist. The importance of surrogate entrepreneurship for incubated Swedish technology ventures. Technovation, 34(2):93–100, February 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213001272.

Lust:1989:RL

[Lüs89] Reimar Lüst. Reimar Lüst. Technovation, 9(2–3):299–300, June 1989. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497289900758.

Luthje:2004:CIU

[Lüt04] C. Lüthje. Characteristics of innovating users in a consumer goods field: an empirical study of sport-related product consumers. Technovation, 24(9):683–695, September 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202001505.

Li:2009:EII

[LV09] Ying Li and Wim Vanhaverbeke. The effects of interindustry and country difference in supplier relationships on pioneering innovations. Technovation, 29(12):843–858, December 2009. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497209001230.

Lafuente:2023:GCO

[LV23] Esteban Lafuente and Yancy Vaillant. Greening as a competitiveness optimizer of servitization. Technovation, 127 (??):??, September 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497223001608.

Lee:2010:FTC

[LVHR10] Jaegul Lee, Francisco M. Veloso, David A. Hounshell, and Edward S. Rubin. Forcing technological change: a case of automobile emissions control technology development in the US. Technovation, 30(4):249–264, April 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209001746.

Luiten:2006:STG

[LvLB06] Esther Luiten, Harro van Lente, and Kornelis Blok. Slow technologies and government intervention: Energy efficiency in industrial process technologies. Technovation, 26 (9):1029–1044, September 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497205001562.

Lafuente:2023:DDO

[LVR23] Esteban Lafuente, Yancy Vaillant, and Rodrigo Rabetino. Digital disruption of optimal co-innovation configurations. Technovation, 125(??):??, July 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000834.

Lafuente:2010:LDK

[LVS10] Esteban Lafuente, Yancy Vaillant, and Christian Serarols. Location decisions of knowledge-based entrepreneurs: Why some Catalan KISAs choose to be rural? Technovation, 30(11–12):590–600, November/December 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210000866.

Lafuente:2023:EPS

[LVVH23] Esteban Lafuente, Yancy Vaillant, and Ferran Vendrell-Herrero. Editorial: Product-service innovation Systems-Opening-up servitization-based innovation to manufacturing industry. Technovation, 120(??):??, February 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222002164.

Littler:1991:SAC

[LW91] Dale Littler and Dominic Wilson. Strategic alliancing in computerized business systems. Technovation, 11(8):457– 473, December 1991. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729190006P.

Liu:1997:RCF

[LW97] Xielin Liu and R. Steven White. The relative contributions of foreign technology and domestic inputs to innovation in Chinese manufacturing industries. Technovation, 17(3):119–125, March 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497296000909.

Lee:2004:TTB

[LW04] J. Lee and H. N. Win. Technology transfer between university research centers and industry in Singapore. Technovation, 24(5):433–442, May 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202001013.

Lin:2010:HDK

[LW10] Bou-Wen Lin and Chia-Hung Wu. How does knowledge depth moderate the performance of internal and external knowledge sourcing strategies? Technovation, 30(11–12):582–589, November/December 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210000830.

Lee:2011:IMI

[LW11] Keon Bong Lee and Veronica Wong. Identifying the moderating influences of external environments on new product development process. Technovation, 31(10–11):598– 612, October/November 2011. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497211000903.

Lill:2021:ACN

[LW21] Philipp A. Lill and Andreas Wald. The agility-control-nexus: a levers of control approach on the consequences of agility in innovation projects. Technovation, 107(??):??, September 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497221000572.

Lai:2009:HKM

[LWC09] Jung-Yu Lai, Chao-Te Wang, and Chun-Yi Chou. How knowledge map fit and personalization affect success of KMS in high-tech firms. Technovation, 29(4):313–324, April 2009. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497208001387.

Lin:2012:AIP

[LWC^{+}12] Chinho Lin, Ya-Jung Wu, ChiaChi Chang, Weihan Wang, and Cheng-Yu Lee. The alliance innovation performance of R&D alliances — the absorptive capacity perspective. Technovation, 32(5):282–292, May 2012. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497212000053.

Li:2022:CEI

[LWL^{+}22] Yingbo Li, Yigang Wei, Yan Li, Zhen Lei, and Alessandra Ceriani. Connecting emerging industry and regional innovation system: Linkages, effect and paradigm in China. Technovation, 111(??):??, March 2022. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001693.

Lundqvist:2024:MWU

[LWM24] Mats Lundqvist and Karen Williams-Middleton. Making the whole university entrepreneurial — decades of legitimacy-building through Chalmers School of Entrepreneurship. Technovation, 132(??):??, April 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000439.

Lee:2014:HDR

[LWP14] Cheng-Yu Lee, Hsueh-Liang Wu, and Huei-Wen Pao. How does R&D intensity influence firm explorativeness? Evidence of R&D active firms in four advanced countries. Technovation, 34(10):582–593, October 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214000546.

Laplume:2015:OAE

[LXODT15] André O. Laplume, Emanuel Xavier-Oliveira, Parshotam Dass, and Ramesh Thakur. The organizational advantage in early inventing and patenting: Empirical evidence from interference proceedings. Technovation, 43–44(??):40–48, September/October 2015. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497215000292.

Lan:1996:ITT

[LY96] Ping Lan and Stephen Young. International technology transfer examined at technology component level: a case study in China. Technovation, 16(6):277–286, June 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497296000053.

Lee:2000:CTH

[LY00] Wen-Hsiung Lee and Wei-Tzen Yang. The cradle of Taiwan high technology industry development — Hsinchu Science Park (HSP). Technovation, 20(1):55–59, January 2000.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299000851.

Lee:2009:ADN

[LYP09] Sungjoo Lee, Byungun Yoon, and Yongtae Park. An approach to discovering new technology opportunities: Keyword-based patent map approach. Technovation, 29 (6–7):481–497, June/July 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497208001326.

Li-Ying:2022:MIE

[LYST22] Jason Li-Ying, Wolfgang Sofka, and Philipp Tuertscher. Managing innovation ecosystems around Big Science Organizations. Technovation, 116(??):??, August 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497222000700.

Li:2022:AIP

[LYY^{+}22] Rebecca Yu Li, Karena Ji Yan, Nengzhi Yao, Kun Tian, Senmao Xia, Xiao hui Yang, and Yu Xiong. Abandoning innovation projects, filing patent applications and receiving foreign direct investment in R&D. Technovation, 114(??): ??, June 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497221002169.

Lu:2024:TFC

[LZD24] Keying Lu, Tengjian Zou, and Jian Du. Two forms of customer involvement and new product development performance in the digital context: the moderating role of new product development stage. Technovation, 134(??):??, June 2024. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497224000737.

Lin:2021:BGF

[LZLL21] Han Lin, Saixing Zeng, Haijian Liu, and Chao Li. Bridging the gaps or fecklessness? A moderated mediating examination of intermediaries’ effects on corporate innovation. Technovation, ??(??):??, ???? 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497218301226.

Moghaddam:2023:LMC

[MAB^{+}23] Ehsan Noorzad Moghaddam, Alireza Aliahmadi, Mehdi Bagherzadeh, Stefan Markovic, Milena Micevski, and Fatemeh Saghafi. Let me choose what I want: the influence of incentive choice flexibility on the quality of crowdsourcing solutions to innovation problems. Technovation, 120(??): ??, February 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497222002309.

Macdonald:1986:HHT

[Mac86a] Stuart Macdonald. Headhunting in high technology. Technovation, 4(3):233–245, June 1986. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0166497286900155.

Mackenzie:1986:ECS

[Mac86b] Kenneth D. Mackenzie. Environmental change and the search for organizational congruency in the USA financial services industry. Technovation, 5(1–3):73–93, October 1986. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497286900453.

Mackenzie:1988:AOC

[Mac88] Kenneth D. Mackenzie. Achieving organizational congruency through organizational design. Technovation, 8(1–3): 179–199, ???? 1988. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497288900600.

MacBryde:1997:CUT

[Mac97a] Jillian MacBryde. Commercialisation of university technology: a case in robotics. Technovation, 17(1):39–46, January 1997. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497296000740.

MacPherson:1997:RET

[Mac97b] Alan D. MacPherson. The role of external technical support in the innovation performance of scientific instruments firms: empirical evidence from New York State. Technovation, 17(3):141–151, March 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497296000818.

Macpherson:2005:LHG

[Mac05] Allan Macpherson. Learning how to grow: resolving the crisis of knowing. Technovation, 25(10):1129–1140, October 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204000707.

Majuri:2022:IFK

[Maj22] Matti Majuri. Inter-firm knowledge transfer in R&D project networks: a multiple case study. Technovation, 115(??):??, July 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000220.

Miocevic:2022:CIF

[MAKKM22] Dario Miocevic, Maja Arslanagic-Kalajdzic, and Selma Kadic-Maglajlic. Competition from informal firms and product innovation in EU candidate countries: a bounded rationality approach. Technovation, 110(??):??, February 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221001462.

Malecki:1987:RLD

[Mal87] E. J. Malecki. The R&D location decision of the firm and “creative” regions — a survey. Technovation, 6(3):205–222, August 1987. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649728790023X.

Malik:2002:ATM

[Mal02] K. Malik. Aiding the technology manager: a conceptual model for intra-firm technology transfer. Technovation, 22(7):427–436, July 2002. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016649720100030X.

Marullo:2022:OII

[MAMD22] Cristina Marullo, Joon Mo Ahn, Irene Martelli, and Alberto Di Minin. Open for innovation: an improved measurement approach using item response theory. Technovation, 109(??):??, January 2022. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722100119X.

Martinez-Alonso:2024:UNO

[MAMRRR24] Rubén Martínez-Alonso, María J. Martínez-Romero, and Alfonso A. Rojo-Ramírez. Unveiling new opportunities: a mixed gamble approach to external search breadth within family firms. Technovation, 129(??):??, January 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223002079.

Mangani:2007:TVS

[Man07] Andrea Mangàni. Technological variety and the size of economies. Technovation, 27(11):650–660, November 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000594.

Marinova:2001:EEP

[Mar01] Dora Marinova. Eastern European patenting activities in the USA. Technovation, 21(9):571–584, September 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000778.

Marsh:2003:DNZ

[Mar03] D. Marsh. Does New Zealand have an innovation system for biotechnology? Technovation, 23(2):103–112, February 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497202001530.

Mason:1986:TSW

[Mas86] David Mason. Two similar worlds divided by innovation: How marketing strength has fuelled the innovative American microchip industry in its continued lead over the inventive British. Technovation, 4(3):205–217, June 1986. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497286900131.

Mahroum:2013:TFF

[MAS13] Sami Mahroum and Yasser Al-Saleh. Towards a functional framework for measuring national innovation efficacy. Technovation, 33(10–11):320–332, October/November 2013. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497213000515.

Massaro:2023:DTH

[Mas23] Maurizio Massaro. Digital transformation in the healthcare sector through blockchain technology. Insights from academic research and business developments. Technovation, 120(??):??, February 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016649722100167X.

Matricano:2022:IGT

[Mat22] Diego Matricano. The influence of gender on technology transfer processes managed in Italian Young Innovative Companies: a stochastic frontier analysis. Technovation, 111(??):??, March 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497221001644.

Menzel:2007:WCE

[MAU07] Hanns C. Menzel, Iiris Aaltio, and Jan M. Ulijn. On the way to creativity: Engineers as intrapreneurs in organizations. Technovation, 27(12):732–743, December 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000648.

Mayne:1986:TCC

[May86] Lucille S. Mayne. Technological change and competition in American banking. Technovation, 4(1):67–83, January 1986. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497286900234.

Mustaffa:2004:EEC

[MB04] S. Mustaffa and N. Beaumont. The effect of electronic commerce on small Australian enterprises. Technovation, 24(2):85–95, February 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497202000391.

Musango:2012:SDA

[MBA^{+}12] Josephine K. Musango, Alan C. Brent, Bamikole Amigun, Leon Pretorius, and Hans Müller. A system dynamics approach to technology sustainability assessment: the case of biodiesel developments in South Africa. Technovation, 32(11):639–651, November 2012. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497212000739.

Mogab:2000:DET

[MC00] John W. Mogab and William E. Cole. Diseconomies of endogenous technological change in the globalizing continuous improvement firm. Technovation, 20(2):61–70, February 2000. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497299001431.

MChirgui:2005:SCI

[M’C05] Zouhaïer M’Chirgui. Smart card industry: a technological system. Technovation, 25(8):929–938, August 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204000318.

McAdam:2004:KCI

[McA04] R. McAdam. Knowledge creation and idea generation: a critical quality perspective. Technovation, 24(9):697–705, September 2004. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497202001694.

Marozzo:2024:IWS

[MCA^{+}24] Veronica Marozzo, Antonio Crupi, Tindara Abbate, Fabrizio Cesaroni, and Vincenzo Corvello. The impact of watching science fiction on the creativity of individuals: the role of STEM background. Technovation, 132(??):??, April 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000440.

McBean:2013:MAD

[McB13] Gordon McBean. Measures against disasters. Technovation, 33(4–5):110, April/May 2013. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497213000333.

Marques:2006:HCU

[MCD06] J. P. C. Marques, J. M. G. Caraça, and H. Diz. How can university–industry–government interactions change the innovation scenario in Portugal? — the case of the University of Coimbra. Technovation, 26(4):534–542, April 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000751.

McGrath:1998:TDM

[McG98] Robert N. McGrath. Technological discontinuities and media patterns: assessing electric vehicle batteries. Technovation, 18(11):677–687, November 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000613.

March-Chorda:2002:PDP

[MCGLA02] Isidre March-Chordà, A. Gunasekaran, and Begoña Lloria-Aramburo. Product development process in Spanish SMEs: an empirical research. Technovation, 22(5):301–312, May 2002. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201000219.

Major:2000:EBS

[MCH00] E. J. Major and M. Cordey-Hayes. Engaging the business support network to give SMEs the benefit of foresight. Technovation, 20(11):589–602, November 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000067.

McMullan:1994:GGS

[McM94a] W. Ed. McMullan. Going global on startup: a case study. Technovation, 14(3):141–143, April 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900515.

McMullan:1994:QAF

[McM94b] W. Ed McMullan. A question about focusing — a comment from the firing line. Technovation, 14(4):219–220, May 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294900078.

McMullan:1996:ETC

[McM96a] W. Ed. McMullan. An entrepreneurial tragi-comedy: Exploring the dark side of venturing. Technovation, 16(9): 515–517, 523, September 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497296000399.

McMullan:1996:JAC

[McM96b] W. Ed McMullan. Just around the corner. Technovation, 16(1):43–45, January 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729500002X.

Molina-Castillo:2009:NPP

[MCMA09] Francisco-Jose Molina-Castillo and José-Luis Munuera-Alemán. New product performance indicators: Time horizon and importance attributed by managers. Technovation, 29(10):714–724, October 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497209000637.

McNamara:1992:GRI

[McN92] John R. McNamara. Government regulation and innovation in local telephone services. Technovation, 12(6):387–396, September 1992. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497292900042.

McQueen:2005:GSR

[McQ05] Douglas H. McQueen. Growth of software related patents in different countries. Technovation, 25(6):657–671, June 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203001950.

Morgan:2006:DIA

[MCT06] Arthur Morgan, David Colebourne, and Brychan Thomas. The development of ICT advisors for SME businesses: an innovative approach. Technovation, 26(8):980–987, August 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205001173.

Mehra:2003:SPO

[MD03] K. Mehra and S. K. Dhawan. Study of the process of organisational learning in software firms in India. Technovation, 23(2):121–129, February 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720100089X.

Merino:2006:YTC

[MdCÁ06] M. Teresa García Merino, M. Liduina Pereira do Carmo, and M. Valle Santos Álvarez. 25 years of Technovation: Characterisation and evolution of the journal. Technovation, 26(12):1303–1316, December 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205001707.

Madanaguli:2023:DUO

[MDT^{+}23] Arun Madanaguli, Amandeep Dhir, Shalini Talwar, Thomas Clauss, Sascha Kraus, and Puneet Kaur. Diving into the uncertainties of open innovation: a systematic review of risks to uncover pertinent typologies and unexplored horizons. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222001298.

Manders:2016:IPI

[MdVB16] Basak Manders, Henk J. de Vries, and Knut Blind. ISO 9001 and product innovation: a literature review and research framework. Technovation, 48–49(??):41–55, February/March 2016. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497215000784.

Mehrez:1988:SRP

[Meh88] Abraham Mehrez. Selecting R&D projects: a case study of the expected utility approach. Technovation, 8(4):299–311, December 1988. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497288900314.

Mehra:2001:ISI

[Meh01] Kavita Mehra. Indian system of innovation in biotechnology — a case study of cardamom. Technovation, 21(1):15–23, January 2001. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497200000201.

Mention:2011:COC

[Men11] Anne-Laure Mention. Co-operation and co-opetition as open innovation practices in the service sector: Which influence on innovation novelty? Technovation, 31(1):44–53, January 2011. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497210000908.

Meredith:1994:BRE

[Mer94a] Sandra Meredith. Book review: Eco-management and auditing: a practical guide to the EC regulation, Joseph Tanega. IFS International Ltd, Bedford, UK (1993). Technovation, 14(6):424, August 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0166497294900256.

Meredith:1994:BRS

[Mer94b] Sandra Meredith. Book review: Science and technology management bibliography 1993, Thomas E. Clarke, Jean Reavley. Stargate Consultants Ltd, Cold Lake, Alberta, Canada (1993). Technovation, 14(6):423, August 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900248.

Meredith:2001:BRC

[Mer01] Sandra Meredith. Book review: Champions of Change: How CEOs and their Companies are Mastering the Skills of Radical Change: David A. Nadler; Jossey-Bass Publishers, San Francisco, USA, Price US$27. Technovation, 21(10): 706, October 2001. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497201000426.

Meredith:2002:BRT

[Mer02] Sandra Meredith. Book review: Team-based Learning: Howard Hill; Gower Publishing Limited, 2001, pp.176, ISBN 0-566-08364-7, Price £42.50 hardback. Technovation, 22(8):535, August 2002. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497202000172.

Merguei:2022:VDE

[Mer22] Nitzan Merguei. Venturing out: Designing effective preacceleration programs. Technovation, 116(??):??, August 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000475.

Maynard:2003:IAW

[MFH03] S. Maynard, S. Foster, and D. J. Hall. ISO 17025 application within racing chemistry: a case study. Technovation, 23(9):773–780, September 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000026.

Maresch:2016:WPM

[MFH16] Daniela Maresch, Matthias Fink, and Rainer Harms. When patents matter: the impact of competition and patent age on the performance contribution of intellectual property rights protection. Technovation, 57–58(??):14–20, November/December 2016. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497215000838.

McGloin:1998:SPSb

[MG98a] E. McGloin and C. Grant. Supporting partnership sourcing in Northern Ireland through advanced technology. Technovation, 18(2):142–143, February 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297000837.

McGloin:1998:SPSa

[MG98b] Eileen McGloin and Caroline Grant. Supporting partnership sourcing in Northern Ireland through advanced technology. Technovation, 18(2):91–99, February 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297001053.

McMullan:1998:DTS

[MG98c] W. E. McMullan and L. M. Gillin. Developing technological start-up entrepreneurs: a case study of a graduate entrepreneurship programme at Swinburne University. Technovation, 18(4):275–286, April 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297001193.

Molero:2008:IAF

[MG08] Jose Molero and Antonio Garcia. The innovative activity of foreign subsidiaries in the Spanish Innovation System: an evaluation of their impact from a sectoral taxonomy approach. Technovation, 28(11):739–757, November 2008. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208000424.

M:2021:LAS

[MG21] Vijaya Sunder M. and L. S. Ganesh. Lean additives in a service factory: a design science approach. Technovation, 104 (??):??, June 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649722100050X.

Mora:2023:SCG

[MGAP23] Luca Mora, Paolo Gerli, Lorenzo Ardito, and Antonio Messeni Petruzzelli. Smart city governance from an innovation management perspective: Theoretical framing, review of current practices, and future research agenda. Technovation, 123(??):??, May 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000287.

Marchi:2011:ELU

[MGdG11] Gianluca Marchi, Claudio Giachetti, and Pamela de Gennaro. Extending lead-user theory to online brand communities: the case of the community Ducati. Technovation, 31(8):350–361, August 2011. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497211000599.

Mwesiumo:2021:IPP

[MGGOS21] Deodat Mwesiumo, Richard Glavee-Geo, Kjetil Magnus Olsen, and Geir Arne Svenning. Improving public purchaser attitudes towards public procurement of innovations. Technovation, 101(??):??, March 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497220300791.

Molas-Gallart:1999:TGT

[MGS99] Jordi Molas-Gallart and Tom Sinclair. From technology generation to technology transfer: the concept and reality of the “Dual-Use Technology Centres”. Technovation, 19(11):661–671, November 1999. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299000632.

Matzler:1998:HMP

[MH98] Kurt Matzler and Hans H. Hinterhuber. How to make product development projects more successful by integrating Kano’s model of customer satisfaction into quality function deployment. Technovation, 18(1):25–38, January 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297000722.

Mark-Herbert:2004:INP

[MH04] Cecilia Mark-Herbert. Innovation of a new product category — functional foods. Technovation, 24(9):713–719, September 2004. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497202001311.

Mellor:2005:MMP

[MH05] Robert Mellor and Paul W. Hyland. Manufacturing management programs: are developing economies bridging the strategic gap? Technovation, 25(8):857–863, August 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204000082.

Meissner:2024:DEC

[MHKZ24] Dirk Meissner, Rainer Harms, Jan Kratzer, and Joseph Zhou. The dead end of classical innovation management and unsustainable innovation. Technovation, 129(??):??, January 2024. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223002274.

Miyazaki:2007:NSI

[MI07] Kumiko Miyazaki and Nazrul Islam. Nanotechnology systems of innovation — an analysis of industry and academia research activities. Technovation, 27(11):661–675, November 2007. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497207000697.

Mian:1994:UUS

[Mia94] Sarfraz A. Mian. US university-sponsored technology incubators: an overview of management, policies and performance. Technovation, 14(8):515–528, October 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294901511.

Miguelez:2019:CPM

[Mig19] Ernest Miguelez. Collaborative patents and the mobility of knowledge workers. Technovation, 86–87(??):62– 74, August/September 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016649721930001X.

Muramatsu:1990:ANA

[MII90] R. Muramatsu, T. Ichimura, and K. Ishii. An analysis of needs assessment and information behaviour in product development based on the fusion model. Technovation, 10(5): 305–317, July 1990. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729090015C.

Mikkola:2001:PMR

[Mik01] Juliana Hsuan Mikkola. Portfolio management of R&D projects: implications for innovation management. Technovation, 21(7):423–435, July 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497200000626.

Miles:1985:BRE

[Mil85] Ian Miles. Book review: The electronic oracle: Computer models and social decisions: D. H. Meadows and J. M. Robinson. John Wiley & Sons, Chichester, 1985. Technovation, 3(4):317–318, October 1985. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497285900288.

Miles:1988:BRO

[Mil88] Ian Miles. Book review: The organisational implications of computer technology for professional work: by Stephen E. Little, Farnborough, 1988, 199 + xvii pages, £19.50. Technovation, 7(2):177–178, May 1988. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900466.

Miller:1994:QRE

[Mil94] Roger Miller. Quality in research: an empirical study. Technovation, 14(6):381–394, August 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0166497294900175.

See [Ano94z].

Milosevic:2002:SCR

[Mil02] D. Z. Milosevic. Selecting a culturally responsive project management strategy. Technovation, 22(8):493–508, August 2002. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201000542.

Minkes:1994:BRG

[Min94] A. L. Minkes. Book review: Global disasters: Inquiries into management ethics: by Robert E. Allinson. Prentice-Hall, 1993, 280 pp., ISBN 0-13-145947-3, £15.00 or US$34.95. Technovation, 14(3):213–214, April 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900620.

Minkes:1995:MEV

[Min95] A. L. Minkes. Managers as entrepreneurs: a view from Asia. Technovation, 15(3):177–182, April 1995. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497295966184.

Minkes:1996:BRC

[Min96] A. L. Minkes. Book review: The Challenger launch decision. Risky technology, culture, and deviance at NASA: By Diane Vaughan. The University of Chicago Press, Chicago and London, 1996, ISBN 0-226-85175-3, US$24.95, 575 pp. Technovation, 16(9):525–526, September 1996. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497296827866.

Miyata:1996:ACR

[Miy96] Yukio Miyata. An analysis of cooperative R&D in the United States. Technovation, 16(3):123–131, March 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729500033X.

Miyata:2000:EAI

[Miy00] Yukio Miyata. An empirical analysis of innovative activity of universities in the United States. Technovation, 20 (8):413–425, August 2000. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497200000249.

Mansell:1992:NIR

[MJ92] Robin Mansell and Michael Jenkins. Networks, industrial restructuring and policy: the Singapore example. Technovation, 12(6):397–406, September 1992. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497292900053.

Macdonald:1994:TTI

[MJ94] Roderick J. Macdonald and Wang Jinliang. Time, timeliness of innovation, and the emergence of industries. Technovation, 14(1):37–53, February 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900698.

Mu:2024:ACV

[MJ24] Wen Mu and Xu Jiang. Absorptive capacity versus competency trap: Experiential knowledge and investment in emerging technologies. Technovation, 131(??):??, March 2024. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497224000233.

Meyer-Krahmer:1987:EIP

[MK87] Frieder Meyer-Krahmer. Evaluating innovation policies: the German experience. Technovation, 5(4):317–330, February 1987. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497287900708.

Murphy:1996:RPA

[MK96] Steven A. Murphy and Vinod Kumar. The role of predevelopment activities and firm attributes in new product success. Technovation, 16(8):431–441, 448–449, August 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497296000284.

Mildenberger:2000:PEF

[MK00] Udo Mildenberger and Anshuman Khare. Planning for an environment-friendly car. Technovation, 20(4):205–214, April 2000. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649729900111X.

McAdam:2005:DIT

[MKGL05] Rodney McAdam, William Keogh, Brendan Galbraith, and Don Laurie. Defining and improving technology transfer business and management processes in university innovation centres. Technovation, 25(12):1418–1429, December 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001385.

Madanmohan:2004:ILT

[MKK04] T. R. Madanmohan, Uma Kumar, and Vinod Kumar. Import-led technological capability: a comparative analysis of Indian and Indonesian manufacturing firms. Technovation, 24(12):979–993, December 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000300.

Mansoori:2019:ILS

[MKL19] Yashar Mansoori, Tomas Karlsson, and Mats Lundqvist. The influence of the lean startup methodology on entrepreneurcoach relationships in the context of a startup accelerator. Technovation, 84–85(??):37–47, June/July 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497217301918.

Mekhilef:2005:PMC

[ML05] Mounib Mekhilef and Julie Stal Le Cardinal. A pragmatic methodology to capture and analyse decision dysfunctions in development projects. Technovation, 25(4): 407–420, April 2005. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203001500.

Ma:2008:PAT

[ML08a] Zhenzhong Ma and Yender Lee. Patent application and technological collaboration in inventive activities: 1980– 2005. Technovation, 28(6):379–390, June 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207001071.

Mansury:2008:IPG

[ML08b] Mica Ariana Mansury and James H. Love. Innovation, productivity and growth in US business services: a firm-level analysis. Technovation, 28(1–2):52–62, January/February 2008. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497207000818.

Meng:2024:PAF

[MLC24] Qiaoyu Meng, Yi Li, and Qinwei Cao. The paradox analysis and functional mechanism between R&D efficiency and transformation effect: Evidence from key universities in China. Technovation, 130(??):??, February 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223002456.

Mian:2016:TBI

[MLF16] Sarfraz Mian, Wadid Lamine, and Alain Fayolle. Technology business incubation: an overview of the state of knowledge. Technovation, 50–51(??):1–12, April/May 2016. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497216000183.

Maine:2012:PBV

[MLG12] Elicia Maine, Sarah Lubik, and Elizabeth Garnsey. Processbased vs. product-based innovation: Value creation by nanotech ventures. Technovation, 32(3–4):179–192, March/ April 2012. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497211001477.

McAdam:2002:SNP

[MM02] Rodney McAdam and John McClelland. Sources of new product ideas and creativity practices in the UK textile industry. Technovation, 22(2):113–121, February 2002. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000025.

McAdam:2008:HTS

[MM08] Maura McAdam and Rodney McAdam. High tech startups in University Science Park incubators: the relationship between the start-up’s lifecycle progression and use of the incubator’s resources. Technovation, 28(5):277–290, May 2008. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497207000971.

Mortara:2011:HDL

[MM11] Letizia Mortara and Tim Minshall. How do large multinational companies implement open innovation? Technovation, 31(10–11):586–597, October/November 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649721100071X.

Mahotra:2024:DIC

[MM24] Arvind Mahotra and Ann Majchrzak. Digital innovations in crowdsourcing using AI tools. Technovation, 133(??):??, May 2024. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497224000476.

Marzi:2023:DDC

[MMC^{+}23] Giacomo Marzi, Mohammad Fakhar Manesh, Andrea Caputo, Massimiliano Matteo Pellegrini, and Bozidar Vlacić.

Do or do not. Cognitive configurations affecting open innovation adoption in SMEs. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222001328.

Mortati:2023:DDH

[MMCD23] Marzia Mortati, Stefano Magistretti, Cabirio Cautela, and Claudio Dell’Era. Data in design: How big data and thick data inform design thinking projects. Technovation, 122 (??):??, April 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497222002395.

Montes:2005:ISL

[MMM05] Fco. Javier Lloréns Montes, Antonia Ruiz Moreno, and Victor García Morales. Influence of support leadership and teamwork cohesion on organizational learning, innovation and performance: an empirical examination. Technovation, 25(10):1159–1172, October 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204000914.

McAdam:2016:SRU

[MMM16] Maura McAdam, Kristel Miller, and Rodney McAdam. Situated regional university incubation: a multi-level stakeholder perspective. Technovation, 50–51(??):69–78, April/ May 2016. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497215000632.

Mariani:2023:AII

[MMMD23] Marcello M. Mariani, Isa Machado, Vittoria Magrelli, and Yogesh K. Dwivedi. Artificial intelligence in innovation research: a systematic review, conceptual framework, and future research directions. Technovation, 122(??):??, April 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497222001705.

McAdam:2012:DUT

[MMMT12] Rodney McAdam, Kristel Miller, Maura McAdam, and Sinead Teague. The development of University Technology Transfer stakeholder relationships at a regional level: Lessons for the future. Technovation, 32(1):57–67, January 2012. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497211001106.

Mittone:2024:ICP

[MMR24] Luigi Mittone, Azzurra Morreale, and Paavo Ritala. Initial conditions and path dependence in explorative and exploitative learning: an experimental study. Technovation, 129(??):??, January 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497223002067.

Moritz:2022:ADC

[MNL22] Alexandra Moritz, Tamara Naulin, and Eva Lutz. Accelerators as drivers of coopetition among early-stage startups. Technovation, 111(??):??, March 2022. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001590.

Magnusson:2014:EHI

[MNW14] Peter R. Magnusson, Johan Netz, and Erik Wästlund. Exploring holistic intuitive idea screening in the light of formal criteria. Technovation, 34(5–6):315–326, May/June 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214000352.

McQueen:2003:GES

[MO03] D. H. McQueen and H. Olsson. Growth of embedded software related patents. Technovation, 23(6):533–544, June 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497202000111.

Mohamed:1995:IIM

[Moh95] M. Zain Mohamed. Innovation implementations in Malaysian firms: Process, problems, critical success factors and working climate. Technovation, 15(6):375–385, August 1995. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729596598N.

Moller:1991:TGM

[Møl91] Kim Møller. Technology and growth — measures and concepts: a case study of Denmark. Technovation, 11(8):475– 481, December 1991. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729190007Q.

Molina:1997:INTb

[Mol97a] A. H. Molina. Insights into the nature of technology diffusion and implementation: the perspective of sociotechnical alignment. Technovation, 17(11–12):723–724, November/ December 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497297904582.

Molina:1997:INTa

[Mol97b] Alfonso H. Molina. Insights into the nature of technology diffusion and implementation: the perspective of sociotechnical alignment. Technovation, 17(11–12):601–626, November/December 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649729700076X.

Molina:1998:URT

[Mol98] Alfonso H. Molina. Understanding the role of the technical in the build-up of sociotechnical constituencies. Technovation, 19(1):1–29, November 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000935.

McAdam:2008:CKS

[MOM08] Rodney McAdam, Tom O’Hare, and Sandra Moffett. Collaborative knowledge sharing in Composite New Product Development: an aerospace study. Technovation, 28(5): 245–256, May 2008. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497207000922.

Montresor:2001:TGT

[Mon01] Sandro Montresor. Techno-globalism, techno-nationalism and technological systems: organizing the evidence. Technovation, 21(7):399–412, July 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0166497200000614.

Montalvo:2006:WTC

[Mon06] Carlos Montalvo. What triggers change and innovation? Technovation, 26(3):312–323, March 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001531.

Moodley:2003:CBS

[Moo03] S. Moodley. The challenge of e-business for the South African apparel sector. Technovation, 23(7):557–570, July 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497202000020.

Morris:1991:BRP

[Mor91] Ann Morris. Book review: Product innovation: idea to exploitation: by J. A. A. Bradbury, Wiley, Chichester, 1989, ISBN 0-471-92169-6, £35.00, 233 pages. Technovation, 11 (2):129–130, March 1991. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497291900526.

Morris:2004:BRI

[Mor04] Neil Morris. Book review: Inside the Virtual Prototype — How Organizations Create Knowledge through Software: Luciana D’Adderio; Edward Elgar Publishing Limited, Cheltenham, UK, 2004, pp. 236 (including Index), ISBN 1-84376-210-2, £55 (US$90). Technovation, 24(8):671–672, August 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204000690.

Mosey:2016:TRO

[Mos16] Simon Mosey. Teaching and research opportunities in technology entrepreneurship. Technovation, 57–58(??):43–44, November/December 2016. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497216302619.

Moultrie:2015:UCR

[Mou15] James Moultrie. Understanding and classifying the role of design demonstrators in scientific exploration. Technovation, 43–44(??):1–16, September/October 2015. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic).  URL http://www.sciencedirect.com/
science/article/pii/S0166497215000437. See correc-

tion [Mou17].

Moultrie:2017:CUC

[Mou17] James Moultrie. Corrigendum to “Understanding and classifying the role of design demonstrators in scientific exploration” [Technovation 43-44 (2015) 1–16]. Technovation, 62–63(??):55, April 2017. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497217302286.

See [Mou15].

Mowery:1988:AED

[Mow88] David C. Mowery. Assessing the effects of divestiture on Bell telephone laboratories. Technovation, 7(4):353–375, October 1988. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497288900077.

Mowshowitz:1989:MTC

[Mow89] Abbe Mowshowitz. On managing technological change. Technovation, 9(8):623–633, December 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http:

//www.sciencedirect.com/science/article/pii/0166497289900035.

Mowery:1990:JCA

[Mow90] David C. Mowery. The Japanese commercial aircraft industry: déjà vu all over again? Technovation, 10(6):419–435, September 1990. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497290900055.

Matsumoto:2002:OTD

[MOWGB02] Kiyofumi Matsumoto, Noritomo Ouchi, Chihiro Watanabe, and Charla Griffy-Brown. Optimal timing of the development of innovative goods with generation — an empirical analysis focusing on Canon’s printer series. Technovation, 22(3):175–185, March 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000037.

Masson:2009:CHS

[MP09a] Sophie Masson and Romain Petiot. Can the high speed rail reinforce tourism attractiveness? The case of the high speed rail between Perpignan (France) and Barcelona (Spain). Technovation, 29(9):611–617, September 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209000790.

Murovec:2009:ACD

[MP09b] Nika Murovec and Igor Prodan. Absorptive capacity, its determinants, and influence on innovation output: Cross-cultural validation of the structural model. Technovation, 29(12):859–872, December 2009. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209000844.

Maine:2005:INM

[MPA05] Elicia Maine, David Probert, and Mike Ashby. Investing in new materials: a tool for technology managers. Technovation, 25(1):15–23, January 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000701.

Masucci:2021:HCV

[MPBC21] Monica Masucci, Simon C. Parker, Stefano Brusoni, and Roberto Camerani. How are corporate ventures evaluated and selected? Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497218307958.

Moncada-Paterno-Castello:2003:EIM

[MPCRB^{+}03] Pietro Moncada-Paternò-Castello, Jaime Rojo, Felix Bellido, Frédéric Fiore, and Alexander Tübke. Early identification and marketing of innovative technologies: a case study of RTD result valorisation at the European Commission’s Joint Research Centre. Technovation, 23(8):655–667, August 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497202000226.

Mu:2009:ERM

[MPM09] Jifeng Mu, Gang Peng, and Douglas L. MacLachlan. Effect of risk management strategy on NPD performance. Technovation, 29(3):170–180, March 2009. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208000904.

Miehe:2023:CSE

[MPO23] Lucas Miehé, Maximilian Palmié, and Pejvak Oghazi. Connection successfully established: How complementors use connectivity technologies to join existing ecosystems — Four archetype strategies from the mobility sector. Technovation, 122(??):??, April 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222002115.

Milesi:2013:IAM

[MPV13] Darío Milesi, Natalia Petelski, and Vladimiro Verre. Innovation and appropriation mechanisms: Evidence from Argentine microdata. Technovation, 33(2–3):78–87, February/ March 2013. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497212001435.

Miles:2003:MCE

[MPW03] M. P. Miles, C. W. Paul, and A. Wilhite. Modeling corporate entrepreneurship as rent-seeking competition. Technovation, 23(5):393–400, May 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202001566.

Murphree:2022:SST

[MPWG22] Michael Murphree, Bui Petersen, Peter Warrian, and Ray Gosine. Scope and scale of technology challenge and MNE subsidiary knowledge sourcing in host countries. Technovation, 116(??):??, August 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222000323.

Muffatto:2000:DPP

[MR00] Moreno Muffatto and Marco Roveda. Developing product platforms:: analysis of the development process. Technovation, 20(11):617–630, November 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001789.

Mohan:2005:STD

[MR05] S. Rama Mohan and A. Ramakrishna Rao. Strategy for technology development in public R&D institutes by partnering with the industry. Technovation, 25(12):1484–1491, December 2005. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497205000830.

Muscio:2019:HDA

[MR19] Alessandro Muscio and Laura Ramaciotti. How does academia influence Ph.D. entrepreneurship? New insights on the entrepreneurial university. Technovation, 82–83 (??):16–24, April/May 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497217307423.

Miorando:2014:EPM

[MRC14] Rogério Feroldi Miorando, José Luis Duarte Ribeiro, and Marcelo Nogueira Cortimiglia. An economic-probabilistic model for risk analysis in technological innovation projects. Technovation, 34(8):485–498, August 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http:

//www.sciencedirect.com/science/article/pii/S0166497214000030.

Merin-Rodriganez:2024:DTF

[MRDA24] Joan Merín-Rodrigáñez, Àngels Dasí, and Joaquín Alegre. Digital transformation and firm performance in innovative SMEs: the mediating role of business model innovation. Technovation, 134(??):??, June 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000774.

Martin-Rojas:2013:ITS

[MRGMBR13] Rodrigo Martín-Rojas, Víctor J. García-Morales, and María Teresa Bolívar-Ramos. Influence of technological support, skills and competencies, and learning on corporate entrepreneurship in European technology firms. Technovation, 33(12):417–430, December 2013. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213000813.

Martinez-Roman:2011:AIS

[MRGT11] Juan A. Martínez-Román, Javier Gamero, and Juan A. Tamayo. Analysis of innovation in SMEs using an innovative capability-based non-linear model: a study in the province of Seville (Spain). Technovation, 31(9):459–475, September 2011. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497211000757.

Martinez-Roman:2021:EAO

[MRGTDG21] Juan A. Martínez-Román, Javier Gamero, Juan A. Tamayo, and Loreto Delgado-González. Empirical analysis of organizational archetypes based on generation and adoption of knowledge and technologies. Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649721830066X.

Mrkajic:2017:BIM

[Mrk17] Boris Mrkajic. Business incubation models and institutionally void environments. Technovation, 68(??):44–55, December 2017. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497217300263.

Martlnez-Ros:2009:IAH

[MROS09] Ester Martlnez-Ros and Francina Orfila-Sintes. Innovation activity in the hotel industry. Technovation, 29(9):632–641, September 2009. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497209000431.

Millen:1998:PPA

[MS98] Robert Millen and Amrik S. Sohal. Planning processes for advanced manufacturing technology by large American manufacturers. Technovation, 18(12):741–750, October 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497298000686.

Momma:1999:DNB

[MS99] Stefan Momma and Margaret Sharp. Developments in new biotechnology firms in Germany. Technovation, 19(5):267– 282, February 1999. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497298001229.

Mitze:2019:DFI

[MS19] Timo Mitze and Falk Strotebeck. Determining factors of interregional research collaboration in Germany’s biotech network: Capacity, proximity, policy? Technovation, 80– 81(??):40–53, February/March 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://

www.sciencedirect.com/science/article/pii/S0166497216301419.

Maine:2015:RED

[MSD15] Elicia Maine, Pek-Hooi Soh, and Nancy Dos Santos. The role of entrepreneurial decision-making in opportunity creation and recognition. Technovation, 39–40(??):53–72, May/June 2015. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649721400025X.

Muller-Seitz:2010:NBS

[MSR10] Gordon Müller-Seitz and Guido Reger. Networking beyond the software code? An explorative examination of the development of an open source car project. Technovation, 30(11–12):627–634, November/December 2010. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649721000088X.

Mosterd:2021:CDT

[MSvdK^{+}21] Lars Mosterd, Vladimir C. M. Sobota, Geerten van de Kaa, Aaron Yi Ding, and Mark de Reuver. Context dependent trade-offs around platform-to-platform openness: the case of the Internet of Things. Technovation, 108(??):??, December 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221001127.

McMillan:2005:FSB

[MT05] G. Steven McMillan and Patrick Thomas. Financial success in biotechnology: company age versus company science. Technovation, 25(5):463–468, May 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001889.

Massa:2008:ISM

[MT08] Silvia Massa and Stefania Testa. Innovation and SMEs: Misaligned perspectives and goals among entrepreneurs, academics, and policy makers. Technovation, 28(7):393– 407, July 2008. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497208000126.

Macrii:2001:SLI

[MTB01] Diego Maria Macriì, Maria Rita Tagliaventi, and Fabiola Bertolotti. Sociometric location and innovation: how the social network intervenes between the structural position of early adopters and changes in the power map. Technovation, 21(1):1–13, January 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000183.

Moerchel:2023:VDI

[MTU23] Alexander Moerchel, Frank Tietze, and Florian Urmetzer. Visualising dynamics in innovation ecosystems: a new method and demonstration in the commercial aircraft MRO ecosystem. Technovation, 128(??):??, December 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001670.

Ma:2002:GMD

[MTY02] Qinhai Ma, Mitchell M. Tseng, and Benjamin Yen. A generic model and design representation technique of service products. Technovation, 22(1):15–39, January 2002. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000857.

Mudd:1990:PIM

[Mud90] Samuel Mudd. The place of innovativeness in models of the adoption process: an integrative review. Technovation, 10(2):119–136, March 1990. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729090032F.

Mudd:1995:KAI

[Mud95] Samuel Mudd. Kirton adaption-innovation theory: organizational implications. Technovation, 15(3):165–175, April 1995. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497295966173.

Mugnozza:1989:GSM

[Mug89] Giantommaso Scarascia Mugnozza. Giantommaso Scarascia Mugnozza. Technovation, 9(2–3):255–261, June 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900655.

Murray:1991:BRP

[Mur91] Hugh Murray. Book review: The portable MBA: Eliza G. C. Collins and Mary Anne Devanna, Wiley, New York. 1990, ISBN 0-471-61997-3, £20.95, 386 pages.

Technovation, 11(8):501–503, December 1991. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729190011R.

Musso:2009:NLO

[Mus09] Christopher Musso. New learning from old plastics: the effects of value-chain-complexity on adoption time. Technovation, 29(4):299–312, April 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497208001107.

Maartensson:2006:UMS

[MV06] Anders Mårtensson and Pablo Valiente. Understanding migration strategies by decoupling application roles and technology generations. Technovation, 26(3):351–358, March 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204002184.

Matos:2022:ICC

[MVS^{+}22] Stelvia Matos, Eric Viardot, Benjamin K. Sovacool, Frank W. Geels, and Yu Xiong. Innovation and climate change: a review and introduction to the special issue. Technovation, 117(??):??, September 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497222001596.

McQueen:1982:SCC

[MW82] D. H. McQueen and J. T. Wallmark. Spin-off companies from Chalmers University of Technology. Technovation, 1(4):305–315, August 1982. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497282900104.

Millson:2006:DNP

[MW06] Murray R. Millson and David Wilemon. Driving new product success in the electrical equipment manufacturing industry. Technovation, 26(11):1268–1286, November 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205001161.

Meyer:2009:MTD

[MW09] Patrick E. Meyer and James J. Winebrake. Modeling technology diffusion of complementary goods: the case of hydrogen vehicles and refueling infrastructure. Technovation, 29(2):77–91, February 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497208000667.

Mangematin:2012:FN

[MW12] Vincent Mangematin and Steve Walsh. The future of nanotechnologies. Technovation, 32(3–4):157–160, March/ April 2012. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497212000041.

Martinez:2014:WCP

[MW14] Marian Garcia Martinez and Bryn Walton. The wisdom of crowds: the potential of online communities as a tool for data analysis. Technovation, 34(4):203–214, April 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214000182.

Meyer-Waarden:2022:BYC

[MWC22] Lars Meyer-Waarden and Julien Cloarec. “Baby, you can drive my car”: Psychological antecedents that drive consumers’ adoption of AI-powered autonomous vehicles. Technovation, 109(??):??, January 2022. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001292.

Madhavan:2023:IVC

[MWJ23] Naveen Madhavan, Gareth R. T. White, and Paul Jones. Identifying the value of a clinical information system during the COVID-19 pandemic. Technovation, 120(??):??, February 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221002273.

McGrath:2002:NSA

[MY02] Robert N. McGrath and Seth B. Young. NASA’s small aircraft costs versus automobile costs and the economic value of traveler time. Technovation, 22(5):325–336, May 2002. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000785.

Mun:2021:FSB

[MYR^{+}21] Changbae Mun, Sejun Yoon, Nagarajan Raghavan, Dongwook Hwang, Subarna Basnet, and Hyunseok Park. Function score-based technological trend analysis. Technovation, 101(??):??, March 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497220300717.

Martinez:2017:CVA

[MZG17] Marian Garcia Martinez, Ferdaous Zouaghi, and Mercedes Sanchez Garcia. Capturing value from alliance portfolio diversity: the mediating role of R&D human capital in high and low tech industries. Technovation, 59(??):55– 67, January 2017. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497216301006.

McKelvey:2015:CIO

[MZL15] Maureen McKelvey, Olof Zaring, and Daniel Ljungberg. Creating innovative opportunities through research collaboration: an evolutionary framework and empirical illustration in engineering. Technovation, 39–40(??):26–36, May/ June 2015. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497214000698.

Narjes:1989:PEI

[Nar89] Karl-Heinz Narjes. Policies and experiences of international organizations for the promotion of enhanced interaction between industries, universities and other research organizations. Technovation, 9(2–3):241–248, June 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900631.

Narula:2004:RCS

[Nar04] Rajneesh Narula. R&D collaboration by SMEs: new opportunities and limitations in the face of globalisation. Technovation, 24(2):153–161, February 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000457.

Naylor:1988:SAS

[Nay88] Thomas H. Naylor. A strategic analysis of the Soviet Union. Technovation, 8(1–3):217–238, ???? 1988. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900624.

Niosi:1992:TAC

[NB92] Jorge Niosi and Maryse Bergeron. Technical alliances in the Canadian electronics industry: an empirical analysis. Technovation, 12(5):309–322, July 1992. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729290039K.

Narayanan:2009:TSD

[NB09] K. Narayanan and Savita Bhat. Technology sourcing and its determinants: a study of basic chemical industry in India. Technovation, 29(8):562–573, August 2009. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209000169.

Narayanan:2011:TSO

[NB11] K. Narayanan and Savita Bhat. Technology sourcing and outward FDI: a study of IT industry in India. Technovation, 31(4):177–184, April 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210000660.

Nylund:2022:ETM

[NBA22] Petra A. Nylund, Alexander Brem, and Nivedita Agarwal. Enabling technologies mitigating climate change: the role of dominant designs in environmental innovation ecosystems. Technovation, 117(??):??, September 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497221000523.

Nugent:2023:OUR

[NC23] Annita Nugent and Ho Fai Chan. Outsourcing university research commercialization to a sophisticated technology transfer office: Evidence from Australian universities. Technovation, 125(??):??, July 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000731.

Nepelski:2015:ITS

[ND15a] Daniel Nepelski and Giuditta De Prato. International technology sourcing between a developing country and the rest of the world. A case study of China. Technovation, 35 (??):12–21, January 2015. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497214001023.

Nokelainen:2015:TAU

[ND15b] Tomi Nokelainen and Ozgur Dedehayir. Technological adoption and use after mass market displacement: the case of the LP record. Technovation, 36–37(??):65–76, February/March 2015. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497214001436.

Navarro:2007:FAU

[NDP07] Juan Gabriel Cegarra Navarro, Frank W. Dewhurst, and Antonio Juan Briones Peñalver. Factors affecting the use of e-government in the telecommunications industry of Spain. Technovation, 27(10):595–604, October 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000387.

Neuhausler:2012:UPI

[Neu12] Peter Neuhäusler. The use of patents and informal appropriation mechanisms — differences between sectors and among companies. Technovation, 32(12):681–693, December 2012. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497212000806.

Newton:2004:BRT

[New04] Becci Newton. Book review: E-tivities: the key to active online learning: Gilly Salmon, Kogan Page: London. ISBN 0-7494-3686-7. Technovation, 24(1):83–84, January 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203001433.

Newbert:2013:HLW

[New13] Scott L. Newbert. Honoring the life’s work of Bruce A. Kirchhoff, Ph.D. Technovation, 33(8–9):247–248, August/ September 2013. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497213000709.

Narduzzo:2024:NIT

[NF24] Alessandro Narduzzo and Valentina Forrer. Nurturing innovation through intelligent failure: the art of failing on purpose. Technovation, 131(??):??, March 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000014.

Negri:1997:QPV

[NG97] Lionello Negri and Massimiliano Galli. Quality policies and value creation strategies in Italian manufacturing industry. Technovation, 17(6):321–328, June 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497296001307.

Naranjo-Gil:2009:IEO

[NG09] David Naranjo-Gil. The influence of environmental and organizational factors on innovation adoptions: Consequences for performance in public sector organizations. Technovation, 29(12):810–818, December 2009. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209001059.

Newbert:2008:LBR

[NGK08] Scott L. Newbert, Shanthi Gopalakrishnan, and Bruce A. Kirchhoff. Looking beyond resources: Exploring the importance of entrepreneurship to firm-level competitive advantage in technologically intensive industries. Technovation, 28(1–2):6–19, January/February 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000910.

Narula:1999:ITS

[NH99] Rajneesh Narula and John Hagedoorn. Innovating through strategic alliances: moving towards international partnerships and contractual agreements. Technovation, 19(5):283– 294, February 1999. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497298001278.

Nafari:2024:PAS

[NHS24] Javid Nafari, Benson Honig, and Ana Cristina O. Siqueira. Promoting academic social intrapreneurship: Developing an international virtual incubator and fostering social impact. Technovation, 133(??):??, May 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000749.

Nguyen:2023:IPP

[NHT23] Thi Phuong Thao Nguyen, Fang Huang, and Xiaowen Tian. Intellectual property protection need as a driver for open innovation: Empirical evidence from Vietnam. Technovation, 123(??):??, May 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000251.

Nijssen:2005:UWC

[NHV05] Edwin J. Nijssen, Bas Hillebrand, and Patrick A. M. Vermeulen. Unraveling willingness to cannibalize: a closer look at the barrier to radical innovation. Technovation, 25(12): 1400–1409, December 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497205000660.

Numprasertchai:2005:MKT

[NI05] Somchai Numprasertchai and Barbara Igel. Managing knowledge through collaboration: multiple case studies of managing research in university laboratories in Thailand. Technovation, 25(10):1173–1182, October 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204000550.

Ng:2004:DBR

[NJ04] Ngai Kheong Ng and Jianxin Jiao. A domain-based reference model for the conceptualization of factory loading allocation problems in multi-site manufacturing supply chains. Technovation, 24(8):631–642, August 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202001256.

Nakagawa:2022:WSI

[NK22] Koichi Nakagawa and Genjiro Kosaka. What social issues do people invest in? An examination based on the empathy-altruism hypothesis of prosocial crowdfunding platforms. Technovation, 114(??):??, June 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497222000554.

Nijhof:2002:ITE

[NKL02] A. Nijhof, K. Krabbendam, and J. C. Looise. Innovation through exemptions: building upon the existing creativity of employees. Technovation, 22(11):675–683, November 2002. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201000888.

Niedergassel:2011:DDK

[NL11] Benjamin Niedergassel and Jens Leker. Different dimensions of knowledge in cooperative R&D projects of university scientists. Technovation, 31(4):142–150, April 2011. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210001136.

Nieto:1998:PAT

[NLC98] Mariano Nieto, Francisco Lopéz, and Fernando Cruz. Performance analysis of technology using the S curve model: the case of digital signal processing (DSP) technologies. Technovation, 18(6–7):439–457, January 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000212.

Ng:2004:SPF

[NLLC04] P. T. Andy Ng, Dawei Lu, C. K. Li, and H. Y. Harry Chan. A strategic perspective on future residential broadband market. Technovation, 24(8):665–669, August 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204000513.

Ng:2005:SLV

[NLLC05] P. T. Andy Ng, Dawei Lu, C. K. Li, and H. Y. Harry Chan. Strategic lessons of value migration in IT industry. Technovation, 25(1):45–51, January 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001282.

Nair:2011:DBP

[NMN11] Shyam Sreekumaran Nair, Mary Mathew, and Dipanjan Nag. Dynamics between patent latent variables and patent price. Technovation, 31(12):648–654, December 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211000940.

Nobelius:2004:LPD

[Nob04] Dennis Nobelius. Linking product development to applied research: transfer experiences from an automotive company. Technovation, 24(4):321–334, April 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000731.

Nikulainen:2010:TSB

[NP10] Tuomo Nikulainen and Christopher Palmberg. Transferring science-based technologies to industry-Does nanotechnology make a difference? Technovation, 30(1):3–11, January 2010. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209001084.

Natalicchio:2014:LRM

[NPG14] A. Natalicchio, A. Messeni Petruzzelli, and A. C. Garavelli. A literature review on markets for ideas: Emerging characteristics and unanswered questions. Technovation, 34(2):65– 76, February 2014. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497213001430.

Natalicchio:2017:IPS

[NPG17] A. Natalicchio, A. Messeni Petruzzelli, and A. C. Garavelli. Innovation problems and search for solutions in crowdsourcing platforms — a simulation approach. Technovation, 64–65(??):28–42, June 2017. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497217303449.

Ndonzuau:2002:SMA

[NPS02] Frédéric Nlemvo Ndonzuau, Fabrice Pirnay, and Bernard Surlemont. A stage model of academic spin-off creation. Technovation, 22(5):281–289, May 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000190.

Nosella:2005:CIB

[NPV05] Anna Nosella, Giorgio Petroni, and Chiara Verbano. Characteristics of the Italian biotechnology industry and new business models: the initial results of an empirical study. Technovation, 25(8):841–855, August 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204000276.

Nieto:2005:ACT

[NQ05] Mariano Nieto and Pilar Quevedo. Absorptive capacity, technological opportunity, knowledge spillovers, and innovative effort. Technovation, 25(10):1141–1157, October 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204000902.

Napolitano:1990:PSE

[NS90] Giovanni Napolitano and Giorgio Sirilli. The patent system and the exploitation of inventions: results of a statistical survey conducted in Italy. Technovation, 10(1):5–16, February 1990. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729090038L.

Nieto:2007:IDC

[NS07] María Jesús Nieto and Lluis Santamaría. The importance of diverse collaborative networks for the novelty of product innovation. Technovation, 27(6–7):367–377, June/July 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497206001064.

Nieto:2023:DFO

[NSB23] María Jesús Nieto, Luis Santamaria, and Yannick Bammens. Digitalization as a facilitator of open innovation: Are family firms different? Technovation, 128(??):??, December 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223001657.

Noh:2018:HIT

[NSYL18] Heeyong Noh, Ju-Hwan Seo, Hyoung Sun Yoo, and Sungjoo Lee. How to improve a technology evaluation model: A data-driven approach. Technovation, 72–73 (??):1–12, April/May 2018. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497217307824.

Nordman:2016:IOC

[NT16] Emilia Rovira Nordman and Daniel Tolstoy. The impact of opportunity connectedness on innovation in SMEs’ foreign-market relationships. Technovation, 57–58(??):47– 57, November/December 2016. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016649721630027X.

Nasiri:2021:MDS

[NUSR21] Mina Nasiri, Juhani Ukko, Minna Saunila, and Tero Rantala. Managing the digital supply chain: the role of smart technologies. Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497220300110.

Nuvolari:2007:BRK

[Nuv07] Alessandro Nuvolari. Book review: Knowledge and Competitive Advantage: the Coevolution of Firms, Technology and National Institutions: J. P. Murmann. Cambridge University Press, New York (2003), ISBN 0-521-68415-

3. Technovation, 27(6–7):414–415, June/July 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720700020X.

Nazarenko:2022:ADT

[NVMD22] Anastasia Nazarenko, Konstantin Vishnevskiy, Dirk Meissner, and Tugrul Daim. Applying digital technologies in technology roadmapping to overcome individual biased assessments. Technovation, 110(??):??, February 2022. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001450.

Nakamura:2003:MEM

[NW03a] Yoshiaki Nakamura and Chihiro Watanabe. Management and the effect of MITI’s R&D project: case study from a supercomputer project. Technovation, 23(3):221–238, March 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201000852.

Nambisan:2003:GSG

[NW03b] Satish Nambisan and David Wilemon. A global study of graduate management of technology programs. Technovation, 23(12):949–962, December 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0166497203001093.

Naveed:2023:ILP

[NW23] Amjad Naveed and Cong Wang. Innovation and labour productivity growth moderated by structural change: Analysis in a global perspective. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222001018.

Nakagawa:2009:CTS

[NWGB09] Masahiro Nakagawa, Chihiro Watanabe, and Charla Griffy-Brown. Changes in the technology spillover structure due to economic paradigm shifts: a driver of the economic revival in Japan’s material industry beyond the year 2000. Technovation, 29(1):5–22, January 2009. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208001077.

Nwoye:1997:RPS

[Nwo97] May Ifeoma Nwoye. The role of the private sector in the promotion of young entrepreneurship in Nigeria. Technovation, 17(9):521–528, 533, September 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297000266.

Nwoye:2002:FGD

[Nwo02] May Ifeoma Nwoye. A focus group discussion approach to the comparative analysis of private and public sector enterprises in Nigeria. Technovation, 22(8):525–534, August 2002. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497200000869.

Oakley:1997:BRF

[Oak97] Paul Oakley. Book review: Finance and technological change: Theory and evidence: by Enrico Santarelli, Macmillan Press Ltd, 1995, 181 + x pp. (hardback), £40. Technovation, 17(1):51, January 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297842228.

Oakey:2003:TEH

[Oak03] R. P. Oakey. Technical entreprenenurship in high technology small firms: some observations on the implications for management. Technovation, 23(8):679–688, August 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000452.

Ortin-Angel:2014:USO

[OÁVH14] Pedro Ortín-Ángel and Ferran Vendrell-Herrero. University spin-offs vs. other NTBFs: Total factor productivity differences at outset and evolution. Technovation, 34(2):101–112, February 2014. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497213001119.

Ottosson:2004:RDS

[OB04] Stig Ottosson and Evastina Björk. Research on dynamic systems-some considerations. Technovation, 24(11):863– 869, November 2004. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203000324.

Onufrey:2020:SWE

[OB20] Ksenia Onufrey and Anna Bergek. Second wind for exploitation: Pursuing high degrees of product and process innovativeness in mature industries. Technovation, 89(??): ??, January 2020. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497216305156.

Onufrey:2021:TMI

[OB21] Ksenia Onufrey and Anna Bergek. Transformation in a mature industry: the role of business and innovation strategies. Technovation, ??(??):??, ???? 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497220300626.

Oztaysi:2009:RFI

[ÖBA09] Basar Öztaysi, Serdar Baysan, and Fatma Akpinar. Radio frequency identification (RFID) in hospitality. Technovation, 29(9):618–624, September 2009. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209000807.

Obel:1986:SCC

[Obe86] Børge Obel. SAS: Changes in competition, strategy, and organization. Technovation, 5(1–3):145–153, October 1986. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497286900490.

Obel:1988:DBI

[Obe88] Børge Obel. Danish business and industry in the nineties: Opportunities and threats. Technovation, 8(1–3):133–142, ???? 1988. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497288900570.

Oberg:2023:TTS

[Öbe23] Christina Öberg. Towards a typology of sharing economy business model transformation. Technovation, 123(??):??, May 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223000330.

Olson:2014:IRU

[OBL14] David L. Olson, John R. Birge, and Jonathan Linton. Introduction to risk and uncertainty management in technological innovation. Technovation, 34(8):395–398, August 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214000649.

Opazo-Basaez:2024:EEG

[OBMOM24] Marco Opazo-Basáez, Juan Carlos Monroy-Osorio, and Josip Marić. Evaluating the effect of green technological innovations on organizational and environmental performance: a treble innovation approach. Technovation, 129(??):??, January 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497223001967.

Omta:1997:MCB

[OBvE97] S. W. F. Omta, L. M. Bouter, and J. M. L. van Engelen. Management control of biomedical research and pharmaceutical innovation. Technovation, 17(4):167–179, 221, April 1997. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497296000995.

Olshfski:1985:IME

[OC85] Dorothy Olshfski and Robert Cunningham. Improving management effectiveness by training — the use of video techniques to assist self appraisal. Technovation, 3(4):235– 242, October 1985. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497285900227.

Oakey:1991:RBP

[OC91] R. P. Oakey and S. Y. Cooper. The relationship between product technology and innovation performance in high technology small firms. Technovation, 11(2):79–92, March 1991. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497291900397.

Oh:2009:ATI

[OCA09] Ka-Young Oh, Doug Cruickshank, and Alistair R. Anderson. The adoption of e-trade innovations by Korean small and medium sized firms. Technovation, 29(2):110– 121, February 2009. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497208001065.

Ostergaard:2022:KTW

[ØD22] Christian Richter Østergaard and Ina Drejer. Keeping together: Which factors characterise persistent universityindustry collaboration on innovation? Technovation, 111 (??):??, March 2022. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649722100170X.

Oduro:2023:DDT

[ODM23] Stephen Oduro, Alessandro De Nisco, and Giada Mainolfi. Do digital technologies pay off? A meta-analytic review of the digital technologies/firm performance nexus. Technovation, 128(??):??, December 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497223001475.

Omsen:1987:INI

[OE87]  Arne H. Omsén and Swen Ekeström.  Implanting

new industrial product ideas into operating companies. Technovation, 7(1):23–37, December 1987. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497287900459.

Oftedal:2024:PVN

[OFR24] Elin M. Oftedal, Lene Foss, and Matthew W. Rutherford. From propriety to validity in new ventures: a nine-year study of three startup companies. Technovation, 131(??): ??, March 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223002602.

ORegan:2005:SPC

[OG05] Nicholas O’Regan and Abby Ghobadian. Strategic planning — a comparison of high and low technology manufacturing small firms. Technovation, 25(10):1107–1117, October 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000919.

ORegan:2006:SDH

[OGG06] Nicholas O’Regan, Abby Ghobadian, and David Gallear. In search of the drivers of high growth in manufacturing SMEs. Technovation, 26(1):30–41, January 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000775.

OLeary:2022:ICC

[OGOF22] Kevin O’Leary, Rob Gleasure, Philip O’Reilly, and Joseph Feller. Introducing the concept of creative ancestry as a means of increasing perceived fairness and satisfaction in online collaboration: an experimental study. Technovation, 110(??):??, February 2022. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001504.

Ogink:2023:MOI

[OGRA23] Ruben H. A. J. Ogink, Martin C. Goossen, A. Georges L. Romme, and Henk Akkermans. Mechanisms in open innovation: a review and synthesis of the literature. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222001687.

ORegan:2006:FTI

[OGS06] Nicholas O’Regan, Abby Ghobadian, and Martin Sims. Fast tracking innovation in manufacturing SMEs. Technovation, 26(2):251–261, February 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497205000040.

OHara:1999:BRR

[O’H99] Suzanne O’Hara. Book review: Rapid Problem Solving with Post-it(R)Notes: David Straker; Gower Publishing, Hampshire, England, 1997, 153 pages, hardback, ISBN 0-566-07836-8, £29.95. Technovation, 19(10):650, October 1999. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299000784.

Oh:2004:BRB

[OH04] Tae-Seog Oh and Mike Hobday. Book review: Beyond late development: Taiwan’s upgrading policies: Alice H. Amsden and Wan-wen Chu, The MIT Press, 2003, 207 pages, ISBN 0-262-01198-0. Technovation, 24(2):176–177, February 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203001755.

Oyebisi:1996:IAR

[OIN96] T. O. Oyebisi, M. O. Ilori, and M. L. Nassar. Industryacademic relations: an assessment of the linkages between a university and some enterprises in Nigeria. Technovation, 16(4):203–209, 215, April 1996. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497295000496.

Ojewale:2001:IAR

[OIOA01] B. A. Ojewale, M. O. Ilori, T. O. Oyebisi, and I. O. Akinwumi. Industry-academic relation: utilization of idle capacities in polytechnics, universities and research organizations by entrepreneurs in Nigeria. Technovation, 21(10):695–704, October 2001. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497200000419.

OKane:2019:TTE

[O’K19] Conor O’Kane. Technology transfer executives’ backwards integration: an examination of interactions between university technology transfer executives and principal investigators. Technovation, 76–77(??):64–77, August/September 2019. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497216302309.

Oerlemans:2013:APD

[OKP13] L. A. G. Oerlemans, J. Knoben, and M. W. Pretorius. Alliance portfolio diversity, radical and incremental innovation: the moderating role of technology management. Technovation, 33(6–7):234–246, June/July 2013. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213000229.

Onsongo:2023:FBM

[OKvB23] Elsie K. Onsongo, Peter Knorringa, and Cees van Beers. Frugal business model innovation in the Base of the Pyramid: the case of Philips Community Life Centres in Africa. Technovation, 121(??):??, March 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222002267.

Oladeji:1998:TPDa

[Ola98a] S. I. Oladeji. Technology policy and the development of small and medium scale enterprises in contemporary Nigeria. Technovation, 18(2):125–132, February 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297000825.

Oladeji:1998:TPDb

[Ola98b] S. I. Oladeji. Technology policy and the development of small and medium scale enterprises in contemporary Nigeria. Technovation, 18(2):144–145, February 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297001065.

Oldsman:1997:MEC

[Old97] Eric S. Oldsman. Manufacturing extension centers and private consultants: collaboration or competition? Technovation, 17(5):237–243, May 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497296000983.

Olleros:2008:LCD

[Oll08] Xavier Olleros. The lean core in digital platforms. Technovation, 28(5):266–276, May 2008. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497207000934.

Olleros:1988:SAM

[OM88] Francisco-Javier Olleros and Roderick J. Macdonald. Strategic alliances: managing complementarity to capitalize on emerging technologies. Technovation, 7(2):155–176, May 1988. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497288900454.

Onida:1989:RCB

[OM89] Fabrizio Onida and Franco Malerba. R&D cooperation between industry, universities and research organizations in Europe. Technovation, 9(2–3):137–195, June 1989. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900552.

Olsson:2000:FIP

[OM00] Henrik Olsson and Douglas H. McQueen. Factors influencing patenting in small computer software producing companies. Technovation, 20(10):563–576, October 2000. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001741.

Obwegeser:2018:IPP

[OM18] Nikolaus Obwegeser and Sune Dueholm Müller. Innovation and public procurement: Terminology, concepts, and applications. Technovation, 74–75(??):1–17, June/ July 2018. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497218301299.

Odgers:1988:ATI

[ON88] John F. Odgers and Norbert Nimmervoll. Accounting for technological innovation: an overview. Technovation, 7(2): 117–129, May 1988. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497288900429.

Onida:1989:FO

[Oni89] Fabrizio Onida. Fabrizio Onida. Technovation, 9(2–3):287– 288, June 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497289900709.

Olajide:2002:PES

[OO02] J. O. Olajide and O. J. Oyelade. Performance evaluation of the Strategic Grain Reserve Storage Programme (SGRSP) in Nigeria. Technovation, 22(7):463–468, July 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000153.

Oyelaran-Oyeyinka:2004:DAU

[OOA04] B. Oyelaran-Oyeyinka and C. N. Adeya. Dynamics of adoption and usage of ICTs in African universities: a study of Kenya and Nigeria. Technovation, 24(10):841–851, October 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497202001700.

Osunbitan:2000:IME

[OOJ^{+}00] J. A. Osunbitan, J. O. Olushina, J. O. Jeje, K. A. Taiwo, M. O. Faborode, and O. O. Ajibola. Information on micro-enterprises involved in cassava and palm oil processing in the Osun and Ondo states of Nigeria. Technovation, 20(10):577–585, October 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001790.

Oehmen:2014:AER

[OOKBD14] Josef Oehmen, Alison Olechowski, C. Robert Kenley, and Mohamed Ben-Daya. Analysis of the effect of risk management practices on the performance of new product development programs. Technovation, 34(8):441–453, August 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213001612.

Oyelaran-Oyeyinka:2006:LNT

[OOL06] Banji Oyelaran-Oyeyinka and Kaushalesh Lal. Learning new technologies by small and medium enterprises in developing countries. Technovation, 26(2):220–231, February 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204001294.

Ohmura:2003:FDS

[OOMW03] A. Ohmura, N. Ouchi, S. Morisaki, and C. Watanabe. Functionality development as a survival strategy for fine ceramics. Technovation, 23(10):833–842, October 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000287.

Oh:2016:IEC

[OPPL16] Deog-Seong Oh, Fred Phillips, Sehee Park, and Eunghyun Lee. Innovation ecosystems: a critical examination. Technovation, 54(??):1–6, August 2016. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497216300062.

Okudan:2006:PBA

[OR06] Gül E. Okudan and Sarah E. Rzasa. A projectbased approach to entrepreneurial leadership education. Technovation, 26(2):195–210, February 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204002044.

Orr:1997:AWA

[Orr97] Stuart C. Orr. Automation in the workplace: an Australasian perspective. Technovation, 17(2):83–89, 103–105, February 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497296000648.

Orr:1998:GMI

[OS98] Stuart Orr and Amrik S. Sohal. Global manufacturing issues: the case of Siemens AG. Technovation, 18(10):605–614, 656–657, October 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000248.

ONeill:1999:BPR

[OS99] Peter O’Neill and Amrik S. Sohal. Business process reengineering: a review of recent literature. Technovation, 19 (9):571–581, September 1999. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497299000590.

ORegan:2008:IHT

[OS08]  Nicholas O’Regan and Martin A. Sims.  Identify-

ing high technology small firms: a sectoral analysis. Technovation, 28(7):408–423, July 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208000308.

ORegan:2004:IMT

[OSG04] Nicholas O’Regan, Martin Sims, and Abby Ghobadian. The impact of management techniques on performances in technology-based firms. Technovation, 24(3):265–273, March 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203001287.

Ozcan:2021:SMM

[OSSC21] Sercan Ozcan, Metin Suloglu, C. Okan Sakar, and Sushant Chatufale. Social media mining for ideation: Identification of sustainable solutions and opinions. Technovation, 107(??):??, September 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497221001036.

Ogwu:2001:AHK

[OTA01] N. E. Ogwu, K. A. Taiwo, and O. O. Ajibola. Assessment of household kitchen equipment: factors influencing its acquisition and its impact on food consumption in Osun State of Nigeria. Technovation, 21(9):613–621, September 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000511.

Ottosson:1997:PYO

[Ott97] Stig Ottosson. Pilot your organization using cash flow planning with frequent feedback. Technovation, 17(9):513–520, 532–533, September 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016649729700031X.

Ottosson:1998:PO

[Ott98a] Stig Ottosson. Planetary organizations. Technovation, 19 (2):81–86, December 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016649729800100X.

Ottosson:1998:SCI

[Ott98b] Stig Ottosson. Strategic considerations of the interplay between R&D and M&S. Technovation, 18(4):235–244, April 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297001223.

Ottosson:2003:DPD

[Ott03a] Stig Ottosson. Dynamic product development of a new Intranet platform. Technovation, 23(8):669–678, August 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000476.

Ottosson:2003:PAR

[Ott03b] Stig Ottosson. Participation action research-: a key to improved knowledge of management. Technovation, 23(2):87– 94, February 2003. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497201000979.

Ottosson:2004:DIP

[Ott04a] S. Ottosson. Dealing with innovation push and market need. Technovation, 24(4):279–285, April 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000780.

Ottosson:2004:DPD

[Ott04b] S. Ottosson. Dynamic product development — DPD. Technovation, 24(3):207–217, March 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0166497202000779.

Obradovic:2021:OIM

[OVD21] Tena Obradović, Bozidar Vlacić, and Marina Dabić. Open innovation in the manufacturing industry: a review and research agenda. Technovation, 102(??):??, April 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722100002X.

Overholm:2015:CCO

[Ove15] Harald Overholm. Collectively created opportunities in emerging ecosystems: the case of solar service ventures.

Technovation, 39–40(??):14–25, May/June 2015. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214000091.

Oakey:1993:BIRa

[OW93a] R. P. Oakey and T. White. Business information and regional economic development: some conceptual observations. Technovation, 13(3):147–159, April 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390037V.

Oakey:1993:BIRb

[OW93b] R. P. Oakey and T. White. Business information and regional economic development: some conceptual observations. Technovation, 13(3):180, April 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390041S.

Orihata:2000:EDP

[OW00a] Motokazu Orihata and Chihiro Watanabe. Evolutional dynamics of product innovation: the case of consumer electronics. Technovation, 20(8):437–449, August 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001522.

Orihata:2000:IBP

[OW00b] Motokazu Orihata and Chihiro Watanabe. The interaction between product concept and institutional inducement: a new driver of product innovation. Technovation, 20(1):11– 23, January 2000. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497299001078.

Ong:2003:FAI

[OWC03] Chin-Huat Ong, David Wan, and Seong-Huat Chng. Factors affecting individual innovation: an examination within a Japanese subsidiary in Singapore. Technovation, 23(7): 617–631, July 2003. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497202000196.

Ooms:2015:ROA

[OWCvdB15] Ward Ooms, Claudia Werker, Marjolein C. J. Caniëls, and Herman van den Bosch. Research orientation and agglomeration: Can every region become a Silicon Valley? Technovation, 45–46(??):78–92, November/December 2015. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497215000589.

Oyebisi:2000:RMM

[Oye00] T. O. Oyebisi. On reliability and maintenance management of electronic equipment in the tropics. Technovation, 20 (9):517–522, September 2000. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497299001662.

Oyebisi:2001:ICS

[Oye01] T. O. Oyebisi. The Internet and the commercial sector of the Nigerian economy: policy and technology management implications. Technovation, 21(6):369–374, June 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720000064X.

Okudan:2006:TPD

[OZ06] Gül E. Okudan and Sarah E. Zappe. Teaching product design to non-engineers: a review of experience, opportunities and problems. Technovation, 26(11):1287–1293, November 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205001653.

Persson:2006:MIM

[PÅ06] Magnus Persson and Pär Åhlström. Managerial issues in modularising complex products. Technovation, 26(11): 1201–1209, November 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497205001525.

Patterson:2015:CAC

[PA15] William Patterson and Véronique Ambrosini. Configuring absorptive capacity as a key process for research intensive firms. Technovation, 36–37(??):77–89, February/ March 2015. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497214001382.

Paff:2004:DAI

[Paf04] Lolita A. Paff. Does the alternate incremental credit affect firm R&D? Technovation, 24(1):41–52, January 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203001445.

Palairet:1993:MIYa

[Pal93a] Michael Palairet. Mismanaging innovation: the Yugo car enterprise (1962–1992). Technovation, 13(3):117–132, April 1993. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729390035T.

Palairet:1993:MIYb

[Pal93b] Michael Palairet. Mismanaging innovation: the Yugo car enterprise (1962–1992). Technovation, 13(3):177–178, April 1993. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729390039X.

Palmer:2000:IAB

[Pal00] Janet J. Palmer. Internet access in Bahrain: business patterns and problems. Technovation, 20(8):451–458, August 2000. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497299001819.

Palmai:2004:IPH

[Pál04] Z. Pálmai. An innovation park in Hungary: INNOTECH of the Budapest University of Technology and Economics. Technovation, 24(5):421–432, May 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000986.

Palmberg:2006:SSI

[Pal06] Christopher Palmberg. The sources and success of innovations — determinants of commercialisation and break-even times. Technovation, 26(11):1253–1267, November 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205001513.

Pandey:1998:PDV

[Pan98] I. M. Pandey. The process of developing venture capital in India. Technovation, 18(4):253–261, April 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000030.

Papadakis:1995:DTL

[Pap95] Maria Papadakis. The delicate task of linking industrial R&D to national competitiveness. Technovation, 15(9):569– 583, November 1995. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729596586I.

Papaioannou:2002:BRK

[Pap02a] Theodoros Papaioannou. Book review: Knowledge, innovation and economic growth: the theory and practise of learning regions: F. Boekema, K. Morgan, S. Bakkers and R. Rutten (Eds.), Edward Elgar, Cheltenham, 2000. Technovation, 22(5):337–338, May 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000700.

Papaioannou:2002:BRM

[Pap02b] Theodoros Papaioannou. Book review: Market-oriented technology management: innovating for profit in entrepreneurial times: F.Y Phillips; Springer-Verlag. Heidelberg, 2001, pp. 417 price £44.50. Technovation, 22(5):338– 339, May 2002. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497201001134.

Papaioannou:2003:BRN

[Pap03] T. Papaioannou. Book review: New technology-based firms in the new millennium: W. During, R. Oakey and S. Kauser (eds.), Pergamon, Oxford 2001. Technovation, 23(7):653– 654, July 2003. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203000257.

Papaioannou:2004:BRG

[Pap04] Theodoros Papaioannou. Book review: The globalising learning economy: D. Archibugi and B.-A. Lundvall (eds), Oxford University Press, Oxford 2002. Technovation, 24 (3):275–277, March 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497203001998.

Papaioannou:2006:BRR

[Pap06] Theodoros Papaioannou. Book review: Regionalisation of Globalised Innovation: Locations for Advanced Industrial Development and Disparities in Participation: U. Hilpert (Ed.). Routledge, London and New York (2003). Technovation, 26(3):416–417, March 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205001318.

Park:2000:NSA

[Par00] Yong-Tae Park. National systems of Advanced Manufacturing Technology (AMT): hierarchical classification scheme and policy formulation process. Technovation, 20(3):151– 159, March 2000. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497299001030.

Pare:2003:BRS

[Par03a] D. J. Paré. Book review: Strategic computing: DARPA and the quest for machine intelligence, 1983–1993: Alex Roland and Philip Shiman (2002). London, MIT Press. pp. 440, ISBN 0-262-18226-2, £33.50. Technovation, 23(8):717, August 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203000269.

Parramore:2003:BRT

[Par03b] K. Parramore. Book review: Technology investment: a game theoretic real options approach: Kuno J. M. Huisman, Kluwer Academic Publishers. Technovation, 23(11): 901, November 2003. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203000671.

Park:2005:ORP

[Par05] John S. Park. Opportunity recognition and product innovation in entrepreneurial hi-tech start-ups: a new perspective and supporting case study. Technovation, 25(7):739–752, July 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204000045.

Pare:2006:BRG

[Par06] Daniel J. Pare. Book review: The Global Internet Economy, Bruce Kogut (Ed.). MIT Press, London (2003),

520. Technovation, 26(7):893–894, July 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000842.

Parker:2007:BRT

[Par07] Mike J. Parker. Book review: Time Strategies, Innovation and Environmental Policy, C. Sartorius, S. Zundel (Eds.). Edward Elgar, Cheltenham (2005). Technovation, 27(4): 233, April 2007. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497206000915.

Patsali:2024:UPL

[Pat24] Sofia Patsali. University procurement-led innovation: Sources, procedures, and effects. Some field-study evidence. Technovation, 130(??):??, February 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223002122.

Paulinyi:1993:TDB

[Pau93] Erno I. Paulinyi. Technological development in Brazilian business: four behavioral types. Technovation, 13(2):83–92, March 1993. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497293900562. See [Ano93n].

Pavon:1989:JP

[Pav89] Julian Pavon. Julian Pavon. Technovation, 9(2–3):289–291, June 1989. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497289900710.

Pawan:1994:GTB

[Paw94] Sikka Pawan. GATT and the technology base in India: Technovation, 14(9) (1994) 613–620. Technovation, 14 (9):623–624, November 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497294900477.

See [Sik94].

Pearson:1993:FMC

[PB93] A. W. Pearson and D. F. Ball. A framework for managing communication at the R&D/marketing interface. Technovation, 13(7):439–447, November 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390027S. See [Ano93f].

Pimenta-Bueno:1993:NLU

[PBC93] J. A. Pimenta-Bueno and R. T. S. Corrêa. A new look at the US pattern of federal R&D spending: preliminary lessons for newly industrializing economies. Technovation, 13(2):101–107, March 1993. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497293900584.

See comment [Ano93l].

Parrilli:2023:EIA

[PBOR23] M. Davide Parrilli, Merima Balavac-Orlić, and Dragana Radicic. Environmental innovation across SMEs in Europe. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497222000888.

Pinto:1989:CFP

[PC89] Jeffrey K. Pinto and Jeffrey G. Covin. Critical factors in project implementation: a comparison of construction and R&D projects. Technovation, 9(1):49–62, May 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900400.

Perunovic:2005:EDI

[PC05] Zoran Perunovic and Thomas B. Christiansen. Exploring Danish innovative manufacturing performance. Technovation, 25(9):1051–1058, September 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204000392.

Paskaleva:2018:OIE

[PC18] Krassimira Paskaleva and Ian Cooper. Open innovation and the evaluation of Internet-enabled public services in smart cities. Technovation, 78(??):4–14, December 2018. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497218304711.

Paskaleva:2021:LLE

[PC21] Krassimira Paskaleva and Ian Cooper. Are living labs effective? Exploring the evidence. Technovation, 106(??): ??, August 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497221000924.

Pan:2022:IPS

[PCG22] Xia Pan, Wenyin Cheng, and Yuning Gao. The impact of privatization of state-owned enterprises on innovation in China: a tale of privatization degree. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222001341.

Pan:2024:PBO

[PCQ24] Xin Pan, Xuanjin Chen, and Shumin Qiu. Pushing boundaries or overstepping? Exploring the paradoxical impact of radical innovation on government subsidies in Chinese SMEs. Technovation, 132(??):??, April 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000385.

Pun:2004:DMS

[PCWG04] K.-F. Pun, K.-S. Chin, A. S. White, and R. Gill. Determinants of manufacturing strategy formulation: a longitudinal study in Hong Kong. Technovation, 24(2):121–137, February 2004. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497202000500.

Pauwels:2016:UNG

[PCWV16] Charlotte Pauwels, Bart Clarysse, Mike Wright, and Jonas Van Hove. Understanding a new generation incubation model: the accelerator. Technovation, 50–51(??):13–24, April/May 2016. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497215000644.

Prodan:2010:CAE

[PD10] Igor Prodan and Mateja Drnovsek. Conceptualizing academic-entrepreneurial intentions: an empirical test. Technovation, 30(5–6):332–347, May/June 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210000313.

Plantec:2023:BDE

[PDHW23] Quentin Plantec, Marie-Alix Deval, Sophie Hooge, and Benoit Weil. Big data as an exploration trigger or problemsolving patch: Design and integration of AI-embedded systems in the automotive industry. Technovation, 124(??):??, June 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223000743.

Philpott:2011:EUE

[PDOL11] Kevin Philpott, Lawrence Dooley, Caroline O’Reilly, and Gary Lupton. The entrepreneurial university: Examining the underlying academic tensions. Technovation, 31(4):161– 170, April 2011. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497210001331.

Pretorius:2000:MAN

[PdW00] M. W. Pretorius and G. de Wet. A model for the assessment of new technology for the manufacturing enterprise. Technovation, 20(1):3–10, January 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299000929.

Pushpananthan:2022:JFC

[PE22] Gouthanan Pushpananthan and Maria Elmquist. Joining forces to create value: the emergence of an innovation ecosystem. Technovation, 115(??):??, July 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497221002340.

Pearson:1990:IS

[Pea90] Alan W. Pearson. Innovation strategy. Technovation, 10(3): 185–192, May 1990. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729090023D.

Purdy:2023:TSD

[PEER23] Lauren Purdy, Hadi Eslami, Kamran Eshghi, and Michel Rod. Technology sourcing and the dark side of open innovation: Evidence from the biopharmaceutical sector. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222000682.

Peneder:2008:PPU

[Pen08] Michael Peneder. The problem of private under-investment in innovation: a policy mind map. Technovation, 28(8):518– 530, August 2008. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497208000199.

Perren:1995:BRB

[Per95] Lew Perren. Book review: Beyond dispute — The invention of team syntegrity: by Stafford Beer. John Wiley & Sons Ltd, Baffins Lane, Chichester, West Sussex PO19 1UD, UK, 1993, 367 pp., ISBN 0-471-94451-3. £24.95 (hardback). Technovation, 15(3):189, April 1995. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729590025X.

Perren:1996:BRE

[Per96] Lew Perren. Book review: European casebook on: Competing through information technology — strategy and implementation: by Tawfik Jelassi. Prentice Hall International (UK) Ltd, Hemel Hempstead, 1994, 298 pp., ISBN 0-13-295304-8, £18.95 (paperback). Technovation, 16(2): 98, February 1996. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497296900097.

Perren:1997:BRSb

[Per97a] Lew Perren. Book review: Small firms: Contributions to economic regeneration: by Robert Blackburn and Peter Jennings. Paul Chapman Publishing Ltd, 144 Liverpool Road, London, N1 1LA, UK, 1996, 161 pp., ISBN 1-85396-339-9, £45.00 (hardback). Technovation, 17(9):535, September 1997. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297903345.

Perren:1997:BRSa

[Per97b] Lew Perren. Book review: The strategy process: European edition: by Henry Mintzberg, James Brian Quinn and Sumantra Ghoshal. Prentice Hall International (UK) Ltd, Campus 400, Maylands Avenue, Hemel Hempstead HP2 7EZ, UK, 985 pp., ISBN 0-13-149626-3, £24.95 (paperback). Technovation, 17(4):225, April 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297846913.

Pereira:2004:BRE

[Per04] Tiago Santos Pereira. Book review: The economics of science and technology: an overview of initiatives to foster innovation, entrepreneurship, and economics growth: Maryann P. Feldman, Albert N. Link, Donald Siegel; Kluwer Academic Publishers, Boston, 2002, pp. 135+xvi, ISBN 1-4020-7000-4. Technovation, 24(1):81–82, January 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203000981.

Petroni:1998:ADC

[Pet98] Alberto Petroni. The analysis of dynamic capabilities in a competence-oriented organization. Technovation, 18 (3):179–189, March 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016649729700093X.

Petruzzelli:2011:ITR

[Pet11] Antonio Messeni Petruzzelli. The impact of technological relatedness, prior ties, and geographical distance on university–industry collaborations: a joint-patent analysis. Technovation, 31(7):309–319, July 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211000228.

Pallister:1998:PPH

[PF98] John G. Pallister and Gordon R. Foxall. Psychometric properties of the Hurt-Joseph-Cook scales for the measurement of innovativeness. Technovation, 18(11):663–675, November 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497298000704.

Pfirrmann:1999:NSN

[Pfi99] Oliver Pfirrmann. Neither soft nor hard — pattern of development of new technology based firms in biotechnology. Technovation, 19(11):651–659, November 1999. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299000802.

Phaal:2006:TMT

[PFP06] Robert Phaal, Clare J. P. Farrukh, and David R. Probert. Technology management tools: concept, development and application. Technovation, 26(3):336–344, March 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000477.

Parvatiyar:1994:SPO

[PG94] Atul Parvatiyar and Yash P. Gupta. Synergistic pay-offs in Indo-US strategic alliances. Technovation, 14(6):395–406, August 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294900183. See [Ano94-37].

Pekovic:2009:QIE

[PG09] Sanja Pekovic and Fabrice Galia. From quality to innovation: Evidence from two French Employer Surveys. Technovation, 29(12):829–842, December 2009. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209001242.

Pries:2011:CIR

[PG11] Fred Pries and Paul Guild. Commercializing inventions resulting from university research: Analyzing the impact of technology characteristics on subsequent business models. Technovation, 31(4):151–160, April 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210000647.

Pundziene:2023:VCE

[PGB^{+}23] Asta Pundziene, Neringa Gerulaitiene, Sea Matilda Bez, Irène Georgescu, Christopher Mathieu, Jordi Carrabina-Bordoll, Josep Rialp-Criado, Hannu Nieminen, Alpo Varri, Susanne Boethius, Mark van Gils, Víctor Giménez-Garcia, Isabel Narbón-Perpiñá, Diego Prior-Jiménez, and Laura Vilutiene. Value capture and embeddedness in social-purpose-driven ecosystems. A multiple-case study of European digital healthcare platforms. Technovation, 124(??):??, June 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223000597.

Perez:2019:NAI

[PGKF19] Jorge A. Heredia Pérez, Cristian Geldes, Martin H. Kunc, and Alejandro Flores. New approach to the innovation process in emerging economies: the manufacturing sector case in Chile and Peru. Technovation, 79(??):35–55, January 2019. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497218301135.

Pan:2021:ETI

[PGLS21] Xiongfeng Pan, Shucen Guo, Mengna Li, and Jinbo Song. The effect of technology infrastructure investment on technological innovation — a study based on spatial Durbin model. Technovation, 107(??):??, September 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000961.

Prajogo:2008:ETP

[PH08] Daniel I. Prajogo and Soon W. Hong. The effect of TQM on performance in R&D environments: a perspective from South Korean firms. Technovation, 28(12):855–863, December 2008. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497208000746.

Phillimore:1999:BLV

[Phi99] John Phillimore. Beyond the linear view of innovation in science park evaluation: an analysis of Western Australian Technology Park. Technovation, 19(11):673–680, November 1999. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497299000620.

Piachaud:2002:OPM

[Pia02] B. S. Piachaud. Outsourcing in the pharmaceutical manufacturing process: an examination of the CRO experience. Technovation, 22(2):81–90, February 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000815.

Pilkington:2004:TPA

[Pil04] A. Pilkington. Technology portfolio alignment as an indicator of commercialisation: an investigation of fuel cell patenting. Technovation, 24(10):761–771, October 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720300004X.

Pistella:1989:MCM

[Pis89] Fabio Pistella. From Madrid to Copenhagen: the Milan Conference within EUREKA developments. Technovation, 9(2–3):205–208, June 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497289900588.

Pandey:1996:VCF

[PJ96] I. M. Pandey and Angela Jang. Venture capital for financing technology in Taiwan. Technovation, 16(9):499–514, 522–523, September 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497296000466.

Park:2006:KMS

[PK06] Yongtae Park and Seonwoo Kim. Knowledge management system for fourth generation R&D: KNOWVATION. Technovation, 26(5–6):595–602, May/June 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001877.

Parteka:2023:AIP

[PK23] Aleksandra Parteka and Aleksandra Kordalska. Artificial intelligence and productivity: global evidence from AI patent and bibliometric data. Technovation, 125(??):??, July 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223000755.

Paradkar:2015:ISU

[PKH15] Amit Paradkar, John Knight, and Paul Hansen. Innovation in start-ups: Ideas filling the void or ideas devoid of resources and capabilities? Technovation, 41–42(??):1–10, July/August 2015. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497215000280.

Petruzzelli:2021:E

[PL21] Antonio Messeni Petruzzelli and Antonio Lerro. Editorial. Technovation, ??(??):??, ???? 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497220300158.

Perez-Luno:2013:LMD

[PLC13] Ana Pérez-Luño and Jesús Cambra. Listen to the market: Do its complexity and signals make companies more innovative? Technovation, 33(6–7):180–192, June/July 2013. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213000230.

Paget:2024:MCI

[PLG24] Nicolas Paget, Pierre-Yves Le Gal, and Frédéric Goulet. Motivations and challenges of intrapreneurship in research organizations. The case of decision support systems in agricultural research for development. Technovation, 130(??): ??, February 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223002353.

Petruzzelli:2024:CRS

[PMN^{+}24] Antonio Messeni Petruzzelli, Luca Mora, Angelo Natalicchio, Federico Platania, and Celina Toscano Hernandez. Consumers’ reaction to sci-fi as a source of information for technological development: an empirical analysis. Technovation, 132(??):??, April 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000208.

Parolin:2024:HCT

[PMP24] Giácomo Parolin, Tim C. McAloone, and Daniela C. A. Pigosso. How can technology assessment tools support sustainable innovation? A systematic literature review and synthesis. Technovation, 129(??):??, January 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722300192X.

Patrucco:2023:RAC

[PMT23] Andrea S. Patrucco, Giacomo Marzi, and Daniel Trabucchi. The role of absorptive capacity and big data analytics in strategic purchasing and supply chain management decisions. Technovation, 126(??):??, August 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001256.

Porter:2011:MER

[PN11] Alan L. Porter and Nils C. Newman. Mining external R&D. Technovation, 31(4):171–176, April 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211000113.

Padula:2015:SIP

[PNC15] Giovanna Padula, Elena Novelli, and Raffaele Conti. SMEs inventive performance and profitability in the markets for technology. Technovation, 41–42(??):38–50, July/August 2015. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497215000036.

Perdomo-Ortiz:2006:TQM

[POGBG06] Jesús Perdomo-Ortiz, Javier González-Benito, and Jesús Galende. Total quality management as a forerunner of business innovation capability. Technovation, 26(10):1170–1185, October 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205001264.

Potter:1993:BRE

[Pot93] Stephen Potter. Book review: Exhausting our options: by Leonie J. Archer. Oxford Institute for energy studies, 1992. Report No. EV 14. ISBN 0-948061-72-3. 132 pp. £14, paperback. Technovation, 13(3):183–184, April 1993. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390043U.

Povey:1997:BRB

[Pov97] Barry Povey. Book review: Benchmarking: a tool for continuous improvement: by C. J. McNair and Kathleen H. J. Leibfried. John Wiley & Sons Ltd, Baffins Lane, Chichester PO19 1UD, UK, 344 pp., ISBN 0-939246-53-8, £14.99. Technovation, 17(4):225–226, April 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297846925.

Powell:2010:ECO

[Pow10] Benjamin C. Powell. Equity carve-outs as a technology commercialization strategy: an exploratory case study of Thermo Electron’s strategy. Technovation, 30(1):37–47, January 2010. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649720900087X.

Poynter:1997:BRE

[Poy97] Gavin Poynter. Book review: Expertise and innovation: Information strategies in the financial services Sector: R. Fincham, J. Fleck, R. Proctor, H. Scarbrough, M. Tierney and R. Williams, Oxford University Press, 1994, 334 pp., ISBN 0-19-828904-9. Technovation, 17(2):107, February 1997. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297841934.

Poznanski:1986:EPC

[Poz86] Kazimierz Z. Poznanski. The extinguishing process: a case study of steel technologies in the world industry. Technovation, 4(4):297–316, August 1986. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497286900325.

Prasad:2002:ESS

[PP02] Rose M. Prasad and Srinivas B. Prasad. Is the enterprise software sector still in transition? A research note. Technovation, 22(12):769–774, December 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000214.

Park:2004:NMT

[PP04] Yongtae Park and Gwangman Park. A new method for technology valuation in monetary value: procedure and application. Technovation, 24(5):387–394, May 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000998.

Pereira:2009:SLT

[PP09] Luciana Pereira and Guilherme Ary Plonski. Shedding light on technological development in Brazil. Technovation, 29 (6–7):451–464, June/July 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497208001521.

Pearson:1989:IMI

[PPB89] G. J. Pearson, A. W. Pearson, and D. F. Ball. Innovation in a mature industry: a case study of warp knitting in the U.K. Technovation, 9(8):657–679, December 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900059.

Papachristos:2024:PSA

[PPM24] George Papachristos, Eleni Papadonikolaki, and Bethan Morgan. Projects as a speciation and aggregation mechanism in transitions: Bridging project management and transitions research in the digitalization of UK architecture, engineering, and construction industry. Technovation, 132 (??):??, April 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497224000178.

Ponta:2024:RRA

[PPMC24] Linda Ponta, Gloria Puliga, Raffaella Manzini, and Silvano Cincotti. Reacting and recovering after an innovation failure. An agent-based approach. Technovation, 129(??):??, January 2024. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223001955.

Phaal:1998:TMMa

[PPP98a] R. Phaal, C. J. Paterson, and D. R. Probert. Technology management in manufacturing business: process and practical assessment. Technovation, 18(8– 9):541–553, 588–589, August/September 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000261.

Phaal:1998:TMMb

[PPP98b] R. Phaal, C. J. Paterson, and D. R. Probert. Technology management in manufacturing business: process and practical assessment. Technovation, 18(8–9):588–589, August/ September 1998. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497298901250.

Panebianco:2006:MST

[PPW06] Silke Panebianco and Claudia Pahl-Wostl. Modelling sociotechnical transformations in wastewater treatment — a methodological proposal. Technovation, 26(9):1090–1100, September 2006. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497205001495.

Panda:1996:TCA

[PR96] H. Panda and K. Ramanathan. Technological capability assessment of a firm in the electricity sector. Technovation, 16(10):561–588, October 1996. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297828969.

Panda:1997:TCA

[PR97] H. Panda and K. Ramanathan. Technological capability assessment as an input for strategic planning: Case studies at Electricité de France and Electricity Generating Authority of Thailand. Technovation, 17(7):359–390, July 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297000102.

Prasad:1986:TTA

[Pra86] S. Benjamin Prasad. Technology transfer: the approach of a Dutch multinational. Technovation, 4(1):3–15, January 1986. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497286900192.

Prabhu:1999:IUI

[Pra99] Ganesh N. Prabhu. Implementing university–industry joint product innovation projects. Technovation, 19(8):495–505, August 1999. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497298001254.

Prabhu:2010:BRI

[Pra10] Jaideep Prabhu. Book review: The Internationalization of Small Firms: a Strategic Entrepreneurship Perspective, Shameen Prashantham. Routledge, London, (2009). 163 pp. (including index), Hardback price: £65, ISBN 10-0-415-43177-8. Technovation, 30(7–8):482, July/August 2010. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209001412.

Park:2022:DSI

[PRA22] Sukyung Park, Eugenia Rosca, and Nivedita Agarwal. Driving social impact at the bottom of the Pyramid through the Internet-of-Things enabled frugal innovations. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497221001620.

Preece:2001:BRM

[Pre01] David Preece. Book review: Managing Technological Innovation: by John Ettlie; Wiley, Chichester and New York, 2000, pp. 432 (incl. Index), £22.50, ISBN 0-471-31546-X (paperback), supplements: Instructor’s Manual, Website. Technovation, 21(7):463, July 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000128.

Primus:2017:CCA

[Pri17] Dirk J. Primus. A configuration and contingency analysis of the development chain. Technovation, 64–65(??):1–15, June 2017. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497217303097.

Pitafi:2023:IVA

[PRID23] Abdul Hameed Pitafi, Muhammad Imran Rasheed, Nazrul Islam, and Amandeep Dhir. Investigating visibility affordance, knowledge transfer and employee agility performance. A study of enterprise social media. Technovation, 128(??):??, December 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497223001852.

Prokop:2021:UEE

[Pro21] Daniel Prokop. University entrepreneurial ecosystems and spinoff companies: Configurations, developments and outcomes. Technovation, 107(??):??, September 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000675.

Power:1997:ELRa

[PS97a] Damien J. Power and Amrik S. Sohal. An examination of the literature relating to issues affecting the human variable in just-in-time environments. Technovation, 17(11–12):649–666, November/December 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297000710.

Power:1997:ELRb

[PS97b] Damien J. Power and Amrik S. Sohal. An examination of the literature relating to issues affecting the human variable in just in time environments. Technovation, 17(11–12):727–728, November/December 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297904612.

Perez:2000:LPS

[PS00a] Manuela P. Pérez and Angel M. Sánchez. Lean production and supplier relations: a survey of practices in the Aragonese automotive industry. Technovation, 20(12):665– 676, December 2000. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497200000110.

Power:2000:HRM

[PS00b] Damien Power and Amrik S. Sohal. Human resource management strategies and practices in Just-In-Time environments: Australian case study evidence. Technovation, 20(7):373–387, July 2000. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497299001510.

Prajogo:2001:TIL

[PS01] Daniel I. Prajogo and Amrik S. Sohal. TQM and innovation: a literature review and research framework. Technovation, 21(9):539–558, September 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000705.

Perez:2003:DUS

[PS03] Manuela Pérez Pérez and Angel Martínez Sánchez. The development of university spin-offs: early dynamics of technology transfer and networking. Technovation, 23(10):823– 831, October 2003. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497202000342.

Prajogo:2004:MTP

[PS04] Daniel I. Prajogo and Amrik S. Sohal. The multidimensionality of TQM practices in determining quality and innovation performance — an empirical examination. Technovation, 24(6):443–453, June 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202001220.

Peng:2014:CCB

[PS14] Yi-Nung Peng and Susan Walsh Sanderson. Crossing the chasm with beacon products in the portable music player industry. Technovation, 34(2):77–92, February 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213001144.

Papazoglou:2021:IKF

[PS21] Michalis E. Papazoglou and Yiannis E. Spanos. “Influential knowledge and financial performance: the role of time and rivals’ absorptive capacity”. Technovation, 102(??):??, April 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221000043.

Perez:2002:BBT

[PSdLC02] M. P. Pérez, A. M. Sánchez, and M. P. de Luis Carnicer. Benefits and barriers of telework: perception differences of human resources managers according to company’s operations strategy. Technovation, 22(12):775–783, December 2002. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201000694.

Perez:2005:DFR

[PSdLCJ05] Manuela Pérez Pérez, Angel Martínez Sánchez, Pilar de Luis Carnicer, and María José Vela Jiménez. The differences of firm resources and the adoption of teleworking. Technovation, 25(12):1476–1483, December 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000726.

Pedersen:2022:EII

[PSG22] Torben Pedersen, Anna Scedrova, and Alina Grecu. The effects of IT investments and skilled labor on firms’ value added. Technovation, 116(??):??, August 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497222000268.

Pramongkit:2000:ATL

[PSS00] Prasopchoke Pramongkit, Teay Shawyun, and Boonmark Sirinaovakul. Analysis of technological learning for the Thai manufacturing industry. Technovation, 20(4):189–195, April 2000. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649729900125X.

Petzold:2023:HOD

[PSS23] Neele Petzold, Alexander Lennart Schmidt, and Laurent Scaringella. How to overcome the disruptor’s dilemma: Exploring strategic alliance reconfiguration of new market entrants. Technovation, 126(??):??, August 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001232.

Paci:1997:IPN

[PSU97] Raffaele Paci, Antonio Sassu, and Stefano Usai. International patenting and national technological specialization. Technovation, 17(1):25–38, January 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649729600065X.

Pilkington:2006:MTT

[PT06] Alan Pilkington and Thorsten Teichert. Management of technology: themes, concepts and relationships. Technovation, 26(3):288–299, March 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497205000465.

Puliga:2023:IDF

[PUFC23] Gloria Puliga, Andrea Urbinati, Enrico Maria Franchin, and Stefano Castegnaro. Investigating the drivers of failure of research-industry collaborations in open innovation contexts. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000906.

Pujari:2006:EIN

[Puj06] Devashish Pujari. Eco-innovation and new product development: understanding the influences on market performance. Technovation, 26(1):76–85, January 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001129.

Purser:1991:RKB

[Pur91] Ronald E. Purser. Redesigning the knowledge-based product development organization: a case study of sociotechnical systems change. Technovation, 11(7):403–416, November 1991. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729190021U.

Purser:1992:SSD

[Pur92] Ronald E. Purser. Sociotechnical systems design principles for computer-aided engineering. Technovation, 12(6): 379–386, September 1992. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729290003Z.

Petroni:2000:DTT

[PV00] Giorgio Petroni and Chiara Verbano. The development of a technology transfer strategy in the aerospace industry: the case of the Italian Space Agency. Technovation, 20(7): 345–351, July 2000. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497299001492.

Petroni:2001:EIR

[PV01] Giorgio Petroni and Chiara Verbano. The evolution of industrial research in Italy: characteristics and perspectives. Technovation, 21(9):585–594, September 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000808.

Pitt:2006:SBS

[PvdMB^{+}06] Leyland Pitt, Rian van der Merwe, Pierre Berthon, Esmail Salehi-Sangari, and Bradley R. Barnes. Swedish BioTech SMEs: the veiled values in online networks. Technovation, 26(5–6):553–560, May/June 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205001252.

Platt:1999:TDP

[PW99] Louise Platt and Gordon Wilson. Technology development and the poor/marginalised: context, intervention and participation. Technovation, 19(6–7):393–402, June 1999. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299000309.

Pandza:2011:CDN

[PWA11] Krsto Pandza, Terry A. Wilkins, and Eva A. Alfoldi. Collaborative diversity in a nanotechnology innovation system: Evidence from the EU Framework Programme. Technovation, 31(9):476–489, September 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211000721.

Pallister:2007:ASI

[PWF07] John G. Pallister, Hui-Chih Wang, and Gordon R. Foxall. An application of the style/involvement model to financial services. Technovation, 27(1–2):78–88, January/ February 2007. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497205001537.

Pao:2020:PIC

[PWL20] Huei-Wen Pao, Hsueh-Liang Wu, and Cheng-Yu Lee. Project interface choice and knowledge creation: Evidence of international science cooperation in Taiwan. Technovation, 89(??):??, January 2020. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497216301778.

Pyka:1997:IN

[Pyk97] Andreas Pyka. Informal networking. Technovation, 17 (4):207–220, 224, April 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497296001058.

Pyka:2000:INI

[Pyk00] Andreas Pyka. Informal networking and industrial life cycles. Technovation, 20(1):25–35, January 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299000838.

Pykalainen:2007:MPS

[Pyk07] Timo Pykäläinen. Model for profiting from software innovations in the new era in computing. Technovation, 27(4): 179–193, April 2007. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497206001118.

Quintana-Garcia:2004:CCI

[QGBV04] Cristina Quintana-García and Carlos A. Benavides-Velasco. Cooperation, competition, and innovative capability: a panel data of European dedicated biotechnology firms. Technovation, 24(12):927–938, December 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000609.

Querci:2024:WAB

[QMP24] Ilaria Querci, Luigi Monsurrò, and Paolo Peverini. When anthropomorphism backfires: Anticipation of negative social roles as a source of resistance to smart object adoption. Technovation, 132(??):??, April 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722400021X.

Queiroz:2006:BRS

[Que06] Sérgio Queiroz. Book review: Systems of Innovation and Development — Evidence from Brazil; José Eduardo Cassiolato, Helena Maria Martins Lastres, Maria Lucia Maciel (Eds.); Edward Elgar (New Horizons in the Economics of Innovation series, edited by Christopher Freeman); Cheltenham UK and Northampton, MA, USA; 2003; ISBN 1-84376-231-5. Technovation, 26(4):543, April 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000969.

Qin:2021:TTC

[QvdR21] Jian Qin and Bo van der Rhee. From trash to treasure: a checklist to identify high-potential NPD projects from previously rejected projects. Technovation, 104(??): ??, June 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497221000407.

Quintas:1992:AIL

[QWM92] Paul Quintas, David Wield, and Doreen Massey. Academicindustry links and innovation: questioning the science park model. Technovation, 12(3):161–175, April 1992. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729290033E.

Robertson:1986:KCP

[RA86] Gregg E. Robertson and David N. Allen. From kites to computers: Pennsylvania’s Ben Franklin partnership. Technovation, 4(1):29–43, January 1986. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497286900210.

Ritala:2017:DEI

[RA17] Paavo Ritala and Argyro Almpanopoulou. In defense of ‘eco’ in innovation ecosystem. Technovation, 60–61(??):39– 42, February 2017. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497216302681.

Rose-Anderssen:2005:IME

[RAATM05] C. Rose-Anderssen, P. M. Allen, C. Tsinopoulos, and I. Mc-Carthy. Innovation in manufacturing as an evolutionary complex system. Technovation, 25(10):1093–1105, October 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204000628.

Reekie:1984:TCT

[RAC84] W. Duncan Reekie, D. E. Allen, and J. N. Crook. On technological change, transfer and business characteristics: Some inferences from twelve case studies emanating from the national engineering laboratory. Technovation, 2(4): 233–254, July 1984. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497284900051.

Radosevic:1997:BRE

[Rad97] Slavo Radosevic. Book review: Educating entrepreneurs in modernising economies: edited by Peter Rosa, Michael G. Scott and Heinz Klandt, Avebury, Aldershot, 1996, £32.50 (hardback), ISBN 1-85972-184-2. Technovation, 17(10):600, October 1997. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649729789089X.

Radosevic:1999:ITT

[Rad99] Slavo Radosevic. International technology transfer policy: from “contract bargaining” to “sourcing”. Technovation, 19(6–7):433–444, June 1999. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497299000528.

Radosevic:2001:STI

[Rad01] Slavo Radosevic. Science, technology and industry outlook 2000: OECD, Paris, p. 258. Technovation, 21(11):757–758, November 2001. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649720100044X.

Raggi:1993:TGIa

[Rag93a] Andrea Raggi. Technological growth in the Italian economy: some indicators compared. Technovation, 13(1):3–15, January 1993. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729390010S.

Raggi:1993:TGIb

[Rag93b] Andrea Raggi. Technological growth in the Italian economy: some indicators compared. Technovation, 13(1):55, January 1993. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729390014M.

Ramos:2011:IST

[RAG11] Encarnación Ramos, Francisco J. Acedo, and M. Rosario Gonzalez. Internationalisation speed and technological patterns: a panel data study on Spanish SMEs. Technovation, 31(10–11):560–572, October/November 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211000915.

Ramo:1981:ATS

[Ram81] Simon Ramo. America’s technology slip — a new political issue. Technovation, 1(1):3–14, February 1981. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497281900031.

Ramanujam:1986:BRD

[Ram86] Vasudevan Ramanujam. Book review: Design and innovation: Edited by Richard Langdon and Roy Rothwell. Francis Pinter Publishers, London, 1986, xxi + 208 pages, Hardback, £18.50. Technovation, 4(4):317–318, August 1986. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497286900337.

Rammer:2023:MPI

[Ram23] Christian Rammer. Measuring process innovation output in firms: Cost reduction versus quality improvement. Technovation, 124(??):??, June 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000640.

Ranjan:2024:UML

[Ran24] Piyush Ranjan. Unraveling the mystery of the link between digital orientation and innovation performance: the interplay of digital business capability and environmental dynamism. Technovation, 131(??):??, March 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000166.

Rasmussen:2008:GIS

[Ras08] Einar Rasmussen. Government instruments to support the commercialization of university research: Lessons from Canada. Technovation, 28(8):506–517, August 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207001526.

Rubin:2015:KFT

[RAS15] Tzameret H. Rubin, Tor Helge Aas, and Andrew Stead. Knowledge flow in Technological Business Incubators: Evidence from Australia and Israel. Technovation, 41–42(??): 11–24, July/August 2015. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497215000139.

Radas:2015:EPS

[RATW15] Sonja Radas, Ivan-Damir Anić, Azra Tafro, and Vanja Wagner. The effects of public support schemes on small and medium enterprises. Technovation, 38(??):15–30, April 2015. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497214001229.

Radas:2009:ASI

[RB09] Sonja Radas and Ljiljana Bozić. The antecedents of SME innovativeness in an emerging transition economy. Technovation, 29(6–7):438–450, June/July 2009. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208001533.

Ruiz:2021:MIE

[RB21] Émilie Ruiz and Michela Beretta. Managing internal and external crowdsourcing: an investigation of emerging challenges in the context of a less experienced firm. Technovation, 106(??):??, August 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000717.

Reger:1998:SPE

[RBBB98] Guido Reger, Andreas Balthasar, Susanne Bührer, and Christoph Bättig. Switzerland’s participation in the European RTD framework programmes: a win-win game? Technovation, 18(6–7):425–438, January 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000327.

Rossetto:2023:MFI

[RBBF23] Dennys Eduardo Rossetto, Felipe Mendes Borini, Roberto Carlos Bernardes, and Gary L. Frankwick. Measuring frugal innovation capabilities: an initial scale proposition. Technovation, 121(??):??, March 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222002255.

Roth:2023:AGC

[RC23] Linus Roth and Simone Corsi. Ambidexterity in a geographic context: a systematic literature review on international exploration and exploitation of knowledge. Technovation, 124(??):??, June 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722300055X.

Ramasamy:2004:MLF

[RCC04] B. Ramasamy, A. Chakrabarty, and M. Cheah. Malaysia’s leap into the future: an evaluation of the multimedia super corridor. Technovation, 24(11):871–883, November 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720300049X.

Rothwell:1992:ETP

[RD92] Roy Rothwell and Mark Dodgson. European technology policy evolution: convergence towards SMEs and regional technology transfer. Technovation, 12(4):223–238, May 1992. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729290044I.

Rothschild:2005:TIS

[RD05] Leora Rothschild and Asaf Darr. Technological incubators and the social construction of innovation networks: an Israeli case study. Technovation, 25(1):59–67, January 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000646.

Raimo:2023:DDT

[RDAV23] Nicola Raimo, Ivano De Turi, Francesco Albergo, and Filippo Vitolla. The drivers of the digital transformation in the healthcare industry: an empirical analysis in Italian hospitals. Technovation, 121(??):??, March 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222001055.

Rodriguez:2017:VEK

[RDS17] Mercedes Rodriguez, David Doloreux, and Richard Shearmur. Variety in external knowledge sourcing and innovation novelty: Evidence from the KIBS sector in Spain. Technovation, 68(??):35–43, December 2017. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649721630284X.

Readman:2006:BRC

[Rea06] Jeff Readman. Book review: The Customer Centric Enterprise, Mitchell M. Tseng, Frank T. Piller (Eds.). Springer, London (2003). Technovation, 26(5–6):751–752, May/June 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205001781.

Reale:2021:MOI

[Rea21] Filippo Reale. Mission-oriented innovation policy and the challenge of urgency: Lessons from Covid-19 and beyond. Technovation, 107(??):??, September 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000870.

Reddy:1990:IST

[Red90] N. Mohan Reddy. International standardization of technical products. Technovation, 10(6):407–417, September 1990. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497290900044.

Reddy:2014:CAS

[Red14] Dan Reddy. Criticality analysis and the supply chain: Leveraging representational assurance. Technovation, 34 (7):362–379, July 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497214000108.

Reece:1999:SCC

[Ree99] J. David Reece. From simple crops to complex ecosystems: agricultural research and the environmental imperative. Technovation, 19(6–7):423–432, June 1999. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299000279.

Reitan:1997:FTE

[Rei97] Bjørnar Reitan. Fostering technical entrepreneurship in research communities: Granting scholarships to would-be entrepreneurs. Technovation, 17(6):287–296, June 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297001211.

Reppy:2008:BMI

[Rep08] Judith Reppy. A biomedical military-industrial complex? Technovation, 28(12):802–811, December 2008. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208001181.

Ranjbari:2024:DEF

[RES^{+}24] Meisam Ranjbari, Zahra Shams Esfandabadi, Peer-Olaf Siebers, Paola Pisano, and Francesco Quatraro. Digitally enabled food sharing platforms towards effective waste management in a circular economy: a system dynamics simulation model. Technovation, 130(??):??, February 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722300250X.

Revilla:2012:RBF

[RF12] Antonio J. Revilla and Zulima Fernández. The relation between firm size and R&D productivity in different technological regimes. Technovation, 32(11):609–623, November 2012. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497212000740.

Rouyre:2024:BPR

[RFB24] Audrey Rouyre, Anne-Sophie Fernandez, and Olga Bruyaka. Big problems require large collective actions: Managing multilateral coopetition in strategic innovation networks. Technovation, 132(??):??, April 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722400018X.

Roberts:2021:PFI

[RFMC21] Ruby Roberts, Rhona Flin, David Millar, and Luca Corradi. Psychological factors influencing technology adoption: a case study from the oil and gas industry. Technovation, 102 (??):??, April 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497220300912.

Rothwell:1985:IIR

[RG85] R. Rothwell and P. Gardiner. Invention, innovation, reinnovation and the role of the user: a case study of British hovercraft development. Technovation, 3(3):167–186, August 1985. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497285900124.

Rice:2006:APD

[RG06] John Rice and Peter Galvin. Alliance patterns during industry life cycle emergence: the case of Ericsson and Nokia. Technovation, 26(3):384–395, March 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720500060X.

Raven:2010:SCE

[RG10] R. P. J. M. Raven and F. W. Geels. Socio-cognitive evolution in niche development: Comparative analysis of biogas development in Denmark and The Netherlands (1973– 2004). Technovation, 30(2):87–99, February 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209001345.

Rauter:2023:ROC

[RGB23] Romana Rauter, Dietfried Globocnik, and Rupert J. Baumgartner. The role of organizational controls to advance sustainability innovation performance. Technovation, 128(??): ??, December 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223001669.

Rodriguez-Gulias:2021:ICC

[RGFLRP21] María Jesús Rodríguez-Gulías, Sara Fernández-López, and David Rodeiro-Pazos. Innovation in cultural and creative industries firms with an academic origin (CCI-USOs): the role of regional context. Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497217300901.

Ryan:2019:MFE

[RGH19] Paul Ryan, Will Geoghegan, and Rachel Hilliard. The microfoundations of firms’ explorative innovation capabilities within the triple helix framework. Technovation, 76–77(??): 15–27, August/September 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497218301287.

Resende:2013:BBT

[RGJ13] David N. Resende, David Gibson, and James Jarrett. BTP — Best Transfer Practices. A tool for qualitative analysis of tech-transfer offices: a cross cultural analysis. Technovation, 33(1):2–12, January 2013. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649721200096X.

Rizova:2018:OEP

[RGMW18] Polly S. Rizova, Samir Gupta, Elliot N. Maltz, and Robert W. Walker. Overcoming equivocality on projects in the fuzzy front end: Bringing social networks back in. Technovation, 78(??):40–55, December 2018. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0166497217301803.

Rajor:1994:ARP

[RGSK94] A. Rajor, R. P. Garg, R. K. Sharma, and R. M. Kothari. Anaerobic recycling of prehydrolysate generated during the manufacture of rayon grade pulp. Technovation, 14(8):549– 557, October 1994. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497294901538.

Ratinho:2010:RSP

[RH10] Tiago Ratinho and Elsa Henriques. The role of science parks and business incubators in converging countries: Evidence from Portugal. Technovation, 30(4):278–290, April 2010. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209001382.

Reinauer:2021:DAO

[RH21] Tobias Reinauer and Ulrich Elmer Hansen. Determinants of adoption in open-source hardware: a review of small wind turbines. Technovation, 106(??):??, August 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000705.

Romme:2023:TTC

[RH23] A. Georges L. Romme and Jan Holmström. From theories to tools: Calling for research on technological innovation informed by design science. Technovation, 121(??): ??, March 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223000032.

Ritala:2009:WIM

[RHL09] Paavo Ritala and Pia Hurmelinna-Laukkanen. What’s in it for me? Creating and appropriating value in innovationrelated coopetition. Technovation, 29(12):819–828, December 2009. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497209001047.

Ruan:2014:GRD

[RHW14] Yi Ruan, Chang Chieh Hang, and Yan Min Wang. Government’s role in disruptive innovation and industry emergence: the case of the electric bike in China. Technovation, 34(12):785–796, December 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214001254.

Ruokolainen:2004:FMF

[RI04] Jari Ruokolainen and Barbara Igel. The factors of making the first successful customer reference to leverage the business of start-up software company — multiple case study in Thai software industry. Technovation, 24(9):673–681, September 2004. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649720200130X.

Rickards:1985:MNT

[Ric85] Tudor Rickards. Making new things happen: an interpretation of observed innovation strategies. Technovation, 3(2): 119–131, June 1985. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497285900057.

Rickards:1995:BRA

[Ric95] Tudor Rickards. Book review: Adaptors and innovators: Styles of creativity and problem-solving: by Michael Kirton (editor). Routledge, London, 1994, 197 pp. ISBN 0-415-1162-7, clothback, £14.99. Technovation, 15(5):335– 336, June 1995. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497295900357.

Riedle:1989:DRA

[Rie89] Klaus Riedle. Demand for R&D activities and the trade off between in-house and external research: a viewpoint from industry with reference to large companies and small and medium-sized enterprises. Technovation, 9(2–3):213–225, June 1989. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497289900606.

Rakshit:2022:IBT

[RIMP22] Sandip Rakshit, Nazrul Islam, Sandeep Mondal, and Tripti Paul. Influence of blockchain technology in SME internationalization: Evidence from high-tech SMEs in India. Technovation, 115(??):??, July 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://

www.sciencedirect.com/science/article/pii/S0166497222000657.

Rycroft:2004:SOI

[RK04] Robert W. Rycroft and Don E. Kash. Self-organizing innovation networks: implications for globalization. Technovation, 24(3):187–197, March 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497203000920.

Ritala:2024:SMD

[RKFR24] Paavo Ritala, Joona Keränen, Jessica Fishburn, and Mika Ruokonen. Selling and monetizing data in B2B markets:

Four data-driven value propositions. Technovation, 130(??): ??, February 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223002468.

Roy:1997:DIS

[RkhR97] Robin Roy and Johann C. k. h. Riedel. Design and innovation in successful product competition. Technovation, 17(10):537–548, 593–594, October 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297000503.

Rocha-Lackiz:2006:BRI

[RL06] A. Rocha-Lackiz. Book review: Imitation of life: how biology is inspiring computing: Nancy Forbes. The MIT Press Cambridge, Massachusetts (2004). Technovation, 26(1):136, January 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205000258.

Radosevic:2009:BVR

[RM09] Slavo Radosevic and Marat Myrzakhmet. Between vision and reality: Promoting innovation through technoparks in an emerging economy. Technovation, 29(10):645–656, October 2009. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497209000467.

Ramani:2014:CBI

[RM14] Shyama V. Ramani and Vivekananda Mukherjee. Can breakthrough innovations serve the poor (bop) and create reputational (CSR) value? Indian case studies. Technovation, 34(5–6):295–305, May/June 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213000758.

Rasmussen:2006:IPC

[RMG06] Einar Rasmussen, Øystein Moen, and Magnus Gulbrandsen. Initiatives to promote commercialization of university knowledge. Technovation, 26(4):518–533, April 2006.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204002196.

Rejeb:2008:MIB

[RMGBA08] Helmi Ben Rejeb, Laure Morel-Guimarães, Vincent Boly, and N’Doli Guillaume Assiélou. Measuring innovation best practices: Improvement of an innovation index integrating threshold and synergy effects. Technovation, 28(12):838– 854, December 2008. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497208001090.

Redoli:2008:MAD

[RMGDLC08] Judith Redoli, Rafael Mompó, Javier García-Díez, and Miguel López-Coronado. A model for the assessment and development of Internet-based information and communication services in small and medium enterprises. Technovation, 28(7):424–435, July 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208000229.

Raja:2024:WTF

[RNFG24] Jawwad Z. Raja, Isabelle Fabienne Neufang, Thomas Frandsen, and Ismail Gölgeci. Working through frame incongruences: a process perspective on (re)framing for digital servitization. Technovation, 129(??):??, January 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722300202X.

Ruokolainen:2023:DHP

[RNJ^{+}23] Jari Ruokolainen, Satu Nätti, Milla Juutinen, Juha Puustinen, Anu Holm, Antti Vehkaoja, and Hannu Nieminen. Digital healthcare platform ecosystem design: a case study of an ecosystem for Parkinson’s disease patients. Technovation, 120(??):??, February 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222000980.

Roca:2022:RRM

[RO22] Jaime Bonnin Roca and Eoin O’Sullivan. The role of regulators in mitigating uncertainty within the Valley of Death. Technovation, 109(??):??, January 2022. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649721830614X.

Robertson:1984:CSI

[Rob84] Andrew Robertson. Characteristics of the successful inventor: Some notes on the nature of creativity and the creative mind. Technovation, 2(2):141–145, May 1984. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497284900178.

Roessner:1987:TPU

[Roe87] J. David Roessner. Technology policy in the United States: Structures and limitations. Technovation, 5(4):229–245, February 1987. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497287900654.

Rogers:1986:RRU

[Rog86] Everett M. Rogers. The role of the research university in the spin-off of high-technology companies. Technovation, 4(3): 169–181, June 1986. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497286900106.

Ritala:2015:KSK

[ROMH15] Paavo Ritala, Heidi Olander, Snejina Michailova, and Kenneth Husted. Knowledge sharing, knowledge leaking and relative innovation performance: an empirical study. Technovation, 35(??):22–31, January 2015. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214001060.

Rosegger:1982:DSD

[Ros82] Gerhard Rosegger. Diffusion and scale dynamics: a case study. Technovation, 1(4):291–304, August 1982. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497282900098.

Rosegger:1985:E

[Ros85] Gerhard Rosegger. Editorial. Technovation, 3(1):1, February 1985. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497285900318.

Rosegger:1987:E

[Ros87] Gerhard Rosegger. Editorial. Technovation, 6(1):1, April 1987. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497287900344.

Rosegger:1988:E

[Ros88] Gerhard Rosegger. Editorial. Technovation, 7(4):275–276, October 1988. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497288900016.

Rosegger:1989:BRF

[Ros89a] Gerhard Rosegger. Book review: Forschung und Entwicklung: Planung und Kontrolle: (Research and Development: Planning and Control), by Klaus Brockhoff, R. Oldenbourg Verlag, 1988, 289 pages. Technovation, 9(1):115–116, May 1989. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497289900461.

Rosegger:1989:BRO

[Ros89b] Gerhard Rosegger. Book review: Organizing business: Trade associations in America and Japan: by Leonard H. Lynn and Timothy J. McKeown, American Enterprise Institute for Public Policy Research, 1988. xviii + 191 pages, hardback, $19.75. Technovation, 9(5):455–456, August 1989. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497289900163.

Rosegger:1989:BRE

[Ros89c] Gerhard Rosegger. Book review: The entrepreneur: Mainstream views and radical critiques: 2nd edn., by Robert F. Hébert and Albert N. Link, Praeger, 1988, xviii + 178 pages, $37.95. Technovation, 9(4):373–375, July 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497287900083.

Rosegger:1990:BRT

[Ros90] Gerhard Rosegger. Book review: Technology and the pursuit of economic growth: By David C. Mowery and Nathan Rosenberg, Cambridge University Press, Cambridge and New York, 1989, Hardback $29.95, viii + 330 pages. Technovation, 10(6):440–441, September 1990. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497290900099.

Rosegger:1992:E

[Ros92] Gerhard Rosegger. Editorial. Technovation, 12(1):iii–iv, February 1992. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729290028G.

Rothwell:1987:BRS

[Rot87a] Roy Rothwell. Book review: Science parks and innovation centres: Their economic and social impact: edited by J. M. Gibb, Elsevier, 1985, 477 pages, Dfl.

175. Technovation, 6(4):303–304, September 1987. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497287900770.

Rothwell:1987:P

[Rot87b] Roy Rothwell. Preface. Technovation, 5(4):209–213, February 1987. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497287900630.

Rothwell:1988:BRS

[Rot88] Roy Rothwell. Book review: Scenarios and strategic management: By Michel Godet, Butterworths, Sevenoaks, Kent, UK, 210 pages, hardcover £27.50, ISBN 0-408-028940-4. Technovation, 7(4):421–423, October 1988. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900132.

Rothwell:1991:ENI

[Rot91] Roy Rothwell. External networking and innovation in small and medium-sized manufacturing firms in Europe. Technovation, 11(2):93–112, March 1991. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729190040B.

Rothwell:1992:IIGa

[Rot92a] Roy Rothwell. Industrial innovation and government environmental regulation: Some lessons from the past. Technovation, 12(7):447–458, November 1992. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729290050R.

Rothwell:1992:IIGb

[Rot92b] Roy Rothwell. Industrial innovation and government environmental regulation: Some lessons from the past. Technovation, 12(7):489–490, November 1992. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729290055M.

Rothwell:1993:CNI

[Rot93] Roy Rothwell. The changing nature of the innovation process. Technovation, 13(1):1–2, January 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390009K.

Roy:1994:EE

[Roy94] Robin Roy. The evolution of ecodesign. Technovation, 14(6):363–380, August 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497294900167.

See [Den94].

Rico-Pena:2023:MUC

[RPASLM23] Juan Jesús Rico-Peña, Raquel Arguedas-Sanz, and Carmen López-Martin. Models used to characterise blockchain features. A systematic literature review and bibliometric analysis. Technovation, 123(??):??, May 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000226.

Robra:2023:CDC

[RPGP23] Ben Robra, Alex Pazaitis, Chris Giotitsas, and Mario Pansera. From creative destruction to convivial innovation — a post-growth perspective. Technovation, 125(??): ??, July 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223000718.

Rhee:2010:DIP

[RPL10] Jaehoon Rhee, Taekyung Park, and Do Hyung Lee. Drivers of innovativeness and performance for innovative SMEs in South Korea: Mediation of learning orientation. Technovation, 30(1):65–75, January 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209000625.

Reid:1988:PPP

[RR88] Stan D. Reid and Laurel Reid. Public policy and promoting manufacturing under licensing. Technovation, 7(4):401–414, October 1988. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497288900090.

Ray:2011:PIP

[RR11a] Sangeeta Ray and Pradeep Kanta Ray. Product innovation for the people’s car in an emerging economy. Technovation, 31(5–6):216–227, May/June 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211000186.

Revilla:2011:TVP

[RR11b] Elena Revilla and Beatriz Rodríguez. Team vision in product development: How knowledge strategy matters. Technovation, 31(2–3):118–127, February/March 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649721000115X.

Robb:2022:EOI

[RRB^{+}22] Anna Robb, Marc Rohrschneider, Alex Booth, Peggy Carter, Richard Walker, and Georgina Andrews. Enhancing organisational innovation capability — a practiceoriented insight for pharmaceutical companies. Technovation, 115(??):??, July 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222000086.

Rodriguez-Repiso:2007:MIP

[RRSS07] Luis Rodriguez-Repiso, Rossitza Setchi, and Jose L. Salmeron. Modelling IT projects success: Emerging methodologies reviewed. Technovation, 27(10):582–594, October 2007. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497207000041.

Reece:2003:MCL

[RS03] J. D. Reece and J. Sumberg. More clients, less resources: toward a new conceptual framework for agricultural research in marginal areas. Technovation, 23(5):409–421, May 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201001225.

Rasmussen:2006:ABE

[RS06] Einar A. Rasmussen and Roger Sørheim. Action-based entrepreneurship education. Technovation, 26(2):185–194, February 2006. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497205000994.

Reilly:2014:SDI

[RS14] M. Reilly and P. Sharkey Scott. Subsidiary driven innovation within shifting MNC structures: Identifying new challenges and research directions. Technovation, 34(3):190– 202, March 2014. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497213001429.

Revilla:2005:TKM

[RSA05] Elena Revilla, Joseph Sarkis, and Juan Acosta. Towards a knowledge management and learning taxonomy for research joint ventures. Technovation, 25(11):1307–1316, November 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204000938.

Raymond:2010:RDI

[RSP10] Louis Raymond and Josée St-Pierre. R&D as a determinant of innovation in manufacturing SMEs: an attempt at empirical clarification. Technovation, 30(1):48–56, January 2010. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209000728.

Rogers:2001:LLA

[RTY01] Everett M. Rogers, Shiro Takegami, and Jing Yin. Lessons learned about technology transfer. Technovation, 21(4): 253–261, April 2001. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497200000390.

Ruberti:1989:F

[Rub89a] Antonio Ruberti. Foreword. Technovation, 9(2–3):127, June 1989. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497289900497.

Ruberti:1989:WR

[Rub89b] Antonio Ruberti. Welcoming remark. Technovation, 9(2–3): 199–202, June 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497289900564.

Ruefli:1986:IDP

[Rue86] Timothy W. Ruefli. Impacts of deregulation on performance and management of the largest American transportation companies. Technovation, 5(1–3):35–60, October 1986. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649728690043X.

Ruefli:1988:OPM

[Rue88] Timothy W. Ruefli. Ordinal performance measures for strategy and control: Application to the U.S. airline industry. Technovation, 8(1–3):43–70, ???? 1988. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900533.

Rantala:2023:SFV

[RUNS23] Tero Rantala, Juhani Ukko, Mina Nasiri, and Minna Saunila. Shifting focus of value creation through industrial digital twins — from internal application to ecosystemlevel utilization. Technovation, 125(??):??, July 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001062.

Ruokolainen:2005:GYS

[Ruo05] Jari Ruokolainen. Gear-up your software start-up company by the first reference customer-nomothetic research study in the Thai software industry. Technovation, 25(2):135– 144, February 2005. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203000737.

Rush:1994:BRF

[Rus94] Howard Rush. Book review: Forecasting for technologists and engineers: a practical guide for better decisions: by Brian C. Twiss. IEE Management of Technology Series 15. Available from Peter Peregrinus Ltd, Michael Faraday House, Six Hills Way, Stevenage, Herts Herts SGI AY, UK. 1992. Paperback (ISBN 0-86341-285-8)£20.00; casebound (ISBN 0-86341-265-3) £40.00. Technovation, 14(8):567–568, October 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294901597.

Rajalo:2017:UII

[RV17] Sigrid Rajalo and Maaja Vadi. University-industry innovation collaboration: Reconceptualization. Technovation, 62–63(??):42–54, April 2017. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497217302481.

Rajalo:2021:CPB

[RV21] Sigrid Rajalo and Maaja Vadi. Collaboration potential between low-capacity SMEs and academic researchers determined by symmetry of motivation. Technovation, 107(??): ??, September 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497221000857.

Romasanta:2022:AKE

[RvdSdE22] Angelo K. S. Romasanta, Peter van der Sijde, and Iwan J. P. de Esch. Absorbing knowledge from an emerging field: the role of interfacing by proponents in big pharma. Technovation, 110(??):??, February 2022. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001449.

Radnejad:2017:MOO

[RVW17] Amir Bahman Radnejad, Harrie Vredenburg, and Jaana Woiceshyn. Meta-organizing for open innovation under environmental and social pressures in the oil industry. Technovation, 66–67(??):14–27, August 2017. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497217300056.

Rothwell:1986:TCP

[RW86] Roy Rothwell and Hans Wissema. Technology, culture and public policy. Technovation, 4(2):91–115, April 1986. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497286900027.

Robson:2019:CNP

[RWP19] Karen Robson, Matthew Wilson, and Leyland Pitt. Creating new products from old ones: Consumer motivations for innovating autonomously from firms. Technovation, 88(??): Article 102075, December 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016649721730336X.

Rongping:2014:SCS

[RY14] Mu Rongping and Fan Yonggang. Security in the cyber supply chain: a Chinese perspective. Technovation, 34(7): 385–386, July 2014. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497214000224.

Ryan:1990:PEA

[Rya90] Neal Ryan. Policy evaluation and Australian support for innovation. Technovation, 10(4):265–272, June 1990. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729090057Q.

Ryan:1996:TSC

[Rya96] Neal Ryan. Technology strategy and corporate planning in Australian high-value-added manufacturing firms. Technovation, 16(4):195–201, 214, April 1996. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497295000518.

Ren:2021:TOD

[RZ21] Haiying Ren and Yuhui Zhao. Technology opportunity discovery based on constructing, evaluating, and searching knowledge networks. Technovation, 101(??):??, March 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497220300687.

Stanko:2022:DCM

[SA22] Michael A. Stanko and B. J. Allen. Disentangling the collective motivations for user innovation in a 3D printing community. Technovation, 111(??):??, March 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001681.

Saad:1998:BRM

[Saa98] M. Saad. Book review: Managing international technology transfer: by Min Chen, International Thomson Business Press, 1996, ISBN 0-41513-323-8. Technovation, 18(2):147, February 1998. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497297000928.

Saemundsson:2005:IBG

[Sae05] Rögnvaldur J. Saemundsson. On the interaction between the growth process and the development of technical knowledge in young and growing technology-based firms. Technovation, 25(3):223–235, March 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203001299.

Saetra:2024:SFS

[Sæt24] Henrik Skaug Sætra. Science fiction, sustainability, and scenario use: Comprehensive scenarios for improved strategy development and innovation. Technovation, 132(??): ??, April 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497224000269.

Salmeron:2003:ESK

[Sal03] Jose L. Salmeron. EIS success: keys and difficulties in major companies. Technovation, 23(1):35–38, January 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000761.

Shaik:2023:DLI

[SAM^{+}23] Aqueeb Sohail Shaik, Safiya Mukhtar Alshibani, Shreya Mishra, Armando Papa, and Maria Teresa Cuomo. Does learning from innovation failure enhance innovation performance? A quantitative investigation of small businesses. Technovation, 127(??):??, September 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001293.

Sanchez:1994:RFS

[Sán94] Angel Martínez Sánchez. R&D and firm size: some evidence from Spain. Technovation, 14(2):63–69, March 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900965. See [BCH94].

Sanchez:1995:ICF

[San95] Angel Martinez Sanchez. Innovation cycles and flexible automation in manufacturing industries. Technovation, 15 (6):351–362, August 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729596596L.

Sandulli:2007:CMP

[San07] Francesco D. Sandulli. CD music purchase behaviour of P2P users. Technovation, 27(6–7):325–334, June/July 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000053.

Sapienza:1989:TTA

[Sap89] Alice M. Sapienza. Technology transfer: an assessment of the major institutional vehicles for diffusion of U.S. biotechnology. Technovation, 9(6):463–478, October 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900278.

Sapienza:1992:IDS

[Sap92] Alice M. Sapienza. Irreconcilable differences? Strategic innovations for a publicly insured market. Technovation, 12(5):333–341, July 1992. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729290041F.

Sapsed:2001:BRK

[Sap01] Jonathan Sapsed. Book review: Knowledge and Strategy: Michael H. Zack (Ed.); Butterworth-Heinemann, 1999 Woburn, MA. Technovation, 21(10):705, October 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000414.

Sapsed:2005:BRI

[Sap05] J. Sapsed. Book review: The Industrial Dynamics of the New Digital Economy: Jens Frøslev Christensen, Peter Maskell (Eds.). Edward Elgar, Cheltenham (2003), ISBN 1-84376-376-1. Technovation, 25(11):1363–1364, November 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205000052.

Sapsed:2010:BRC

[Sap10] Jonathan Sapsed. Book review: Creativity, Innovation and the Cultural Economy: Andy C. Pratt, Paul Jeffcutt (Eds.). Routledge, London (2009). 280 pp., (hardback), $160 publishers, £76 Amazon, ISBN 978-0-415-41975-8. Technovation, 30(5–6):386–388, May/June 2010. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209001424.

Sarkis:1995:MSE

[Sar95] Joseph Sarkis. Manufacturing strategy and environmental consciousness. Technovation, 15(2):79–97, March 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729596612W.

Sarkissian:2008:IPR

[Sar08] Alfred Sarkissian. Intellectual property rights for developing countries: Lessons from Iran. Technovation, 28(11):786– 798, November 2008. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497208000515.

Sarbu:2022:III

[Sar22] Miruna Sarbu. The impact of industry 4.0 on innovation performance: Insights from German manufacturing and service firms. Technovation, 113(??):??, May 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497221001966.

Saiki:2006:NDP

[SAWT06] Tomoko Saiki, Yuji Akano, Chihiro Watanabe, and Yuji Tou. A new dimension of potential resources in innovation: a wider scope of patent claims can lead to new functionality development. Technovation, 26(7):796–806, July 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000908.

Sambasivan:2009:IPQ

[SAY09] Murali Sambasivan, Mohani Abdul, and Yuzliani Yusop. Impact of personal qualities and management skills of entrepreneurs on venture performance in Malaysia: Opportunity recognition skills as a mediating factor. Technovation, 29(11):798–805, November 2009. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209000479.

Steiner:1991:TLC

[SB91] Michael Steiner and Alex Belschan. Technology life cycles and regional types: an evolutionary interpretation and some stylized facts. Technovation, 11(8):483–498, December 1991. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729190008R.

Savioz:2002:SFT

[SB02] P. Savioz and M. Blum. Strategic forecast tool for SMEs: how the opportunity landscape interacts with business strategy to anticipate technological trends. Technovation, 22(2):91–100, February 2002. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497201000827.

Salimi:2015:DWI

[SBF15] Negin Salimi, Rudi Bekkers, and Koen Frenken. Does working with industry come at a price? A study of doctoral candidates’ performance in collaborative vs. non-collaborative Ph.D. projects. Technovation, 41–42(??):51–61, July/ August 2015. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497215000310.

Solomon:2013:SFT

[SBMP13] George T. Solomon, Andrew Bryant, Kevin May, and Vanessa Perry. Survival of the fittest: Technical assistance, survival and growth of small businesses and implications for public policy. Technovation, 33(8–9):292–301, August/ September 2013. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497213000643.

Sarpong:2023:TPR

[SBOB23] David Sarpong, Derrick Boakye, George Ofosu, and David Botchie. The three pointers of research and development (R&D) for growth-boosting sustainable innovation system. Technovation, 122(??):??, April 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222001286.

Santoro:2021:CMC

[SBP21] Gabriele Santoro, Stefano Bresciani, and Armando Papa. Collaborative modes with Cultural and Creative Industries and innovation performance: the moderating role of heterogeneous sources of knowledge and absorptive capacity. Technovation, ??(??):??, ???? 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649721730398X.

Solleiro:2005:CIS

[SC05] José Luis Solleiro and Rosario Castañón. Competitiveness and innovation systems: the challenges for Mexico’s insertion in the global context. Technovation, 25(9):1059–1070, September 2005. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649720400032X.

Scillitoe:2010:RII

[SC10] Joanne L. Scillitoe and Alok K. Chakrabarti. The role of incubator interactions in assisting new ventures. Technovation, 30(3):155–167, March 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209001734.

Saemundsson:2017:ACI

[SC17] Rognvaldur J. Saemundsson and Marina Candi. Absorptive capacity and the identification of opportunities in new technology-based firms. Technovation, 64–65(??):43–49, June 2017. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497216304096.

Shankar:2020:SQF

[SC20] Raj K. Shankar and Tommy H. Clausen. Scale quickly or fail fast: an inductive study of acceleration. Technovation, 98(??):??, December 2020. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497220300468.

Siepel:2017:NFH

[SCC17] Josh Siepel, Marc Cowling, and Alex Coad. Non-founder human capital and the long-run growth and survival of hightech ventures. Technovation, 59(??):34–43, January 2017. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497216303121.

Seip:2018:TTA

[SCFD18] Marcel Seip, Carolina Castaldi, Meindert Flikkema, and Ard-Pieter De Man. The timing of trademark application in innovation processes. Technovation, 72–73(??): 34–45, April/May 2018. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497218301172.

Schwartz:1985:IBD

[Sch85] Hugh Schwartz. Industrialization bottlenecks and the development of industrial technology. Technovation, 3(4):297– 313, October 1985. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497285900264.

Scheutz:1986:CES

[Sch86] Curt Scheutz. Critical events for Swedish entrepreneurs in entrepreneurial spin-offs. Technovation, 5(1–3):169–182, October 1986. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497286900519.

Schulz:1992:ERT

[Sch92] William C. Schulz. The emergence of the real-time computer reservation system as a competitive weapon in the U.S. airline industry 1958–1989. A paper on strategic innovation. Technovation, 12(2):65–74, March 1992. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729290016B.

Seaton:1993:DAIa

[SCH93a] R. A. F. Seaton and M. Cordey-Hayes. The development and application of interactive models of industrial technology transfer. Technovation, 13(1):45–53, January 1993. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390013L.

Seaton:1993:DAIb

[SCH93b] R. A. F. Seaton and M. Cordey-Hayes. The development and application of interactive models of industrial technology transfer. Technovation, 13(1):58–59, January 1993. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390017P.

Schulz:1994:BSB

[Sch94] William C. Schulz III. Building the skill-based corporation — the case of Combustion Engineering, 1982– 1990. Technovation, 14(7):461–476, September 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900043.

Schmidt:2009:NPR

[Sch09] Richard S. Schmidt. NASA pressure-relieving foam technology is keeping the leading innerspring mattress firms awake at night. Technovation, 29(3):181–191, March 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208000874.

Son:2019:DTE

[SCH19] Hosung Son, Yanghon Chung, and Heeju Hwang. Do technology entrepreneurship and external relationships always promote technology transfer? Evidence from Korean public research organizations. Technovation, 82–83 (??):1–15, April/May 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497216300724.

Shen:2023:PTT

[SCH23] Huijun Shen, Wim Coreynen, and Can Huang. Prestige and technology-transaction prices: Evidence from patentselling by Chinese universities. Technovation, 123(??):??, May 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223000214.

Sciberras:1983:NCT

[Sci83] E. Sciberras. New competition and technical change in the computer industry. Technovation, 2(1):17–26, February 1983. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497283900044.

Spithoven:2010:BAC

[SCK10] André Spithoven, Bart Clarysse, and Mirjam Knockaert. Building absorptive capacity to organise inbound open innovation in traditional industries. Technovation, 30(2):130– 141, February 2010. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497209001333.

Spithoven:2011:BAC

[SCK11] André Spithoven, Bart Clarysse, and Mirjam Knockaert. Building absorptive capacity to organise inbound open innovation in traditional industries. Technovation, 31(1):10– 21, January 2011. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497210001112.

Subramanian:2011:CRI

[SCM11] Annapoornima M. Subramanian, Kah-Hin Chai, and Shifeng Mu. Capability reconfiguration of incumbent firms: Nintendo in the video game industry. Technovation, 31 (5–6):228–239, May/June 2011. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497211000137.

Singh:2024:AGA

[SCM24] Kuldeep Singh, Sheshadri Chatterjee, and Marcello Mariani. Applications of generative AI and future organizational performance: the mediating role of explorative and exploitative innovation and the moderating role of ethical dilemmas and environmental dynamism. Technovation, 133 (??):??, May 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497224000713.

Scott:1989:BRP

[Sco89] Ron Scott. Book review: Product Innovation: Idea to Exploitation: by J. A. A. Bradbury, Wiley, Chichester, 1989, 181 pages, £18.95, ISBN 0-471-92169-

6. Technovation, 9(5):456–459, August 1989. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900175.

Su:2006:LIP

[SCS06] Chao-Ton Su, Yung-Hsin Chen, and D. Y. Sha. Linking innovative product development with customer knowledge: a data-mining approach. Technovation, 26(7):784– 795, July 2006. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497205000799.

Saemundsson:2022:IPF

[SCS22] Rögnvaldur Saemundsson, Marina Candi, and Throstur Olaf Sigurjonsson. The influence of performance feedback and top management team orientation on decisions about R&D in technology-based firms. Technovation, 113(??):??, May 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497221002017.

Seo:2017:RCU

[SCY17] Hangyeol Seo, Yanghon Chung, and Hyungseok (David) Yoon. R&D cooperation and unintended innovation performance: Role of appropriability regimes and sectoral characteristics. Technovation, 66–67(??):28–42, August 2017. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497217301906.

Son:2022:HCU

[SCY22] Hosung Son, Yanghon Chung, and Sangpil Yoon. How can university technology holding companies bridge the Valley of Death? Evidence from Korea. Technovation, 109(??): ??, January 2022. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497218305716.

Simoni:2014:PDS

[SCZ14] Michele Simoni, Cabirio Cautela, and Francesco Zurlo. Product design strategies in technological shifts: an explorative study of Italian design-driven companies. Technovation, 34(11):702–715, November 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214000844.

Swinnerton-Dyer:1989:EPC

[SD89a] Peter Swinnerton-Dyer. Existing and potential capabilities for R&D activities in industry-related projects: a viewpoint from universities and other research organizations. Technovation, 9(2–3):227–233, June 1989. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900618.

Swinnerton-Dyer:1989:PSD

[SD89b] Peter Swinnerton-Dyer. Peter Swinnerton-Dyer. Technovation, 9(2–3):295–296, June 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497289900734.

Sweeting:1995:IIP

[SD95] R. C. Sweeting and R. E. Davies. Industrial innovation and parallel accounting developments. Technovation, 15(5): 289–302, June 1995. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729596601O.

Swierczek:2004:LIP

[SD04] F. W. Swierczek and G. P. Dhakal. Learning and its impact on the performance of manufacturing joint ventures in developing countries. Technovation, 24(1):53–62, January 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497202000494.

Sun:2010:DII

[SD10] Yifei Sun and Debin Du. Determinants of industrial innovation in China: Evidence from its recent economic census. Technovation, 30(9–10):540–550, September/October 2010. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210000702.

Shaygan:2023:TMM

[SD23] Amir Shaygan and Tugrul Daim. Technology management maturity assessment model in healthcare research centers. Technovation, 120(??):??, February 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722100225X.

Sohal:2001:RRP

[SDFN01] Amrik S. Sohal, Brian D’Netto, Paul Fitzpatrick, and Hamid Noori. The roles and responsibilities of production/operations managers in SMEs: evidence from Canada. Technovation, 21(7):437–448, July 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000602.

Stojcic:2024:CCC

[SDK24] Nebojsa Stojcić, Marina Dabić, and Sascha Kraus. Customisation and co-creation revisited: Do user types and engagement strategies matter for product innovation success? Technovation, 134(??):??, June 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000956.

Suominen:2023:UVT

[SDN23] Arho Suominen, Matthias Deschryvere, and Rumy Narayan. Uncovering value through exploration of barriers — a perspective on intellectual property rights in a national innovation system. Technovation, 123(??):??, May 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000305.

Salerno:2015:IPW

[SdO^{+}15] Mario Sergio Salerno, Leonardo Augusto de Vasconcelos Gomes, Débora Oliveira da Silva, Raoni Barros Bagno, and Simone Lara Teixeira Uchôa Freitas. Innovation processes: Which process for which project? Technovation, 35(??):59– 70, January 2015. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497214001072.

Shankar:2022:BFW

[SDT^{+}22] Amit Shankar, Amandeep Dhir, Shalini Talwar, Nazrul Islam, and Piyush Sharma. Balancing food waste and sustainability goals in online food delivery: Towards a comprehensive conceptual framework. Technovation, 117 (??):??, September 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222001535.

Silva:2024:IPE

[SdVGdFB24] Lucas Emmanuel Nascimento Silva, Leonardo Augusto de Vasconcelos Gomes, Aline Mariane de Faria, and Felipe Mendes Borini. Innovation processes in ecosystem settings: an integrative framework and future directions. Technovation, 132(??):??, April 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000348.

Silvestre:2014:CAI

[Se14] Bruno S. Silvestre and Romeu e Silva Neto. Capability accumulation, innovation, and technology diffusion: Lessons from a Base of the Pyramid cluster. Technovation, 34(5–6):270–283, May/June 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213001120.

Segal:1986:UTE

[Seg86] N. S. Segal. Universities and technological entrepreneurship in Britain: Some implications of the Cambridge phenomenon. Technovation, 4(3):189–204, June 1986. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649728690012X.

Senker:1985:SHT

[Sen85] Jacqueline Senker. Small high technology firms: some regional implications. Technovation, 3(4):243–262, October 1985. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497285900239.

Senker:1996:NSI

[Sen96] Jacqueline Senker. National systems of innovation, organizational learning and industrial biotechnology. Technovation, 16(5):219–229, 265, May 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0166497295000690.

Seror:1996:ARI

[Sér96] Ann C. Séror. Action research for international information technology transfer: a methodology and a network model. Technovation, 16(8):421–429, 447–448, August 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497296000326.

Serageldin:2013:HDW

[Ser13] Ismail Serageldin. How do we move forward to maintain sustainability for the future of humankind? Technovation, 33(4–5):105, April/May 2013. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016649721300028X.

Szmigin:1998:TFI

[SF98] Isabelle Szmigin and Gordon Foxall. Three forms of innovation resistance: the case of retail payment methods. Technovation, 18(6–7):459–468, January 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000303.

Shefer:2005:RFS

[SF05] Daniel Shefer and Amnon Frenkel. R&D, firm size, and innovation: an empirical analysis. Technovation, 25(1):25– 32, January 2005. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203001524.

Siqueira:2023:EER

[SFBK23] Erica H. S. Siqueira, Bruno Brandão Fischer, Adriana Bin, and Jill Kickul. Entrepreneurial ecosystems’ readiness towards knowledge-intensive sustainable entrepreneurship: Evidence from Brazil. Technovation, 126(??):??, August 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223001311.

Shi:2014:DMG

[SFC14] Xiaohui Shi, Kiran Fernandes, and Pattarin Chumnumpan. Diffusion of multi-generational high-technology products. Technovation, 34(3):162–176, March 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213001466.

Santos:2023:IEE

[SFdO^{+}23] Juliana Bonomi Santos, Aline Rodrigues Fernandes, Pedro Tonhozi de Oliveira, Laís Moltene Maia, and Raul Beal Partyka. Increasing entrepreneurial ecosystemlevel outcomes through orchestration: a proposed framework. Technovation, 128(??):??, December 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001840.

Simms:2021:FER

[SFF21] Christopher Simms, Johan Frishammar, and Nicholas Ford. The front end in radical process innovation projects: Sources of knowledge problems and coping mechanisms. Technovation, ??(??):??, ???? 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497220300869.

Suhada:2021:MIC

[SFVI21] Thontowi A. Suhada, Jerad A. Ford, Martie-Louise Verreynne, and Marta Indulska. Motivating individuals to contribute to firms’ non-pecuniary open innovation goals. Technovation, 102(??):??, April 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000146.

Samsom:1993:USE

[SG93] Karel J. Samsom and Michael A. Gurdon. University scientists as entrepreneurs: a special case of technology transfer and high-tech venturing. Technovation, 13(2):63–71, March 1993. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729390054Y. See [Ano93s].

Shah:2022:SES

[SG22] Muhammad Umair Shah and Paul D. Guild. Stakeholder engagement strategy of technology firms: a review and applied view of stakeholder theory. Technovation, 114(??):??, June 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000074.

Sundriyal:2024:ECF

[SG24] Vivek Kumar Sundriyal and Jonas Gabrielsson. The employment consequences of founding an incorporated business among STEM founders: Evidence from Swedish microdata. Technovation, 133(??):??, May 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000695.

Samara:2012:IIP

[SGB12] Elpida Samara, Patroklos Georgiadis, and Ioannis Bakouros. The impact of innovation policies on the performance of national innovation systems: a system dynamics analysis. Technovation, 32(11):624–638, November 2012. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497212000727.

Stock:2002:FSD

[SGF02] Gregory N. Stock, Noel P. Greis, and William A. Fischer. Firm size and dynamic technological innovation. Technovation, 22(9):537–549, September 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720100061X.

Sohal:1999:MPC

[SGFS99] Amrik S. Sohal, John Gordon, Greg Fuller, and Alan Simon. Manufacturing practices and competitive capability: an Australian study. Technovation, 19(5):295–304, February 1999. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497298001217.

Sgobbi:1995:IAD

[Sgo95] Francesca Sgobbi. An integrated approach to the diffusion of technological innovation: the non-manufacturing advanced robotics field. Technovation, 15(1):11–24, February 1995. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729596608V.

Situmeang:2016:RTB

[SGWL16] Frederik B. I. Situmeang, Gerda Gemser, Nachoem M. Wijnberg, and Mark A. A. M. Leenders. Risk-taking behavior of technology firms: the role of performance feedback in the video game industry. Technovation, 54(??):22–34, August 2016. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497216300165.

Stepan:1995:OLC

[SH95] Adolf Stepan and Christian Hillinger. The optimal lifetime of capital goods and pricing policies for replacement components and repair costs. Technovation, 15(2):111–120, March 1995. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729596614Y.

Sung:1999:DPN

[SH99] Chang Sup Sung and Sa Kyun Hong. Development process of nuclear power industry in a developing country: Korean experience and implications. Technovation, 19(5):305–316, February 1999. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497298001242.

Sun:2002:ABM

[SH02] Hongyi Sun and Cui Hong. The alignment between manufacturing and business strategies: its influence on business performance. Technovation, 22(11):699–705, November 2002. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201000669.

Salmeron:2006:MRE

[SH06] Jose L. Salmeron and José M. Hurtado. Modelling the reasons to establish B2C in the fashion industry. Technovation, 26(7):865–872, July 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000945.

Schwartz:2008:SSB

[SH08] Michael Schwartz and Christoph Hornych. Specialization as strategy for business incubators: an assessment of the Central German Multimedia Center. Technovation, 28(7): 436–449, July 2008. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497208000187.

Schwartz:2010:CPI

[SH10a] Michael Schwartz and Christoph Hornych. Cooperation patterns of incubator firms and the impact of incubator specialization: Empirical evidence from Germany. Technovation, 30(9–10):485–495, September/October 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210000635.

Suh:2010:AES

[SH10b] Dukrok Suh and Junseok Hwang. An analysis of the effect of software intellectual property rights on the performance of software firms in South Korea. Technovation, 30(5–6):376– 385, May/June 2010. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497209001321.

Shafique:2022:LUT

[SH22] Muhammad Shafique and John Hagedoorn. Look at U: Technological scope of the acquirer, technological complementarity with the target, and post-acquisition R&D output. Technovation, 115(??):??, July 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497222000803.

Shi:2023:REI

[SH23] Yuwei Shi and John Herniman. The role of expectation in innovation evolution: Exploring hype cycles. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222000062.

Shaw:1986:ATI

[Sha86] Brian Shaw. Appropriation and transfer of innovation benefit in the U.K. medical equipment industry. Technovation, 4(1):45–65, January 1986. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497286900222.

Sharp:1987:NPT

[Sha87] Margaret Sharp. National policies towards biotechnology. Technovation, 5(4):281–304, February 1987. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649728790068X.

Shaw:1991:BRS

[Sha91a] Brian Shaw. Book review: Successful product development: A. Johne and P. Snelson, Basil Blackwell, Oxford, 1990, £29.95, 240 pages. Technovation, 11(8):499–500, December 1991. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729190009S.

Shaw:1991:BRU

[Sha91b] Brian Shaw. Book review: Understanding business markets: Edited by D. Ford, Academic press, London, UK, 1990, Softback £16.95, Hardback £35.00, 556 pages. Technovation, 11(8):500–501, December 1991. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497291900102.

Shaw:1993:FINa

[Sha93a] Brian Shaw. Formal and informal networks in the UK medical equipment industry. Technovation, 13(6):349–365, September 1993. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497293900768.

Shaw:1993:FINb

[Sha93b] Brian Shaw. Formal and informal networks in the UK medical equipment industry. Technovation, 13(6):400–403, September 1993. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729390080F.

Shapiro:1995:BRB

[Sha95] Gillian Shapiro. Book review: Bringing technology home: Gender and technology in a changing Europe: edited by Cynthia Cockburn and Ruza Fürst Dilić. Open University Press, Celtic Court, 22 Ballmoor, Buckingham MK18 1XW, UK, 1994, 187 pp., ISBN 0-335-19158-4. £12.99 (paperback), £37.50 (hardback). Technovation, 15(2):129–130, March 1995. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497295900195.

Sharma:2003:RSA

[Sha03] B. Sharma. R&D strategy and Australian manufacturing industry: an empirical investigation of emphasis and effectivness. Technovation, 23(12):929–937, December 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000354.

Sheng:2005:WM

[She05] Margaret L. Sheng. The wired mother. Technovation, 25 (9):1071–1077, September 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497204000367.

Sherif:2009:ISI

[She09] Mostafa Hashem Sherif. Introduction to the special issue on technology management in the service economy. Technovation, 29(5):325–326, May 2009. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209000388.

Shi:2024:ACE

[Shi24] Jincheng Shi. Adaptive change: Emerging economy enterprises respond to the international business environment challenge. Technovation, 133(??):??, May 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000488.

Si:2023:TEI

[SHS^{+}23] Steven Si, Jeremy Hall, Roy Suddaby, David Ahlstrom, and Jiang Wei. Technology, entrepreneurship, innovation and social change in digital economics. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222000311.

Sime:2023:IPL

[SHT23] Serge Sime, John Hagedoorn, and Hui Tan. Innovation performance and licensing: the effect of the compositional quality of direct and indirect network ties. Technovation, 127(??):??, September 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001372.

Shu:2022:PFC

[Shu22] Ei (Emily) Shu. Paradoxical framing and coping process on sustainable new product development. Technovation, 111 (??):??, March 2022. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497221001735.

Salami:1998:LDM

[SI98] Ayobami T. Salami and M. O. Ilori. Locational decisionmaking and pattern of diffusion of frozen fish depots in Ibadan city, Nigeria. Technovation, 18(3):207–213, March 1998. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297000886.

Sikka:1991:STD

[Sik91] Pawan Sikka. Strategies for technology development in India. Technovation, 11(7):445–452, November 1991. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729190025Y.

Sikka:1994:GTB

[Sik94] Pawan Sikka. GATT and the technology base in India. Technovation, 14(9):613–620, November 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900434. See [Paw94].

Sikka:1996:IDA

[Sik96] Pawan Sikka. Indigenous development and acquisition of technology: an Indian perspective. Technovation, 16(2):85–90, 94–95, February 1996. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497295000224.

Sikka:1997:TSFa

[Sik97a] Pawan Sikka. Technology support and financing system for development and commercialisation — Perspectives in India. Technovation, 17(11–12):707–714, November/ December 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497297000527.

Sikka:1997:TSFb

[Sik97b] Pawan Sikka. Technology support and financing system for development and commercialisation — perspectives in India. Technovation, 17(11–12):731, November/December 1997. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649729790465X.

Sikka:1999:TIS

[Sik99] Pawan Sikka. Technological innovations by SME’s in India. Technovation, 19(5):317–321, February 1999. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298001266.

Sanni:2001:NTP

[SIOO01] S. A. Sanni, M. O. Ilori, A. O. Opaleye, and A. A. Oyewale. Nigeria’s technology policy: is it adequate in the globalizing world? Technovation, 21(4):237–243, April 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000444.

Sirilli:1991:TBP

[Sir91] Giorgio Sirilli. The technological balance of payments as an indicator of technology transfer in OECD countries.

The case of Italy. Technovation, 11(1):3–25, January 1991. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729190003M.

Soetanto:2016:IUB

[SJ16] Danny Soetanto and Sarah Jack. The impact of universitybased incubation support on the innovation strategy of academic spin-offs. Technovation, 50–51(??):25–40, April/ May 2016. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497215000759.

Slavova:2021:UAF

[SJ21] Kremena Slavova and Simcha Jong. University alliances and firm exploratory innovation: Evidence from therapeutic product development. Technovation, 107(??):??, September 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221000912.

Sun:2024:CBS

[SJCT24] Yutao Sun, Lin Jiang, Cong Cao, and Fang-Mei Tseng. From contributors to boundary spanners: Evolving roles of government agencies in China’s innovation policy network (1980–2019). Technovation, 132(??):??, April 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000245.

Sharma:2023:DWN

[SJP23] Rohit Sharma, Geetika Jain, and Justin Paul. Does the world need to change its vaccine distribution strategy for COVID-19? Technovation, 126(??):??, August 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649722300130X. See corrigen-

dum [SJSP23].

Sharma:2023:CDW

[SJSP23] Rohit Sharma, Geetika Jain, Archana Shrivastava, and Justin Paul. Corrigendum to “Does the World need to change its vaccine distribution strategy for COVID-19?”

[Technovation 126 (2023) 102819]. Technovation, 127(??): ??, September 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223001451.  See

[SJP23].

Sharma:1992:IAC

[SK92] Rajesh K. Sharma and R. M. Kothari. Innovative application of corn steep liquor for the increased production of food grains. Technovation, 12(4):213–221, May 1992. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729290043H.

Standing:2011:HCO

[SK11] Craig Standing and Sarah Kiniti. How can organizations use wikis for innovation? Technovation, 31(7):287–295, July 2011. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649721100040X.

Samant:2023:BFF

[SK23] Shantala Samant and Jongwook Kim. Best foot forward? The importance of contractual governance mechanisms for innovation from alliances. Technovation, 127(??): ??, September 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223001396.

Skipton:1984:IIM

[Ski84] M. D. Skipton. Industrial innovation — Matching technological offering to customer need. Technovation, 2(4): 277–291, July 1984. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497284900075.

Song:2017:DNT

[SKL17] Kisik Song, Karp Soo Kim, and Sungjoo Lee. Discovering new technology opportunities based on patents: Textmining and F-term analysis. Technovation, 60–61(??):1–14, February 2017. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497217301682.

Skowronski:1987:TTI

[Sko87] Stanisław Skowroński. Transfer of technology and industrial cooperation. Technovation, 7(1):17–22, December 1987. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497287900447.

Shibata:2008:DER

[SKTM08] Naoki Shibata, Yuya Kajikawa, Yoshiyuki Takeda, and Katsumori Matsushima. Detecting emerging research fronts based on topological measures in citation networks of scientific publications. Technovation, 28(11):758–775, November 2008. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497208000436.

Stoop:1998:PDRa

[SL98a] Mat. L. M. Stoop and Alfred J. D. Lambert. Processing of discarded refrigerators in The Netherlands. Technovation, 18(2):101–110, February 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497297000709.

Stoop:1998:PDRb

[SL98b] Mat L. M. Stoop and Alfred J. D. Lambert. Processing of discarded refrigerators in the Netherlands: some historical, energitical and environmental aspects. Technovation, 18(2):143–144, February 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497297000916.

Shluzas:2014:IVP

[SL14a] Lauren M. Aquino Shluzas and Larry J. Leifer. The insight-value-perception (iVP) model for user-centered design. Technovation, 34(11):649–662, November 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497212000946.

Sun:2014:SPS

[SL14b] Hongyi Sun and Jonathan D. Linton. Structuring papers for success: Making your paper more like a high impact publication than a desk reject. Technovation, 34(10):571– 573, October 2014. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497214001035.

Sun:2014:VDC

[SL14c] Hongyi Sun and Jonathan D. Linton. Value of design for competitiveness — Introduction to the special issue. Technovation, 34(11):647–648, November 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649721400128X.

Solomon:2016:SCK

[SL16] George T. Solomon and Jonathan D. Linton. Standing at the crossroad of knowledge: Technology, innovation, entrepreneurship and the small business management — Policy perspectives. Technovation, 57–58(??):1–3, November/ December 2016. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497216302425.

SantosDalenogare:2023:BDS

[SLA^{+}23a] Lucas Santos Dalenogare, Marie-Anne Le Dain, Néstor F. Ayala, Giuditta Pezzotta, and Alejandro G. Frank. Building digital servitization ecosystems: an analysis of interfirm collaboration types and social exchange mechanisms among actors. Technovation, 124(??):??, June 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000676.

Slavova:2023:LUT

[Sla23b] Kremena Slavova. In-licensing university technology and firm innovation. Technovation, 125(??):??, July 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722300069X.

Sohn:2013:VAP

[SLJ13] So Young Sohn, Won Sang Lee, and Yong Han Ju. Valuing academic patents and intellectual properties: Different perspectives of willingness to pay and sell. Technovation, 33(1):13–24, January 2013. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016649721200106X.

Seclen-Luna:2024:EIP

[SLSCS24] Jean Pierre Seclen-Luna, Jubalt Alvarez Salazar, Christian A. Cancino, and Valentina Schmitt. The effects of innovations on Peruvian companies’ sales: the mediating role of KIBS. Technovation, 129(??):??, January 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001888.

Sheen:1995:ICA

[SM95] Margaret R. Sheen and Jillan C. MacBryde. The importance of complementary assets in the development of smart technology. Technovation, 15(2):99–109, March 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729596613X.

Schoonwinkel:1997:MMD

[SM97] Arnold Schoonwinkel and Garth W. Milne. Managing a microsatellite development programme within the university environment. Technovation, 17(1):1–9, January 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497296000764.

Stock:2001:OSP

[SM01] Gregory N. Stock and Christopher M. McDermott. Organizational and strategic predictors of manufacturing technology implementation success: an exploratory study. Technovation, 21(10):625–636, October 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000517.

Sun:2005:PMB

[SM05] Hongyi Sun and Tianchao Ma. A packing-multiple-boxes model for R&D project selection and scheduling. Technovation, 25(11):1355–1361, November 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204001166.

Sainio:2009:LRL

[SM09] Liisa-Maija Sainio and Emma Marjakoski. The logic of revenue logic? Strategic and operational levels of pricing in the context of software business. Technovation, 29(5):368– 378, May 2009. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649720800148X.

Sarpong:2012:MDV

[SM12a] David Sarpong and Mairi Maclean. Mobilising differential visions for new product innovation. Technovation, 32 (12):694–702, December 2012. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016649721200079X.

Sushandoyo:2012:TWR

[SM12b] Dedy Sushandoyo and Thomas Magnusson. A two-way relationship between multi-level technological change and organisational characteristics-cases involving the development of heavy hybrid buses. Technovation, 32(7–8):477–486, July/August 2012. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497212000545.

Smalheiser:2001:PET

[Sma01] N. R. Smalheiser. Predicting emerging technologies with the aid of text-based data mining: the micro approach. Technovation, 21(10):689–693, October 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000487.

Sokolov:2014:SCC

[SMC14] Alexander Sokolov, Vladimir Mesropyan, and Alexander Chulok. Supply chain cyber security: a Russian outlook. Technovation, 34(7):389–391, July 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214000054.

Stroh:2023:IEP

[SMD23] Tim Stroh, Anne-Laure Mention, and Cameron Duff. The impact of evolved psychological mechanisms on innovation and adoption: a systematic literature review. Technovation, 125(??):??, July 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000706.

Scuotto:2024:IIF

[SMGPP24] Veronica Scuotto, Domitilla Magni, Alexeis Garcia-Perez, and Marco Pironti. The impact of innovation failure: Entrepreneurship adversity or opportunity? Technovation, 131(??):??, March 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223002559.

Smith:1985:NAO

[Smi85] Stuart L. Smith. Notes for an address to the Ottawa association for applied economics. Technovation, 3(3):221–225, August 1985. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497285900161.

Smith:2007:PIW

[Smi07] D. J. Smith. The politics of innovation: Why innovations need a godfather. Technovation, 27(3):95–104, March 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497206000538.

Spano:2023:BVC

[SMI23] Rosanna Spanò, Maurizio Massaro, and Silvia Iacuzzi. Blockchain for value creation in the healthcare sector. Technovation, 120(??):??, February 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221002212.

Schiefer:2024:CAM

[SMvFM24] Tom Schiefer, Dominik Mahr, Paul C. van Fenema, and Kars Mennens. A collaborative approach to manage continuous service innovation. Technovation, 134(??):??, June 2024. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497224000798.

Sullivan:2021:LHM

[SMW21] Diane M. Sullivan, Matthew R. Marvel, and Marcus T. Wolfe. With a little help from my friends? How learning activities and network ties impact performance for high tech startups in incubators. Technovation, 101(??): ??, March 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649722030081X.

Swierczek:1992:TDT

[SN92] Fredric William Swierczek and Cary Nourie. Technology development in Thailand: a private sector view. Technovation, 12(3):145–159, April 1992. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729290032D.

Snaddon:1990:PWSa

[Sna90a] D. R. Snaddon. Productivity within sections of firms — I. The design of a productivity measure for a section of a firm. Technovation, 10(4):235–251, June 1990. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729090055O.

Snaddon:1990:PWSb

[Sna90b] D. R. Snaddon. Productivity within sections of firms — II. Measuring productivity in a section of a firm. Technovation, 10(5):291–304, July 1990. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729090014B.

Snaddon:1990:PWSc

[Sna90c] D. R. Snaddon. Productivity within sections of firms — III. Departmental productivity through information: the case of maintenance. Technovation, 10(6):363–377, September 1990. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729090001Z.

Snaddon:1990:PWSd

[Sna90d] D. R. Snaddon. Productivity within sections of firms — IV. Calibration of a sectional productivity measure using a cross-sectional study. Technovation, 10(7):453–465, October 1990. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729090025F.

Snaddon:1996:MM

[Sna96] D. Roy Snaddon. The manufacturing mix. Technovation, 16 (8):385–396, 444, August 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497296000193.

Snaddon:1998:MMC

[Sna98] D. Roy Snaddon. The manufacturing mix and consulting firms — how different are they in associating tasks with objectives? Technovation, 18(10):627–638, 658, October 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000315.

Snaddon:2001:CTL

[Sna01] D. R. Snaddon. Competition in transportation — a literature analysis. Technovation, 21(6):375–383, June 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000535.

Santamaria:2012:SIM

[SNM12] Lluís Santamaría, María Jesús Nieto, and Ian Miles. Service innovation in manufacturing firms: Evidence from Spain. Technovation, 32(2):144–155, February 2012. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211001258.

Salami:2002:ITL

[SO02] Abiodun Olusola Salami and Oluwole Osonubi. Improving the traditional landuse system through agrobiotechnology: a case study of adoption of vesicular arbuscular mycorrhiza (VAM) by resource-poor farmers in Nigeria. Technovation, 22(11):725–730, November 2002. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000948.

Sims:2006:SGU

[SO06] Martin A. Sims and Nicholas O’Regan. In search of gazelles using a research DNA model. Technovation, 26(8):943–954, August 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205001045.

Senyo:2020:UAF

[SO20] PK Senyo and Ellis L. C. Osabutey. Unearthing antecedents to financial inclusion through FinTech innovations. Technovation, 98(??):??, December 2020. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497220300365.

Smania:2024:UPT

[SOA^{+}24] Guilherme Sales Smania, Lauro Osiro, Néstor Fabián Ayala, Wim Coreynen, and Glauco H. S. Mendes. Unraveling paradoxical tensions in digital servitization ecosystems: an analysis of their interrelationships from the technology provider’s perspective. Technovation, 131(??):??, March 2024. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497224000075.

Soderlund:2008:BRB

[Sod08] Jonas Soderlund. Book review: The Business of Projects: Managing Innovation in Complex Products and Systems, Andrew Davies, Michael Hobday. Cambridge University Press, Cambridge. Technovation, 28(10):709–710, October 2008. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497207001332.

Sohal:1996:AAI

[Soh96] Amrik S. Sohal. Assessing AMT implementations: an empirical field study. Technovation, 16(8):377–384, 443–444, August 1996. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497296000454.

Sohal:1997:CPT

[Soh97a] Amrik S. Sohal. Computerised parts traceability: an implementation case study. Technovation, 17(10):583–591, 596– 598, October 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497297000394.

Sohal:1997:MRV

[Soh97b] Amrik S. Sohal. Manufacturing reorganisation at Varian Australia: a case study on value-added management. Technovation, 17(6):329–339, June 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497296001319.

Sohal:1999:INT

[Soh99] Amrik S. Sohal. Introducing new technology in to a small business: a case study. Technovation, 19(3):187–193, January 1999. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497298001102.

Scholten:2015:BTR

[SOKE15] Victor Scholten, Onno Omta, Ron Kemp, and Tom Elfring. Bridging ties and the role of research and start-up experience on the early growth of Dutch academic spin-offs. Technovation, 45–46(??):40–51, November/December 2015. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497215000425.

Souder:1983:OMT

[Sou83a] Wm. E. Souder. Organizing for modern technology and innovation: a review and synthesis. Technovation, 2(1):27– 44, February 1983. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497283900056.

Soundararajan:1983:RDT

[Sou83b] P. Soundararajan. Research development and transfer of technology approach by the National Research Development Corporation of India. Technovation, 2(1):55–60, February 1983. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649728390007X.

Souitaris:2001:ECD

[Sou01] Vangelis Souitaris. External communication determinants of innovation in the context of a newly industrialised country: a comparison of objective and perceptual results from Greece. Technovation, 21(1):25–34, January 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000146.

Sharifi:1996:PDM

[SP96] Sudi Sharifi and Kulwant S. Pawar. Product design as a means of integrating differentiation. Technovation, 16(5):255–264, 267–268, May 1996. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497296891255.

Sanchez:1998:ENHa

[SP98a] Angel Martinez Sanchez and Olga Urbina Perez. Entrepreneurship networks and high technology firms: the case of Aragon. Technovation, 18(5):335–345, ???? 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000078.

Sanchez:1998:ENHb

[SP98b] Angel Martinez Sanchez and Olga Urbina Perez. Entrepreneurship networks and high technology firms: the case of Aragon. Technovation, 18(5):365, ???? 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298800308.

Sanchez:2003:FNP

[SP03] A. M. Sánchez and M. P. Pérez. Flexibility in new product development: a survey of practices and its relationship with the product’s technological complexity. Technovation, 23(2):139–145, February 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000967.

Shin:2010:EOT

[SP10] Juneseuk Shin and Yongtae Park. Evolutionary optimization of a technological knowledge network. Technovation, 30(11–12):612–626, November/December 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210000672.

Scarra:2022:ITT

[SP22a] Deepa Scarrà and Andrea Piccaluga. The impact of technology transfer and knowledge spillover from Big Science: a literature review. Technovation, 116(??):??, August 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016649721930015X.

Seo:2022:WIL

[SP22b] Ribin Seo and Ji-Hoon Park. When is interorganizational learning beneficial for inbound open innovation of ventures? A contingent role of entrepreneurial orientation. Technovation, 116(??):??, August 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S016649722200061X.

Sanzo-Perez:2022:PBS

[SPÁG22] María José Sanzo-Pérez and Luis I. Álvarez-González. Partnerships between Spanish social enterprises and nonprofits: a rich hybridity-based setting for social innovation. Technovation, 110(??):??, February 2022. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001577.

Schwartz:2012:WDI

[SPFG12] Michael Schwartz, François Peglow, Michael Fritsch, and Jutta Günther. What drives innovation output from subsidized R&D cooperation? — project-level evidence from Germany. Technovation, 32(6):358–369, June 2012. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497212000338.

Saarenketo:2008:LKI

[SPKK08] Sami Saarenketo, Kaisu Puumalainen, Kalevi Kyläheiko, and Olli Kuivalainen. Linking knowledge and internationalization in small and medium-sized enterprises in the ICT sector. Technovation, 28(9):591–601, September 2008. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207001563.

Sick:2019:NFA

[SPLB19] Nathalie Sick, Nina Preschitschek, Jens Leker, and Stefanie Bröring. A new framework to assess industry convergence in high technology environments. Technovation, 84– 85(??):48–58, June/July 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497217300238.

Saura:2023:EBO

[SPMRS23] Jose Ramon Saura, Daniel Palacios-Marqués, and Domingo Ribeiro-Soriano. Exploring the boundaries of open innovation: Evidence from social media mining. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497221002285.

Sawers:2008:SSD

[SPO08] Jill L. Sawers, Marthinus W. Pretorius, and Leon A. G. Oerlemans. Safeguarding SMEs dynamic capabilities in technology innovative SME — large company partnerships in South Africa. Technovation, 28(4):171–182, April 2008. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207001290.

Sohal:1998:DPN

[SPP98] Amrik S. Sohal, Marcia Perry, and Trevor Pratt. Developing partnerships and networks: learning from practices in Australia. Technovation, 18(4):245–251, April 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297001181.

Shehabuddeen:2006:TPC

[SPP06] Noordin Shehabuddeen, David Probert, and Robert Phaal. From theory to practice: challenges in operationalising a technology selection framework. Technovation, 26(3):324– 335, March 2006. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497204002081.

Sestino:2020:ITB

[SPPG20] Andrea Sestino, Maria Irene Prete, Luigi Piper, and Gianluigi Guido. Internet of Things and Big Data as enablers for business digitalization strategies. Technovation, 98(??): ??, December 2020. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497220300456.

Sohal:2001:EAP

[SPS01] Amrik S. Sohal, Trevor Pratt, and Richard Schroder. An evaluation of the 1984–1996 Australian Passenger Motor Vehicle Plan. Technovation, 21(8):479–488, August 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000717.

Shaikh:2022:MRM

[SR22] Ibrahim Shaikh and Krithika Randhawa. Managing the risks and motivations of technology managers in open innovation: Bringing stakeholder-centric corporate governance into focus. Technovation, 114(??):??, June 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497221002182.

Sengupta:2023:RBU

[SR23] Abhijit Sengupta and Federica Rossi. The relationship between universities’ funding portfolios and their knowledge exchange profiles: a dynamic capabilities view. Technovation, 121(??):??, March 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222002371.

Sanchez:2006:IPF

[SRDS06] José Ignacio López Sánchez, Beatriz Minguela Rata, Antonio Rodríguez Duarte, and Francesco D. Sandulli. Is the Internet productive? A firm-level analysis. Technovation, 26(7):821–826, July 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000647.

Sainio:2012:CRI

[SRHL12] Liisa-Maija Sainio, Paavo Ritala, and Pia Hurmelinna-Laukkanen. Constituents of radical innovation-exploring the role of strategic orientations and market uncertainty. Technovation, 32(11):591–599, November 2012. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497212000764.

Saenz-Royo:2023:AVP

[SRLR23] Carlos Sáenz-Royo and Álvaro Lozano-Rojo. Authoritarianism versus participation in innovation decisions. Technovation, 124(??):??, June 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497223000524.

Saastamoinen:2018:SSP

[SRT18] Jani Saastamoinen, Helen Reijonen, and Timo Tammi. Should SMEs pursue public procurement to improve innovative performance? Technovation, 69(??):2–14, January 2018. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497216303674.

Supapol:1994:RIP

[SS94] Atipol Bhanich Supapol and Fredric William Swierczek. The role of intellectual property rights in stimulating commercialization in ASEAN: lessons from Canada. Technovation, 14(3):181–195, April 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497294900558.

See [Ano94-30].

Sohail:2003:UTP

[SS03] M. S. Sohail and A. S. Sohal. The use of third party logistics services: a Malaysian perspective. Technovation, 23(5):401– 408, May 2003. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497202000032.

Siriram:2004:LTM

[SS04] R. Siriram and D. R. Snaddon. Linking technology management, transaction processes and governance structures. Technovation, 24(10):779–791, October 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000245.

Siriram:2005:VLT

[SS05] R. Siriram and D. R. Snaddon. Verifying links in technology management, transaction processes and governance structures. Technovation, 25(4):321–337, April 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203001494.

Smith:2007:UAT

[SS07] Roger Smith and Nawaz Sharif. Understanding and acquiring technology assets for global competition. Technovation, 27(11):643–649, November 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497207000491.

Slayton:2016:RIS

[SS16] Rebecca Slayton and Graham Spinardi. Radical innovation in scaling up: Boeing’s Dreamliner and the challenge of socio-technical transitions. Technovation, 47(??):47–58, January 2016. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497215000619.

Sasaki:2021:IPT

[SS21] Hajime Sasaki and Ichiro Sakata. Identifying potential technological spin-offs using hierarchical information in international patent classification. Technovation, 100(??):??, February 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649722030064X.

Sheth:2022:AFC

[SS22] Ananya Sheth and Joseph V. Sinfield. An analytical framework to compare innovation strategies and identify simple rules. Technovation, 115(??):??, July 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497222000815.

Schiuma:2023:ILO

[SS23] Giovanni Schiuma and Francesco Santarsiero. Innovation labs as organisational catalysts for innovation capacity development: a systematic literature review. Technovation, 123(??):??, May 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497223000019.

Shekhar:2023:RCI

[SSC23] Himanshu Shekhar, Krishna Satyanarayana, and Deepak Chandrashekar. Role and contributions of an incubator in academic intrapreneurship — an examination. Technovation, 126(??):??, August 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001323.

Santarsiero:2023:DTH

[SSCH23] Francesco Santarsiero, Giovanni Schiuma, Daniela Carlucci, and Nina Helander. Digital transformation in healthcare organisations: the role of innovation labs. Technovation, 122 (??):??, April 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497222001912.

Salavisa:2012:TIN

[SSF12] Isabel Salavisa, Cristina Sousa, and Margarida Fontes. Topologies of innovation networks in knowledge-intensive sectors: Sectoral differences in the access to knowledge and complementary assets through formal and informal ties. Technovation, 32(6):380–399, June 2012. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497212000296.

Sohag:2023:FCC

[SSGN23] Kazi Sohag, S. M. Riad Shams, Anna Gainetdinova, and Fabio Nappo. Frequency connectedness and cross-quantile dependence among Medicare, medicine prices and healthtech equity. Technovation, 120(??):??, February 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722200030X.

Soltanzadeh:2024:CCD

[SSKB24] Javad Soltanzadeh, Navid Sahebjamnia, Elnaz Mesma Khosroshahi, and Abderaouf Bouguerra. Commercializing Covid-19 diagnostic technologies: a review of challenges, success factors, and insights from the profiting from innovation framework. Technovation, 130(??):??, February 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223002146.

Scaglione:2009:WAS

[SSM09] Miriam Scaglione, Roland Schegg, and Jamie Murphy. Website adoption and sales performance in Valais’ hospitality industry. Technovation, 29(9):625–631, September 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209000819.

Suomala:2004:CAS

[SSP04] P. Suomala, M. Sievänen, and J. Paranko. Customization from the after sales point of view-implications of product and item customization for spare-part business. Technovation, 24(10):831–840, October 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000117.

Stornelli:2024:EDC

[SSRO24] Aldo Stornelli, Christopher Simms, Wiebke Reim, and Sercan Ozcan. Exploring the dynamic capabilities of technology provider ecosystems: a study of smart manufacturing projects. Technovation, 130(??):??, February 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223002365.

Simeone:2017:ADA

[SSS17] Luca Simeone, Giustina Secundo, and Giovanni Schiuma. Adopting a design approach to translate needs and interests of stakeholders in academic entrepreneurship: the MIT Senseable City Lab case. Technovation, 64–65(??):58–67, June 2017. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497216304886.

Sharma:2022:DIO

[SSW22] Abhijit Sharma, Cristina Sousa, and Richard Woodward. Determinants of innovation outcomes: the role of institutional quality. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222001092.

Sanchez:1995:UIR

[ST95] Angel Martinez Sanchez and Ana-Clara Pastor Tejedor. University-industry relationships in peripheral regions: the case of Aragon in Spain. Technovation, 15(10):613–625, December 1995. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729599329E.

Soares:2021:TST

[ST21] Thiago J. Soares and Ana L. V. Torkomian. TTO’s staff and technology transfer: Examining the effect of employees’ individual capabilities. Technovation, 102(??):??, April 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497220300857.

Staropoli:1998:CRP

[Sta98] Carine Staropoli. Cooperation in R&D in the pharmaceutical industry — The network as an organizational innovation governing technological innovation. Technovation, 18(1):13–23, January 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497297001077.

Saad:2024:MLP

[STA24] Elie Abi Saad, Nathalie Tremblay, and Marine Agogué. A multi-level perspective on innovation intermediaries: the case of the diffusion of digital technologies in healthcare. Technovation, 129(??):??, January 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223002109.

Sternberg:1989:ICT

[Ste89] Rolf Sternberg. Innovation centres and their importance for the growth of new technology-based firms: experience gained from the Federal Republic of Germany. Technovation, 9(8):681–694, December 1989. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900060.

Sternberg:2014:SFU

[Ste14a] Rolf Sternberg. Success factors of university-spin-offs: Regional government support programs versus regional environment. Technovation, 34(3):137–148, March 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213001399.

Stevens:2014:FFE

[Ste14b] Eric Stevens. Fuzzy front-end learning strategies: Exploration of a high-tech company. Technovation, 34(8):431– 440, August 2014. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497213001624.

Suikki:2006:PMC

[STH06] Raija Suikki, Raija Tromstedt, and Harri Haapasalo. Project management competence development framework in turbulent business environment. Technovation, 26(5– 6):723–738, May/June 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497204002172.

Stoop:1999:AMC

[Sto99] M. L. M. Stoop. Application of a mathematical calculation model to reduce slaughterhouse (water) pollution in developing countries. Technovation, 19(5):323–331, February 1999. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497298001230.

Stoop:2003:WMP

[Sto03] M. L. M. Stoop. Water management of production systems optimised by environmentally oriented integral chain management:: case study of leather manufacturing in developing countries. Technovation, 23(3):265–278, March 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201001171.

Stojcic:2021:SPO

[Sto21] Nebojsa Stojcić. Social and private outcomes of green innovation incentives in European advancing economies. Technovation, 104(??):??, June 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497221000511.

Solgaard:1988:ASR

[STS88] Hans S. Solgaard and Erik Trolle-Schultz. An analysis of the strategic responsiveness of firms. Technovation, 8(1–3): 155–178, ???? 1988. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497288900594.

Sohal:2003:TBS

[STZ03] Amrik S. Sohal, Milé Terziovski, and Ambika Zutshi. Team-based strategy at Varian Australia: a case study. Technovation, 23(4):349–357, April 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201001146.

Sawang:2011:WAN

[SU11] Sukanlaya Sawang and Kerrie L. Unsworth. Why adopt now? Multiple case studies and survey studies comparing small, medium and large firms. Technovation, 31(10–11):554–559, October/November 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211000769.

Subrahmanya:2005:PTI

[Sub05] M. H. Bala Subrahmanya. Pattern of technological innovations in small enterprises: a comparative perspective of Bangalore (India) and Northeast England (UK). Technovation, 25(3):269–280, March 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000944.

Suchard:1987:WII

[Suc87] Hazel T. Suchard. Worker involvement in implementing new technology — a comment. Technovation, 6(1):69–71, April 1987. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649728790040X.

Sundbo:1996:BES

[Sun96] Jon Sundbo. The balancing of empowerment. A strategic resource based model of organizing innovation activities in service and low-tech firms. Technovation, 16(8): 397–409, 445–446, August 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497296000296.

Sun:2000:CFP

[Sun00] Hongyi Sun. Current and future patterns of using advanced manufacturing technologies. Technovation, 20(11):631–641, November 2000. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497200000079.

Surrey:1992:TCP

[Sur92] John Surrey. Technical change and productivity growth in the British coal industry, 1974–1990. Technovation, 12(1):15–39, February 1992. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729290030L.

Sun:2001:HSH

[SV01] Hongyi Sun and Patri K. Venuvinod. The human side of holonic manufacturing systems. Technovation, 21(6):353– 360, June 2001. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497200000559.

Santos-Vijande:2007:IOI

[SVÁG07] María Leticia Santos-Vijande and Luis Ignacio Álvarez-González. Innovativeness and organizational innovation in total quality oriented firms: the moderating role of market turbulence. Technovation, 27(9):514–532, September 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000752.

Soetanto:2015:GRB

[SvG15] Danny Soetanto and Marina van Geenhuizen. Getting the right balance: University networks’ influence on spin-offs’ attraction of funding for innovation. Technovation, 36–37(??):26–38, February/March 2015. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214001576.

Sala-Vilar:2024:HDI

[SVLYT24] Lluís Ròmul Sala-Vilar, Jason Li-Ying, and Tim Traunecker. How do innovation intermediaries’ business models cope with their need to develop new digital services? Technovation, 131(??):??, March 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223002614.

Spanos:2015:AII

[SVV15] Yiannis E. Spanos, Nicholas S. Vonortas, and Irini Voudouris. Antecedents of innovation impacts in publicly funded collaborative R&D projects. Technovation, 36–37(??):53–64, February/March 2015. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214001059.

Sonneborn:1991:RCS

[SW91] Marcene Sonneborn and David Wilemon. R&D’s contributions to strategic decision-making: Rationale, content, and process. Technovation, 11(5):267–281, July 1991. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729190026Z.

Salter:2004:BRS

[SW04] A. Salter and J. Whyte. Book review: Serious play: how the world’s best companies simulate to innovate: Michael Schrage, Harvard Business School Press, Boston, Massachusetts. Technovation, 24(3):277–278, March 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203001780.

Sun:2005:CSF

[SW05] Hongyi Sun and Wong Chung Wing. Critical success factors for new product development in the Hong Kong toy industry. Technovation, 25(3):293–303, March 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720300097X.

Saghiri:2021:ESD

[SW21] Soroosh Saghiri and Richard Wilding. On the effectiveness of supplier development programs: the role of supplyside moderators. Technovation, 103(??):??, May 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000158.

Steele:2022:MTI

[SW22] Logan M. Steele and Logan L. Watts. Managing technical innovation: a systematic review of 11 leader functions. Technovation, 115(??):??, July 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://

www.sciencedirect.com/science/article/pii/S0166497222000864.

Swierczek:1992:SBIa

[Swi92a] Fredric William Swierczek. Strategies for business innovation: Evaluating the prospects of incubation in Thailand. Technovation, 12(8):521–533, December 1992. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729290075S. See [Swi92b].

Swierczek:1992:SBIb

[Swi92b] Fredric William Swierczek. Strategies for business innovation: Evaluating the prospects of incubation in Thailand: Technovation, 12:8 (1992) 521–533. Technovation, 12 (8):537–538, December 1992. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729290079W.

See [Swi92a].

Schumann:2012:TMS

[SWW12] Jan H. Schumann, Nancy V. Wünderlich, and Florian Wangenheim. Technology mediation in service delivery: a new typology and an agenda for managers and academics. Technovation, 32(2):133–143, February 2012. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649721100143X.

Shao:2024:LCD

[SXS24] Yanmin Shao, Kunliang Xu, and Yuan George Shan. Leveraging corporate digitalization for green technology innovation: the mediating role of resource endowments. Technovation, 133(??):??, May 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722400049X.

Sher:2005:EIC

[SY05] Peter J. Sher and Phil Y. Yang. The effects of innovative capabilities and R&D clustering on firm performance: the evidence of Taiwan’s semiconductor industry. Technovation, 25(1):33–43, January 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000683.

Schlecht:2014:IPR

[SY14] Lisa Schlecht and Maria Yang. Impact of prototyping resource environments and timing of awareness of constraints on idea generation in product design. Technovation, 34(4): 223–231, April 2014. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497213001284.

Soenksen:2017:SGP

[SY17] Luis R. Soenksen and Youseph Yazdi. Stage-gate process for life sciences and medical innovation investment. Technovation, 62–63(??):14–21, April 2017. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497216302474.

Su:2023:TEU

[SYCH23] Taoyong Su, Yuzhu Yu, Yongheng Chen, and Wanrong Hou. On or off: the triggering effect of underperformance duration on cooperative innovation. Technovation, 126(??):??, August 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223001281.

Sharma:1994:IRV

[SYK94] Rajesh K. Sharma, Karm Raj Yadav, and P. M. Kothari. Innovative recycling of vegetable protein waste for increased crop productivity. Technovation, 14(1):31–36, February 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729490068X.

Szakonyi:1990:EDS

[Sza90] Robert Szakonyi. Establishing discipline in the selection, planning, and carrying out of R&D projects. Technovation, 10(7):467–486, October 1990. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729090026G.

Szanto:1992:EEW

[Szá92] Borisz Szántó. Eastern Europe and the world in the light of forecasting scientific and technological activity. Technovation, 12(6):369–377, September 1992. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729290002Y.

Szanto:1994:ICHa

[Szá94a] Borisz Szántó. Innovation in crisis: Hungary before and after the watershed of 1989. Technovation, 14(9):601–611, November 1994. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497294900426. See [Szá94b].

Szanto:1994:ICHb

[Szá94b] Borisz Szántó. Innovation in crisis: Hungary before and after the watershed of 1989: Technovation, 14(9) (1994) 601–611. Technovation, 14(9):623, November 1994. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497294900469. See [Szá94a].

Szanto:1996:SPV

[Szá96] Borisz Szántó. Science policy vs. technology policy? Technovation, 16(8):411–420, 446–447, August 1996. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497296000430.

Szanto:2001:PG

[Szá01] Borisz Szántó. The paradigm of globalism. Technovation, 21(10):673–687, October 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497200000821.

Szanto:2005:UCI

[Szá05] Borisz Szántó. The unconventional challenges of InfoTech. Technovation, 25(5):469–476, May 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203001949.

Samra:2019:CMN

[SZLR19] Yassir M. Samra, Haisu Zhang, Gary S. Lynn, and Richard R. Reilly. Crisis management in new product development: a tale of two stories. Technovation, 88(??): Article 102038, December 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497217306302.

Song:2023:IGT

[SZSR23] Yang Song, Zhiyuan Zhang, Jean-Michel Sahut, and Ofir Rubin. Incentivizing green technology innovation to confront sustainable development. Technovation, 126(??):??, August 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223000998.

Szwejczewski:2007:BRO

[Szw07] Marek Szwejczewski. Book review: Operations Management: a Strategic Approach: Alison Bettley, David Mayle, Tarek Tantoush (Eds.), SAGE Publications, Beverley Hills, 2005 (328 pp., Paperback £24.99), ISBN 1-4129-1903-7. Technovation, 27(8):489, August 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000193.

Tomes:1998:MTUa

[TA98a] Anne Tomes and Peter Armstrong. Managing tensions in user groups. Technovation, 18(2):79–89, February 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297000801.

Tomes:1998:MTUb

[TA98b] Anne Tomes and Peter Armstrong. Managing tensions in user groups. Technovation, 18(2):141, February 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298800114.

Tolstoy:2010:NRC

[TA10] Daniel Tolstoy and Henrik Agndal. Network resource combinations in the international venturing of small biotech firms. Technovation, 30(1):24–36, January 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720900090X.

Tomes:1996:UGA

[TAC96] Anne Tomes, Peter Armstrong, and Murray Clark. User groups in action: the management of user inputs in the NPD process. Technovation, 16(10):541–551, October 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297828945.

Taggart:1997:RCU

[Tag97] James H. Taggart. R&D complexity in UK subsidiaries of manufacturing multinational corporations. Technovation, 17(2):73–82, 103, February 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497296000879.

Thomas:2022:PEE

[TAG22] Llewellyn D. W. Thomas, Erkko Autio, and David M. Gann. Processes of ecosystem emergence. Technovation, 115(??):??, July 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497221002224.

Taiwo:1998:PCH

[Tai98] Kehinde A. Taiwo. The potential of cowpea as human food in Nigeria. Technovation, 18(6–7):469–481, January 1998.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000157.

Tickle:2011:DAB

[TAM11] Matthew Tickle, Dotun Adebanjo, and Zenon Michaelides. Developmental approaches to B2B virtual communities. Technovation, 31(7):296–308, July 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211000563.

Tamura:2016:NIP

[Tam16] Suguru Tamura. A new intellectual property metric for standardization activities. Technovation, 48–49(??):87– 98, February/March 2016. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497216000080.

Tanaka:1987:PPI

[Tan87] Yoshihiko Tanaka. A proposal for a product innovation method through fusion strategy in the intelligent computer aided management (ICAM) age. Technovation, 6(1):3–23, April 1987. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497287900356.

Tan:1995:ETT

[Tan95] Raykun R. Tan. Establishing technology transfer infrastructure as a strategy for promoting manufacturing automation in Taiwan. Technovation, 15(7):407–421, September 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729596591G.

Tang:1998:IMIb

[Tan98a] H. K. Tang. An integrative model of innovation in organisations. Technovation, 18(5):361, ???? 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298800278.

Tang:1998:IMIa

[Tan98b] H. K. Tang. An integrative model of innovation in organizations. Technovation, 18(5):297–309, ???? 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000091.

Tang:1998:IOI

[Tan98c] H. K. Tang. An inventory of organizational innovativeness. Technovation, 19(1):41–51, November 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000777.

Tang:2008:BRD

[Tan08] Puay Tang. Book review: Digital Era Governance, Patrick Dunleavy, Helen Margetts, Simon Bastow, Jane Tinkler, Oxford University Press, Great Clarendon Street, Oxford (2006). 289 pp., ISBN 0-19-929619-7, 978-0-19-929619-4. Technovation, 28(9):630–631, September 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000612.

Tatham:2004:BRM

[Tat04] Rod Tatham. Book review: Managing Web Projects: J. Rodney Turner, Gower Publishing, Gower House, Croft Road, Aldershot, Hampshire GU11 3HR, UK, 228 pages, ISBN 0-566-08567-4. Technovation, 24(12):1011, December 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204001415.

Tyebjee:1984:VCI

[TB84] Tyzoon T. Tyebjee and Albert V. Bruno. Venture capital: Investor and investee perspectives. Technovation, 2(3):185– 208, June 1984. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497284900038.

Teirlinck:2024:ECI

[TB24] Peter Teirlinck and Anneleen Bruylant. Extended cyclic innovation model as a tool for failure identification in innovation management. Case study of 25 years of CMOS image sensor technology in Belgium. Technovation, 129(??): ??, January 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223002110.

Bae:1986:TDP

[tBL86] Zong tae Bae and Jinjoo Lee. Technology development patterns of small and medium sized companies in the Korean machinery industry. Technovation, 4(4):279–296, August 1986. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497286900313.

Thomas:2021:EUS

[TBSM21] V. J. Thomas, Martin Bliemel, Cynthia Shippam, and Elicia Maine. Endowing university spin-offs pre-formation: Entrepreneurial capabilities for scientist-entrepreneurs. Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497218307302.

Ta-Cheng:1997:CDM

[TC97] Hsiao Ta-Cheng. Capability development and management of R&D professionals in a developing context, Taiwan. Technovation, 17(10):569–581, 595–596, October 1997. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297000473.

Tsao:2022:IDF

[TCC^{+}22] Yu-Chung Tsao, Yu-Kai Chen, Shih-Hao Chiu, Jye-Chyi Lu, and Thuy-Linh Vu. An innovative demand forecasting approach for the server industry. Technovation, 110(??): ??, February 2022. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497221001528.

Trott:1995:ITT

[TCHS95] P. Trott, M. Cordey-Hayes, and R. A. F. Seaton. Inward technology transfer as an interactive process. Technovation, 15(1):25–43, February 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729596609W.

Tian:2022:PBS

[TCMS22] Jiamian Tian, Wim Coreynen, Paul Matthyssens, and Lei Shen. Platform-based servitization and business model adaptation by established manufacturers. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497221000031.

Tien:2005:ESC

[TCT05] Shiaw-Wen Tien, Yi-Chan Chung, and Chih-Hung Tsai. An empirical study on the correlation between environmental design implementation and business competitive advantages in Taiwan’s industries. Technovation, 25(7):783–794, July 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204000021.

Torimiro:2005:CIE

[TDA05] Dixon O. Torimiro and Erlinda A. Dionco-Adetayo. Children involvement in entrepreneurship in rural communities: an attitudinal analysis. Technovation, 25(6):683–689, June 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203001779.

Therrien:2011:INC

[TDC11] Pierre Therrien, David Doloreux, and Tyler Chamberlin. Innovation novelty and (commercial) performance in the service sector: a Canadian firm-level analysis. Technovation, 31(12):655–665, December 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211001088.

Tongur:2014:BMD

[TE14] Stefan Tongur and Mats Engwall. The business model dilemma of technology shifts. Technovation, 34(9):525–535, September 2014. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497214000248.

Tomes:2000:TEI

[TEA00] Anne Tomes, Rosie Erol, and Peter Armstrong. Technological entrepreneurship: Integrating technological and product innovation. Technovation, 20(3):115–127, March 2000. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001169.

Teichert:1993:SPI

[Tei93] Thorsten Teichert. The success potential of international R&D cooperation. Technovation, 13(8):519–532, December 1993. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497293900665. See [Ano93m].

Teirlinck:2022:ERE

[Tei22] Peter Teirlinck. Enhancing R&D employment slack during environmental turbulence: Triggers and firm performance consequences for R&D-active SMEs. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222001699.

Tether:2000:SFI

[Tet00] Bruce S. Tether. Small firms, innovation and employment creation in Britain and Europe: a question of expectations …. Technovation, 20(2):109–113, February 2000. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299000930.

Tortorella:2022:CHD

[TFS^{+}22] Guilherme L. Tortorella, Flavio S. Fogliatto, Tarcísio A. Saurin, Leandro M. Tonetto, and Duncan McFarlane. Contributions of Healthcare 4.0 digital applications to the resilience of healthcare organizations during the COVID-19 outbreak. Technovation, 111(??):??, March 2022. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001607.

Temiz:2022:ODL

[THBW22] Serdar Temiz, Marcus Holgersson, Joakim Björkdahl, and Martin W. Wallin. Open data: Lost opportunity or unrealized potential? Technovation, 114(??):??, June 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000827.

Toroslu:2023:OIN

[THC^{+}23] Adrian Toroslu, Andrea M. Herrmann, Maryse M. H. Chappin, Brita Schemmann, and Carolina Castaldi. Open innovation in nascent ventures: Does openness influence the speed of reaching critical milestones? Technovation, 124 (??):??, June 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223000433.

Tsouri:2022:KRE

[THHS22] Maria Tsouri, Teis Hansen, Jens Hanson, and Markus Steen. Knowledge recombination for emerging technological innovations: the case of green shipping. Technovation, 114(??):??, June 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222000013.

Thomas:1988:C

[Tho88a] Steve Thomas. Comment. Technovation, 7(4):277–279, October 1988. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649728890003X.

Thomas:1988:DANa

[Tho88b] Steve Thomas. The development and appraisal of nuclear power. Part I. Regulatory and institutional aspects. Technovation, 7(4):281–304, October 1988. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900041.

Thomas:1988:DANb

[Tho88c] Steve Thomas. The development and appraisal of nuclear power. Part II. The role of technical change. Technovation, 7(4):305–339, October 1988. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497288900053.

Thomas:1999:WMR

[Tho99] Christine Thomas. Waste management and recycling in Romania: a case study of technology transfer in an economy in transition. Technovation, 19(6–7):365–371, June 1999. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299000280.

Thongpapanl:2012:CLT

[Tho12] Narongsak (Tek) Thongpapanl. The changing landscape of technology and innovation management: an updated ranking of journals in the field. Technovation, 32(5):257–271, May 2012. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497212000028.

Thumm:2005:PGI

[Thu05] Nikolaus Thumm. Patents for genetic inventions: a tool to promote technological advance or a limitation for upstream inventions? Technovation, 25(12):1410–1417, December 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497204001154.

Thun:2018:IMP

[Thu18] Eric Thun. Innovation at the middle of the pyramid: State policy, market segmentation, and the Chinese automotive sector. Technovation, 70–71(??):7–19, ???? 2018. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016649721830124X.

Tidd:2002:KEL

[TI02] Joe Tidd and Yasuhiko Izumimoto. Knowledge exchange and learning through international joint ventures: an Anglo-Japanese experience. Technovation, 22(3):137–145, March 2002. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201000062.

Tikas:2023:TIC

[Tik23] Gaurav Dilip Tikas. Team innovation capability: Scale development and validation. Technovation, 126(??):??, August 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223000846.

Tither:1990:CST

[Tit90a] Denis Tither. A case study of technology transfer and funding mechanisms in an industrially supported multi-centred university research initiative. Technovation, 10(1):39–46, February 1990. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729090041H.

Tither:1990:RAC

[Tit90b] Denis Tither. Research activity in carbon and diamond film technology in the United Kingdom; January-December 1988. Technovation, 10(1):31–38, February 1990. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729090040Q.

Tither:1994:PFCa

[Tit94a] Denis Tither. The people factor in collaboration and technology transfer. Technovation, 14(5):283–286, June 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729490071X.

Tither:1994:PFCb

[Tit94b] Denis Tither. The people factor in collaboration and technology transfer: observations and profile. Technovation, 14 (5):345, June 1994. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497294900760.

Tither:1994:TESa

[Tit94c] Denis Tither. Time for elitism in small and medium-size manufacturing companies. Technovation, 14(10):689–692, December 1994. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497294901147.

Tither:1994:TESb

[Tit94d] Denis Tither. Time for elitism in small and medium-size manufacturing companies. Technovation, 14(10):697, December 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294901198.

Tahi:2022:PPI

[TKBC22] Sofiane Tahi, Wafa Khlif, Khaled Belghoul, and Vanessa Casadella. Public-private innovation networks in services: Revisiting PPPs with servitization. Technovation, 118 (??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497221001176.

Tynnhammar:2018:CIM

[TKS18] Marcus Tynnhammar, Olga Kokshagina, and Sabrina Schneider. Composing the innovation management symphony — a note on the 2017 ISPIM Innovation Conference in Vienna. Technovation, 74–75(??):66, June/July 2018. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497218301597.

Takeda:2008:AGA

[TKSM08] Yoshiyuki Takeda, Yuya Kajikawa, Ichiro Sakata, and Katsumori Matsushima. An analysis of geographical agglomeration and modularized industrial networks in a regional cluster: a case study at Yamagata prefecture in Japan. Technovation, 28(8):531–539, August 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208000023.

Treibich:2013:DVI

[TKT13] Tania Treibich, Kornelia Konrad, and Bernhard Truffer. A dynamic view on interactions between academic spin-offs and their parent organizations. Technovation, 33(12):450– 462, December 2013. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497213000746.

Talaei-Khoei:2024:HDI

[TKYM24] Amir Talaei-Khoei, Alan T. Yang, and Masialeti Masialeti. How does incorporating ChatGPT within a firm reinforce agility-mediated performance? The moderating role of innovation infusion and firms’ ethical identity. Technovation, 132(??):??, April 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000257.

Todtling:2009:DDT

[TLK09] Franz Tödtling, Patrick Lehner, and Alexander Kaufmann. Do different types of innovation rely on specific kinds of knowledge interactions? Technovation, 29(1):59–71, January 2009. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497208000576.

To:2007:SMI

[TLL07] Pui-Lai To, Chechen Liao, and Tzu-Hua Lin. Shopping motivations on Internet: a study based on utilitarian and hedonic value. Technovation, 27(12):774–787, December 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000144.

Telg:2023:HFI

[TLL23] Nina Telg, Boris Lokshin, and Wilko Letterie. How formal and informal intellectual property protection matters for firms’ decision to engage in coopetition: the role of environmental dynamism and competition intensity. Technovation, 124(??):??, June 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000627.

Therrien:2003:HIC

[TM03] Pierre Therrien and Pierre Mohnen. How innovative are Canadian firms compared to some European firms? A comparative look at innovation surveys. Technovation, 23(4): 359–369, April 2003. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497201001122.

Terziovski:2006:MPS

[TM06] Milé Terziovski and John P. Morgan. Management practices and strategies to accelerate the innovation cycle in the biotechnology industry. Technovation, 26(5–6):545–552, May/June 2006. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649720400210X.

Todorovic:2011:EUE

[TMG11] Zelimir William Todorovic, Rod B. McNaughton, and Paul Guild. ENTRE-U: an entrepreneurial orientation scale for universities. Technovation, 31(2–3):128–137, February/ March 2011. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497210001173.

Tajzadeh-Namin:1996:BIT

[TNSG96] Ardeshir Abolfazle Tajzadeh-Namin, Ebrahim Shayan, and Murray Gillin. Barriers to international trade between Iran and Australia. Technovation, 16(7):317–325, July 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497296000181.

Takata:2022:NEH

[TNY^{+}22] Megumi Takata, Koichi Nakagawa, Mari Yoshida, Terumasa Matsuyuki, Toshihiko Matsuhashi, Kosuke Kato, and Ashley J. Stevens. Nurturing entrepreneurs: How do technology transfer professionals bridge the Valley of Death in Japan? Technovation, 109(??):??, January 2022. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497218305984.

Taiwo:2001:TCU

[TOA01] K. A. Taiwo, E. N. Ogwu, and O. O. Ajibola. Technological considerations in the utilisation and maintenance of household kitchen equipment. Technovation, 21(11):747– 755, November 2001. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497200000651.

Todtling:1994:RNHb

[Töd94a] Franz Tödtling. Regional networks of high-technology firms — the case of Greater Boston. Technovation, 14(5):349, June 1994. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497294900922.

Todtling:1994:RNHa

[Töd94b] Franz Tödtling. Regional networks of high technology firms — the case of the Greater Boston region. Technovation, 14(5):323–343, June 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497294900752.

Tomer:1990:DWC

[Tom90] John F. Tomer. Developing world class organization: Investing in organizational capital. Technovation, 10(4):253– 263, June 1990. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729090056P.

Tomes:2003:UGS

[Tom03] A. Tomes. UK government science policy: the ‘enterprise deficit’ fallacy. Technovation, 23(10):785–792, October 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203000142.

Tonurist:2015:FAR

[Tõn15] Piret Tõnurist. Framework for analysing the role of state owned enterprises in innovation policy management: the case of energy technologies and Eesti Energia. Technovation, 38(??):1–14, April 2015. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214001217.

Taiwo:2000:TAP

[TOS^{+}00] K. A. Taiwo, O. K. Owolarafe, L. A. Sanni, J. O. Jeje, K. Adeloye, and O. O. Ajibola. Technological assessment of palm oil production in Osun and Ondo states of Nigeria. Technovation, 20(4):215–223, April 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001108.

Tekic:2024:CDS

[TP24] Anja Tekic and Diana Vilma Alfonzo Pacheco. Contest design and solvers’ engagement behaviour in crowdsourcing: the neo-configurational perspective. Technovation, 132(??): ??, April 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497224000361.

Theodorakopoulos:2012:TTU

[TPB12] Nicholas Theodorakopoulos, Deycy Janeth Sánchez Preciado, and David Bennett. Transferring technology from university to rural industry within a developing economy context: the case for nurturing communities of practice. Technovation, 32(9–10):550–559, September/October 2012. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497212000533.

Trabucchi:2023:TNG

[TPBM23] Daniel Trabucchi, Andrea S. Patrucco, Tommaso Buganza, and Giacomo Marzi. Is transparency the new green? How business model transparency influences digital service adoption. Technovation, 126(??):??, August 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001141.

Terjesen:2013:NRV

[TPFD13] Siri Terjesen, Pankaj C. Patel, James O. Fiet, and Rodney D’Souza. Normative rationality in venture capital financing. Technovation, 33(8–9):255–264, August/September 2013. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649721200123X.

Tietze:2015:OOH

[TPH15] Frank Tietze, Thorsten Pieper, and Cornelius Herstatt. To own or not to own: How ownership impacts user innovation — an empirical study. Technovation, 38(??):50–63, April 2015. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497214001588.

Thanasopon:2016:ROF

[TPV16] Bundit Thanasopon, Thanos Papadopoulos, and Richard Vidgen. The role of openness in the fuzzy front-end of service innovation. Technovation, 47(??):32–46, January 2016. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497215000814.

Tranfield:2000:BRV

[Tra00] David Tranfield. Book review: Valuing Technology — Organisation, Cultures and Change: Janice McLaughlin, Paul Rosen, David Skinner and Andrew Webster; Routledge, hardback, ISBN 0-415-19210-2. Technovation, 20(6):343– 344, June 2000. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497200000055.

Tremblay:1998:ITH

[Tre98] Pierre J. Tremblay. Informal thinkering — how is it important?: Case studies of technical change processes: an industrialized-industrializing country comparison. Technovation, 18(12):751–763, October 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000807.

Tehrani:2024:DAR

[TRR^{+}24] Ali N. Tehrani, Subhasis Ray, Sanjit K. Roy, Richard L. Gruner, and Francesco P. Appio. Decoding AI readiness: an in-depth analysis of key dimensions in multinational corporations. Technovation, 131(??):??, March 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223002596.

Tao:2023:LLN

[TRW23] Xiangming (Tommy) Tao, Paul J. A. Robson, and Catherine L. Wang. To learn or not to learn from new product development project failure: the roles of failure experience and error orientation. Technovation, 127(??):??, September 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223001414.

Trolle-Schultz:1988:RRD

[TS88] Erik Trolle-Schultz. The re-regulation of the Danish capital market. Technovation, 8(1–3):71–86, ???? 1988. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900545.

Terziovski:2000:ACI

[TS00] Milé Terziovski and Amrik S. Sohal. The adoption of continuous improvement and innovation strategies in Australian manufacturing firms. Technovation, 20(10):539–550, October 2000. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649729900173X.

Tsoutsos:2005:SDR

[TS05] Theocharis D. Tsoutsos and Yeoryios A. Stamboulis. The sustainable diffusion of renewable energy technologies as an example of an innovation-focused policy. Technovation, 25(7):753–761, July 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203002207.

Teirlinck:2013:RCR

[TS13] Peter Teirlinck and André Spithoven. Research collaboration and R&D outsourcing: Different R&D personnel requirements in SMEs. Technovation, 33(4–5):142–153, April/May 2013. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497212001393.

Tsai:2005:RPF

[Tsa05a] Kuen-Hung Tsai. R&D productivity and firm size: a nonlinear examination. Technovation, 25(7):795–803, July 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203002190.

Tsang:2005:GIE

[Tsa05b] Denise Tsang. Growth of indigenous entrepreneurial software firms in cities. Technovation, 25(11):1331–1336, November 2005. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497204000689.

Thurik:2013:REE

[TSA13a] A. R. Thurik, E. Stam, and D. B. Audretsch. The rise of the entrepreneurial economy and the future of dynamic capitalism. Technovation, 33(8–9):302–310, August/September 2013. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497213000771.

Tsai:2013:MDM

[Tsa13b] Bi-Huei Tsai. Modeling diffusion of multi-generational LCD TVs while considering generation-specific price effects and consumer behaviors. Technovation, 33(10–11):345– 354, October/November 2013. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016649721300062X.

Terziovski:2002:BPP

[TSH02] Mile Terziovski, Amrik Sohal, and Andrea Howell. Best practice in product innovation at Varian Australia. Technovation, 22(9):561–569, September 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000608.

Thomson:2023:CBM

[TSPJ23] Linus Thomson, David Sjödin, Vinit Parida, and Marin Jovanovic. Conceptualizing business model piloting: an experiential learning process for autonomous solutions. Technovation, 126(??):??, August 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001268.

Tsuji:2001:PDJ

[Tsu01] Youichirou S. Tsuji. Product development in the Japanese and US printer industries. Technovation, 21(5):325–332, May 2001. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497200000365.

Tsuji:2002:OBR

[Tsu02] Youichirou S. Tsuji. Organizational behavior in the R&D process based on patent analysis: Strategic R&D management in a Japanese electronics firm. Technovation, 22(7): 417–425, July 2002. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497201000323.

Teng:2006:CCS

[TTC06] Ching-I Teng, Fan-Chen Tseng, and David M. Chiang. Customer-capturing strategies: the way to replace existing technology characterized by network effects. Technovation, 26(12):1384–1389, December 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205001136.

Tuzi:2005:USG

[Tuz05] Fabrizio Tuzi. Useful science is good science: empirical evidence from the Italian National Research Council. Technovation, 25(5):505–512, May 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203001925.

Takayama:2002:MMN

[TW02] Makoto Takayama and Chihiro Watanabe. Myth of market needs and technology seeds as a source of product innovation — an analysis of pharmaceutical new product development in an anti-hypertensive product innovation. Technovation, 22(6):353–362, June 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000924.

Thakur-Wernz:2021:ARP

[TWBC21] Pooja Thakur-Wernz, Olga Bruyaka, and Farok Contractor. Antecedents and relative performance of sourcing choices for new product development projects. Technovation, ?? (??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497217306041.

Takayama:2002:RIS

[TWGB02a] M. Takayama, C. Watanabe, and C. Griffy-Brown. Remaining innovative without sacrificing stability: an analysis of strategies in the Japanese pharmaceutical industry that enable firms to overcome inertia resulting from successful market penetration of new product development. Technovation, 22(12):747–759, December 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000670.

Takayama:2002:ASC

[TWGB02b] Makoto Takayama, Chihiro Watanabe, and Charla Griffy-Brown. Alliance strategy as a competitive strategy for successively creative new product development: the proof of the co-evolution of creativity and efficiency in the Japanese pharmaceutical industry. Technovation, 22(10):607–614, October 2002. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201000621.

Udell:1982:ENP

[UB82] Gerald G. Udell and Kenneth G. Baker. Evaluating new product ideas … systematically. Technovation, 1(3):191– 202, February 1982. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497282900025.

Urbinati:2019:CCV

[UBCF19] Andrea Urbinati, Marcel Bogers, Vittorio Chiesa, and Federico Frattini. Creating and capturing value from Big Data: a multiple-case study analysis of provider companies. Technovation, 84–85(??):21–36, June/July 2019. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497218300361.

Uddin:2006:RDI

[Udd06] M. Kamal Uddin. The role of diffusion of innovations for incremental development in small enterprises. Technovation, 26(2):274–284, February 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000714.

Uyarra:2014:BIT

[UEGE^{+}14] Elvira Uyarra, Jakob Edler, Javier Garcia-Estevez, Luke Georghiou, and Jillian Yeow. Barriers to innovation through public procurement: a supplier perspective. Technovation, 34(10):631–645, October 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214000388.

Ukobitz:2022:RII

[UF22] Desirée Valeria Ukobitz and Rita Faullant. The relative impact of isomorphic pressures on the adoption of radical technology: Evidence from 3D printing. Technovation, 113(??): ??, May 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497221001991.

Ughanwa:1988:BWM

[Ugh88] Davidson Oyemeka Ughanwa. Better ways of managing design: the Queen’s award winners’ experience. Technovation, 7(4):377–399, October 1988. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497288900089.

Ulhoi:2005:SDE

[Ulh05] J. P. Ulhøi. The social dimensions of entrepreneurship. Technovation, 25(8):939–946, August 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204000306.

Ugwu:2000:OII

[UOIA00] L. O. Ugwu, T. O. Oyebisi, M. O. Ilori, and E. R. Adagunodo. Organisational impact of information technology on the banking and insurance sector in Nigeria. Technovation, 20(12):711–721, December 2000. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000134.

Un:2018:LRO

[UR18] C. Annique Un and Alicia Rodríguez. Learning from R&D outsourcing vs. learning by R&D outsourcing. Technovation, 72–73(??):24–33, April/May 2018. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://

www.sciencedirect.com/science/article/pii/S0166497217309057.

Upadhayay:2024:BTR

[URG^{+}24] Neha B. Upadhayay, Silvia Rocchetta, Shivam Gupta, Sachin Kamble, and Rebecca Stekelorum. Blazing the trail: the role of digital and green servitization on technological innovation. Technovation, 130(??):??, February 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722300233X.

Uzun:2001:TIA

[Uzu01] Ali Uzun. Technological innovation activities in Turkey: the case of manufacturing industry, 1995–1997. Technovation, 21(3):189–196, March 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016649720000033X.

deMacedo-Soares:1994:BNQa

[vAdMSC94a] T. Diana L. v. A. de Macedo-Soares and S. G. R. Chamone. The Brazilian National Quality Award: sharing some best practices of the winner in 1993. Technovation, 14(10):657– 678, December 1994. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497294901120.

deMacedo-Soares:1994:BNQb

[vAdMSC94b] T. Diana L. v. A. de Macedo-Soares and S. G. R. Chamone. The Brazilian national quality award: sharing some best practices of the winner in 1993. Technovation, 14(10):695– 696, December 1994. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497294901171.

deMacedo-Soares:1995:ETQ

[vAdMSL95] T. Diana L. v. A. de Macedo-Soares and Débora C. Lucas. Empowerment and total quality: Comparing research findings in the USA and Brazil. Technovation, 15(8):475– 495, October 1995. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729596587J.

VandenPanhuyzen:1985:TTT

[Van85] Wim Van den Panhuyzen. Technology transfer through technological attachés: the Belgian experience. Technovation, 3(1):69–72, February 1985. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497285900380.

VanWyk:2010:TAP

[Van10] Rias Johann Van Wyk. Technology assessment for portfolio managers. Technovation, 30(4):223–228, April 2010. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209000911.

Vavakova:2006:RIP

[Vav06] Blanka Vavakova. Reconceptualizing innovation policy. The case of France. Technovation, 26(4):444–462, April 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720400152X.

Varela:2005:NPD

[VB05] José Varela and Leandro Benito. New product development process in Spanish firms: typology, antecedents and technical/marketing activities. Technovation, 25(4):395– 405, April 2005. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203001548.

Vazquez-Bustelo:2006:AMI

[VBA06] Daniel Vázquez-Bustelo and Lucía Avella. Agile manufacturing: Industrial case studies in Spain. Technovation, 26 (10):1147–1161, October 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497205001768.

vanBurg:2021:WDA

[vBDK21] Elco van Burg, Jingshu Du, and Jannigje Gerdien Kers. When do academics patent outside their university? An in-depth case study. Technovation, 107(??):??, September 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497221000687.

Verspagen:2004:SWS

[VD04] Bart Verspagen and Geert Duysters. The small worlds of strategic technology alliances. Technovation, 24(7):563–571, July 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497202001232.

vandenBrink:2002:TSR

[vdBS02] J. van den Brink and A. Szirmai. The Tanzanian scrap recycling cycle. Technovation, 22(3):187–197, March 2002. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000882.

vandenEnde:2008:ISF

[vdEJG08] Jan van den Ende, Ferdinand Jaspers, and Donald Gerwin. Involvement of system firms in the development of complementary products: the influence of novelty. Technovation, 28(11):726–738, November 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720800031X.

vandeVrande:2009:OIS

[vdVdJVdR09] Vareska van de Vrande, Jeroen P. J. de Jong, Wim Vanhaverbeke, and Maurice de Rochemont. Open innovation in SMEs: Trends, motives and management challenges. Technovation, 29(6–7):423–437, June/July 2009.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208001314.

vandeVrande:2011:ROI

[vdVdM11] Vareska van de Vrande and Ard-Pieter de Man. A response to “Is open innovation a field of study or a communication barrier to theory development?”. Technovation, 31(4):185– 186, April 2011. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497211000125.  See

[Lic11].

vanDuren:1989:FPA

[vDvM89] A. J. van Duren and M. van Manen. Flexible production automation: a description and definition. Technovation, 9(5):389–399, August 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497289900126.

vanderValk:2009:CPD

[vdVMM09] Tessa van der Valk, Ellen H. M. Moors, and Marius T. H. Meeus. Conceptualizing patterns in the dynamics of emerging technologies: the case of biotechnology developments in the Netherlands. Technovation, 29(4):247–264, April 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208001223.

vanderWaal:2023:MBS

[vdWFvdB23] Mark B. van der Waal, Jelle J. Feddema, and Linda H. M. van de Burgwal. Mapping the broad societal impact of patents. Technovation, 128(??):??, December 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001876.

vonDelft:2021:BMP

[vDZ21] Stephan von Delft and Yang Zhao. Business models in process industries: Emerging trends and future research. Technovation, ??(??):??, ???? 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497220300675.

Vedin:1981:MIC

[Ved81] Bengt-Arne Vedin. Media industry creativity management applied to technology. Technovation, 1(2):147–155, August 1981. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649728190016X.

Vedovello:1997:SPU

[Ved97] Conceição Vedovello. Science parks and university–industry interaction: Geographical proximity between the agents as a driving force. Technovation, 17(9):491–502, 530–531, September 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497297000278.

Vekstein:1999:DCT

[Vek99] Daniel Vekstein. Defense conversion, technology policy and R&D networks in the innovation system of Israel. Technovation, 19(10):615–629, October 1999. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299000668.

Velu:2015:BMI

[Vel15] Chander Velu. Business model innovation and third-party alliance on the survival of new firms. Technovation, 35(??): 1–11, January 2015. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497214001291.

Venanzi:1996:NAI

[Ven96] Daniela Venanzi. A new approach to the identification and selection of international technology transfer modes: logical framework and empirical evidence. Technovation, 16(6): 287–300, June 1996. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497296000235.

Venter:2014:SIS

[Ven14] Hein S. Venter. Security issues in the security cyber supply chain in South Africa. Technovation, 34(7):392–393, July 2014. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497214000236.

Vermaak:1990:BRH

[Ver90a] Leon Vermaak. Book review: High technology Europe: Strategic issues for global competitiveness: By Philippe De Woot, Basil Blackwell, Oxford, 1990, Hardback £35.00, 233 pages, ISBN 0-631-17104-5. Technovation, 10(6):441–442, September 1990. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729090010H.

Vermaak:1990:BRM

[Ver90b] Leon Vermaak. Book review: Managing across borders: the transnational solution: by Christopher A. Bartlett and Sumantra Ghoshal, Harvard Business Books, 1989, Hardback, 265 pages, ISBN 0-09174-1181. Technovation, 10(4): 282–284, June 1990. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729090060W.

Vermaak:1991:BRH

[Ver91a] Leon Vermaak. Book review: High technology Europe: strategic issues for global competitiveness: Philippe de Wool, Basil Blackwell, Oxford, 1990, ISBN 0-631-17104-5, £35.00, 233 pages. Technovation, 11(2):129, March 1991. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497291900515.

Vermaak:1991:BRM

[Ver91b] Leon Vermaak. Book review: Managing technology in the decentralized firm: by A. H. Rubenstein, Wiley, Chichester, 1989, ISBN 0-471-61024-0, £44.80, 476 pages. Technovation, 11(2):125–126, March 1991. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497291900478.

Vertova:2001:NTS

[Ver01] G. Vertova. National technological specialisation and the highest technological opportunities historically. Technovation, 21(9):605–612, September 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0166497201000347.

Versluis:2005:ITI

[Ver05] Cokki Versluis. Innovations on thin ice. Technovation, 25 (10):1183–1192, October 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497204000720.

Vasileiou:2024:RIP

[VGAL24] Efi Vasileiou, Nikolaos Georgantzis, Giuseppe Attanasi, and Patrick Llerena. The role of innovation portfolio in green innovation decisions: a study of French and Italian firms. Technovation, 130(??):??, February 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223002328.

Vega-Gonzalez:2010:TVS

[VGQK^{+}10] L. R. Vega-González, N. Qureshi, O. V. Kolokoltsev, R. Ortega-Martínez, and J. M. Saniger Blesa. Technology valuation of a scanning probe microscope developed at a university in a developing country. Technovation, 30(9–10):533–539, September/October 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210000659.

vanGeenhuizen:2009:ASO

[vGS09] Marina van Geenhuizen and Danny P. Soetanto. Academic spin-offs at different ages: a case study in search of key obstacles to growth. Technovation, 29(10):671–681, October 2009. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497209000820.

Vuola:2006:MBJ

[VH06] Olli Vuola and Ari-Pekka Hameri. Mutually benefiting joint innovation process between industry and bigscience. Technovation, 26(1):3–12, January 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000659.

vonHippel:2010:COI

[vH10] Eric von Hippel. Comment on ‘is open innovation a field of study or a communication barrier to theory development?’. Technovation, 30(11–12):555, November/December 2010. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210000933.

Vitali:1990:DIT

[Vit90] Giampaolo Vitali. Determinants of international trade of machine tools during the eighties: a shift-share analysis. Technovation, 10(8):507–519, December 1990. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729090047N.

Veer:2012:CMT

[VJ12] Theresa Veer and Florian Jell. Contributing to markets for technology? A comparison of patent filing motives of individual inventors, small companies and universities. Technovation, 32(9–10):513–522, September/October 2012. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497212000314.

Varshney:2023:TAF

[VJ23a] Mayank Varshney and Amit Jain. Technology acquisition following inventor exit in the biopharmaceutical industry. Technovation, 126(??):??, August 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001335.

Varshney:2023:URK

[VJ23b] Mayank Varshney and Amit Jain. Understanding “reverse” knowledge flows following inventor exit in the semiconductor industry. Technovation, 121(??):??, March 2023.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222001882.

Vilkamo:2003:STP

[VK03] T. Vilkamo and T. Keil. Strategic technology partnering in high-velocity environments — lessons from a case study. Technovation, 23(3):193–204, March 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201001080.

vonKrogh:2011:OIF

[vK11] Georg von Krogh. Is open innovation a field of study or a communication barrier to theory development? A commentary. Technovation, 31(7):286, July 2011. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211000514.

Voudouris:2012:ETI

[VLIC12] Irini Voudouris, Spyros Lioukas, Maria Iatrelli, and Yannis Caloghirou. Effectiveness of technology investment: Impact of internal technological capability, networking and investment’s strategic importance. Technovation, 32(6):400–414, June 2012. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497212000442.

Vaillant:2023:AIP

[VLVH23] Yancy Vaillant, Esteban Lafuente, and Ferran Vendrell-Herrero. Assessment of industrial pre-determinants for territories with active product-service innovation ecosystems. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497222002097.

Vrakking:1992:IWC

[VM92] W. J. Vrakking and P. Mulders. The implementation of ‘world class manufacturing’ principles in smaller industrial companies: a case study from consulting practice. Technovation, 12(5):297–308, July 1992. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729290038J.

Vesper:1997:NVS

[VM97] Karl Vesper and Ed McMullan. New venture scholarship versus practice: When entrepreneurship academics try the real thing as applied research. Technovation, 17(7):349– 358, July 1997. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497297000047.

Verdegem:2011:RDI

[VM11] Pieter Verdegem and Lieven De Marez. Rethinking determinants of ICT acceptance: Towards an integrated and comprehensive overview. Technovation, 31(8):411–423, August 2011. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497211000393.

Vanderstraeten:2012:SBD

[VM12] Johanna Vanderstraeten and Paul Matthyssens. Servicebased differentiation strategies for business incubators: Exploring external and internal alignment. Technovation, 32 (12):656–670, December 2012. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497212000971.

VanLancker:2016:OIS

[VMWV16] Jonas Van Lancker, Koen Mondelaers, Erwin Wauters, and Guido Van Huylenbroeck. The Organizational Innovation System: a systemic framework for radical innovation at the organizational level. Technovation, 52–53(??):40–50, June/ July 2016. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497215000826.

Visnjic:2018:POD

[VNJ18] Ivanka Visnjic, Andy Neely, and Marin Jovanovic. The path to outcome delivery: Interplay of service market strategy and open business models. Technovation, 72–73 (??):46–59, April/May 2018. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497216303406.

Vittori:2022:BMI

[VNP^{+}22] Davide Vittori, Angelo Natalicchio, Umberto Panniello, Antonio Messeni Petruzzelli, and Francesco Cupertino. Business Model Innovation between the embryonic and growth stages of industry lifecycle. Technovation, 117 (??):??, September 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222001390.

Vittori:2024:FOH

[VNP^{+}24] Davide Vittori, Angelo Natalicchio, Umberto Panniello, Antonio Messeni Petruzzelli, Vito Albino, and Francesco Cupertino. Failure is an option: How failure can lead to disruptive innovations. Technovation, 129(??):??, January 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223002080.

Voss:1984:MII

[Vos84] C. A. Voss. Multiple independent invention and the process of technological innovation. Technovation, 2(3):169– 184, June 1984. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497284900026.

Vos:2005:DSS

[Vos05] Jan-Peter Vos. Developing strategic self-descriptions of SMEs. Technovation, 25(9):989–999, September 2005. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204000410.

Visintin:2005:ISB

[VOTH05] Francesca Visintin, Burak Ozgen, Andrew Tylecote, and Robert Handscombe. Italian success and British survival: case studies of corporate governance and innovation in a mature industry. Technovation, 25(6):621–629, June 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203001974.

Visintin:2014:FTC

[VP14] Francesca Visintin and Daniel Pittino. Founding team composition and early performance of university-Based spinoff companies. Technovation, 34(1):31–43, January 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213001090.

vanRijnsoever:2021:MDM

[vRE21] Frank J. van Rijnsoever and Chris P. Eveleens. Money don’t matter? How incubation experience affects start-up entrepreneurs’ resource valuation. Technovation, 106(??): ??, August 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497221000754.

vonRaesfeld:2012:IPD

[vRGJ^{+}12] Ariane von Raesfeld, Peter Geurts, Mark Jansen, Johannes Boshuizen, and Regina Luttge. Influence of partner diversity on collaborative public R&D project outcomes: a study of application and commercialization of nanotechnologies in the Netherlands. Technovation, 32(3–4):227–233, March/ April 2012. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497211001684.

Vujanovic:2022:FSE

[VRS^{+}22] Nina Vujanović, Slavo Radosević, Nebojsa Stojcić, Mehtap Hisarciklilar, and Iraj Hashi. FDI spillover effects on innovation activities of knowledge using and knowledge creating firms: Evidence from an emerging economy. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222000591.

Vonortas:2006:MLR

[VS06] Nicholas S. Vonortas and Richard N. Spivack. Managing large research partnerships: Examples from the Advanced Technology Program’s Information Infrastructure for Healthcare program. Technovation, 26(10):1101–1110, October 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205001306.

Viardot:2016:MIS

[VSC16] Eric Viardot, Mostafa Hashem Sherif, and Jin Chen. Managing innovation with standardization: an introduction to recent trends and new challenges. Technovation, 48–49(??):1–3, February/March 2016. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497216000079.

Vasi:2024:ECS

[VSE24] Melissa Vasi, Giuliano Sansone, and Vincent English. Exogenous crises and SMEs resilience: the Dynamic Open Innovation Funnel. Technovation, 129(??):??, January 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001979.

Verdu:2012:MEE

[VTRM12] Antonio J. Verdu, Ignacio Tamayo, and Antonia Ruiz-Moreno. The moderating effect of environmental uncertainty on the relationship between real options and technological innovation in high-tech firms. Technovation, 32(9–10):579–590, September/October 2012. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497212000715.

Vanderstraeten:2020:OSS

[VvWM20] Johanna Vanderstraeten, Arjen van Witteloostuijn, and Paul Matthyssens. Organizational sponsorship and service co-development: a contingency view on service co-development directiveness of business incubators. Technovation, 98(??):??, December 2020. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497218304826.

vanWyk:1984:PST

[vW84] Rias J. van Wyk. Panoramic scanning and the technological environment. Technovation, 2(2):101–120, May 1984. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497284900154.

vanWyk:1988:MTN

[vW88] Rias J. van Wyk. Management of technology: New frameworks. Technovation, 7(4):341–351, October 1988. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900065.

vanWetten:2021:GIS

[vWGdG21] Sanne J. L. van Wetten, Ruud Gerards, and Andries de Grip. Are graduates’ intrapreneurial skills optimally used for innovation? Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497218303985.

vanWeele:2017:YCA

[vWvRN17] Marijn van Weele, Frank J. van Rijnsoever, and Frans Nauta. You can’t always get what you want: How entrepreneur’s perceived resource needs affect the incubator’s assertiveness. Technovation, 59(??):18–33, January 2017. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497216302498.

Vonortas:1997:PIS

[VX97] Nicholas S. Vonortas and Lan Xue. Process innovation in small firms: Case studies on CNC machine tools. Technovation, 17(8):427–438, August 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297000163.

vanZee:2014:ROV

[vZS14] Roger D. van Zee and Stefan Spinler. Real option valuation of public sector R&D investments with a down-and-out barrier option. Technovation, 34(8):477–484, August 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213000679.

Watanabe:2003:DIB

[WA03] Chihiro Watanabe and Behrooz Asgari. Dynamic interactions between assimilation capacity, technology spillovers, sales and R&D intensity — the case of electrical machinery industry in Japan. Technovation, 23(1):15–34, January 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201000876.

Watanabe:2004:CVC

[WA04a] Chihiro Watanabe and Bernadetta Kwintiana Ane. Constructing a virtuous cycle of manufacturing agility: concurrent roles of modularity in improving agility and reducing lead time. Technovation, 24(7):573–583, July 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202001189.

Watanabe:2004:IFD

[WA04b] Chihiro Watanabe and Behrooz Asgari. Impacts of functionality development on dynamism between learning and diffusion of technology. Technovation, 24(8):651–664, August 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497202001268.

Woodson:2019:PII

[WAdN19] Thomas Woodson, Julia Torres Alcantara, and Milena Silva do Nascimento. Is 3D printing an inclusive innovation?: an examination of 3D printing in Brazil. Technovation, 80–81(??):54–62, February/March 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497217302420.

Wallace:1985:DAE

[Wal85] William Wallace. Democracy and adaptation to economic and technological change: Is the British political system capable of adjusting? Technovation, 3(2):133–138, June 1985. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497285900069.

Wallmark:1997:IPU

[Wal97] J. Torkel Wallmark. Inventions and patents at universities: the case of Chalmers University of Technology. Technovation, 17(3):127–139, March 1997. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297000941.

Walwyn:2007:FMP

[Wal07] David Walwyn. Finland and the mobile phone industry: a case study of the return on investment from governmentfunded research and development. Technovation, 27(6–7): 335–341, June/July 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497207000260.

Walsh:2012:INI

[Wal12] Philip R. Walsh. Innovation Nirvana or Innovation Wasteland? Identifying commercialization strategies for small and medium renewable energy enterprises. Technovation, 32(1):32–42, January 2012. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497211001271.

Wan:1989:FHT

[Wan89] Victor Wan. Financing high technology: the Australian venture capital market. Technovation, 9(4):337–355, July 1989. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497287900058.

Wan:1991:AVC

[Wan91] Victor Wan. Australian venture capital market revisited. Technovation, 11(6):327–337, September 1991. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729190016W.

Watanabe:2003:VCB

[WAN03] C. Watanabe, B. Asgari, and A. Nagamatsu. Virtuous cycle between R&D, functionality development and assimilation capacity for competitive strategy in Japan’s hightechnology industry. Technovation, 23(11):879–900, November 2003. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649720200041X.

Wang:2004:BRC

[Wan04] Qing Wang. Book review: China’s leap into the information age: innovation and organisation in the computer industry: Qiwen, Oxford University Press, Oxford, ISBN 0-19-8295375. Technovation, 24(2):173–174, February 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203001482.

Wang:2023:OSC

[Wan23a] Tao Wang. The ownership structure of corporate venture capital financing and innovation. Technovation, 123(??):??, May 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223000470.

Wang:2023:TUI

[Wan23b] Tao Wang. Toward an understanding of innovation failure: the timing of failure experience. Technovation, 125(??):??, July 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223000986.

Wang:2024:DRD

[Wan24] Fang Wang. Does the recombination of distant scientific knowledge generate valuable inventions? An analysis of pharmaceutical patents. Technovation, 130(??):??, February 2024. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223002584.

Watkins:1990:BGB

[Wat90] Todd A. Watkins. Beyond guns and butter: managing dualuse technologies. Technovation, 10(6):389–406, September 1990. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497290900033.

Wittfoth:2022:RID

[WBM22] Sven Wittfoth, Theo Berger, and Martin G. Moehrle. Revisiting the innovation dynamics theory: How effectivenessand efficiency-oriented process innovations accompany product innovations. Technovation, 112(??):??, April 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001917.

Woods:1985:PTM

[WC85] Mike Woods and Steve Carter. The perception of teaching methods and the retraining of mature managers. Technovation, 3(2):77–87, June 1985. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497285900021.

Wu:2004:NFO

[WC04] Chih Wei Wu and Chyong Ling Chen. A new focus on overcoming the improvement failure. Technovation, 24(7): 585–591, July 2004. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203002001.

Wang:2006:FIP

[WC06a] Tai-Yue Wang and Shih-Chien Chien. Forecasting innovation performance via neural networks — a case of Taiwanese manufacturing industry. Technovation, 26(5–6):635–643, May/June 2006. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497204002135.

Wu:2006:ISM

[WC06b] Chih Wei Wu and Chyong Ling Chen. An integrated structural model toward successful continuous improvement activity. Technovation, 26(5–6):697–707, May/June 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205000805.

Wang:2007:ITD

[WC07] Tai-Yue Wang and Shih-Chien Chien. The influences of technology development on economic performance — the example of ASEAN countries. Technovation, 27(8):471–488, August 2007. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497207000570.

Wang:2010:MRI

[WCT10] Chun-Hsien Wang, Yang-Chieh Chin, and Gwo-Hshiung Tzeng. Mining the R&D innovation performance processes for high-tech firms based on rough set theory. Technovation, 30(7–8):447–458, July/August 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209001606.

Wang:2014:ERN

[WCW^{+}14] Fangrui Wang, Jin Chen, Yuandi Wang, Ning Lutao, and Wim Vanhaverbeke. The effect of R&D novelty and openness decision on firms’ catch-up performance: Empirical evidence from China. Technovation, 34(1):21–30, January 2014. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497213001107.

Wang:2014:MBM

[WCZZ14] Zhiqiang Wang, Lingfeng Chen, Xiande Zhao, and Wenhui Zhou. Modularity in building mass customization capability: the mediating effects of customization knowledge utilization and business process improvement. Technovation, 34(11):678–687, November 2014. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214001242.

Wallin:2006:SSO

[WD06] Martin W. Wallin and Åsa Lindholm Dahlstrand. Sponsored spin-offs, industrial growth and change. Technovation, 26(5–6):611–620, May/June 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0166497205000441.

Webb:1996:PDP

[Web96] G. Kent Webb. Product development and price trends for fiber optic components. Technovation, 16(3):133–141, March 1996. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497295000046.

Webster:1999:TTP

[Web99] Andrew Webster. Technologies in transition, policies in transition: foresight in the risk society. Technovation, 19 (6–7):413–421, June 1999. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497299000310.

Weckowska:2015:LUT

[Wec15] Dagmara M. Weckowska. Learning in university technology transfer offices: transactions-focused and relationsfocused approaches to commercialization of academic research. Technovation, 41–42(??):62–74, July/August 2015. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214001710.

Werker:2003:IMP

[Wer03] C. Werker. Innovation, market performance, and competition: lessons from a product life cycle model. Technovation, 23(4):281–290, April 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497201001092.

Wessel:1993:TTIa

[Wes93a] V. William Wessel. Technology transfer and intellectual property: an analysis of the NASA approach. Technovation, 13(3):133–146, April 1993. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729390036U.

Wessel:1993:TTIb

[Wes93b] V. William Wessel. Technology transfer and intellectual property: an analysis of the NASA approach. Technovation, 13(3):178–180, April 1993. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497293900403.

Wetzel:1981:TII

[Wet81] William E. Wetzel, Jr. Technovation and the informal investor. Technovation, 1(1):15–30, February 1981. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497281900043.

Wu:2010:ECC

[WF10] Sou-Chin Wu and WenChang Fang. The effect of consumer-to-consumer interactions on idea generation in virtual brand community relationships. Technovation, 30(11–12):570–581, November/December 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497210000878.

Wognum:2002:BRM

[WFW02] P. M. Wognum, Olaf A. M. Fisscher, and Suzanne A. J. Weenink. Balanced relationships: management of clientsupplier relationships in product development. Technovation, 22(6):341–351, June 2002. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497201000311.

Wang:2021:IDI

[WFWC21] Quan-Jing Wang, Gen-Fu Feng, Hai-Jie Wang, and Chun-Ping Chang. The impacts of democracy on innovation: Revisited evidence. Technovation, 108(??):??, December 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001140.

Walrave:2023:GSG

[WG23] Bob Walrave and Victor A. Gilsing. Game of skill or game of luck? Distant search in response to performance feedback. Technovation, 121(??):??, March 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222002322.

Woolgar:1998:ACRb

[WGV^{+}98] Steve Woolgar, Paula Gomes, Janet Vaux, Jean-Noel Ezingeard, and Robert Grieve. Abilities and competencies required, particularly by small firms, to identify and acquire new technology. Technovation, 18(8–9):591–592, August/ September 1998. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497298901286.

Wang:2022:DES

[WGW22] Yanping Wang, Jie Gao, and Zelong Wei. The double-edged sword of servitization in radical product innovation: the role of latent needs identification. Technovation, 118(??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497221000651.

Woiceshyn:1996:SPC

[WH96] Jaana Woiceshyn and Diana Hartel. Strategies and performance of Canadian biotechnology firms: an empirical investigation. Technovation, 16(5):231–243, 265–266, May 1996. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497295000666.

Watanabe:2004:CFS

[WH04a] C. Watanabe and M. Hobo. Creating a firm self-propagating function for advanced innovation-oriented projects: lessons from ERP. Technovation, 24(6):467–481, June 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202001037.

Watanabe:2004:CEB

[WH04b] Chihiro Watanabe and Masayo Hobo. Co-evolution between internal motivation and external expectation as a source of firm self-propagating function creation. Technovation, 24 (2):109–120, February 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497202000433.

Watanabe:2004:RRS

[WH04c] Chihiro Watanabe and Jae Yong Hur. Resonant R&D structure for effective technology development amidst megacompetition — an empirical analysis of smart cooperative R&D structure in Japan’s transport machinery industry. Technovation, 24(12):955–969, December 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203000506.

Wallberg-Henriksson:2013:GH

[WH13] Harriet Wallberg-Henriksson. Global health. Technovation, 33(4–5):106, April/May 2013. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497213000291.

Wang:2021:KSS

[WH21] Changfeng Wang and Qiying Hu. Knowledge sharing in supply chain networks: Effects of collaborative innovation activities and capability on innovation performance. Technovation, ??(??):??, ???? 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497217308994.

Woodson:2021:ENB

[WHB21] Thomas S. Woodson, Elina Hoffmann, and Sophia Boutilier. Evaluating the NSF broader impacts with the Inclusion-Immediacy Criterion: a retrospective analysis of nanotechnology grants. Technovation, 101(??):??, March 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497220300821.

Wheatley:1986:CPC

[Whe86] Jeffery J. Wheatley. Competition, privatization and change at British telecom. Technovation, 5(1–3):115–124, October 1986. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497286900477.

Wheatley:1988:PMT

[Whe88] Jeffery J. Wheatley. Performance measures in telecommunications: Changing needs. Technovation, 8(1–3):201–216, ???? 1988. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497288900612.

Whiston:1989:MOI

[Whi89a] Thomas G. Whiston. Managerial and organizational integration needs arising out of technical change and U.K. commercial structures. Part I: Introduction and overview. Technovation, 9(7):577–605, November 1989. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497289900229.

Whitman:1989:MTD

[Whi89b] Charles I. Whitman. Managing technology development: perspectives of a practitioner. Technovation, 9(7):551–560, November 1989. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166497289900205.

Whiston:1990:MOIa

[Whi90a] Thomas G. Whiston. Managerial and organizational integration needs arising out of technical change and U.K. commercial structures. Part II: a managerial and organizational perspective. Technovation, 10(1):47–58, February 1990. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729090042I.

Whiston:1990:MOIb

[Whi90b] Thomas G. Whiston. Managerial and organizational integration needs arising out of technical change and U.K. commercial structures. Part III: Managerial and organizational integration within the organic enterprise. Technovation, 10(2):95–118, March 1990. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729090031E.

Whiston:1990:MOIc

[Whi90c] Thomas G. Whiston. Managerial and organizational integration needs arising out of technical change and U.K. commercial structures. Part IV: Remaining policy issues and organizational questions. Technovation, 10(3):143–161, May 1990. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016649729090021B.

Whiston:1994:BRB

[Whi94] T. G. Whiston. Book review: Being an effective expert witness: the technologist in the courtroom, Derek Smith. Thames Publishing, London (1993). Technovation, 14 (6):424–425, August 1994. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497294900264.

White-Hunt:1985:DSF

[WHP85] K. White-Hunt and A. Pomykalski. A description of some factors affecting technological innovation in industry in the United Kingdom and some remedies to improve the climate. Technovation, 3(2):139–151, June 1985. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497285900070.

Wei:2022:WDS

[WHWS22] Zelong Wei, Wengao Huang, Yanping Wang, and Lulu Sun. When does servitization promote product innovation? The moderating roles of product modularization and organization formalization. Technovation, 117(??):??, September 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222001419.

Whyte:2002:BRD

[Why02] J. Whyte. Book review: Design in business: strategic innovation design: Margaret Bruce and John Bessant; Pearson Education Limited, Harlow £29.99, 277 pages. Technovation, 22(12):807–808, December 2002. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000810.

Whyte:2005:BRD

[Why05] Jennifer Whyte. Book review: Design Management: Using Design to Build Brand Value and Corporate Innovation: Brigitte Borja de Mozota. Design Management Institute and Allworth Press, New York (2003). Technovation, 25(3):305–306, March 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497204002068.

Williams:1983:EIE

[Wil83] Edward E. Williams. Entrepreneurship, innovation and economic growth. Technovation, 2(1):3–15, February 1983.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497283900032.

Willcocks:1990:BRI

[Wil90] Leslie Willcocks. Book review: Innovation in technology and organization: by Peter Clark and Neil Staunton, Routledge, London, 1989, Hardback £35.00, ISBN 0415004225, 242 pages. Technovation, 10(5):355–357, July 1990. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729090019G.

Wilson:1993:IAV

[Wil93] Heather I. M. Wilson. An interregional analysis of venture capital and technology funding in the UK. Technovation, 13(7):425–438, November 1993. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390026R. See [Ano93j].

Williams:2014:SCS

[Wil14] Colin Williams. Security in the cyber supply chain: Is it achievable in a complex, interconnected world? Technovation, 34(7):382–384, July 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497214000212.

Winiecki:1987:STE

[Win87] Jan Winiecki. Soviet-type economies’ strategy for catchingup through technology imports — an anatomy of failure. Technovation, 6(2):115–145, June 1987. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497287900150.

Wittingslow:1985:RAE

[Wit85] G. E. Wittingslow. A report on an Australian experiment at attempting to increase innovation and entrepreneurship within the community: Results of efforts over four years. Technovation, 3(1):39–47, February 1985. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497285900355.

Woo:2015:EIP

[WJK15] Seokkyun Woo, Pilseong Jang, and Yeonbae Kim. Effects of intellectual property rights and patented knowledge in innovation and industry value added: a multinational empirical analysis of different industries. Technovation, 43–44(??):49–63, September/October 2015. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497215000279.

waKabecha:1997:CJQ

[wK97] Wanjau wa Kabecha. Consumer judgement of the quality of informal sector products lessons for innovative microentrepreneurs. Technovation, 17(2):91–100, 105, February 1997. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297841922.

waKabecha:1998:TCM

[wK98] Wanjau wa Kabecha. Technological capability of the microenterprises in Kenya’s informal sector. Technovation, 19 (2):117–126, December 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497298000674.

Watanabe:2003:IET

[WK03] Chihiro Watanabe and Reiko Kondo. Institutional elasticity towards IT waves for Japan’s survival — the significant role of an IT testbed. Technovation, 23(4):307–320, April 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201000736.

Wolf:2024:CFF

[WKF24] Patricia Wolf, Ute Klotz, and Marianne Harbo Frederiksen. Consumer flash fiction: a methodology to support the early sensing of far-future innovation opportunities. Technovation, 133(??):??, May 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000646.

Watanabe:2003:POD

[WKN03] C. Watanabe, R. Kondo, and A. Nagamatsu. Policy options for the diffusion orbit of competitive innovations — an application of Lotka–Volterra equations to Japan’s transition from analog to digital TV broadcasting. Technovation, 23(5):437–445, May 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000044.

Watanabe:2004:ELG

[WKN04a] C. Watanabe, M. Kishioka, and A. Nagamatsu. Effect and limit of the government role in spurring technology spillover — a case of R&D consortia by the Japanese government. Technovation, 24(5):403–420, May 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000755.

Watanabe:2004:RSS

[WKN04b] Chihiro Watanabe, Miharu Kishioka, and Akira Nagamatsu. Resilience as a source of survival stategy for high-technology firms experiencing megacompetition. Technovation, 24(2):139–152, February 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000482.

Watanabe:2003:FIF

[WKOW03] Chihiro Watanabe, Reiko Kondo, Noritomo Ouchi, and Haihong Wei. Formation of IT features through interaction with institutional systems-empirical evidence of unique epidemic behavior. Technovation, 23(3):205–219, March 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201001067.

Willcocks:1991:ISI

[WL91] Leslie Willcocks and Stephanie Lester. Information systems investments: evaluation at the feasibility stage of projects. Technovation, 11(5):283–302, July 1991. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497291900272.

Wang:2021:UII

[WL21] Wenjing Wang and Shan Lu. University–industry innovation community dynamics and knowledge transfer: Evidence from China. Technovation, 106(??):??, August 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000869.

Woolthuis:2005:SFF

[WLG05] Rosalinde Klein Woolthuis, Maureen Lankhuizen, and Victor Gilsing. A system failure framework for innovation policy design. Technovation, 25(6):609–619, June 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203002037.

Wang:2010:POR

[WLH10] Juite Wang, Willie Lin, and Yu-Hsiang Huang. A performance-oriented risk management framework for innovative R&D projects. Technovation, 30(11–12):601–611, November/December 2010. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497210000854.

Watanabe:2009:FIT

[WLO09] Chihiro Watanabe, Shanyu Lei, and Noritomo Ouchi. Fusing indigenous technology development and market learning for greater functionality development — an empirical analysis of the growth trajectory of Canon printers. Technovation, 29(4):265–283, April 2009. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208001375.

Wang:2023:GDI

[WLTH23] Xincheng Wang, Yuan Li, Longwei Tian, and Ye Hou. Government digital initiatives and firm digital innovation: Evidence from China. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016649722200092X.

Wang:2014:WDI

[WLY14] Yuandi Wang and Jason Li-Ying. When does inward technology licensing facilitate firms’ NPD performance? A contingency perspective. Technovation, 34(1):44–53, January 2014. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497213001077.

Weinberg:1988:GUI

[WM88] Mark L. Weinberg and Mary Ellen Mazey. Government– university–industry partnerships in technology development: a case study. Technovation, 7(2):131–142, May 1988. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497288900430.

Watanabe:2001:DDT

[WMGB01] Chihiro Watanabe, Kiyofumi Matsumoto, and Charla Griffy-Brown. Development and diffusion trajectory of innovative products in the light of institutional maturity — a comparative empirical analysis of the laser beam printer and optical cards. Technovation, 21(10):637–647, October 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000791.

Whittaker:2023:MDL

[WML^{+}23] Lucas Whittaker, Rory Mulcahy, Kate Letheren, Jan Kietzmann, and Rebekah Russell-Bennett. Mapping the deepfake landscape for innovation: a multidisciplinary systematic review and future research agenda. Technovation, 125 (??):??, July 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223000950.

Watanabe:2003:SSS

[WN03] C. Watanabe and A. Nagamatsu. Sources of structural stagnation in R&D intensity in Japan’s electrical machinery industry — an analysis of mismatching with IT functionality development. Technovation, 23(7):571–591, July 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000081.

Watanabe:2003:BTR

[WNGB03] C. Watanabe, A. Nagamatsu, and C. Griffy-Brown. Behavior of technology in reducing prices of innovative goods

— an analysis of the governing factors of variance of PV module prices. Technovation, 23(5):423–436, May 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497201001213.

Wu:2006:CGR

[WNLY06] N. C. Wu, M. A. Nystrom, T. R. Lin, and H. C. Yu. Challenges to global RFID adoption. Technovation, 26(12): 1317–1323, December 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016649720500146X.

Walsh:1995:SFF

[WNM95] Vivien Walsh, Jorge Niosi, and Philippe Mustar. Small-firm formation in biotechnology: a comparison of France, Britain and Canada. Technovation, 15(5):303–327, June 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729596602P.

Wu:2008:MIT

[WO08] Liang-Chuan Wu and Chorng-Shyong Ong. Management of information technology investment: a framework based on a Real Options and Mean-Variance theory perspective. Technovation, 28(3):122–134, March 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000715.

Wan:2005:DFI

[WOL05] David Wan, Chin Huat Ong, and Francis Lee. Determinants of firm innovation in Singapore. Technovation, 25(3):261– 268, March 2005. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497203000968.

Wonglimpiyarat:2004:ASL

[Won04] J. Wonglimpiyarat. Amex’s strategies for launching the smart card innovation. Technovation, 24(10):773–777, October 2004. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497202001815.

Wonglimpiyarat:2005:DCA

[Won05a] Jarunee Wonglimpiyarat. Does complexity affect the speed of innovation? Technovation, 25(8):865–882, August 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497204000094.

Wonglimpiyarat:2005:NRS

[Won05b] Jarunee Wonglimpiyarat. The nano-revolution of Schumpeter’s Kondratieff cycle. Technovation, 25(11):1349–1354, November 2005. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497204001087.

Wonglimpiyarat:2006:DEE

[Won06] Jarunee Wonglimpiyarat. The dynamic economic engine at Silicon Valley and US Government programmes in financing innovations. Technovation, 26(9):1081–1089, September 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205001215.

Wonglimpiyarat:2007:MGV

[Won07] Jarunee Wonglimpiyarat. Management and governance of venture capital: a challenge for commercial bank. Technovation, 27(12):721–731, December 2007. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000673.

Wonglimpiyarat:2008:BRH

[Won08] Jarunee Wonglimpiyarat. Book review: High-Tech Entrepreneurship in Asia-Innovation, Industry and Institutional Dynamics in Mobile Payments: Marina Yue Zhang, Mark Dodgson. Edward Elgar Publishing Ltd, Cheltenham, UK; Northampton, MA, USA. 352 pp., £79.95, ISBN 978-1-84720-056-3. Technovation, 28(6):391, June 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720700082X.

Wonglimpiyarat:2013:REF

[Won13] Jarunee Wonglimpiyarat. The role of equity financing to support entrepreneurship in Asia-The experience of Singapore and Thailand. Technovation, 33(4–5):163–171, April/ May 2013. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497213000023.

Wonglimpiyarat:2016:GPT

[Won16] Jarunee Wonglimpiyarat. Government policies towards Israel’s high-tech powerhouse. Technovation, 52–53(??):18– 27, June/July 2016. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497216300049.

Woods:1985:DLO

[Woo85] Mike Woods. Development, learning and organizational change. Technovation, 3(2):75–76, June 1985. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649728590001X.

Woodruff:1987:BTG

[Woo87] Alan Woodruff. British technology group — technology transfer on a broad front. Technovation, 6(3):151–153, August 1987. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0166497287900198.

Wang:2006:IIDa

[WPF06a] Hui-Chih Wang, John G. Pallister, and Gordon R. Foxall. Innovativeness and involvement as determinants of website loyalty: I. A test of the style/involvement model in the context of Internet buying. Technovation, 26(12):1357–1365, December 2006. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497205001732.

Wang:2006:IIDb

[WPF06b] Hui-Chih Wang, John G. Pallister, and Gordon R. Foxall. Innovativeness and involvement as determinants of website loyalty: II. Determinants of consumer loyalty in B2C e-commerce. Technovation, 26(12):1366–1373, December 2006. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497205001744.

Wang:2006:IIDc

[WPF06c] Hui-Chih Wang, John G. Pallister, and Gordon R. Foxall. Innovativeness and involvement as determinants of website loyalty: III. Theoretical and managerial contributions. Technovation, 26(12):1374–1383, December 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205001756.

Wareham:2022:SSB

[WPR^{+}22] Jonathan Wareham, Laia Pujol Priego, Angelo Kenneth Romasanta, Thomas Wareham Mathiassen, Markus Nordberg, and Pablo Garcia Tello. Systematizing serendipity for big science infrastructures: the AT-TRACT project. Technovation, 116(??):??, August 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497221001553.

Wen:2021:EEI

[WQZ21] Jinyan Wen, William J. Qualls, and Deming Zeng. To explore or exploit: the influence of inter-firm RD network diversity and structural holes on innovation outcomes. Technovation, 100(??):??, February 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649722030050X.

Wield:1995:RCS

[WR95] David Wield and Robin Roy. R&D and corporate strategies in UK materials-innovating companies. Technovation, 15(4):195–210, May 1995. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729596603Q.

Watanabe:2014:TST

[WS14] Chihiro Watanabe and M. Paz Salmador. Technology strategy and technology policy. Technovation, 34(12):731–733, December 2014. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497214001412.

Wanzenbock:2013:HDF

[WSF13] Iris Wanzenböck, Thomas Scherngell, and Manfred M. Fischer. How do firm characteristics affect behavioural additionalities of public R&D subsidies? Evidence for the Austrian transport sector. Technovation, 33(2–3):66–77, February/March 2013. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649721200140X.

Weidner:2023:ICF

[WSH23] Nadia Weidner, Oliver Som, and Djerdj Horvat. An integrated conceptual framework for analysing heterogeneous configurations of absorptive capacity in manufacturing firms with the DUI innovation mode. Technovation, 121 (??):??, March 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497222001857.

Wang:2018:EIP

[WSN^{+}18] Delu Wang, Dylan Sutherland, Lutao Ning, Yuandi Wang, and Xin Pan. Exploring the influence of political connections and managerial overconfidence on R&D intensity in China’s large-scale private sector firms. Technovation, 69(??):40–53, January 2018. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497217307812.

Watanabe:2019:TDR

[WT19] Chihiro Watanabe and Yuji Tou. Transformative direction of R&D-lessons from Amazon’s endeavor. Technovation, 88 (??):Article 102081, December 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://

www.sciencedirect.com/science/article/pii/S0166497219300409.

Wahla:1997:PRM

[WTAW97] N. Wahla, D. Tither, W. Ahmed, and D. Whitaker. Preliminary results from a market analysis of the two-stroke combustion engine industry. Technovation, 17(7):403–408, July 1997. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497297000059.

Seguin-Dulude:1991:AIC

[wtcoLSDD91] Denyse L. Dagenais with the collaboration of Louise Séguin-Dulude and Claude Desranleau. Analysis of individual Canadian inventor behavior. Technovation, 11(6):357–372, September 1991. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729190018Y.

Watanabe:2001:PSD

[WTGB01] Chihiro Watanabe, Youichirou S. Tsuji, and Charla Griffy-Brown. Patent statistics: deciphering a ‘real’ versus a ‘pseudo’ proxy of innovation. Technovation, 21(12):783– 790, December 2001. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497201000256.

Wan:2023:DDD

[WTJ23] Qunchao Wan, Shiyi Tang, and Zhenwu Jiang. Does the development of digital technology contribute to the innovation performance of China’s high-tech industry? Technovation, 124(??):??, June 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000494.

Wahla:1998:RIIb

[WTN98a] N. Wahla, D. Tither, and P. Naude. A research investigation to identify the key players in the 2-stroke engine industry. Technovation, 18(2):145–146, February 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298800126.

Wahla:1998:RIIa

[WTN98b] N. Wahla, D. Tither, and P. Naude. A research investigation to identify the key players in the two-stroke engine industry. Technovation, 18(2):133–140, February 1998. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497297000862.

Watanabe:2002:TSC

[WTN^{+}02] Chihiro Watanabe, Makoto Takayama, Akira Nagamatsu, Takashi Tagami, and Charla Griffy-Brown. Technology spillover as a complement for high-level R&D intensity in the pharmaceutical industry. Technovation, 22(4):245–258, April 2002. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497201000049.

Wurmseher:2017:EHO

[Wür17] Martin Würmseher. To each his own: Matching different entrepreneurial models to the academic scientist’s individual needs. Technovation, 59(??):1–17, January 2017. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497216303480.

Willemstein:2007:DBM

[WvdVM07] L. Willemstein, T. van der Valk, and M. T. H. Meeus. Dynamics in business models: an empirical analysis of medical biotechnology firms in the Netherlands. Technovation, 27(4):221–232, April 2007. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497206000964.

Walrave:2024:KDT

[WvdWG24] Bob Walrave, Nino van de Wal, and Victor Gilsing. Knowledge diversity and technological innovation: the moderating role of top management teams. Technovation, 131(??): ??, March 2024. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649722400004X.

Woolgar:1998:ACRa

[WVG^{+}98] Steve Woolgar, Janet Vaux, Paula Gomes, Jean-Noel Ezingeard, and Robert Grieve. Abilities and competencies required, particularly by small firms, to identify and acquire new technology. Technovation, 18 (8–9):575–584, 591–592, August/September 1998. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497298000492.

Williams:2023:UPP

[WvT23] Christopher Williams and Sander van Triest. Understanding performance in professional services for innovation intermediation: Technology consultants vs. management consultants. Technovation, 126(??):??, August 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001359.

Wang:2004:EIU

[WW04] Clement K. Wang and Poh-Kam Wong. Entrepreneurial interest of university students in Singapore. Technovation, 24(2):163–172, February 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497202000160.

Wu:2014:IRM

[WW14] Jie Wu and Zefu Wu. Integrated risk management and product innovation in China: the moderating role of board of directors. Technovation, 34(8):466–476, August 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213001442.

Wang:2021:CRE

[WW21] Jimin Wang and Cong Wang. Can religions explain cross country differences in innovative activities? Technovation, 107(??):??, September 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497221000663.

Wu:2015:CUI

[WWH15] Yonghong Wu, Eric W. Welch, and Wan-Ling Huang. Commercialization of university inventions: Individual and institutional factors affecting licensing of university patents. Technovation, 36–37(??):12–25, February/March 2015. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214001266.

Watanabe:2000:IDC

[WWM00] Chihiro Watanabe, Kouji Wakabayashi, and Toshinori Miyazawa. Industrial dynamism and the creation of a “virtuous cycle” between R&D, market growth and price reduction: the case of photovoltaic power generation (PV) development in Japan. Technovation, 20(6):299–312, June 2000. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299001467.

Wan:2015:AID

[WWY15] Feng Wan, Peter J. Williamson, and Eden Yin. Antecedents and implications of disruptive innovation: Evidence from China. Technovation, 39–40(??):94–104, May/June 2015. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214000820.

Wen:2023:IEI

[WWYZ23] Jun Wen, Siqin Wang, Xiuyun Yang, and Xiaozhou Zhou. Impacts of epidemics on innovation: an empirical analysis. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497222001845.

Wyatt:2000:BRV

[Wya00] Sally Wyatt. Book review: Virtual Geographies. Bodies, space and relations: Mike Crang, Phil Crang and Jon May (Eds); Routledge, London, 1999, pp. 322 (incl. index), £15.99, ISBN 0-415-16828-7 (paperback). Technovation, 20 (6):343, June 2000. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497299001832.

Wyatt:2004:BRI

[Wya04] Sally Wyatt. Book review: Internet governance in transition. Who is the master of this domain?: Daniel J Paré; Rowman & Littlefield, Lanham, Maryland, 2003, pp. 192 (incl. index), £20.95, ISBN 0-7425-1846-9 (paperback). Technovation, 24(2):175–176, February 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203001743.

Wyatt:2007:BRB

[Wya07] Sally Wyatt. Book review: B.-K. Kim, Internationalizing the Internet: the Co-evolution of Influence and Technology, Edward Elgar, Cheltenham, UK and Northampton, MA, USA, ISBN 1-84376-497-0, 2005 (300p. including references and index, £75.00, hardback only). Technovation, 27(5): 319, May 2007. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497206000277.

Watanabe:2001:GTS

[WZGBA01] Chihiro Watanabe, Bing Zhu, Charla Griffy-Brown, and Behrooz Asgari. Global technology spillover and its impact on industry’s R&D strategies. Technovation, 21(5): 281–291, May 2001. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497200000481.

Wang:2015:TEC

[WZNC15] Yuandi Wang, Zhao Zhou, Lutao Ning, and Jin Chen. Technology and external conditions at play: a study of learning-by-licensing practices in China. Technovation, 43–44(??):29–39, September/October 2015. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497215000309.

Winterhalter:2017:BMF

[WZNG17] Stephan Winterhalter, Marco B. Zeschky, Lukas Neumann, and Oliver Gassmann. Business models for frugal innovation in emerging markets: the case of the medical device and laboratory equipment industry. Technovation, 66–67(??):

3–13, August 2017. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016649721730531X.

Wang:2016:ESI

[WZSZ16] Zhiqiang Wang, Min Zhang, Hongyi Sun, and Guilong Zhu. Effects of standardization and innovation on mass customization: an empirical investigation. Technovation, 48–49(??):79–86, February/March 2016. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497216000043.

Wang:2023:OPL

[WZY23] Ling Wang, Yujia Zhang, and Yushan Yan. Offensive patent litigation strategic choice: an organizational routine perspective. Technovation, 122(??):??, April 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222002140.

Chen:2022:DNL

[xCZZ^{+}22] Jian xun Chen, Bo Zhang, Wu Zhan, Piyush Sharma, Pawan Budhwar, and Hui Tan. Demystifying the nonlinear effect of high commitment work systems (HCWS) on firms’ strategic intention of exploratory innovation: an extended resource-based view. Technovation, 116(??):??, August 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000463.

Xie:2016:SER

[XHM^{+}16] Zongjie Xie, Jeremy Hall, Ian P. McCarthy, Martin Skitmore, and Liyin Shen. Standardization efforts: the relationship between knowledge dimensions, search processes and innovation outcomes. Technovation, 48–49(??):69–78, February/March 2016. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497215000929.

Xie:2004:TLC

[Xie04] Wie Xie. Technological learning in China’s colour TV (CTV) industry. Technovation, 24(6):499–512, June 2004. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202000767.

Xing:2023:IPS

[XL23] Yijun Xing and Yipeng Liu. Integrating product-service innovation into green supply chain management from a life cycle perspective: a systematic review and future research directions. Technovation, 126(??):??, August 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001360.

Xie:2023:MAR

[XLC23] Xuemei Xie, Xiaojie Liu, and Jialing Chen. A metaanalysis of the relationship between collaborative innovation and innovation performance: the role of formal and informal institutions. Technovation, 124(??):??, June 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000512.

Xing:2023:SIS

[XLD23] Yijun Xing, Yipeng Liu, and Philip Davies. Servitization innovation: a systematic review, integrative framework, and future research directions. Technovation, 122(??):??, April 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497222001924.

Xue:2023:WWD

[XLG^{+}23] Jinjie Xue, Junqi Liu, Zizhen Geng, Hongping Yuan, and Lei Chao. Why and when do paradoxical management capabilities matter to paradoxical pressure? An empirical investigation of the role of coopetition. Technovation, 120(??): ??, February 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497222002334.

Xu:2021:MNE

[XPC21] Jin Xu, Biyu Peng, and Joep Cornelissen. Modelling the network economy: a population ecology perspective on network dynamics. Technovation, 102(??):??, April 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497220300845.

Xia:2008:CCA

[XR08] Tianjiao Xia and Stephen Roper. From capability to connectivity-Absorptive capacity and exploratory alliances in biopharmaceutical firms: a US-Europe comparison. Technovation, 28(11):776–785, November 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208000448.

Xia:2021:SIE

[XW21] Dan Xia and Xiang-Qian Wang. The synergetic impact of environmental and innovation information disclosure on corporate financial performance: an empirical study based on China coal listed companies. Technovation, 100(??):??, February 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497220300511.

Xie:2023:IOC

[XWZ23] Xuemei Xie, Lijun Wang, and Tao Zhang. Involving online community customers in product innovation: the doubleedged sword effect. Technovation, 123(??):??, May 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222002383.

Xiong:2021:MBC

[XX21] Yu Xiong and Senmao Xia. Mechanisms behind China’s innovation achievements: a multi-level view. Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497220300134.

Yildirim:2011:FFF

[YA11] Nihan Yildirim and Hacer Ansal. Foresighting FLOSS (free/libre/open source software) from a developing country perspective: the case of Turkey. Technovation, 31(12):666– 678, December 2011. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497211001052.

Yamashita:2021:ECI

[Yam21] Nobuaki Yamashita. Economic crisis and innovation capacity of Japan: Evidence from cross-country patent citations. Technovation, 101(??):??, March 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497220300808.

Yasuda:2005:FSA

[Yas05] Hiroshi Yasuda. Formation of strategic alliances in hightechnology industries: comparative study of the resourcebased theory and the transaction-cost theory. Technovation, 25(7):763–770, July 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497204000069.

Yang:2005:GIP

[YC05] Deli Yang and Peter Clarke. Globalisation and intellectual property in China. Technovation, 25(5):545–555, May 2005. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203001986.

Yi:2022:STO

[YCL22] Yaqun Yi, Yu Chen, and Dan Li. Stakeholder ties, organizational learning, and business model innovation: a business ecosystem perspective. Technovation, 114(??):??, June 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497221002261.

Yeh:2013:NEP

[Yeh13] Nai-Chang Yeh. Nanotechnology for electronics & photonics. Technovation, 33(4–5):108, April/May 2013. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649721300031X.

Yamin:1999:RBG

[YGM99] Shahid Yamin, A. Gunasekaran, and Felix T. Mavondo. Relationship between generic strategies, competitive advantage and organizational performance: an empirical analysis. Technovation, 19(8):507–518, August 1999. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497299000243.

Yusuf:2006:IER

[YGW06] Yahaya Yusuf, Angappa Gunasekaran, and Canglin Wu. Implementation of enterprise resource planning in China. Technovation, 26(12):1324–1336, December 2006. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497205001823. See erratum [YGW08].

Yusuf:2008:EIE

[YGW08] Yahaya Yusuf, Angappa Gunasekaran, and Canglin Wu. Erratum to “Implementation of enterprise resource planning in China” [Technovation 26 (2006) 1324– 1336]. Technovation, 28(12):864, December 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497208001211. See [YGW06].

Yang:2009:MDI

[YH09] Heng-Li Yang and Shu-Ling Hsiao. Mechanisms of developing innovative IT-enabled services: a case study of Taiwanese healthcare service. Technovation, 29(5):327–337, May 2009. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497209000042.

Yu:2011:CTC

[YH11] Dan Yu and Chang Chieh Hang. Creating technology candidates for disruptive innovation: Generally applicable R&D strategies. Technovation, 31(8):401–410, August 2011.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497211000411.

Yi:2017:RSO

[YHcHW17] Jingtao Yi, Junjie Hong, Wen chung Hsu, and Chengqi Wang. The role of state ownership and institutions in the innovation performance of emerging market enterprises: Evidence from China. Technovation, 62–63(??):4–13, April 2017. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649721730247X.

Yi:2021:RRS

[YHcHW21] Jingtao Yi, Junjie Hong, Wen chung Hsu, and Chengqi Wang. Reprint of “The role of state ownership and institutions in the innovation performance of emerging market enterprises: Evidence from China”. Technovation, ??(??): ??, ???? 2021. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497219300604.

Young:2008:IPM

[YHDR08] Beth Young, Nola Hewitt-Dundas, and Stephen Roper. Intellectual Property management in publicly funded R&D centres — a comparison of university-based and companybased research centres. Technovation, 28(8):473–484, August 2008. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497208000205.

Yang:2022:EAV

[YHL22] Jialei Yang and Pia Hurmelinna-Laukkanen. Evolving appropriability — variation in the relevance of appropriability mechanisms across industries. Technovation, 118 (??):??, December 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222001407.

Yasuda:2005:LBS

[YI05] Hiroshi Yasuda and Junichi Iijima. Linkage between strategic alliances and firm’s business strategy: the case of semiconductor industry. Technovation, 25(5):513–521, May 2005. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497203001706.

Yu:2021:CCD

[YIKC21] Yan Yu, Julio E. Ibarra, Kuldeep Kumar, and Vasilka Chergarova. Coevolution of cyberinfrastructure development and scientific progress. Technovation, 100(??):??, February 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497220300523.

Yoon:2020:SAP

[YJLL20] Byungun Yoon, Yujin Jeong, Keeeun Lee, and Sungjoo Lee. A systematic approach to prioritizing R&D projects based on customer-perceived value using opinion mining. Technovation, 98(??):??, December 2020. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497218306874.  See corrigendum

[YJLL21].

Yoon:2021:CSA

[YJLL21] Byungun Yoon, Yujin Jeong, Keeeun Lee, and Sungjoo Lee. Corrigendum to “A systematic approach to prioritizing R&D projects based on customer-perceived value using opinion mining” [Technovation 98 (2020) 102164]. Technovation, 102(??):??, April 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000018. See [YJLL20].

Yang:2008:SAG

[YK08] Suechin Yang and Hsin-Hong Kang. Is synergy always good? Clarifying the effect of innovation capital and customer capital on firm performance in two contexts. Technovation, 28(10):667–678, October 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720800014X.

Yanez:2010:IED

[YKW10] Mario Yanez, Tarek M. Khalil, and Steven T. Walsh. IAMOT and education: Defining a technology and innovation management (TIM) body-of-knowledge (BoK) for graduate education (TIM BoK). Technovation, 30(7–8): 389–400, July/August 2010. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497210000489.

Youtie:2006:IUR

[YLB06] Jan Youtie, Dirk Libaers, and Barry Bozeman. Institutionalization of university research centers: the case of the National Cooperative Program in Infertility Research. Technovation, 26(9):1055–1063, September 2006. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720500177X.

Yoon:2022:ECN

[YLH22] Jungsub Yoon, Jeong-Dong Lee, and Seogwon Hwang. Episodic change: a new approach to identifying industrial transition. Technovation, 115(??):??, July 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383

(electronic). URL https://www.sciencedirect.com/
science/article/pii/S0166497222000219.

Yildiz:2021:ICD

[YMK^{+}21] H. Emre Yildiz, Adis Murtic, Magnus Klofsten, Udo Zander, and Anders Richtnér. Individual and contextual determinants of innovation performance: a micro-foundations perspective. Technovation, ??(??):??, ???? 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497218307843.

Yildiz:2024:ILA

[YMMYK24] H. Emre Yildiz, Adis Murtic, Sergey Morgulis-Yakushev, and Magnus Klofsten. Individual-level absorptive capacity: Unveiling the interplay between dispositions and work context. Technovation, 131(??):??, March 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000154.

Yakob:2018:CFA

[YNS18] Ramsin Yakob, H. Richard Nakamura, and Patrik Ström. Chinese foreign acquisitions aimed for strategic assetcreation and innovation upgrading: The case of Geely and Volvo Cars. Technovation, 70–71(??):59–72, ???? 2018. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497218301160.

Yousafzai:2003:PMT

[YPF03] Shumaila Y. Yousafzai, John G. Pallister, and Gordon R. Foxall. A proposed model of e-trust for electronic banking. Technovation, 23(11):847–860, November 2003. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497203001305.

Yap:1993:FST

[YS93] Chee Meng Yap and Wm. E. Souder. A filter system for technology evaluation and selection. Technovation, 13(7): 449–469, November 1993. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016649729390028T.

See [Ano93e].

Yun:2021:SJI

[YSKL21] Siyeong Yun, Kisik Song, Chulhyun Kim, and Sungjoo Lee. From stones to jewellery: Investigating technology opportunities from expired patents. Technovation, 103(??):??, May 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649722100016X.

Yu:2009:UBL

[YT09] Chian-Son Yu and Yu-Hui Tao. Understanding businesslevel innovation technology adoption. Technovation, 29 (2):92–109, February 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497208000886.

Yanez-Valdes:2024:DID

[YVG24] Claudia Yáñez-Valdés and Maribel Guerrero. Determinants and impacts of digital entrepreneurship: a pre- and post-COVID-19 perspective. Technovation, 132(??):??, April 2024. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497224000336.

Yoshikawa:2008:SSE

[YW08] Gentoku Yoshikawa and Chihiro Watanabe. Structural source enabling firm revitalization innovation of sector — an empirical analysis of Japanese 31 industrial Sectors. Technovation, 28(1–2):37–51, January/February 2008. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497207000788.

Yang:2009:IQI

[YWC09] Mei-Ling Yang, Andrew Ming-Long Wang, and Kuo-Chih Cheng. The impact of quality of IS information and budget slack on innovation performance. Technovation, 29(8):527– 536, August 2009. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497209000054.

Ying:2022:MBS

[YWL22] Ying Ying, Shixiang Wang, and Yang Liu. Make bricks without straw: Eco-innovation for resource-constrained firms in emerging markets. Technovation, 114(??):??, June 2022. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222000645.

Yuan:2021:BSA

[YXH21] Chun Yuan, Doudou Xue, and Xin He. A balancing strategy for ambidextrous learning, dynamic capabilities, and business model design, the opposite moderating effects of environmental dynamism. Technovation, 103(??):??, May 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000067.

Yuan:2010:CSF

[YZY10] Li Yuan, Su Zhongfeng, and Liu Yi. Can strategic flexibility help firms profit from product innovation?

Technovation, 30(5–6):300–309, May/June 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497209001096.

Zaffaroni:1981:AEB

[Zaf81] Alejandro Zaffaroni. ALZA: an enterprise in biomedical innovation. Technovation, 1(2):135–146, August 1981. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497281900158.

Zirar:2023:WWA

[ZAI23] Araz Zirar, Syed Imran Ali, and Nazrul Islam. Worker and workplace Artificial Intelligence (AI) coexistence: Emerging themes and research agenda. Technovation, 124(??):??, June 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223000585.

Zakaryan:2023:OKN

[Zak23] Arusyak Zakaryan. Organizational knowledge networks, search and exploratory invention. Technovation, 122(??): ??, April 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497222002310.

Zhang:2019:CTI

[ZC19] Haisu Zhang and Weizhi Chen. Crowdfunding technological innovations: Interaction between consumer benefits and rewards. Technovation, 84–85(??):11–20, June/ July 2019. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497217305837.

Zhang:2022:NGD

[ZC22] Yi Zhang and Kaihua Chen. Network growth dynamics: the simultaneous interaction between network positions and research performance of collaborative organisations. Technovation, 115(??):??, July 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222000852.

Zhang:2019:SET

[ZCF19] Yi Zhang, Kaihua Chen, and Xiaolan Fu. Scientific effects of Triple Helix interactions among research institutes, industries and universities. Technovation, 86–87(??):33– 47, August/September 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497218300890.

Zhang:2023:RID

[ZCHC23] Stephen X. Zhang, Jiyao Chen, Liangxing He, and Afreen Choudhury. Responsible Innovation: the development and validation of a scale. Technovation, 124(??):??, June 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223000652.

Zhong:2022:EPS

[ZCR22] Xi Zhong, Weihong Chen, and Ge Ren. The effects of performance shortfalls on firms’ exploitation and exploration R&D internationalization decisions: Does industry environmental matter? Technovation, 112(??):??, April 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001899.

Zeschky:2014:CGR

[ZDWG14] Marco Zeschky, Michael Daiber, Bastian Widenmayer, and Oliver Gassmann. Coordination in global R&D organizations: an examination of the role of subsidiary mandate and modular product architectures in dispersed R&D organizations. Technovation, 34(10):594–604, October 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497214000832.

Zegveld:1987:TCI

[Zeg87] Walter Zegveld. Technology and change in industrial societies: Implications for public policy. Technovation, 5 (4):215–227, February 1987. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0166497287900642.

Zou:2023:AHS

[ZGSG23] Bo Zou, Jinyu Guo, Sunny Li Sun, and Feng Guo. Achieving harmony: Social identification in academic entrepreneurs’ role transition. Technovation, 128(??):??, December 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001700.

Zhou:1992:SQCa

[Zho92a] Sanduo Zhou. Status quo of China’s economy and new progress in enterprise reform. Technovation, 12(8):515–520, December 1992. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016649729290074R. See [Zho92b].

Zhou:1992:SQCb

[Zho92b] Sanduo Zhou. Status quo of China’s economy and new progress in enterprise reform: Technovation, 12:8 (1992) 515–520. Technovation, 12(8):536–537, December 1992. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729290078V. See [Zho92a].

Zhang:2022:EDA

[ZHY22] Wenjie Zhang, Lingling He, and Hongping Yuan. Enterprises’ decisions on adopting low-carbon technology by considering consumer perception disparity. Technovation, 117(??):??, September 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497221000195.

Zhu:2021:ETD

[ZHZL21] Shanshan Zhu, John Hagedoorn, Shuhui Zhang, and Fengchao Liu. Effects of technological distance on innovation performance under heterogeneous technological orientations. Technovation, 106(??):??, August 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221000821.

Zhang:2001:MPD

[ZI01] Wei Zhang and Barbara Igel. Managing the product development of China’s SPC switch industry as an example of CoPS. Technovation, 21(6):361–368, June 2001. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000341.

Zabala-Iturriagagoitia:2021:PNI

[ZIAO^{+}21] Jon Mikel Zabala-Iturriagagoitia, Juan Aparicio, Lidia Ortiz, Elias G. Carayannis, and Evangelos Grigoroudis. The productivity of national innovation systems in Europe: Catching up or falling behind? Technovation, 102(??):??, April 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497220300870.

Zamborsky:2023:KCC

[ZIB23] Peter Zámborský, Igor Ingrst, and Krishna Raj Bhandari. Knowledge creation capability under different innovationinvestment motives abroad: the knowledge-based view of international innovation management. Technovation, 127(??): ??, September 2023. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497223001402.

Zahoor:2023:VAI

[ZKW^{+}23] Nadia Zahoor, Zaheer Khan, Jie Wu, Shlomo Y. Tarba, Francis Donbesuur, and Huda Khan. Vertical alliances and innovation: a systematic review of the literature and a future research agenda. Technovation, 122(??):??, April 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222001353.

Zhang:2001:ISC

[ZL01] Michael J. Zhang and Augustine A. Lado. Information systems and competitive advantage: a competency-based view. Technovation, 21(3):147–156, March 2001. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497200000304.

Zhou:2024:DGD

[ZL24] Chao Zhou and Feng Lin. Does global diversification promote or hinder green innovation? Evidence from Chinese multinational corporations. Technovation, 129(??):??, January 2024. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016649722300216X.

Zhou:2018:OVW

[ZLC18] Michael Zhou, Mark A. A. M. Leenders, and Ling Mei Cong. Ownership in the virtual world and the implications for long-term user innovation success. Technovation, 78(??):56–65, December 2018. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497216303327.

Zobel:2017:FIA

[ZLH17] Ann-Kristin Zobel, Boris Lokshin, and John Hagedoorn. Formal and informal appropriation mechanisms: the role of openness and innovativeness. Technovation, 59(??):44– 54, January 2017. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497216303339.

Zhou:2021:MFB

[ZLJ^{+}21] Jianghua Zhou, Jizhen Li, Hao Jiao, Hang Qiu, and Zixu Liu. The more funding the better? The moderating role of knowledge stock on the effects of different governmentfunded research projects on firm innovation in Chinese cultural and creative industries. Technovation, ??(??):??, ???? 2021. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497217303565.

Zahra:2023:HDT

[ZLS23] Shaker A. Zahra, Wan Liu, and Steven Si. How digital technology promotes entrepreneurship in ecosystems. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222000049.

Zheng:2022:IEB

[ZMH22] Li Zheng, Pengcheng Ma, and Jacky Fok Loi Hong. Internal embeddedness of business group affiliates and innovation performance: Evidence from China. Technovation, 116(??):??, August 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497222000414.

Zhu:2023:TMS

[ZMJM23] Peiyu Zhu, Xiaoming Miao, Shumo Jin, and Robert Moehler. Transactive memory system, boundary-spanning search and business model innovation: the moderating role of environmental dynamism. Technovation, 128(??):??, December 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0166497223001633.

Zamani-Miandashti:2014:ITR

[ZMPRP14] Naser Zamani-Miandashti, Gholamreza Pezeshki-Rad, and Jaber Pariab. The influence of telecenters on rural life and their success correlates: Lessons from a case study in Iran. Technovation, 34(5–6):306–314, May/June 2014. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497213000734.

Zhang:2019:EBP

[ZQW^{+}19] Min Zhang, Yinan Qi, Zhiqiang Wang, Xiande Zhao, and Kulwant S. Pawar. Effects of business and political ties on product innovation performance: Evidence from China and India. Technovation, 80–81(??):30–39, February/March 2019. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497216302887.

Zhao:1993:MITa

[ZR93a] Liming Zhao and N. Mohan Ready. Managing international technology transfer negotiation: a social exchange perspective. Technovation, 13(6):383–397, September 1993. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016649729390078A.

Zhao:1993:MITb

[ZR93b] Liming Zhao and N. Mohan Reddy. Managing international technology transfer negotiation: a social exchange perspective. Technovation, 13(6):404–405, September 1993.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0166497293900827.

Zutshi:2004:EMS

[ZS04a] A. Zutshi and A. Sohal. Environmental management system adoption by Australasian organisations: Part 1: reasons, benefits and impediments. Technovation, 24(4):335– 357, April 2004. CODEN ???? ISSN 0166-4972 (print),

1879-2383 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0166497202000536.

Zutshi:2004:SEM

[ZS04b] Ambika Zutshi and Amrik Sohal. A study of the environmental management system (EMS) adoption process within Australasian organisations 2. Role of stakeholders. Technovation, 24(5):371–386, May 2004. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497202001153.

Zhou:2016:PTT

[ZSMB16] Haibo Zhou, Philipp G. Sandner, Simon Luca Martinelli, and Joern H. Block. Patents, trademarks, and their complementarity in venture capital funding. Technovation, 47(??):14–22, January 2016. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497215000796.

Zhang:2024:DES

[ZSX24] Xun Zhang, Kathy Ning Shen, and Biao Xu. Double-edged sword of knowledge inertia: Overcoming healthcare professionals’ resistance in innovation adoption. Technovation, 133(??):??, May 2024. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497224000610.

Zhou:2021:RUK

[ZT21] Ruoying Zhou and Puay Tang. The role of university Knowledge Transfer Offices: Not just commercialize research outputs! Technovation, ??(??):??, ???? 2021.

CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497217307897.

Zenobia:2009:AMR

[ZWD09] Brent Zenobia, Charles Weber, and Tugrul Daim. Artificial markets: a review and assessment of a new venue for innovation research. Technovation, 29(5):338–350, May 2009. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720800120X.

Zhang:2021:HDN

[ZWDZ21] Gupeng Zhang, Xiao Wang, Hongbo Duan, and Leven J. Zheng. How do new entrants’ pre-entry technological backgrounds impact their cross-industry innovation performances? A retrospective study of the mobile phone vendors. Technovation, 100(??):??, February 2021. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497220300481.

Zhang:2023:HPB

[ZWZ^{+}23] Xi Zhang, Xin Wei, Te Zhang, Yahe Tan, Dongming Xu, and Patricia Ordóñez de Pablos. How platform-based Internet hospital innovation affects doctors’ active stress coping efforts: the conservation of resource theory perspective. Technovation, 121(??):??, March 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497222001031.

Zhu:2019:FBF

[ZXDG19] Xiaoxuan Zhu, Zhenxin Xiao, Maggie Chuoyan Dong, and Jibao Gu. The fit between firms’ open innovation and business model for new product development speed: a contingent perspective. Technovation, 86–87(??):75–85, August/September 2019. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0166497217307642.

Zeng:2010:RBC

[ZXT10] S. X. Zeng, X. M. Xie, and C. M. Tam. Relationship between cooperation networks and innovation performance of SMEs. Technovation, 30(3):181–194, March 2010. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016649720900131X.

Zhan:2023:CMC

[ZXX23] Yuanzhu Zhan, Yu Xiong, and Xinjie Xing. A conceptual model and case study of blockchain-enabled social media platform. Technovation, 119(??):??, January 2023. CODEN ???? ISSN 0166-4972 (print), 1879-

2383 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0166497222001572.

Zhao:2021:SOA

[ZYSL21] Mengli Zhao, Jianjun Yang, Chengli Shu, and Jinxin Liu. Sustainability orientation, the adoption of 3D printing technologies, and new product performance: a crossinstitutional study of American and Indian firms. Technovation, 101(??):??, March 2021. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0166497220300699.

Zhang:2024:DER

[ZYZZ24] Feng Zhang, Xile Yin, Haina Zhang, and Xiangcen Zhan. Does environmental regulation enhance servitization in aspirant markets? Evidence from China’s manufacturing sector. Technovation, 131(??):??, March 2024. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497224000026.

Zhangqi:2022:TII

[ZZL22] Zhong Zhangqi, Chen Zhuli, and He Lingyun. Technological innovation, industrial structural change and carbon emission transferring via trade — an agent-based modeling approach. Technovation, 110(??):??, February 2022. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497221001310.

Zhao:2023:QQR

[ZZL^{+}23] Shengchao Zhao, Deming Zeng, Jian Li, Ke Feng, and Yao Wang. Quantity or quality: the roles of technology and science convergence on firm innovation performance. Technovation, 126(??):??, August 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001013.

Zhang:2023:MRE

[ZZXW23] Feng Zhang, Lei Zhu, Zhi Xu, and Yingying Wu. Moving from reverse engineering to disruptive innovation in emerging markets: the importance of knowledge creation. Technovation, 125(??):??, July 2023. CO-DEN ???? ISSN 0166-4972 (print), 1879-2383 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0166497223001025.